A Complete Bibliography of Statistics in Medicine (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 January 2021 Version 1.01

Title word cross-reference

$ [1550]. (Y + X)/2 [2687]. $100.95 [3222]. 100 log_{e} [295, 841]. $105.00 [2502]. $114.50 [2503]. $119.95 [2345]. $123.00 [2859]. $124.50 [1787]. $13.95 [455]. $130.00 [3181, 3138]. $139-95 [1503]. $155.00 [871]. $175 [201]. $175.00 [1786, 1438]. $195.00 [495]. 2 [1151]. 25 [2248]. $28.00 [1698]. $29.29 [1802]. 2 × 2 [3135, 3357, 2804, 1292, 1544, 276, 1266, 3179, 1007, 1543]. 2 × 3 [716]. 2 × K [778, 395]. 2 × m [2093]. $31.95 [3139]. $34.99 [1671]. $39.95 [843]. 3 × 3 [1703]. 4 [1317]. $44.00 [2961]. $44.50 [2503]. $44.95 [2399, 1839]. $45.00 [2006]. $49.46 [2755]. $49.50 [94]. $49.95 [1751, 1437, 872]. $50.00 [2519]. $54.00 [2857]. $54.95 [2718]. $59.95 [543, 1533, 2720]. $60.00 [1825]. $60.40 [285]. $65.00 [1802]. $69.0 [603]. $69.50 [1787]. $69.95 [2754, 526, 3204, 2409, 3347, 1840, 2858, 2005, 1610]. $71.96 [2841]. 74 [1470]. $74.50 [91]. $74.94 [1750]. $74.95 [2501]. $78.00 [1774, 2109]. $79.95 [1594, 1460, 1647, 1873, 2004, 3312, 3057, 2140, 3180, 821, 1609, 869, 1546, 2139, 2719, 2295, 2034, 210, 1416]. $80.00 [3182, 1671]. $83.95 [3311]. $84.95 [2138, 566, 1684]. $85.00 [1752]. $85.90 [41]. $89.85 [2718]. $89.95 [2824, 2604, 2959, 2960, 3122, 1516]. $90 [258]. $95.00 [820]. $96.00 [2856]. $99.50 [542]. $99.95 [3203, 1801, 545]. ^{13}C [1430]. _{3} [1028]. _{50} [836]. A + B [2177]. C [1434]. χ^{2} [1229]. D [2809, 2808]. E/O [2868]. F [2939, 2337]. G [1061, 3266]. γ [135]. K [2087, 1165, 1678, 3235]. K ≥ 3 [1743]. λ [2669]. M [3200, 1724]. n [1791]. P [2937, 4, 3310, 588, 3213, 2423, 3386, 2532, 1266, 977, 857]. P (X < Y) [2648]. P (Y_{1} > Y_{2}) [2934]. q [1204]. R [1812, 1617]. R^{2} [15, 39, 3063, 1997, 135]. r × c [2536]. r × k [193, 283]. S [73, 762]. t [2737, 1998, 3258, 2401]. T^{2} [1229]. τ [3005]. X [2687].

-computation [3266]. -distributed [2401]. -distributions [73, 762]. -estimation [1061]. -in-a-row [3235]. -interval [1204]. -like [1229]. -Margin [2669]. -of-1 [1791]. -optimal [2809, 2808]. -sample [2087]. -squared [1812]. -statistic [2737]. -structures [3200]. -tabulated [2470]. -tests [3258]. -to-one [1724]. -urea [1430]. -value [3386, 3310, 2532, 2937, 588]. -values [2423, 977, 857, 4, 3213]. -year [1317, 2248]. -year-old [1470].

0 [1530, 91, 341, 819, 817, 1157, 1647, 93, 1565, 454, 1592, 421, 542, 198, 509, 2138, 1752, 1786, 1351, 601, 144, 455, 360, 2824, 526, 885, 434, 1158, 497, 1608, 286, 782, 361, 143, 200, 127, 183, 1750, 301, 145, 402, 843, 362, 288, 1415, 129, 1566, 1550, 90, 1134, 543, 1402, 358, 600, 2004, 199, 399, 42, 1461, 420, 1438, 602, 201, 3311, 303, 1245, 511, 1180, 1548, 1194, 1182, 541, 211, 1352, 182, 566, 1018, 1195, 342, 844, 1774, 3182, 439, 196, 1310, 417, 1144, 1181, 436, 92, 438, 1311, 2140, 1159, 1309, 104]. 0 [435, 1133, 495, 1142, 130, 300, 302, 1118, 1531, 524, 299, 1595, 1156, 221, 132, 962, 142, 1801, 359, 357, 525, 3056, 1193, 1363, 440, 1143, 1145, 2858, 544, 869, 2519, 872, 419, 1192, 43, 257, 1546, 146, 1471, 2139, 2502, 287, 870, 260, 1648, 545, 1516, 2841, 1206, 2345, 128, 298, 400, 418, 820, 103, 540, 1517, 1065, 510, 187, 284, 1384, 1825, 1698, 147, 105, 259, 184, 496, 842, 1066, 433, 210, 401, 1416, 871, 94]. 0-0893-7168-6 [869]. 0-11-620917-8 [818]. 0-11-621049-4 [818]. 0-12-274065-3 [1065]. 0-1264-2350-4 [435]. 0-19-262313-3 [146]. 0-19-262745-7 [540]. 0-19-262919-0 [420]. 0-19-262941-7 [1142]. 0-19-262959-X [1531]. 0-19-262992-1 [1180]. 0-19-514489-9 [1143]. 0-19-514525-9 [1245]. 0-19-515296-4 [1595]. 0-19-517179-9 [1802]. 0-19-517180-2 [1802]. 0-19-850485-3 [302]. 0-19-850765-8 [1592]. 0-19-850984-7 [1787]. 0-19-850994-4 [1415]. 0-19-852436-6 [187]. 0-19-852615-6 [1608]. 0-19-852894-9 [1549]. 0-19-852895-7 [1549]. 0-19-856582-8 [1787]. 0-19-856701-4 [2519]. 0-19-856760-X [2503]. 0-19-856761-8 [2503]. 0-1926-2785-6 [144]. 0-262-19440-6 [1133]. 0-306-45956-6 [542]. 0-316-75780-2 [91]. 0-335-19885-6 [399]. 0-340-61461-7 [211]. 0-340-64555-5 [92]. 0-340-64629-2 [286].

0-340-66230-1 [299]. 0-340-66241-7 [145]. 0-340-67783-X [1825]. 0-340-691530 [817]. 0-340-69185-9 [142]. 0-340-69228-6 [600]. 0-340-69243-X [361]. 0-340-71940-0 [128]. 0-340-71995-8 [842]. 0-340-71997-4 [357]. 0-340-740590 [454]. 0-340-74120-1 [288]. 0-340-74125-2 [962]. 0-340-74149X [436]. 0-340-76112-1 [1193]. 0-340-761431 [602]. 0-340-80655-9 [1753]. 0-387-00727 [1546]. 0-387-00749-0 [1684]. 0-387-098819-X [1384]. 0-387-20275-7 [2140]. 0-387-20279-X [2138]. 0-387-40081-8 [1750]. 0-387-94803-1 [525]. 0-387-94909-7 [196]. 0-387-94996-8 [301]. 0-387-95157-1 [872]. 0-387-95232-2 [1157]. 0-387-95277-2 [1517]. 0-387-95284-5 [1310]. 0-387-95370-1 [1647]. 0-387-95389-2 [1548]. 0-387-95399-X [1352]. 0-387-95577-1 [1516]. 0-387-98216-7 [129]. 0-387-98503-4. [419]. 0-387-98504-2 [434]. 0-387-98563-8 [1156]. 0-387-98671-5 [497]. 0-387-98784-3 [526]. 0-387-98814-9 [1182]. 0-387-98828-9 [510]. 0-387-98844-0 [1565]. 0-387-98863-7 [1195]. 0-387-98873-4 [566]. 0-387-98997-7 [543]. 0-412-04061-1 [104]. 0-412-04411-0 [90]. 0-412-04731-4 [439]. 0-412-31780-X [184]. 0-412-60650-X [130]. 0-412-71630-5 [43]. 0-412-72860-5 [105]. 0-412-75410-X [42]. 0-412-81820-5 [93]. 0-412-9828-1 [198]. 0-412-99631-6 [342]. 0-412-99718-8 [200]. 0-444-50079-0 [1786]. 0-470-84145-1 [1194]. 0-470-84337-3 [1416]. 0-471-13404-X [287]. 0-471-13773-1 [438]. 0-471-15410-5 [182]. 0-471-15575-6 [496]. 0-471-16386-4 [359]. 0-471-17082-8 [298]. 0-471-18386-5 [1363]. 0-471-19045-4 [147]. 0-471-19365-8 [524]. 0-471-21487-6 [2502]. 0-471-22490-1 [1145]. 0-471-22615-7 [1311]. 0-471-24029-X [103]. 0-471-29008-4 [402]. 0-471-31711-X [400]. 0-471-34991-7 [418]. 0-471-35632-8 [819]. 0-471-36357-X [1530]. 0-471-36914-4 [1566]. 0-471-39405-X [1471]. 0-471-39685-0 [1134]. 0-471-48971-90 [1774]. 0-471-48986-7 [1402]. 0-471-49066-0 [1206]. 0-471-49975-7 [1752]. 0-471-52629-0 [1801]. 0-471-86094-8 [1351]. 0-471-95816-6 [260]. 0-471-96554-5 [358]. 0-471-96583-9 [360]. 0-471-96861-7 [544]. 0-471-97572-9 [1144]. 0-471-97576-1 [127]. 0-471-98370-5 [1309]. 0-471-98608-9 [1018]. 0-471-98611-9 [303]. 0-471-98634-8 [417]. 0-471-98731-2 [509]. 0-471-99890-7 [820]. 0-4719-7645-8 [183]. 0-512-63567-5 [433]. 0-521-47266-0 [259]. 0-521-47860-X [259]. 0-521-54023-2 [1698]. 0-521-54027-5 [1671]. 0-521-59693-9 [440]. 0-521-64079-2 [885]. 0-521-66945-6 [1158]. 0-521-77362-8 [843]. 0-521-83261-6 [1671]. 0-632-052570 [1118]. 0-691-05782-6 [94]. 0-7169-6927-0 [1156]. 0-7169-7385-0 [1156]. 0-7279-1599-1 [1159]. 0-7619-0257-0 [257]. 0-7619-0937-0 [132]. 0-7619-1323-8 [455]. 0-7619-1440-4 [221]. 0-7619-1621-0 [870]. 0-7619-2026-9 [1461]. 0-7619-6151-8 [284]. 0-7619-8533-6 [511]. 0-7923-4159-7 [421]. 0-79234-257-7 [1066]. 0-8039-7398-5 [199]. 0-8176-3901-2 [300]. 0-8247-0130-5 [201]. 0-8247-0467-3 [1192]. 0-8247-5469-7 [2345]. 0-8247-7572-4 [495].

0-8247-9025-1 [1438]. 0-8247-9030-8 [871]. 0-849-30316-8 [541]. 0-849-30387-7 [844]. 0-849-30389-3 [143]. 0-8493-0385-0 [601]. 0-8493-0386-9 [362]. 0-8493-0398 [341]. 0-8493-2805-5 [545]. 0-8493-3132-8 [210]. 0-8493-3799-2 [2824]. 0-86542-871-9 [1648]. 0-89603-951-X [1550]. 00727 [1546]. 034 [453]. 034-067668 [453]. 0340 [186]. 0340-6775-9 [186]. 0398 [341]. 0471 [452, 781]. 0471-34912-7 [452]. 0471-36996-9 [781]. 052570 [1118]. 067668 [453]. 07-115 [199]. 07-124 [221]. 07-126 [455]. 0962 [141]. 0962-2802 [141].

1 [1594, 1460, 1647, 198, 1425, 1364, 1076, 782, 1751, 127, 1873, 288, 2006, 603, 2004, 1119, 1438, 1011, 961, 1180, 1067, 1194, 1205, 2399, 438, 2409, 1533, 1159, 2299, 1593, 104, 1090, 780, 299, 2959, 821, 1609, 3192, 1840, 1437, 525, 1940, 1193, 1145, 872, 2976, 2005, 2109, 2139, 2295, 3122, 1516, 185, 2841, 3138, 2034, 1547, 1610, 1504, 1839, 1244, 1503, 2683, 2755]. 1-5-488-195-X [780]. 1-57808-304-4 [2006]. 1-58025-397-0 [782]. 1-584-88145-3 [1425]. 1-5848-8216-6 [1119]. 1-58488-075-9 [1244]. 1-58488-094-5 [821]. 1-58488-160-7 [1205]. 1-58488-165-8 [1437]. 1-58488-170-4 [961]. 1-58488-175-5 [1593]. 1-58488-1852 [1547]. 1-58488-208-5 [2295]. 1-58488-265-4 [1504]. 1-58488-266-2 [1364]. 1-58488-307-3 [1460]. 1-58488-308-1 [1594]. 1-58488-321-9 [2034]. 1-58488-335-9 [1609]. 1-58488-369-3 [1839]. 1-58488-388-X [1533]. 1-58488-404-5 [1751]. 1-58488-405-3 [1873]. 1-58488-408-8 [1610]. 1-58488-415-0 [2004]. 1-58488-425-8 [2005]. 1-58488-432-0 [2139]. 1-58488-433-9 [1840]. 1-58488-440-1 [1503]. 1-58488-4509 [2399]. 1-58488-480-0 [2841]. 1-58488-587-4 [2858]. 1-58488-619-6 [2959]. 1-58488-756-7 [2755]. 1-58488-776-1 [2960]. 1-58488-982-9 [3122]. 1-59047-500-3 [2604]. 1-59047-504-6 [2409]. 1-59718-008-4 [2857]. 1-86094-153-2 [1067]. 1-86094-441-8 [2109]. 1-898298-59-9 [185]. 1.Principles [1065]. 10 [414]. 10.1002 [2751, 2820, 3070, 2450, 3379, 2653, 2489, 2452, 2770, 2839]. 10.1002/sim.2538 [2450]. 10.1002/sim.2618 [2452]. 10.1002/sim.2655 [2476, 2489]. 10.1002/sim.2822 [2476]. 10.1002/sim.2832 [2751]. 10.1002/sim.2855 [2653]. 10.1002/sim.2929 [2638, 2641, 2640, 2643, 2642, 2637, 2639]. 10.1002/sim.2974 [2770]. 10.1002/sim.3034 [2839]. 10.1002/sim.3069 [2820]. 10.1002/sim.3122 [3070]. 10.1002/sim.3722 [3379]. 10.2002 [2344]. 10.2002/sim.2600 [2344]. 100 [284]. 1060 [714]. 11 [857]. 1130 [1591, 1709]. 1193 [2002, 1645]. 1196 [813]. 12 [297]. 135 [1728]. 1359 [3190]. 138DM [300]. 1430 [2676]. 1454 [815]. 1553 [3386, 3310]. 16 [2379, 125, 13, 10, 11, 29]. 1610. [1362]. 1623 [1799]. 1627 [2245]. 1637 [1715, 1714, 1716]. 1679 [3382]. 17 [1850]. 1702 [3381]. 1704 [297]. 1712 [2183]. 1723 [2052]. 1732 [181]. 1752 [397]. 176 [1589]. 1773 [414]. 1785 [1053]. 18 [355, 39, 387, 78, 354, 397]. 1852 [1547]. 1864 [416]. 19 [538, 2000, 841, 386, 757, 960, 416, 537, 283]. 1905 [125]. 1912 [259]. 1914 [283]. 1950s [1225]. 1960s [1225]. 1967 [757]. 1980s [1226]. 1988 [3221]. 1997 [1470]. 1999 [305].

2 [91, 1157, 509, 2824, 885, 434, 1364, 286, 1011, 1067, 1548, 300, 962, 1941, 1517, 1698, 2720, 2684, 41]. 20 [1374, 803, 813, 1337, 1925, 714, 1016, 958, 2911, 759]. 2007 [2912]. 2012 [2160]. 203 [2327]. 2040 [2817]. 2083 [1117, 1545, 1669]. 21 [1168, 1424, 1589, 1054, 1056, 2558, 1169, 1799, 1644, 2911, 2160, 1053, 1494]. 2100 [538]. 2126 [1649]. 2150 [2134]. 22 [1117, 1545, 1649, 1561, 1696, 1563, 1362, 1669]. 2266 [1682]. 23 [1591, 2002, 1784, 1682, 1645, 1728, 1725, 1711, 1450]. 231 [3384]. 24 [1951, 2052, 2033, 2134, 1715, 1714, 2817, 2183, 1851, 2186, 1716, 2032]. 2401 [2379]. 2438 [2033]. 2439 [1925]. 25 [2440, 2327, 2245, 2439, 2499, 3384, 2437, 2855, 2559]. 2523 [3106]. 2530 [3380]. 2535 [3135]. 25th [1979, 1939, 1914]. 26 [2751, 2910, 2907, 2735, 3024, 3190, 2676, 3385, 3136, 2819, 2734, 2787, 3302, 2476]. 2608 [1169]. 261 [1056]. 27 [3041, 3105, 3135, 3202, 3108, 3381, 3219, 3106, 3137, 3288]. 273 [3385]. 2736 [1561]. 28 [3386, 3380, 3310, 3383, 3382]. 2802 [141]. 2962 [960]. 2976 [1016]. 2A [299]. 2nd [1530, 91, 819, 3222, 2824, 2604, 361, 1750, 2004, 961, 1548, 1352, 844, 3181, 1533, 818, 1133, 495, 821, 1437, 1193, 1363, 1532, 2858, 1471, 3122, 1648, 540, 2720, 1066, 1549, 41, 2755].

3 [1460, 526, 2604, 1425, 143, 1873, 3204, 343, 40, 3312, 1181, 302, 1872, 494, 285, 1192, 146, 106, 3231, 1065, 197, 1839, 1416, 41]. 3-437-21056-4 [197]. 3-486-23146-4 [494]. 3-540-61609-8 [40]. 3-540-62201-2 [41]. 3-540-63737-0 [1181]. 3-631-34610-7 [343]. 3-7643-6818-7 [1872]. 3-7908-0982-9 [285]. 3003 [387]. 3009 [1424]. 3045 [2735, 2734]. 3047 [2855, 2559]. 3094 [2910, 3024, 2909]. 3125 [841]. 3151 [1644]. 3159 [1168]. 3207 [1784]. 3222 [2186]. 3289 [537]. 33 [2000]. 3373 [958]. 3489 [3202]. 35- [1470]. 351 [386]. 358 [1851]. 3596 [1696]. 3611 [3105]. 3629 [2050]. 3682 [3383]. 370 [3041]. 370. [1054]. 3751 [1374]. 3754 [3108]. 3774 [1494]. 3786 [2437]. 379 [1532]. 3817 [3302]. 3914’ [3137]. 3rd [2138, 3203, 1751, 1245, 1753, 2959, 3347, 1801, 2345, 298, 496, 210, 1503].

4 [402, 1415, 2006, 1566, 961, 566, 439, 3181, 435, 495, 1595, 221, 359, 357, 494, 2519, 437, 1504, 147, 197]. 4. [419]. 4292 [2499]. 4351 [2787]. 4385 [2907]. 4509 [2399]. 465 [1711]. 4th [1594, 3311, 1118, 2961, 842].

5 [93, 3222, 497, 1751, 358, 199, 201, 131, 182, 1310, 92, 1593, 1309, 2857, 821, 1363, 43, 2295, 545, 105, 433]. 5045 [2819]. 5880-5889 [3162]. 5974 [3288].

6 [542, 144, 1158, 1608, 3203, 600, 1119, 399, 511, 3057, 342, 2409, 1133, 2959, 3347, 869, 2502, 260, 187, 496, 94, 302]. 602 [2198]. 654 [803]. 656 [867]. 6611 [3219]. 672.Price [842]. 687 [2558, 2911]. 691530 [817].

7 [421, 1752, 145, 2501, 129, 543, 343, 1402, 2856, 211, 452, 1195, 1205, 844, 196, 1311, 2140, 1142, 1872, 544, 2345, 418, 820, 540, 1066, 2755, 2404]. 70 [2440, 2439]. 708 [1951]. 726 [1563]. 731 [1337]. 740 [2032]. 740590 [454]. 74149X [436]. 75 [1450]. 761431 [602]. 798 [355, 354]. 7th [359].

8 [819, 1592, 3139, 1351, 2754, 455, 200, 183, 1750, 301, 843, 40, 541, 417, 884, 524, 2960, 1437, 2005, 2109, 298, 1610, 284, 842, 210, 871, 436]. 8-10 [1156]. 8-tetrachlorodibenzo-p-dioxin [709]. 857 [1725]. 872 [1850]. 879 [857]. 882 [759].

9 [360, 362, 1461, 303, 1245, 1182, 1018, 1144, 867, 1609, 142, 781, 1840, 440, 285, 1143, 2961, 106, 2859, 3122, 1648, 185, 815, 2034, 510, 186, 1244]. 9-8102-1388-3 [106]. 90 [1774]. 92 [603, 131, 884, 437, 401]. 92-4-156199-8 [884]. 92-832-0405-0 [401]. 92-832-2143-5 [131]. 92-832-2145-1 [603]. 92-832-2149-4 [437]. 92832-2144-3 [437]. 964 [3136]. 97 [3057]. 97-1-584-88767-6 [3057]. 978 [3139, 3222, 2754, 2604, 3203, 2501, 3204, 2856, 3311, 3312, 3182, 3181, 2857, 3347, 3192, 2960, 3056, 2858, 2961, 2859, 2719, 3138, 3231, 2720]. 978-0-19-859966-1 [3192]. 978-0-19-929659-0 [3056]. 978-0-19-929660-6 [3056]. 978-0-387-22538-8 [2718]. 978-0-387-23530-1 [2718]. 978-0-387-30059-7 [2501]. 978-0-470-01877-4 [3181]. 978-0-470-02842-1 [3138]. 978-0-470-02864-3 [3231]. 978-0-470-31971-0 [3182]. 978-0-471-22618-5 [3222]. 978-0-471-65396-7 [2856]. 978-0-471-74583-9 [2859]. 978-1-58488-469-X [2719]. 978-1-58488-587-0 [2858]. 978-1-58488-624-2 [2720]. 978-1-58488-630-3 [3312]. 978-1-58488-668-6 [3203]. 978-1-58488-697-6 [3347]. 978-1-58488-701-0 [3311]. 978-1-58488-724-9 [3180]. 978-1-58488-776-8 [2960]. 978-1-59047-500-3 [2604]. 978-1-59047-886-8 [2754]. 978-1-59718-008-5 [2857]. 978-1-59994-165-3 [3204]. 978-1-59994-298-8 [3139]. 978-3-8055-8189-9 [2961].

A-Z [128]. Aalen [1762]. Abdelmonem [1594]. Abdi [221]. abdominal [621]. aberration [1626]. ability [2281, 2638, 3125, 2641, 2640, 2643, 2642, 2636, 2637, 2639, 3017]. abnormal [3258]. abnormality [631]. abortion [1093, 1278]. Abrams [1752, 1206]. Abramson [1245]. absence [876, 766, 3321]. absences [1667]. absenteeism [2180]. absolute [1121, 1799, 1277, 1301, 790]. abstraction [1181]. Academic [421, 435, 1065, 1066, 2262, 985]. Accelerated [2568, 36, 928, 2614, 1510, 925, 942, 1664, 2192, 2447]. acceleration [2474]. access [463]. accommodating [3013]. Accommodation [1900]. according [849]. account [2373, 1026]. Accountability [460]. Accounting [1525, 71, 3224, 2954, 656, 2778, 2412, 2829, 2827, 3270, 2117, 1865]. accounts [3000]. accrual [2086, 2793]. accumulating [702, 1475]. accuracies [1988, 173]. Accuracy [610, 3155, 18, 1272, 3270, 792, 387, 1466, 2457, 2749, 1581, 1987, 1023, 1395, 2182, 621, 1933, 1723, 1920, 2118, 2291, 591, 1138, 1445, 1989, 2732].

Accurate [2093, 2487]. Accutane [865]. achieve [2367]. Acion [2198]. Acknowledgement [1577, 1902, 2229, 2625, 2968]. acoustic [117]. acquisition [2978]. across [2513, 1428, 2012, 1343, 3201, 2662, 3031]. Act [460]. action [2162, 2078]. activation [1782]. Active [2655, 3229, 1980, 2334, 1038, 2400, 988, 3362, 1756, 2896, 2159, 1485, 513, 989, 3158, 3041]. Active-control [2655, 2896, 3041]. active-controlled [2159]. activity [2084]. Actuarial [143]. acuity [1555]. Acute [2371, 2041, 2899, 1563, 621, 1470, 2722, 1334, 1022]. AD [1286]. Adam [300]. adapt [2059]. Adaptation [1320, 1908, 476, 1140, 2577]. Adapting [3355]. Adaptive [1778, 2158, 2736, 3183, 2423, 3340, 3205, 2737, 2739, 2147, 1436, 805, 2579, 2440, 2432, 2738, 3206, 2439, 1907, 1700, 1518, 2740, 3333, 2148, 2151, 664, 2856, 209, 3226, 3304, 2413, 2330, 1333, 1776, 2674, 2394, 1563, 2611, 1265, 2368, 1699, 944, 1878, 1485, 1967, 1740, 3029, 1022, 1995, 2077, 2960]. add [3152, 1377]. added [2638, 2641, 2640, 2643, 2642, 2636, 2637, 2639, 604]. adding [2178]. additional [1900]. Additive [2083, 1889, 88, 2432, 1572, 1762, 1633, 1674, 1070, 2691, 3291, 2495, 592, 2887, 2084, 2901, 1449, 3240, 1048, 3110]. additivity [1569]. Addressing [3238]. Adelchi [130]. adenoma [2236]. adenomas [2338]. ADEPT [3011, 3010]. adherence [2053, 1798]. adjunctive [2778]. adjust [1703, 1761, 407, 2799]. Adjusted [1828, 2540, 2623, 2871, 2876, 2425, 2724, 1398, 2602, 534, 1012, 3019, 1936, 1990, 2564, 1678, 2273, 3249, 3197, 1618, 34, 1740, 3185, 3173, 2518, 997]. Adjusting [2800, 1160, 1704, 1523, 1122, 1815, 1429, 1026, 1662, 337, 1426, 1427, 1289, 160, 3002, 1540, 2746, 2663, 1649]. Adjustment [948, 414, 1327, 2579, 1328, 2738, 3, 2750, 2449, 1696, 2521, 664, 1909, 634, 1821, 3216, 2330, 3045, 3177, 1246, 476, 1634, 2237, 2415, 2352, 1247, 1451, 506, 1780, 3076, 3098, 1479, 690, 902, 1756, 1208, 2530, 3279, 3099, 2971, 2364, 1609]. Adjustments [1086, 3265, 1171, 775, 3039]. adjuvant [2170]. Administration [1221]. Administrative [549, 1706, 1540]. Admissible [1314]. admission [909, 1424]. admitting [264]. ado [2234]. adolescent [1814, 3309]. adolescents [1002]. adoptees [1262]. Adrian [1873]. Adult [1734, 251, 1536, 2011, 1262]. adulthood [3161]. adults [1106]. Advanced [889, 713, 299, 1027]. advancements [1979]. Advances [1786, 2123, 984, 579]. Advantages [1446, 1682]. adversary [2666]. adverse [2757, 1398, 2335, 2124, 2940, 2201, 1579, 2202, 1420, 3012, 825]. advice’ [2644]. aerosol [2597]. Aerts [1547]. affect [909, 1424, 1246]. affected [702, 809]. affecting [2926, 2729]. affective [828]. Afifi [1594]. African [2262, 3131, 2253]. after [1613, 2606, 515, 1894, 1320, 1405, 2445, 1576, 282, 588, 1470, 1098, 2170, 390]. against [1792, 937, 723, 2333]. Age [1127, 2436, 3161, 2723, 1101, 334, 2244, 44, 352, 3241, 2924, 2472, 1830, 120, 2261, 2470, 1971, 1560, 960, 1820, 1678, 1665, 1408, 278, 2253, 2016, 2925, 73, 3355, 2620, 734, 1555, 521, 2132, 2889, 3215, 2735, 2734]. Age- [3161]. age-adjusted [1678]. Age-conditional [1101, 3215]. age-dependent [2016]. Age-period-cohort [334, 1971, 2734]. age-related [2924, 1408, 73, 1555]. Age-specific [1127, 2723, 44, 1830, 2261, 960, 278]. aged [552, 2070]. ageing [158, 863, 652]. Agency [603]. agent [2394]. agents [2073]. ages [637]. Aggregate [932, 1958, 2760, 3062, 2679]. aggregate-level [2679]. aggregated [2547]. aggregation [948, 3359, 2836, 3360, 3103, 1075, 2996, 2254]. aging [1220]. ago [306]. agonism [2114]. agreement [1690, 75, 191, 800, 38, 2273, 2294, 799, 2690, 3127, 2275, 3175, 2830, 2034]. Agresti [3222, 2032, 1450]. Ahn [3135]. aided [2363, 1594]. aiding [432]. AIDS [2770, 352, 45, 522, 3229, 1308, 116, 2967, 177, 396, 33, 2979, 2688, 102, 2965, 32, 3214, 1618, 2927, 2301, 1406, 2192]. AIDS-epidemic [2770, 2688]. aims [3095]. air [2448, 1321, 1604, 1667, 2215, 2171]. Aitchison [2295]. Akazawa [757]. al [2440, 3202, 2439, 2638, 2641, 2643, 3134, 2637, 2908, 2036, 3137, 2639]. al. [2826]. Alan [3222, 299, 2032]. Albert [1597, 181]. alcohol [554, 475, 2883, 3240]. alcohol-dependence [2883]. alcoholism [1657]. alerting [3333]. Alex [2754, 2409, 1206]. Alfred [183]. Alfredo [1872]. algorithm [3333, 682, 485, 1589, 2597, 1474, 2509, 2593, 2572, 2252, 3266]. algorithms [2810]. Alho [2718]. Alison [818]. Alistair [1156]. Allan [817]. allele [702]. allelic [1260]. allergic [548]. Allison [511]. Allocation [609, 17, 1778, 2540, 1862, 1700, 3339, 1368, 1563, 2368, 2726, 705, 397, 1022, 2957, 69]. allocations [1411]. allogeneic [1576]. allografting [282]. allow [3324, 3201]. Allowing [1366, 2683, 2684, 2086, 673, 3154]. allows [2554]. alone [220]. Alonzo [387]. Alpha [81, 80, 83, 82, 79]. alter [2393]. Alternative [3270, 980, 806, 423, 638, 2698, 2087, 271, 135, 643]. Alternatives [954, 1032, 138, 3054, 723]. Altman [2499, 537, 2687]. Alvan [1119]. always [1761]. Alzheimer [155, 1283, 157, 159, 158, 1287, 1284, 160, 152, 149, 163, 2397, 1297, 1298, 1296, 167, 168, 150, 153, 172, 1289, 151]. am [343]. amalgamating [2989]. ambiguous [1767]. ambulatory [659]. American [2485]. AMI [2060, 2432]. Amir [2787]. amniotic [1632]. among [2878, 1204, 2756, 504, 1260, 2927, 2485, 2922, 1995]. amount [2233]. Amsterdam [1786]. anaemia [121]. Anaesthesia [146]. analogue [175]. analyse [192, 1521, 749, 2523, 951, 1934]. analysed [217]. Analyses [844, 3036, 330, 1361, 2297, 1374, 875, 3085, 1331, 899, 3284, 1204, 3174, 1751, 1456, 3030, 3210, 174, 393, 1732, 558, 698, 2335, 318, 2698, 2164, 2680, 3380, 1777, 476, 1634, 2237, 2415, 1513, 2300, 3276, 3136, 2227, 727, 3362, 1546, 3012, 2685, 672, 655, 2974, 860, 1391, 971, 109, 788, 1486, 2090, 2433, 647, 1839, 2741, 2755, 2995]. analysi [2112, 3383]. Analysing [1638, 1853, 2223, 1278, 2323, 313, 121, 2305, 207, 1179, 953, 2470, 1320, 1796, 136, 1299, 477, 1379, 52, 690, 338, 1439, 443, 2053, 2696, 2531, 1189, 1666, 1147, 2489, 2476]. Analysis [1924, 1792, 645, 341, 817, 2527, 2321, 311, 2513, 1554, 601, 2504, 332, 1044, 1365, 2816, 402, 1443, 992, 1148, 666, 808, 2992, 5, 42, 778, 1888, 1067, 3144, 1779, 74, 1194, 1352, 209, 307, 92, 3224, 2705, 659, 684, 2061, 1893, 2835, 1156, 660, 3055, 2996, 359, 1318, 2264, 2456, 720, 1363, 254, 1143, 1145, 544, 253, 2562, 3295, 287, 1251, 2786, 422, 1417, 1512, 2210, 418, 2469, 153, 2250, 371, 186, 433, 1416, 41, 94, 633, 957, 2969, 3110, 1130, 2252, 1530, 1886, 3236, 2716, 3081, 1647, 1317, 1168, 2000, 559, 1064, 2944, 2481, 527, 1498, 1652, 2919]. analysis [2119, 875, 941, 157, 36, 803, 1163, 100, 45, 522, 2623, 2606, 3166, 3061, 710, 2332, 2172, 2473, 709, 2189, 403, 2550, 950, 323, 2181, 702, 1094, 1746, 3241, 77, 1514, 2907, 1482, 1630, 2621, 2891, 134, 516, 813, 772, 1183, 3344, 1249, 1844, 3203, 777, 296, 755, 108, 2334, 1782, 3270, 1814, 2758, 3364, 3184, 281, 1072, 1382, 2937, 444, 2538, 2939, 2598, 831, 832, 833, 1492, 1258, 789, 1655, 1009, 935, 120, 907, 2471, 1738, 3210, 2463, 2426, 174, 2196, 2847, 1155, 644, 3375, 933, 2008, 1412, 1520, 2378, 3077, 3078, 2420, 2267, 1215, 2175, 1097, 3176]. analysis [1227, 1968, 2612, 863, 2757, 350, 101, 3075, 201, 1279, 2029, 111, 181, 612, 1992, 2335, 318, 1558, 1077, 665, 501, 219, 1975, 610, 1656, 546, 787, 3060, 2272, 1863, 1821, 968, 1493, 3103, 2168, 2557, 630, 3217, 1137, 331, 2983, 3254, 1186, 1419, 2097, 3120, 2330, 1993, 3111, 921, 3027, 2042, 2366, 2733, 865, 297, 3142, 1640, 2044, 1925, 2076, 692, 21, 1701, 2769, 1551, 2249, 528, 956, 654, 2461, 1173, 1012, 2306, 495, 2392, 1557, 923, 1574, 1285, 451, 2826, 1202, 3019, 2794, 2417, 845, 3349, 1473, 2013, 2492, 2697, 3385, 955, 851, 1052, 180, 940, 1422, 1663, 2028]. analysis [2308, 3034, 3343, 1780, 807, 2491, 3148, 2558, 846, 1595, 1007, 385, 27, 465, 380, 2188, 719, 1028, 3271, 2395, 3101, 26, 1104, 1212, 1887, 1354, 2596, 348, 1820, 1852, 1356, 1728, 1927, 1256, 1537, 1139, 761, 3023, 1047, 2182, 14, 1721, 392, 1932, 1799, 1243, 942, 3091, 2373, 560, 681, 1420, 243, 1434, 1043, 646, 3350, 370, 2613, 2072, 1741, 902, 1439, 1059, 1614, 2388, 383, 2458, 3068, 2881, 753, 754, 2109, 1471, 1897, 2622, 312, 1937, 2235, 2760, 3062, 1335, 2719, 2761, 990, 2711, 2232, 3152, 591, 2236, 1516, 1259, 2269, 2686, 2270, 2309, 2710, 2572]. analysis [1984, 1455, 839, 1005, 582, 3138, 1945, 57, 837, 1206, 2280, 2574, 491, 89, 2228, 1149, 313, 1985, 1271, 3267, 2539, 2908, 920, 2796, 2434, 901, 1686, 1783, 1129, 2277, 2396, 3262, 2066, 691, 3242, 1511, 1623, 1610, 667, 31, 110, 78, 532, 1063, 335, 624, 2435, 2403, 2679, 2678, 1377, 389, 643, 838, 572, 1685, 1573, 1301, 1763, 1527, 2938, 734, 2524, 3173, 325, 373, 658, 1080, 2233, 3278, 1912, 3137, 3092, 680, 723, 1748, 2511, 2485, 2647, 790, 2825, 445, 1250, 604, 1885, 2683, 2684, 547, 620, 1588, 932, 824, 1705, 3248, 337, 517, 2390, 2852, 2192]. analysis [2337, 3275, 2830, 2706, 2904, 3309, 1048, 1692, 2406, 2670, 3194, 2101, 3218, 327, 713, 1809, 3017, 2002, 1645, 1157, 455, 361, 961, 182, 440, 260, 1594, 3222, 1786, 2004, 1548, 196, 1533, 3347, 2502, 298, 1517, 566, 438, 2409, 1309, 104, 2959, 1532, 1244, 871, 3312, 3231]. Analytic [731, 2111, 2357]. Analytical [1295, 259, 2234, 3307]. analyze [3094]. Analyzing [2894, 2548, 3378, 2915, 2731, 3280, 3059, 2844, 3139, 600, 199, 1684]. ancestor [1079]. ancillary [3038]. ANCOVA [1696, 1247]. and/or [1031, 1779]. Anders [2857]. Anderson [434, 2452]. Andrea [2499]. Andreas [1182]. Andrew [1364, 1067, 1194, 1195, 417, 1144, 1533, 2109, 259, 3150]. Andrzej [2841]. Anemona [3108]. Angela [600]. angina [1026]. angiography [1191]. angular [281, 2017]. animal [201, 1165]. Ann [399]. Anne [1750, 1309]. Annette [1437]. Annibale [417]. Annie [1750]. Anniversary [1979, 1939, 1914]. Annotated [1565]. annual [3364, 251, 3383]. anomalies [1112]. anomalous [1958]. ANOVA [2181, 772, 2882]. anovulation [1422]. answer [1473]. antagonism [2703]. antedependence [192]. antenatal [1280, 2253]. anthracis [2597]. anthrax [1625]. anti [987, 570, 2973]. anti-bacterial [570]. anti-HIV [2973]. anti-infective [987]. antibiotic [2172, 1578]. antibody [2119, 937, 1190, 3239]. anticancer [1030, 2073]. anticoagulation [2217]. antidepressant [2756, 923, 2078, 218]. antidepressants [2757]. antigen [1111]. antigenic [1856]. antihypertensive [1815]. antimicrobial [1868, 1080]. antiretroviral [3104, 564, 3229, 1166, 2300, 1454, 2264]. antiretroviral-resistant [3104]. antivirals [3031]. any [1463]. anything [3152]. appendices [359]. Application [3323, 1769, 3104, 3020, 2567, 3273, 3238, 2740, 2610, 2608, 2979, 1975, 630, 2882, 1575, 657, 2707, 169, 994, 3068, 2778, 2435, 2746, 195, 2301, 1792, 645, 2000, 3359, 1064, 2725, 311, 875, 36, 100, 3212, 3289, 1641, 2172, 651, 158, 2181, 2525, 2363, 2766, 1539, 1869, 1691, 76, 755, 2041, 12, 1492, 119, 1151, 3070, 2048, 598, 2179, 2544, 1257, 1070, 1520, 2420, 970, 2017, 230, 2317, 2975, 3150, 652, 1558, 2587, 1320, 3007, 922, 194, 732, 209, 2951, 1004, 2777, 3304, 1822, 1604, 2371, 865, 1559, 1834, 1764, 21, 3037, 282, 1012, 1557, 1510, 2943, 2213, 940, 2308, 3034]. application [2883, 3167, 1893, 2201, 2545, 1480, 1016, 1028, 1454, 3094, 1165, 854, 2393, 1665, 2616, 3080, 2318, 1336, 2416, 2881, 753, 754, 874, 3114, 834, 1135, 1343, 2214, 2270, 1908, 2309, 422, 507, 2759, 2121, 2259, 901, 3353, 2396, 691, 2351, 1896, 628, 171, 1583, 981, 390, 734, 2007, 658, 3253, 3117, 2193, 2485, 1057, 862, 2132, 1931, 500, 1406, 2192, 1200, 1427, 486, 2904, 3240, 721, 2969, 2459, 489]. applicationin [1321]. Applications [1157, 601, 455, 1003, 129, 2659, 541, 3035, 221, 525, 2295, 3351, 2824, 244, 613, 2731, 3280, 1038, 199, 1000, 2011, 576, 2733, 1467, 48, 386, 38, 1990, 2835, 2920, 28, 2377, 1164, 1439, 118, 208, 2682, 861, 1153, 3232, 496, 2524, 685, 1824, 1142, 106, 2139, 1461, 1774]. applied [1355, 827, 731, 750, 1898, 1600, 728, 65, 629, 534, 1054, 696, 845, 1998, 3271, 2271, 2362, 1063, 3064, 1549, 1267, 819, 1425, 2006, 358, 3311, 1182, 182, 1595, 2502, 298, 400]. Applying [573, 3099]. Appraisal [540, 2771, 2775, 2772, 2773, 1863, 2774, 128]. Approach [434, 2110, 1943, 3236, 1916, 563, 1772, 559, 2321, 352, 1498, 2225, 1268, 2111, 509, 2313, 2431, 3079, 3186, 3143, 2255, 2800, 1103, 750, 1539, 3340, 3344, 3159, 87, 2984, 2802, 2312, 767, 1152, 3132, 2952, 3280, 204, 2288, 2065, 2426, 61, 1505, 3375, 2357, 1307, 1923, 1288, 2658, 766, 2315, 252, 262, 2242, 404, 2056, 3216, 294, 3217, 274, 1419, 853, 1835, 2602, 2042, 3177, 1925, 1796, 2631, 1895, 718, 1884, 2492, 3388, 1998, 2906, 2955, 2558, 2343, 1278, 2078, 1739, 2834, 719, 2994, 1887, 2798, 3363, 409, 1047, 1479, 3387, 52, 2887, 2603, 2580, 642, 1059, 1857]. approach [835, 3198, 2761, 2217, 1464, 591, 3054, 2896, 3138, 2198, 1105, 2668, 2675, 1644, 3381, 3214, 2844, 2799, 479, 643, 1763, 1435, 1722, 1837, 2594, 2296, 2971, 2807, 2429, 1410, 3052, 3288, 919, 445, 2310, 2191, 2852, 1854, 1490, 1988, 2790, 2729, 320, 2744, 2459, 2009, 713, 454, 1134, 1143, 2501, 1549]. Approach.Giovanni [1018]. Approaches [1971, 2416, 646, 259, 3066, 1752, 275, 733, 176, 644, 265, 1576, 1378, 2147, 1212, 555, 3094, 1176, 3281, 939, 2780, 1211, 1442, 2045, 313, 3267, 1254, 3358, 389, 1025, 2071, 1334, 604]. approaching [2378, 3077]. appropriate [1387, 3284, 1767, 466]. Appropriateness [2307, 3243, 1108]. approval [2657, 904]. Approximate [1088, 538, 1103, 813, 444, 213, 3145, 776]. approximating [2680]. approximation [2324, 3120, 2675]. approximations [47]. Arani [1925]. arbitrarily [1954, 3016]. arbitrary [1099]. Archimedean [2948]. Arcsine [2685, 3152]. area [551, 2166, 2638, 1536, 2641, 916, 2640, 2643, 3123, 1957, 23, 1095, 2920, 2817, 1940, 1096, 2636, 2637, 233, 1470, 1745, 2639, 237]. areas [1813, 3007, 3321, 2646, 2749, 1987, 1665, 2418, 721, 2645]. arising [732, 2428, 1063]. arithmetic [1561, 1174]. arm [2086, 3265, 2116, 2667, 1107, 2249, 976, 3385, 3362, 1034, 1135, 2896, 1426]. armed [355, 714, 16, 426, 354]. Armitage [127, 2842, 1118, 3296, 2075, 1534, 2129]. arms [976]. Armstrong [1424]. Arndt [2198]. Arnold [817, 454, 286, 361, 145, 288, 600, 602, 1753, 211, 436, 92, 453, 299, 962, 142, 357, 1193, 141, 128, 1825, 842, 186]. arrangement [631]. Array [1145, 3376]. arrests [2460]. ART [1931]. artefacts [230]. arterial [490]. arteriosclerosis [1125]. artery [1191, 565, 515, 1558, 994]. Artherosclerosis [3038]. arthritis [2316, 326]. Arthur [1802]. article [2476]. artificial [651, 767, 549, 63]. artificially [3252]. asbestos [1204, 2310]. ascertained [2290, 2381, 411, 809]. ascertainment [1771, 2050, 2996, 1867]. aseptic [1420]. Ash [39]. Ashwini [10]. asked [2305]. Asleep [2665, 2664]. Aspects [183, 362, 1067, 2576, 3388, 2798, 2109]. assault [3240]. Assay [1848, 1600, 1968, 595, 2404]. assays [2073, 1080, 1250]. assess [18, 249, 3125, 387, 274, 1822, 615, 2881, 1627, 3038, 2619, 1231]. Assessing [1690, 1035, 1747, 1108, 12, 2610, 413, 2420, 2658, 318, 2667, 247, 934, 1185, 3291, 1834, 2921, 1596, 864, 621, 799, 1034, 938, 2132, 2310, 2892, 2917, 1810, 173, 1886, 3186, 490, 2094, 3040, 2048, 3287, 292, 350, 732, 1626, 165, 1521, 1075, 2947, 2630, 2307, 271, 1723, 2084, 2778, 785, 2765, 1399, 3346, 1870, 1597, 1598]. Assessment [144, 360, 417, 718, 781, 1154, 2012, 544, 614, 3374, 1308, 2724, 160, 1986, 1929, 1882, 1478, 865, 1576, 2564, 2078, 2997, 3227, 2273, 806, 233, 2682, 2421, 3100, 3355, 1553, 171, 1315, 873, 2336, 326, 825, 489, 542]. assessments [1155, 27, 662, 2365]. Assigning [475]. assignment [2326]. assisted [2240, 601]. associate [1827, 55]. Associated [1456, 559, 1624, 1925, 2588, 1618, 1128]. association [2567, 3051, 3079, 1094, 3089, 3118, 1382, 1284, 726, 1767, 1339, 3287, 2958, 2058, 1522, 3335, 2124, 1292, 1544, 1560, 1285, 3343, 1689, 1266, 2536, 3179, 1887, 1797, 799, 2128, 1543, 2686, 3038, 2421, 1910, 1685, 2275, 1798, 1338, 2917, 1231, 3240, 3102, 1167, 1653, 2129, 2130, 2995]. associations [1044, 1642, 2994, 2182, 2269, 3358, 1440, 320]. assume [473]. assuming [1331]. assumption [1259, 1644, 919]. assumptions [772, 2590, 3334, 2033, 2615, 560, 1775, 2956, 2763]. asthma [1554, 750, 2011, 1259]. asymptomatic [801, 1053]. asymptotes [2378, 3077]. Asymptotic [19, 931, 1214, 1941]. atazanavir [2318]. Atkinson [3056, 397]. Atlas [239]. atmospheric [2597]. atrial [3238]. ATSDR [2346, 2347, 2912]. attacks [1624]. attained [1919]. attempts [475]. attitudes [3246]. Attributable [414, 2728, 3, 3336, 99, 2414, 2011, 112, 503, 1150, 1817, 2455, 1736, 2569]. attribute [2365]. attrition [3272, 2412, 3099]. atypical [2838]. auditory [1782]. augmentation [558, 1256, 2853]. aureus [1369]. Austin [3024, 2909, 2772, 2773, 2452, 2774]. Australian [1502]. Auswertung [494]. Author [2199, 398, 2453, 3373, 3245, 415, 1529, 2246, 2752, 1607, 814, 1710, 587, 959, 1017, 816, 2380, 2184, 1590, 3109, 3011, 1727, 858, 1323, 1325, 1670, 2135, 1683, 539, 2001, 758, 2821, 3220, 3107, 2840, 1055, 1697, 2137, 126, 804, 30, 1375, 1712, 388, 1170, 2477, 2490, 1938, 356, 1562, 1483, 2051, 2823, 715, 2677, 3332, 1515, 1646, 2003, 1281, 1800, 1785, 2185, 2451, 2818, 1564, 868, 2476]. autocorrelated [1999]. Automated [2529, 2216, 1370]. autoregressive [2508, 2378, 3077, 3078, 238, 2834, 2009]. auxiliary [1477, 2168, 2283]. availability [1786]. available [2441]. average [2990, 3273, 3159, 3364, 599, 2315, 2797, 3383, 270, 3274]. averaged [1059]. averaging [1947, 618, 1538]. avoidance [1377]. Avoiding [3084, 207]. axillary [1336]. Azzalini [130].

B [1364, 497, 1873, 3070, 3190, 3379, 2640, 2134, 1144, 436, 1533, 1925, 1054, 1531, 1595, 1156, 221, 958, 1363, 141, 2734, 2841, 103, 2160, 3221, 44, 1177, 2545, 412, 3087, 3031]. B-spline [1177, 2545]. Bacillus [2597]. Back [44, 1127, 45, 522, 767, 1914, 33, 2784, 1896, 2980]. Back-calculating [44]. back-calculation [45, 522, 33, 2784, 1896]. back-projection [1127]. back-propagation [767]. backwards [2865]. bacteria [1307]. bacterial [726, 570]. bad [1911]. Baddeley [1873]. Bagging [1276]. Bailer [439]. Balakrishnan [1786]. Balance [3320, 366, 2281, 1198, 1879, 2772, 3201]. Balanced [161, 2487]. balancing [1880]. Balding [1351]. band [1200]. bands [60, 2422]. bandwidth [826]. Bang [3136]. banking [861]. Barbara [1159]. Barnett [183, 1774]. barrier [25, 622]. Barry [438]. BARS [1861]. Bartholomew [361]. Based [130, 1916, 967, 1772, 1117, 1545, 2440, 909, 1424, 315, 470, 1374, 53, 157, 1554, 2894, 1051, 2871, 2876, 965, 2122, 3153, 750, 2423, 2439, 2768, 2624, 1907, 351, 1580, 1853, 1737, 1382, 2284, 1982, 176, 666, 1155, 2179, 2942, 933, 682, 3097, 918, 177, 1589, 1116, 1180, 2898, 2693, 265, 2056, 2085, 840, 2427, 2973, 2848, 353, 1757, 704, 1186, 2737, 1040, 2366, 1764, 692, 3164, 2461, 1804, 2156, 923, 1285, 451, 3261, 1286, 1032, 2749, 3298, 2126, 2541, 1987, 2591, 706, 1199, 277, 2630, 2307, 1007, 1671, 345, 3119, 735, 977, 1901, 1678, 1381, 2855, 3363, 1940, 1941]. based [2559, 1702, 2318, 973, 1262, 939, 321, 993, 2387, 2291, 2084, 2274, 2217, 2572, 1669, 655, 3138, 2974, 376, 2632, 3239, 1960, 762, 3115, 3355, 31, 1831, 2986, 1445, 2383, 2403, 1120, 1322, 1431, 424, 2561, 2145, 3173, 2649, 2708, 1555, 1315, 3092, 3269, 391, 2843, 1906, 1102, 1200, 1860, 2831, 3102, 2645, 1870, 2872, 1047, 1479, 1723, 233]. Basel [1872, 2961]. Baseline [330, 2614, 2588, 760, 1171, 1886, 1328, 3320, 1806, 1044, 2471, 2895, 2289, 2612, 1160, 1327, 1742, 3380, 3276, 660, 902, 2228, 1662]. baselines [1626]. Basic [842, 155, 156, 2824]. basis [651, 1333, 1380, 2657, 2721]. batches [2422]. Bayarri [1608]. Bayes [961, 72, 1964, 558, 294, 1269, 3264, 3387, 237, 961]. Bayesian [1317, 2440, 93, 2439, 1134, 1018, 3388, 3180, 1609, 2858, 1355, 3233, 3323, 1772, 315, 2174, 551, 827, 1498, 3351, 2111, 36, 100, 3352, 1771, 710, 709, 3206, 2801, 2791, 3259, 2864, 2800, 1539, 1869, 2131, 2116, 2795, 3340, 3159, 1907, 2802, 729, 2320, 2049, 2065, 120, 2329, 393, 2196, 1444, 2037, 3058, 2038, 3287, 1307, 2088, 1923, 2175, 2554, 2793, 2315, 2797, 248, 252, 612, 99, 558, 2576, 3015, 1909, 404, 3195, 2905, 1372, 2597, 1821, 3216, 1582, 3217, 217, 576, 2042, 3142, 1628, 2044, 1796, 2631, 246, 1895, 1965, 718, 1202, 2794, 1884, 845, 2484, 2492, 2985]. Bayesian [1422, 3343, 2100, 1904, 1095, 2201, 2564, 2545, 3196, 1278, 465, 910, 1680, 2949, 2395, 3101, 2950, 1887, 1820, 1089, 761, 1665, 2798, 2468, 3281, 392, 1932, 1744, 1799, 2377, 2603, 2580, 1043, 1614, 1857, 835, 1947, 1937, 2214, 2269, 2712, 2244, 2367, 2556, 2577, 2635, 2035, 2039, 491, 1367, 2668, 2796, 901, 2277, 2396, 3262, 3100, 372, 1254, 1818, 2799, 2036, 2863, 2435, 3064, 1896, 823, 2160, 1025, 2561, 1301, 1758, 1734, 1334, 373, 2007, 658, 1080, 618, 1538, 1602, 790, 3029, 1885, 1906, 1930, 2792, 2310, 3290, 2790, 3050, 1921, 2670, 3102, 3194, 2729, 3049, 339, 589]. Bayesian [2009, 1608, 3057, 1533, 302, 3347, 1752, 1517]. Bayesian-decision [404]. Bayesian-frequentist [2668]. Bayesianly [1770]. Bayesians [2664, 2665]. Bayesschen [343]. be [1146, 1386, 1855, 3380, 2676, 1232, 3276, 660, 1579, 3201, 2328, 3193, 788, 1046, 2862]. Bearden [1461]. because [1020]. beer [998]. before [1613, 1098]. before-after [1098]. Beginner [1156]. Beginning [2230, 469]. behavioral [3264]. behaviour [3131, 1239]. behavioural [2206, 770, 150, 908]. behaviours [1814, 2393]. behind [464]. Beispielen [197]. beliefs [382]. Bellavance [813]. Bellomo [300]. below [3149]. Belson [1794]. Benedetti [342]. beneficial [2940]. benefit [330, 980, 2631, 670, 1906, 484, 997]. benefits [451]. Benichou [414]. Bent [200]. Berger [1608, 302]. Bergsma [1544]. Berlin [40, 1548, 196]. Bernardo [1608, 302]. Bernoulli [2806, 1865]. Bernsen [416]. Berry [1118, 871]. Bertollini [417]. Best [1142, 740, 953, 3304, 739, 1237, 1376]. Beta [2870, 2321, 2805, 1882, 277]. Beta-binomial [2870, 2321, 2805, 1882, 277]. Beta-binomial/Poisson [2870]. better [256, 889]. Betty [1648]. between [998, 1209, 538, 2281, 2771, 2775, 3320, 2111, 1944, 1453, 1076, 552, 3238, 3128, 1738, 154, 3285, 1111, 3287, 2450, 2757, 3354, 1522, 2772, 3335, 2773, 1899, 732, 2951, 213, 2614, 752, 2464, 1313, 1689, 1617, 1480, 1901, 270, 592, 2551, 2167, 98, 3012, 3295, 3172, 1470, 1464, 3038, 765, 313, 2517, 3267, 2421, 2434, 2799, 1623, 2774, 329, 3230, 1440, 3306, 1865, 2260, 172, 2275, 1798, 873, 2917]. between-cluster [3230]. between-study [3238, 2551]. beyond [2638, 2641, 2640, 2643, 2636, 2637, 2639]. biallelic [1260]. Bias [487, 3131, 1009, 3030, 1733, 1968, 2237, 690, 1114, 1859, 862, 1615, 1649, 1126, 1970, 803, 1806, 2389, 2945, 697, 3083, 1862, 2189, 2800, 2937, 1291, 1086, 2050, 1279, 318, 2437, 2187, 2438, 3123, 927, 1704, 2112, 1604, 1173, 2100, 385, 2102, 441, 1728, 2551, 1867, 2685, 860, 2280, 2811, 1122, 394, 1312, 1705, 173, 1289, 3047, 3162, 1809]. Bias-corrected [1859, 1615]. Biased [2472, 17, 1778, 2282, 2427, 3255, 2452, 1717, 397]. biased-coin [17, 1778, 397]. Biases [23]. Biennial [2346, 2347]. bigger [889]. Biggeri [417]. Bihar [854]. bilateral [1796]. Binary [936, 832, 1321, 114, 1275, 717, 3155, 3272, 1272, 1615, 3041, 2481, 2540, 3351, 471, 2324, 1207, 13, 1869, 516, 3159, 3131, 1807, 756, 3342, 3280, 789, 2629, 2284, 2471, 2099, 393, 2781, 2008, 3208, 2079, 3312, 74, 2164, 665, 1060, 1656, 968, 1004, 534, 518, 1788, 3142, 1640, 2249, 800, 3116, 1476, 3188, 3305, 2352, 2061, 2466, 3385, 1616, 2947, 687, 2884, 2100, 512, 1761, 11, 2446, 2558, 345, 719, 7, 1357, 3098, 2368, 1395, 1171, 2528, 1799, 3059, 254, 2104, 1304, 912, 1614, 1857, 2388, 2562, 2173, 2593, 1211, 3114, 1667, 1603, 2685, 1859, 582, 1791, 860, 2280]. binary [1711, 1115, 2542, 3267, 2655, 1129, 2066, 1006, 667, 2863, 29, 2469, 3232, 1301, 3230, 2531, 3306, 2938, 325, 373, 371, 2467, 2069, 790, 632, 2741, 1450, 1002, 1147, 2706]. binless [2515]. binocular [2459]. binomial [3105, 3135, 2321, 2122, 2178, 931, 2805, 1487, 736, 2888, 2, 1811, 1882, 1582, 1341, 2804, 1049, 1936, 3298, 277, 504, 85, 2855, 2559, 3021, 1013, 815, 1818, 628, 2628, 1810, 2353, 535]. binomial/Poisson [2870]. bio [1624]. bio-terrorist [1624]. bioassay [836, 1694]. bioavailability [3203, 495, 1654]. bioavailability/bioequivalence [1654]. bioelectromagnetic [2509]. bioequivalance [584]. bioequivalence [1035, 269, 3357, 3203, 2314, 2984, 716, 978, 1036, 272, 2315, 2797, 262, 265, 273, 1185, 274, 267, 266, 495, 1654, 38, 271, 977, 270, 1702, 806, 263, 264, 268]. bioinformatics [3296]. biologic [2316]. biological [1802]. biologically [2394]. Biologist [1145]. Biology [439, 2043]. biology-progression [2043]. Biomarker [3104, 1600, 2600, 307, 2965, 136, 3154, 3237]. biomarkers [2327, 2802, 1155, 2987, 2366, 1581, 2742, 1395, 973, 68, 2573, 2649, 1915]. biomedical [1471, 69]. Biopharmacy [2570]. biopsy [2468, 326]. biostatistical [1217, 2279, 216, 1234, 2161, 1566, 1550, 781]. biostatistician [668]. Biostatisticians [2161, 2162]. Biostatistics [77, 181, 78, 881, 728, 1226, 1849, 3057, 2140, 2398, 2031, 1568, 1876, 2212, 127, 3180, 359]. biosurveillance [2529, 3333]. bioterrorism [2362]. Biplot [2759]. Biplots. [43]. Birkhauser [1872, 300]. birth [470, 561, 76, 1258, 1123, 1239, 2308, 1820, 818]. birth-and-death [2308]. births [1429]. birthweight [99]. Bishop [1351]. Bivariate [2180, 2317, 2055, 2254, 3127, 2814, 710, 3051, 3079, 564, 505, 1869, 2285, 3078, 652, 64, 3293, 3035, 2870, 2615, 1164, 2235, 3232, 1583, 2303]. BK [3383]. Black [603]. Blackwell [1118, 1648]. bladder [1320, 1260, 1343]. Bland [1180, 2687]. bleeding [66]. Blinded [1041, 3227, 682, 3097, 1589, 1246, 2102, 1731]. blindness [2002, 1365, 1645]. Block [940, 3097, 1704, 3227, 2726]. blocks [138]. Blomqvist [154]. blood [553, 1076, 251, 1536, 2203, 1000, 659, 1480, 1815, 1200]. bloodstream [2606]. BMD [1665]. BMJ [1159]. Board [1826]. Boca [1594, 1460, 2824, 1364, 3203, 1751, 1873, 362, 2004, 3311, 3312, 3057, 1205, 2399, 1533, 780, 3180, 2959, 821, 1609, 3347, 2960, 1840, 2858, 2005, 2139, 2719, 2295, 3122, 2841, 2034, 1610, 1504, 2720, 1839, 1244, 1503, 2755]. Bock [494]. Bodart [884]. body [1106, 1502]. Böhning [601, 3312, 417]. Boisson [2344]. bone [970, 1576, 508, 3035]. bone-marrow [970, 1576]. Bonett [2855, 2559]. Bonferroni [3153, 2986]. Bonferroni-based [3153, 2986]. Book [1530, 1594, 91, 341, 819, 817, 1157, 1460, 1647, 1317, 93, 1565, 454, 1592, 421, 542, 3139, 198, 509, 3222, 2138, 1752, 1786, 1351, 2754, 601, 144, 455, 360, 2824, 526, 2604, 885, 1425, 434, 1158, 1364, 497, 1608, 286, 3203, 782, 361, 1751, 143, 200, 127, 183, 1750, 1873, 301, 145, 402, 843, 362, 288, 2501, 3204, 1415, 1151, 2006, 129, 603, 1566, 1550, 90, 1134, 543, 343, 1402, 358, 600, 2004, 1119, 199, 399, 2856, 42, 1461, 420, 1438, 602, 258, 40, 201, 1787, 961, 3311, 303, 1245, 511, 1180, 1067, 1548, 1753, 131]. Book [3312, 1194, 1182, 541, 3057, 211, 1352, 452, 182, 566, 1018, 1195, 1205, 342, 844, 1774, 1802, 3182, 439, 196, 3181, 2399, 1310, 417, 1144, 1181, 436, 92, 438, 2409, 1533, 2503, 1311, 2140, 453, 1159, 1593, 1309, 104, 818, 435, 1133, 884, 495, 1142, 2857, 780, 130, 300, 302, 1118, 3180, 1531, 524, 299, 1595, 2959, 1156, 221, 1671, 821, 1609, 132, 3347, 962, 142, 1872, 1801, 3192, 359, 2960, 781, 357, 494, 1840, 1437, 525, 3056, 1193, 1363, 440, 285, 141, 1143, 1532, 1145, 2858, 544, 869, 2519, 872, 419, 1192, 2961, 43, 257, 1546]. Book [2005, 146, 1470, 106, 2109, 1471, 2139, 2502, 2859, 2719, 287, 2295, 870, 3122, 260, 1648, 545, 1516, 185, 2841, 3138, 3231, 437, 1206, 2345, 128, 2034, 298, 400, 418, 820, 1547, 103, 540, 1517, 1065, 1610, 510, 187, 1504, 2718, 284, 1384, 1825, 1698, 147, 105, 2248, 2720, 259, 184, 496, 197, 842, 1066, 186, 433, 210, 1684, 1839, 401, 1549, 1244, 1503, 1416, 871, 41, 94, 2755, 1169]. Books [1159]. Boos [3302]. boosting [2429]. Bootstrap [115, 492, 2514, 1526, 1852, 46, 2084, 3359, 2865, 311, 2138, 1033, 3015, 2602, 2697, 3298, 238, 1678, 2173, 1015, 2581, 263, 1153, 1263, 1666, 452]. Bootstrap-based [2084, 1678]. Bootstrap-type [492]. Borgdorff [416]. Borgelt [1416]. Boris [421, 2183]. born [1267]. Boston [300, 1066]. both [731, 1497, 116, 3000, 818, 3343, 2207, 2077]. boundaries [485]. boundary [1228]. bounded [914, 3094, 2221]. Bounds [2590, 3334, 3141, 2498, 2449, 3093, 3219, 2986, 3302]. bouts [2210]. Bouvier [1750]. bovine [1093]. Bowling [399]. Box [1558, 1000, 1821, 1501]. brackets [3350]. Bradford [305]. Bradley [961, 3347]. Brady [2841]. Brain [3005, 2048, 1298, 2837, 1299, 633, 1370]. branching [3361]. Branson [3379]. Braunstein [2787]. Brazil [551, 473]. BRCA [2357]. BRCAPRO [2954]. breached [2002, 1365, 1645]. break [1201]. break-points [1201]. breaking [2374]. breast [640, 1591, 651, 2915, 1830, 1151, 2261, 2357, 849, 922, 934, 996, 2954, 2493, 1089, 1336, 338, 2416, 2271, 321, 3114, 334, 3004, 1303, 2170, 981, 734, 1358, 3253, 1048, 1709]. breast-feeding [640]. breath [1430]. Brian [1751, 1067, 844, 357, 2109, 2719, 186, 1839, 2755]. bridge [1083]. Bridging [1623]. Briggs [1142]. Brillinger [129, 94]. Broemeling [3057]. bronchiolitis [679]. bronchodilation [953]. bronchodilator [953]. Brown [2245, 358]. Bruce [541, 1801, 2718]. Bruno [649]. BSE [362]. BUGS [3314, 3315, 3313, 3319, 3316, 3317, 3318]. building [2618, 3138]. Bulpitt [1066]. bump [2325]. Buncher [2345]. burden [2117, 2564]. Bureau [1219]. Burnham [434]. Burr [2222]. Burst [103]. Burton [2499]. Burwalls [738]. Business [2718]. bypass [1191, 515].

C [3135, 1649, 2327, 542, 198, 1786, 1351, 2820, 455, 2604, 1425, 1158, 3386, 1566, 2909, 3311, 2140, 3310, 1142, 2452, 3056, 43, 1648, 437, 815, 2345, 397, 1494, 433, 210, 1151, 3087]. c-erbB- [1151]. C. [1728]. C30 [667, 971]. CABG [1191, 1179]. Cabrera [1195]. Cai [3219]. calcification [1558]. calculating [44, 96, 930]. calculation [45, 522, 926, 33, 3028, 2784, 1835, 2869, 35, 2530, 1896, 2980, 3160, 2663]. calculations [730, 381, 2095, 19, 3259, 2093, 3159, 367, 2088, 1107, 2087, 1882, 518, 368, 1819, 2113, 2607, 1688, 3122, 255, 2661, 1098, 2403, 1045, 2221, 1410, 3140, 3248]. calculus [81, 80, 83, 82]. calibrating [3231]. calibration [1489, 125, 3174, 10, 2781, 3002, 6, 3076, 340, 3163, 324]. California [1667, 2915, 140, 477]. call [2162]. Cambridge [885, 1158, 843, 1133, 1671, 440, 1698, 259, 433]. Cameron [433]. Campbell [2751, 509]. Can [2895, 590, 2282, 473, 906, 2452, 1312]. Canadian [2201, 239, 172]. Cancer [1387, 603, 258, 922, 962, 1808, 437, 239, 3323, 1591, 529, 1051, 1771, 2474, 651, 2621, 1869, 3252, 1204, 2915, 1830, 12, 175, 204, 1738, 2926, 3210, 2261, 1101, 2781, 2357, 1894, 849, 219, 2056, 913, 2117, 934, 732, 2998, 2776, 2777, 576, 996, 1314, 980, 48, 386, 1973, 2213, 1032, 3388, 2308, 2493, 1531, 1832, 28, 477, 1318, 1176, 1089, 2798, 1260, 3308, 1336, 973, 338, 2416, 764, 2271, 805, 321, 2780, 3114, 334, 1135, 1343, 2244, 2778, 220, 3004, 1729, 691, 663, 1303, 2170, 981, 1440, 1758, 734, 2108, 1358, 943, 3253, 2310, 1426, 1427]. cancer [1048, 1921, 3337, 2729, 2889, 121, 1709, 489, 437, 603, 131, 401, 1151]. Cancer-specific [922]. cancers [2779, 922, 2954, 2700]. candidate [2553]. candidiasis [102]. Cannings [1351]. cannot [1232]. canonical [1012]. capsaicin [911]. capsaicin-induced [911]. capture [1746, 613, 915, 2920, 2305]. capture-recapture [1746, 613, 915, 2305]. capture/recapture [2920]. captured [729]. CAR [3228]. carcinogenesis [49, 1524]. carcinogenicity [1165]. carcinoma [3035, 486]. cardiac [1346, 2907, 2460, 2539, 2908, 3137]. cardiovascular [998, 899, 837, 3022]. care [1866, 1643, 1752, 751, 1833, 367, 328, 878, 2242, 1794, 704, 1369, 3037, 364, 498, 694, 972, 2722, 2365, 721, 493]. Careful [1812]. cares [899]. caries [3351, 2525, 3354]. Carlin [961, 1533, 3347]. Carlo [93, 2452, 2858, 2281, 2282, 228, 3260, 2906, 2294, 1091]. Carole [545]. Carpenter [3137, 2908]. carriage [726]. carrier [2954]. carriers [2357, 801, 1053]. carry [1514, 833, 616, 2089, 1149]. carry-forward [1514, 1149]. carry-over [833, 616, 2089]. Carstensen [2734]. Cary [3139, 2754, 2604, 782, 3204, 2409]. Case [982, 414, 3384, 3111, 1560, 2704, 1143, 268, 3323, 3178, 1601, 1643, 2060, 1652, 1970, 205, 3, 975, 2093, 2293, 1319, 2116, 1844, 2815, 1784, 3161, 1191, 79, 1456, 1009, 1151, 1443, 1767, 2897, 674, 2241, 554, 2756, 2898, 2693, 1182, 1883, 913, 1899, 732, 2514, 1604, 2371, 1993, 980, 2349, 774, 3108, 336, 475, 2803, 1261, 1293, 22, 3343, 112, 1150, 1817, 548, 2835, 2207, 2950, 2384, 2456, 2092, 1587, 1420, 1262, 974, 1614, 98, 233, 1897, 1211, 1937, 1959, 2128, 3054, 2022, 861, 1910, 2066, 1511, 2620, 3106, 1189, 1734, 618, 607, 1948, 2031, 1399, 2917, 2827, 1653, 2129]. case [2130, 2995, 327, 489, 1570]. Case-cohort [2704, 1319, 1262]. Case-Control [414, 982, 3111, 3323, 1601, 1652, 1970, 3, 2093, 1844, 1784, 1151, 1767, 674, 2898, 2693, 913, 732, 1993, 2349, 774, 475, 2803, 1261, 22, 3343, 112, 1150, 1817, 2207, 2384, 2456, 1587, 1614, 98, 1897, 1211, 1937, 1959, 2128, 3054, 2022, 861, 1910, 1511, 3106, 1734, 618, 2827, 1653, 2129, 2130, 2995, 327, 489, 1570]. case-crossover [1604, 2835, 607]. case-deletion [1948]. case-only [1456]. Case-Oriented [1143]. case-series [1899]. Case-sibling [1560]. case-study [2060]. cases [982, 1746, 736, 177, 2241, 2998, 165, 1900, 2867, 807, 3064]. CASI [1291]. casualty [2360]. catchability [2305]. Categorical [402, 579, 3222, 1600, 256, 1768, 1492, 969, 778, 248, 1863, 395, 533, 1634, 84, 1665, 3295, 3325, 3262, 1684, 685, 3158]. categories [1078, 1633, 2526]. categorized [1361, 3342]. Categorizing [28, 936, 1010]. category [2698, 394]. Catherine [962]. Causal [1060, 3341, 2069, 2894, 3212, 3289, 2332, 2590, 3334, 2740, 796, 921, 3297, 3093, 2840, 2709, 1495, 2354, 468, 2232, 3219, 1759, 1816, 1315, 2839, 3266]. Causality [843]. Causation [1133]. cause [1082, 2606, 651, 175, 2696, 1267]. cause-of-death [1082]. cause-specific [2606, 651, 175, 2696, 1267]. caused [2597]. causes [175, 784, 2799]. caution [74]. cautionary [2374, 168]. CD [301, 3204, 3180, 1671]. CD-ROM [301]. CD4 [796, 773, 624, 650, 1583]. CDC [2346, 2347, 2912]. cDNA [1975]. celebration [2263]. cell [2606, 1632, 2116, 666, 3256, 2072, 624, 1583, 2978, 1427, 2404]. cells [631, 3152]. cellular [2598]. Censored [1352, 1302, 1848, 53, 941, 1554, 689, 965, 3051, 3079, 2287, 3022, 49, 1636, 2592, 351, 1580, 1737, 2952, 3016, 597, 2276, 2608, 1986, 396, 1522, 2029, 1001, 1359, 1432, 3144, 1558, 2206, 1372, 2283, 2951, 2861, 307, 1526, 549, 865, 956, 2627, 67, 1735, 1509, 506, 2013, 724, 2239, 1893, 595, 333, 1605, 2997, 683, 3023, 250, 20, 2615, 2458, 834, 507, 2853, 449, 1128, 2533, 1527, 2071, 1250, 1588, 2192, 957, 2670, 1400, 2549, 2252]. censoring [2874, 3229, 1215, 3129, 1433, 549, 1540, 2202, 1829, 422, 335, 2561, 1583, 2406]. censorings [703]. Census [600, 1219, 2920]. cent [2919, 3364, 1109, 617, 3383, 3221]. Center [2350, 155]. centers [2614]. centile [1501, 3299]. central [668]. centralized [161]. centre [114, 1879, 2099, 353, 1884, 910, 2613, 346, 993, 486]. centres [1242, 1329]. certain [1435]. cervical [1141]. Cessation [2193, 3273]. CGH [3376]. Chain [93, 2858, 1197, 2695, 2624, 1768, 228, 3260, 2291, 1091]. chains [2435]. Challenge [2312, 2736, 911, 1230]. challenge-escalation [911]. challenges [157, 1284, 1624, 2659, 654, 469]. challenging [108]. Chambers [419]. chance [220, 1698]. Chang [2960]. Change [2228, 1886, 1290, 1044, 1853, 3364, 3040, 154, 2992, 1520, 2450, 152, 612, 163, 652, 3335, 1014, 3380, 1452, 724, 2906, 3276, 1595, 1028, 150, 3383, 743, 1050, 1976, 1442, 422, 712, 742, 653, 2260]. change-point [2906, 1050, 1976]. changepoint [2376]. changes [2977, 2878, 2598, 2176, 3224, 1090, 167, 3004, 2746, 1555, 1315]. changing [1463, 2301, 888]. Chapman [1594, 341, 1460, 93, 198, 601, 2824, 1425, 1364, 3203, 1751, 143, 200, 1873, 362, 90, 2004, 1119, 42, 961, 3311, 3312, 541, 3057, 1205, 342, 844, 439, 2399, 1533, 104, 780, 130, 3180, 2959, 821, 1609, 3347, 2960, 1840, 1437, 2858, 43, 2005, 2139, 2719, 2295, 3122, 2841, 2345, 2034, 1547, 1610, 1504, 105, 2720, 184, 1839, 1503, 2755]. characteristic [3139, 24, 1682, 3007, 3123, 1707, 1446, 807, 762, 763, 2408]. characteristics [3335, 2156, 2588, 3038, 785, 1885]. characterization [2306]. characterize [1642, 3189]. Characterizing [1166, 2505, 2597, 2516]. Charles [1647, 1158, 2140]. CHARM [1885]. charts [1996]. Chaudhary [2959]. check [3126]. Checking [2030, 3008, 1601, 2609, 596, 2714, 2383]. checks [1088, 235]. Chemical [545]. chemicals [1779]. chemotherapy [970, 805]. Chen [538, 1517]. Cheng [211]. Chernick [452]. Cheung [2440, 2439]. Chevret [2344]. CHF [2961]. Chi [2488, 211, 1715, 1714, 534, 2751]. Chi-Lun [211]. chi-square [534]. Chi-squared [2488, 2751]. Chichester [509, 1752, 1351, 360, 127, 1402, 358, 303, 1194, 452, 1018, 1774, 3182, 3181, 417, 1144, 1311, 1309, 544, 287, 260, 3138, 3231, 1206, 418, 820, 147, 1416]. chickenpox [119]. child [1638, 640, 2694, 1919, 639, 2969]. childbearing [2253]. Childbirth [818]. Childhood [437, 2000, 473, 3161, 251, 1536, 21, 1239, 477, 233, 2700, 1555, 633]. Children [1667, 982, 2120, 1430, 1332, 432, 2070, 2413, 1480, 3355, 691]. Chinn [420]. CHMP [2020, 2021, 2245, 2019]. Cho [1801]. Choice [935, 64, 2536, 2322, 118, 1573, 2669, 1057, 2020, 2245, 1584, 2019, 1964, 836, 955, 3014, 1744, 1376, 3193, 1759, 3190]. choices [2456, 3308]. Choosing [1984, 3213, 3354, 3267, 3386, 3310]. choropleth [244]. Chow [2002, 3203, 1645, 201, 495, 2960, 287, 3122, 1644]. Chris [402]. Christian [1416]. Christl [362]. Christopher [541, 1563, 1066]. Christopoulos [1802]. Christy [2754, 2409]. chromosomal [1112]. Chronic [1535, 158, 2610, 99, 2448, 934, 293, 3266]. Chuang [2754, 2409].

Chuang-Stein [2754, 2409]. Chung [2002, 3203, 1645, 201, 495, 1056, 2960, 287, 3122]. cigarette [2900]. circadian [2195]. circular [2252]. cirrhosis [66]. Cities [259]. City [633]. cladistic [3043]. Clara [140]. Clare [600]. Clarence [3136]. Clarendon [187]. Clark [1594, 1133, 1054]. Clarke [92, 1471, 842]. Class [455, 3178, 1923, 766, 3354, 2522, 2600, 136, 3044, 701, 2950, 1537, 2388, 3068, 384, 2759, 2956, 3242]. classes [1166]. Classic [103]. Classical [2430, 299, 2534, 1334, 1593]. Classics [418]. Classification [421, 360, 3376, 1082, 223, 2953, 65, 1787, 2242, 1297, 2602, 940, 2188, 1877, 2074, 2027, 837, 63]. classifications [1793]. classifier [1818]. classifiers [2307]. classifying [2721]. Claude [884]. Clegg [3383]. Cliff [259]. Clinical [1568, 1876, 2212, 144, 2570, 887, 1402, 1438, 602, 1067, 541, 770, 1311, 1309, 569, 1192, 287, 2295, 1233, 418, 540, 1065, 1066, 896, 717, 967, 1209, 2002, 330, 1778, 1706, 2331, 529, 527, 2540, 2158, 2297, 1918, 2738, 355, 137, 1752, 2313, 275, 2332, 379, 3183, 895, 975, 1942, 2791, 3259, 750, 382, 2665, 1487, 1696, 3340, 2454, 3065, 108, 1365, 767, 1108, 1235, 3282, 831, 832, 833, 79, 789, 585, 2610, 728, 1146, 2065, 2501, 2402, 2875, 1188, 2724, 2148, 2329, 2586, 2026, 3327, 894, 933, 900, 161, 2521, 2856, 5, 2895, 2088, 2793, 1097]. clinical [1929, 1645, 2033, 1300, 485, 181, 1099, 303, 2756, 2576, 1042, 404, 665, 95, 2056, 3170, 1880, 898, 968, 2479, 605, 2848, 3226, 307, 212, 1715, 2777, 576, 1040, 996, 2409, 297, 1423, 1314, 1341, 2124, 3380, 1521, 2631, 3264, 3164, 202, 1777, 81, 1714, 2113, 336, 2392, 573, 1264, 1389, 1983, 976, 718, 406, 1247, 626, 1451, 3328, 2985, 867, 1805, 1990, 2601, 3276, 1789, 2110, 512, 3329, 2558, 3136, 1611, 465, 910, 577, 345, 428, 1688, 2707, 719, 1028, 407, 757, 1563, 442, 441, 555, 2960, 1265, 2368, 727, 80, 83, 2664, 189, 416, 1248, 3023, 2468, 1851]. clinical [1612, 392, 82, 2578, 2580, 1462, 2553, 1043, 974, 642, 1034, 1741, 902, 805, 346, 16, 1546, 2689, 253, 1756, 993, 2109, 882, 3122, 1135, 1464, 2712, 1208, 582, 2581, 1380, 422, 89, 1569, 1871, 3332, 3326, 1367, 798, 2668, 514, 1485, 1967, 1716, 1713, 2675, 1775, 2421, 2911, 2480, 705, 1021, 3067, 3193, 797, 663, 556, 1401, 3087, 354, 218, 823, 2160, 1754, 109, 572, 107, 424, 153, 1440, 1758, 2401, 2248, 2720, 1841, 2628, 2145, 2971, 325, 669, 658, 2219, 1664, 1599, 2265, 670, 3092, 513, 889, 2829, 3029, 1022, 604, 344, 759, 1020, 2792, 2741, 484]. clinical [997, 1730, 1881, 888, 3331, 1397, 337, 1184, 2333, 425, 1200, 1426, 1427, 2831, 486, 1, 2353, 3194, 3337, 1289, 3330, 342, 1159, 1531]. clinically [3081, 2743, 2089]. clinician [149, 55]. Clinicians [440, 155, 3349, 156, 1158]. clinics [2253]. Clogg [191]. clonal [1524]. closed [207, 2278, 2986]. closed-form [2278]. closer [1349, 504]. closing [483]. closure [3115]. cloth [1461]. Cluster [817, 1364, 369, 1957, 234, 2689, 3278, 1792, 2481, 1163, 1862, 2231, 3265, 1204, 3157, 969, 907, 375, 699, 1964, 1969, 3028, 1882, 3170, 2998, 374, 3027, 368, 654, 1452, 769, 1965, 1953, 1342, 2100, 656, 277, 2091, 364, 1171, 3059, 254, 3307, 2373, 370, 1059, 366, 220, 372, 1961, 2523, 1301, 3230, 373, 1366, 2007, 2250, 371, 964, 1494, 2467, 2999, 2671, 1147, 1963, 365, 2410]. cluster-level [1163]. cluster-randomization [2100]. cluster-randomized [1792, 3157, 3059]. cluster-unit [1059]. clustered [717, 1554, 279, 2082, 1492, 992, 1148, 3086, 1213, 1300, 1448, 1656, 1433, 2717, 1926, 534, 518, 395, 1640, 1113, 802, 3268, 3225, 1360, 3350, 1823, 1547, 1129, 1486, 2090, 1931, 1184, 3382]. Clustering [2707, 1641, 599, 140, 2998, 735, 3307, 2373, 1959, 2635, 34, 981, 1962, 3275, 2936, 421]. clusters [1954, 472, 3265, 808, 2828, 2838, 3119, 2208, 2933, 1823, 2565, 1962, 365]. CML [700]. Cnattingius [1711]. co [2312, 1682, 1446, 2254]. co-aggregation [2254]. co-ordinates [1682, 1446]. co-primary [2312]. coarse [3292]. coarsened [2385, 3046]. cocaine [2204]. Cochran [3061, 2129]. Cochrane [891]. code [19, 28]. coding [2029, 588, 2517, 2620]. Cody [3204]. coefficient [3048, 1103, 1898, 1161, 617, 1969, 2557, 533, 2444, 595, 2854, 1435, 1845, 1760, 2054, 1366, 2007, 964, 1263, 1494, 2845]. coefficients [2865, 369, 1975, 1821, 2717, 830, 1895, 2208, 944, 3172, 2173, 2274, 3127, 1722, 722, 394, 1048]. cognitive [2120, 1291, 3040, 152, 163, 652, 1822, 1299, 150, 2258, 171, 151, 3017]. Cohen [129, 75, 2441]. coherence [2512]. cohort [2030, 3020, 2436, 2735, 1319, 3131, 1525, 3184, 2731, 430, 1767, 152, 1011, 2470, 3335, 1899, 1971, 2776, 2541, 166, 2704, 1820, 1262, 334, 2309, 2244, 2808, 1573, 2734]. cohorts [2414]. coin [17, 1778, 397]. cokeworks [1508]. Collaboration [891, 1622]. collapsed [225]. collected [3059, 2305]. Collection [639]. College [1067, 2857, 2109]. Colombia [972]. colon [1204, 1048]. colonized [1369]. Colorado [3020]. colorectal [2621, 981, 1921]. colour [1399]. Colton [127]. Combination [1506, 18, 3277, 2847, 2658, 387, 212, 3115, 1120, 2057, 324]. combinations [1581, 2394, 2214, 2974]. combine [3386, 3310, 3213]. Combined [906, 2313, 3257, 1085, 1362, 2310, 2390]. Combining [1064, 3118, 251, 726, 3060, 619, 2682, 3038, 2359, 1397, 1988, 77, 1453, 1108, 831, 832, 833, 864, 504, 14, 2881, 2760, 3231, 2370, 1910, 78, 1553, 2722]. Comment [1561, 3070, 3383, 2770, 2817, 1322, 3302, 2327, 928, 3150, 4, 857]. Commentaries [3373, 3332, 3319, 83]. Commentary [3368, 3327, 1597, 3369, 2772, 3328, 3329, 82, 3370, 794, 3371, 3372, 3331, 3330, 2151, 81, 1598, 80]. Comments [3041, 2751, 3105, 3135, 2440, 2820, 2907, 2665, 3386, 2638, 2641, 3314, 2626, 3190, 3379, 3315, 2909, 2640, 2643, 2772, 2773, 2653, 3108, 3310, 2642, 2676, 2489, 2840, 2826, 3385, 3388, 3136, 2819, 2855, 2559, 2644, 2637, 1726, 3316, 2734, 3219, 2787, 3106, 3317, 3318, 3137, 2639, 2476, 3382, 2839, 265, 2908]. Committee [1328, 2020, 2021, 1385, 892, 1475, 1387]. Committees [2021, 1402, 457, 2245, 2019, 1388, 571, 1386, 2346]. Common [305, 3234, 2282, 1371, 2452, 2786, 1488, 1845, 1760]. commonsense [509]. Community [157, 72, 1554, 706, 1129, 2412]. Community-based [157, 1554]. Companion [128]. company [265]. comparable [980]. Comparative [6, 3073, 730, 3105, 2888, 2, 2218, 2479, 1368, 2366, 234, 1495, 1688, 1023, 815, 1946, 2506, 69, 125, 10]. compare [429, 1988]. compared [2619]. Comparing [2225, 2814, 205, 70, 2311, 1111, 2872, 2494, 194, 1524, 1342, 1654, 2256, 2078, 1212, 727, 1248, 942, 1096, 1115, 2662, 3064, 218, 2807, 227, 114, 1275, 1639, 3320, 1813, 731, 279, 1743, 1411, 1883, 1062, 1269, 1085, 85, 1362, 2227, 2832, 1901, 1395, 1442, 1528, 1644, 2032, 1675, 919, 2057, 1057, 3158, 1450, 1404]. Comparison [969, 176, 61, 292, 1011, 1656, 2188, 2273, 2528, 1262, 2045, 2516, 3115, 2810, 650, 1890, 2510, 1334, 2999, 1253, 1538, 1948, 2067, 1022, 2512, 2995, 826, 3233, 3178, 3041, 1179, 2060, 2281, 2432, 2481, 355, 803, 2234, 1578, 403, 813, 1183, 1316, 444, 2382, 627, 1679, 3002, 664, 3208, 2267, 970, 779, 2898, 2397, 634, 1297, 140, 1707, 2272, 791, 3129, 983, 2602, 1850, 1788, 3037, 2392, 2428, 2415, 1032, 626, 1936, 112, 1780, 3044, 1266, 385, 958, 2634, 555, 3094, 773, 1176, 2368, 2424, 1927, 9, 409, 608, 1408, 3308, 16, 2572, 2713, 313, 1967, 3267, 2370, 2655, 2025]. comparison [1006, 110, 3358, 389, 354, 1754, 2523, 2722, 1189, 1210, 2250, 964, 1494, 647, 3031, 1020, 2741, 824, 2790, 721, 2645, 589, 1563]. Comparisons [782, 446, 3012, 1197, 198, 1037, 49, 1600, 1497, 3238, 290, 1337, 2698, 2532, 688, 1056, 1506, 2491, 846, 555, 2456, 1725, 2316, 2377, 1723, 902, 2027, 1865, 2971, 2485, 391, 1026, 50]. compartment [2425]. Compatible [2986, 1349]. Competing [644, 882, 389, 645, 3081, 2344, 2606, 3166, 2311, 755, 2288, 2426, 1767, 1719, 2058, 292, 3217, 2371, 1576, 2223, 1240, 1593, 1973, 1519, 2489, 2298, 2607, 2940, 240, 1047, 2183, 930, 2323, 2398, 2047, 672, 643, 838, 3082, 2476]. competing-risks [1240, 1047]. competitors [446]. complementary [109]. Complete [685, 1009, 1832, 2700]. complete-case [1009]. completely [1342]. complex [1290, 2332, 2563, 2276, 1794, 2414, 1493, 822, 1473, 2837, 2996, 2455, 2182, 1499, 2622, 1254]. compliance [2297, 3032, 1207, 2449, 25, 1883, 1060, 921, 1551, 1451, 2588, 656, 760, 2446, 2558, 2630, 3136, 719, 3283, 189, 416, 686, 2710, 255, 582, 1985, 2911]. complications [2556]. component [332, 2891, 272, 2420, 2566, 940, 1748, 1824, 2706]. Components [105, 1842, 619, 2562]. Composite [3185, 2583, 2562]. compound [2522, 3268]. comprehensive [2213]. Computation [2708, 1467, 406, 759, 3266]. Computational [580, 1629, 285, 1181]. computations [1982]. compute [429]. computed [391]. Computer [2363, 19, 2780, 1594, 601]. Computer-aided [2363, 1594]. Computer-Assisted [601]. Computerized [2413, 3114]. Computing [159, 2695, 400, 2343, 762]. concentration [1076, 1654]. concentrations [2815, 1200]. concept [1732, 3116, 2596]. conception [2240]. Concepts [2537, 2295, 2787, 984, 1979, 985, 2788, 1872]. conceptual [2914]. concern [3090]. concerning [563, 585, 62, 1569]. conclusions [2069]. Concomitant [1694, 1431, 2845]. Concomitant-based [1431]. Concordance [1898, 1693, 329, 809, 533, 2444, 2932]. concordances [2290, 2381]. Conditional [2113, 1540, 3154, 2169, 2125, 580, 3128, 575, 1101, 2781, 558, 64, 2812, 1640, 1804, 2156, 2803, 3322, 84, 2591, 238, 2989, 3363, 3242, 2810, 3106, 3215]. conditional-logistic [558]. conditionally [2950]. conditioned [244].

Conditioning [2282, 2452]. conditions [2513, 275, 1853, 2688, 519, 2770, 1096]. Conduct [1311, 889]. conducting [364]. conduction [2512]. Confidence [2122, 2292, 1109, 1116, 1405, 2428, 1990, 277, 7, 2384, 1175, 1940, 1941, 3021, 3172, 1013, 1322, 1864, 2934, 2866, 2233, 1946, 1639, 1168, 2865, 2919, 2498, 2178, 1031, 115, 1487, 736, 2386, 755, 3112, 1033, 617, 1359, 492, 274, 474, 60, 1526, 1850, 3053, 3298, 2422, 112, 2464, 2591, 2611, 2455, 9, 1434, 46, 2167, 806, 3325, 2896, 1233, 920, 2032, 3381, 479, 2986, 3145, 776, 1845, 3302, 964, 1263, 1494, 2708, 3052, 3288, 1102, 1200, 1810, 1619, 2441, 2744, 1117, 1545, 13, 11, 1669, 29, 3221, 2855, 2559]. confidentiality [463, 894]. Confirmatory [3206, 3183, 572]. Conflicts [1389]. conformity [811]. confounded [3141]. confounder [1203, 3219]. confounders [3184, 3093, 23, 1150]. confounding [1361, 1572, 1331, 2590, 2733, 1718, 2395, 2251, 2215, 2190]. congestive [432]. conjunctivitis [2041]. Connections [270]. cons [137]. consensus [640]. consent [867, 1]. Consequences [1203, 1428, 2266]. conservative [3105, 2888]. conservativeness [2896]. consider [1328]. consideration [1715, 1966, 1714, 1716, 1713]. considerations [1652, 269, 574, 2576, 273, 81, 80, 83, 694, 82, 370, 1013, 570, 57]. considering [3264]. consistency [1197, 2897, 3108]. consistent [2532, 2495]. CONSORT [375]. constant [2903, 2422, 722, 3017]. constant-width [2422]. constrained [2278, 3148, 2726, 1586]. constraints [2584, 2819]. Constructing [2178, 2589, 2007]. Construction [1919, 1260, 1343, 1380, 662, 2744, 3153, 1000, 1474, 695, 3172, 1864, 3381, 360]. consultant [1221, 668]. consultants [985]. consulting [2279, 216, 1195]. consumer [2365]. consumption [554]. contact [119, 840, 242]. containing [2802]. contamination [1862]. context [1747, 2094, 639, 3075, 1927]. Continents [131]. contingency [3135, 2172, 1630, 778, 395, 2804, 3053, 1266, 3042]. continual [576, 406, 675, 3308, 759, 3033, 589]. continuation [322]. continued [1979]. continuing [579]. continuity [704, 1377]. Continuous [1907, 831, 2587, 936, 2257, 810, 139, 1209, 1361, 2225, 2871, 2876, 1003, 1674, 1944, 2891, 1939, 1097, 485, 1992, 546, 2777, 3111, 533, 2415, 3167, 606, 1418, 1537, 1933, 3147, 2613, 677, 2760, 938, 188, 1464, 1911, 2618, 3138, 3262, 762, 372, 556, 2523, 3232, 2408, 3269, 632, 2497, 2830, 1619, 3006, 3218, 2009, 2439, 2440]. continuous-scale [1933, 1619, 3006]. contraception [641]. contraceptive [645, 755, 622]. contrast [2265]. Contrasting [465]. contrasts [931, 1188, 1557, 97]. contribution [1610]. contributions [1220, 1223, 1222, 1221]. contributors [947]. Control [414, 2190, 3323, 139, 1601, 982, 1652, 1970, 2575, 3, 2093, 2924, 1844, 1497, 1980, 1784, 256, 1151, 1038, 1767, 674, 2898, 2693, 913, 732, 1993, 3111, 2673, 2349, 96, 774, 475, 2532, 2803, 1190, 1261, 22, 3343, 112, 1150, 1817, 2207, 771, 2384, 2456, 3362, 1587, 1614, 98, 1897, 1211, 1937, 1959, 2128, 3054, 1259, 2896, 2022, 861, 2655, 1910, 1511, 3106, 1758, 1734, 618, 3158, 2827, 1653, 2129, 2130, 2995, 327, 489, 1570, 3041, 2530]. Controlled [602, 1309, 1066, 355, 2563, 792, 2724, 2164, 665, 1107, 1880, 791, 988, 374, 1718, 380, 2102, 657, 2092, 16, 2159, 1485, 1063, 354, 2720, 647, 513, 989, 2031]. Controlling [1508, 79, 1403, 1019, 2656, 2543, 2929, 1765]. controls [982, 2113, 475, 3064, 2333]. Controversies [1569, 579, 571, 894]. controversy [2517]. convergence [228]. converting [296]. convolution [1781]. Cook [524, 400]. Cooke [842]. Cooperative [1387]. Coordinating [155]. Coping [750, 3251]. CoPlot [2785]. copula [2948, 3079, 923, 3080, 2080]. copula-based [923]. copy [590]. core [352]. corneal [631, 1796]. Cornfield [1231]. coronary [1541, 1094, 1191, 565, 515, 1558, 2616, 994, 2746]. Correa [813]. correct [1806, 2189, 2898, 3047, 3162]. Corrected [2762, 1615, 1859]. Correcting [3174, 719, 2024, 632, 2437, 2187, 2438, 1599, 1315, 2558]. Correction [2478, 1472, 2877, 1650, 2389, 2860, 1803, 3223, 2500, 3342, 123, 2981, 3026, 3151, 1484, 2651, 2962, 1414, 2142, 386, 2753, 883, 588, 124, 1874, 2928, 3071, 2141, 2963, 2560, 1350, 2982, 2098, 1978, 1494, 122, 1952, 2585, 2391, 2964, 1970, 2945, 3202, 3229, 2789, 2274, 2879, 1114, 1829, 1105, 1717]. Corrections [1132, 1131, 1235, 1377]. correlate [937]. Correlated [129, 1717, 3232, 471, 1163, 3202, 13, 473, 2284, 2779, 3375, 3205, 74, 3293, 2042, 1796, 1819, 2352, 2464, 2493, 1085, 2372, 512, 11, 2959, 7, 1362, 596, 2430, 2689, 2235, 2274, 2879, 2555, 1783, 2696, 2103, 29, 1189, 2885, 2193, 1267, 2023, 2645]. correlates [2010]. correlating [2132]. Correlation [2694, 595, 2524, 2702, 349, 1346, 369, 1103, 332, 1898, 1076, 2082, 3157, 2276, 617, 1969, 2693, 3146, 533, 2444, 1012, 2515, 3019, 1693, 656, 3055, 3005, 3172, 2622, 2173, 2274, 1859, 2778, 1512, 3267, 1435, 1845, 1722, 1366, 2007, 964, 1263, 1494, 2904]. correlations [2702, 2225, 768, 1339, 1214, 3039, 2090, 3240]. correlative [2306]. Correspondence [3024, 2029]. cortex [1782, 2084]. Cortical [2508]. cosine [2195]. cosinor [1212]. Cost [2621, 1741, 620, 1179, 311, 926, 2638, 281, 2049, 699, 2643, 1042, 1372, 474, 2015, 2631, 451, 1052, 3167, 1373, 392, 2887, 2458, 2714, 1254, 2533, 1588, 1490, 2549]. cost-benefit [2631]. cost-effective [2638, 2643]. Cost-effectiveness [2621, 1741, 620, 926, 281, 2049, 1372, 474, 2015, 451, 392, 2458, 1254, 1588]. cost-related [1490]. costs [899, 1833, 2176, 1821, 1526, 2239, 2419, 493]. coughing [911]. Count [433, 41, 1769, 2575, 777, 2599, 708, 3354, 796, 2044, 1481, 2308, 2343, 3268, 1999, 2547, 1394, 1783, 1583, 2807, 2023]. counterfactuals [3374]. counties [461, 477, 981]. counting [1632]. countries [895, 1176, 334, 897, 896, 3299, 603]. Country [638, 1795]. Counts [818, 2482, 1073, 2870, 773, 3307, 2236, 624, 2900, 1963]. County [140]. Cours [40]. Course [842, 2738, 1235, 1040, 1463, 1379]. courses [881, 107]. Courtroom [543]. Covariance [681, 2706, 3236, 1087, 627, 331, 1012, 180, 958, 560, 3039, 2228, 1063, 109, 572]. covariance-adjusted [1012]. covariances [2680]. Covariate [1451, 506, 3098, 2434, 1740, 2971, 2440, 17, 2540, 3324, 1163, 249, 3186, 2449, 1696, 3210, 3002, 2289, 2899, 3075, 1160, 1001, 612, 3216, 3177, 3380, 202, 1172, 2460, 1889, 1509, 1247, 3276, 757, 1605, 2091, 2118, 856, 902, 188, 3201, 397, 390, 1527, 1045, 1657, 3247, 121, 661]. Covariate-adjusted [1740]. Covariates [2386, 2382, 2257, 1209, 1328, 3320, 2750, 2293, 2439, 1907, 3334, 1814, 2952, 935, 1738, 900, 2267, 2289, 1888, 248, 1327, 2568, 2973, 3045, 3244, 3037, 2794, 1261, 955, 2028, 3196, 2104, 930, 677, 2593, 2240, 1464, 1259, 1829, 1523, 507, 2701, 1783, 2810, 3099, 2469, 1734, 1865, 2708, 723, 2069, 500, 1406, 2462, 3110, 2762]. Covariates-dependent [2386]. Coverage [2518, 1631, 2178, 464, 2913, 2100, 2920, 2350, 2925, 479]. Coverage-adjusted [2518]. Cox [2052, 526, 42, 780, 821, 3106, 1498, 1720, 1051, 2871, 2876, 249, 3125, 2952, 2285, 2494, 1558, 1000, 1821, 2803, 1261, 2417, 2256, 2906, 1248, 1409, 1501, 3274, 1890, 1573, 1410, 1948, 3247, 1931, 2670, 661, 2133]. Coxian [2419]. CPMP [1328, 570]. Cramér [218]. Crans [3105]. CRC [1594, 341, 1460, 601, 2824, 1425, 1364, 3203, 1751, 1873, 362, 2004, 1119, 961, 3311, 3312, 541, 3057, 1205, 844, 2399, 1533, 1593, 780, 3180, 2959, 821, 1609, 3347, 2960, 1840, 1437, 2858, 869, 2005, 2139, 2719, 2295, 3122, 545, 2841, 2345, 2034, 1610, 1504, 2720, 210, 1839, 1244, 1503, 2755]. CRC/Chapman [541]. create [236]. Creutzfeldt [1584, 2268]. Criteria [2434, 1916, 2330, 266, 771, 1870]. criterion [1396, 269, 1968, 986, 1981, 264, 1854, 227]. Critical [3388, 2798, 128, 2771, 2775, 1905, 2772, 2773, 2774, 540]. critique [3314, 3315, 1965, 2110, 3313, 3319, 3316, 3317, 3318, 1399]. CRM’ [3011, 2059, 3010]. Cross [1489, 2163, 2133, 2702, 2086, 2293, 332, 813, 772, 1183, 444, 831, 832, 833, 430, 1703, 1571, 3028, 3293, 1185, 1474, 2076, 911, 2100, 616, 503, 1007, 3055, 1088, 1171, 1702, 2318, 1420, 2089, 806, 170, 2309, 1871, 73, 2250, 371, 2081]. cross-component [332]. cross-cultural [170]. Cross-over [2163, 2086, 813, 772, 1183, 444, 831, 832, 833, 1703, 1571, 3028, 1185, 1474, 911, 616, 3055, 1702, 2318, 1420, 2089, 806, 2309, 1871, 2250]. cross-platform [2702]. cross-section [2076]. cross-sectional [430, 3293, 2100, 503, 1007, 1171, 73, 371, 2081]. Cross-validated [2133]. cross-validation [2293]. cross-validatory [1088]. crossing [2256]. crossover [3357, 3304, 1604, 2835, 607, 1474]. Crowder [1593]. Crowley [1438, 342, 1016]. crude [175]. Cruickshante [146]. Cryptosporidium [950]. cubic [659, 118]. Cullen [542]. cultural [170]. culture [2404]. cultures [1632]. Cumulative [175, 548, 2171, 190, 1004, 3291, 2526, 1442, 3036, 3046, 3345, 2988]. Cure [2627, 691, 1407, 289, 3132, 925, 3001, 218, 2301, 3050, 2447]. cured [1735, 502, 478, 3346]. Curious [1909]. Current [2652, 654, 1632, 3344, 2816, 3293, 282, 902, 3248, 1692, 2653]. curriculum [745, 668]. Curtailed [3065, 2331]. curtailment [1368, 1020]. Curvature [2425]. Curvature-adjusted [2425]. curve [2944, 2801, 877, 1942, 2166, 1672, 2638, 1536, 2641, 2285, 1033, 61, 916, 1155, 1682, 2640, 2643, 3015, 1000, 3123, 1802, 853, 1446, 1585, 2195, 2854, 1436, 2817, 2636, 2637, 3237, 2084, 1523, 2385, 390, 763, 1745, 2647, 2639, 2486, 1870]. curves [1813, 3139, 2513, 1919, 3284, 3121, 1673, 24, 3375, 918, 1411, 3007, 1106, 3321, 60, 2548, 718, 2646, 2749, 1654, 1987, 215, 1501, 1251, 762, 1189, 2649, 2408, 562, 721, 826, 3006, 2645, 450, 3299]. CUSCORE [1204]. CUSUM [2871, 2876, 2872, 2565]. CUSUM-based [2872]. cut [2119, 1950, 1430]. cut-off [2119, 1430]. cut-offs [1950]. cutpoint [1405]. cuts [2520]. Cuzick [1591, 1709]. cystic [76, 765]. Cytotoxic [2073, 1090]. cytotoxicity [3287].

D [1530, 1951, 509, 803, 601, 2820, 144, 2604, 13, 1608, 2907, 782, 129, 603, 42, 3379, 258, 511, 131, 3057, 2134, 2503, 3380, 1142, 780, 524, 11, 1595, 221, 1016, 494, 2817, 2855, 2559, 3056, 1799, 544, 43, 437, 400, 2718, 2160, 29, 3302, 842, 3137, 94, 3382, 3221, 283, 1028, 1351]. D.Tobias [782]. D’Agostino [2754, 3190, 2640, 2642]. daily [552]. Dale [600, 1440]. Dalene [871]. Daley [885]. DALYs [3196]. Dan [1671]. Dani [93, 454, 2858]. Daniel [714, 359, 2519, 187]. Danish [1267]. Dankmar [601, 3312, 417]. Daryl [885]. Data [341, 2021, 526, 460, 2504, 155, 1385, 129, 2463, 961, 558, 1245, 1352, 182, 1310, 1156, 419, 170, 2620, 1416, 892, 94, 936, 114, 579, 1275, 645, 1769, 967, 3048, 948, 1082, 3041, 1317, 2002, 1951, 2000, 2245, 1892, 559, 3359, 2944, 2030, 1706, 1652, 3351, 470, 311, 1848, 3324, 1695, 1374, 1905, 53, 1127, 1163, 3222, 1554, 45, 697, 2324, 689, 3166, 710, 965, 3051, 3079, 232, 2122, 249, 3186, 2172, 2189, 1346, 1207, 2785, 931, 2019, 2864, 702, 2292, 2525, 279, 473, 750, 1898, 1242, 113, 1539, 2891, 1869, 2880, 2040, 1672]. data [134, 516, 613, 2082, 2695, 49, 2592, 2795, 2915, 3273, 3344, 3357, 1249, 2454, 777, 2180, 351, 1580, 2286, 755, 1365, 1477, 767, 1349, 256, 295, 841, 1768, 3184, 1737, 708, 1108, 2816, 2931, 1191, 3342, 3280, 1087, 1492, 1258, 1496, 565, 1388, 2923, 1009, 2610, 935, 2288, 2779, 969, 3016, 2471, 597, 1673, 992, 1148, 3210, 726, 2426, 3333, 571, 1386, 2261, 393, 2004, 598, 463, 423, 2008, 2608, 676, 1413, 2521, 1986, 2992, 2378, 3077, 3078, 2420, 2267, 1215, 2175, 1929, 970, 2083, 1645, 162, 1288, 778, 1227, 1475, 2058, 101, 3354, 3075, 1300, 1522, 2029, 111, 181]. data [929, 1448, 1635, 1359, 848, 1794, 2756, 2483, 3144, 2522, 1042, 3312, 1558, 2470, 2693, 1320, 1077, 2206, 2325, 3335, 1252, 1975, 2218, 1372, 2812, 3293, 546, 3060, 1707, 1863, 1802, 1493, 1957, 194, 2283, 2557, 630, 3217, 1433, 2290, 2381, 2861, 3376, 3338, 1742, 353, 2717, 3254, 2063, 853, 996, 1926, 2371, 3234, 1993, 865, 1533, 2503, 1423, 534, 518, 395, 1062, 2340, 2869, 1640, 2044, 2965, 2127, 2223, 2882, 1925, 3256, 2893, 21, 1551, 956, 2705, 2627, 533, 3037, 1777, 1956, 67, 2461, 1781, 1476, 966, 573, 1389, 3035, 1113, 659, 923, 1574, 581, 1981, 2489, 1190, 684, 1481, 2943]. data [3188, 1261, 3305, 1451, 241, 506, 2061, 2013, 2697, 3257, 724, 1052, 1998, 2730, 2947, 22, 84, 583, 2126, 180, 2633, 940, 687, 2308, 3034, 3074, 3167, 2884, 1085, 1893, 1958, 512, 3148, 917, 1313, 1949, 1595, 2959, 548, 2835, 2307, 3044, 1007, 333, 3268, 1278, 660, 465, 504, 1680, 32, 203, 428, 711, 1016, 2106, 1660, 1299, 1688, 1454, 3101, 3347, 735, 26, 280, 1362, 1104, 1605, 749, 555, 168, 1354, 1318, 2704, 2264, 348, 1820, 1852, 1356, 2854, 2456, 683, 1927, 1089, 2509, 1256, 2855, 3301, 3225, 1665, 1999, 1360, 3363, 2559, 1047, 3080, 498, 1171, 1373, 52, 1851, 1721]. data [1612, 392, 893, 2547, 1243, 254, 2887, 250, 20, 596, 2615, 560, 690, 681, 856, 2104, 243, 1394, 2933, 3069, 1304, 1262, 1043, 2519, 939, 2323, 1637, 912, 1614, 478, 2388, 383, 2458, 253, 2562, 2941, 3295, 1847, 233, 753, 754, 3292, 1499, 1897, 2622, 1937, 1501, 2760, 3062, 2240, 2719, 1959, 1627, 2362, 1667, 2761, 2274, 2217, 2045, 2027, 834, 1330, 2682, 3054, 1516, 1859, 2253, 2686, 2572, 2728, 3038, 1005, 2581, 655, 1421, 1823, 2786, 623, 2280, 2635, 1711, 3165, 2574, 1542, 2385, 3046, 2121, 700, 2016, 1512, 3267, 3258, 1391, 2434, 2655, 1686, 2925, 1783]. data [3214, 1129, 1488, 2853, 3365, 1006, 691, 667, 31, 110, 532, 2863, 1063, 1401, 449, 2435, 2679, 1377, 650, 2696, 236, 1586, 2103, 3145, 109, 628, 2469, 2523, 1573, 1583, 153, 1301, 1837, 2533, 2303, 3306, 1189, 653, 1210, 2071, 2885, 2524, 2433, 373, 2250, 172, 943, 1912, 3253, 3117, 408, 763, 703, 723, 1748, 2511, 3175, 1071, 391, 685, 1250, 1457, 2683, 2684, 547, 2792, 1588, 521, 932, 3082, 2671, 2132, 1450, 2476, 3088, 2081, 1931, 2862, 1406, 3066, 2023, 1828, 2390, 2892, 2852, 3290, 1147, 633, 2192, 2337, 2497, 2692, 2831, 3382, 957, 3050, 2706, 1407, 2115, 721, 1692, 2506, 2406, 2670]. data [1400, 2549, 2727, 133, 2101, 227, 1130, 2252, 2305, 1091, 2459, 327, 3378, 427, 3299, 1809, 2133, 3017, 1387, 402, 1245, 1194, 566, 104, 1363, 1145, 187, 186, 433, 1244, 41, 1530, 1532, 1471, 1547, 1610, 1684, 1402, 285]. data-adaptive [3333]. data-application [1492]. data-revisited [1227]. database [161, 2182]. date [470, 396, 1123, 2980]. dates [116]. Davey [2839]. David [819, 2245, 1565, 1752, 360, 434, 1364, 361, 1402, 182, 2140, 1531, 2961, 545, 1206, 2787, 1825, 3302, 1503]. Davis [1647, 3136]. Dawid [1608, 302]. Day [2245, 303, 1866]. day-care [1866]. Dayton [455]. DC [821]. Dealing [542, 930, 3208, 669, 2219]. death [1082, 3229, 2907, 1287, 1443, 1077, 1240, 2308, 1262, 3249, 1843, 2701, 2539, 2908, 1698, 3137, 3378]. deaths [328, 2414]. Debajyoti [1517]. Debbie [2839]. decade [377]. Decennial [2920]. deciding [2434]. deciphering [2654, 259]. Decision [2026, 1018, 1355, 315, 3206, 3340, 1475, 404, 1628, 1629, 675, 28, 2147, 2578, 3134, 2084, 1623, 3133, 1906, 3194]. decision-making [315, 28, 2084]. Decision-theoretic [2026, 2578]. decisions [1622]. deck [2633, 972, 2632]. decline [163, 166, 2258]. Decomposition [2046, 1278, 2485]. defect [432]. deficiency [2220]. defined [955, 1164, 650]. defining [2914, 771, 590]. Definition [267, 1308]. definitions [980]. degree [3307]. degrees [2272]. Dekker [1438, 201, 495, 1192, 871]. delay [1305, 2867, 1618]. delay-reported [2867]. delayed [76]. deletion [1948]. delivered [578]. delivery [1280]. dementia [1287, 162, 1288, 1077, 1297, 165, 1293, 169, 2270, 2258]. DeMets [1402]. demography [2718]. demonstrating [1715, 1714, 1716, 1713]. Denison [1364]. Denmark [1470]. Dennis [3302]. denominator [2272]. density [3051, 3256, 477, 854, 109, 3240]. dental [3351, 2525, 1492, 3354, 3080, 3355, 1931]. depend [1829]. dependence [3085, 3128, 3285, 732, 488, 2883, 1318, 2759, 2530, 3306, 2511]. Dependencies [42]. Dependency [1513, 2460, 2669, 607]. dependent [2257, 3048, 1892, 2948, 3083, 3168, 2166, 2386, 51, 1738, 850, 2289, 1888, 3084, 3129, 1926, 3380, 1889, 1261, 2417, 2457, 2028, 3276, 2955, 27, 1680, 2950, 1820, 3172, 2711, 507, 2701, 2016, 2810, 3345, 2807, 445, 1988, 661]. depressive [1002]. derived [2307, 1190]. deriving [2698]. Describing [1141, 1834]. Description [2584, 19, 2819]. descriptor [1641]. Design [341, 1652, 1482, 2099, 1067, 92, 1311, 336, 780, 2593, 990, 418, 210, 2647, 1184, 1813, 2540, 157, 323, 2224, 2423, 1319, 3340, 2552, 3282, 985, 2382, 2425, 2629, 2499, 2582, 3000, 992, 2004, 2847, 161, 5, 2895, 1571, 1882, 2085, 2479, 605, 307, 1742, 212, 2983, 1419, 2055, 3304, 1604, 297, 1701, 528, 654, 2583, 3305, 3349, 1904, 1832, 714, 2835, 2394, 1023, 2200, 2997, 2960, 364, 1463, 3362, 1699, 409, 3235, 2578, 1420, 2580, 370, 1059, 2109, 874, 2232, 1908, 2712, 426, 991, 3279, 3067, 1008, 1120, 1058, 3306, 1189, 2628, 1366, 3092, 562, 889, 2057, 344, 1020]. design [3158, 1730, 2165, 337, 3066, 2077, 3337, 2130, 493, 2101, 817, 3203, 201, 884, 495, 287]. design-based [2085]. designed [1237, 3201]. Designing [461, 204, 1294, 822, 1507, 1039]. Designs [3368, 3373, 3367, 3369, 3370, 3371, 3372, 405, 17, 1778, 2158, 2738, 2736, 289, 2157, 3206, 3183, 2093, 3286, 2411, 3357, 3065, 108, 716, 1743, 1518, 2471, 2918, 2148, 2329, 2586, 2026, 3011, 3009, 2144, 3205, 2011, 1185, 209, 2776, 1757, 979, 2177, 1966, 1474, 1314, 96, 2851, 1983, 1342, 867, 2541, 2674, 2100, 1789, 2372, 3055, 1306, 949, 1436, 466, 3056, 2528, 2809, 1878, 806, 882, 1135, 1380, 2022, 2668, 1967, 3010, 861, 340, 1717, 705, 1021, 556, 1098, 823, 2160, 2136, 397, 2808, 424, 908, 1486, 2090, 2433, 1930, 2930, 2929, 1, 2645, 339, 2226]. desired [768]. details [2537, 2788, 2787]. detect [803, 1806, 3170, 140, 385, 2089]. detectability [999, 1843, 1092]. detectable [1380]. detected [1782, 3255, 1303]. Detecting [931, 2806, 637, 346, 97, 3358, 730, 3089, 1632, 1795, 2597, 2951, 812, 707, 1995]. detection [3233, 1954, 2363, 1747, 2094, 2802, 2598, 1624, 2872, 1964, 848, 2216, 1957, 2998, 1626, 1956, 1965, 1953, 2906, 1958, 2949, 3281, 3237, 939, 3114, 1959, 1627, 860, 2811, 2565, 1961, 171, 981, 2573, 1406, 3149, 1130, 2936]. detector [2597]. deteriorating [669]. deterioration [2662]. determination [931, 2791, 2131, 2465, 1430, 2576, 1411, 1521, 2631, 1765, 802, 3148, 2630, 2147, 1059, 2618, 1818, 1120, 2790, 535]. determinations [2951]. determine [3264, 2342]. determining [2634, 2455, 2291, 1500]. Developing [3286, 603, 2774, 1268, 895, 1101, 897, 896]. Development [3333, 2699, 1454, 640, 777, 3070, 578, 2899, 2975, 1106, 2070, 2413, 169, 1699, 310, 239, 313, 1019, 663, 2930, 1490, 3215, 3150, 201, 3181]. developmental [3280, 2204]. Developments [2231, 1953, 186, 574, 2678, 904]. deviance [227]. deviates [3066]. deviation [1561, 38, 1174]. deviations [1561, 2833, 1174, 2269]. device [1333, 994, 450]. devices [643]. diabetes [1820, 233]. diabetic [1317, 1268, 710, 2354, 2396]. diagnosed [2780, 2700]. diagnoses [116, 1618]. diagnosing [621]. Diagnosis [1950, 1127, 2801, 2363, 2454, 432, 678, 1222, 3156, 2729, 2889]. diagnosis-aiding [432]. Diagnostic [1445, 326, 3155, 18, 1126, 2119, 2750, 731, 70, 2800, 3284, 3270, 3128, 61, 1444, 1811, 766, 1279, 801, 387, 520, 3057, 56, 1793, 1173, 1466, 2457, 2749, 786, 1581, 1987, 2742, 807, 1199, 1313, 864, 62, 1378, 1023, 2950, 1728, 1154, 2528, 590, 1933, 2118, 1336, 46, 1947, 215, 3062, 591, 2778, 63, 2759, 2542, 762, 2799, 1553, 1053, 763, 2408, 3269, 1948, 2067, 2165, 1989, 1988, 1653, 1619, 2681, 450]. Diagnostics [811, 1585, 2251, 1775, 3320, 59, 2179, 1957, 3126, 1824, 2033]. diagram [2436, 2735, 2734, 3215]. diagrams [428]. dialling [1086]. diaries [2206]. diarrhoea [3086]. dichotomization [2472]. dichotomized [2473]. Dichotomizing [1911, 1418, 2219]. dichotomous [15, 731, 39, 1107, 1981, 1202, 3251, 2254, 327]. Dicing [1698]. Dictionary [1415, 357, 303]. Diego [1065]. Diener [2820]. diet [3058]. diet/disease [3058]. dietary [3363]. difference [2878, 1453, 736, 2209, 2386, 2, 3321, 2913, 2646, 3298, 1990, 2674, 2464, 1724, 1949, 2446, 2475, 592, 1799, 2933, 2167, 3172, 3152, 815, 376, 1864, 3145, 2296, 2999, 790, 193, 283]. differences [1643, 1813, 1613, 13, 2590, 295, 841, 3113, 2044, 774, 11, 1617, 7, 1579, 592, 1096, 1380, 1995, 2988]. different [2281, 2481, 2443, 2910, 1853, 176, 2946, 3024, 947, 2909, 652, 2392, 1176, 2424, 338, 2780, 2214, 2686, 837, 329, 389, 1447, 3378]. differential [1291, 927, 2973, 171, 1426, 1427]. differentially [2131]. differentiating [1297]. differing [731]. differnces [29]. differs [1689]. diffuse [1749]. Diffusion [551, 436]. Diggle [129, 1532]. digit [1086]. digital [2127]. dilution [2389, 2945]. Dimension [1072]. dimensional [2946, 1419, 1141, 3256, 3044, 839, 1488, 3033, 3110]. dioxin [709]. dioxins [2784]. Direct [3123, 1520, 1718, 1506, 3141]. direction [3314, 3315, 2532, 3319, 3316, 3317, 3318]. direction-mixed [2532]. directional [1764]. directions [319, 3313, 129]. Directly [22, 1654]. Disabilities [427]. Disability [143, 430, 2564, 2603]. disability-adjusted [2564]. disaggregate [264]. disattenuated [2702]. disciplines [668]. discordance [2951]. Discovering [701]. discovery [3104, 2768, 1939, 578, 2543, 3387, 2765, 2929]. Discrete [187, 2715, 1638, 1898, 1242, 2891, 2695, 795, 793, 3144, 2777, 1974, 3254, 1369, 2803, 3018, 606, 2519, 794, 2385, 3046, 291, 3106, 1759, 2221, 1458, 2762]. discrete-continuous [2777]. discrete-state [1974, 3254]. Discrete-time [2715, 1638, 2695, 795, 793, 1974, 3254, 794]. discretization [3277]. discriminant [203]. discrimination [1892, 1677, 431, 1434, 2644]. discussants [458]. Discussion [740, 2775, 2149, 2150, 2155, 748, 2773, 2152, 81, 746, 743, 744, 2153, 741, 2774, 2154, 586, 375]. Disease [1082, 601, 155, 840, 536, 1965, 259, 730, 948, 1283, 1954, 1584, 157, 72, 159, 100, 731, 245, 223, 876, 1641, 1633, 1508, 2095, 2801, 950, 158, 1003, 472, 1094, 2363, 225, 3089, 2924, 3270, 2268, 1287, 565, 1795, 410, 1284, 160, 1444, 1855, 3187, 3058, 152, 599, 1213, 214, 2017, 149, 248, 2241, 1522, 99, 1992, 163, 2698, 3335, 3195, 2905, 2216, 1297, 140, 934, 1793, 3255, 246, 2361, 1075, 1298, 1054, 1575, 234, 240, 696, 1953, 1123, 3361, 231, 955, 2046, 3156, 3034, 3343, 1958, 1991, 695, 1095, 2201, 2564, 2545, 3196, 1296, 1088, 2339, 2834, 167]. disease [523, 168, 150, 2456, 1797, 2616, 2547, 243, 3069, 1027, 1096, 1535, 1344, 1499, 1442, 2269, 2309, 411, 422, 229, 68, 2565, 235, 206, 34, 1128, 1890, 153, 1334, 172, 3163, 1736, 1457, 2978, 809, 2132, 1963, 237, 1995, 1289, 2936, 3378, 151, 3266, 417]. disease-risk [1995]. disease-specific [1890]. disease-state [3195]. diseased [1141]. diseases [2000, 158, 464, 2954, 21, 1560, 621, 308, 837, 3239, 1025, 2301, 669, 1267, 1082]. disequilibrium [1910, 2191, 3042]. disjunction [1112]. diskette [494]. disorder [828]. disorders [76, 605, 2732]. Disparities [2914, 2915, 2916, 1794, 2925, 2927, 2922]. dispersed [1688]. Dispersion [1994, 200]. Displaying [428]. disposal [2354]. Disproportionate [2348]. dissection [1336]. distance [1641, 747, 748, 560, 2632]. distance-based [2632]. distinct [718]. distinctive [374]. distributed [1079, 161, 501, 1465, 2405, 2997, 1825, 2401, 2649]. distribution [2321, 3320, 352, 941, 2814, 3061, 1641, 2891, 552, 2538, 2939, 1456, 1307, 2017, 396, 2612, 2317, 519, 2557, 492, 3255, 2913, 2614, 1781, 23, 2903, 2850, 2601, 2419, 26, 2475, 2770, 1479, 1439, 3249, 1140, 1501, 650, 3253]. distribution-free [2903]. distributional [2615]. distributions [2725, 3259, 1747, 2094, 2463, 2793, 2688, 2609, 1773, 2660, 488, 3147, 499, 3292, 3201, 672, 73, 762]. district [854]. divergence [1036]. divergent [1881]. dividing [2833]. DKK [782]. DM [434, 343, 494, 1517, 197, 41]. DM29 [40]. Dmitrienko [2754, 2409]. DN [386]. DNA [1145, 1575, 945, 2635, 946, 3042]. Do [3246, 58, 909, 1424, 1146, 2033, 56, 2503, 1246, 1775]. Dobson [1437]. doctors [3246, 3246]. documenting [382]. Dodge [1415]. Does [740, 1396, 1761, 739, 172, 55]. DOI [2751, 2344, 2820, 2638, 2641, 3070, 2450, 3379, 2640, 2643, 2653, 2642, 2489, 2452, 2770, 2637, 2639, 2476, 2839]. Doing [1156]. Donald [1533, 1363, 871]. donation [2622]. donation-transfusion [2622]. Donev [3056]. Donnelly [362, 525]. Donner [817, 3381]. donor [282, 2632]. Dorfman [1707]. dosage [1355, 2375]. Dose [3011, 1757, 3164, 2985, 3010, 1758, 3368, 3373, 3367, 931, 290, 1033, 3000, 2026, 2847, 1109, 3009, 1790, 3369, 2056, 2445, 2776, 625, 1419, 2055, 2177, 2737, 3304, 1441, 2278, 3116, 923, 626, 2300, 1904, 1552, 550, 2394, 1436, 1100, 52, 3370, 805, 3371, 2556, 3372, 514, 97, 556, 1008, 572, 2573, 2265, 1906, 1930, 3033, 2077, 589, 2226, 3221]. Dose-escalation [3011, 3010, 3368, 3373, 3367, 3009, 3369, 625, 3370, 3371, 3372, 1906, 1930]. dose-finding [3000, 2026, 1419, 2055, 2177, 3304, 2077]. dose-ranging [589]. dose-related [550]. dose-response [931, 290, 1033, 1790, 2445, 2737, 2278, 626, 1436, 52, 556, 572, 2573]. Dose-schedule [2985]. doses [1031, 1790, 836, 212, 476, 1634, 2237, 1756]. dosing [529]. double [2867]. double-nonparametric [2867]. Doubly [3165, 1737, 865, 2027, 2670]. doubly-censored [865]. Douglas [2499, 537, 147]. down [979, 3235, 1008, 862]. Dr [1598]. draft [265, 273]. Draper [437, 298]. drift [1856]. drinking [172]. driven [3260]. drop [1161, 1770, 1443, 423, 1883, 1863, 2134, 1383, 1476, 1540, 1755, 1028, 749, 1354, 3073, 912, 383, 667, 653, 445, 337]. drop-out [1161, 1770, 1443, 1883, 2134, 1383, 1476, 1540, 1755, 1028, 749, 3073, 383, 667, 337]. drop-outs [423, 1863, 1354, 912, 653, 445]. dropout [3131, 2931, 2079, 2809, 3147, 1583, 2852, 3290]. dropouts [3077, 3142, 2462]. Dror [782]. Drug [2492, 1221, 1030, 2880, 2795, 2815, 1178, 2375, 578, 3287, 252, 3182, 3181, 212, 2769, 3322, 2422, 2394, 3341, 2182, 2846, 1394, 2970, 384, 310, 2974, 507, 514, 1019, 1136, 1196, 3115, 1120, 1312, 2485, 2930, 1200, 1490, 825]. Drug-drug [2492, 2846]. drugs [2334, 2847, 1166, 2335, 1779, 2987, 2124, 2580, 805, 2057]. Dryden [260]. DSMC [892]. dual [1215, 1904, 2394]. due [1584, 1009, 318, 220, 653, 2683, 2684, 1267, 2827]. Duffy [1591, 962, 1709]. Duley [1159]. Dunn [453, 141, 186]. Dunnett [2739]. Dunning [3150]. duplication [1232]. duration [252, 1843]. during [37, 2241, 1280, 1463, 2170, 2401, 3378]. Durrleman [1715, 1714]. Dutch [1808]. dwelling [2364]. Dynamic [1051, 2886, 1880, 911, 608, 2595, 2037, 2176, 476, 2970, 2217, 623, 2035, 2039, 2036, 2462, 301]. dynamical [835]. dynamics [2041, 2038, 2484, 2837, 879, 2354, 3087, 1500, 1934, 500, 725, 952, 2191, 300].

E9 [569, 974, 975]. ear [1307]. Early [2541, 797, 2353, 852, 289, 2059, 2598, 611, 99, 1298, 1239, 1293, 2949, 3154, 1417, 1555]. early- [2059]. EARS’ [2872]. easy [28]. ebb [909, 1424]. EBM [740, 739]. ECG [432]. ecologic [698]. ecological [697, 2693, 2063, 23, 3196]. Econometric [1194, 41]. Economic [898, 899, 1108, 2580]. economist [649]. ECRU [2360]. Ed [600, 286, 1415, 1550, 543, 1438, 1181, 1311, 1609, 106, 1065, 836]. Edelman [221]. edge [2789]. Edition [2824, 1245, 1156, 1118, 210, 1352, 2519]. editor [1708, 1872, 538, 1117, 1168, 1545, 1649, 1591, 2002, 1951, 2000, 2245, 2344, 2379, 1424, 2910, 355, 1374, 803, 2052, 125, 13, 39, 77, 2735, 2439, 414, 813, 1696, 1784, 2499, 3384, 10, 1682, 2450, 2050, 1645, 2033, 181, 1337, 1589, 387, 2437, 2134, 1715, 297, 1850, 1925, 3380, 1606, 1714, 1054, 1056, 867, 11, 714, 2558, 1016, 2452, 757, 1563, 958, 1362, 960, 1169, 1728, 416, 537, 1725, 2183, 1851, 1799, 2186, 1669, 815, 1711, 857, 2198, 2822, 1716, 1324, 1644, 2032, 3381, 78, 354, 2160, 29, 1053, 397, 3288, 759, 1450, 3221, 283, 1709]. Editorial [1979, 1326, 1903, 3348, 1826]. Editors [2626]. Edn [361, 1530, 1594, 91, 819, 3222, 2138, 2604, 3203, 1751, 1750, 2004, 961, 3311, 1548, 1753, 844, 3181, 1533, 818, 1133, 2959, 821, 3347, 1801, 359, 1437, 1193, 1363, 1532, 2858, 2961, 2109, 1471, 3122, 1648, 2345, 298, 540, 2720, 496, 842, 1066, 1549, 1503, 41, 2755, 495]. eds [1786, 1351, 2754, 144, 1364, 127, 183, 129, 603, 201, 131, 1194, 1182, 417, 1533, 1159, 884, 1142, 300, 302, 1531, 141, 1516, 2345, 820, 1547, 510, 186, 433, 871, 94]. education [2279, 2537, 2788, 2787]. Edward [438, 2183]. Edwards [1565, 3383]. EEG [2510, 2512]. Effect [1944, 3255, 2569, 1082, 2282, 2894, 697, 859, 2550, 1942, 2878, 3229, 1236, 3273, 487, 296, 1747, 2094, 2937, 729, 585, 3384, 2563, 2724, 24, 3208, 3190, 1307, 2050, 1288, 2317, 2698, 2448, 914, 501, 796, 1405, 474, 1040, 1966, 3297, 1576, 1551, 2614, 2428, 847, 615, 2466, 1616, 760, 512, 1552, 3268, 657, 2452, 407, 3283, 168, 1356, 1940, 1941, 2220, 2809, 1394, 2322, 2689, 1917, 1140, 135, 312, 1867, 188, 3152, 2745, 3325, 2369, 2574, 2619, 313, 514, 3381, 2657, 1511, 2746, 2215, 2999, 1315, 1657, 2512, 2310, 1946, 3160, 2441, 2744, 3266]. effect-adjusted [2724]. effect-measure [2878]. Effective [2014, 1031, 2638, 2643, 2219]. Effectiveness [2320, 2060, 926, 899, 2621, 2880, 87, 281, 2049, 1372, 1715, 474, 2015, 1714, 451, 476, 1634, 2237, 2415, 392, 1741, 2458, 215, 1716, 1713, 1254, 2351, 620, 1588]. effectiveness-one [1715, 1714, 1716, 1713]. Effects [1275, 560, 2257, 1943, 1706, 2297, 3243, 2407, 1977, 1163, 3061, 2703, 2181, 2449, 3265, 3277, 2907, 3334, 2802, 1658, 708, 1191, 1087, 795, 793, 1655, 1146, 120, 2747, 3063, 2048, 1505, 829, 900, 1719, 1257, 2378, 3077, 3078, 673, 2079, 2289, 3113, 1556, 2757, 33, 1894, 2164, 95, 219, 1074, 3084, 2068, 1677, 770, 1478, 331, 3120, 3234, 921, 1467, 1718, 528, 1172, 2614, 3093, 1264, 3035, 1510, 1889, 2660, 3188, 2417, 1634, 2237, 2415, 845, 2405, 2903, 2730, 2850, 2239, 2849, 2883, 2100, 706, 3136, 1495, 2319, 2920, 1480, 2106, 1579, 3227, 3094, 3341, 1379, 189, 416, 1479, 2183]. effects [498, 924, 1799, 2377, 2603, 1702, 2318, 226, 383, 2881, 1847, 2387, 874, 2593, 2235, 794, 1667, 2027, 834, 468, 2232, 2270, 2572, 322, 2244, 2556, 2974, 2198, 1105, 2258, 1985, 2539, 2908, 2421, 3141, 3219, 2053, 785, 3345, 2808, 1890, 390, 1815, 1816, 2571, 1841, 1458, 317, 1664, 2763, 647, 3278, 1736, 3137, 2193, 790, 1885, 2018, 2978, 1881, 2132, 1450, 500, 725, 952, 2845, 2171, 1241, 1854, 1426, 2932, 486, 594, 3102, 121, 2009, 1450]. Efficacy [1208, 717, 640, 1807, 1477, 1496, 622, 3000, 850, 350, 1042, 86, 1661, 3304, 1559, 954, 3264, 3164, 1032, 1110, 2316, 2970, 2053, 3115, 2146, 424, 1841, 2546, 2077]. efficacy/tolerability [3304]. Efficiency [1571, 391, 2895, 2983, 1611, 2809, 2458, 2722, 2410, 1337, 1725]. Efficient [2738, 1633, 1966, 3305, 3301, 2726, 2121, 340, 2473, 2093, 1846, 2382, 2144, 1474, 3013, 2591, 1124, 2057]. efficiently [3324]. Egypt [645]. EIC [1854]. Eight [1371, 2919]. Einführung [197]. EJ [386, 3383]. elderly [2835]. electroencephalograms [2181]. Electronic [1333, 1312, 590]. Elementary [106, 3180]. Elements [1310]. Elena [3231]. elevated [410]. elicitation [406, 759]. Eliciting [1885]. eliminate [2916]. eliminating [637]. elimination [2865, 1743]. ELISA [1190]. ELISA-derived [1190]. Elizabeth [1516, 510]. Ellenberg [1402]. Elliott [1142]. ellipse [3052, 3288]. elliptic [2197]. Elsevier [1786]. Elston [3386, 3310]. Elteren [2096]. elusive [294]. Elwood [540]. EM-based [1047, 1479]. embryo [2240]. embryo-specific [2240]. emergence [965]. emergencies [309]. emergent [3012]. emerging [464]. Emmanuel [1696, 1750, 417]. emphasis [1811, 2384, 2786]. Empirical [3085, 961, 294, 3023, 72, 767, 367, 174, 1571, 779, 224, 2973, 1269, 1176, 2854, 3387, 3100, 237]. enabled [577]. encompassing [1972]. encounter [2356]. encounters [985]. encouragement [2101]. Encyclopedia [127]. end [3034]. endothelium [631]. endpoint [2575, 585, 3070, 2975, 3150, 1757, 615, 3385, 1904, 2562, 1208, 2577, 3158, 1184, 1731, 3194]. endpoints [2327, 1396, 2313, 2312, 79, 1597, 1412, 1097, 2164, 1137, 2249, 1596, 1598, 2352, 3013, 1832, 906, 345, 2997, 606, 614, 1756, 1915, 1810]. Enfield [2006]. England [1425]. English [2926, 368]. English- [2926]. enhance [2359]. enhanced [1691, 3044]. Enhancing [578]. enoxaparin [874]. enrollment [2926]. entitled [569]. entropy [2518]. entry [1795, 2486]. Environment [183, 2125, 161, 674, 2204, 2214, 1347, 2827, 1570, 183]. environment-induced [2204]. Environmental [439, 1203, 935, 2267, 247, 2126, 1304, 2682, 1711, 2917, 1774, 1169, 869]. EORTC [667, 971]. EPIC [3058]. epidemic [2268, 850, 666, 2979, 879, 2770, 2843, 521, 3088, 2688, 885]. Epidemics [259, 3130, 32, 2949, 242, 1457, 362]. epidemiologic [1317, 710, 1286, 1872, 1230, 1573, 1143]. Epidemiological [1245, 540, 613, 3174, 1812, 2812, 2733, 626, 2214, 2421, 785]. Epidemiology [2064, 1142, 308, 401, 36, 747, 1203, 244, 915, 1070, 748, 1225, 927, 2709, 1991, 1609, 88, 854, 608, 467, 188, 208, 1233, 557, 2839, 437, 2840, 91, 1144, 103, 1566, 1840, 2004, 103]. Epidemiology. [341]. epilepsy [2922, 873]. epileptic [2880]. episode [1432]. episodes [66, 102]. epistatic [2191]. equal [2012, 1644, 919, 2410]. equality [1861, 2513, 2286, 51, 1659, 493]. equalizing [477]. equals [2826, 2825]. equation [553, 1035, 2599, 1188, 2973, 1202, 2794, 856, 3267, 2396, 2844, 3262]. equations [3157, 120, 627, 1339, 1858, 3146, 630, 3019, 2208, 958, 1125, 2186, 775, 912, 3039, 3165, 2701, 1512, 2210, 1836, 593, 1338, 320, 1130, 1460]. equilibrium [1844, 2269]. Equivalence [2830, 1041, 2897, 1024, 528, 3108, 687, 1987, 1313, 2630, 590, 370, 642, 2654, 614, 993, 1985, 1058, 1329, 1810, 2549, 1205]. eradication [1249]. erbB [1151]. Eric [2140]. erroneous [637]. Error [1162, 2702, 1361, 3324, 553, 3202, 2953, 1203, 1079, 530, 1044, 3174, 1536, 24, 2781, 3058, 3002, 2175, 612, 2898, 2568, 927, 3338, 2602, 3045, 1551, 96, 1246, 528, 2614, 1465, 2920, 1609, 1357, 3076, 2854, 2534, 3237, 2243, 215, 2274, 2879, 1603, 2369, 860, 1019, 376, 3353, 2555, 2656, 2595, 2853, 1717, 2401, 1045, 1734, 2649, 408, 2569, 632, 1347, 2929, 2827, 2969, 1653, 211, 185]. errors [1121, 2414, 3216, 752, 3196, 2430, 1717, 1406, 3003, 2009]. errors-in-variables [752, 3003]. Escalation [2177, 3368, 3373, 3367, 3011, 3009, 3369, 1757, 625, 911, 3370, 3371, 3372, 3010, 1758, 1906, 1930]. ESRD [3034]. Essays [94]. essential [2772, 1671, 1648]. establishing [2119, 1037, 733]. Esteve [962]. Estimate [1174, 2379, 2702, 3051, 1453, 516, 552, 795, 793, 565, 8, 3187, 177, 1522, 2015, 853, 1884, 2601, 112, 2316, 2387, 794, 1843, 2165, 1731, 1426, 1561]. estimated [272, 2421]. Estimates [3081, 1098, 72, 159, 2356, 2209, 2049, 410, 1733, 33, 638, 2698, 2448, 927, 3084, 2557, 980, 2515, 2890, 84, 555, 2393, 1100, 1277, 1114, 2710, 229, 1618, 347, 1431, 2215, 2569, 3006]. Estimating [717, 3155, 640, 1302, 1541, 2111, 1977, 2623, 1928, 876, 3051, 859, 2953, 2990, 1076, 414, 1807, 1273, 1477, 3364, 116, 430, 1496, 160, 2114, 2357, 214, 162, 1432, 1668, 796, 2445, 1004, 165, 2965, 1551, 774, 1735, 1232, 2419, 2854, 1201, 2534, 3341, 189, 416, 1256, 1125, 592, 3281, 2118, 216, 499, 2970, 764, 2976, 2243, 1140, 2780, 3292, 468, 2253, 411, 1196, 2700, 73, 1303, 1008, 593, 408, 2018, 521, 2081, 2486, 1338, 3299, 1460, 2702, 827, 1706, 2060, 2443, 2910, 352, 2297, 2993, 3, 768, 564, 1079, 2766, 1453, 2815, 2802, 3157, 120, 627, 3058, 3024, 900]. estimating [1339, 2895, 3097, 2176, 396, 2909, 1858, 163, 3146, 294, 630, 2838, 2602, 921, 2392, 3019, 2867, 812, 2493, 3136, 2920, 2208, 958, 1436, 775, 856, 912, 3039, 3165, 1105, 2701, 1512, 2210, 3267, 3214, 479, 1890, 1666, 1989, 339, 320, 1130, 2226, 2459, 3383]. Estimation [2157, 2166, 3273, 1830, 1305, 119, 850, 916, 2414, 3144, 652, 2206, 2688, 282, 752, 2466, 1616, 2730, 2126, 1817, 2994, 441, 1165, 1243, 1304, 3001, 1947, 2972, 2217, 185, 3239, 1092, 1831, 2746, 628, 722, 943, 2649, 1457, 1347, 2901, 2418, 2546, 489, 2257, 2874, 1272, 2282, 3025, 2865, 470, 53, 3104, 2945, 75, 1633, 2095, 2801, 950, 877, 2864, 1674, 369, 1103, 1846, 76, 755, 2758, 2268, 281, 729, 3121, 627, 915, 3024, 1811, 682, 2289, 1571, 574, 1589, 1099, 2203, 3228, 3015, 1676, 2789, 2056, 3123, 1372, 2812, 86, 2068, 3321, 2338, 2427, 836, 3254, 2371]. estimation [3297, 1061, 1062, 1341, 1441, 2954, 1718, 1694, 800, 2299, 3116, 2461, 1781, 2609, 2156, 769, 1113, 3261, 2646, 724, 851, 2850, 2304, 2813, 706, 503, 917, 1150, 1724, 1681, 2446, 1495, 595, 3196, 2452, 1028, 949, 280, 531, 958, 3098, 1265, 854, 3301, 3023, 2468, 250, 1050, 1462, 2933, 1434, 3154, 1878, 3249, 2458, 835, 3198, 2526, 2205, 2274, 2045, 834, 672, 1421, 1823, 3274, 2359, 448, 2385, 2121, 700, 1271, 377, 1967, 1124, 1187, 110, 449, 650, 2696, 557, 2355, 1435, 1722, 1345, 1816, 1210, 2571, 2628, 2467, 2518, 1253, 1657, 670, 3052, 3288, 2843, 862, 2062, 1184, 1854, 237]. estimation [1407, 2194, 2506, 2447, 2096, 2129, 2762, 1711, 2770, 105]. estimations [840, 376]. estimator [1651, 2356, 530, 970, 2777, 3298, 775, 1859, 1828, 1447]. Estimators [159, 1615, 245, 2311, 351, 487, 1033, 1640, 760, 616, 1371, 3283, 1262, 3039, 2370, 653, 2549]. Ethical [895, 469, 374]. Ethnic [3038, 2918, 947, 2925, 2922]. EUR [1786, 1872, 2961, 1546, 1516, 3231, 1684]. EUR54.00 [3182]. EUR67.50 [1752]. EUR69.50 [2501]. EUR74.95 [1750]. EUR82.50 [1774]. European [903]. Euston [436]. Eva [1873]. evaluate [553, 2992, 953, 619, 2780, 412, 2668, 572, 3163, 594]. evaluated [2404]. Evaluating [852, 1126, 899, 1632, 947, 673, 2975, 2998, 1956, 2940, 2674, 1739, 2636, 1877, 2074, 641, 2956, 1128, 107, 77, 1242, 937, 1525, 1518, 1036, 766, 262, 847, 1820, 1356, 14, 514, 1136, 78, 3140, 1490, 2638, 2641, 3070, 3150, 2640, 2643, 2642, 2637, 2639]. Evaluation [2086, 3243, 2750, 3022, 212, 534, 1559, 1261, 962, 606, 1702, 236, 3031, 1866, 3105, 2327, 2800, 473, 1076, 767, 256, 2888, 622, 3333, 2329, 2586, 2450, 1787, 224, 898, 1957, 3120, 2349, 2124, 266, 336, 2803, 234, 2742, 1296, 657, 62, 1089, 1941, 874, 2235, 57, 2542, 2925, 1129, 556, 3106, 2260, 1915, 1930, 2845, 237, 1752]. evaluations [591, 2146]. Evans [958]. Event [597, 182, 2076, 132, 3048, 2000, 2440, 909, 1424, 1613, 2181, 2907, 2439, 1907, 1782, 1191, 935, 2842, 1428, 2757, 2335, 1368, 625, 954, 21, 2705, 1333, 241, 851, 488, 1693, 1595, 2835, 2201, 2837, 1913, 1920, 973, 3012, 3295, 1823, 2385, 3046, 2539, 2908, 1686, 1610, 2810, 2696, 109, 1128, 1440, 1210, 1912, 3137, 703, 932, 2969]. event-based [909, 1424]. event-related [2181, 2837]. event-time [2385, 3046]. Events [3171, 1209, 2919, 941, 379, 2169, 2234, 1636, 2552, 3282, 2608, 190, 2335, 2600, 2514, 2124, 2609, 2902, 2633, 1680, 1913, 2202, 2182, 2294, 1420, 2259, 1831, 2746, 1334, 1664, 828, 1906]. Everitt [1751, 1067, 844, 357, 141, 2109, 2719, 186, 1839, 2755]. everything [890]. Evidence [315, 90, 2403, 1197, 866, 2431, 472, 2049, 1475, 4, 1180, 1715, 2503, 1714, 1506, 1671, 321, 2214, 857, 1716, 1713, 1254, 1906, 873, 1870, 2913, 3231]. Evidence-based [315, 2403, 1180, 1671, 1870]. Evolution [3314, 3315, 3313, 3319, 3316, 3317, 3318, 328, 1583, 525]. EWMA [3049]. Exact [579, 2464, 991, 291, 705, 3145, 2441, 405, 3105, 3061, 736, 580, 2888, 1811, 2646, 2884, 1266, 2353]. Examination [167, 326]. examinations [807]. Examining [1920, 1643]. example [1631, 3212, 3289, 2915, 565, 3204, 927, 1106, 307, 954, 321, 667, 650, 1657, 1026]. examples [19, 505, 1048, 1750]. excess [707, 1890]. excessive [3354]. exchangeable [3280, 2299, 1147]. exclusion [976]. exclusive [2938]. exist [56]. existence [626]. expansion [2284, 1524]. expectancies [430]. expectancy [2913, 2046, 3023, 1303]. expectation [1349]. expectations [2659]. expected [617, 2046, 2412]. Expenditure [2364]. experience [903]. experiences [2820, 2617]. experiment [2598, 209]. Experimental [2847, 1120, 210, 2513, 3305, 1699, 3056]. experiments [2131, 2203, 2983, 2278, 1766, 504, 3119, 3294, 606, 1237, 2764, 1586, 2765, 92, 780, 418]. expert [3343, 1885]. Expertensysteme [197]. expertise [3092]. expertise-based [3092]. expiratory [1574]. Explained [1620, 1236, 615, 2369, 1138, 1241]. Explaining [2927, 1096]. explains [2904]. explanatory [2781, 588, 771, 2534]. exploration [1072, 2463]. exploratory [219, 1104, 2342, 2792]. Exploring [2029, 786, 2837, 226, 1091, 875]. exponential [552, 934, 1585, 1501, 2277, 2238, 2247]. exponentially [2997]. exposed [1316, 1855]. Exposure [542, 394, 3020, 877, 1203, 2125, 3336, 3342, 1009, 2494, 1909, 2698, 2448, 2784, 2204, 2366, 475, 1990, 3343, 1150, 2207, 98, 1499, 1211, 2214, 2619, 2555, 3100, 2662, 1734, 2301, 607, 1736, 828, 2569, 2310]. exposure-disease [3343]. exposure-response [2494]. exposure-time-response [3020]. exposures [3202, 1674, 1944, 3058, 1604, 1140, 2881, 2879, 3345, 1664, 1599, 2827, 545]. expressed [2131]. expression [1827, 1269, 1409, 1516, 1005, 2764, 2390, 2904, 2133, 3358]. expressions [3295]. Extended [1854, 2470, 1961]. Extending [526, 375, 1749, 3156, 2182]. extension [2400, 675, 2274, 2418]. extensions [1723]. external [2820, 2617, 688, 1056, 3124]. extra [2379, 8, 1582, 3268]. extra-binomial [1582]. extra-Poisson [2379, 8]. extract [2581]. Extracting [1412]. extraction [983]. extraneous [3208]. extrapolate [2049]. extrapolation [2972]. extreme [2389, 2945]. extremely [3135, 2804]. eye [2924, 3112]. eyes [1522].

F [1565, 1608, 813, 2856, 2134, 3380, 302, 821, 2858, 1471, 128]. facilitate [3226]. Facilitating [2698]. Factor [361, 160, 3208, 1405, 732, 1342, 2994, 1857, 98, 2084, 1105, 1488, 971]. factor-by-curve [2084]. factorial [3286, 212, 1189]. Factors [2926, 998, 1591, 1258, 1738, 1964, 1796, 67, 2126, 1422, 2580, 1304, 2881, 2387, 1230, 57, 1711, 3064, 1128, 1890, 1358, 873, 725, 2729, 1130, 489]. faculty [728]. failing [1795]. Failure [1244, 1530, 645, 36, 2178, 279, 1308, 2592, 2426, 644, 2083, 292, 432, 2568, 732, 3217, 2861, 2614, 1510, 925, 760, 660, 1852, 1395, 942, 293, 2941, 1421, 507, 2696, 2561, 1458, 1664, 2486, 2192, 3050, 2447, 2406, 1400, 2936]. failure-time [2696]. failures [1367, 838]. fallacies [1024]. fallacy [207]. falls [2835]. false [1615, 2768, 190, 2543, 2468, 1947, 2765, 2929]. false-discovery-rate-based [2768]. false-negative [2468]. false-positive [190]. falsification [1024]. Familial [2767, 217, 320, 948, 1591, 3360, 1339, 1075, 3019, 2996, 1318, 1358]. families [1497, 3103, 1900, 2532, 2704]. Family [1285, 552, 3118, 3187, 732, 2557, 488, 2126, 1480, 2424, 1304, 1711, 2277, 2620, 2706, 2929, 3102, 1972]. Family-based [1285, 3102]. family-wise [2929]. Faraggi [1016, 3221]. Faraway [2005]. Farewell [2961]. Farrell [1159]. Farrington [815]. Fast [1954, 1019]. fatal [442]. fatality [2371]. favour [289]. Fay [2379, 386, 2418]. Fayers [144, 544]. FDA [273, 570, 2146, 1390]. FDG [1297]. FDG-PET [1297]. FDR [1765, 2530]. FDR-control [2530]. FDR-controlling [1765]. feature [983, 837, 2516]. features [364]. Fedorov [2593]. feeding [640, 639]. Feifang [2856]. Feinstein [1119]. fellows [881, 728]. females [1814]. fenestration [1307]. Feng [958]. Ferguson [362]. Feridun [183]. Ferlay [258, 131]. Fernholz [94]. fertility [827, 2575, 71, 413, 643]. Feuer [2379, 386, 3383]. few [676, 1413, 2717, 2366]. fibrillation [3238]. fibrodysplasia [865]. fibrosis [76, 765]. Fiddian [1925]. Fidler [416]. field [460, 2508, 2360, 2512]. Fieldhouse [600]. fielding [309]. Fields [2139]. Fieller [946]. Fifth [2570, 3296]. figures [305]. fill [2551]. film [590]. filtering [3044]. filtering-enhanced [3044]. filters [236]. final [3378]. financed [1389]. Finding [3064, 3000, 2026, 1419, 2055, 2177, 2737, 3304, 2985, 1904, 805, 3033, 2077]. findings [1387, 1403]. fine [1910]. fine-mapping [1910]. Finetti [649]. fingerprint [2635]. Finite [3039, 1430, 3189, 1445]. Finnish [233]. Firebaugh [199]. firing [2513]. firing-rate [2513]. First [207, 2696, 3243, 2389, 2945, 2603, 2163, 838]. First-event [2696]. Fischer [197]. Fisher [2751, 3105, 1715, 1714, 3243, 2488, 2888, 874]. Fishing [1379]. fit [810, 563, 564, 3300, 1679, 2935, 191, 880, 966, 1381, 1587, 1154, 1408, 677, 1348, 278, 1891, 2782, 723, 133, 960]. fitness [2299]. Fitting [2989, 3022, 3169, 1673, 228, 1935, 1971, 1802, 1802]. Fitzmaurice [2502]. Five [131]. Fixed [3113, 847, 2539, 2908, 3137, 3063, 1368, 2548, 2794, 845, 2167, 2765, 647, 1619, 2907]. Fixed-effect [847]. FL [2824, 3057, 2399, 2960, 2858, 2719, 3122, 2841, 2720, 2755]. Fleiss [1801, 418]. Fleming [1402, 2653]. Flexibility [3352]. Flexible [2993, 3377, 2582, 2660, 2616, 2377, 834, 3345, 1758, 1458, 3096, 3085, 2157, 2276, 3216, 2614, 2674, 714, 1039, 1878, 426, 2971, 2429]. floating [1277]. floor [1556]. floor-effects [1556]. Florida [2917]. flow [909, 1424, 1574]. flowgraph [957]. flu [1856]. fluid [1632]. fluster [2689]. fMRI [1782, 3275, 2506]. focus [332]. foetal [37, 629, 1093, 678, 217, 753, 754]. fold [2072]. follow [1317, 2086, 3168, 25, 164, 996, 2491, 1507, 2783, 1303, 1130, 3378]. follow-up [1317, 2086, 3168, 25, 164, 996, 2491, 1507, 2783, 1303, 1130, 3378]. following [1487, 1240, 282, 1491, 2603, 1462, 1008, 2722]. Food [1221]. force [2016]. Forecasting [2261, 2529, 608, 2718]. forensic [2993]. forerunner [649]. Foreword [671, 995, 148, 1282, 737]. forgotten [2537, 2788, 2887, 2787]. form [2321, 1821, 2278, 2618, 809]. form-36 [2321]. Formal [275]. formula [1519]. formulae [2, 2832, 815]. Formulating [3093, 77, 14, 78, 3219]. formulation [272]. forward [1514, 1149]. Foster [1156]. Foundation [359]. Four [2072, 355, 477, 16, 806, 389, 354, 2741]. four-armed [355, 16, 354]. Four-fold [2072]. four-period [806]. Fourth [2842]. fr [884]. fraction [2816, 1112, 1215, 99, 2011, 478, 1947, 491, 2569, 1407]. fractional [1051, 3286, 1889, 1015, 1464, 3138, 2016]. fractions [2728]. frailties [924]. Frailty [3086, 2105, 2169, 3344, 3375, 2566, 3293, 2557, 1796, 2813, 2493, 1680, 1318, 1994, 2024, 3001, 2222, 478, 2941, 2205, 1267, 633, 3050]. frames [3211]. framework [3153, 2768, 2599, 2899, 2204, 1776, 2564, 1238, 1512, 2210, 2565, 325]. frameworks [2914]. Framingham [159, 1299, 1401]. France [603, 401, 119]. Francois [417]. Frank [1157]. Franke [1340]. Frankfurt [343]. free [1518, 1855, 3285, 56, 282, 2903, 2541, 1887, 2711, 264]. Freedman [1598]. freedom [2095, 2272]. Freidlin [2183]. French [40, 2570, 3214]. frequencies [1111, 2072]. frequency [1660, 2837]. Frequentist [2586, 2100, 2668, 2561]. frequently [3030]. Frey [542]. Friede [1589, 3385]. Friedman [1310]. Frison [297]. fronto [1297]. fronto-temporal [1297]. Fuerst [125, 10]. Fujian [1206]. Full [1211, 3032, 2583]. Full-likelihood [1211]. full-scale [2583]. Fully [3298, 3363]. function [53, 651, 1651, 1674, 2120, 3040, 2289, 252, 1359, 2777, 1369, 1491, 724, 1299, 250, 765, 449, 722, 2506]. Functional [2181, 1821, 753, 754, 2337, 2508, 2261, 3360, 2882, 166, 150, 2603, 939, 2618, 2556, 2505, 2692]. functioning [1291, 169, 171]. Functions [102, 1861, 2513, 279, 2286, 2538, 3016, 1177, 2206, 2557, 3129, 3291, 1764, 1660, 3036, 1421, 3046, 3365, 291, 1401, 2988, 489]. fundamental [988]. Fundamentals [1546, 872]. Funnel [1675]. further [852, 732]. futility [1732, 1804, 2156, 3154, 1020]. Future [129, 319, 3314, 1914, 3095, 3315, 1226, 654, 3313, 3319, 3316, 3317, 3318, 2161]. fuzzy [2029].

G [3105, 1364, 2499, 3070, 129, 199, 3190, 1461, 2134, 92, 1850, 3380, 1142, 1118, 494, 2770, 2855, 537, 2559, 1799, 141, 257, 3231, 437, 815, 1547, 1517, 197, 759, 2839, 842]. G. [3381]. Gad [210]. gain [864]. Galecki [2841].

Gallin [1065]. Gallois [2819]. Gamerman [93, 454, 2858]. gamma [2105, 2538, 3375, 3293, 3001, 2222, 2941, 1267]. gamma-frailty [3375, 1267]. Gani [885]. Ganju [3108]. gap [2609, 1421, 1623]. gaps [2705, 2633]. Garrett [2502, 1516]. Garthwaite [1937]. Gastric [1808, 1051]. Gastwirth [543]. Gatekeeping [1146, 1691, 2402, 2875, 2520]. gauging [1942]. Gaussian [113, 2522, 2813, 1721, 2852, 2139]. Gaussianization [2284]. Gaussianization-based [2284]. GBP [1415, 1461]. GEE [1679, 3113, 74, 1062, 1788, 2947, 1378, 479, 1657]. Geert [2409]. Geisser [510]. Gelman [1533]. Genant [125, 10]. gender [1643, 441, 901]. Gene [1409, 1748, 2125, 1382, 1456, 674, 1827, 2769, 1269, 1560, 940, 3119, 3295, 1516, 2214, 2269, 1005, 1347, 2390, 2904, 2827, 2133, 1570]. gene-association [1560]. gene-disease [2269]. gene-drug [2769]. gene-environment [2125, 674, 1347, 2827, 1570]. gene-gene [1382]. gene-gene-environment [2214]. gene-treatment [1456]. General [1188, 563, 1076, 423, 2117, 140, 840, 368, 488, 1463, 1940, 1976, 3381, 2782, 2744]. generalizability [2747]. generalization [3253, 193, 283]. Generalized [2105, 1339, 95, 1014, 1012, 1437, 3237, 1400, 2432, 1572, 3352, 1633, 1674, 490, 134, 2538, 2939, 3157, 627, 191, 99, 1858, 3146, 2861, 2733, 3177, 533, 684, 3019, 2850, 917, 2343, 2208, 977, 531, 958, 1125, 775, 321, 2084, 2556, 1523, 3165, 1512, 2210, 3267, 2025, 1435, 1722, 1837, 2594, 2296, 1045, 2342, 2901, 2892, 1449, 133, 178, 320, 2404, 1460]. Generalizing [2059, 2756]. generate [2810]. generated [1899]. Generating [1720, 2553, 2052]. generation [2162, 3354, 1913, 2726]. generic [3354]. genes [2131, 2204, 2840, 2709, 2707, 3294, 2839]. Genetic [120, 1284, 3089, 3241, 2766, 3118, 3287, 2958, 863, 1548, 196, 2826, 3361, 2126, 1887, 1110, 1304, 2128, 2686, 2974, 1711, 2620, 1763, 2825, 2132, 2904, 3102, 1167, 1653, 2129, 2130, 2995]. Genetics [1351, 1456, 525, 505, 2123, 319, 145, 510]. Geneva [884]. GenHAT [1456]. genital [2815]. genome [2124, 2390]. genome-wide [2124, 2390]. genomes [1934]. genomewide [2769]. genomic [3007]. Genotype [3019, 2125, 1456, 1178, 1166, 1642, 2828]. Genotype-adjusted [3019]. genotype-phenotype [1642]. genotypes [3091, 2686]. genotypic [982, 2544]. geo [2083]. geo-referenced [2083]. geoadditive [1638, 2244]. geocode [2936]. Geoff [2719]. Geoffrey [2452]. Geographic [2693]. Geographic-based [2693]. geographical [854]. Geographically [1797]. Geography [259, 1962]. geometric [2040, 3052, 3288]. geometrically [1164]. George [2604]. Geostatistical [3275]. Gerber [1156]. geriatric [2419]. germ [3256]. Germany [1182, 525, 1517]. Gerstman [103]. gestational [637]. Getting [2127]. Geys [1547]. Gibbs [217]. Gicquel [2819]. Gilbert [3070]. Gilthorpe [2450]. Gini [2854]. Giovanni [1516]. Girling [1531]. GIS [2916, 233]. GIS-based [233]. given [3283, 1602]. Glaz [1384]. Glidden [2140]. GLM [2431]. Global [3300, 2691, 966, 259, 2571, 3051, 2998, 1173, 169, 1260, 662]. Globocan [258]. glucose [1772, 2354]. glycoprotein [1094]. Glymour [1133]. Gnann [1925]. go [473]. goal [229]. Godalming [419]. Goehring [1728]. going [3097, 1731]. goitre [854]. gold [876, 125, 10, 766, 1676, 520, 3321, 6, 2997, 1933, 1818]. Goldberg [714]. Goldstein [1753, 820]. Gönen [3139]. Good [2138, 1311, 1725, 986, 1981]. Goodman [857]. Goodness [880, 1348, 278, 1891, 810, 563, 3300, 1679, 2935, 966, 1587, 1154, 1408, 677, 2782, 133, 960]. Goodness-of-fit [880, 1348, 278, 810, 563, 3300, 2935, 966, 1587, 1154, 677, 2782, 960]. Gopal [284]. Gordon [2501, 1937]. Gower [43]. grading [205, 983]. Graduates [286]. graft [1191, 1576, 1796]. graft-versus-leukaemia [1576]. Graham [1117, 1545, 453, 1669, 186]. Grambsch [526]. graphic [70]. Graphical [878, 3126, 596, 875, 249, 3143, 1072, 1279, 1320, 968, 217, 1739, 1728, 409, 1243, 1416]. Graphics [524, 1181, 524, 400]. graphs [597]. Graubard [438]. Gray [722]. greater [2786, 620]. greatest [1230]. Green [342]. Greenfield [286]. Greenhouse [1217, 1220, 1223, 1222, 1219, 1218, 1221]. Greenland [91]. Greenwood [2064]. Gregoire [129]. Grootendorst [2452]. Group [289, 2570, 2155, 1777, 1229, 1023, 555, 683, 1967, 691, 513, 352, 1613, 697, 1487, 3340, 3300, 367, 2329, 2586, 627, 3338, 1742, 2177, 1966, 1141, 1788, 1983, 822, 1228, 35, 1663, 2419, 958, 2147, 2611, 2832, 1265, 2368, 466, 1927, 1878, 2572, 1380, 1006, 705, 3067, 3029, 1946, 517, 2353, 2645, 1387, 541]. group-level [697]. group-randomized [466, 1927]. group-sequential [1265, 2368, 705, 3067]. grouped [1113, 2355, 1931, 1407]. grouping [3043]. groupings [1860]. groups [3320, 1316, 1428, 1111, 947, 3189, 2055, 980, 1081, 1269, 812, 338, 3201, 2214, 429, 1644, 1865, 919, 2922, 1057, 1989]. Growth [1308, 2120, 2979, 2694, 2407, 37, 1672, 777, 2285, 611, 1668, 1226, 629, 1106, 678, 1585, 3341, 1164, 1251, 390, 1996, 450, 2732, 1870, 420]. Guiard [494]. Guidance [286, 265, 273]. Guide [1158, 452, 1311, 1156, 440, 1145, 419, 115, 1750, 1802, 2841, 3231, 1193, 2754, 3204, 2409]. guideline [2020, 2021, 975, 1228, 569, 974]. guidelines [2245, 2019, 905, 2765]. Gunnarsson [1053]. Gustafson [1609].

H [3135, 1649, 1565, 542, 1374, 2820, 125, 497, 782, 1245, 1925, 386, 221, 1169, 2770, 440, 106, 1937, 2502, 1644, 298, 820, 1547]. H-J [386]. H.-C [2820]. H.-H [1374]. HAART [2770, 2688]. Håvard [2139]. Häävikko [3355]. Haberman [143]. haematological [26]. haematology [2420, 1893]. haemodialysis [1242]. haemoglobin [1399]. Haemophilus [44]. haemorrhagic [2041]. Haenszel [845]. Haggett [259]. Hájek [435]. Hal [1533]. Hall [1594, 341, 1460, 93, 198, 601, 2824, 1425, 1364, 3203, 1751, 143, 200, 1873, 362, 90, 2004, 1119, 42, 961, 3311, 3312, 541, 3057, 1205, 342, 844, 439, 2399, 1533, 104, 780, 130, 3180, 2959, 1156, 821, 1609, 3347, 2960, 1840, 1437, 2858, 43, 2005, 2139, 2719, 2295, 3122, 2841, 2345, 2034, 1547, 1610, 1504, 105, 2720, 184, 1839, 1503, 2755]. Hall/CRC [1594, 341, 1460, 601, 2824, 1425, 3203, 1751, 1873, 362, 2004, 1119, 961, 3311, 3312, 3057, 1205, 844, 2399, 1533, 780, 3180, 2959, 821, 1609, 3347, 2960, 1840, 1437, 2858, 2005, 2139, 2719, 2295, 3122, 2841, 2345, 2034, 1610, 1504, 2720, 1839, 1503, 2755].

Halloran [510]. Hampshire [2006]. Hand [360, 43, 1825]. Handbook [542, 844, 1751, 1839, 2755, 1786, 1351, 1503]. Handling [1383, 3073, 2911, 3066]. handsearching [791]. Hankey [3383]. Hansheng [3122]. haplotype [1767, 3043]. Harbord [2839]. hard [590]. hard-copy [590]. Hardback [1787, 259, 1549, 1753, 1671, 3056]. Hardcover [1415, 1352, 1873, 1802, 3182, 2503, 2140, 3180, 1801, 1840, 2139, 3122, 2718, 1684]. Hardin [1460]. Hardy [1844, 2269, 1910]. Hardy-Weinberg [1844, 2269, 1910]. Harel [2437]. Hargrove [301]. harmonically [688, 1056]. Harrell [1157]. Harry [10]. Harvey [1753, 1802]. Hastie [1310]. Hate [870]. Haudiquet [975]. having [652, 2620]. Haybittle [1983]. Hays [144]. haystack [2736]. hazard [2344, 53, 2538, 795, 793, 2557, 1369, 2882, 954, 2298, 2256, 660, 2326, 3008, 2107, 794, 1455, 3274, 507, 3345, 2081]. hazards [2086, 1613, 1720, 2052, 689, 2606, 3083, 1762, 651, 3125, 2886, 1009, 2942, 928, 2691, 2289, 1177, 558, 3244, 2223, 1895, 1519, 1884, 2256, 2906, 2813, 3018, 924, 1620, 942, 2107, 1808, 834, 507, 2714, 1124, 2696, 109, 1890, 317, 1972, 2762]. Hazim [1158]. Hbk [1549]. HCV [2970]. head [697, 121]. headbanging [230]. headlines [464]. Heagarty [1532]. Health [817, 3289, 1919, 460, 2570, 3022, 183, 399, 420, 1194, 438, 359, 1502, 2350, 1667, 820, 871, 1643, 3352, 1752, 232, 3212, 2836, 751, 473, 497, 2915, 2599, 1833, 2916, 301, 1492, 430, 1518, 2723, 2918, 463, 2267, 2176, 111, 1794, 247, 1957, 1626, 1628, 1081, 2913, 2921, 884, 451, 1239, 1660, 3211, 2393, 694, 642, 1535, 1627, 2362, 3038, 2359, 1115, 2358, 1623, 1831, 2435, 236, 3099, 2215, 1666, 2171, 1824, 2917, 2365, 721, 480, 493, 3049, 399, 417, 1698]. health-care [1752]. health-related [3212, 1628, 1831, 1666]. health-services [3099]. Healthy [1908, 2723, 1930, 2916]. heaping [2900]. hearing [1307]. Heart [1908, 1541, 1094, 432, 2616, 293, 753, 754, 837]. Heckerman [1608]. Hedibert [2858]. Heejung [3136]. Heidelberg [434, 1182, 1181, 2140, 525, 285, 1517]. Heisterkamp [2770]. Heitjan [2134]. Held [2139]. Helen [358]. Helicobacter [1430, 1900]. Helio [454]. help [462, 2033, 1761, 1775]. hemodynamic [2506]. hemophiliacs [3207]. hepatitis [412, 3087]. hepatocytes [3271]. hepatotoxicity [2484]. heritability [2493, 1267]. Herman [2787]. herpes [559, 1925]. Herrendörfer [494]. Herriot [2418]. Hertfordshire [601, 961]. Hesketh [1751, 844, 2857, 2755]. Heterogeneity [174, 406, 1649, 875, 3061, 3238, 2563, 1428, 3375, 784, 3195, 787, 2097, 2614, 2392, 1884, 786, 3268, 910, 3283, 3251, 2227, 2551, 646, 2761, 2764, 2370, 2434, 1686, 3279, 1122, 2233, 2511, 2829, 2957, 1091, 759]. heterogeneous [1273, 3157, 770, 2306, 659, 1805, 2351, 1962, 2305]. heterogeneously [1886]. heteroscedastic [3003]. heteroscedasticity [2667, 1557]. heuristic [1334]. Hew [2006]. HGLM [1640]. HGLMs [1994]. Hidden [679, 1769, 2527, 1003, 294, 1092, 2782, 3117]. Hierarchical [1106, 1095, 879, 321, 2796, 1025, 2647, 2684, 3240, 645, 2864, 2766, 2795, 3287, 2554, 1213, 2216, 2042, 659, 1574, 684, 2564, 1820, 1932, 686, 2173, 591, 2236, 2555, 3100, 3306, 951, 2365, 2670, 2729, 227, 535, 1470]. hierarchically [2402]. High [1660, 3359, 3131, 2946, 1185, 3044, 2319, 523, 1488, 1500, 3110]. high-dimensional [2946, 3044, 1488, 3110]. High-frequency [1660]. high-order [1185]. high-resolution [1500]. high-throughput [3359]. highly [471, 3229, 3377, 102, 1092]. Hilbe [1460]. Hill [305, 962]. Hip [3263, 1257, 2587, 1513]. hippocampus [1298]. Hisayama [2882]. histograms [2556]. histological [2567, 326]. histology [663]. Historical [259, 2113, 2792, 2333]. histories [933, 219, 488, 132]. history [1565, 597, 3187, 177, 2076, 2213, 1872, 2424, 2271, 2360, 2620]. HIV [2770, 640, 352, 1127, 45, 522, 3104, 564, 1079, 1076, 1525, 116, 1178, 2967, 829, 2544, 639, 177, 2317, 2376, 350, 1011, 252, 2979, 1332, 2688, 1661, 294, 2973, 1137, 1004, 307, 3207, 2965, 1834, 2299, 1090, 2300, 333, 32, 1454, 999, 2634, 773, 2318, 2970, 2976, 835, 2253, 2974, 507, 3214, 1618, 1896, 650, 2980, 1583, 2927, 2301, 2524, 1798, 1500, 1934, 2978, 521, 2081, 500, 725, 952, 2191, 2462, 2192, 2969]. HIV-1 [1076, 1011, 2299, 1090, 2976]. HIV-2 [1011]. HIV-infected [294, 773, 1798, 3207]. HIV-RNA [2317, 1332]. HIV-specific [1090]. HIV/AIDS [352, 2967, 177, 2979, 2965, 32, 1618, 2927, 2301, 2192]. HLA [1111]. Hoboken [1530, 3222, 2856, 2502, 2859]. Hochberg [782]. Hodder [1825]. Hoffman [3302]. Holder [2499]. Holdsworth [600]. Holford [141]. Hollander [147]. Holm [2986]. home [1923, 2175]. homeostasis [1772]. homes [2050, 1867]. homogeneity [3135, 2804, 1557, 276, 1949, 2012]. homogeneous [934, 2236, 2435, 1860]. Hong [1305, 1321]. Honor [94]. hormone [848, 560]. hormones [1164]. Hosmer [819, 182]. hospital [2725, 673, 1821, 1186, 3037, 1470, 2722, 3185, 951, 3088]. hospitalization [1096]. hospitalizations [1049]. host [725]. hot [2633, 972, 2632]. hot-deck [2633, 2632]. hotspots [808]. Hougaard [566]. household [1977, 1539, 1496]. Howard [1156]. HPV [1130, 1559]. Hsu [198]. Hu [2856]. Hua [2437]. Hubben [2770]. Huet [1750]. Hui [1561, 10, 1517]. Hul [125]. Human [145, 1550, 525, 2020, 2021, 1220, 71, 413, 2123, 666, 3103, 1298, 1164, 3214, 1934, 1130, 2732, 545]. hunting [2325]. Huntington [2132]. Hurricane [2916]. hybrid [1651, 2668]. Hygiene [306]. Hypertension [1456]. Hypotheses [1205, 1497, 585, 1983, 3014, 993, 1485, 513, 2138]. Hypothesis [2692, 3003, 289, 1188, 2589, 2, 629, 2087, 2208, 743, 560, 2128, 815, 742, 1435, 1722, 2727].

I/II [2985]. Ian [2295]. IARC [258, 131, 437, 401]. Ibrahim [1517, 759]. ICH [569, 975, 974, 905]. ICIDH [427]. idea [1911]. Ideas [524]. identical [1275, 1037, 1450]. Identifiability [228, 1816]. Identification [792, 2510, 1705, 725, 1203, 3146, 2983, 1190, 971, 668]. identifier [1976]. identifiers [1746]. identify [791, 3119]. Identifying [1031, 2973, 1687, 67, 2131, 2838, 3304, 2394, 523, 1334]. if [801, 1053]. IgG [1000]. Ignorability [2763]. ignorable [3272, 1768, 1161, 1770, 423, 2079, 162, 1288, 1883, 968, 3037, 1476, 1451, 1540, 2405, 1028, 761, 3147, 2104, 2080, 2277, 2852, 1289]. ignoring [2680]. II [2160, 405, 1283, 2331, 2158, 1487, 2411, 3340, 3065, 832, 947, 2088, 2479, 2777, 1314, 2673, 3164, 2985, 1663, 1789, 1832, 407, 754, 2712, 2577, 2259, 3279, 797, 823, 2136, 1025, 2265, 2829, 3158, 3066, 1427, 3194, 437]. II/III [3164, 2577]. III [1158, 833, 996, 3164, 2577, 1021]. IIIa [1094]. IL6 [3271]. illicit [1196]. illicit-drug-using [1196]. illness [1287, 1077, 3249, 3266]. illness-death [1287, 1077, 3249]. Illustrated [1193, 1520, 230]. Illustration [2702, 473, 367, 1042, 3293, 934]. ILSI [545]. IMA [525]. image [2255, 1370]. images [590]. imaginary [2178]. imaging [1346, 1520, 1297, 1298, 2507, 1243]. imbalance [202, 757, 3307, 1260]. imbalances [1806]. immigration [1843]. immigration-death [1843]. immune [300]. Immunization [1153, 1631]. immunodeficiency [3214]. immunologic [2376]. immunological [564, 2010]. immunotherapy [2264]. Impact [1964, 2680, 2890, 2318, 2574, 653, 3083, 1079, 3210, 99, 247, 165, 202, 1773, 569, 2850, 757, 1622, 2925, 971, 828, 1457, 1705, 1653]. Impacts [1609, 3004, 2917]. impairment [1822]. imperfect [3155, 1771, 520, 27, 1843, 1092, 340]. imperfectly [1002]. Imperial [1067, 2109]. Implementation [1153, 369, 351, 2442, 884, 28, 1887]. Implementing [2578]. implication [2204]. Implications [1110, 1632, 2915, 699, 5, 2112, 297, 368, 2936]. importance [1913, 1096]. Important [432, 2618, 2573]. imprecision [1366]. improve [2865, 3118, 251, 1581, 1312]. Improved [1487, 979, 1718, 1172, 2208, 1277, 776, 1619, 2732, 1639, 3187, 3338, 2182, 514, 700, 2032]. improvement [3134, 2662, 3133]. improvements [2644]. Improving [1498, 564, 3089, 2242, 2063, 2319, 135, 3228, 2458]. Imputation [554, 353, 1488, 2831, 1918, 53, 2324, 1770, 2284, 2923, 2828, 3002, 3075, 1522, 2437, 2187, 2397, 2442, 2438, 2168, 2283, 3129, 2861, 581, 2061, 2633, 2994, 3076, 1502, 773, 3363, 20, 972, 2072, 3198, 3036, 1083, 2632, 2980, 1754, 445, 685, 500, 2115, 480, 3047, 3162]. Imputation-based [2831]. imputed [2027, 2862]. Imputing [1456, 3247]. inappropriateness [682, 1589]. INCA [2838]. Incidence [1287, 258, 131, 437, 239, 1792, 2030, 44, 159, 522, 158, 3252, 2915, 116, 1305, 119, 176, 2873, 3086, 1856, 162, 164, 3291, 165, 2965, 3171, 234, 1990, 1958, 1176, 1256, 334, 3036, 2253, 2713, 3046, 3214, 1618, 1831, 109, 2927, 2108, 521, 2988]. incidences [3239, 1921]. incident [3229]. incidents [2360]. include [3077]. included [3204, 3193]. includes [301]. Including [400, 1511, 159, 1072, 23, 1034, 1458]. inclusion [1968, 177, 771]. Incomplete [104, 2324, 2864, 1249, 2284, 2923, 2085, 1004, 2705, 581, 2061, 1052, 3074, 2264, 773, 1820, 1851, 1612, 138, 2027, 1488, 1586, 3145, 1210, 3290, 957]. incompletely [3273, 411]. inconsistency [3238]. inconsistent [3131]. inconsistently [3031]. Incorporating [2893, 2304, 2028, 3018, 2741, 2390, 1591, 522, 732, 659, 1889, 1605, 3091, 2580, 1021, 1358, 2792, 2077]. increase [2895, 2806, 619]. Increasing [1353, 2401]. Incremental [484, 997]. incubation [1625, 1855, 396, 33, 2688, 1781, 2770, 3292, 650]. Independence [1391, 2125, 1899]. Independent [571, 1302, 2225, 1387, 473, 736, 3112, 1388, 1386, 292, 3298, 3021, 429, 1810]. Index [2650, 1943, 1892, 2952, 191, 1106, 2134, 2213, 38, 2699, 1755, 271, 1502, 2198, 408]. India [854]. indicator [3212, 3289, 1123, 3117]. indices [432, 56, 194, 2273]. Indirect [3254, 3384, 1974, 1506, 846, 2316, 1917]. Individual [697, 716, 1870, 1163, 564, 2766, 1035, 269, 2624, 3184, 272, 1855, 262, 265, 546, 273, 1185, 274, 2063, 267, 266, 1333, 38, 977, 2832, 270, 2580, 2526, 2760, 3062, 2761, 263, 2434, 1686, 31, 2679, 547, 268]. individual- [2679]. individual-level [1163, 2063]. individual-subject [1333]. Individualized [2553, 2894]. Individualizing [2375]. individuals [731, 3238, 2779, 294, 809, 2418]. Indonesian [2120]. induce [927]. Induced [2287, 2204, 911]. Industry [905, 1182, 1389, 308, 2345]. ineffectiveness [1020]. inequalities [2915]. inequality [1231]. Infant [2407, 639, 2699]. infants [99]. infarction [2907, 910, 1470, 2539, 2908, 2722, 3137]. infarctions [1737]. infected [1332, 519, 3207, 1090, 294, 773, 1798]. infection [44, 3104, 2606, 666, 2544, 252, 3207, 1834, 1900, 999, 1744, 2253, 1843, 507, 2016, 1092, 2980, 1128, 1583, 1130]. infections [1792, 328, 773, 1256, 3309]. infectious [2000, 1795, 2241, 2216, 21, 3361, 3069, 1499, 308, 3239, 2301, 1457]. infective [987]. infectives [1273]. infectivity [1011]. Inference [454, 225, 2758, 3112, 1001, 1310, 1842, 436, 130, 2207, 3046, 1972, 579, 3323, 563, 1868, 1951, 2550, 926, 2425, 795, 793, 2740, 3016, 99, 3007, 2063, 2851, 2840, 299, 2307, 333, 238, 1680, 531, 3341, 2202, 3237, 1637, 2689, 2868, 794, 2745, 1367, 1729, 3365, 1818, 1504, 2103, 1759, 1602, 1406, 3290, 2648, 93, 434, 843, 2858]. Inferences [1760, 1837, 2054, 2594, 2296, 349, 2058, 2709, 1445, 2885, 2839]. inferential [2850]. Inferenz [343]. inferiority [984, 538, 3041, 2020, 2245, 2313, 1578, 1037, 2019, 1482, 1980, 2552, 985, 1982, 1703, 992, 2652, 1986, 1041, 2658, 2667, 988, 213, 2249, 2653, 1776, 1983, 986, 1981, 2749, 3385, 687, 1987, 1199, 1313, 2997, 3362, 3225, 987, 1034, 993, 1528, 990, 1984, 2896, 2159, 2668, 1485, 991, 2655, 2656, 2657, 2669, 513, 989, 1810, 1404, 1979, 3382]. infinite [3084]. inflated [777, 2180, 2120, 3354, 1837, 2341, 2892, 594]. inflating [2401]. Inflation [1361, 860, 2893, 1481, 1394, 2023]. Influence [1655, 1824, 1258, 2420, 616, 2047, 1455, 785, 172, 1885, 2892, 825]. influenced [2994]. influences [384]. influencing [3064]. influenza [1539, 2037, 2038, 954, 2949, 879, 2035, 2039, 2036, 3031]. influenzae [44]. informant [1493]. informants [2991]. Information [434, 2145, 17, 3243, 877, 639, 1676, 732, 3226, 1694, 884, 3361, 3343, 760, 577, 1605, 3227, 2118, 1336, 874, 255, 2581, 2359, 2516, 1303, 3358, 397, 1730, 2390, 425, 1854, 227]. Information-based [2145]. Information-Theoretic [434]. informative [2931, 2088, 1883, 1433, 2202, 690, 383, 422, 2574, 2811, 335, 1583, 703, 2671, 1692, 2406]. informatively [2207, 2615, 2385, 3046]. infrequent [1559]. infusion [282]. ingredients [2334]. inherent [1208]. inheritance [2686]. initial [2450, 2541, 2260]. initiate [1525]. injecting [507]. injury [2564, 594]. innovation [578]. Innovations [2570]. Innovative [908, 2201]. Inputs [542, 2833, 1239]. inseparable [246]. insertion [557]. inspections [1374, 655]. instability [59]. Institute [3139, 2604, 782, 3204, 545]. institutional [3224, 2377, 1675]. Instruction [1245]. instrument [554, 2826, 340]. Instrumental [2332, 2297, 1061, 2733, 2826, 3136, 3283, 2825]. instruments [2840, 2709, 2825, 2839]. insulin [1772, 1820, 2354]. insulin-dependent [1820]. Insurance [460, 2913, 143]. intake [998]. intakes [552, 1312]. Integrated [454, 2781, 2335, 1059, 2644]. Integrative [2110]. integrity [1390]. intensity [381, 704, 2746]. Intensive [881, 328, 878, 1369, 972]. intent [2297, 1578, 3136, 1984, 1985]. intent-to-treat [2297, 1578, 3136, 1984, 1985]. Intention [3027, 3283, 669, 3142, 2710]. Intention-to-treat [3027, 669, 3142, 2710]. intentional [1658]. Inter [1923, 2554, 1690, 673, 2294, 2354, 2014]. inter-hospital [673]. Inter-rater [1923, 2554, 2294, 2014]. inter-temporal [2354]. interaction [1382, 2629, 272, 2847, 674, 1779, 3170, 331, 2769, 2492, 592, 2846, 1702, 2318, 346, 98, 2214, 206, 172, 873, 1329, 2827]. interactions [2332, 249, 3186, 1944, 2768, 2125, 155, 1456, 1749, 2534, 2084, 1464, 156, 901, 1934, 1347, 489, 1570]. interactivity [1181]. intercept [2375]. intercourse [71, 2367]. interest [1389]. interfacing [1455]. interference [2509]. Interim [1097, 345, 2675, 1809, 1374, 1905, 1387, 1353, 2065, 2330, 655, 2577, 2146, 2741]. intermediate [941, 1613, 1718, 3141, 2741]. Internal [2522, 1152, 96, 688, 1056, 216, 340]. internally [2351]. International [1082, 2570, 1608, 603, 569, 1568, 1876, 2212, 437]. interobserver [3127, 2034]. interpoint [1641]. interpolation [176, 562]. Interpretability [350, 135]. Interpretation [42, 1245, 544, 2320, 2448, 604]. interpreted [788]. interpreting [2489, 2323, 2476, 2336]. interrater [800, 938]. Interscience [1530, 2502]. Interval [75, 2209, 915, 1359, 2068, 3321, 917, 1150, 1724, 1681, 2446, 1100, 2933, 2526, 2274, 1722, 2628, 2062, 2874, 1168, 2993, 53, 941, 1554, 689, 965, 3051, 3079, 1103, 2695, 1204, 1636, 2592, 351, 755, 2276, 2608, 617, 1522, 1001, 2206, 2283, 2861, 274, 474, 2070, 1835, 956, 2627, 1781, 1509, 2646, 724, 2674, 112, 1893, 503, 1371, 3301, 250, 20, 1434, 3172, 2896, 507, 920, 2853, 449, 1864, 1128, 2561, 1435, 1527, 964, 1494, 2843, 2192, 2406, 1400, 2441, 2252]. Interval-censored [1359, 965, 3079, 1636, 2592, 351, 2276, 2608, 1522, 1001, 2861, 956, 1509, 724, 1893, 507, 2853, 449, 1128, 1527, 2192, 1400, 2252]. intervals [1639, 2865, 2919, 2122, 2178, 1031, 13, 2292, 115, 1487, 736, 2386, 3112, 1033, 1109, 1359, 1971, 1405, 492, 1526, 1850, 1834, 2428, 3053, 3298, 1990, 11, 277, 1617, 7, 2611, 2384, 960, 2855, 2455, 9, 1175, 1940, 1941, 2559, 3059, 46, 3021, 2167, 806, 1507, 1013, 278, 3325, 2032, 479, 3145, 29, 776, 1845, 2934, 2866, 2007, 1263, 2233, 2708, 1102, 1946, 1810, 1619, 3221]. Intervention [1257, 1866, 852, 473, 367, 1894, 2266, 3170, 1842, 1293, 1616, 2588, 498, 1129, 2412, 2999, 2845]. interventions [1792, 640, 3286, 1747, 2094, 12, 2563, 770, 921, 2654, 1140, 3099, 1026, 1046]. intra [1690, 2318, 2764, 643]. intra-subject [2318]. intra-tumor [2764]. intra-uterine [643]. Intraclass [2173, 768, 1103, 3157, 617, 1969, 3172, 1859, 1435, 1845, 1722, 964, 1263, 1494]. intracluster [369, 3039, 1366, 2007]. intrarater [800, 938]. intrauterine [678]. intravenous [252]. Intrinsic [1630]. Introduction [602, 1224, 636, 142, 1437, 103, 3222, 2266, 885, 92, 2859, 2720, 1549]. introductory [2245, 2019, 1327, 2399]. intuitive [1943, 2198]. invariant [2407, 3121]. Inverse [3025, 3229, 3024, 795, 793, 794, 1908, 2866, 1828]. investigate [3210, 2017]. Investigating [399, 784, 2761, 1686, 2932, 2998, 1123]. investigation [1970, 697, 1079, 2218, 2543, 2596, 570, 1015, 3353]. investigators [1546]. Invited [738]. involvement [384]. involving [3282, 1883, 534, 548, 595, 555]. Iowa [2555]. Irizarry [1516]. irregularly [332, 136, 661]. Irwig [803]. Irwin [2751, 2488]. Isabel [401]. ISBN [1530, 1594, 91, 341, 819, 817, 1157, 1460, 1647, 93, 1565, 454, 1592, 421, 542, 3139, 198, 509, 3222, 2138, 1752, 1786, 1351, 2754, 601, 144, 455, 360, 2824, 526, 2604, 885, 1425, 434, 1158, 1364, 497, 1608, 286, 3203, 782, 361, 1751, 143, 200, 127, 183, 1750, 1873, 301, 145, 402, 843, 362, 288, 2501, 3204, 1415, 2006, 129, 603, 1566, 1550, 90, 1134, 543, 343, 1402, 358, 600, 2004, 1119, 199, 399, 2856, 42, 1461, 420, 1438, 602, 40, 201, 1787, 961, 3311, 303, 1245, 511, 1180, 1067, 1548, 1753, 131, 3312, 1194, 1182]. ISBN [541, 3057, 211, 1352, 452, 182, 566, 1018, 1195, 1205, 342, 844, 1774, 1802, 3182, 439, 196, 3181, 2399, 1310, 417, 1144, 1181, 436, 92, 438, 2409, 1533, 2503, 1311, 2140, 453, 1159, 1593, 1309, 104, 818, 435, 1133, 884, 495, 1142, 2857, 780, 130, 300, 302, 1118, 3180, 1531, 524, 299, 1595, 2959, 1156, 221, 1671, 821, 1609, 132, 3347, 962, 142, 1872, 1801, 3192, 359, 2960, 781, 357, 494, 1840, 1437, 525, 3056, 1193, 1363, 440, 285, 1143, 1145, 2858, 544, 869, 2519, 872, 419, 1192, 2961, 43, 257, 1546, 2005, 146, 106, 2109, 1471, 2139]. ISBN [2502, 2859, 2719, 287, 2295, 870, 3122, 260, 1648, 545, 1516, 185, 2841, 3138, 3231, 437, 1206, 2345, 128, 2034, 298, 400, 418, 820, 1547, 103, 540, 1517, 1065, 1610, 510, 187, 1504, 2718, 284, 1384, 1825, 1698, 147, 105, 2720, 259, 184, 496, 197, 842, 1066, 186, 433, 210, 1684, 1839, 401, 1549, 1244, 1503, 1416, 871, 41, 94, 2755]. ISBN-13 [3180]. ischaemic [1049, 1479]. ischemic [1382]. Isfahan [1908]. isn’t [890]. isotonic [1209, 2055, 2985, 88]. ISSN [141]. Issue [2504, 833, 3327, 3354, 3328, 3329, 3332, 3326, 2036, 838, 3331, 3330]. Issues [789, 907, 2148, 1461, 3090, 1838, 1039, 2362, 3232, 205, 895, 323, 985, 228, 1295, 1624, 894, 2652, 520, 3181, 307, 988, 2653, 1294, 1513, 583, 2200, 1940, 243, 1208, 239, 584, 3092, 889, 873, 337, 268]. Italian [1278, 3355]. Item [2535, 680, 1291, 554, 169, 700, 662, 171, 427]. items [1293, 3242]. iterative [1117, 1545, 1116, 3301, 1669, 2632, 1322]. ITT [3012, 2101]. IUDs [643]. ix [454, 542, 2754, 1425, 143, 40, 1437, 1066, 401, 2755]. Izikson [3221]. Izu [3108].

J [1530, 91, 538, 3105, 1168, 1649, 1591, 2000, 2327, 509, 1752, 1351, 2820, 144, 360, 885, 1158, 1608, 2907, 813, 841, 2641, 288, 129, 1134, 3190, 2050, 3379, 2033, 420, 602, 258, 2640, 2643, 1245, 131, 297, 1925, 386, 1593, 104, 3108, 1142, 300, 302, 1118, 1531, 714, 1156, 1016, 1563, 142, 1169, 781, 494, 2770, 1437, 1363, 1799, 1532, 2637, 43, 2005, 1937, 2295, 870, 437, 1206, 2198, 418, 540, 3106, 1825, 1066, 3137, 1549, 1244, 2639, 1503, 759, 3382, 2839, 1709]. Jacqueline [342]. Jacques [414]. Jakob [1584, 2268]. James [1460, 538, 2824, 301, 2502]. Jan [1053]. Janet [2501, 1180, 3192]. Jaques [962]. Jaroslav [435]. Jason [198]. Javier [1195]. Jean [417]. Jean-Francois [417]. Jeffrey [1684]. Jen [3203, 201, 495, 287]. Jen-Pei [495, 287, 3203]. Jennison [541]. Jensen [1873]. Jeremy [1156]. Jeroen [132]. Jerome [1310]. Jersey [1530, 3222, 1550, 1801, 2859]. Jesper [1504]. Jewell [1840]. Ji [1561, 3302]. Jia [2345]. Jia-Yeong [2345]. Jim [2295]. Joe [885]. John [3222, 1438, 211, 1352, 342, 439, 3181, 2399, 1533, 1595, 1363, 419, 2859, 2295, 3231, 1065, 2720, 94]. Johnson [13, 1682, 11, 29]. joinpoint [2731, 48, 386]. Joint [2257, 45, 2842, 1505, 1992, 3034, 3167, 442, 2748, 3068, 1124, 1583, 2573, 3082, 2889, 2610, 2426, 2946, 2600, 1000, 2812, 3217, 1032, 851, 973, 2388, 312, 2235, 2259, 1445, 2980, 1736, 1912, 680]. Jointly [765, 1884, 1210]. Jolivet [1750]. Jonathan [1752, 2033, 1648]. Jonckheere [2544]. Jones [2907, 1194, 1362, 1206, 3137]. Jong [1610]. Jonkman [1168]. Jørgensen [200]. Joseph [1460, 543, 1801, 1517, 1384]. journals [314, 537]. journey [377]. Jr [2640, 2642]. Judea [843]. Judith [1595]. Juha [2718]. Julian [2005, 437]. Jun [1157, 2002, 1645, 3122]. June [2452].

K.-Y [1532]. Kalbfleisch [1530]. Kang [538, 3135, 2450]. Kanji [284]. Kaplan [1651, 530, 1580, 970, 1828]. Kappa [830, 563, 75, 51, 2012, 593]. Karger [2961]. Kastner [1340]. Kathleen [2841]. Katrina [2916]. Katsumi [2006]. Katz [440]. Kay [3182, 2295]. Keith [1752, 299, 1206]. Kelly [2000]. Kempson [92]. Kendall [361, 211, 3005, 436, 299]. Kennedy [1563]. Kenneth [91, 434, 1548]. Kent [435]. Kept [103]. Keratoconus [1796]. Kernel [2789, 826]. Kerr [1156]. Kerry [3192]. key [1633]. Keying [10]. Keynes [399]. kidney [1510]. Kieser [1589, 3385]. Kim [1610, 386]. Kimmel [545]. kindreds [809]. kinetic [1306, 1243]. kinetics [2815]. Kinukawa [757]. Kirkwood [1648]. kit [1947]. Klar [817]. Klaus [40]. Klein [1352]. Klinke [285]. Kluwer [421, 1066]. Knott [361]. know [3349]. known [2473, 1855, 1956]. Knox [2339]. Knudsen [1145]. Kobayashi [2006]. Kong [1305, 1321]. König [2820]. Koornwinder [106]. Korn [1597, 438, 2183]. Kozub [1156]. Kramárová [437]. Krause [1182]. Krickeberg [40]. Kringle [3302]. Kristin [1156]. Kruse [1416]. Krzanowski [142]. Kuhnert [3312]. Kulinskaya [3231]. Kullback [1036, 425]. Kunio [2859]. Kuroki [3219]. kurtotic [1501]. Kwon [3382].

L [1530, 3041, 1117, 1545, 2000, 803, 13, 77, 2907, 301, 2499, 603, 543, 1402, 1597, 131, 1352, 1715, 436, 438, 297, 1925, 104, 1714, 11, 1801, 2183, 1532, 260, 1516, 1669, 2198, 418, 1547, 1065, 78, 29, 3137, 94]. laboratory [38, 595, 2786]. Lachin [781, 3106]. Lack [2747, 1746, 1381, 3201]. lack-of-fit [1381]. lag [1738, 1140, 3160]. lagged [625, 3119]. Laird [2502]. Lan [2501]. Landau [1839]. Lang [343]. Lange [1548, 196]. Language [419]. Lanke [1053]. large [551, 1103, 2915, 2033, 1107, 913, 1294, 1229, 1575, 1299, 727, 3023, 1775, 2746, 1329, 2337, 721]. large-sample [1103]. large-scale [1299]. larger [1715, 1714, 1716, 1713]. Larson [3190]. Lasso [2495]. Last [1514, 1149, 2123]. Late [981, 2059]. late-onset [2059]. latency [941]. Latent [1814, 2758, 1491, 384, 1376, 2303, 2815, 1658, 3128, 1188, 2463, 1923, 766, 2600, 2991, 1321, 3338, 136, 2699, 701, 2950, 1439, 2388, 3068, 1114, 2574, 2759, 2277, 3242, 1253, 455, 361]. Latent-variable [2303, 3338]. Latouche [2344]. Lau [1649]. Lauder [2295]. Lavoisier [185]. Lawlor [2839]. Lawson [1364, 417, 1144, 1054]. Lead [1187, 980]. Lead-time [1187]. leadership [1386]. leadtime [1525]. Learning [1310, 747, 748, 3337, 3204]. least [986, 1981, 2067, 3110]. LEBEL [914]. LeBlanc [1016]. Lecture [129, 305, 2842, 3296, 2075]. Lectures [1534]. Lees [1563]. left [1848, 3022, 2952, 1558, 595, 333, 3023, 1583]. left-censored [1848, 3022, 1558, 333]. left-censoring [1583]. left-truncated [3023]. leg [611]. Legedza [759]. Legionella [1256]. Lehmann [2087]. Leibler [1036, 425]. Leiden [3103]. Leishmaniasis [551]. Lelia [1159]. Lemeshow [819, 182, 496]. Lemuel [2665, 3180, 1546]. length [909, 1424, 2725, 611, 3255, 3227, 951]. lengths [1432]. Leonhard [2139]. Lesaffre [1696, 417]. lesion [1769]. lesions [852]. Lesley [1531]. less [1198]. lessons [1502]. lethal [475]. Letter [538, 1117, 1168, 1545, 1649, 1591, 2002, 1951, 2000, 2245, 2344, 2379, 1424, 2910, 355, 1374, 803, 2052, 125, 13, 39, 77, 2735, 2439, 414, 813, 1696, 1784, 1708, 2499, 3384, 10, 1682, 2450, 2050, 1645, 2033, 181, 1337, 1589, 387, 2437, 2134, 1715, 297, 1850, 1925, 3380, 1606, 1714, 1054, 1056, 867, 11, 714, 2558, 1016, 2452, 757, 1563, 958, 1362, 960, 1169, 1728, 416, 537, 1725, 2183, 1851, 1799, 2186, 1669, 815, 1711, 857, 2198, 2822, 1716, 1324, 1644, 2032, 3381, 78, 354, 2160, 29, 1053, 397, 3288, 759, 1450, 3221, 283, 1709, 3303]. leukaemia [2610, 666, 970, 1576, 282, 2555, 1734]. leukaemia-free [282]. level [1126, 1163, 697, 2356, 154, 2050, 2698, 914, 2063, 1574, 3034, 2883, 498, 254, 2354, 2167, 1847, 1667, 1867, 3038, 2844, 2679, 2364, 2506, 1619]. levels [2119, 3070, 2975, 3150, 99, 914, 1827, 774, 1150, 1689]. Levin [1801]. Levy [496]. Lexis [2735, 2734, 2436]. Leyland [820]. Li [3070]. Liang [1532]. Libman [2787]. Library [361, 211, 436, 418]. License [1379]. Lichtenstein [1711]. Life [2262, 144, 3289, 545, 275, 3212, 1768, 2723, 2724, 611, 1320, 2913, 2943, 2013, 2046, 619, 2564, 749, 1901, 683, 3023, 1027, 312, 448, 667, 1303, 3185, 828, 1666, 680, 544]. life-threatening [275]. life-years [2046]. lifetime [159, 3249]. like [1229, 2510]. Likelihood [866, 90, 1960, 2071, 1555, 1102, 1117, 1545, 941, 3104, 2945, 689, 245, 70, 2431, 2255, 2125, 2984, 3132, 2781, 1682, 599, 1116, 2991, 1957, 2951, 836, 853, 1446, 1781, 3053, 231, 724, 851, 2601, 2849, 2591, 1007, 333, 238, 280, 3076, 3098, 2528, 2104, 2976, 1211, 2205, 2761, 3054, 1859, 1669, 1105, 1322, 1345, 2708, 2843, 1854, 3006, 2681, 2459, 2732, 130, 1592, 3312]. Likelihood-based [1555, 1102, 1007, 2708, 2843]. likelihoods [2989]. Lim [2000]. limb [490]. limit [1600, 3281, 3237, 3149]. limited [1626, 2219]. limits [1117, 1545, 1747, 2094, 2802, 1116, 2464, 2591, 1669, 3381, 1322, 2708, 1406, 2744]. Lindsay [191]. Lindsey [288, 1549]. line [1426, 1427, 3215]. Linear [1157, 3324, 203, 1437, 1836, 2257, 998, 1302, 3352, 2945, 3265, 490, 1076, 1672, 134, 1750, 1655, 2375, 1188, 1935, 3063, 627, 2847, 829, 423, 2378, 3077, 3078, 101, 1160, 1001, 2522, 629, 1849, 1106, 1802, 2733, 3177, 1467, 2140, 684, 2903, 1998, 2730, 2850, 1581, 299, 2343, 2106, 26, 958, 1212, 2487, 169, 2534, 608, 2809, 1702, 2318, 856, 686, 973, 2430, 874, 1335, 208, 2572, 2619, 2025, 2808, 2934, 195, 1527, 2342, 2275, 562, 500, 725, 952, 1241, 1854, 2692, 2194, 133, 178, 2404, 2009, 2186, 132, 2005, 2841]. linearization [3243]. linearly [1881]. link [1674, 2907, 2539, 2908, 3137]. linkage [3241, 1767, 779, 3103, 2191, 3042]. Linked [2358, 2977, 244, 461, 475, 1232]. Linking [467, 427]. lipid [1339, 320]. Lippeveld [884]. Lippincott [91]. Lippincott-Raven [91]. Lise [2452]. List [457, 1924, 909, 1424, 1602]. literacy [2926]. literature [2327, 2771, 2775, 1412, 2772, 2773, 31, 2774, 1915]. Littell [2604]. Little [1363, 437]. Liu [3203, 201, 3380, 3310, 495, 287]. liver [66]. lives [2921]. LLC [869]. Lloyd [402]. LMS [3355]. load [1332, 2951, 1137, 307, 999, 1583, 1798]. Local [941, 2047, 1455, 689, 3051, 2508, 2420, 1964, 2134, 1957, 1240, 1755, 2339, 2759, 2852, 121]. localized [1958]. locating [808]. location [2390]. loci [1500, 2191]. locus [1934]. Log [2275, 2814, 2891, 1561, 295, 841, 2847, 1307, 3060, 2861, 1559, 1174, 2121, 776, 2594, 195, 1102, 2062, 1828, 3160, 1400, 132]. Log-linear [2275, 2847, 195, 132]. log-normal [2814, 1307, 2121, 776, 1102]. log-rank [2861, 1559, 1828, 3160, 1400]. log-regression [2594, 2062]. log-skew-normal [2891]. log-transformed [1561, 3060, 1174]. logic [1687]. Logistic [819, 1157, 471, 3277, 431, 1951, 1361, 2432, 2623, 1078, 2292, 2624, 1807, 580, 2731, 2465, 2629, 1679, 2935, 612, 558, 1749, 1060, 1074, 855, 2218, 880, 2338, 1997, 1478, 3376, 2717, 1822, 2140, 966, 2803, 1261, 588, 1761, 27, 1381, 1587, 1050, 47, 1637, 677, 754, 1908, 2053, 340, 31, 110, 1468, 2383, 3064, 3106, 1410, 2219, 1538, 2193]. logit [3252, 548, 3080, 2526]. lognormal [2209, 1837]. logrank [3045, 202, 757, 531]. London [341, 817, 93, 454, 542, 198, 1425, 361, 183, 145, 288, 90, 600, 1119, 42, 1461, 602, 511, 1067, 1753, 211, 342, 844, 439, 92, 1159, 1593, 104, 818, 130, 299, 1156, 821, 132, 962, 142, 357, 1437, 1193, 141, 869, 43, 106, 2109, 870, 128, 1547, 284, 1825, 184, 186, 306]. Long [3035, 1808, 486, 1037, 2166, 900, 928, 1097, 515, 2266, 2587, 849, 898, 996, 2493, 345, 2416, 1542, 734, 888, 952, 2715, 3346]. long- [734]. Long-term [3035, 1808, 486, 1037, 2166, 900, 928, 1097, 515, 2266, 2587, 849, 898, 996, 2493, 345, 2416, 1542, 888, 952, 2715, 3346]. LongCriSP [2325]. longevity [3103]. Longitudinal [751, 2293, 129, 181, 2100, 1502, 150, 936, 3272, 139, 1489, 1290, 3202, 564, 1539, 134, 3273, 3377, 1807, 1768, 2931, 1087, 2923, 2610, 1443, 1673, 154, 726, 2426, 393, 627, 2946, 1505, 423, 2008, 2992, 1520, 2378, 3077, 3078, 3190, 1097, 1288, 101, 3144, 2600, 1320, 1077, 2325, 1252, 3170, 1863, 2134, 1383, 1842, 630, 3217, 1004, 853, 3234, 1777, 2461, 1476, 923, 581, 2943, 3188, 2237, 2415, 2857, 1540, 2061, 2697, 851, 2588, 2730, 2947, 488, 2633, 35, 3074, 3167, 1893, 3148, 1595, 1755, 203, 280, 958, 1605, 749, 168, 1354, 2264, 3363, 1721, 3147, 1532, 2615, 690, 2104]. longitudinal [973, 1027, 2322, 3073, 912, 2388, 383, 3068, 253, 3295, 1470, 170, 2502, 2879, 2045, 2270, 2572, 2786, 422, 623, 3165, 2701, 1115, 2258, 3267, 68, 2053, 2844, 3365, 1006, 667, 2863, 1754, 1583, 153, 2303, 653, 1210, 1486, 2524, 2433, 1912, 723, 1162, 3082, 1406, 2845, 2852, 3290, 2337, 2497, 2692, 2831, 133, 3378]. longitudinally [332]. look [504]. Looking [2736]. Looney [1550, 1362]. Lopes [2858]. Lorenz [2854]. loss [2086, 25, 2768, 1658, 1307, 217, 839]. lost [3032, 2826, 2046]. lot [2897, 3108]. Louis [961, 3347]. low [99]. lower [490, 611, 2986]. lower-limb [490]. lowess [1078]. Ltd [3181, 3231]. Lu [125, 3386, 10, 3380]. Lun [211]. lung [2474, 1738, 2781, 1668, 99, 219, 913, 2117, 679, 2213, 2778, 1334, 2310, 1427]. LVAR [2028]. LX [3383]. Lyle [3057]. lymph [1336]. lymphocyte [282, 1090]. lymphocytic [2610]. lymphoma [1204]. Lyon [603, 258, 131, 401].

M [1460, 1951, 2379, 355, 509, 1351, 144, 526, 39, 1608, 2641, 3384, 362, 603, 1134, 1597, 3379, 258, 1589, 2640, 2643, 387, 131, 92, 1925, 3380, 2642, 2489, 3385, 867, 302, 714, 2959, 1016, 2452, 960, 781, 2770, 416, 2855, 2559, 2637, 544, 419, 1192, 2502, 437, 1711, 2034, 3219, 1547, 510, 2718, 354, 3106, 184, 842, 2639, 2839, 431]. M. [283]. MA [1133, 2134]. Macaskill [803]. Macfarlane [818]. Machin [509, 544]. macro [159]. macroscopic [1298]. magnetic [2255, 983, 3005]. magnitude [553, 1347]. Mahesh [1531]. Main [343, 3208, 340]. maintenance [3250]. Major [2064, 2204, 1393]. Making [2356, 315, 1475, 1628, 1629, 2840, 2709, 28, 2084, 1623, 1390, 2839, 1018, 1245]. malaria [1843, 195]. male [2575, 2815, 1164]. Malley [2820]. Mallick [3380]. mammography [1920, 2336]. management [364, 170, 2595]. Manager [1311]. Managing [1299]. Manfred [197]. Mann [3209, 2096, 2663, 1939, 1940, 1941]. Manning [815]. Mantel [845]. Manual [1245]. many [2293, 461, 617]. map [536, 477]. Mapping [601, 417, 2564, 100, 245, 223, 225, 3169, 230, 248, 2905, 1054, 240, 696, 1991, 2201, 2545, 3196, 1088, 2834, 3005, 523, 1797, 2547, 243, 229, 1910, 235, 206, 1500, 2191]. maps [244, 695, 236]. MAR [1863]. Mara [1610]. Marcel [1438, 201, 495, 1192, 871]. Mardia [260]. Margaret [1787]. Margin [2669, 2020, 2245, 2019, 225, 1980, 3014, 2896]. Marginal [1000, 672, 623, 2443, 3025, 2910, 1331, 279, 3024, 2909, 796, 2557, 2609, 2890, 84, 2304, 1378, 720, 3301, 799, 384, 3365, 2696, 1345, 1426]. Marginalized [1476, 2943, 3188]. margins [1982]. Marie [1750]. Marie-Anne [1750]. Marion [2819]. Marion-Gallois [2819]. Mark [341, 2004, 2450, 2960, 540]. marked [3361]. marker [1916, 1236, 2638, 2641, 729, 177, 2640, 2643, 2642, 2636, 2637, 2639]. marker-based [177]. markers [564, 1914, 3384, 413, 61, 1154, 1260, 1917, 215, 2369, 1945, 2650]. marketing [2786]. Markov [2139, 810, 1769, 2527, 1003, 926, 3022, 2621, 2695, 1768, 228, 393, 2608, 2079, 3260, 1320, 934, 679, 1090, 2405, 2949, 2603, 338, 1376, 1114, 2045, 1259, 2270, 1092, 2435, 650, 2782, 3117, 1091, 93, 2858]. Markus [1726]. marrow [970, 1576, 508, 3035]. Martin [361, 3190, 1180, 1593]. martingale [2132]. mass [1106, 1502, 2360, 1945, 2650, 3004, 663]. mass-screening [3004]. Masuyer [437]. matched [1275, 1639, 3320, 1652, 1151, 992, 1148, 1767, 970, 674, 2898, 2772, 2983, 2803, 1261, 3305, 22, 3343, 2884, 2200, 2456, 720, 3301, 3225, 1614, 98, 1211, 1937, 991, 2032, 1511, 1058, 3106, 1450, 3382, 2130, 327, 489, 1570]. matched-pair [992, 1148, 2983, 720, 3301, 3225, 1058, 3382]. matched-pairs [1275, 2884, 991, 1450]. matches [2374]. matching [2771, 2775, 1151, 2772, 2068, 2773, 1817, 1724, 1681, 2774]. maternal [638, 1280]. maternity [951]. mathematical [666, 2955, 2195, 421, 1548, 196]. mathematician [649]. Mathematics [525, 146]. Mathur [125, 10]. matrices [2695]. matrix [3243, 1087, 874]. Matsuyama [2558, 2911]. Matthew [259]. Matthews [602, 1118, 2961, 2720]. Maurice [144]. max [1167]. Maximally [472]. maximization [2516]. maximizing [2350]. Maximum [2945, 724, 851, 333, 280, 3076, 3104, 245, 2125, 2056, 2991, 836, 231, 2601, 3098, 1100, 2205, 2761, 1859, 1105, 1854, 3006]. maximum-likelihood [3104]. Maxwell [128]. May [1594, 13, 11, 29, 1579, 1046]. Mayer [1671]. MC [3353]. MC-SIMEX [3353]. McCulloch [2140]. McDougall [1195]. MCF [2404]. MCF-7 [2404]. MCMC [1539, 1273, 1773]. McNemar [1448]. McShane [1597]. mean [58, 5, 3144, 853, 3224, 297, 267, 2903, 1175, 972, 776, 1837, 2594, 2296, 2342, 3173, 2569, 1102, 2062]. mean-variance [2342]. meaningful [2743]. means [2814, 1561, 2209, 1188, 1085, 1362, 1659, 1174, 2761, 2121, 1644, 2807, 2219, 919, 493]. measles [850]. measure [139, 15, 75, 2878, 39, 2179, 1794, 3335, 704, 1292, 1544, 2515, 512, 169, 1940, 1941, 2220, 1933, 1434, 135, 1330, 3152, 1543, 2745, 670]. measured [2702, 2281, 3058, 1551, 2555, 1406]. Measurement [3327, 2781, 149, 211, 3328, 3329, 2294, 2879, 1603, 185, 3332, 3326, 3331, 3330, 3324, 553, 3202, 3212, 3289, 1203, 1044, 3174, 1536, 24, 3002, 2175, 612, 2898, 2568, 927, 3216, 3338, 528, 752, 1465, 3076, 169, 2534, 592, 3237, 2243, 215, 2274, 2369, 3353, 2853, 1717, 1045, 2649, 632, 1399, 1347, 3003, 3202, 1609, 1825]. measurements [192, 2389, 2945, 3168, 251, 2842, 2471, 2426, 61, 1668, 652, 1883, 811, 2951, 3217, 307, 2366, 1062, 1788, 84, 802, 1693, 917, 2446, 62, 1379, 1537, 560, 938, 2080, 2210, 3239, 335, 371, 1555, 3269, 1162, 445, 1662, 2830, 3149, 1647]. Measures [87, 3125, 3230, 3175, 3081, 1064, 909, 1424, 2282, 2623, 3051, 1331, 472, 1944, 113, 1853, 1152, 1536, 3040, 5, 1520, 181, 2242, 1992, 629, 1812, 796, 297, 2140, 2444, 2457, 2405, 180, 166, 2300, 3034, 2078, 2452, 1579, 2227, 2118, 681, 412, 443, 3325, 2359, 89, 3381, 712, 1910, 1120, 1586, 2721, 171, 2722, 3185, 2647, 391, 1397, 337, 2932, 2337, 2904, 2744, 2034]. Measuring [1191, 152, 1661, 2350, 215, 1989, 1241, 1943, 1082, 866, 2918, 554, 2413, 38, 2091, 621, 2198, 2662]. Mechanism [2973].

Mechanism-based [2973]. mechanisms [1215, 1929, 2786]. Mechanistic [913, 2815, 467]. Media [2718, 464]. Median [1252, 2386, 3298, 3358]. medians [1046]. Mediation [2716, 3176, 2466, 3140]. mediational [1478]. mediator [2466]. mediators [2743]. Medical [745, 1119, 2064, 1018, 1793, 1118, 141, 1206, 128, 186, 2364, 730, 314, 2262, 568, 740, 2771, 2775, 1126, 2736, 2356, 2224, 115, 2499, 65, 1787, 1279, 929, 1398, 2772, 2537, 2788, 2773, 2015, 1526, 2503, 1629, 1173, 830, 1052, 2239, 3167, 2201, 1378, 3192, 1728, 739, 537, 1373, 994, 590, 2961, 2654, 2719, 118, 1648, 2759, 2542, 2787, 2238, 2247, 2774, 2533, 1657, 3117, 2647, 2549, 509]. medication [2300, 2994, 604]. medicinal [1328, 2020, 2021]. Medicine [306, 3024, 871, 315, 2174, 2993, 2785, 305, 1877, 2074, 188, 208, 647, 1979, 538, 3041, 2751, 3105, 3135, 1117, 1168, 1545, 1649, 1591, 2002, 1951, 2000, 2440, 2327, 2245, 2344, 2379, 1424, 2910, 355, 1374, 803, 2052, 3202, 2820, 125, 13, 39, 77, 2907, 2735, 2439, 414, 813, 1561, 1696, 3386, 2638, 1784, 841, 2641, 2499, 3384, 10, 3070, 1682, 2450, 3190, 2050, 3379, 1645, 2033, 181, 1337, 1589, 2909, 2640, 2643, 387, 2437, 2134, 1715, 297, 1850, 1925, 3380, 386, 1714, 2653, 3108, 3310, 2642, 2676, 2489, 1054, 1056, 3385, 867, 11, 714, 2558, 3136, 1016, 2452, 757]. Medicine [2819, 958, 1362, 960, 1169, 1728, 3383, 2770, 2817, 416, 2855, 537, 1725, 2559, 2183, 1851, 1799, 2186, 2637, 2734, 1669, 815, 1711, 857, 2198, 1716, 1644, 2032, 3381, 2911, 3219, 2787, 78, 2036, 354, 2160, 3106, 29, 1053, 397, 3302, 1494, 3137, 3288, 2639, 759, 1450, 2476, 3382, 2839, 3221, 283, 1709, 2665, 1939, 1914, 358, 2772, 2773, 2295, 2774, 1180, 3057, 1671, 2231, 1270, 2163]. medicines [641]. Medine [1563]. Medizin [197]. MEDLINE [791]. Meehl [257]. Meeting [2570, 483, 1608]. MEG [2510, 2512]. MEGA [2746]. Mehrotra [3380]. Meier [1651, 530, 1580, 970, 1828]. meiosis [1112]. Melania [2476]. melanoma [2766, 2126]. mellitus [1820]. Melvin [1352]. memorandum [2064]. Memorial [305]. memory [1217, 2515]. men [1470]. Mendelian [2826, 2709, 3091, 1763, 2825, 2840, 2826, 2839]. Mengersen [1117, 1545, 1669]. Menière [2309]. meningitis [1420]. meningococcal [1344]. menopausal [2320]. menopause [120]. menstrual [2170]. Mental [167, 1518, 2435]. Menten [1306]. MEPS [2364]. merged [2686]. merits [2374]. message [1286]. messy [253]. Meta [1886, 2944, 601, 77, 516, 3270, 831, 832, 833, 3375, 546, 3060, 2491, 14, 3091, 2760, 3062, 2686, 1206, 1271, 78, 2679, 1763, 647, 547, 824, 1168, 330, 1064, 2111, 875, 803, 697, 3061, 2189, 2234, 403, 2550, 1094, 2907, 1630, 3284, 296, 3174, 3184, 2937, 789, 907, 2463, 174, 1412, 2357, 1968, 2612, 318, 2698, 2164, 665, 501, 787, 1403, 2680, 2112, 2097, 3120, 1172, 2392, 1557, 845, 2492, 846, 385, 1887, 2227, 1356, 727, 1139, 2316, 1799, 646, 2072, 1439, 2235, 2761, 2685, 3152, 591, 2236, 2269, 860, 2280, 313, 2539, 2908, 920, 2421, 2434, 1686]. meta [785, 2403, 2678, 1377, 1685, 788, 2938, 325, 2233, 3137, 790, 2683, 2684, 932, 1705, 713, 3312, 1309, 3231, 871]. Meta-analyses [647, 875, 3284, 3174, 174, 2698, 2164, 2680, 2227, 727, 2685, 860]. meta-analysi [2112]. Meta-Analysis [601, 1886, 2944, 77, 516, 3270, 831, 832, 833, 3375, 546, 3060, 2491, 14, 3091, 2760, 3062, 2686, 1206, 1271, 78, 2679, 1763, 547, 824, 1168, 1064, 875, 803, 3061, 2189, 403, 2550, 1094, 2907, 1630, 296, 3184, 2937, 789, 907, 2463, 174, 1412, 1968, 2612, 318, 665, 501, 787, 2097, 3120, 2392, 1557, 845, 2492, 846, 385, 1887, 1356, 1139, 1799, 646, 1439, 2235, 2761, 3152, 591, 2236, 2269, 2280, 313, 2539, 2908, 920, 2434, 1686, 2403, 2678, 1377, 1685, 2938, 325, 2233, 3137, 790, 2683, 2684, 932, 1705, 713, 1309, 871, 3312, 3231]. meta-analytic [2111, 2357]. meta-analytical [2234]. meta-epidemiological [2421, 785]. meta-re-analyses [330]. meta-regression [1403, 1172, 2316, 2269, 2434, 788, 713]. meta-regressions [697]. metabolic [1316]. metabolism [2366]. metabolism-based [2366]. metabonomic [2899]. Metal [2881]. methamphetamine [2485]. Methicillin [1369]. Methicillin-resistant [1369]. method [1117, 1545, 875, 3202, 249, 1862, 2550, 3022, 2423, 736, 3386, 2125, 296, 1456, 1935, 3097, 230, 953, 1116, 1794, 1398, 665, 3007, 2848, 576, 3120, 1062, 1467, 3264, 3310, 2687, 2460, 406, 675, 615, 3013, 2947, 2028, 2849, 1085, 1991, 3044, 407, 1362, 1379, 3301, 2455, 1360, 3308, 3199, 3001, 2551, 2879, 582, 514, 2911, 3353, 1910, 2480, 762, 2746, 1322, 1685, 1435, 2938, 2265, 989, 2031, 2843, 759, 1020, 824, 3213, 500, 3290, 1731, 2790, 1995, 3033, 721, 2447, 2353, 589]. Methodologic [2915]. Methodological [3090, 370, 239]. methodology [984, 1979, 640, 1283, 256, 2872, 576, 3023, 2846, 1238, 443, 2721, 2248]. Methods [2346, 3, 144, 158, 2570, 1746, 286, 1235, 2471, 129, 65, 399, 766, 396, 961, 541, 1144, 1452, 1142, 1118, 1010, 962, 781, 141, 1723, 1192, 1516, 496, 1399, 1416, 493, 3233, 1647, 1121, 3041, 2060, 2481, 2443, 2910, 803, 1771, 731, 223, 866, 1919, 2234, 403, 25, 1183, 1636, 1273, 2382, 1178, 2967, 1566, 3024, 2872, 664, 2267, 1223, 1929, 2017, 230, 292, 2909, 848, 1548, 2218, 1656, 1297, 1707, 2272, 1863, 1774, 196, 2400, 2998, 1626, 2602, 1850, 1628, 1629, 1524, 2349, 2140, 1269, 1551, 266, 2543, 573, 1804, 1075, 1286, 234, 1123, 626, 2697, 2100, 2201, 385]. methods [2920, 2188, 62, 28, 3347, 1563, 1872, 1801, 2634, 1913, 2960, 1901, 2704, 773, 1927, 2393, 2273, 9, 1940, 1941, 2182, 1851, 1612, 1799, 3307, 2373, 2104, 3073, 2430, 1459, 1471, 2173, 2291, 2084, 1959, 1208, 1669, 2367, 1083, 837, 2713, 2655, 1153, 2956, 1129, 3100, 372, 1717, 1006, 785, 110, 532, 3064, 1754, 109, 2523, 171, 1301, 1758, 2722, 1189, 1210, 373, 2250, 964, 1494, 669, 1315, 1666, 703, 1538, 1948, 790, 1996, 1022, 2683, 2741, 1588, 2115, 3149, 826, 713, 414, 2347, 2912, 1425, 452, 872, 147, 401, 198, 1550, 3311, 1206]. methylation [3359]. metrics [1264, 1370]. metropolitan [551]. Metz [1707]. Mexico [2041]. MIC [1868]. Michael [1158, 2501, 3190, 452, 821]. Michaelis [437, 1306]. Michele [185]. Micro [1145]. Micro-Array [1145]. microarray [2131, 1975, 2602, 1766, 945, 940, 2307, 3119, 3294, 2530, 2764, 2765, 1748, 2133]. Microarrays [1104, 2306, 1765, 2707]. microbiologic [801, 1053]. Microdata [600]. micromap [244, 461]. microscopic [1298]. microsimulation [12]. Mid [1040, 344, 2738, 1235, 1266]. Mid-course [1040, 2738, 1235]. Mid-trial [344]. midstream [888]. Midthune [386]. Migon [454]. migraine [2527]. migrant [2252]. Millard [1182, 869]. millennium [887, 81, 80, 83, 82, 469]. Miller [1158]. Milliken [2604]. Milton [399]. MIMP [3198]. Min [2032, 1450, 1167]. Miners [3020]. Ming [1517]. Ming-Hui [1517]. Mini [167]. minimal [1772]. Minimax [861]. Minimization [1879, 1528, 1404, 3339]. Minimization-reducing [1879]. minimize [1862]. Minimizing [1198]. Minimum [85, 1121, 1816]. Mining [1310, 1071, 1416]. Minmax [2022]. Minnesota [981, 1429]. minority [1081]. minute [293]. Miranda [818]. Mirkin [421]. mis [1969, 2557, 2890]. mis-specification [1969, 2890]. mis-specified [2557]. misaligned [227]. misclassification [948, 2567, 3342, 160, 1909, 528, 3343, 1761, 1609, 2188, 1857, 1211, 1114, 2045, 2799, 394, 1736, 2569, 2827, 1151]. misclassified [1003, 3159, 2207, 1614, 1937, 1603, 2762]. Mises [2017, 218]. misleading [3171]. mismatches [1602]. mismeasured [1616, 3018]. missed [165]. Missing [1929, 1007, 3363, 1363, 750, 113, 3184, 2382, 1009, 2288, 71, 3210, 2175, 162, 1288, 3075, 554, 1668, 968, 353, 3037, 2461, 1451, 2405, 84, 2630, 2207, 711, 2106, 555, 2118, 856, 2104, 912, 3198, 2562, 2574, 2701, 2080, 2925, 2277, 2863, 2469, 653, 2433, 669, 685, 1662, 2683, 2684, 3247, 500, 2462, 2831, 2101, 121]. missingness [1768, 2304, 690, 3198, 2194]. Misspecification [2805, 3277, 3338, 921, 3126, 2615]. Misspecified [2298, 2850, 1657, 2344]. Mistake [3384]. Misunderstandings [1871]. misuse [2547]. misuses [63]. Mitchell [455, 440]. Mithat [3139]. Mixed [490, 2880, 358, 1642, 2828, 2757, 168, 1376, 3031, 1197, 3324, 3243, 3352, 2703, 2604, 3265, 1076, 3252, 3238, 1655, 1935, 3063, 627, 829, 1597, 2378, 3077, 3078, 947, 2079, 778, 101, 2522, 629, 1849, 1074, 1467, 1834, 2223, 1596, 1598, 2532, 1049, 1634, 2237, 2415, 1540, 2903, 136, 1998, 2730, 2850, 2100, 1506, 2491, 3055, 701, 958, 1212, 606, 2487, 1927, 2316, 994, 2809, 1702, 2318, 226, 383, 1847, 993, 874, 2593, 2027, 2572, 2841, 2556, 2574, 1105, 2258, 2661, 2025, 3262, 2808, 2429, 3278, 500, 725, 952, 2341, 2892, 1241, 1854, 2497, 2194, 1400, 2404, 3299]. mixed-effect [2574]. Mixed-effects [2757, 2703, 1655, 1074, 1634, 2237, 2415, 1702, 2318, 1847, 874, 2572, 2258, 2808, 500, 725, 952]. mixed-model [1927]. mixing [519, 2601]. Mixture [2119, 2008, 2725, 232, 1747, 2094, 3132, 1161, 1770, 1430, 423, 2088, 3354, 1779, 1093, 3189, 2777, 2893, 2940, 1054, 696, 2013, 26, 749, 523, 3341, 1047, 1479, 1027, 499, 3001, 764, 637, 951, 1824, 486, 1407, 2447, 3346]. Mixtures [601, 3259, 2688, 812, 2770, 431, 1251, 2649]. ML [1113]. MLE [3281]. MLEs [2278]. MMR [1420]. modality [2647]. Model [1601, 2731, 1969, 2566, 2085, 921, 3261, 706, 2025, 3323, 3236, 3272, 58, 967, 3048, 1772, 1591, 2344, 15, 551, 827, 1584, 2105, 2119, 3243, 2407, 2948, 205, 381, 689, 3085, 2871, 2876, 249, 379, 2169, 950, 1331, 2864, 39, 926, 3022, 3277, 750, 1035, 2621, 1076, 1672, 3252, 25, 1319, 2592, 3344, 2180, 87, 3131, 1807, 2815, 3270, 1768, 3125, 1287, 3157, 2952, 2285, 204, 119, 1009, 3128, 2375, 3040, 622, 1430, 2010, 1935, 3063, 328, 666, 627, 1679, 2946, 2276, 2048, 2037, 3058, 2038, 423, 3285, 2608, 1112, 3287, 2114, 2378, 3077, 3078, 1307]. model [2554, 2079, 2691, 1213, 2176, 766, 2317, 863, 101, 1177, 1888, 1001, 2979, 3228, 1749, 638, 163, 2600, 2470, 1093, 1077, 2056, 1074, 3293, 796, 1821, 678, 2991, 194, 2338, 2445, 3376, 3338, 3189, 704, 1626, 3291, 1974, 3254, 1280, 996, 3142, 1524, 1585, 3244, 1834, 2223, 2893, 1796, 2627, 2299, 282, 3116, 2614, 2461, 752, 2609, 2803, 3035, 659, 923, 1574, 1895, 1509, 2298, 2940, 2890, 1261, 1884, 2417, 2495, 845, 925, 1465, 2013, 2405, 136, 2239, 2906, 3074, 2883, 3126, 2493, 2496, 277, 299, 2564, 910, 1739, 657, 2195, 2106, 1454, 3271, 958, 1212, 1887, 2832, 1927, 1587, 1047]. model [3080, 608, 1744, 254, 1920, 596, 2603, 2615, 3308, 560, 1434, 3001, 2222, 637, 2271, 1439, 3249, 2388, 383, 2354, 3068, 2868, 1140, 1847, 1947, 2387, 312, 1667, 1230, 2045, 1603, 2027, 208, 1343, 2236, 1843, 2572, 322, 839, 2618, 3138, 507, 2035, 2039, 491, 1105, 1542, 2080, 2258, 700, 2259, 2555, 2956, 1488, 2844, 691, 971, 235, 2036, 2383, 1896, 650, 2980, 218, 3106, 1573, 2594, 2531, 2533, 1440, 1345, 3306, 1734, 2301, 2108, 325, 1358, 317, 722, 2573, 618, 3117, 951, 723, 1538, 1500, 1934, 3029, 3247, 1267, 1390, 1931, 2957, 2062, 2845, 2341, 2418, 1963, 1860, 486, 1921]. model [2447, 2670, 3110, 3346, 178, 1167, 2762, 1709, 3003, 3017, 526, 434]. Model-based [3261, 706, 750, 2056]. model-building [3138]. Model-checking [1601]. model-free [1887]. Modeling [1157, 3079, 3128, 3195, 182, 3361, 2745, 2764, 2511, 2900, 2922, 2671, 3351, 3020, 2703, 2508, 3168, 2795, 3377, 301, 2828, 2797, 2899, 3256, 2857, 3257, 3167, 3341, 3147, 2748, 3127, 3345, 2978, 3088, 3240, 3102, 2729, 3049, 2889, 526, 129, 1018]. modelization [2725]. Modelldiagnose [343]. modelled [1501, 1210, 725]. Modelling [1209, 1866, 2579, 965, 651, 1346, 1625, 505, 1242, 2040, 2041, 1087, 2563, 1856, 829, 611, 1556, 190, 2043, 3354, 248, 1300, 1894, 1332, 3335, 2812, 1973, 2460, 2213, 231, 180, 166, 1805, 1595, 3268, 1357, 167, 1479, 117, 478, 1459, 1499, 2690, 2581, 2016, 820, 2351, 3087, 2170, 1440, 1798, 1147, 2527, 2407, 3032, 1051, 1771, 3083, 232, 1508, 885, 2181, 2525, 561, 2599, 1833, 12, 2610, 2842, 726, 3086, 1505, 1642, 515, 2612, 2376, 929, 1042, 913, 2134, 1321, 2366, 1141, 1452, 67, 1333, 246, 1090, 1519, 234, 718, 1049, 1491, 2484, 22, 583, 911]. modelling [2256, 2955, 1755, 695, 1095, 1617, 26, 879, 1373, 1744, 2318, 3069, 2107, 2416, 467, 1470, 384, 334, 188, 834, 1464, 3138, 765, 2635, 2259, 2421, 1547, 2595, 1254, 3242, 110, 1583, 242, 2007, 1458, 1555, 1549, 2193, 3082, 1931, 2364, 2115, 450, 1364, 142, 185]. Models [1157, 542, 361, 143, 843, 362, 358, 42, 132, 1376, 2240, 1416, 1638, 1197, 810, 2716, 1769, 998, 66, 2527, 2694, 1179, 1302, 2281, 2432, 3324, 2993, 553, 3243, 1290, 522, 2894, 1720, 2052, 1572, 3352, 289, 1928, 1762, 3212, 3289, 1633, 2836, 2169, 702, 1003, 1674, 3265, 2805, 2766, 2436, 2735, 490, 2880, 937, 134, 613, 2082, 3300, 3159, 2624, 777, 2120, 2268, 1737, 1382, 3132, 2886, 2425, 1655, 1161, 1770, 1274, 2779, 120, 1738, 228, 2873, 2261, 393, 2935, 915, 2942, 829, 928, 1597, 3208, 2083, 878, 350, 3354, 1160, 2494, 2979, 224]. models [2522, 629, 1849, 2566, 1320, 1060, 2568, 2905, 247, 2216, 64, 1106, 86, 1802, 1971, 1957, 934, 1478, 2973, 3338, 217, 1793, 679, 549, 2366, 2733, 3177, 1467, 2882, 508, 1596, 2392, 1476, 446, 447, 1510, 1889, 1054, 240, 696, 684, 847, 1202, 2660, 2943, 2794, 3188, 241, 1540, 2903, 1998, 2730, 2947, 84, 2850, 2304, 300, 2699, 2849, 2813, 2870, 548, 2343, 238, 3018, 2920, 3055, 203, 1088, 2949, 1378, 701, 2950, 749, 523, 168, 88, 1201, 2989, 2487, 2534, 466, 2507, 1665, 1248, 2251, 1999, 2616, 498, 1994, 1721, 994, 1932, 924, 1620, 2377, 2024, 942, 2809, 2887]. models [1702, 2318, 681, 1050, 686, 973, 1027, 799, 3008, 764, 226, 2976, 321, 478, 677, 1348, 2458, 2005, 835, 2398, 2526, 993, 384, 2387, 412, 874, 2593, 2084, 3114, 2205, 2734, 118, 834, 1015, 2686, 2270, 1455, 672, 2244, 2841, 2556, 623, 2574, 2759, 2254, 712, 1124, 68, 2053, 1092, 2661, 2025, 2277, 2396, 3262, 3365, 3355, 1891, 235, 31, 1468, 3064, 1586, 1759, 2808, 1890, 2934, 195, 390, 2782, 2303, 1527, 1045, 734, 2342, 2429, 1410, 2275, 1664, 647, 3278, 1912, 680, 2647, 1046, 2684, 2901, 500, 725, 952, 2462, 1824, 2892, 1241, 633, 1854, 1426, 2497, 2692]. models [957, 1449, 3050, 2706, 1407, 2194, 2365, 2715, 133, 3218, 227, 535, 1972, 2404, 2009, 1870, 324, 1598, 1437, 2604, 200, 1753, 2140, 2519, 187]. modelwith [1318]. moderate [1951, 2033, 1637, 1775]. moderate-sized [1951, 1637]. moderating [616]. moderators [2743]. Modern [872, 103, 171, 91]. Modification [655, 425, 2878, 1040, 1374]. modifications [1858]. Modified [1811, 1983, 1103, 2164, 1046]. modifiers [697]. modifying [1296]. Modulation [2873]. Moeschberger [1352]. Mogg [3380]. Mohamed [2959, 2034]. Mohammad [2959]. molecular [557, 1685]. molecularly [1757, 2308]. Molenberghs [2409, 1547]. Møller [1504]. moment [3002]. money [2091]. monitor [2204, 659]. monitored [2969]. Monitoring [2021, 1387, 1385, 585, 1402, 894, 2266, 1442, 2238, 2247, 892, 2245, 2694, 275, 2172, 323, 2019, 1388, 2065, 2501, 571, 1386, 2151, 878, 1475, 485, 1398, 2587, 1186, 2124, 1333, 1389, 718, 345, 2339, 683, 893, 1238, 686, 3197, 2145, 658, 1312, 2333, 1870]. monotone [3121, 2701, 2080, 2258, 2194]. monotonic [1033, 1014, 626, 208, 2556]. monotonicity [3334]. Monte [93, 2452, 2858, 2281, 2282, 228, 3260, 2906, 2294, 1091]. months [1191, 2925]. Mor [2050]. Morabia [1872]. morbidity [1280, 2699]. Morgenthaler [3231, 94]. Morphometric [2598]. Morris [3386, 3310]. Mortality [258, 1638, 66, 2432, 1482, 561, 1287, 175, 2731, 1738, 2261, 673, 515, 111, 638, 2117, 194, 2588, 2239, 32, 442, 226, 990, 3004, 3124, 2799, 1890, 1267]. Morton [1117, 1545, 1669]. mosaicism [1632]. most [1079, 2069, 2226]. mother [640, 639, 2969]. mother-to-child [640, 639, 2969]. Motulsky [1802]. mouth [2471, 3349]. movement [1141]. moves [1270]. Moyé [2665, 3180, 1546]. MP [386]. MRI [2505]. Much [2234]. Mugford [818]. Müller [1374]. Multi [66, 2924, 508, 2596, 3038, 2485, 3017, 1197, 114, 2086, 1879, 87, 2099, 328, 2942, 2566, 1107, 353, 1559, 976, 3034, 2883, 910, 254, 2613, 346, 2398, 993, 1135, 2270, 839, 672, 1136, 2259, 662, 1896, 1740, 1189, 2647, 486, 2929]. multi-arm [2086, 1107, 976, 1135]. multi-case [1189]. multi-centre [114, 1879, 2099, 353, 910, 2613, 346, 993, 486]. multi-component [2566]. multi-drug [1136]. Multi-Ethnic [3038]. multi-item [662]. multi-level [3034, 2883, 254]. multi-modality [2647]. multi-parameter [1197]. multi-plane [839]. multi-reader [1189, 2647]. Multi-rule [2924]. Multi-sample [2485]. multi-stage [672, 1740, 2929]. Multi-state [66, 508, 3017, 87, 328, 2942, 1559, 2398, 2270, 2259, 1896]. Multi-syndrome [2596]. multicenter [3058]. multicentre [2579, 1300, 2223, 2410]. Multicollinearity [971]. multicomponent [3286]. multidimensional [1627, 2571, 821]. multifactorial [1793]. Multilevel [929, 390, 2193, 1931, 2525, 3058, 698, 1141, 2832, 3059, 325, 1753, 2857, 820]. Multilocus [2191]. multinational [1108, 1741]. Multinomial [522, 2935, 349, 2292, 3252, 1074, 1899, 2396]. multinomials [1860]. Multiple [1918, 53, 49, 3357, 1497, 1580, 782, 1770, 1070, 1522, 2187, 2397, 2442, 2438, 2283, 2954, 1502, 2430, 2128, 2632, 2853, 1865, 3117, 2115, 480, 3047, 3162, 3233, 3155, 1769, 1064, 355, 3243, 2513, 2313, 3143, 3153, 2334, 1273, 2312, 2931, 1258, 3128, 585, 2779, 2402, 2589, 2026, 598, 2608, 2828, 3075, 554, 2335, 922, 2272, 2520, 1493, 2991, 2168, 3129, 2861, 2848, 212, 1827, 1269, 800, 2532, 581, 2061, 3257, 851, 1998, 588, 2304, 2633, 1780, 512, 2491, 714, 1150, 1817, 1681, 27, 1016, 2707, 26, 3076, 3211, 1381, 2456, 1537, 3363, 3387, 1721, 621, 20, 972, 16, 3068, 3198, 614, 1756]. multiple [2783, 2362, 1911, 3036, 2682, 3054, 1908, 1208, 426, 1083, 2359, 514, 1512, 1254, 838, 354, 34, 2721, 107, 1025, 1253, 3031, 50, 500, 1447, 2437, 198, 782, 511, 1546]. multiple-class [1537]. multiple-race [1083]. multiple-record [1447]. multiple-related [2721]. multiplicative [381, 704, 1889, 1048]. Multiplicity [1756, 2356, 3261, 2352, 1208]. multiply [2803, 3106, 2862]. multirater [2042]. Multiregion [242]. Multiscale [1958, 1991]. multisera [2812]. multistage [1280, 1341]. multistate [1768, 2608, 863, 282, 3214]. Multivariable [2421, 3085, 1010, 1015, 2618, 440, 3138]. Multivariate [1290, 2335, 566, 865, 1441, 104, 2361, 241, 3271, 1139, 3080, 3069, 3242, 1995, 1594, 2432, 3351, 2324, 965, 2785, 332, 490, 1044, 1636, 2592, 3344, 2314, 3174, 2610, 1188, 2276, 2048, 1505, 878, 2797, 2029, 3228, 2557, 2680, 3234, 956, 2543, 1229, 923, 3188, 3074, 3196, 2106, 1852, 3281, 994, 1932, 2388, 2881, 1442, 2027, 1421, 1488, 2844, 1401, 2696, 1763, 1345, 3278, 3050, 2706, 713, 489, 42, 257]. München [494]. muscle [326]. musculoskeletal [2955]. mutation [939, 2620]. mutational [808]. mutations [808, 2357, 2974]. Muth [2824]. mutual [3358]. mutually [2938]. Muzaffarpur [854]. myeloid [1734]. myeloma [598, 1016, 2707]. Myles [1752, 147]. myocardial [2907, 1737, 910, 1470, 2539, 2908, 2722, 3137]. myopia [1842]. Myunghee [1801].

N [1168, 1649, 1786, 1425, 42, 602, 3310, 1142, 780, 300, 1118, 757, 494, 416, 3056, 857, 298, 2720, 2839, 2452]. n/a [2452]. Nagaraj [869]. Nagelkerke [416]. Nakamura [757]. Nam [3382]. Nan [2502]. Nathan [3386]. National [1631, 420, 1153]. natural [177, 2271, 3141]. nature [2946]. Naus [1384]. NC [3139, 2754, 2604, 3204]. near [2178, 1508]. nearly [2218, 475]. necessity [2820, 2617]. neck [121]. need [2638, 3354, 2144, 2643, 2789, 3246]. needed [1733, 1838, 429, 408, 670]. needle [2736]. needs [1393, 1535, 2362]. Neerchal [869]. negative [731, 1049, 2468, 1947]. negatively [74]. neglected [3327, 3328, 3329, 3332, 3326, 3331, 3330]. Neglia [437]. neighborhood [3038, 2763, 3240]. neighborhood-level [3038]. Neil [817, 362, 870]. neonatal [561]. nephropathy [1268, 2396]. Ness [211]. Nested [2306, 2205, 3054, 1842, 2854, 681, 1426]. Net [3134, 3133, 451, 2644, 484, 997]. Netemeyer [1461]. Netherlands [1256]. Netta [2787]. Network [846, 2508, 2904]. networks [651, 1869, 767, 598, 1016, 1335, 63, 221]. Neuenschwander [3379]. Neuhäuser [1726]. Neuilly [185]. Neumann [545]. neural [1498, 651, 1869, 767, 598, 1016, 1335, 2084, 63, 2516, 221]. neural-Bayesian [1498]. neuroblastoma [3189, 1187]. neuromotor [1556]. Neuronal [2504, 2511]. neurons [1764]. neurophysiology [209]. neuroreceptor [3169]. Neurostatistics [2659]. newborn [37]. Newcombe [1850]. newly [2780]. Newman [1566]. next [2230, 2162]. Neyman [2692]. NHS [1424, 207, 909]. NIA [155]. Nicholas [1840]. Niels [257]. Nigel [3311]. Nigeria [1638]. Nils [1682]. Nilsson [1711]. NINDS [512]. NJ [2856, 2502]. NNT [408, 670, 2623]. No [1530, 1594, 91, 341, 819, 817, 1157, 1460, 1647, 93, 1565, 454, 1592, 421, 542, 3139, 198, 509, 3222, 2138, 1752, 1786, 1351, 2754, 601, 144, 455, 360, 2824, 526, 2604, 885, 1425, 434, 1158, 1364, 497, 1608, 286, 3203, 782, 361, 1751, 143, 200, 127, 183, 1750, 1873, 301, 145, 402, 843, 362, 288, 2501, 3204, 1415, 2006, 129, 603, 1566, 1550, 90, 1134, 543, 343, 1402, 358, 600, 2004, 1119, 399, 2856, 42, 1461, 420, 40, 201, 1787, 961, 3311, 303, 1245, 511, 1180, 1067, 1548, 1753, 131, 3312, 1194, 1182, 541, 3057, 211]. No [1352, 452, 182, 566, 1018, 1195, 1205, 342, 844, 1774, 1802, 3182, 439, 196, 3181, 2399, 1310, 417, 1144, 1181, 436, 92, 438, 2409, 1533, 2503, 1311, 2140, 453, 1159, 1593, 1309, 104, 818, 435, 1133, 884, 495, 1142, 2857, 780, 130, 300, 302, 1118, 3180, 1531, 524, 299, 1595, 2959, 1156, 1671, 821, 1609, 132, 3347, 962, 142, 1801, 3192, 359, 2960, 781, 357, 494, 1840, 1437, 525, 3056, 1193, 1363, 440, 285, 141, 1143, 1532, 1145, 2858, 544, 869, 2519, 872, 419, 1192, 2961, 43, 257, 1546, 2005, 146, 106, 2109, 1471, 2139, 2502, 2859, 2719]. No [287, 2295, 870, 3122, 260, 1648, 545, 1516, 185, 2841, 3138, 3231, 437, 1206, 2345, 128, 298, 400, 418, 820, 103, 540, 1517, 1065, 1610, 510, 187, 1504, 2718, 284, 1384, 1825, 1698, 147, 105, 2720, 259, 184, 496, 197, 842, 1066, 186, 433, 210, 1684, 1839, 401, 1244, 1503, 1416, 871, 41, 94, 2755, 199, 1676, 914, 221, 1818]. no-effect [914]. No.of [1549]. node [2468, 1336]. nodes [756]. NOEL [914]. nominal [2178, 2343, 2690]. nomogram [1240]. Non [984, 998, 245, 1674, 2449, 755, 2552, 985, 1178, 1033, 1986, 2175, 2289, 2371, 2609, 986, 1981, 2106, 1605, 3301, 3225, 2202, 138, 686, 1335, 1421, 347, 109, 572, 1263, 703, 2009, 2257, 1943, 1275, 3272, 538, 3041, 1117, 1545, 2020, 2245, 2297, 311, 2086, 3032, 2313, 1207, 1578, 1037, 877, 2019, 1944, 1482, 813, 1319, 1696, 1980, 1768, 3238, 444, 1750, 1161, 1770, 1982, 1703, 992, 1428, 61, 2, 829, 423, 2652, 1112, 1086, 1041, 1215, 2079, 918, 214, 162, 1288, 2658, 1177, 252, 1359, 1116, 2483, 2576, 1883, 2667, 1060, 519, 1802, 3182, 968, 2011]. non [934, 2168, 988, 1419, 3045, 921, 1141, 213, 1467, 1764, 1551, 2249, 3037, 1900, 2653, 2306, 1776, 2743, 1476, 1264, 1983, 923, 1247, 1451, 506, 1540, 2405, 2749, 3385, 1616, 1084, 687, 1987, 2849, 1199, 1313, 2446, 2558, 2630, 3136, 1552, 1095, 2195, 719, 1028, 2200, 442, 3283, 2997, 2264, 1356, 3362, 189, 431, 416, 761, 1665, 1851, 1612, 3147, 1702, 2318, 2104, 1462, 973, 499, 2107, 987, 1034, 1808, 1140, 993, 874, 3114, 1528, 1230, 990, 208, 3036, 2236, 2572, 1984, 2896, 582, 1669, 815, 2159, 2574, 2198, 2668, 2701, 2080, 2258, 1985, 1485, 2259, 991, 2655, 2911, 2053, 2656, 2657]. non [2277, 2351, 2435, 1058, 1322, 2531, 1841, 2071, 2669, 2275, 513, 989, 1026, 1450, 500, 725, 952, 2364, 2852, 1427, 1988, 1231, 2194, 1810, 721, 2727, 1289, 2252, 1404, 489, 1979, 3382]. non-adherence [2053]. non-compliance [2297, 3032, 1207, 2449, 1883, 1060, 921, 1551, 1451, 2446, 2558, 2630, 3136, 719, 3283, 189, 416, 582, 1985, 2911]. non-diseased [1141]. non-disjunction [1112]. non-effects [1275, 1450]. non-fatal [442]. non-Gaussian [2852]. non-homogeneous [934, 2236, 2435]. non-ignorable [3272, 1768, 1161, 1770, 423, 2079, 162, 1288, 1883, 968, 3037, 1476, 1451, 1540, 2405, 1028, 761, 3147, 2104, 2080, 2277, 2852, 1289]. Non-inferiority [984, 2552, 985, 986, 1981, 3225, 538, 3041, 2020, 2245, 2313, 1578, 1037, 2019, 1482, 1980, 1982, 1703, 992, 2652, 1986, 1041, 2658, 2667, 988, 213, 2249, 2653, 1776, 1983, 2749, 3385, 687, 1987, 1199, 1313, 2997, 3362, 987, 1034, 993, 1528, 990, 1984, 2896, 2159, 2668, 1485, 991, 2655, 2656, 2657, 2669, 513, 989, 1810, 1404, 1979, 3382]. non-informative [2574, 703]. Non-iterative [3301, 1117, 1545, 1116, 1669, 1322]. Non-linear [998, 2106, 686, 1335, 2009, 2257, 1750, 829, 1802, 1467, 1702, 2318, 973, 874, 208, 2572, 500, 725, 952]. non-migrant [2252]. non-monotone [2701, 2080, 2258, 2194]. Non-normal [2175, 2483, 923, 2849, 2264]. non-normality [1264]. non-null [2727]. Non-parametric [245, 1674, 2449, 755, 1178, 1033, 1986, 2289, 2371, 2609, 1605, 2202, 138, 1421, 347, 109, 572, 1263, 703, 1943, 311, 1944, 813, 1696, 444, 1703, 61, 918, 252, 1359, 2168, 1419, 3045, 1764, 2306, 2743, 1247, 506, 1356, 1851, 1612, 1462, 1140, 3114, 1230, 3036, 2198, 2071, 1988, 721, 489]. non-proportional [2086, 1177, 2107, 1808]. non-random [877, 519]. non-randomized [1428, 1616, 2200, 2531, 1026]. non-randomly [1900]. non-rare [1095]. non-responders [1552]. non-response [214, 2011, 1084, 761, 1665, 499, 2258, 2351, 2364]. non-rheumatic [3238]. non-sequential [2576]. non-significant [1231]. non-sinusoidal [2195]. non-small-cell [1427]. non-specific [431]. non-standard [1319]. non-statisticians [3182]. non-susceptible [1215]. non-telephone [1086]. non-terminal [2259]. non-transient [1841]. non-uniform [2086, 2275]. non-unity [2, 815, 1058]. non-zero [2, 815]. noncompliance [3229, 3334, 3027]. nonignorable [2134, 1755, 2326, 3290, 2831]. Noninferiority [3014, 2654, 2884]. nonlinear [3243, 2703, 3287, 2366, 1212, 2388, 2593, 3240]. Nonparametric [1425, 3007, 3129, 2742, 3362, 2868, 3149, 3121, 3311, 2951, 2867, 2947, 2601, 2859, 1503, 147]. nonrandomized [2601]. nonrespondent [2783]. nonsusceptible [2816]. Nordic [1176, 334]. Norleans [1192]. Normal [1423, 1778, 2814, 2324, 2891, 2767, 1307, 2175, 2317, 2483, 1585, 2428, 923, 2730, 2849, 1085, 1362, 1354, 2264, 3281, 637, 1439, 2556, 2121, 2675, 1488, 776, 2934, 1555, 1102]. normality [1264]. normally [501]. Normand [77, 2452, 78]. Norovirus [1744]. Northern [2915]. Nose [757]. nosocomial [328]. note [2245, 1762, 2019, 772, 1696, 2268, 2875, 3379, 1327, 519, 1247, 168, 3356, 2896, 2911]. Notes [129, 503]. nothing [2234, 3152, 1377]. notion [55]. Notzer [2787]. Nova [76]. Novel [1135, 2955]. NRC [947]. NRC-II [947]. null [289, 2, 609, 815, 2727]. Number [603, 602, 670, 1733, 2414, 519, 294, 836, 2838, 1369, 1838, 2867, 504, 2014, 1823, 3193, 408, 2843, 1329, 365]. numbers [2241, 2337]. Numerical [266, 1112]. nursing [1923, 2050, 2175, 1867]. nutritional [3202, 2879].

O [1608, 1461, 302]. O. [1494]. Oaks [455]. OBEL [914]. obese [1106]. obesity [364]. objectives [2402]. obliterans [679, 1125]. observable [1704]. observation [1514, 1974, 2705, 1149, 1692]. observational [3184, 565, 1257, 2772, 3335, 534, 2395, 3101, 2326, 694, 2217, 2232, 3201, 1115]. observations [1275, 852, 2225, 1432, 1656, 2427, 1526, 1559, 752, 1229, 136, 1199, 280, 3008, 3350, 2355, 3173, 1450, 268]. observed [332, 2695, 3273, 1496, 1992, 1040, 760, 1901, 376, 1664, 1599, 1162, 1046, 1002, 661]. observer [2272]. obsolete [2672]. Obtaining [84, 1453]. Obuchowski [1707]. occasion [267]. occasion-to-occasion [267]. occupational [2784, 3100, 594]. occurrence [1613, 1831, 2969, 1595]. occurrences [2339]. Odds [1418, 3041, 2443, 3025, 2910, 2878, 1674, 2472, 2592, 296, 160, 3024, 2909, 927, 1292, 1544, 3126, 2591, 512, 2496, 917, 2989, 2220, 834, 1543, 2796, 2655, 2662, 2467, 2999, 2219, 1810, 1972, 193, 283, 825]. O’Donnell [1194]. Ofer [2437]. off [2119, 1430]. Offen [2409]. offer [740, 739]. Office [818, 2955]. offs [1950, 619]. Offspring [1299]. often [466]. old [1470]. Oldenbourg [494]. Oliver [2604]. omitted [723, 1231]. on-going [3097, 1731]. onco [1893]. onco-haematology [1893]. oncological [849]. oncology [2567, 3206, 1482, 3011, 3009, 2672, 2987, 990, 63, 3010, 1438, 342]. One [2411, 1789, 3158, 3061, 3265, 1497, 367, 609, 1715, 1714, 2464, 2591, 1724, 1716, 1713, 685, 2506]. One- [2411, 1789, 3158]. one-level [2506]. one-sample [685]. one-sided [1497, 2464, 2591]. one-way [3061]. ongoing [2521]. Online [2452, 1270, 1186]. only [1456, 1579]. onset [2059, 3241, 76, 2017, 163, 2011, 679, 1560, 1285, 2126, 2078, 1820, 68, 218, 828, 2132]. Open [399, 658]. Opening [2570, 1069, 2347, 459, 2912]. Operating [2156, 3139, 24, 1682, 3007, 3123, 1707, 1446, 807, 762, 763, 2408]. operational [352, 2114]. operations [571]. operative [192, 515]. ophthalmologic [2103]. ophthalmology [505, 2459]. opinions [1885]. opportunistic [773]. opportunities [1284]. opportunity [740, 739]. optical [2507]. Optimal [3009, 485, 2372, 1306, 949, 3193, 557, 424, 562, 1881, 69, 2226, 1355, 3386, 2425, 2629, 1405, 3310, 2394, 3387, 2809, 874, 2972, 2217, 2367, 2650, 2808, 2573, 1730, 3213, 339]. Optimization [2593].

Optimized [2929]. Optimizing [2895, 3029]. Optimum [17, 2203, 2957, 397, 3056]. Oral [481, 645, 3352, 102, 1115]. Ord [299]. Order [550, 3243, 3157, 1258, 1935, 393, 1185, 800, 2460, 1920]. Order-restricted [550]. Ordered [1537, 256, 2402, 2589, 2055, 395, 1634, 1032, 2456, 138, 3295, 3325, 1421, 2469, 3158]. ordering [933]. orderings [3178]. Ordinal [3359, 1970, 710, 2172, 505, 2880, 2923, 65, 2554, 1992, 2340, 2893, 1819, 2943, 476, 2237, 3074, 1313, 548, 1689, 2343, 910, 657, 1462, 1439, 383, 1348, 1847, 2745, 3325, 322, 1098, 1063, 2303, 2885, 2275, 1555, 547, 1397, 1989]. ordinates [1682, 1446]. Organization [884, 2508, 1919]. organizations [2351]. organs [2953, 62]. Oriented [1143]. origin [955, 946]. original [2476]. origins [629]. orthodontic [3350]. oscillations [2508]. oscillatory [2514]. ossificans [865]. osteosarcoma [389]. other [37, 175, 2382, 1258, 2112, 2954, 1264, 2188, 3355, 2986, 2301, 1486, 1130]. our [3095]. outbreak [950, 2041, 3130, 2596, 1256, 1744, 3088]. outbreaks [1977, 2363, 2241, 2216, 2597, 1956, 1344]. outcome [1064, 15, 39, 3168, 1044, 1600, 1477, 2895, 3208, 2267, 665, 546, 968, 2479, 3111, 508, 528, 1761, 1689, 657, 442, 3094, 1418, 1395, 254, 2373, 972, 1857, 2388, 1499, 1897, 443, 2080, 1301, 3306, 1210, 2221, 373, 1705, 2889, 3218]. outcome-dependent [3168]. outcomes [2716, 139, 2481, 349, 379, 1003, 332, 3377, 3159, 1807, 756, 1235, 831, 832, 789, 2842, 2946, 2026, 2175, 1097, 1929, 2079, 1398, 2587, 1107, 2848, 2717, 1186, 3142, 1819, 2940, 2415, 2300, 2100, 512, 2955, 2558, 2630, 1689, 3018, 719, 2200, 1357, 606, 3098, 994, 1799, 3059, 3147, 973, 2613, 2689, 3068, 2881, 2173, 2235, 2760, 1135, 2685, 3325, 3054, 3038, 582, 860, 2701, 1115, 2542, 3353, 1686, 2277, 1098, 3232, 3230, 2531, 2934, 2303, 2571, 2938, 1865, 325, 2467, 2219, 1666, 790, 3031, 1057, 1046, 547, 932, 2917]. outlet [3240]. outliers [474]. outputs [1239]. outs [423, 1863, 1354, 912, 653, 445]. ovarian [560, 681, 1135]. ovary [2954]. over-dispersed [1688]. Over-ruling [1228]. Overall [1434, 3096, 2550, 3285, 501, 2387]. overdiagnosis [1187]. overdispersed [552, 3307, 1783]. overdispersion [2341]. overdose [1758]. overestimation [1969]. overexpression [1151]. Overlap [1604, 470, 3112, 347]. overlapping [1447]. overview [1121]. overweight [1106, 2418]. overweight/obese [1106]. ovulation [413]. Owen [1194]. oxcarbazepine [2251]. Oxford [1592, 144, 1608, 1415, 420, 1787, 1245, 1180, 1802, 2503, 1142, 302, 1118, 1531, 1595, 3192, 3056, 1143, 1532, 2519, 146, 1648, 437, 540, 187, 1825, 1549].

P [1117, 1545, 2000, 1424, 803, 1752, 1425, 434, 1608, 10, 129, 3024, 1682, 1597, 258, 181, 961, 1182, 1352, 1925, 1142, 302, 1118, 1531, 3347, 1362, 960, 1840, 1532, 544, 869, 1669, 815, 1711, 105, 496, 433, 3221, 709, 2452]. P.-S [181]. PA [512, 2922]. paediatric [1308, 621]. Pages [421, 3222, 1752, 1872, 2034, 1530, 1594, 91, 341, 819, 817, 1157, 1460, 1647, 93, 1565, 454, 1592, 542, 3139, 198, 509, 2138, 1786, 1351, 2754, 601, 144, 455, 360, 2824, 526, 2604, 885, 1425, 434, 1158, 1364, 497, 1608, 286, 3203, 782, 361, 1751, 143, 200, 127, 183, 1750, 1873, 301, 145, 402, 843, 362, 288, 2501, 3204, 1415, 2006, 129, 603, 1566, 1550, 90, 1134, 543, 343, 1402, 358, 600, 2004, 1119, 199, 399, 2856, 42, 1461, 420, 602, 40, 201, 1787, 961, 3311, 303, 1245, 511, 1180, 1067, 1548, 1753, 131, 3312, 1194]. pages [1182, 541, 3057, 211, 1352, 452, 182, 566, 1018, 1195, 1205, 342, 844, 1774, 1802, 3182, 439, 196, 3181, 2399, 1310, 417, 1144, 1181, 436, 92, 438, 2409, 1533, 2503, 1311, 2140, 453, 1159, 1593, 1309, 104, 818, 435, 1133, 884, 495, 1142, 2857, 780, 130, 300, 302, 1118, 3180, 1531, 524, 299, 1595, 2959, 1156, 221, 1671, 821, 1609, 132, 3347, 962, 142, 1801, 3192, 359, 2960, 781, 357, 494, 1840, 1437, 525, 3056, 1193, 1363, 440, 285, 141, 1143, 1532, 1145, 2858, 544, 869, 2519, 872, 419, 1192, 2961, 43, 257, 1546, 2005, 146, 106, 2109]. pages [1471, 2139, 2502, 2859, 2719, 287, 2295, 870, 3122, 260, 1648, 545, 1516, 185, 2841, 3138, 3231, 437, 1206, 2345, 128, 298, 400, 418, 820, 103, 540, 1517, 1065, 1610, 510, 187, 1504, 2718, 284, 1384, 1825, 1698, 147, 105, 2720, 259, 184, 496, 197, 1066, 186, 433, 210, 1684, 1839, 401, 1549, 1244, 1503, 1416, 871, 41, 94, 2755, 842]. Paik [1801]. pain [192, 559, 1925, 621, 383]. pair [702, 992, 1148, 2983, 720, 3301, 3225, 98, 1058, 3382]. pair-matched [98]. paired [1275, 2482, 1813, 1695, 2122, 2953, 1453, 811, 3321, 732, 2869, 2646, 2749, 687, 1987, 1199, 1313, 62, 1605, 683, 2855, 409, 2559, 2528, 1864, 628, 1450, 1810, 493]. paired-sample [1313]. pairs [1275, 1977, 3241, 3375, 3305, 2884, 991, 1450]. Pairwise [1703, 1780, 933]. País [1924]. Pak [145]. Panel [2364, 156]. Paper [221, 199, 1461]. Paperback [1787, 259, 1549, 1802, 2503, 1671, 3192, 3056, 1698]. Papers [455, 1939, 1914, 1979]. papilloma [1130]. Paradigm [577, 90]. parallel [1691, 831, 832, 833, 1946]. parallel-group [1946]. parallels [2232]. parameter [1197, 827, 1487, 2268, 627, 1811, 1288, 2203, 1676, 2112, 958, 2393, 835, 2941, 1013, 1114, 1859, 1912]. parameterization [423]. parameterizations [3270]. parameters [2919, 2431, 950, 2990, 2590, 2048, 2114, 2812, 3053, 3281, 1243, 2970, 1586, 1457]. Parametric [72, 561, 2538, 2779, 928, 2600, 1510, 238, 1373, 3249, 2885, 1315, 1943, 2379, 311, 2138, 245, 1674, 1944, 2449, 813, 1696, 755, 1833, 444, 2939, 2320, 1178, 1033, 1703, 61, 2196, 8, 1986, 918, 2289, 2612, 2315, 252, 1359, 3216, 2168, 331, 1419, 996, 2371, 3045, 1764, 1269, 1694, 2306, 2743, 2609, 1509, 2660, 1247, 506, 925, 3055, 1605, 1356, 1248, 2616, 1047, 2202, 1851, 1612, 2377, 2615, 138, 1462, 2458, 1140, 2941, 3114, 1230, 834, 3036, 1421, 2198, 1542, 762, 347, 624, 109, 572, 2071, 734, 2807, 1263, 703, 1503, 952, 2192, 1988, 3309, 721, 1921, 3006, 489]. parametric [1851]. parasite [1843]. parasites [1092]. Paris [185]. Parkin [603, 258, 131, 437]. Parmer [1531]. Parmigiani [1018, 1516]. Parsimonious [2417]. part [1695, 446, 447, 2883, 2533, 1940, 1941, 2683, 2684, 1426, 1427]. Partial [2011, 445, 877, 3007, 2749, 2817, 2956, 1745, 721, 3110, 3047, 3162]. partially [2695, 1600, 2589, 1662]. participant [3184, 3251]. participants [1128]. particular [1103, 641, 1057]. partition [248, 2905, 2188]. partitioning [936, 497, 1822]. partnerships [2252]. Pasant [1193]. Path [2495, 2175]. Pathways [353]. Patient [529, 26, 697, 1207, 2356, 933, 2242, 3335, 2588, 2201, 2545, 407, 2760, 3062, 2761, 2778, 1791, 2434, 1686, 3279, 31, 1602, 2829, 1885, 547]. patient- [697]. patient-level [2356]. Patient-specific [26]. patients [1268, 651, 859, 564, 175, 2926, 328, 1111, 152, 849, 922, 2117, 546, 1369, 282, 1090, 1735, 1513, 2419, 1454, 1579, 150, 773, 621, 2603, 2318, 502, 972, 293, 2354, 2780, 765, 1303, 389, 195, 1798, 3378]. Patricia [526]. Patrick [2499, 3138]. Pattern [1027, 1161, 1770, 2013, 166, 1104, 749, 3198, 1499]. pattern-mixture [1161, 1770, 2013, 749]. patterns [1641, 3089, 2731, 1007, 226, 2690, 1442, 839]. Paul [511, 1609, 2452, 257, 1937]. Pawitan [1592, 1711]. Pbk [1549]. PCA [2596]. PCR [1723]. PCR-based [1723]. PD [2987, 2251]. Peacock [1180, 3192]. peak [1574]. Pearl [843]. Pearson [810]. Pearson-type [810]. pediatric [2756]. pedigrees [1214]. Pei [201, 495, 287, 3203]. Penalized [853, 2255, 1308, 3132, 1895, 238, 2205, 2761]. penalties [2573]. Pencina [2638, 2641, 3190, 2640, 2643, 2642, 2637, 2639]. penicillin [1866]. penicillin-resistant [1866]. People [870, 552]. Pepe [1787, 387]. per-protocol [1578]. percentage [295, 841]. percentile [1000]. percentiles [1436]. Pereira [128]. Pereira-Maxwell [128]. perfect [2297, 3136]. perform [1412, 2316]. Performance [2330, 3307, 2551, 912, 2336, 15, 2443, 2910, 2234, 2311, 39, 1161, 2179, 3024, 766, 3075, 1279, 2909, 1858, 194, 3224, 3037, 1173, 38, 2850, 2674, 812, 3043, 2307, 864, 1739, 1728, 608, 2027, 1675, 2722, 1962, 2404]. performances [2780]. performed [2862]. perinatal [2916, 647]. period [1625, 2436, 2735, 1703, 2895, 2470, 1971, 806, 3292, 2734, 334, 2244]. periodic [1764]. periodically [753, 754]. Periodograms [848]. periodontal [3353]. periods [1280, 3055]. Permutation [279, 48, 386, 812, 1901, 1951, 1813, 137, 1337, 2543, 1927, 1725, 1637, 3099, 391, 2138]. permutations [2294]. Perrier [1728]. person [3171]. person-time [3171]. personal [1591, 1746, 1358]. Perspective [464, 2262, 745, 462, 149, 308, 1005, 2146, 3042, 1402]. Perspectives [1217, 1905, 463, 893, 905]. pertussis [2119]. PET [1297, 1243]. Peter [2245, 144, 782, 127, 3070, 2909, 3311, 2772, 2773, 1133, 2452, 525, 2774, 259, 1244]. Peters [1794]. Petersen [1169]. Peterson [958, 2198]. Petit [40]. Peto [1983]. PH [3106, 2803]. pharmaceutical [2824, 578, 201, 1182, 265, 308, 2786, 2345, 2892, 2754]. pharmacodynamic [3243, 1312]. pharmacodynamics [1030]. pharmacoeconomic [1043]. pharmacogenetic [3090, 1701]. pharmacogenetics [3043]. pharmacokinetic [3243, 2880, 3357, 2425, 2203, 1558, 336, 583, 1702, 2318, 2121, 1312, 425]. pharmacokinetic/pharmacodynamic [3243, 1312]. pharmacokinetics [1030, 1467, 1654, 874]. pharmacokinetics/pharmacodynamics [1030]. pharmacological [2114, 1376]. pharmacology [2006]. Pharmacovigilance [309, 3233]. Phase [559, 3065, 2672, 2987, 2136, 759, 405, 2331, 529, 2158, 1487, 2411, 3340, 1767, 2088, 214, 162, 3009, 2397, 2056, 2479, 2776, 2777, 576, 979, 996, 1314, 2673, 3164, 406, 2985, 3388, 1663, 1904, 1789, 1832, 2394, 2419, 407, 2798, 3308, 1462, 3154, 805, 882, 2712, 2022, 2577, 448, 2258, 1417, 3279, 1021, 797, 556, 823, 2160, 1025, 1758, 2265, 2829, 3158, 3066, 425, 1200, 3194, 2995, 1925]. phase-type [2419]. phenomena [1909]. phenotype [1178, 1642, 2306]. Philadelphia [2922]. Philip [566]. Phillip [2138, 1311, 1725]. Phillips [975]. phylogenetic [2976]. Physica [285]. Physica-Verlag [285]. physician [621]. physicians [881]. Pickles [1067, 2109]. PIE [159]. piecewise [1672, 338, 722, 3017]. piecewise-constant [722, 3017]. Piegorsch [439]. pill [645]. Pillai [2006]. pilot [96, 1417, 1440]. pilots [1152, 2522]. Pintilie [2489, 2476]. Pisani [258]. Pitacco [143]. pitfalls [1505]. pituitary [560]. pivotal [1715, 1714, 1716, 1713]. PK [2251]. PK/PD [2251]. PlA [1094]. Placebo [1518, 2724, 108, 1456, 1038, 1107, 2102, 3362, 3341, 1034, 2159, 2057]. placebo- [2159]. placebo-challenging [108]. placebo-controlled [2724, 1107, 2102]. Placebo-free [1518]. Plackett [1440]. plane [839]. Planning [617, 1790, 2249, 2521, 35, 2022, 2412, 3385]. Plateau [3020]. platelet [1094]. platform [2702]. play [705]. play-the-winner [705]. Playing [1390]. Plenum [542]. plot [450]. plots [244, 461, 2687, 1675]. Plus [1750, 3180, 494, 1182, 1169, 869]. PMI [2993]. Pneumonia [1256]. pneumoniae [1866]. Pocock [297]. point [2119, 1641, 631, 3040, 1430, 612, 163, 610, 2548, 724, 2541, 2906, 2319, 2634, 1050, 1976, 1504]. points [1328, 1795, 652, 1014, 1201, 2066, 1881]. Poisson [2379, 2919, 2180, 935, 915, 8, 3375, 2757, 1899, 1368, 1333, 1895, 1465, 2405, 2372, 2870, 1371, 3268, 1688, 1797, 1659, 2236, 2799, 394, 951, 1824, 2341, 2892, 2497, 594]. Poisson-distributed [2405]. Poisson-generated [1899]. Poisson-Poisson [2180]. Policy [904, 1632, 2992, 1622, 3004, 871]. pollution [2448, 247, 1321, 1604, 1667, 2215, 2171]. Poly [1678, 1165]. Poly- [1678, 1165]. polymerase [2624, 2291]. polymorphic [1092]. polymorphism [1094]. polynomial [1935, 1015]. polynomials [1051, 1889, 1464, 3138, 2016]. pool [2681]. pooled [1155, 532, 3042]. Pooling [2422, 1575, 2649]. Population [1030, 709, 1286, 525, 3096, 1772, 207, 470, 1846, 1844, 2314, 1444, 262, 99, 2203, 265, 2117, 2348, 336, 1292, 1232, 3261, 2126, 977, 2455, 1434, 1059, 874, 1543, 1984, 1196, 429, 2555, 1445, 2108, 1312, 2569, 1602, 264, 268, 1544]. population-averaged [1059]. Population-based [1286, 1772, 2126]. populations [640, 2225, 72, 731, 158, 1099, 294, 1687, 2428, 1344, 1818, 2807, 1962, 1447, 496]. Porcher [2344]. Portability [460]. portfolios [2580]. Portnoy [1610]. posed [1726]. Position [2064]. positive [1615, 190, 2717, 1454, 949]. possibilities [3242]. possible [2456]. post [192, 1249, 634, 1616, 1394, 2060, 286]. post-AMI [2060]. post-drug [1394].

Post-Graduates [286]. post-mismeasured [1616]. post-operative [192]. post-randomization [634]. post-treatment [1249]. Poster [482]. Posterior [235, 2793]. posteriori [2584, 2819]. Postma [2770]. postmortem [2993]. postpartum [1280]. potency [1441]. Potential [837, 2716, 2508, 2363, 2590, 318, 927, 1535, 1809]. potentially [3013, 2264, 3258]. potentials [2181]. Potter [1951]. Poursat [1750]. Power [290, 444, 1982, 1107, 2087, 1521, 447, 945, 1266, 2091, 2089, 1045, 2221, 1486, 2090, 2433, 3248, 3160, 2481, 2086, 3032, 931, 19, 2093, 1453, 3159, 3118, 2094, 575, 3379, 1160, 1448, 56, 2097, 3045, 3224, 2113, 1804, 2156, 2352, 802, 2343, 555, 2487, 3356, 2941, 2593, 1501, 2577, 1098, 1063, 1961, 2765, 1410, 1962, 178, 813]. Powerful [2520, 2958, 2884, 1579, 1910, 1995]. pp [1460]. Practical [159, 3212, 3289, 434, 1158, 1402, 2617, 440, 2644, 2765, 2754, 323, 115, 1750, 1295, 262, 574, 1802, 2777, 2409, 2343, 2841, 1644, 2774, 919, 268, 2820]. practicalities [364]. Practice [1193, 1065, 94, 568, 840, 368, 498, 1096, 1825, 365, 144, 1531]. practiceand [902]. practices [1632]. Practitioner [452]. pragmatic [771, 364, 3138]. Prakasa [436]. Pranab [435]. pre [1293, 1616, 498, 1394, 2786, 3066]. pre- [1616]. pre-drug [1394]. pre-intervention [498]. pre-marketing [2786]. pre-screening [1293]. pre-specified [3066]. precancerous [852]. precise [1453, 2676, 2328]. Precision [2874, 768, 2698, 1551, 619, 1210, 1312]. preclinical [163, 2087]. predict [1178, 972]. predictability [1198, 1879]. Predicting [527, 2793, 2241, 2117, 973, 1344, 3004, 15, 2432, 39, 1536, 1239, 1395, 3374]. Prediction [756, 849, 1133, 1689, 773, 1176, 2579, 1591, 3359, 551, 2166, 2638, 1737, 251, 176, 328, 1444, 3187, 2942, 3354, 1073, 2643, 2688, 1822, 2602, 508, 2492, 2730, 711, 2770, 1877, 2074, 3124, 663, 2620, 1358, 3163, 1253, 1709, 1310, 1787]. predictions [610, 2319]. Predictive [2977, 2544, 1138, 3259, 2638, 1950, 2641, 795, 793, 2065, 2793, 2640, 2643, 801, 2445, 2642, 1088, 949, 2384, 2636, 2637, 794, 2712, 2577, 235, 1468, 1053, 2067, 2639, 178]. predictor [330, 471, 1078, 1633, 3234, 2373]. predictors [2285, 3190, 2304, 1689, 2322, 1911, 2618]. Preface [1567, 1029, 363, 693, 261, 316, 783, 2605, 2143, 1216, 2966, 567, 3072, 886, 222, 1068, 1621, 963, 635, 304, 54, 2211, 1255, 456, 1875, 648]. preference [2179, 1791]. preference-based [2179]. preferences [847]. prefrontal [2084]. Pregnancy [818, 37, 2367, 1130]. preliminary [2583]. premature [1262]. prenatal [2179, 2204]. Prentice [1530, 1916, 1396]. preparing [1387]. Prescott [358, 1937]. prescribing [712]. presence [645, 730, 1649, 1302, 1126, 3022, 2931, 175, 1719, 2175, 2667, 3123, 1677, 2097, 2733, 1718, 1551, 2461, 2607, 1552, 711, 2854, 3341, 189, 416, 2183, 502, 2551, 2243, 215, 2369, 2786, 422, 3141, 2351, 2136, 3145, 1122, 1347, 173, 661]. Present [2064, 801, 849, 1173, 1053]. Presentation [1401, 295, 841]. Presentations [481, 482]. presented [2698]. Presenters [458]. presenting [1387, 3192]. Preservation [2657]. preserving [1390]. prespecified [2480]. Press [1592, 542, 2754, 601, 144, 885, 1158, 1608, 843, 1415, 1550, 1134, 399, 420, 258, 1787, 1245, 1180, 1067, 1802, 2409, 2503, 1593, 435, 1133, 1142, 2857, 1531, 1595, 1671, 3192, 3056, 440, 1143, 1532, 869, 2519, 146, 2109, 545, 540, 1065, 187, 1825, 1698, 259, 433, 210, 401, 1549, 94, 462]. pressure [553, 323, 251, 1536, 2554, 659, 1480, 1815]. pressures [1000, 2976]. Prevalence [56, 2457, 72, 731, 876, 2095, 158, 76, 3270, 2723, 1444, 214, 840, 2012, 764, 2253, 411, 2016, 2700, 1831, 1445, 943, 521, 2081]. Prevalence-dependent [2457]. Prevalence-free [56]. prevalent [2030, 177, 166, 2980]. prevent [640]. preventable [3196]. prevented [2569]. prevention [323, 1737, 3238, 204, 1295, 1294, 1535, 2746, 594]. previously [1316]. Price [1530, 1594, 91, 341, 819, 817, 1460, 1647, 93, 1565, 454, 1592, 421, 542, 3139, 198, 509, 3222, 2138, 1752, 1786, 2754, 601, 144, 455, 360, 2824, 526, 2604, 885, 1425, 434, 497, 1608, 286, 3203, 782, 361, 1751, 143, 200, 127, 183, 1750, 1873, 301, 145, 402, 843, 362, 288, 2501, 3204, 1415, 2006, 129, 603, 1566, 1550, 90, 543, 343, 1402, 358, 600, 2004, 1119, 199, 399, 2856, 42, 1461, 420, 1438, 602, 258, 40, 201, 1787, 961, 3311, 303, 511, 1180, 1067, 1548, 1753, 131, 3312, 1182, 541, 3057, 211, 452, 182, 566]. Price [1018, 342, 844, 1774, 1802, 3182, 196, 3181, 2399, 417, 1181, 92, 2409, 1533, 2503, 2140, 453, 1159, 1593, 104, 818, 884, 495, 2857, 780, 130, 300, 302, 1118, 3180, 1531, 524, 299, 1595, 2959, 221, 1671, 821, 1609, 132, 3347, 962, 142, 1872, 1801, 3192, 359, 2960, 781, 357, 494, 1840, 2855, 1437, 525, 2559, 3056, 285, 141, 2858, 544, 869, 2519, 872, 419, 2961, 43, 1546, 2005, 146, 106, 2109, 1471, 2139, 2502, 2859, 2719, 287, 2295, 870, 3122, 260, 1648, 545, 1516, 185, 2841, 3138, 3231, 437, 2345, 128, 2034, 298, 400, 418, 820]. Price [1547, 103, 540, 1517, 1065, 1610, 510, 187, 1504, 2718, 284, 1825, 147, 105, 2720, 259, 184, 496, 197, 1066, 186, 210, 1684, 401, 1549, 1503, 1416, 871, 41, 94, 2755, 1839, 439, 436, 438, 435, 440, 1698, 433]. primary [2312, 585, 1146, 367, 1668, 3013, 1513, 364, 498, 2373, 3038, 2746, 1184]. primer [511]. Princeton [94]. principal [2420, 3176, 3297, 940, 1748]. principle [1463, 3115]. principled [2971]. Principles [569, 1531, 401, 975, 374, 974, 288, 1119]. Prior [2483, 1078, 3259, 382, 3379, 1964, 1676, 2299, 1773, 406, 3343, 3356, 3147, 759, 2390]. priors [2088, 558, 2483, 2299, 237]. prison [507]. Probabilistic [1943, 1268, 470, 233, 2509, 1370, 2198, 542]. probabilities [1950, 1101, 1004, 1663, 1125, 720, 3301, 1528, 1303, 3306, 1404]. Probability [781, 2355, 3025, 3032, 3229, 1453, 1795, 3024, 1112, 1341, 2954, 1100, 499, 220, 798, 1008, 3246, 1828]. probability-of-censoring [3229]. probably [1432]. probit [1928, 2554, 2274]. problem [381, 1658, 855, 1473, 1726, 2071, 946]. problems [2320, 2288, 955, 47, 902, 2360]. procedure [355, 1453, 676, 1413, 682, 1589, 1099, 1185, 1827, 3322, 2867, 2633, 3119, 1356, 3387, 16, 2159, 263, 700, 429, 2986, 354, 2333]. Procedures [1949, 563, 1198, 3143, 1580, 2875, 2589, 627, 2151, 1461, 779, 2520, 474, 1246, 1776, 2609, 2532, 1264, 1765, 2850, 112, 1313, 2307, 958, 2238, 2247, 264, 1906, 257, 1503].

Proceedings [1608]. process [3022, 2040, 935, 863, 1559, 3361, 2308, 1688, 1913, 338, 1114]. Processes [436, 1290, 100, 631, 880, 2863, 1504]. product [3020, 2905]. production [2354]. products [1328, 2020, 2021, 570]. profile [1227, 2899, 3312, 3304, 3053, 271, 2708]. profiles [751, 1505, 1339, 2378, 3077, 1269, 999, 320]. Profiling [2722, 2050, 1409, 1867]. Prognosis [55, 1051, 565, 1454, 1409, 1470]. Prognostic [31, 110, 936, 58, 967, 3125, 102, 1405, 67, 2307, 2707, 1454, 28, 1010, 2387, 834, 1343, 1330, 57, 63, 971, 1890, 324]. Program [1908]. programme [412, 594]. programmer [3204]. programmes [1019]. Programming [419]. programs [1490]. Progress [525]. progressing [891]. progression [3285, 149, 2043, 1522, 2979, 934, 679, 2541, 2271, 2711, 3266]. progression-free [3285, 2541, 2711]. progressiva [865]. progressive [3249, 422]. progressively [669]. project [3314, 3315, 3313, 3319, 3316, 3317, 3318]. Projecting [530, 158]. projection [1127, 477]. projections [1584, 943]. promising [1353]. Promoting [155, 156]. promotion [2301]. Proof [3116, 1732]. propagation [767]. Propensity [3198, 3191, 3200, 1706, 2060, 2281, 2282, 2443, 2771, 2775, 2910, 3320, 1806, 3024, 2909, 2772, 2068, 2773, 2015, 3297, 2302, 476, 1634, 2237, 2415, 1495, 2452, 3101, 3199, 2619, 2774, 2901, 2364]. propensity-matched [2772]. propensity-score [2771, 2775, 3320, 2772, 2773]. Proper [2851, 1264]. propertie [2819]. Properties [1152, 272, 1997, 2373, 763, 1868, 2584, 2227, 169, 479]. Proportion [1236, 1778, 1846, 2120, 729, 1811, 1850, 2543, 1735, 615, 1681, 2630, 761, 9, 1864, 3145, 1128, 1431, 2628, 2418, 3346]. Proportional [924, 507, 317, 2219, 2086, 1720, 2052, 689, 3083, 2592, 3125, 1009, 2942, 928, 1177, 558, 3244, 2223, 1895, 1519, 1884, 2256, 2906, 2813, 3126, 2496, 3018, 2989, 1620, 942, 2107, 1808, 834, 1455, 2714, 1124, 109, 1890, 1972, 2762]. proportional-hazards [558, 834]. proportional-odds [834]. proportions [1275, 1639, 538, 3135, 2122, 13, 1037, 1453, 736, 213, 2804, 1936, 3298, 2464, 11, 1724, 7, 2475, 2855, 592, 2559, 3021, 2032, 29, 1450, 1426, 1427, 2727, 1801]. proposal [1965, 2110, 3192]. proposals [3158]. proposed [367]. proprietary [1328]. Pros [137]. Proschan [355, 2501, 354]. Prospective [711, 658, 1737, 1011, 2673, 1466, 2339, 1857, 327]. Prospects [2064]. prostate [3323, 1869, 732, 973, 2108, 2729, 2889]. prostatectomy [1491]. prostheses [2587]. protection [937, 2010, 292, 1635]. protein [2764]. protocol [1578, 3283, 1984]. provider [1508, 2050, 1867]. provider-level [2050, 1867]. providers [751]. providing [3187]. Pseudo [1862, 2104, 1345, 2189, 1812, 3008, 2523, 3173]. pseudo-data [2189]. Pseudo-likelihood [2104, 1345]. pseudo-observations [3008, 3173]. Pseudoscore [2427]. Pseudoscore-based [2427]. psychiatric [1222, 1579]. psychiatry [1225, 1012, 2102, 453]. psychological [2563]. Psychologists [357]. psychometric [171]. psychometrics [3263]. psychopharmacology [1222]. psychosocial [2524]. PTE [1236]. pteryomaxillary [560]. pubertal [1814]. Public [417, 232, 463, 111, 1626, 2921, 1535, 1627, 2362, 1233, 1623, 2171, 1824, 480]. publication [1649, 803, 2189, 2937, 2112, 385, 2551, 2685, 860, 2280, 1122, 3192]. Publications [603, 1461, 131, 302, 1156, 132, 257, 870, 437, 187, 284]. publicly [1389]. Published [2452, 1412]. Publishers [421, 2006, 1193, 1066]. Publishing [185, 897]. pulmonary [765]. pulsatile [848]. pulse [848]. Puri [1193]. putative [1456, 1038, 707]. Putting [365, 578]. pylori [1430, 1900].

Q [3310]. QAL [3249]. Qin [3070]. Qing [3386]. QLQ [667, 971]. QLQ-C30 [667, 971]. QT [3302, 2498, 1473, 1105]. Qu [3384]. Quadratic [3365]. quadrature [2813]. qualitative [1360, 1329]. Quality [3289, 1190, 139, 3212, 2924, 1768, 2724, 2242, 2445, 3090, 2943, 2013, 619, 749, 1901, 694, 1027, 3249, 312, 2421, 667, 2722, 3185, 3173, 1666, 680, 997, 144, 544]. quality-adjusted [3185, 3173, 997]. quality-of-life [1768, 2013]. Quan [1561]. quantification [2624, 1825]. quantify [3290, 3266]. Quantifying [2703, 382, 787, 2987, 3207, 1395, 1917, 1230, 2369, 230, 1260, 2999, 3384]. Quantile [1462, 1996, 1449, 1643, 1078, 1633, 1308, 918, 2777, 3120, 2415, 1469]. quantile-categories [1633]. quantiles [492, 339]. Quantitative [455, 221, 3105, 3368, 3373, 3367, 2888, 199, 1279, 3369, 811, 630, 1728, 3370, 46, 2291, 3371, 3372, 1815, 1080, 1500, 1934, 1250, 2191, 2904, 962]. quarantine [1855]. quasi [736, 2284, 238]. quasi-exact [736]. quasi-imputation [2284]. quasi-likelihood [238]. questionnaire [2011]. questions [1180, 2531]. quintile [1634]. quintile-stratified [1634]. Qureshi [437].

R [1530, 2327, 2344, 1752, 1786, 2820, 434, 1750, 129, 603, 90, 1402, 1119, 42, 2640, 452, 2399, 92, 1925, 3380, 2653, 3310, 2642, 780, 524, 1563, 3383, 2770, 416, 2855, 2559, 3056, 141, 872, 1471, 1648, 1206, 298, 400, 105, 41, 94, 2839, 2959, 2005]. Rabe [1751, 844, 2857, 2755]. Rabe-Hesketh [1751, 844, 2857, 2755]. race [1083]. Racial [2925, 2918, 2927]. Racial/ethnic [2925, 2918]. radial [651]. radiation [2048, 2372, 2556, 628]. radical [1491]. radiologists [2336]. radon [2781, 927, 2555, 1734]. Rafael [1516]. Rainer [884]. raking [2302]. Ralph [2754, 3190, 2345]. Ramon [2604]. Rand [872]. Randall [782]. Random [2376, 2139, 2053, 2407, 3061, 379, 877, 2181, 2907, 1242, 1349, 1087, 1161, 2375, 120, 3040, 2048, 24, 1505, 1086, 1288, 3113, 2317, 519, 1821, 1014, 3120, 1172, 2614, 3035, 1510, 847, 2660, 3188, 2405, 2730, 2850, 2239, 2849, 2883, 2319, 3268, 2920, 657, 2106, 3094, 1479, 498, 924, 1799, 2377, 2186, 2603, 912, 312, 2235, 2270, 322, 2539, 2908, 1836, 776, 317, 647, 3137, 2193, 790, 1147, 486, 594, 3102, 2009]. random-coefficient [1161]. random-digit-dialling [1086]. random-effect [847]. random-effects [2048, 1505, 2660, 2405, 2850, 2377, 2603, 2235]. Randomised [1066].

Randomization [817, 2584, 3339, 498, 2440, 2481, 1198, 1862, 2439, 1907, 2856, 3097, 634, 1880, 3226, 654, 1452, 769, 2840, 2826, 2709, 2100, 380, 3227, 3091, 370, 1569, 1740, 1763, 1315, 2825, 2930, 2839, 2819]. randomization-based [1315]. Randomized [602, 2673, 604, 1, 1792, 527, 2540, 2297, 311, 2894, 323, 3229, 2231, 1696, 3334, 708, 3157, 2740, 969, 907, 2563, 1428, 792, 375, 627, 699, 2521, 1969, 3028, 1060, 3170, 2479, 791, 1742, 374, 2983, 1704, 1926, 3027, 980, 202, 976, 1247, 1342, 1616, 656, 2446, 2558, 2630, 3136, 1266, 2102, 657, 719, 1028, 757, 2200, 958, 2091, 606, 364, 3098, 466, 1927, 189, 416, 1171, 3059, 254, 3307, 2689, 253, 366, 2232, 582, 1367, 2421, 2053, 372, 705, 797, 1063, 1754, 3099, 109, 572, 2523, 1301, 3230, 2531, 2720, 2971, 373, 1366, 2007, 2250, 371, 964, 1494]. randomized [2467, 2999, 658, 1599, 647, 3092, 889, 1026, 1662, 2018, 484, 997, 1946, 365, 2101, 2410, 867]. randomizing [367]. randomly [1900]. randomness [1620, 1960]. range [2780]. ranges [3161, 1408, 1555]. ranging [589]. Rank [1926, 435, 993, 2287, 2942, 1337, 2087, 2861, 2848, 2869, 1559, 1725, 2107, 2274, 1431, 391, 1828, 3160, 1400]. Rank-based [993, 2848]. Ranked [1846, 2274]. Ranking [1631, 3294]. ranks [2881]. Rao [1786, 436, 105]. rapid [2363, 1624]. rare [2919, 2234, 2907, 140, 605, 1095, 2182, 2539, 2908, 3137]. Rasch [494, 700]. Rasmus [1504]. rate [2874, 2030, 1361, 2513, 289, 2768, 119, 1428, 154, 2895, 2806, 96, 1246, 2627, 774, 3171, 1990, 1371, 2468, 1659, 1096, 2243, 753, 754, 837, 860, 1019, 2517, 2656, 2401, 2765, 2929]. Rater [191, 2102, 1923, 2554, 2294, 799, 2014]. raters [800, 2014]. rates [1792, 1615, 3048, 1117, 1545, 2379, 2953, 1316, 410, 8, 673, 1213, 1668, 1073, 1116, 2979, 48, 386, 1232, 231, 1084, 2256, 695, 1095, 2188, 1028, 1801, 1096, 1669, 2713, 1092, 557, 1322, 653, 1429, 521, 2081, 3240]. ratings [1556, 329, 2275]. ratio [3041, 1117, 1545, 3025, 2919, 2122, 2878, 2209, 2386, 487, 296, 281, 1151, 3024, 1682, 599, 2909, 1116, 927, 2371, 1446, 3256, 954, 1292, 1544, 2591, 512, 277, 276, 917, 1371, 1681, 2630, 3005, 2326, 720, 2855, 3301, 2220, 2559, 1659, 3021, 2868, 1543, 322, 1669, 2655, 1322, 2866, 2467, 2999, 3052, 3288, 1810, 193, 283]. rationale [571]. ratios [2443, 2910, 70, 1031, 1674, 2472, 160, 111, 474, 774, 688, 1056, 1418, 2528, 3274, 2662, 2681, 2732, 825]. Raton [1594, 1460, 2824, 1364, 3203, 1751, 1873, 362, 2004, 3311, 3312, 3057, 1205, 2399, 1533, 780, 3180, 2959, 821, 1609, 3347, 2960, 1840, 2858, 2005, 2139, 2719, 2295, 3122, 2841, 2034, 1610, 1504, 2720, 1839, 1244, 1503, 2755]. Rattanasiri [3312]. Raven [91]. raw [3060]. Raymond [131]. Raynor [259]. Re [3097, 96, 1897, 330, 682, 574, 1589, 1099, 464, 3120, 2156, 769, 1265, 377, 1967, 376]. Re-calculating [96]. re-emerging [464]. Re-estimating [3097]. re-estimation [682, 574, 1589, 1099, 2156, 769, 1265, 377, 1967]. re-estimations [376]. re-evaluation [3120]. Re-use [1897]. reaction [2291]. reaction-based [2291]. reactions [2624, 431, 825]. reactivation [2515]. reader [1189, 2647]. readily [2441]. Reading [3112]. readings [1565, 1690]. real [1956, 1189, 2843]. real-time [2843]. realistic [2894]. really [2033, 2503, 1775]. Reardon [1158]. rears [697]. Reasonable [2110, 264]. Reasoning [843]. reassessment [1905, 2582, 1041, 576, 406, 675, 3308, 759, 3033, 589]. rebound [829]. Rebuttal [2908, 3137]. recalculation [575, 2159]. recall [1970]. recapture [1746, 613, 915, 2920, 2305]. Receiver [807, 3139, 24, 1682, 3007, 3123, 1707, 1446, 762, 763, 2408]. recipients [679, 1334]. reclassification [2638, 2641, 2640, 2643, 2644, 2636, 3134, 2637, 3133, 2639]. recognition [1104]. Recommendation [947, 2751]. recommendations [2488, 2774]. Recommended [3179]. Reconsidering [1292, 1544, 1543]. reconstruct [32]. reconstruction [2255, 3002, 536]. record [779, 2925, 1447]. Records [259, 2356, 792, 1232]. recovery [1491, 1092, 390]. Recruitment [699, 2579, 909, 1424, 1784]. rectangle [428]. recurrence [2338, 1240, 973, 121]. recurrent [1792, 3048, 2000, 44, 1636, 2552, 3282, 935, 3086, 2267, 2600, 21, 2705, 2609, 2633, 2835, 660, 1680, 1913, 1823, 703]. recursive [808, 1822, 497]. recycling [3153]. redistribution [1618]. reduce [1572, 1536, 1312]. Reduced [2942, 2107]. Reduced-rank [2942, 2107]. reduces [2515]. reducing [1879]. reduction [2880, 487, 1072, 2320, 2237, 2535, 735, 2620]. refer [1447]. Reference [918, 3155, 18, 3161, 387, 2520, 1957, 1835, 1834, 2687, 488, 960, 1408, 1164, 278, 641, 340, 2708, 1996, 450, 3215]. referenced [2083]. referral [1870]. refined [665, 3005]. Reflections [306]. reformulation [2849]. Regaining [3032]. regarding [894, 3024, 265, 1544, 1871]. regimens [1136]. regimes [2217]. Regional [60, 3299]. regions [1164]. register [948]. registers [2704]. registrations [840, 1447, 2305]. Registry [2350, 565, 3210, 1820, 943]. Regression [3048, 1157, 702, 1274, 933, 211, 182, 1493, 1075, 3257, 1052, 2545, 735, 348, 1721, 400, 1783, 3173, 1588, 517, 2497, 1048, 2406, 810, 3323, 1209, 1121, 1951, 2344, 1179, 1302, 1361, 1643, 2432, 2865, 471, 2993, 1163, 2623, 2389, 2945, 1762, 2431, 1078, 2287, 2292, 1944, 3277, 332, 1308, 3344, 3300, 3159, 777, 1807, 3174, 1072, 580, 2886, 1750, 2731, 3342, 2285, 2465, 2629, 1009, 1935, 726, 393, 2935, 598, 2781, 2942, 3002, 1070, 2691, 918, 2017, 2833, 2757, 1177, 3075, 1001, 612, 558, 511, 1749, 1252, 1074, 855, 2218, 1403, 1802, 1014, 2991, 880, 934, 2557]. regression [2338, 1997, 2973, 3376, 2717, 3254, 1822, 1687, 549, 3224, 1524, 1764, 2882, 48, 386, 1333, 1172, 752, 966, 2803, 1889, 1895, 2298, 1049, 1261, 588, 2304, 3126, 2100, 2870, 1761, 3196, 27, 657, 1016, 1378, 26, 3076, 1381, 88, 1201, 2534, 1927, 1587, 1797, 1360, 1047, 2316, 944, 2603, 856, 1050, 1409, 3008, 2107, 2430, 47, 1637, 677, 1348, 2458, 1140, 1847, 1947, 753, 754, 384, 1469, 2387, 3114, 1230, 1911, 591, 2269, 1908, 839, 3138, 3274, 623, 2619, 2434, 298, 1124, 340, 2799, 31, 110, 1468, 2383, 2863, 624, 3064, 3106, 109, 788, 2594, 1527, 1045]. regression [1410, 2219, 394, 1253, 951, 1162, 1538, 1948, 2193, 1996, 1046, 1931, 2062, 1824, 2341, 2192, 2337, 2692, 594, 1449, 1995, 3309, 3218, 661, 713, 2133, 819, 2859, 1157, 2140, 524, 433]. Regression-based [735]. regression-effective [2219]. regression-to-the-mean [3224]. Regressions [524, 697]. regressive [1888]. regressors [1360]. regret [2217]. regret-based [2217]. regular [1457]. regularization [2505]. regulate [2191]. regulated [578]. regulation [643, 3182]. Regulatory [893, 269, 1038, 1295, 2576, 903, 1296, 2657, 904]. Reid [780]. Reilly [1711]. Reinforcement [3337]. Reinhard [343]. Reiser [3221]. rejection [1796]. rejective [3143]. Rejoinder [3373, 795, 2438, 1598, 3319, 2666, 3332, 378, 1340, 3011]. relapse [282]. relapses [1894]. relate [3240]. Related [3289, 3212, 2181, 2924, 1782, 3161, 1628, 550, 2837, 1408, 2084, 2710, 73, 1831, 1303, 2721, 1555, 1666, 1490]. relating [3208, 1299]. relation [998, 561, 2450, 2260]. relationship [1209, 1076, 154, 3058, 2757, 626, 2300, 3295, 765, 313, 2342, 1995]. Relationships [2878, 1030, 3020, 2767, 2043, 2494, 428, 572]. Relative [2474, 2624, 531, 781, 2410, 1272, 1615, 982, 2878, 2766, 2886, 175, 1274, 1982, 2, 1177, 1888, 3075, 3228, 2698, 609, 1441, 2392, 1889, 1054, 696, 512, 1949, 2616, 1799, 1459, 2387, 2778, 815, 1891, 1058, 1890, 790, 50, 2115, 173]. relaxation [2222]. Release [2521, 2597]. relevance [1521]. relevant [2089]. Reliability [2830, 151, 2473, 792, 1923, 2554, 2242, 1975, 2012, 938, 2014, 1399]. remain [2537, 2788, 2787]. remains [1230]. Remarks [3199, 1069, 2347, 483, 459, 2912, 2570]. remedies [59]. remission [2603]. REML [1854, 1113, 3281]. removal [1320]. remove [1078, 1264]. renal [1924, 2608, 3034, 422]. reorientation [2638, 2643]. repair [1809]. Repeatability [2482]. Repeated [2169, 5, 1822, 2902, 2300, 2611, 412, 89, 712, 1586, 1647, 1331, 3168, 113, 1152, 2842, 3040, 61, 199, 1520, 2079, 190, 181, 1883, 796, 2366, 1062, 1788, 2140, 2444, 2405, 84, 180, 802, 3034, 2100, 2870, 917, 2446, 2558, 2078, 719, 1579, 3094, 1379, 1171, 560, 681, 443, 322, 582, 2080, 2210, 2066, 2662, 335, 371, 1664, 445, 391, 685, 337, 2337, 297]. Repeated-measures [412, 1586]. repeatedly [1992]. replace [590]. replacements [1257, 1513]. Replacing [914]. replicated [1690, 1269]. replicates [2389, 2945, 617, 632]. replication [3324, 3342, 4, 857, 946]. Reply [398, 415, 1529, 2752, 1607, 814, 1710, 587, 959, 816, 1590, 3109, 1727, 1325, 2135, 539, 2001, 758, 2821, 1544, 3107, 1055, 1697, 126, 804, 30, 1375, 1712, 388, 1170, 1938, 356, 1562, 1483, 2823, 715, 2677, 1515, 1646, 2003, 1800, 1785, 2818, 868, 2199, 2453, 3245, 2246, 1017, 2380, 2184, 858, 1323, 1670, 1683, 3220, 2137, 2477, 2490, 2051, 1281, 2185, 2451, 1564, 2476]. reported [1886, 3030, 3060, 2867, 2686, 2900, 3031]. reporters [462]. reporting [77, 1305, 1386, 318, 14, 902, 1083, 1618, 78, 825]. reports [3131, 2757, 791]. representations [2837]. Reproducibility [293, 798, 617, 117]. reproductive [1164, 642, 2843]. require [1146]. required [374, 50]. requirements [768, 674, 3170, 447]. Resampling [2974, 2307, 264]. Resampling-based [2974, 2307]. resampling-free [264]. rescue [604]. Research [1349, 603, 141, 1065, 1283, 1643, 352, 2736, 3352, 369, 728, 1284, 161, 265, 2773, 319, 2733, 2349, 1222, 830, 1298, 2110, 2201, 1535, 2362, 2360, 897, 1623, 785, 2524, 2763, 896, 480, 817, 1118, 1206, 3122, 286, 399]. research-based [265]. researching [2360]. residence [352]. residential [2781]. Residual [692, 663, 1527, 1572]. Residual-based [692]. Residuals [134, 2461, 2132]. residuals-based [2461]. resistance [2172]. resistant [1866, 3104, 1369]. resolution [520, 1500]. resonance [2255, 983]. respect [1037, 2117, 2330]. respecting [56]. Respiratory [2371, 1508, 1096, 1267]. Respondent [3260]. Respondent-driven [3260]. responder [1521]. responders [676, 1413, 1190, 1552]. Response [3202, 3303, 975, 3150, 83, 3134, 545, 114, 3359, 3368, 3373, 1970, 3243, 3020, 931, 937, 87, 1700, 290, 1033, 2099, 1109, 2856, 2554, 214, 1790, 2494, 3144, 3369, 1060, 2011, 2445, 2973, 3226, 2737, 2893, 2278, 2840, 626, 1084, 2372, 1832, 2994, 1265, 2368, 1436, 169, 3341, 761, 1665, 52, 3370, 596, 1462, 499, 1348, 2243, 1847, 3114, 3371, 3372, 2258, 700, 2956, 97, 556, 2351, 218, 572, 1583, 2594, 2505, 2573, 680, 2062, 2364, 2506, 2353, 3221, 427]. response-adaptive [1700, 2856, 1265]. responses [717, 3272, 2527, 1778, 3367, 2540, 505, 2599, 2284, 2847, 2376, 74, 501, 1004, 3116, 1090, 1616, 2343, 2106, 3251, 2104, 1439, 1603, 322, 2844, 2066, 372, 3242, 291, 2469, 1440, 2741]. restorations [1931]. restricted [3178, 2048, 485, 659, 550, 1303, 1854]. restriction [678]. restrictions [800]. restrictive [1198]. Restructuring [1826]. result [2282, 1353, 2452]. resulting [1968, 2925]. results [3085, 1578, 3060, 880, 732, 2965, 1521, 727, 1808, 2370, 172]. retaining [1198, 1879]. retardation [37]. retention [2372]. retest [938]. retinopathy [1268, 710, 1317]. Retrospective [1844, 76, 1737, 1186]. revalidation [3094]. revascularization [2796]. Revealing [288]. reverse [1806, 1231]. Review [1530, 1594, 91, 341, 819, 817, 1157, 1460, 1647, 1317, 93, 1565, 454, 1592, 421, 542, 3139, 198, 509, 3222, 2138, 1752, 1786, 1351, 2754, 601, 144, 455, 360, 2824, 526, 885, 1425, 434, 1158, 1364, 497, 1608, 286, 414, 3203, 782, 361, 1751, 143, 200, 127, 183, 1750, 1873, 301, 145, 402, 843, 362, 288, 2501, 3204, 1415, 1151, 2006, 129, 603, 1566, 1550, 90, 1134, 543, 343, 1402, 358, 600, 2004, 1119, 199, 399, 2856, 42, 1461, 420, 1438, 602, 258, 40, 201, 1787, 961, 3311, 303, 1245, 511, 1180, 1067, 1548, 1753, 131]. Review [3312, 1194, 1182, 541, 3057, 211, 1352, 452, 182, 566, 1018, 1195, 1205, 342, 844, 1774, 1802, 3182, 439, 196, 3181, 2399, 1310, 417, 1144, 1181, 436, 92, 438, 2409, 1533, 1311, 2140, 453, 1159, 1593, 1309, 104, 818, 435, 1133, 884, 495, 1142, 780, 130, 300, 302, 1118, 3180, 1531, 524, 299, 1595, 2959, 221, 1671, 821, 1609, 132, 3347, 962, 142, 1872, 1801, 3192, 359, 1169, 781, 357, 494, 1840, 1437, 525, 3056, 1193, 1363, 440, 285, 141, 1143, 1532, 1145, 544, 869, 2519, 872, 419, 1192, 2961, 43, 257, 1546, 2005, 146, 1470, 106]. Review [2109, 1471, 2139, 2502, 2859, 2719, 287, 2295, 870, 3122, 260, 1648, 545, 1516, 185, 2841, 3138, 3231, 1206, 2345, 128, 2034, 298, 400, 418, 820, 1547, 103, 540, 1517, 1065, 1610, 510, 187, 1504, 2718, 284, 1384, 1825, 1698, 147, 105, 2248, 2720, 184, 496, 197, 842, 1066, 186, 433, 210, 1684, 1839, 401, 1549, 1244, 1503, 1416, 871, 41, 94, 2755, 2327, 2174, 205, 3, 2450, 698, 2442, 1804, 28, 3073, 882, 2927, 2260, 1915]. Reviews [2604, 2503, 2857, 1156, 2960, 2858, 437, 259, 792, 784, 786, 647, 344, 2681]. revised [2519, 2109]. Revision [1082, 427, 324]. revisited [3233, 710, 2465, 154, 1227, 616, 3235, 2228, 3047, 3162]. Revisiting [2260, 2450]. reweighting [1587]. rhetoric [305]. rheumatic [3238]. rheumatoid [2316, 326]. rhinitis [548]. rhinometry [117]. rhythms [2195]. Richard [1461, 3182, 1133, 1531, 714]. right [49, 1580, 2286, 2206, 3023, 2071]. right-censored [49, 1580, 3023, 2071]. Risk [414, 252, 417, 3197, 998, 1591, 2344, 330, 159, 3, 2474, 2871, 2876, 1633, 2836, 1508, 2878, 1094, 2166, 2590, 487, 3131, 160, 2, 900, 2357, 3208, 190, 2612, 2756, 1398, 219, 2789, 913, 1403, 1821, 732, 3189, 3291, 1796, 774, 246, 2392, 2489, 2298, 1054, 240, 696, 2495, 1422, 112, 512, 277, 276, 503, 1150, 1817, 2446, 2319, 2994, 1357, 1579, 85, 531, 2455, 1799, 2933, 338, 2323, 1277, 1857, 98, 233, 2047, 1230, 2682, 815, 313, 2555, 1618, 2620, 1401, 3064, 643, 1058, 1128, 1698, 1301, 2866, 172, 1358, 2999, 3163, 1736, 2569, 790, 862, 2476, 1995, 3110]. risk [193, 283, 489, 1158]. Risk-adjusted [3197, 2871, 2876, 1398]. risks [645, 3081, 982, 2606, 3166, 37, 2311, 2766, 2590, 755, 2288, 2426, 644, 3187, 1767, 1719, 2058, 292, 3228, 609, 3217, 2371, 1576, 2223, 1240, 1973, 1519, 2607, 2110, 523, 1047, 2183, 930, 2398, 672, 389, 838, 707, 3082, 237, 3374, 781, 1593]. RJMCMC [234]. RNA [1076, 2317, 1332, 333, 1583]. Robert [3386, 1310, 1850, 1471, 3231, 3302]. Roberto [420, 417]. Robin [358]. Robust [1355, 2791, 3169, 1444, 3130, 1214, 2194, 2130, 2431, 813, 772, 2937, 444, 101, 1858, 630, 1998, 3294, 2509, 2430, 3165, 2620, 2807, 2885, 2418]. Robustness [2324, 1099, 836, 2615, 1063, 350, 2557, 1551]. ROC [2638, 2641, 24, 1682, 2640, 2643, 1707, 1446, 807, 2817, 2637, 2639, 1813, 2801, 1942, 2166, 3284, 3121, 61, 2196, 916, 1155, 3015, 2272, 3321, 60, 2042, 2646, 2749, 1987, 1537, 1932, 1933, 1238, 2636, 3237, 215, 1523, 1945, 1189, 2649, 1745, 3269, 2647, 721, 826, 3006, 2645]. ROC-type [1933]. Rockette [1707]. Roderick [1363]. Roger [2499]. Roland [2819]. Role [892, 1387, 2279, 900, 2772, 467, 1392]. roles [1239, 2581, 109]. ROM [301, 782]. Ron [144, 3204]. Rona [420]. Ronny [3312, 1053]. Rony [2787]. Rosenberger [2856]. rosiglitazone [2907, 2908, 3137, 2539]. Rosner [3202]. Rothman [91]. route [1744]. routine [948, 2704]. row [3235]. Roy [2050]. Royall [90]. Royston [2499, 960, 3138]. Rubin [1533, 1363]. Rudas [1544, 191]. Rudolf [1416]. Rue [2139]. rule [2924, 2116, 1228, 2027, 663]. rules [2894, 2368, 2553, 2783, 570, 2367, 2957, 589, 360]. ruling [1228]. run [2895, 3119]. run-based [3119]. run-in [2895]. rural [2485]. Russek [129]. Russek-Cohen [129]. Russell [2604, 782]. Ryan [1547].

S [1647, 3041, 3135, 1591, 2344, 454, 803, 125, 77, 2907, 143, 1750, 1134, 1402, 1597, 2450, 602, 181, 2640, 1067, 131, 1715, 436, 1533, 297, 1925, 1714, 3310, 1118, 1362, 357, 2770, 2817, 1799, 141, 1532, 419, 2109, 2719, 1516, 437, 1711, 857, 2198, 1644, 400, 1610, 78, 2160, 105, 2720, 496, 186, 1684, 1839, 3137, 94, 2755, 1709, 1182, 1169, 869]. S-L [77, 78]. S-Plus [1750, 1182, 1169, 869].

S.-H [3135]. S.-J [1925]. Sabine [1839]. sacrifice [2278]. Sadasivan [2006]. Saddlepoint [47]. SAEM [2572]. safe [1031, 1390]. Safety [892, 275, 733, 3000, 3264, 3164, 923, 3322, 2201, 2545, 2182, 1528, 2786, 641, 424, 1404]. Safi [1158]. Sage [455, 199, 1461, 511, 1156, 221, 132, 257, 870, 284]. Salkind [870]. Sally [1531, 3192]. Salmon [2907, 3137]. Sam [1220, 1223, 1222, 1219, 1221]. same [329, 662, 2067]. Sample [405, 730, 381, 768, 2095, 3265, 1453, 2094, 2465, 2629, 575, 367, 2918, 3250, 2897, 674, 3028, 1448, 574, 1411, 1882, 3170, 1423, 518, 1062, 1788, 2869, 1819, 2803, 769, 1519, 2607, 35, 802, 3148, 2630, 1688, 3294, 2832, 1265, 2092, 694, 3307, 98, 255, 57, 2530, 377, 184, 3140, 2096, 2663, 1272, 2751, 3105, 1918, 2086, 2738, 1374, 3186, 2311, 931, 1037, 2791, 3259, 926, 369, 2488, 1103, 736, 1183, 1353, 1316, 1477, 2888, 2582, 1982, 2, 1811, 2521, 664, 682, 1041, 2088, 3097, 1589, 1099, 2576, 1858, 404, 3084, 2087, 2848, 274, 2330, 1822, 1835, 1040, 1467, 368]. sample [96, 1246, 1521, 2631, 3264, 573, 447, 1229, 2156, 1735, 945, 1466, 1765, 3257, 2674, 1313, 2536, 1579, 2147, 1605, 1587, 20, 775, 2580, 930, 47, 1059, 3039, 1528, 655, 815, 2159, 1417, 1967, 376, 429, 1818, 1468, 1098, 2403, 1120, 1864, 2401, 2765, 1045, 2071, 2971, 2221, 1410, 1666, 2067, 2485, 685, 2792, 3248, 50, 1184, 2790, 3160, 535, 1404, 3108, 3106, 3122]. Sample-size [1819, 1818]. sampled [2414, 1900, 2294, 2933]. samples [3320, 72, 113, 1668, 2085, 1493, 1805, 1085, 2208, 1362, 1100, 291, 517, 2190]. Sampling [3366, 1081, 2360, 2567, 1846, 1349, 2425, 162, 3260, 2203, 217, 3255, 2348, 1214, 3322, 277, 503, 1371, 1007, 1908, 1445, 1431, 2866, 425, 496]. Samuel [1217, 1218]. San [1065]. Sander [91]. sandwich [775]. Sankarnarayanan [603]. Santa [140]. Santos [401]. sarcoma [1240, 1740]. sarcoma-specific [1240]. Sargent [714]. SARIMA [608]. SARS [2371, 2040, 1855, 1781]. SAS [3139, 2604, 782, 3204, 2409, 2959, 2754, 2409, 3056, 2719]. Sasivimol [3312]. SAS(R) [2754, 3139, 2604]. Sato [2911]. Sato/Matsuyama [2911]. Sauerborn [884]. Sauerbrei [3138]. saved [2921, 683]. Scale [427, 1316, 295, 841, 3190, 2070, 2583, 2535, 1299, 592, 1933, 2322, 118, 662, 1573, 1301, 2275, 1399, 1619, 3006]. scaled [428, 1063]. scales [3060, 1799, 2690, 329, 2999, 790]. Scaling [455, 2833, 821, 1461]. scan [1955, 1186, 2806, 2340, 3177, 2197, 2361, 1922, 2565, 1384]. scattering [2925]. scenarios [2780]. Schabenberger [2604]. Schafer [104, 1374]. schedule [1374, 2985, 2078, 655]. scheduled [136]. schedules [827, 425]. Scheines [1133]. scheme [625]. schemes [2996]. schizophrenia [2111, 1298, 1043]. Schmid [1649]. Schnabel [197]. scholarly [1220]. School [306, 3131, 353, 1667]. school-based [353]. Schwartz [39]. Science [2006, 543, 302, 1118, 1648, 187, 2718, 578, 309, 2161]. Science-Business [2718]. Sciences [455, 221, 545, 1217, 199, 359, 497, 301]. Scientific [603, 131, 106, 437]. Scientists [1134, 155, 156, 1234]. sclerosis [1769, 1998]. scope [3095]. Score [3241, 3053, 1552, 707, 1706, 2060, 2281, 2282, 2443, 2771, 2775, 2910, 3320, 1806, 1207, 1037, 2382, 3024, 2958, 2909, 1858, 2772, 2068, 2773, 2803, 1481, 476, 1495, 2452, 3101, 1260, 3199, 3198, 2619, 2066, 2774, 1401, 2746, 3106, 2023, 2341, 3346, 2762, 2404]. scores [1117, 1545, 2750, 2544, 1116, 2015, 3297, 1239, 548, 3094, 2424, 1669, 3191, 3200, 3124, 2774, 1322, 2221, 2901, 1167]. scoring [65, 2302]. Scotia [76]. Scott [1516]. screen [1629, 3255, 2182, 1303]. screen-detected [1303]. screening [1272, 1615, 2073, 731, 2454, 1830, 12, 792, 190, 2043, 1676, 2397, 2672, 2987, 934, 1687, 980, 1293, 2535, 714, 1089, 1920, 321, 426, 3004, 2811, 1187, 862, 2336, 2165, 1810, 173, 2732, 962]. scrutiny [3209]. seamless [3164, 2577]. search [485, 2188, 3064, 1133]. searching [791, 2367]. Seasonal [2215, 119, 2873, 2460, 1784]. seasonality [1541, 2017, 1429]. Sebastiani [2036]. second [3157, 214, 98, 2022, 1426, 1427, 2075]. second-line [1426, 1427]. second-order [3157]. Secondary [1993, 79, 3254, 2190, 3266]. secretion [681]. section [2076]. sectional [430, 3293, 2100, 503, 1007, 1171, 73, 371, 2081]. see [1786]. segmentation [808, 1370]. SEIR [119]. seizures [2880]. selected [2865, 1928, 472, 3241, 3103]. Selecting [953, 1293, 1237, 3114]. Selection [434, 708, 749, 2618, 3236, 3272, 1121, 2948, 1806, 3085, 1919, 1862, 2293, 2768, 3270, 2731, 789, 2899, 2566, 3195, 2117, 2085, 1677, 1405, 194, 2445, 983, 1704, 2302, 3164, 2299, 2739, 2495, 2496, 3044, 3294, 407, 348, 2024, 2976, 1878, 312, 1230, 1015, 1208, 2896, 2786, 837, 491, 2675, 2632, 2956, 1021, 3067, 971, 797, 110, 3064, 618, 2265, 1538, 1748, 3029, 1705, 2862, 826]. Selective [976, 318, 934]. selectivity [2516]. Self [3070, 1245, 1928, 2092, 2900, 2031]. self-controlled [2092, 2031]. Self-Instruction [1245]. self-reported [2900]. self-selected [1928]. Selvin [1143]. semen [1076, 2881]. Semi [1833, 2612, 1612, 2458, 734, 2192, 3006, 2379, 3022, 2621, 2320, 2288, 2196, 8, 2608, 2058, 2315, 558, 331, 1694, 925, 3167, 3055, 1047, 2377, 2941, 624, 952, 3309, 1921, 1851]. semi-Bayes [558]. semi-competing [2288, 2058]. semi-continuous [3167]. semi-Markov [3022, 2621, 2608]. Semi-parametric [1833, 2612, 1612, 2458, 734, 2192, 3006, 2379, 2320, 2196, 8, 2315, 331, 1694, 925, 3055, 1047, 2377, 2941, 624, 952, 3309, 1921, 1851]. seminar [728]. Semiparametric [3168, 2058, 2794, 3269, 3102, 2836, 2801, 2973, 3291, 3256, 52, 3001, 2661, 3262, 2863, 2429, 2408, 2506, 2447]. Sen [435]. Senegal [1011]. Senn [1696, 3181, 2676, 1698]. Sense [1245, 1599, 305]. sensitive [1813, 2531, 2069]. sensitivities [2167]. Sensitivity [1331, 3210, 1481, 1028, 1354, 2326, 2711, 2710, 667, 1250, 1830, 2747, 1968, 179, 2134, 968, 921, 3142, 2921, 1173, 1466, 1755, 380, 2395, 949, 409, 3154, 46, 2748, 1627, 2778, 479, 1553, 1599, 2165, 2852, 3290, 1231, 1619, 2459]. sensor [2509]. sentinel [2468]. separated [2218]. Separating [1480]. Separation [1163, 855, 1330]. septal [432]. sequences [1079, 2726, 557]. Sequential [1743, 2740, 541, 2124, 2574, 2542, 1021, 1841, 3033, 327, 1570, 17, 1652, 2225, 289, 323, 1487, 3340, 2329, 2586, 485, 2576, 1966, 1777, 1229, 1983, 822, 1228, 3322, 35, 1663, 2742, 1023, 26, 2147, 2611, 1265, 2368, 683, 1878, 1380, 1967, 2066, 1006, 291, 705, 3067, 1445, 397, 2561, 2248, 513, 3029, 344, 2165, 2333, 2353, 3194, 2645]. Sequentially [2769, 2894, 3143]. serial [852, 1781, 3055, 2253, 2843]. Series [455, 221, 781, 2529, 728, 2873, 199, 878, 2029, 2216, 1899, 276, 1278, 2596, 2092, 2089, 1667, 2215, 2031, 3275]. seroconversion [177, 396]. serological [3293, 1457]. service [216]. Services [399, 2358, 3099]. session [156]. sessions [2662]. set [3359, 1846, 2698, 2634, 1431, 3253, 339]. Sets [1398, 1951, 470, 2864, 2915, 767, 2589, 3044, 1637, 110, 1189, 172, 721]. setting [3233, 3202, 1387, 1038, 1010, 720, 2879, 556]. Seung [538]. Seung-Ho [538]. seven [1850, 9]. Seventh [1608]. several [1179, 2513, 2311, 413, 1444, 1411, 1966, 626, 112, 1436, 1343, 3267, 3115, 2807, 1397]. Severe [2371]. Severity [427, 2111, 1992, 548]. sexual [1491, 2252]. sexually [3309]. Seymour [510]. SF-36 [2321]. Sham [145]. Shao [2002, 1645, 3122]. Shape [260, 2407, 2767]. shaped [1954]. shapes [718, 2195]. Shared [2270, 1912, 1288, 3293, 2239, 2941]. sharing [702]. Sharma [1461]. Sharon [2452]. Sharon-Lise [2452]. Shayne [210]. Shein [2002, 3203, 1645, 201, 495, 2960, 287, 3122]. Shein-Chung [2002, 3203, 1645, 201, 495, 2960, 287, 3122]. Sheldon [1206]. shelf [448]. shelf-life [448]. Sheskin [1503]. Shiboski [2140]. shifts [577]. short [881, 2321, 1037, 2695, 2520, 3226, 1140, 1470, 734]. short-cuts [2520]. short-interval [2695]. short-term [1037, 3226, 1140, 1470, 734]. Shoukri [2959, 2034]. Should [1386, 3276, 3201, 1703, 2687, 3349, 660, 788, 2862, 2681, 3380]. show [2378, 3077]. shrinkage [1360, 1468, 2390]. Shun [1715, 1714]. Shuster [3105, 2907, 2033, 3137]. sib [702]. sibling [3241, 1560]. Sidák [435]. sided [1497, 2464, 2591, 1266, 9, 1850]. Sidik [1168]. SIDP [1260]. SIDS [1429]. sieve [350]. Sigbert [285]. sigmoidally [2195]. signal [3233, 3271, 1627]. Significance [1766, 174, 3190, 2514, 3387, 944, 2322, 1469]. significant [1146, 2973, 2707, 743, 742, 725, 1231]. silent [1782, 2969]. Silva [401]. sim.2538 [2450]. sim.2600 [2344]. sim.2618 [2452]. sim.2655 [2476, 2489]. sim.2822 [2476]. sim.2832 [2751]. sim.2855 [2653]. sim.2929 [2638, 2641, 2640, 2643, 2642, 2637, 2639]. sim.2974 [2770]. sim.3034 [2839]. sim.3069 [2820]. sim.3122 [3070]. sim.3722 [3379]. SIMEX [3353]. similar [1444]. similarity [271, 2654]. Simon [2245, 303, 2503, 525, 2851, 3066]. Simonoff [1684]. Simple [1639, 1673, 1246, 103, 3124, 1168, 53, 2945, 70, 1651, 2550, 296, 176, 2033, 2414, 1473, 555, 2455, 1469, 1976, 1829, 89, 1775, 920, 429, 2342, 1410, 3290, 450, 3266, 2032]. simplify [295, 841]. Simplifying [967]. simulate [1720, 2052]. simulated [1927]. Simulating [3166]. Simulation [93, 351, 967, 2703, 2224, 2499, 2942, 3009, 2898, 880, 1626, 3045, 1773, 1261, 1342, 1913, 2227, 1702, 2615, 502, 2858, 1627, 1504, 1890, 1573, 653, 2467, 2191, 1200]. simulation-based [2942, 2898, 1200]. simulations [336, 2780]. Simultaneous [515, 800, 409, 3325, 1406, 1982, 2422, 2986, 1553, 2510]. simultaneously [1776, 443]. since [177]. Singer [497, 1595]. Single [1791, 405, 2513, 2178, 2116, 585, 1850, 1172, 9, 1083, 291, 662]. single-race [1083]. single-stage [405]. single-unit [2513]. Sinha [1517]. sinusoidal [2195]. Siqueira [3041]. situations [521]. Siu [10]. six [293]. six-minute [293]. Size [890, 405, 1943, 730, 470, 2086, 2738, 1374, 381, 768, 2095, 931, 2791, 3259, 926, 369, 3265, 1453, 1353, 561, 296, 2094, 3161, 2465, 2629, 575, 2582, 1982, 367, 2, 3250, 1811, 2521, 664, 682, 1041, 2088, 2897, 3097, 674, 3028, 1448, 574, 1589, 1099, 2576, 1411, 404, 1882, 3170, 2848, 2330, 1835, 1040, 518, 1062, 1788, 2869, 1467, 368, 96, 1246, 1521, 2631, 3264, 1819, 202, 3108, 447, 2156, 2803, 769, 1232, 3261, 1519, 2607, 945, 1342, 1466, 1765, 35, 802, 2674, 3148, 2630, 1688, 757, 2147, 555, 2832, 1265, 2487, 1940, 1941, 2220, 694, 3307, 2373, 2580]. size [930, 1059, 98, 135, 1528, 3122, 2745, 255, 655, 57, 815, 2159, 2198, 2530, 377, 1967, 376, 1196, 429, 1818, 1468, 1098, 2403, 1120, 3106, 2401, 2765, 1045, 2221, 1410, 3140, 1666, 1602, 2792, 2671, 3248, 50, 1946, 1184, 2790, 3160, 2096, 2663, 535, 2441, 1404]. size-related [3161]. sized [1951, 1637]. sizes [3265, 1966, 1423, 368, 3294, 2092, 2613, 1417, 1147, 2410]. Sizing [2393]. skew [2891, 2317, 2730, 1501]. skew-normal [2317, 2730]. skewed [3377, 3060, 1057]. skewing [1778]. skill [3156]. Skrondal [2857]. sliding [219]. slopes [1242, 1062]. Small [1272, 1858, 2920, 775, 2751, 1951, 1918, 3186, 2311, 2488, 113, 736, 1183, 1477, 2918, 2164, 3084, 1957, 274, 1766, 2536, 1095, 504, 2208, 1665, 1100, 1096, 2613, 47, 1637, 233, 1470, 1528, 1823, 2713, 1417, 110, 1864, 2418, 1427, 237, 365, 1404]. small- [1951, 1637]. small-area [1095]. Small-sample [1858, 775, 1183, 3084, 1864]. small-study [2164]. smaller [609, 1715, 1714, 1716, 1713]. Smallman [259]. Smallman-Raynor [259]. Smeeton [1425, 3311]. Smith [1608, 302, 298, 1244, 2839]. Smoking [37, 2193, 3273, 219, 913, 2213, 468, 172, 2310, 141]. Smooth [2697, 250, 1501, 2408, 3051, 3121, 2463, 853, 1458, 3016]. smoothed [522, 1770, 410, 3376, 236]. Smoothing [1369, 564, 2287, 1308, 2494, 1849, 2947, 2545, 118, 2244, 2796, 3355, 2238, 2247, 650, 981, 826, 3049]. smoothly [505]. SNP [2127, 3358]. SNP-expression [3358]. Social [455, 221, 2915, 199]. Society [1568, 1876, 2212, 2570]. socioeconomic [1738]. soft [756]. softcover [1872, 2718, 2718]. Software [1635, 2442, 1516, 2841]. sojourn [1830, 3255]. solution [855, 2517]. solutions [2297, 2278, 3136]. Some [323, 1223, 520, 732, 273, 988, 1123, 583, 2200, 3296, 314, 59, 2241, 1432, 2576, 2307, 1579, 555, 537, 2118, 2430, 1013, 2581, 2577, 2355, 604, 2549, 2305]. sometimes [462]. Song [1206]. Sons [3222, 3181, 1363, 2859, 3231]. Soong [1925]. Sophia [1751, 844, 2857, 2755]. sore [323]. source [1493, 3207, 2304]. sources [1079, 2931, 247, 786, 1721, 2682, 1512, 1254, 707, 2365, 3299]. South [2262, 3131]. Southern [2915, 1667]. Space [232, 1738, 551, 827, 247, 246, 2339, 879, 2507, 2565, 1963]. Space-time [232, 1738, 551, 247, 246, 2565, 1963]. Spacing [164]. Spain [1924]. Spanish [2926]. Spanish-speaking [2926]. sparse [2292, 966, 1949, 1499, 2280, 1377]. Spatial [1073, 1144, 240, 696, 2936, 3351, 100, 1641, 1346, 2599, 1191, 631, 1955, 228, 599, 194, 2998, 2340, 3177, 1956, 2197, 231, 241, 1991, 1095, 1442, 1959, 2244, 1960, 1504, 236, 34, 981, 2510, 3240, 3049, 2889, 1054, 1364, 1142]. Spatially [129, 1079, 2779, 1213, 2555]. Spatio [1782, 695, 206, 2834, 226, 227]. Spatio-temporal [1782, 695, 206, 2834, 226]. spatio-temporally [227]. spatiotemporal [2915, 2796]. speaking [2926]. Special [2504, 1939, 1914, 1979, 1811, 2520, 488]. specific [559, 44, 529, 1127, 2606, 651, 2990, 3252, 1830, 175, 2723, 1673, 2261, 1790, 922, 1240, 1925, 26, 960, 431, 1479, 1434, 98, 2240, 278, 1783, 2696, 1890, 394, 1267, 1090]. specification [1969, 2890, 3363]. specificity [2747, 179, 2921, 1466, 949, 409, 3154, 2748, 2167, 1627, 479, 1553, 2165, 1619, 2459]. specified [2557, 3298, 2577, 3066]. specimens [326]. spectra [808]. spectroscopy [983]. Spectrum [1279, 1728]. spells [3022]. Spencer [2718]. spending [291]. Sperm [2575]. Spiegelhalter [1752, 3379, 1799]. spike [2514]. spine [2767, 1141]. spirits [998]. spirometry [3161]. Spirtes [1133]. Spline [1764, 1382, 1177, 1849, 2545, 1660, 1140, 118]. Spline-based [1764]. splines [3048, 2432, 3020, 564, 1935, 659, 348, 2796, 3355, 1995]. split [2471, 1882, 1822, 3349]. split-cluster [1882]. split-mouth [2471, 3349]. sponsor [1386, 1392]. spontaneous [825]. spread [1866, 3361, 2512]. Sprenger [2770]. Sprent [1425, 3311]. Springer [1157, 1647, 1565, 2138, 526, 434, 497, 1750, 301, 2501, 129, 543, 40, 1548, 1182, 1352, 566, 1195, 196, 1310, 1181, 2140, 1156, 525, 285, 872, 419, 1546, 1516, 1517, 510, 2718, 1384, 1684, 41]. Springer-Verlag [1157, 1647, 526, 434, 301, 129, 543, 1352, 1195, 872, 41]. SPSS [1156, 1839, 1193]. spurious [1403]. square [534]. squared [2488, 1812, 2569, 2751]. squares [2067, 3110]. Sr [2640]. SROC [763]. SSVEP [2510]. Stability [2334, 1015, 2795, 59, 2422, 2763]. stably [2649]. staff [457, 728, 2955]. Stage [3252, 3215, 405, 2540, 2158, 2157, 2411, 3065, 3366, 3282, 2582, 3000, 3250, 605, 1314, 1524, 1776, 2851, 3034, 1789, 949, 1027, 2271, 2712, 672, 2159, 1485, 1967, 861, 3279, 532, 823, 2160, 424, 1740, 2561, 2628, 734, 2683, 3158, 2930, 3066, 2929, 2729]. Stage-specific [3252]. staging [205, 2554]. stains [947]. Stallard [1784]. Standard [2151, 3155, 876, 125, 530, 1561, 1319, 10, 766, 2833, 2414, 1676, 520, 3321, 2097, 6, 1266, 2997, 1933, 1174, 1818, 3064, 3358]. Standardization [688, 407, 2485, 1056]. standardized [639, 111, 2428, 2296]. standards [1919, 460]. Stangl [871]. Stanley [819, 182]. Staphylococcus [1369]. Staquet [144]. Start [2916]. Stata [2857, 1751, 844, 2857, 2755]. State [167, 2507, 66, 1631, 87, 328, 2942, 3195, 1974, 3254, 1559, 508, 854, 2398, 1114, 2045, 2270, 2259, 1896, 3378, 3017]. State-space [2507]. statement [375]. states [2111, 461, 2723, 955, 2456, 650, 2485, 3378, 1082, 2965]. Statictics [3024]. statin [2060]. Station [2857]. stationarity [2030]. stationary [810, 207]. Stationery [818]. statistic [563, 3061, 1939, 789, 1955, 3063, 2838, 1186, 2737, 2806, 534, 2340, 3171, 2197, 2339, 531, 1940, 1941]. Statistical [1868, 1283, 2346, 349, 1351, 2431, 950, 2504, 1425, 2570, 1636, 183, 288, 2967, 1415, 733, 2914, 463, 1624, 2576, 629, 2537, 2788, 1205, 1677, 844, 1310, 2400, 1478, 436, 1701, 451, 2352, 3043, 2347, 1617, 62, 142, 1622, 2487, 466, 2475, 1999, 1420, 243, 974, 872, 2622, 2291, 1959, 570, 584, 1729, 1713, 68, 2912, 785, 335, 2103, 284, 2721, 147, 3092, 1915, 873, 3088, 1692, 2648, 1530, 559, 2481, 157, 866, 2824, 403, 937, 108, 1980, 2094, 1751, 622, 2471, 2123, 578, 174, 3285, 1223, 1787, 3311, 1548, 1753, 3007, 1656, 196, 319, 273, 1925, 1294, 1123, 583]. statistical [3271, 1913, 879, 1463, 2424, 2846, 2354, 310, 1005, 2841, 3231, 860, 1623, 3358, 1839, 1962, 1500, 1934, 1503, 2755, 2404, 2327, 1715, 1714, 569, 1716, 975, 2787, 1647, 454, 1750, 402, 362, 2501, 90, 543, 1180, 1067, 211, 1195, 3182, 3181, 1144, 2503, 130, 1118, 1801, 1363, 141, 1192, 2109, 1471, 2719, 2295, 260, 1504, 2718, 186]. Statistician [1385, 2262, 1387, 1388, 578, 649, 1391, 1392]. statisticians [115, 1386, 3327, 3182, 3349, 3328, 3329, 1393, 3332, 3326, 3331, 3330, 1873]. Statistics [314, 2570, 1608, 361, 129, 2064, 211, 436, 818, 1156, 1563, 1169, 781, 357, 141, 869, 146, 128, 842, 306, 568, 740, 2174, 745, 1695, 1786, 2754, 2824, 747, 2224, 1580, 985, 2499, 51, 2006, 2, 1412, 5, 748, 3205, 1225, 1182, 1774, 3103, 1997, 297, 2869, 3177, 1777, 2361, 302, 1609, 3192, 960, 739, 2012, 1922, 2362, 278, 815, 3296, 2565, 3246, 593, 2401, 2265, 1549, 391, 1979, 538, 3041, 2751, 3105, 3135, 1117, 1168, 1545, 1649, 1591, 2002, 1951, 2000, 2440, 2327, 2245, 2344, 2379, 1424, 2910, 355, 1374, 803, 2052, 3202, 2820]. Statistics [125, 13, 39, 77, 2907, 2735, 2665, 2439, 414, 813, 1561, 1696, 3386, 2638, 1784, 841, 2641, 1939, 1914, 2499, 3384, 10, 3070, 1682, 2450, 3190, 2050, 3379, 1645, 2033, 181, 1337, 1589, 2909, 2640, 2643, 387, 2437, 2772, 2773, 2134, 1715, 297, 1850, 1925, 3380, 386, 1714, 2653, 3108, 3310, 2642, 2676, 2489, 1054, 1056, 3385, 867, 11, 714, 2558, 3136, 1016, 2452, 757, 2819, 958, 1362, 960, 1169, 1728, 3383, 2770, 2817, 416, 2855, 537, 1725, 2559, 2183, 1851, 1799, 2186, 2637, 2734, 1669, 815, 1711, 857, 2198, 1716, 1644, 2032, 3381, 2911, 3219, 2787, 78, 2036, 2774]. Statistics [354, 2160, 3106, 29, 1053, 397, 3302, 1494, 3137, 3288, 2639, 759, 1450, 2476, 3382, 2839, 3221, 283, 1709, 1119, 285, 870, 820, 1384, 1565, 509, 2399, 299, 2961, 1648, 2231, 183, 1270, 145, 1438, 439, 453, 1840, 870, 2163, 2345, 510, 210]. Statistique [40]. status [3344, 2816, 1444, 2176, 849, 3293, 475, 2207, 2710, 411, 2170, 3248, 1692]. Staudte [3231]. stay [2725, 951]. STD [2206, 1128, 242]. Steen [1145]. Stefan [1205]. Stein [2754, 183, 2409]. Steinberg [2787]. stem [2606, 2116]. stem-cell [2606]. Stenning [1531]. stenoses [490]. stenosis [1125]. stenting [994]. step [1453]. Stephan [3231]. Stephanie [342, 2819]. Stephen [1696, 1566, 1550, 3181, 2503, 2140, 2676, 1156, 962, 1610, 1698]. Stephens [1531]. STEPP [3186]. steps [207]. stepwise [1031, 1538]. Stereology [1873]. stereotype [3323]. Stern [1533]. Sterne [1648, 2839]. steroid [681]. Steve [1143]. Steven [3070, 1182, 299, 869, 146]. Stewart [1531]. Stiller [437]. stimulated [753, 754]. stimulation [3005]. Stochastic [93, 1319, 204, 1077, 2858, 3064, 1020]. Stochastically [2331]. Stopping [2783, 1020, 289, 2116, 2330, 1804, 2156, 1228, 2541, 2353, 589]. straightforward [409]. strain [292, 350]. strata [2200]. strata-matched [2200]. Strategic [310]. Strategies [114, 1157, 3059, 564, 3089, 1691, 1518, 2740, 2284, 1146, 699, 2520, 353, 2415, 2352, 3043, 1699, 2972, 2128, 89, 2516, 3115, 2561, 513, 2831, 326]. strategy [3096, 3085, 2923, 1715, 1714, 581, 2387, 188, 1485, 1716, 1713]. stratifed [72]. Stratification [1158, 1495, 2619, 935, 3176, 1676, 2415, 1563, 2136, 1816, 1022]. stratified [1601, 3125, 969, 1634, 1936, 1342, 1789, 1371, 660, 85, 2989, 370, 1129, 2096, 2727]. stratifying [1342]. stratum [2990]. stratum-specific [2990]. Streamlined [2661]. streams [2362]. Strength [1475, 2544]. Streptococcus [1866]. stressful [828]. stroke [1382, 3238, 1049, 512, 1563, 3094, 1479, 2603, 390, 1022]. stroke-specific [1479]. Strothotte [1181]. Stroup [2604]. Structural [1664, 1426, 553, 1331, 1035, 2599, 1188, 796, 3338, 1202, 2794, 2890, 760, 2396, 2844, 3262]. structure [2598, 2276, 2175, 3146, 180, 1958, 1299, 701, 2510]. Structured [2082, 2402, 692, 1344]. Structures [285, 1339, 248, 3267, 3200, 1338]. Stuart [299]. student [745, 3258]. students [2537, 2788, 2787]. studie [3190, 3302]. Studies [414, 540, 259, 730, 139, 1489, 948, 1601, 982, 1977, 157, 3, 1771, 2498, 859, 702, 1094, 2224, 1220, 926, 3089, 3277, 2907, 2411, 1183, 1249, 1844, 3118, 3270, 767, 1784, 1525, 1830, 290, 3280, 2049, 2499, 2926, 1443, 71, 1284, 1295, 272, 627, 2847, 2781, 3250, 1767, 1257, 674, 177, 201, 953, 2898, 318, 2448, 2693, 2772, 1182, 2397, 1107, 3007, 913, 3084, 1812, 2087, 1899, 1383, 732, 307, 164, 1715, 1835, 3090, 534, 2349, 954, 2124, 1819, 774, 1973, 1714, 1560, 475, 495, 3093, 1298, 23, 1286, 1575, 945, 2943, 3349, 626, 1540, 1466, 112, 1150, 1817, 1007, 2207, 1480]. studies [2707, 1306, 2996, 1023, 2200, 2395, 949, 958, 2384, 1913, 1887, 150, 1165, 2092, 1089, 1587, 2012, 1248, 1373, 694, 3059, 3091, 2373, 1027, 2322, 3073, 1857, 321, 2089, 2387, 1211, 3062, 2128, 1343, 2232, 3201, 2014, 2686, 1984, 2530, 2701, 1417, 1716, 1713, 861, 2539, 2908, 2370, 2811, 3219, 1910, 73, 1120, 1864, 2103, 1685, 1763, 1815, 1189, 653, 2215, 618, 607, 1312, 3137, 2647, 1250, 1930, 824, 1397, 1406, 1946, 2845, 2790, 268, 2827, 3102, 1653, 2129, 2995, 1289, 2101, 69, 589, 327, 489, 1570, 3203, 2248]. Study [341, 420, 3103, 771, 1502, 3038, 3323, 967, 1317, 1643, 2060, 2281, 2282, 852, 1970, 205, 1554, 3085, 710, 2703, 3186, 975, 1003, 2766, 1539, 2924, 1869, 2880, 516, 49, 772, 2116, 3273, 777, 351, 3131, 2815, 3161, 3238, 1191, 1456, 1151, 1443, 992, 1428, 174, 2004, 3058, 2420, 2897, 3009, 1011, 617, 1790, 2756, 2164, 1883, 1368, 2204, 1321, 353, 625, 2680, 2514, 3304, 3045, 96, 1900, 3108, 336, 2583, 2803, 3035, 923, 1773, 1261, 1342, 1293, 1654, 1616, 911, 166, 2422, 2883, 3343, 1495, 548, 2201, 1480, 2452, 1299, 2707, 2950, 2326, 2227, 3094, 2615, 502, 3154, 2551, 974]. study [1614, 98, 233, 1470, 1937, 170, 2270, 2309, 2022, 1115, 2258, 2555, 2595, 340, 1717, 1511, 785, 1468, 2620, 1098, 2351, 2746, 3106, 1128, 1890, 1573, 1734, 653, 2301, 2090, 2433, 2467, 2193, 2057, 1399, 952, 2917, 2506, 493, 489, 159, 3022, 1667, 1401]. study/internal [340]. Studying [524, 2125, 901, 2729]. Stuttgart [197]. style [1096]. styrene [1316]. sub [1428, 1614, 1937]. sub-groups [1428]. sub-study [1614, 1937]. subclasses [1000]. subclinical [44, 1256]. subdistribution [2344, 2606, 2223, 2298, 2047]. subfertility [3030]. Subgroup [902, 318, 770, 2480]. subgroups [3096, 2918, 2348, 1995]. Subhash [1461]. Subject [1368, 3272, 1044, 3131, 1768, 1673, 272, 1444, 617, 1790, 2568, 1333, 659, 1357, 749, 2534, 3281, 2318, 3237, 383, 1783, 2561, 2054, 2649, 607, 408]. subject-by-formulation [272]. subject-specific [1673, 1790]. Subjective [3147]. Subjectivity [1134]. subjects [2281, 2389, 2945, 2899, 617, 609, 2319, 2014, 1136, 2067]. submission [1270]. subpopulation [701]. subpopulations [3366]. subset [944, 901]. subsets [249, 2299, 3018]. substance [1814]. substudies [1929]. subtypes [165, 1025]. subunit [1126]. subunits [1926]. subverting [380]. success [2178, 3306]. successful [2226]. successive [1617]. suggested [2911]. suicidality [2756]. suicide [1305, 2757, 475]. Sullivan [1787]. sum [1121, 2423, 1442]. Summarizing [1672, 2988, 133, 178, 954]. summary [2944, 3284, 789, 1412, 5, 56, 297, 1777, 2817, 156, 89, 483, 763, 1745, 2647, 391]. sums [3097]. Sung [1610]. Sungho [3386]. superficial [1343]. superiority [3041, 2313, 1776, 1983, 2668, 1485, 2655, 513]. supplemental [2693, 1007]. Supplementary [1663]. suppression [2509]. Sure [379]. surfaces [2789, 3269]. surgery [1179, 515]. surgical [2266, 1320, 2595]. Surrey [510]. surrogacy [2610]. Surrogate [1914, 1916, 2327, 2575, 1396, 1236, 729, 3384, 3070, 1929, 2975, 3150, 615, 1465, 606, 1917, 1915]. surrogates [1597, 1596, 1598, 1717]. surveillance [45, 2864, 2037, 2038, 2872, 3130, 1626, 2965, 1956, 2921, 2361, 3322, 2545, 1922, 2846, 3069, 1627, 2362, 2035, 2039, 2713, 3214, 2036, 3049]. Survey [1631, 2349, 1153, 2364, 2454, 2984, 1794, 1493, 573, 300, 465, 2351, 1457]. surveys [2095, 3366, 430, 1086, 1923, 214, 1081, 2348, 3257, 3211, 2455, 2783, 3038, 2359, 2358, 438, 184, 199]. Survival [1157, 526, 877, 3336, 603, 182, 566, 2168, 21, 2887, 3350, 2385, 1244, 2257, 1638, 3081, 1892, 2944, 2030, 1498, 2086, 1374, 1613, 1720, 2052, 1786, 3166, 3085, 3186, 1651, 2169, 859, 1037, 2287, 279, 1242, 49, 2386, 87, 1580, 2286, 1072, 1235, 2886, 2538, 2939, 175, 2049, 2610, 2582, 1274, 2779, 3016, 597, 3210, 2276, 598, 3375, 3285, 928, 1412, 676, 1413, 1257, 1215, 515, 1177, 1888, 3075, 1300, 1359, 1411, 1093, 2206, 849, 922, 3084, 634, 64, 610, 1276, 2557, 3129, 1433, 3226, 3189, 492, 1926, 865, 1827, 2869, 3255, 2140, 692, 956, 2278, 282, 2548]. survival [2614, 822, 1113, 1510, 1889, 1285, 1509, 718, 506, 1513, 955, 724, 2541, 2028, 3034, 3167, 1780, 1739, 1016, 1454, 2200, 1605, 2704, 773, 348, 1356, 683, 2616, 3023, 1479, 942, 250, 1434, 338, 2416, 764, 478, 1459, 1808, 2780, 2387, 312, 1335, 443, 2761, 834, 1135, 1330, 2711, 2728, 2244, 655, 1421, 765, 2577, 57, 3197, 491, 1542, 1517, 1891, 449, 1890, 1345, 734, 3173, 722, 680, 997, 2671, 3248, 2132, 1931, 2486, 1184, 1828, 957, 1407, 2115, 2715, 3194, 2889, 1972, 324, 1809, 2000, 1352, 1610]. survival/sacrifice [2278]. survive [2262]. surviving [2954, 491]. survivor [1359, 3035, 486, 3346]. survivors [1037, 2493, 1542, 2715]. Susan [1402, 420, 1156]. Susanne [1594]. susceptibility [3089, 1178, 2893, 2974]. susceptible [1215]. suspected [384]. Sutton [1206]. Sw [884]. Sweden [3130]. switch [859, 2665, 2664]. switchability [267]. switching [2949, 1729, 2018]. Sylvan [1384]. Sylvie [1750]. symmetric [295, 841, 2522, 2195]. symmetrical [339]. Sympercents [295, 841]. Symposium [2346, 2347, 2912, 1217]. symptom [2111]. symptomatology [1002]. synchronous [2514]. Syndrome [2371, 679, 2596, 862]. syndromic [2872]. synergism [2703, 1120]. synergism/antagonism [2703]. Synthesis [3218, 1197, 2049, 321, 2235, 2214]. Synthetic [616]. system [2723, 666, 300, 782]. systematic [1078, 792, 784, 786, 2927, 647, 2681]. Systematically [3184]. systems [205, 65, 1626, 242, 1447, 825, 884]. systolic [1815].

T [125, 1364, 77, 841, 129, 1589, 3385, 2452, 757, 141, 2961, 2841, 2911, 94, 666, 1090, 2978, 512]. T-cell [666, 2978]. T-lymphocyte [1090]. t-PA [512]. T1DM [1820]. table [3135, 3053, 1007, 2072, 3042]. Tableman [1610]. tables [405, 2751, 2172, 2093, 2488, 1630, 2873, 778, 395, 2804, 276, 1266, 2536, 3179, 193, 283]. tabular [1635]. tabulated [2470]. TACT [989]. tail [1940]. tail-area-based [1940]. tailored [1360]. Taiwan [552, 934]. taken [1853]. Takezawa [2859]. taking [1175]. tale [2374]. talk [738]. Tamas [1544]. Tango [2186]. Tanur [1134]. target [1687, 1008]. Targeted [3098, 3104, 3206, 1757, 2308, 1611]. task [812]. Tavaré [525]. Taxometric [257]. taxonomy [2538]. Teaching [747, 748, 743, 742, 740, 745, 770, 739]. Technical [2784]. technique [854]. Techniques [542, 1352, 1601, 1918, 709, 176, 2494, 2898, 972, 2619, 650, 1253]. technologies [1833, 1877, 2074]. technology [577]. telephone [1086, 2348]. tell [462, 2503]. templates [244]. Temple [2198]. Temporal [1922, 1782, 1738, 3333, 2029, 219, 1297, 695, 2834, 226, 2354, 2517, 206]. temporally [1079, 227]. ten [767]. tendency [217]. tenofovir [2318]. tensor [3020]. Tenth [1082, 2346, 2347]. term [1037, 2166, 900, 928, 1097, 515, 2266, 2587, 849, 898, 3226, 996, 3035, 2493, 345, 2416, 1808, 1140, 1470, 1542, 734, 888, 952, 486, 2715, 3346]. terminal [2608, 2259]. termination [1663, 3154]. Terms [1415]. Terrin [1649]. terrorist [1624]. Terry [526]. Test [1291, 2, 1394, 1329, 810, 192, 18, 538, 3105, 1868, 1951, 1813, 3212, 3289, 2473, 1207, 3143, 1037, 2800, 3241, 1204, 3118, 2314, 1700, 2888, 1430, 154, 2589, 3209, 2179, 2544, 1811, 2958, 599, 1279, 1448, 3205, 1858, 387, 2164, 2325, 2087, 609, 3321, 2951, 1433, 2290, 2381, 2861, 3338, 2514, 2097, 213, 2965, 2278, 202, 1173, 1776, 1557, 2902, 2803, 626, 1239, 2749, 2903, 3156, 2884, 1199, 1266, 2536, 757, 1165, 1678, 1381, 1728, 944, 20, 138, 987, 293, 46, 1637, 98, 1469, 2593, 3062, 938, 2685, 591, 2778, 2280, 2542, 2066, 1445, 3358, 218, 3106, 34, 2401, 2782, 763]. test [3269, 513, 1057, 2165, 2023, 2341, 1828, 1426, 2337, 2830, 3160, 1400, 2096, 2663, 3346, 1619, 2404, 815]. test-retest [938]. testicular [3256, 1318, 663]. Testing [1861, 2513, 2836, 3277, 2286, 256, 2931, 631, 51, 410, 3360, 1719, 2958, 1899, 331, 2717, 1313, 1587, 1659, 502, 1878, 723, 3042, 2904, 3096, 3041, 1762, 3265, 1453, 3300, 3357, 1580, 716, 978, 676, 1413, 2520, 3338, 988, 1793, 2413, 1827, 2769, 1766, 1983, 1765, 1949, 2208, 442, 2534, 3387, 590, 930, 2941, 2972, 2084, 2128, 991, 2655, 2480, 3115, 1445, 1120, 34, 1435, 1722, 1841, 989, 264, 517, 2692, 1653, 493, 3003, 2183, 1205]. Tests [1613, 2567, 113, 782, 1767, 395, 2804, 687, 1987, 276, 977, 2727, 2977, 3155, 18, 730, 1272, 1615, 3178, 2751, 1126, 137, 2138, 731, 70, 3153, 279, 2488, 2805, 813, 1316, 3300, 351, 1980, 290, 444, 3128, 1982, 2402, 969, 1188, 1703, 1444, 1679, 2935, 2691, 766, 1787, 1337, 801, 1676, 387, 2772, 2397, 501, 520, 140, 1582, 1926, 3045, 3224, 1559, 48, 386, 1172, 2739, 966, 1229, 1285, 1481, 1936, 506, 786, 812, 1552, 1266, 3179, 864, 550, 2837, 1378, 1579, 2950, 1605, 2487, 1154, 2012, 3225, 1725, 409, 2528, 590, 1723, 743, 1408, 1702, 2318, 560, 677, 1348, 993]. tests [2291, 2572, 860, 2668, 3258, 991, 1960, 762, 291, 3124, 742, 2986, 1961, 1058, 1553, 3099, 707, 1053, 1962, 2408, 3269, 2067, 2647, 391, 1988, 3382, 1810, 1167, 2129, 2130, 173, 3006, 151, 1570, 3135, 435, 284]. tetrachlorodibenzo [709]. tetrad [3212, 3289]. Thailand [195]. their [1079, 3277, 2393, 608, 109, 2057, 604]. them [906]. Theo [884]. Theodore [127]. Theoretic [434, 3340, 2026, 404, 2147, 2578]. Theory [568, 780, 106, 2139, 1355, 198, 200, 2856, 698, 2442, 1214, 675, 299, 169, 3134, 1825, 3133, 365, 427, 435, 184]. Therapeutic [1024, 528, 847, 2265, 1906]. therapies [899, 1525, 1883, 2015, 1296]. therapy [1643, 3206, 3229, 2048, 3250, 282, 2308, 1454, 628, 1815]. there [1966, 1989]. Therneau [526]. Think [870]. thinking [3182, 310]. Thomas [10, 1402, 961, 1181, 2653, 3347]. Thompson [1799, 184]. those [727]. Thousand [455]. threatening [275]. Three [2776, 2532, 1189, 1370, 3178, 330, 355, 2723, 3070, 2975, 3150, 2667, 140, 2479, 605, 1141, 2249, 2392, 1574, 3019, 3385, 3094, 3362, 3091, 1034, 1393, 16, 1847, 1667, 839, 2066, 354, 172, 2485, 3158, 1989]. three- [355, 16, 354]. three-arm [2667, 2249, 3385, 3362, 1034]. three-class [3178]. three-dimensional [1141, 839]. Three-dose-cohort [2776]. three-level [1574, 1847, 1667]. three-outcome [2479]. three-stage [605]. threshold [2040, 2131, 1779, 348, 2650, 2546, 3309]. thresholding [2692]. thrombolytic [1643]. throughout [578]. throughput [3359]. thymoma [205]. tibolone [2320]. Tibshirani [1310]. ties [1829, 2932, 2663]. tightest [3093, 3219]. Tikhonov [2505]. Til [357]. Time [182, 828, 1530, 2257, 1638, 810, 645, 3048, 1892, 909, 1424, 551, 352, 2111, 36, 1051, 3020, 3083, 2871, 2876, 232, 1078, 1003, 2529, 1079, 279, 3229, 1242, 2166, 2695, 2592, 1807, 1830, 1191, 1496, 795, 793, 2842, 1738, 850, 2873, 2426, 3190, 673, 2083, 2289, 878, 177, 396, 33, 292, 1888, 1359, 2203, 2335, 3335, 219, 2568, 3084, 247, 2216, 2688, 3170, 209, 3217, 2861, 1974, 3254, 2330, 980, 954, 2548, 246, 2614, 1090, 3171, 475, 2609, 2902, 1510, 1889, 1895, 1261, 2417, 925, 955, 851, 2541, 1693, 2028, 760, 2955, 1278, 1680, 2339, 2837, 3119, 1454]. time [2596, 1852, 1395, 879, 1463, 2770, 1379, 1479, 2547, 3059, 942, 743, 973, 2322, 2941, 794, 1667, 672, 1421, 507, 2385, 3046, 2701, 1686, 2565, 2066, 1187, 1511, 1610, 742, 2810, 3345, 650, 2696, 218, 109, 1573, 1440, 1210, 2215, 2301, 3173, 1458, 722, 1664, 1657, 2843, 932, 1881, 2845, 2192, 1963, 3275, 3050, 1048, 2447, 2406, 2969, 3160, 1400, 2715, 661, 2009]. time-changing [2301]. time-continuous [810]. time-dependent [2257, 3048, 1892, 3083, 2166, 1738, 850, 2289, 1888, 3084, 1889, 1261, 2417, 2028, 2955, 1680, 507, 2701, 2810, 3345, 661]. time-frequency [2837]. time-independent [292]. time-lagged [3119]. time-linked [475]. time-of-year [1511]. time-points [1881]. time-scale [1573]. time-series [1667, 2215]. time-to-admission [909, 1424]. time-to-event [2842, 2335, 954, 851, 109, 1440, 1210, 932]. time-variations [1051]. Time-varying [828, 3229, 795, 793, 673, 1895, 794, 1458, 722, 1657, 2845, 1048]. timed [332]. timeliness [1643]. timepoint [2577]. times [2440, 207, 527, 1720, 2052, 2287, 2439, 2386, 1907, 2425, 2276, 1855, 644, 732, 3255, 2614, 1285, 2491, 660, 1829, 507, 3197, 335, 2810, 1128, 2561, 1912, 2486, 1184, 1692]. times-to-onset [1285]. Timing [1732, 1814, 2300, 2367, 1162]. Timpson [2839]. tissue [2556, 2764, 861]. TITE [2059]. Titrating [1136]. titration [923]. Tiwari [3383]. Tobias [3056]. Todd [3041]. tolerability [3304]. tolerance [1834]. tolerated [2056, 1100]. Tom [106]. Tombs [1784]. Tony [286]. tool [2785, 1627, 2759, 3163, 1399]. tools [3206, 580, 1750]. tooth [965, 839]. tooth-loss [839]. Topics [1547]. Toshiro [2186]. Total [38, 1690, 1257, 2902, 1832]. Totowa [1550]. toxicities [2059]. toxicity [3280, 485, 1882, 1032, 1100, 1008, 2077]. toxicokinetic [709, 1455]. toxicological [2598]. Toxicologists [210]. toxicology [2006, 467, 439]. trachoma [1792]. track [1019]. tracking [3215]. tract [2815]. Trade [2350, 619]. trade-offs [619]. traditional [2078]. Training [2162, 767, 1234]. trains [2514]. trait [1767, 800, 1500, 1934, 2191]. trait-haplotype [1767]. traits [3360, 2290, 2381, 1110, 1304, 1711, 2254, 1815, 2706]. trajectories [2285, 2204, 1491, 136]. trajectory [2393]. transcranial [3005]. transduction [3271]. Transformation [3121, 281, 1558, 1000, 2985, 1932, 1459, 2274, 3050]. Transformation-invariant [3121]. transformations [1264, 1015]. transformed [1561, 295, 841, 3060, 2195, 1174]. transforming [1682, 1446]. transfusion [2622]. transient [2294, 1841]. Transition [1050, 2695, 1814, 2758, 3195, 2461, 1476, 2583, 2405, 1083]. transitional [25]. Transitions [1002, 2111, 2270]. translation [2826]. transmission [640, 950, 1539, 2041, 1273, 639, 1499, 2969]. transmitted [3309]. transplant [1924, 970, 679, 3035, 1510, 1334]. transplantation [2606, 2116, 2608, 1576, 508]. transposition [557]. trauma [2899, 3037]. treat [2297, 1578, 1733, 3027, 1838, 3142, 3136, 3283, 2710, 1984, 1985, 669, 408, 670]. treated [2281, 1454, 3283]. Treatment [1600, 106, 1943, 330, 17, 1706, 2282, 3320, 2297, 1613, 2894, 3104, 697, 249, 3186, 1862, 1508, 859, 564, 1942, 2449, 3265, 3277, 1236, 3273, 1249, 87, 1807, 2802, 1700, 256, 3238, 2937, 729, 1456, 1518, 2740, 585, 3384, 2747, 2563, 900, 1719, 1929, 3339, 501, 2667, 95, 1060, 2068, 2973, 331, 1040, 865, 1551, 2739, 2614, 976, 2890, 1884, 476, 2415, 1032, 615, 911, 2674, 760, 512, 1506, 2491, 3136, 846, 1495, 2452, 1563, 3227, 2326, 1356, 1699, 189, 416, 2183, 2316, 1394, 499, 2553, 1034, 1237, 1878, 346, 1917, 1756, 3012, 1376, 2972, 2761, 834, 1464, 3201, 3152, 2369, 1380]. treatment [1791, 2198, 2619, 1367, 1729, 313, 376, 2675, 2421, 901, 2053, 1021, 3067, 785, 2662, 2170, 2746, 397, 107, 2561, 1583, 1815, 2971, 1315, 1657, 1022, 3031, 1026, 1885, 2018, 1020, 1881, 2957, 1828, 1426, 1427, 3160, 69]. treatment-by-centre [346]. treatment-by-gender [901]. treatment-covariate [249, 3186]. treatment-effect [2563]. treatments [114, 1928, 1743, 3238, 795, 793, 1518, 2563, 634, 451, 28, 85, 2832, 1901, 2316, 794, 2657, 3193, 1740, 3158]. Tree [1382, 59, 2402, 2383, 2875, 2242, 692, 1381, 1860]. Tree-based [2383, 1860]. Tree-structured [2402, 692]. trees [2432, 756, 598, 1070, 1276, 1016]. trend [3364, 290, 2544, 1936, 550, 3383, 1271, 1167, 2129, 2130]. trends [314, 1078, 209, 1820, 537, 73]. Trevor [1310, 821, 1206]. triads [3202, 2879]. Trial [2570, 1808, 3105, 529, 1905, 2332, 323, 1387, 1487, 2552, 2888, 79, 585, 2610, 969, 367, 1386, 2148, 2329, 2586, 699, 1597, 2521, 1929, 2897, 3097, 3009, 2144, 2756, 1042, 404, 2667, 605, 1842, 1742, 1757, 1521, 3264, 2653, 3108, 1596, 1598, 336, 1294, 1663, 512, 2630, 910, 2102, 1028, 1563, 364, 1463, 1927, 2393, 2468, 392, 254, 1043, 370, 1034, 902, 1393, 253, 1528, 2761, 2896, 422, 2668, 514, 2421, 3193, 797, 556, 2136, 1754, 109, 1740, 153, 2250, 658, 1664, 2265, 1022, 1026, 344, 3158, 1730, 888, 3066, 1731, 1200, 1427, 2831, 486, 365, 1404]. Trials [817, 144, 1402, 342, 569, 1886, 984, 1979, 717, 1792, 3041, 2579, 1778, 3368, 2481, 2331, 527, 3373, 3367, 2540, 2158, 2297, 311, 1918, 2086, 2738, 355, 1374, 875, 2894, 137, 1752, 2313, 275, 379, 3183, 1578, 895, 975, 1879, 1942, 2791, 3259, 369, 2231, 3265, 750, 1482, 382, 2665, 813, 1696, 3340, 3065, 3334, 2984, 108, 1980, 708, 1108, 1235, 3282, 444, 3157, 831, 832, 833, 985, 204, 2740, 887, 789, 2582, 3030, 1146, 2065, 2501, 2402, 2875, 1188, 907, 2099, 1703, 2563, 1038, 792, 2724, 375, 367, 627, 2026, 3327, 2, 894, 2652, 900, 1041, 5]. trials [2895, 2088, 2793, 1097, 2033, 2897, 2043, 1969, 3028, 1300, 485, 181, 1732, 1099, 2576, 3369, 2164, 665, 95, 2056, 3170, 1880, 898, 968, 2672, 2987, 2479, 770, 1368, 791, 2400, 2848, 2776, 3226, 1137, 1004, 307, 212, 988, 2777, 374, 576, 979, 1419, 1704, 2177, 2737, 1040, 996, 3027, 2409, 2503, 297, 1423, 1314, 1341, 2673, 980, 954, 1701, 3380, 368, 2249, 2631, 528, 3164, 202, 1777, 654, 1452, 81, 3108, 2113, 2392, 573, 1264, 1389, 1983, 822, 769, 986, 1981, 976, 718, 406, 1247, 1032, 1451, 3328, 1342, 1293, 2985, 3385, 3388, 867, 35, 1805, 1990, 2601, 3276, 656, 1904]. trials [1789, 616, 706, 1531, 512, 1832, 3329, 714, 2446, 2558, 3136, 1611, 465, 380, 657, 577, 345, 1688, 719, 271, 407, 757, 442, 771, 1579, 85, 958, 2611, 441, 555, 2091, 2832, 2264, 3098, 1265, 2368, 727, 80, 83, 2664, 466, 189, 416, 3023, 2798, 1171, 1851, 1612, 994, 82, 3370, 1799, 3307, 2578, 1702, 2318, 3308, 2580, 1462, 2553, 974, 987, 642, 2613, 1741, 1039, 805, 346, 2654, 1059, 16, 1546, 2689, 1756, 366, 993, 882, 3371, 990, 1135, 1464, 2232, 2712, 1208, 426, 582, 2581, 655, 1380, 1791, 2577, 815, 2159, 1233, 89, 1569, 1871, 2163]. trials [3332, 3372, 3326, 1367, 798, 1729, 1985, 1485, 1967, 2675, 1775, 2421, 2655, 2911, 2656, 2657, 1129, 2480, 3115, 372, 3279, 705, 1021, 3067, 1063, 3087, 643, 2412, 354, 218, 823, 2160, 3099, 572, 2523, 2146, 107, 424, 1025, 908, 1301, 3230, 1440, 1758, 2401, 1841, 2628, 2145, 2971, 2669, 1486, 325, 373, 1366, 2007, 371, 964, 1494, 2467, 2999, 669, 1599, 647, 3092, 513, 989, 889, 790, 2829, 3029, 3031, 604, 1662, 1885, 2018, 344, 759, 1020, 2792, 2741, 484, 997, 873, 1881, 3331, 337, 1184, 2333, 425, 1426, 1427, 1329, 896, 1, 2353, 3194, 3337, 2410, 3330]. Trials [602, 1067, 541, 1311, 1159, 1309, 1192, 287, 540, 1066, 2856, 303, 2960, 2109, 2720]. triangular [1700]. tribute [1218]. trichotomous [2290, 2381]. trigger [852]. triggers [2309]. trim [2551]. Triple [229]. Triple-goal [229]. Trivedi [433]. Tropical [306]. Troxel [2134]. true [1615, 411]. Truncated [1352, 1049, 2286, 724, 280, 3023]. truncation [2948, 2692]. truth [462]. Trying [2328, 2676]. Tsay [2345]. Tsong [3303]. Tu [2450]. tuberculin [431]. tuberculosis [431, 2635]. Tukey [94]. tumor [300, 2764]. tumor-immune [300]. Tumour [983, 205, 87, 2043, 1586, 633, 1091]. tumours [2048, 1668, 1370]. tuning [1764]. Turk [2501]. Turkman [183]. Turnbull [541, 351]. Tutorial [1849, 2398, 2031, 77, 181, 78]. Twin [2290, 2381, 349]. twins [2254, 1267]. Two [1695, 244, 1316, 3282, 3000, 3256, 1298, 1032, 9, 1164, 939, 677, 448, 1485, 532, 2683, 2930, 1615, 538, 2751, 3105, 3135, 1117, 1545, 2225, 2540, 2158, 1861, 381, 731, 2157, 1037, 2864, 2488, 1453, 2411, 736, 2209, 2386, 1497, 2286, 3065, 3366, 1853, 2888, 2582, 51, 1038, 3250, 1111, 214, 162, 2833, 1116, 1676, 2397, 634, 1899, 194, 3129, 2951, 1715, 1419, 1314, 1788, 1524, 213, 2804, 1714, 2543, 1776, 2851, 2428, 446, 447, 451, 3298, 2256, 2883, 2464, 1789, 1085, 1313, 1266, 864, 657, 1579, 949, 1362, 1605, 1089, 2475, 409, 1659, 2528, 20, 1096, 3021, 2167, 3172]. two [2235, 1528, 1114, 2045, 2712, 2896, 1669, 2022, 2159, 860, 2258, 1967, 1716, 1713, 861, 429, 1644, 2844, 3279, 1818, 2799, 1098, 823, 2160, 1322, 1553, 424, 2561, 2533, 1210, 1865, 2071, 2628, 2971, 2275, 1253, 3052, 3288, 3269, 919, 2067, 2057, 1057, 3158, 3066, 1824, 3033, 2506, 2995, 173, 2645, 1404, 1850]. two-arm [2896]. two-by-two [2751, 2488]. two-component [1824]. Two-dimensional [3256, 1419, 3033]. two-group [1788]. two-level [2844, 2506]. Two-part [1695, 446, 447, 2883, 2533]. Two-phase [448, 214, 162, 2397, 2022, 2258, 2995]. Two-sample [1316, 3105, 381, 1037, 2888, 1579, 1605, 20, 1098, 2071, 2971]. Two-sided [9, 1497, 1266, 1850]. Two-stage [3282, 3000, 1485, 532, 2683, 2930, 2540, 2158, 2411, 3065, 3366, 2582, 3250, 1314, 1524, 1776, 2851, 1789, 949, 2712, 2159, 1967, 3279, 823, 2160, 424, 2561, 2628, 3158, 3066]. two-stage-designs [861]. two-state [1114, 2045]. two-step [1453]. Two-treatment [1032]. Type [436, 810, 44, 1361, 666, 492, 3045, 96, 1246, 1557, 2308, 2419, 860, 1019, 2656, 2396, 2401, 376, 1933]. Tyrer [1709, 1591].

U.K [885, 143, 1402, 602, 961, 303, 1194, 541, 342, 844, 436, 1311, 453, 435, 962, 419, 437, 1206, 418, 187, 105]. U.K. [3207]. U.S. [1387, 2162, 265, 2920, 2927]. U.S.A [91, 782, 1550, 257, 1825, 210, 1244, 94]. ugly [697]. UK [1351, 259]. ulcer [2554, 684]. ULTRA [2351]. umbrella [2939]. unadjusted [2999]. unbalanced [3135, 25, 3339, 212, 2804, 203]. unbiased [3298]. unblinded [1353, 1391, 1392]. uncertain [2554]. Uncertainties [1618]. Uncertainty [1584, 542, 33, 1336, 522, 12, 2784, 1834, 3116, 1254, 2219, 2683, 2684]. Unconditional [2591, 538, 213, 991, 3145]. unconfounded [3277]. uncorrelated [1085, 1362]. under-reporting [825]. undergraduate [745]. underlying [313]. underpowered [3013]. Understanding [598, 2961, 1016]. undertaken [788]. undesirable [1906]. undiagnosed [2241]. Unequal [368, 1411, 1971, 429, 2410]. unexposed [1316]. unification [3100]. unified [3280, 2501, 2899, 1185]. uniform [2874, 2086, 1120, 2275]. unifying [1916]. uninterpretable [807]. unit [2513, 328, 1369, 972, 1059, 2364]. United [1082, 2965]. units [2838, 2721]. unity [2, 815, 1058]. univariate [1152, 3333, 1060, 1442]. universe [2360]. University [1592, 144, 455, 885, 1158, 1608, 843, 1415, 199, 399, 420, 1787, 1245, 1180, 1802, 2503, 1142, 1531, 1595, 221, 1671, 3192, 3056, 440, 1143, 1532, 2519, 146, 540, 1825, 1698, 259, 433, 1549, 94]. unknown [1674, 1456, 393, 612, 1201, 2360, 1128, 1167]. unlinked [2486]. unmatched [2093, 475]. unmeasured [1331, 2733, 1718, 3093, 2395, 3219]. unobservable [2828]. unobserved [3255, 2392, 2710, 2069]. unordered [2536, 3262]. unpaired [1453]. unrelated [3118]. unreported [71]. unresectable [1427]. UNRRA [1219]. untreated [2281, 1090]. unusual [2836]. unverified [731]. up-and-down [979, 3235, 1008]. update [867, 1]. updated [2403]. Updating [3095, 955, 1468]. upon [2403]. Uranium [3020]. urea [1430]. Urn [86, 1743, 1436]. Use [3202, 2891, 937, 1249, 2454, 281, 2733, 512, 2339, 477, 3023, 972, 642, 2780, 2879, 1167, 645, 2020, 2021, 1706, 1643, 2060, 1198, 2766, 516, 296, 1814, 1700, 1939, 1518, 1038, 2148, 1968, 252, 219, 1812, 3297, 2893, 475, 2687, 1773, 2813, 906, 2994, 523, 498, 1897, 215, 310, 1259, 3355, 2774, 1401, 3246, 1759, 1666, 2485, 1250, 1046, 425, 1231, 661, 1377]. used [1626, 466, 2182, 805, 151]. useful [3081, 15, 39, 208]. usefulness [2987]. users [2756, 507, 2485]. Using [192, 18, 2865, 553, 1572, 3020, 1078, 2189, 1942, 552, 782, 1536, 3157, 2916, 2285, 795, 793, 565, 1794, 844, 3226, 1742, 2015, 2840, 2709, 32, 3211, 2316, 1376, 794, 1627, 2365, 1209, 948, 3041, 2527, 3359, 2030, 982, 2225, 3243, 2407, 72, 205, 1861, 3104, 2389, 2945, 381, 2606, 2754, 70, 3079, 2703, 249, 1633, 709, 931, 2356, 1674, 3022, 1035, 2621, 1076, 3286, 2040, 1561, 2915, 1950, 1273, 1477, 1751, 1287, 1191, 116, 3342, 1456, 1009, 575, 631, 2375, 120, 410, 413, 3210, 2261, 393, 1444, 598, 808, 829, 1597, 5, 2088, 2317, 1177, 2242, 629]. using [3312, 2693, 1182, 2206, 3335, 520, 2905, 247, 2688, 2087, 1297, 2068, 2773, 1802, 1014, 934, 2168, 3376, 307, 1974, 3254, 2063, 1687, 2806, 2409, 1141, 297, 1788, 1269, 2299, 1777, 2921, 1333, 2614, 1596, 1598, 336, 1781, 2460, 2803, 659, 1574, 1773, 2940, 1190, 684, 688, 1056, 3019, 2213, 406, 3361, 231, 2857, 675, 2061, 2985, 2126, 2028, 2422, 2699, 2100, 1085, 1832, 1266, 1156, 3268, 1278, 2078, 1680, 3055, 428, 1016, 2419, 3005, 531, 1362, 1318, 2596, 3341, 1089, 2770, 1256, 1587, 1125, 2273, 1922, 1479, 1659, 2294, 2377, 2024, 254, 1920, 775, 1420, 499, 1304, 1409, 799, 3008]. using [764, 226, 2976, 478, 1459, 3198, 1174, 233, 993, 2593, 312, 1501, 3062, 1335, 2719, 2362, 2205, 1230, 2027, 1464, 3325, 2253, 2686, 2841, 1083, 1711, 2574, 89, 1105, 2016, 2675, 2632, 2796, 1196, 2655, 1092, 2956, 3214, 2565, 1488, 73, 291, 1610, 667, 971, 329, 1896, 218, 2746, 3106, 1763, 1440, 3185, 2007, 2505, 2573, 3278, 1839, 2265, 1748, 2067, 632, 1399, 1885, 547, 759, 3158, 1931, 2901, 952, 633, 2337, 2755, 957, 2670, 3110, 2129, 2130, 227, 1130, 2839, 2252, 2404, 1091, 387, 2399, 1156, 2961]. usual [552]. Utah [734]. uterine [643]. utility [1676, 56, 2921, 23]. utility-respecting [56]. Utilization [27, 2358, 1657, 3117]. utilized [3054]. utilizing [3227].

V [2344, 975, 2050, 2140, 3380, 958, 494, 416, 260, 362]. vaccinated [519]. vaccination [412, 2925, 2546]. vaccine [1977, 926, 487, 1477, 1496, 733, 3070, 850, 2897, 292, 350, 2975, 3150, 86, 1661, 1137, 1004, 1559, 3108, 1190, 706, 1110, 1420, 2546]. vaccines [2978]. vague [1773]. vaguely [955]. vaguenes [2676]. vagueness [2328]. Valencia [1608, 1924]. Valentin [221]. valid [111]. validate [1396]. validated [2133]. validating [58]. Validation [1468, 324, 1916, 1848, 2820, 2293, 1477, 1822, 2617, 2699, 3018, 1454, 1614, 2868, 1937, 1343, 340, 3124, 663, 1370]. validatory [1088]. Validity [179, 305, 2273, 1399]. value [3386, 2937, 578, 2450, 801, 3310, 2532, 588, 949, 216, 1336, 1434, 1053, 2722, 2260, 1730]. valued [2837]. values [2423, 554, 4, 2405, 977, 2384, 3363, 1175, 3198, 857, 669, 2067, 3247, 3213, 2831, 121]. vanishing [3212, 3289]. Variability [542, 545, 3040, 673, 1279, 1582, 2515, 1728, 2318, 837, 3230]. Variable [361, 2302, 2496, 618, 936, 1302, 1361, 2865, 1633, 2293, 1044, 1561, 2815, 1658, 3128, 1188, 1151, 2781, 3208, 3195, 634, 2991, 1321, 983, 3338, 3111, 3380, 1560, 2826, 1491, 588, 2699, 3276, 3044, 595, 27, 28, 1010, 735, 3283, 1418, 1175, 1174, 2574, 2277, 329, 662, 3064, 776, 1435, 1722, 1837, 2594, 2296, 2303, 1253, 2825, 1002, 2862, 1231]. variables [1121, 2702, 2281, 2297, 471, 2332, 1078, 2473, 1899, 2168, 2283, 1061, 2733, 1718, 752, 1202, 851, 3136, 2534, 3281, 2186, 1897, 3114, 1015, 2618, 2574, 3141, 2396, 3262, 1836, 3052, 3288, 632, 2305, 3003, 3138]. Variance [2983, 1823, 449, 3236, 1651, 2311, 516, 272, 627, 1842, 958, 3281, 775, 2045, 2516, 2370, 1644, 2661, 1816, 2342, 919, 2957, 1731, 2129, 105]. variances [2814, 752, 659, 429]. variant [1584, 2268]. variants [2766]. variation [2379, 8, 611, 3287, 246, 1470, 1138, 1760, 2054, 1312, 1241, 2904, 2365]. variations [1051, 292, 350, 1096]. varicella [937]. various [237]. vary [752]. Varying [3265, 827, 3229, 795, 793, 1935, 673, 1895, 1913, 3059, 794, 1458, 722, 828, 1657, 2845, 1048]. varying-coefficient [2845]. Vasan [2640, 2642]. vasculitis [326]. VAT [1750]. vCJD [362]. vector [1449]. vectorcardiography [3175]. Vedel [1873]. velocity [1308]. ventricular [432]. Verfahrensbibliothek [494]. Verification [1615, 1126, 2800, 2437, 2187, 2438, 3123, 1173, 173, 1289, 3047, 3162]. Verlag [1157, 1647, 526, 434, 497, 301, 129, 543, 40, 1182, 1352, 1195, 1872, 494, 285, 872, 41]. Vermunt [132]. Vernon [2961]. Versions [1156, 1266].

Versuchsplanung [494]. versus [1706, 697, 2151, 2872, 1000, 791, 1640, 1576, 267, 1389, 1228, 847, 1096, 2232, 2410]. vertically [1332]. very [99]. Verzani [2399]. veterinary [36]. vi [221, 357]. via [2716, 3032, 3384, 2168, 3129, 2861, 1834, 2278, 234, 2495, 1495, 20, 1917, 1251, 3036, 1058, 2390, 1449]. Vic [183, 1774]. Vickers [3134]. Victor [494]. Viel [417]. view [1905, 1226, 1225, 1286, 2786]. viewpoint [1296]. VII [131, 455, 145, 199, 92, 2409, 435, 302, 1156, 106, 187]. viii [509, 420, 453, 285]. violators [771]. Viral [999, 1249, 829, 1332, 2951, 1137, 307, 2299, 2634, 2970, 3087, 1583, 1798, 2462]. Virginia [1594]. virologic [2300]. virological [564]. virus [666, 3214, 1934, 1130]. Visual [1071, 1555]. visualisation [1181]. Visualizing [3244, 2785]. vital [2379, 8, 2889]. vitamin [99, 1028]. vitro [978]. Vittinghoff [2140]. Voelkl [1156]. Vol [437, 129, 525, 131]. Volkers [2245]. Volume [299, 818, 2373, 2512]. volume-outcome [2373]. Volumes [127, 818, 525]. voluntary [1278]. volunteers [1930]. Vontheim [343]. Vries [2770]. vs [2907, 1715, 1714, 1716, 1713, 2539, 2908, 3137].

W [819, 1460, 3135, 1217, 1591, 1565, 1424, 2604, 13, 129, 1550, 541, 211, 182, 439, 1925, 1218, 1156, 1362, 359, 2295, 29, 94, 1709]. Waagepetersen [1504]. waiting [1924, 207, 909, 1424, 672]. Wakefield [1142]. Wald [2593]. walking [293]. Wallenstein [1384]. Waller [257]. Walley [2327]. Walter [803, 2604, 439, 2409, 2817]. Wang [3122, 1644]. Ware [2502]. Warn [1799]. Warren [129, 11]. Washington [821]. wave [2510]. wave-like [2510]. wavelet [2692]. Wavelets [106]. waves [164, 2783]. way [3061, 2342]. Wayne [359]. wealth [2913]. weather [2460]. weekly [1856]. Wehby [2826]. Weibull [718, 3253, 3017]. weight [1658]. Weighted [3103, 2601, 2701, 3025, 2990, 351, 1580, 2382, 3024, 970, 599, 2338, 2869, 688, 1056, 1797, 856, 912, 3274, 2067, 1130]. Weighting [1061, 795, 793, 1084, 1495, 794, 2956, 1828]. weights [3229, 1084, 85, 3099]. Weimar [2820]. Weinberg [1844, 2269, 1910]. Weir [2327, 1563]. Weisberg [400]. Weiss [1925]. Welch [1557, 2841]. Welch-type [1557]. Welcome [459, 2912]. Wellek [1205]. Welsh [368]. Wen [1056]. Wermuth [42]. West [1608, 2841]. Westfall [782]. Whelan [131]. where [2895, 2577]. whether [256, 2904]. which [859, 115, 516, 2378, 3077]. while [1198, 1865]. whilst [1879]. Whitehall [2259]. Whitehead [3041, 1309]. Whitley [1925]. Whitney [1939, 3209, 1940, 1941, 2096, 2663]. Who [870, 899, 282, 1393, 1391, 2182, 1399]. whose [2969]. Wicher [1544]. wide [2124, 2780, 2390]. width [617, 2422]. Wilcox [872]. Wilcoxon [3209, 3045, 2096, 2663, 2645]. Wilcoxon-based [2645]. Wiley [1530, 819, 509, 3222, 1752, 1351, 360, 127, 183, 402, 1566, 1134, 1402, 358, 2856, 303, 1194, 452, 182, 1018, 1774, 3182, 3181, 417, 1144, 438, 1311, 1309, 524, 1801, 359, 781, 1363, 1145, 544, 1471, 2502, 2859, 287, 260, 3138, 3231, 1206, 298, 400, 418, 820, 103, 147, 496, 1416]. Wiley-Interscience [1530, 2502]. Willett [1595]. Willi [3138]. William [2856, 1461, 11, 1471].

WinBUGS [1773]. window [2411]. Windows [1156, 2203, 219, 1156]. wine [998]. Winkelmann [41]. winner [705]. Wisconsin [1317, 710]. wise [2929]. withdrawal [2711]. withdrawals [2710]. Within [607, 1879, 516, 3075, 318, 2680, 659, 23, 3361, 812, 1480, 3055, 1238, 2778, 1136, 2054]. within-area [23]. within-patient [2778]. within-study [2680]. Within-subject [607, 659, 2054]. without [125, 1950, 10, 432, 2789, 6, 1418, 2689, 1644, 2401, 919]. Wittes [2501]. Wojtele [142]. Wolfe [147]. Wolfinger [2604, 782, 129]. women [2253, 2927, 3309, 1502]. Won [3386, 3310]. Wood [1925]. Woodward [341, 2004]. Woolson [1471]. work [1390]. workers [1204]. workflow [1955]. Working [3146, 3267]. Working-correlation-structure [3146]. World [884, 2350, 106, 259, 1825, 1919]. Worldwide [258]. Worm [450]. worth [2091]. Wright [960]. wrong [473]. WTA [620]. WTP [620]. www.elsevier.com [1786]. Wykes [357]. Wynn [2593].

X [2503, 1192, 817, 3139, 144, 129, 603, 1134, 1144, 1159, 495, 130, 300, 1671, 142, 494, 1193, 2519, 2005, 146, 185, 1825, 1530, 2138, 361, 1550, 42, 1352, 1533, 780, 130, 1531, 1471, 2719, 287, 400, 103, 1384, 1825, 259, 184]. xenograft [1586]. xi [1194, 1018, 844, 1774, 884, 1609, 132, 525, 1532, 287, 128, 197, 186]. Xiao [2437]. Xiao-Hua [2437]. xic [3231]. xii [819, 360, 2604, 885, 286, 402, 1195, 3192, 544, 1698, 1839]. xiii [1460, 93, 1592, 526, 1608, 1751, 1402, 600, 2856, 3057, 182, 1145, 872, 2109, 437, 147]. xiv [198, 1158, 1364, 782, 1750, 2501, 602, 1067, 211, 1311, 1309, 104, 299, 821, 359, 1840, 1143, 3122, 437, 418, 1517, 184, 1549]. xix [2138, 1351, 2824, 434, 3181, 417, 2502, 2841]. xv [1565, 421, 183, 399, 1753, 1352, 2140, 1363, 440, 419, 1065, 1610, 1384, 433, 1684]. xvi [1594, 90, 1787, 1245, 1205, 2399, 1310, 436, 1156, 3056, 43, 103]. xvii [961, 566, 2503, 1142, 1156, 1872, 2858, 260, 1206, 298, 820, 105]. xviii [1182, 541, 3182, 524, 2859]. xx [1595, 2961]. xxi [3180, 870]. xxii [543, 2004]. xxiii [1181, 2857, 1546, 2345, 259, 842]. xxiv [1647]. xxv [1786, 1533, 818]. xxvi [400]. xxvii [1801, 2718]. xxx [818, 1471]. xxxi [496]. xxxii [3204]. xxxiv [131].

Yadolah [1415]. Ye [125, 10]. year [1317, 2174, 1470, 1511, 2248]. years [738, 306, 2263, 2230, 2723, 2123, 2070, 1219, 2046, 2564, 683, 2163, 3185]. Yeong [2345]. Ying [10]. Yongyi [2032]. York [819, 1157, 1647, 1565, 542, 2138, 144, 526, 497, 1750, 301, 843, 2501, 129, 1566, 1134, 543, 1438, 201, 1352, 182, 566, 1195, 1310, 438, 495, 524, 1595, 1156, 821, 1363, 1145, 2519, 872, 1192, 1546, 1471, 1516, 2345, 298, 400, 103, 1384, 1825, 496, 1684, 871, 633]. Yosef [782]. Youden [2650]. Young [131, 251, 1536, 1430, 2413]. Youth [2193]. youthful [384]. Yu [2450]. Yu-Kang [2450]. Yudi [1592]. Yutaka [2558].

Z [1245, 1715, 1714, 958, 3219, 128]. Zbynek [435]. Zeger [1532, 1516]. Zelen [867]. Zelterman [2519, 187]. Zero [777, 2893, 594, 2891, 2180, 2120, 2, 3354, 1481, 1394, 3152, 1945, 2650, 815, 1837, 2023, 2341, 2892]. Zero-inflated [777, 594, 2180, 2120, 3354, 1837, 2341, 2892]. zero-inflation [1481, 1394, 2023]. zeros [2802, 3354, 3268]. Zhang [497, 1561, 3303]. Zhong [3303]. Zhou [2437]. zidovudine [796]. Ziegler [2820, 1340]. Zipi [2787]. zoster [559, 1925]. Zou [3381].

References

Zelen:1990:RCD

[1] M. Zelen. Randomized consent designs for clinical trials: an update. Statistics in Medicine, 9(6):645–656, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [867].

Farrington:1990:TSS

[2] Conor P. Farrington and Godfrey Manning. Test statistics and sample size formulae for comparative binomial trials with null hypothesis of nonzero risk difference or non-unity relative risk. Statistics in Medicine, 9 (12):1447–1454, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [815].

Benichou:1991:MAE

[3] Jacques Benichou. Methods of adjustment for estimating the attributable risk in case-control studies: a review. Statistics in Medicine, 10(11): 1753–1773, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [414].

Goodman:1992:CRV

[4] Steven N. Goodman. A comment on replication, P -values and evidence. Statistics in Medicine, 11(7):875–879, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [857].

Frison:1992:RMC

[5] Lars Frison and Stuart J. Pocock. Repeated measures in clinical trials: Analysis using mean summary statistics and its implications for design. Statistics in Medicine, 11(13):1685–1704, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [297].

Lu:1997:CCG

[6] Ying Lu, Keying Ye, Ashwini K. Mathur, Siu Hui, Thomas P. Fuerst, and Harry K. Genant. Comparative calibration without a gold standard. Statistics in Medicine, 16(16):1889–1905, August 30, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [10, 125].

May:1997:CID

[7] Warren L. May and William D. Johnson. Confidence intervals for differences in correlated binary proportions. Statistics in Medicine, 16(18): 2127–2136, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [13, 11, 29].

Fay:1997:SPE

[8] Michael P. Fay and Eric J. Feuer. A semi-parametric estimate of extra-Poisson variation for vital rates. Statistics in Medicine, 16(21):2389– 2401, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2379].

Newcombe:1998:TSC

[9] Robert G. Newcombe. Two-sided confidence intervals for the single proportion: comparison of seven methods. Statistics in Medicine, 17(8): 857–872, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1850].

Dunn:1998:LEC

[10] Graham Dunn. Letter to the Editor: Comparative calibration without a gold standard, by Ying Lu, Keying Ye, Ashwini K. Mathur, Siu Hui, Thomas P. Fuerst and Harry K. Genant, Statistics in Medicine, 16, 1889– 1905 (1997). Statistics in Medicine, 17(11):1294–1296, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [6].

Lui:1998:LEC

[11] Kung-Jong Lui. Letter to the Editor: Confidence intervals for differences in correlated binary proportions, by Warren L. May and William D. Johnson, Statistics in Medicine, 16, 2127–2136 (1997). Statistics in Medicine, 17(17):2017–2020, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [7, 13, 29].

Cronin:1998:AUM

[12] Kathleen A. Cronin, Julie M. Legler, and Ruth D. Etzioni. Assessing uncertainty in microsimulation modelling with application to cancer screening interventions. Statistics in Medicine, 17(21):2509–2523, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Broemeling:1999:LEC

[13] Lyle D. Broemeling. Letter to the Editor: Confidence intervals for differences in correlated binary proportions, by W. L. May and W. D. Johnson, Statistics in Medicine, 16, 2127–2136 (1997). Statistics in Medicine, 18 (1):109, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [7, 11, 29].

Normand:1999:MAF

[14] Sharon-Lise T. Normand. Meta-analysis: formulating, evaluating, combining, and reporting. Statistics in Medicine, 18(3):321–359, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [78].

Ash:1999:UMM

[15] Arlene Ash and Michael Shwartz. R^{2}: a useful measure of model performance when predicting a dichotomous outcome. Statistics in Medicine, 18(4):375–384, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [39].

Proschan:1999:MCP

[16] Michael A. Proschan. A multiple comparison procedure for three- and four-armed controlled clinical trials. Statistics in Medicine, 18(7):787– 798, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [355, 354].

Atkinson:1999:OBC

[17] Anthony C. Atkinson. Optimum biased-coin designs for sequential treatment allocation with covariate information. Statistics in Medicine, 18 (14):1741–1752, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [397].

Alonzo:1999:UCR

[18] Todd A. Alonzo and Margaret S. Pepe. Using a combination of reference tests to assess the accuracy of a new diagnostic test. Statistics in Medicine, 18(22):2987–3003, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [387].

Brown:1999:APC

[19] Barry W. Brown, James Lovato, and Kathy Russell. Asymptotic power calculations: description, examples, computer code. Statistics in Medicine, 18(22):3137–3151, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2000:TST

[20] Wei Pan. A two-sample test with interval censored data via multiple imputation. Statistics in Medicine, 19(1):1–11, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kelly:2000:SAR

[21] Patrick J. Kelly and Lynette L-Y. Lim. Survival analysis for recurrent event data: an application to childhood infectious diseases. Statistics in Medicine, 19(1):13–33, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2000].

Lindsey:2000:DMM

[22] J. K. Lindsey. Directly modelling matched case-control data. Statistics in Medicine, 19(1):35–44, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lasserre:2000:BES

[23] Virginie Lasserre, Chantal Guihenneuc-Jouyaux, and Sylvia Richardson. Biases in ecological studies: utility of including within-area distribution of confounders. Statistics in Medicine, 19(1):45–59, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraggi:2000:ERM

[24] David Faraggi. The effect of random measurement error on receiver operating characteristic (ROC) curves. Statistics in Medicine, 19(1): 61–70, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2000:TMB

[25] Pai-Lien Chen, Emelita Wong, Rosalie Dominik, and Markus J. Steiner. A transitional model of barrier methods compliance with unbalanced loss to follow-up. Statistics in Medicine, 19(1):71–82, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McLaren:2000:PSA

[26] C. E. McLaren, E. L. Kambour, G. J. McLachlan, H. C. Lukaski, X. Li, G. M. Brittenham, and G. D. McLaren. Patient-specific analysis of sequential haematological data by multiple linear regression and mixture distribution modelling. Statistics in Medicine, 19(1):83–98, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Magder:2000:UMI

[27] Laurence S. Magder, Michael A. Sloan, Show-Hong Duh, Joseph F. Abate, and Steven J. Kittner. Utilization of multiple imperfect assessments of the dependent variable in a logistic regression analysis. Statistics in Medicine, 19(1):99–111, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mazumdar:2000:CPV

[28] Madhu Mazumdar and Jill R. Glassman. Categorizing a prognostic variable: review of methods, code for easy implementation and applications to decision-making about cancer treatments. Statistics in Medicine, 19 (1):113–132, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2000:LEC

[29] T. Tango. Letter to the Editor: Confidence intervals for differnces in correlated binary proportions by W. L. May and W. D. Johnson, Statistics in Medicine, 16, 2127–2136 (1997). Statistics in Medicine, 19(1): 133–136, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [7, 13, 11].

May:2000:AR

[30] W. L. May and W. D. Johnson. Authors’ reply. Statistics in Medicine, 19(1):136–139, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steyerberg:2000:PMB

[31] E. W. Steyerberg, M. J. C. Eijkemans, J. C. Van Houwelingen, K. L. Lee, and J. D. F. Habbema. Prognostic models based on literature and individual patient data in logistic regression analysis. Statistics in Medicine, 19(2):141–160, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mariotto:2000:UAM

[32] Angela B. Mariotto and Arduino Verdecchia. Using AIDS mortality data to reconstruct HIV/AIDS epidemics. Statistics in Medicine, 19(2):161– 174, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gigli:2000:UAI

[33] Anna Gigli and Arduino Verdecchia. Uncertainty of AIDS incubation time and its effects on back-calculation estimates. Statistics in Medicine, 19(2):175–189, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2000:TSD

[34] Toshiro Tango. A test for spatial disease clustering adjusted for multiple testing. Statistics in Medicine, 19(2):191–204, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2000:SSC

[35] Aiyi Liu, James M. Boyett, and Xiaoping Xiong. Sample size calculation for planning group sequential longitudinal trials. Statistics in Medicine, 19(2):205–220, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bedrick:2000:BAF

[36] Edward J. Bedrick, Ronald Christensen, and Wesley O. Johnson. Bayesian accelerated failure time analysis with application to veterinary epidemiology. Statistics in Medicine, 19(2):221–237, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bouckaert:2000:SDP

[37] Andre Bouckaert. Smoking during pregnancy: foetal growth retardation and other risks for the newborn. Statistics in Medicine, 19(2):239–254, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2000:TDI

[38] Lawrence I-Kuei Lin. Total deviation index for measuring individual agreement with applications in laboratory performance and bioequivalence. Statistics in Medicine, 19(2):255–270, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:2000:LEU

[39] Marc Buyse. Letter to the Editor: R^{2}: a useful measure of model performance when predicting a dichotomous outcome, by A. Ash and M. Schwartz, Statistics in Medicine, 18, 375–384 (1999). Statistics in Medicine, 19(2):271–274, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [15].

Gijbels:2000:BRP

[40] Irène Gijbels. Book review: Petit Cours de Statistique. Klaus Krickeberg, Springer Verlag, Berlin, 1996 (in French). No. of pages: ix + 149. Price:

DM29. ISBN 3-540-61609-8. Statistics in Medicine, 19(2):275–276, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Windmeijer:2000:BRE

[41] Frank Windmeijer. Book review: Econometric Analysis of Count Data. 2nd edn, R. Winkelmann, Springer-Verlag, 1997. No. of pages: 193. Price: £55.50, $85.90, DM 128. ISBN 3-540-62201-2. Statistics in Medicine, 19 (2):276–277, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frydenberg:2000:BRM

[42] Morten Frydenberg. Book review: Multivariate Dependencies. Models, Analysis and Interpretation. D. R. Cox and N. Wermuth, Chapman and Hall, London, 1996, No. of pages: 272. Price: £39. ISBN 0-412-75410-X. Statistics in Medicine, 19(2):277–278, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Postelnicu:2000:BRB

[43] Tiberiu Postelnicu. Book review: Biplots. J. C. Gower and D. J. Hand, Chapman and Hall, London, 1996. No. of pages: xvi + 277. Price: £32. ISBN 0-412-71630-5. Statistics in Medicine, 19(2):278–280, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Auranen:2000:BCA

[44] Kari Auranen. Back-calculating the age-specific incidence of recurrent subclinical Haemophilus influenzae type B infection. Statistics in Medicine, 19(3):281–296, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bellocco:2000:JAH

[45] Rino Bellocco and Ian C. Marschner. Joint analysis of HIV and AIDS surveillance data in back-calculation. Statistics in Medicine, 19(3):297– 311, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Platt:2000:BCI

[46] Robert W. Platt, James A. Hanley, and Hong Yang. Bootstrap confidence intervals for the sensitivity of a quantitative diagnostic test. Statistics in Medicine, 19(3):313–322, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Platt:2000:SAS

[47] Robert W. Platt. Saddlepoint approximations for small sample logistic regression problems. Statistics in Medicine, 19(3):323–334, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2000:PTJ

[48] Hyune-Ju Kim, Michael P. Fay, Eric J. Feuer, and Douglas N. Midthune. Permutation tests for joinpoint regression with applications to cancer rates. Statistics in Medicine, 19(3):335–351, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [386].

Chen:2000:MCC

[49] Yuh-Ing Chen. Multiple comparisons in carcinogenesis study with rightcensored survival data. Statistics in Medicine, 19(3):353–367, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Witte:2000:RSS

[50] John S. Witte, Robert C. Elston, and Lon R. Cardon. On the relative sample size required for multiple comparisons. Statistics in Medicine, 19 (3):369–372, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:2000:TET

[51] Allan Donner, Mohamed M. Shoukri, Neil Klar, and Emma Bartfay. Testing the equality of two dependent kappa statistics. Statistics in Medicine, 19(3):373–387, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nottingham:2000:SAA

[52] Quinton J. Nottingham and Jeffrey B. Birch. A semiparametric approach to analysing dose-response data. Statistics in Medicine, 19(3):389–404, February 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bebchuk:2000:MIS

[53] Judith D. Bebchuk and Rebecca A. Betensky. Multiple imputation for simple estimation of the hazard function based on interval censored data. Statistics in Medicine, 19(3):405–419, February 15, 2000. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schafer:2000:P

[54] Helmut Schäfer, Mitchell H. Gail, and Patrick Royston. Preface. Statistics in Medicine, 19(4):421, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Windeler:2000:PWD

[55] Jürgen Windeler. Prognosis — what does the clinician associate with this notion? Statistics in Medicine, 19(4):425–430, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:2000:PFU

[56] Jørgen Hilden. Prevalence-free utility-respecting summary indices of diagnostic power do not exist. Statistics in Medicine, 19(4):431–440, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmoor:2000:SSC

[57] Claudia Schmoor, Willi Sauerbrei, and Martin Schumacher. Sample size considerations for the evaluation of prognostic factors in survival analysis. Statistics in Medicine, 19(4):441–452, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:2000:WDW

[58] Douglas G. Altman and Patrick Royston. What do we mean by validating a prognostic model? Statistics in Medicine, 19(4):453–473, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dannegger:2000:TSD

[59] Felix Dannegger. Tree stability diagnostics and some remedies for instability. Statistics in Medicine, 19(4):475–491, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jensen:2000:RCB

[60] Katrin Jensen, Hans-Helge Müller, and Helmut Schäfer. Regional confidence bands for ROC curves. Statistics in Medicine, 19(4):493–509, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emir:2000:CDM

[61] Birol Emir, Sam Wieand, Sin-Ho Jung, and Zhiliang Ying. Comparison of diagnostic markers with repeated measurements: a non-parametric ROC curve approach. Statistics in Medicine, 19(4):511–523, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martus:2000:SME

[62] P. Martus. Statistical methods for the evaluation of diagnostic measurements concerning paired organs. Statistics in Medicine, 19(4):525–540, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwarzer:2000:MAN

[63] Guido Schwarzer, Werner Vach, and Martin Schumacher. On the misuses of artificial neural networks for prognostic and diagnostic classification in oncology. Statistics in Medicine, 19(4):541–561, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henderson:2000:CCM

[64] Robin Henderson and Helen Prince. Choice of conditional models in bivariate survival. Statistics in Medicine, 19(4):563–574, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feldmann:2000:MOC

[65] Uwe Feldmann and Ingo Steudel. Methods of ordinal classification applied to medical scoring systems. Statistics in Medicine, 19(4):575–586, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2000:MSM

[66] Per Kragh Andersen, Sille Esbjerg, and Thorkild I. A. Sørensen. Multistate models for bleeding episodes and mortality in liver cirrhosis. Statistics in Medicine, 19(4):587–599, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klinger:2000:IMP

[67] Artur Klinger, Felix Dannegger, and Kurt Ulm. Identifying and modelling prognostic factors with censored data. Statistics in Medicine, 19 (4):601–615, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Slate:2000:SML

[68] Elizabeth H. Slate and Bruce W. Turnbull. Statistical models for longitudinal biomarkers of disease onset. Statistics in Medicine, 19(4):617–637, February 29, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2000:OTA

[69] Wei Zhu and Weng Kee Wong. Optimal treatment allocation in comparative biomedical studies. Statistics in Medicine, 19(5):639–648, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biggerstaff:2000:CDT

[70] Brad J. Biggerstaff. Comparing diagnostic tests: a simple graphic using likelihood ratios. Statistics in Medicine, 19(5):649–663, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunson:2000:AUM

[71] David B. Dunson and Clarice R. Weinberg. Accounting for unreported and missing intercourse in human fertility studies. Statistics in Medicine, 19(5):665–679, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beckett:2000:PEB

[72] Laurel A. Beckett and Daniel J. Tancredi. Parametric empirical Bayes estimates of disease prevalence using stratifed samples from community populations. Statistics in Medicine, 19(5):681–695, March 15, 2000. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sorribas:2000:EAR

[73] A. Sorribas, J. March, and E. O. Voit. Estimating age-related trends in cross-sectional studies using S-distributions. Statistics in Medicine, 19(5):697–713, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanley:2000:GAN

[74] James A. Hanley, Abdissa Negassa, and Michael D. deB. Edwardes. GEE analysis of negatively correlated binary responses: a caution. Statistics in Medicine, 19(5):715–722, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blackman:2000:IEC

[75] Nicole J.-M. Blackman and John J. Koval. Interval estimation for Cohen’s kappa as a measure of agreement. Statistics in Medicine, 19(5): 723–741, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chong:2000:REB

[76] Mei Yin Chong, David C. Hamilton, and David E. C. Cole. Retrospective estimation of the birth prevalence for delayed onset disorders: application to cystic fibrosis in Nova Scotia. Statistics in Medicine, 19(5):743–751, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carlin:2000:LET

[77] John B. Carlin. Letter to the Editor: Tutorial in Biostatistics. Metaanalysis: formulating, evaluating, combining, and reporting, by S-L. T. Normand, Statistics in Medicine, ‘bf 18, 321–359 (1999). Statistics in Medicine, 19(5):753–759, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stijnen:2000:LET

[78] Theo Stijnen. Letter to the Editor: Tutorial in Biostatistics. Metaanalysis: formulating, evaluating, combining, and reporting, by S-L. Normand, Statistics in Medicine, 18, 321–359 (1999). Statistics in Medicine, 19(5):759–761, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [14].

DAgostino:2000:CAC

[79] Ralph B. D’Agostino, Sr. Controlling alpha in a clinical trial: the case for secondary endpoints. Statistics in Medicine, 19(6):763–766, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moye:2000:ACC

[80] Lemuel A. Moyé. Alpha calculus in clinical trials: considerations and commentary for the new millennium. Statistics in Medicine, 19(6):767– 779, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:2000:DAC

[81] Gary G. Koch. Discussion for ‘alpha calculus in clinical trials: considerations and commentary for the new millennium’. Statistics in Medicine, 19(6):781–784, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ONeill:2000:CAC

[82] Robert T. O’Neill. Commentary on ‘alpha calculus in clinical trials: considerations and commentary for the new millennium’. Statistics in Medicine, 19(6):785–793, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moye:2000:RCA

[83] Lemuel A. Moyé. Response to commentaries on ‘alpha calculus in clinical trials: considerations for the new millennium’. Statistics in Medicine, 19(6):795–799, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:2000:OME

[84] J. K. Lindsey. Obtaining marginal estimates from conditional categorical repeated measurements models with missing data. Statistics in Medicine, 19(6):801–809, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2000:MRW

[85] Devan V. Mehrotra and Radha Railkar. Minimum risk weights for comparing treatments in stratified binomial trials. Statistics in Medicine, 19(6):811–825, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hernandez-Suarez:2000:UMV

[86] Carlos M. Hernández-Suárez and Carlos Castillo-Chavez. Urn models and vaccine efficacy estimation. Statistics in Medicine, 19(6):827–835, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chevret:2000:MTE

[87] Sylvie Chevret, Michel Leporrier, and Claude Chastang. Measures of treatment effectiveness on tumour response and survival: a multi-state model approach. Statistics in Medicine, 19(6):837–848, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morton-Jones:2000:AIR

[88] Tony Morton-Jones, Peter Diggle, Louise Parker, Heather O. Dickinson, and Keith Binks. Additive isotonic regression models in epidemiology. Statistics in Medicine, 19(6):849–859, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2000:RMC

[89] Stephen Senn, Lynda Stevens, and Nish Chaturvedi. Repeated measures in clinical trials: simple strategies for analysis using summary measures. Statistics in Medicine, 19(6):861–877, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:2000:BRS

[90] A. W. F. Edwards. Book review: Statistical Evidence: a Likelihood Paradigm. R. Royall, Chapman and Hall, London, 1997. No. of pages: xvi + 191. Price: £39.95. ISBN 0-412-04411-0. Statistics in Medicine, 19(6):879–880, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2000:BRM

[91] Per Kragh Andersen. Book review: Modern Epidemiology. 2nd edn. Kenneth J. Rothman and Sander Greenland, Lippincott-Raven, U.S.A., 1998. No. of pages: 737. Price: $74.50. ISBN 0-316-75780-2. Statistics in Medicine, 19(6):881–882, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jarrett:2000:BRI

[92] Richard Jarrett. Book review: Introduction to the Design and Analysis of Experiments. G. M. Clarke and R. E. Kempson, Arnold, London, 1997. No. of pages. vii + 334. Price: £19.99. ISBN 0-340-64555-5. Statistics in Medicine, 19(6):882–883, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arjas:2000:BRM

[93] Elja Arjas. Book review: Markov Chain Monte Carlo: Stochastic Simulation for Bayesian Inference. Dani Gamerman, Chapman and Hall, London, 1997. No. of pages: xiii + 245. Price: £24.95. ISBN 0-412-81820-5. Statistics in Medicine, 19(6):883–884, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:2000:BRP

[94] Weng Kee Wong. Book review: The Practice of Data Analysis: Essays in Honor of John W. Tukey. D. R. Brillinger, L. T. Fernholz and S. Morgenthaler (eds), Princeton University Press, U.S.A., 1998. No. of pages: 337. Price: $49.50. ISBN 0-691-05782-6. Statistics in Medicine, 19(6):884–885, March 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauck:2000:GTE

[95] Walter W. Hauck, Terry Hyslop, and Sharon Anderson. Generalized treatment effects for clinical trials. Statistics in Medicine, 19(7):887–899, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kieser:2000:RCS

[96] Meinhard Kieser and Tim Friede. Re-calculating the sample size in internal pilot study designs with control of the type I error rate. Statistics in Medicine, 19(7):901–911, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stewart:2000:DDR

[97] William H. Stewart and Stephen J. Ruberg. Detecting dose response with contrasts. Statistics in Medicine, 19(7):913–921, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qiu:2000:SST

[98] Peihua Qiu, Melvin L. Moeschberger, Glen E. Cooke, and Pascal J. Goldschmidt-Clermont. Sample size to test for interaction between a specific exposure and a second risk factor in a pair-matched case-control study. Statistics in Medicine, 19(7):923–935, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graham:2000:BIG

[99] Patrick Graham. Bayesian inference for a generalized population attributable fraction: the impact of early vitamin a levels on chronic lung disease in very low birthweight infants. Statistics in Medicine, 19(7): 937–956, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bell:2000:BAS

[100] B. Sue Bell and Lyle D. Broemeling. A Bayesian analysis for spatial processes with application to disease mapping. Statistics in Medicine, 19 (7):957–974, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gill:2000:RML

[101] Paramjit S. Gill. A robust mixed linear model analysis for longitudinal data. Statistics in Medicine, 19(7):975–987, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilton:2000:FOC

[102] Joan F. Hilton. Functions of oral candidiasis episodes that are highly prognostic for AIDS. Statistics in Medicine, 19(7):989–1004, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sorensen:2000:BRE

[103] Thorkild I. A. Sørensen. Book review: Epidemiology Kept Simple: an Introduction to Classic and Modern Epidemiology. B. Burst Gerstman, Wiley, New York, 1998. No. of pages: xvi + 301. Price: £25.95. ISBN 0-471-24029-X. Statistics in Medicine, 19(7):1005–1006, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knorr-Held:2000:BRA

[104] Leonhard Knorr-Held. Book review: Analysis of Incomplete Multivariate Data. J. L. Schafer, Chapman and Hall, London, 1997. No. of pages: xiv + 430. Price: £39.95. ISBN 0-412-04061-1. Statistics in Medicine, 19(7): 1006–1008, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tjur:2000:BRV

[105] Tue Tjur. Book review: Variance Components Estimation. P. S. R. S. Rao, Chapman and Hall, U.K. 1997. No. of pages: xvii + 199. Price: £42.95. ISBN 0-412-72860-5. Statistics in Medicine, 19(7):1008–1009, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raz:2000:BRW

[106] Jonathan A. Raz. Book review: Wavelets: an Elementary Treatment of Theory and Applications. Tom H. Koornwinder (ed.), World Scientific, London, 1995. No. of pages: vii + 240. Price: £49. ISBN 9-8102-1388-3. Statistics in Medicine, 19(7):1009, April 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:2000:EMT

[107] Peter F. Thall, Randall E. Millikan, and Hsi-Guang Sung. Evaluating multiple treatment courses in clinical trials. Statistics in Medicine, 19(8): 1011–1028, April 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2000:SAP

[108] Shein-Chung Chow, Jun Shao, and Hsin-Tsung Ho. On statistical analysis for placebo-challenging designs in clinical trials. Statistics in Medicine, 19(8):1029–1037, April 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tangen:2000:NPC

[109] Catherine M. Tangen and Gary G. Koch. Non-parametric covariance methods for incidence density analyses of time-to-event data from a randomized clinical trial and their complementary roles to proportional hazards regression. Statistics in Medicine, 19(8):1039–1058, April 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steyerberg:2000:PML

[110] Ewout W. Steyerberg, Marinus J. C. Eijkemans, Frank E. Harrell, Jr., and J. Dik F. Habbema. Prognostic modelling with logistic regression analysis: a comparison of selection and estimation methods in small data sets. Statistics in Medicine, 19(8):1059–1079, April 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldman:2000:SMR

[111] Daniel A. Goldman and Jean D. Brender. Are standardized mortality ratios valid for public health data analysis? Statistics in Medicine, 19(8): 1081–1088, April 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Llorca:2000:CSP

[112] Javier Llorca and Miguel Delgado-Rodríguez. A comparison of several procedures to estimate the confidence interval for attributable risk in case-control studies. Statistics in Medicine, 19(8):1089–1099, April 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Catellier:2000:TGR

[113] Diane J. Catellier and Keith E. Muller. Tests for Gaussian repeated measures with missing data in small samples. Statistics in Medicine, 19(8):1101–1114, April 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agresti:2000:SCT

[114] Alan Agresti and Jonathan Hartzel. Strategies for comparing treatments on a binary response with multi-centre data. Statistics in Medicine, 19 (8):1115–1139, April 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carpenter:2000:BCI

[115] James Carpenter and John Bithell. Bootstrap confidence intervals: when, which, what? A practical guide for medical statisticians. Statistics in Medicine, 19(9):1141–1164, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cui:2000:EHI

[116] Jisheng Cui and Niels G. Becker. Estimating HIV incidence using dates of both HIV and AIDS diagnoses. Statistics in Medicine, 19(9):1165–1177, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nurminen:2000:MRA

[117] Markku Nurminen, Maija Hytönen, and Eeva Sala. Modelling the reproducibility of acoustic rhinometry. Statistics in Medicine, 19(9):1179– 1189, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2000:CSC

[118] Patrick Royston. Choice of scale for cubic smoothing spline models in medical applications. Statistics in Medicine, 19(9):1191–1205, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deguen:2000:ECR

[119] Séverine Deguen, Guy Thomas, and Nguyen Phong Chau. Estimation of the contact rate in a seasonal SEIR model: application to chickenpox incidence in France. Statistics in Medicine, 19(9):1207–1216, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Do:2000:GAA

[120] K-A. Do, B. M. Broom, P. Kuhnert, D. L. Duffy, A. A. Todorov, S. A. Treloar, and N. G. Martin. Genetic analysis of the age at menopause by using estimating equations and Bayesian random effects models. Statistics in Medicine, 19(9):1217–1235, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhuang:2000:AEA

[121] Dongliang Zhuang, Nathaniel Schenker, Jeremy M. G. Taylor, Veronique Mosseri, and Bernard Dubray. Analysing the effects of anaemia on local recurrence of head and neck cancer when covariate values are missing. Statistics in Medicine, 19(9):1237–1249, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whittemore:2000:C

[122] Alice S. Whittemore. Correction. Statistics in Medicine, 19(9):1251, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dyba:2000:C

[123] Tadeusz Dyba. Correction. Statistics in Medicine, 19(9):1251, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2000:C

[124] Kung-Jong Lui. Correction. Statistics in Medicine, 19(9):1251, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bolfarine:2000:LEC

[125] Heleno Bolfarine and Filidor Vilca-Labra. Letter to the Editor: Comparative calibration without a gold standard Y. Lu, K. Ye, A. Mathur, S. Hul, T. Fuerst and H. Genant, Statistics in Medicine 1997; 16:1889– 1905. Statistics in Medicine, 19(9):1253–1254, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [6].

Lu:2000:AR

[126] Ying Lu, Keying Ye, and Aswini Mathur. Authors’ reply. Statistics in Medicine, 19(9):1254, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2000:BRE

[127] D. R. Cox. Book review: Encyclopedia of Biostatistics. Peter Armitage and Theodore Colton (eds), 6 Volumes, Wiley, Chichester, 1998. No. of pages: 4898. Price: £1495.00. ISBN 0-471-97576-1. Statistics in Medicine, 19(9):1255–1257, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shore:2000:BRZ

[128] Sarah Shore. Book review: A-Z of Medical Statistics: a Companion for Critical Appraisal. F. Pereira-Maxwell, Arnold, London, 1998. No. of pages: xi + 118, Price: £15.99. ISBN 0-340-71940-0. Statistics in Medicine, 19(9):1257–1258, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunsmuir:2000:BRM

[129] William T. M. Dunsmuir. Book review: Modeling Longitudinal and Spatially Correlated Data: Methods, Applications and Future Directions. T. G. Gregoire, D. R. Brillinger, P. J. Diggle, E. Russek-Cohen, W. G. Warren and R. D. Wolfinger (eds). Lecture Notes in Statistics, vol 122, Springer-Verlag, New York, 1997. No. of pages: x + 402. Price: £30.50. ISBN 0-387-98216-7. Statistics in Medicine, 19(9):1258–1260, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:2000:BRS

[130] J. K. Lindsey. Book review: Statistical Inference Based on the Likelihood. Adelchi Azzalini, Chapman and Hall, London, 1996. No. of pages: x + 341. Price: £39.95. ISBN 0-412-60650-X. Statistics in Medicine, 19(9): 1260–1261, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hakulinen:2000:BRC

[131] Timo Hakulinen. Book review: Cancer Incidence in Five Continents Vol. VII. D. M. Parkin, S. L. Whelan, J. Ferlay, L. Raymond and J. Young (eds), IARC Scientific Publications No. 143, Lyon, 1997. No. of pages: xxxiv + 1240. Price: £129. ISBN 92-832-2143-5. Statistics in Medicine, 19(9):1261–1263, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martinussen:2000:BRL

[132] Torben Martinussen. Book review: Log-Linear Models for Event Histories. Jeroen K. Vermunt, Sage Publications, London, 1997. No. of pages: xi + 348. Price: £37. ISBN 0-7619-0937-0. Statistics in Medicine, 19(9): 1263–1264, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2000:SGF

[133] Beiyao Zheng. Summarizing the goodness of fit of generalized linear models for longitudinal data. Statistics in Medicine, 19(10):1265–1275, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2000:RAG

[134] Yue-Cune Chang. Residuals analysis of the generalized linear models for longitudinal data. Statistics in Medicine, 19(10):1277–1293, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reidpath:2000:IIA

[135] Daniel D. Reidpath, Mark R. Diamond, Gunter Hartel, and Paul Glasziou. Improving interpretability: γ as an alternative to R^{2} as a measure of effect size. Statistics in Medicine, 19(10):1295–1302, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2000:LCM

[136] Haiqun Lin, Charles E. McCulloch, Bruce W. Turnbull, Elizabeth H. Slate, and Larry C. Clark. A latent class mixed model for analysing biomarker trajectories with irregularly scheduled observations. Statistics in Medicine, 19(10):1303–1318, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2000:PCP

[137] Vance W. Berger. Pros and cons of permutation tests in clinical trials. Statistics in Medicine, 19(10):1319–1328, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2000:NPT

[138] Eunsik Park and Young Jack Lee. Non-parametric test of ordered alternatives in incomplete blocks. Statistics in Medicine, 19(10):1329–1337, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ambrosius:2000:QCM

[139] Walter T. Ambrosius and Siu L. Hui. A quality control measure for longitudinal studies with continuous outcomes. Statistics in Medicine, 19(10):1339–1362, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hill:2000:CTT

[140] Elizabeth G. Hill, Li Ding, and Lance A. Waller. A comparison of three tests to detect general clustering of a rare disease in Santa Clara County, California. Statistics in Medicine, 19(10):1363–1378, May 30, 2000. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osler:2000:BRS

[141] Merete Osler and Eva Prescott. Book review: Statistical Methods in Medical Research. The Statistics of Smoking. B. S. Everitt, G. Dunn and T.

R. Holford (eds), Arnold, London, 1998. No. of pages: 123. Price: £57.00. ISSN 0962-2802. Statistics in Medicine, 19(10):1379–1380, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McKnight:2000:BRI

[142] Barbara McKnight. Book review: An Introduction to Statistical Modelling. Wojtele J. Krzanowski, Arnold, London, 1998. No. of pages: x + 252. Price: £19.99. ISBN 0-340-69185-9. Statistics in Medicine, 19(10): 1380–1381, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Commenges:2000:BRA

[143] Daniel Commenges. Book review: Actuarial Models for Disability Insurance. S. Haberman and E. Pitacco, Chapman and Hall, U.K., 1999. No. of pages: ix + 280. Price: £35.00. ISBN 0-849-30389-3. Statistics in Medicine, 19(10):1381–1382, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bjorner:2000:BRQ

[144] Jakob Bjørner. Book review: Quality of Life Assessment in Clinical Trials; Methods and Practice. Maurice J. Staquet, Ron D. Hays and Peter M. Fayers (eds), Oxford University Press, New York, 1998. No. of pages: x + 360. Price: £29.50. ISBN 0-1926-2785-6. Statistics in Medicine, 19(10):1382–1384, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Datta:2000:BRS

[145] Susmita Datta. Book review: Statistics in Human Genetics. Pak Sham, Arnold, London, 1998. No. of pages: vii + 290. Price: £24.99. ISBN 0-340-66241-7. Statistics in Medicine, 19(10):1384–1385, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rasmussen:2000:BRM

[146] Lars S. Rasmussen. Book review: Mathematics and Statistics in Anaesthesia. Steven Cruickshante, Oxford University Press, Oxford, 1998. No. of pages: x + 258. Price: £49.50. ISBN 0-19-262313-3. Statistics in Medicine, 19(10):1385–1386, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tiit:2000:BRN

[147] Ene-Margit Tiit. Book review: Nonparametric Statistical Methods. Myles Hollander and Douglas A. Wolfe, Wiley, Chichester, 1999. No. of pages: xiii + 779. Price: £39.95. ISBN 0-471-19045-4. Statistics in Medicine, 19(10):1386–1388, May 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kryscio:2000:F

[148] Richard J. Kryscio and Frederick A. Schmitt. Foreword. Statistics in Medicine, 19(11-12):1389–1391, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelb:2000:MPA

[149] Douglas J. Gelb. Measurement of progression in Alzheimer’s disease: a clinician’s perspective. Statistics in Medicine, 19(11-12):1393–1400, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mohs:2000:LSC

[150] Richard C. Mohs, James Schmeidler, and Mosen Aryan. Longitudinal studies of cognitive, functional and behavioural change in patients with Alzheimer’s disease. Statistics in Medicine, 19(11-12):1401–1409, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBelle:2000:RCT

[151] Gerald van Belle and Alice Arnold. Reliability of cognitive tests used in Alzheimer’s disease. Statistics in Medicine, 19(11-12):1411–1420, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galasko:2000:MCC

[152] Douglas R. Galasko, Robert L. Gould, Ian S. Abramson, and David P. Salmon. Measuring cognitive change in a cohort of patients with Alzheimer’s disease. Statistics in Medicine, 19(11-12):1421–1432, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2000:ALD

[153] Ronald G. Thomas, Julie D. Berg, Mary Sano, and Leon Thal. Analysis of longitudinal data in an Alzheimer’s disease clinical trial. Statistics in Medicine, 19(11-12):1433–1440, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edland:2000:BRH

[154] Steven D. Edland. Blomqvist revisited: how and when to test the relationship between level and longitudinal rate of change. Statistics in Medicine, 19(11-12):1441–1452, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cronin-Stubbs:2000:PIB

[155] Diane Cronin-Stubbs, Steven T. DeKosky, John C. Morris, and Denis A. Evans. Promoting interactions with basic scientists and clinicians: the NIA Alzheimer’s Disease Data Coordinating Center. Statistics in Medicine, 19(11-12):1453–1461, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmitt:2000:PIB

[156] Frederick A. Schmitt, J. Philip Miller, and Richard J. Kryscio. Promoting interactions with basic scientists and clinicians: summary of the panel session. Statistics in Medicine, 19(11-12):1463–1468, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beckett:2000:CBS

[157] Laurel A. Beckett. Community-based studies of Alzheimer’s disease: statistical challenges in design and analysis. Statistics in Medicine, 19 (11-12):1469–1480, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:2000:MPI

[158] Ron Brookmeyer and Sarah Gray. Methods for projecting the incidence and prevalence of chronic diseases in ageing populations: application to Alzheimer’s disease. Statistics in Medicine, 19(11-12):1481–1493, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beiser:2000:CEI

[159] Alexa Beiser, Ralph B. D’Agostino, Sr., Sudha Seshadri, Lisa M. Sullivan, and Philip A. Wolf. Computing estimates of incidence, including lifetime risk: Alzheimer’s disease in the Framingham Study. The Practical Incidence Estimators (PIE) macro. Statistics in Medicine, 19(11-12):1495–1522, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emsley:2000:EOR

[160] Christine L. Emsley, Sujuan Gao, Kathleen S. Hall, and Hugh C. Hendrie. Estimating odds ratios adjusting for misclassification in Alzheimer’s disease risk factor assessment. Statistics in Medicine, 19(11-12):1523– 1530, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foster:2000:BCD

[161] Norman L. Foster, Eszter Gombosi, Cheryl Teboe, and Roderick J. A. Little. Balanced centralized and distributed database design in a clinical research environment. Statistics in Medicine, 19(11-12):1531–1544, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2000:EID

[162] Sujuan Gao and Siu L. Hui. Estimating the incidence of dementia from two-phase sampling with non-ignorable missing data. Statistics in Medicine, 19(11-12):1545–1554, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hall:2000:CPM

[163] Charles B. Hall, Richard B. Lipton, Martin Sliwinski, and Walter F. Stewart. A change point model for estimating the onset of cognitive decline in preclinical Alzheimer’s disease. Statistics in Medicine, 19(11-12):1555–1566, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hui:2000:SFW

[164] Siu L. Hui and Sujuan Gao. Spacing of follow-up waves in incidence studies. Statistics in Medicine, 19(11-12):1567–1575, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Izmirlian:2000:EID

[165] Grant Izmirlian, Dwight Brock, and Lon White. Estimating incidence of dementia subtypes: assessing the impact of missed cases. Statistics in Medicine, 19(11-12):1577–1591, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2000:MDP

[166] Xinhua Liu, Jeanne A. Teresi, and Christine Waternaux. Modelling the decline pattern in functional measures from a prevalent cohort study.

Statistics in Medicine, 19(11-12):1593–1606, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mendiondo:2000:MMM

[167] Marta S. Mendiondo, J. Wesson Ashford, Richard J. Kryscio, and Frederick A. Schmitt. Modelling mini mental state examination changes in Alzheimer’s disease. Statistics in Medicine, 19(11-12):1607–1616, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Milliken:2000:MEM

[168] J. Kevin Milliken and Steven D. Edland. Mixed effect models of longitudinal Alzheimer’s disease data: a cautionary note. Statistics in Medicine, 19(11-12):1617–1629, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mungas:2000:AIR

[169] Dan Mungas and Bruce R. Reed. Application of item response theory for development of a global functioning measure of dementia with linear measurement properties. Statistics in Medicine, 19(11-12):1631–1644, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2000:DML

[170] R. John Roberts, Beverly S. Musick, Bola Olley, Kathleen S. Hall, Hugh C. Hendrie, and Allen B. O. O. Oyediran. Data management in a longitudinal cross-cultural study. Statistics in Medicine, 19(11-12):1645– 1649, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teresi:2000:MPM

[171] Jeanne A. Teresi, Marjorie Kleinman, and Katja Ocepek-Welikson. Modern psychometric methods for detection of differential item functioning: application to cognitive assessment measures. Statistics in Medicine, 19 (11-12):1651–1683, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tyas:2000:DIB

[172] Suzanne L. Tyas, John J. Koval, and Linda L. Pederson. Does an interaction between smoking and drinking influence the risk of Alzheimer’s disease? Results from three Canadian data sets. Statistics in Medicine, 19(11-12):1685–1696, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2000:ARA

[173] Xiao-Hua Zhou and Richard E. Higgs. Assessing the relative accuracies of two screening tests in the presence of verification bias. Statistics in Medicine, 19(11-12):1697–1705, June 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Engels:2000:HSS

[174] Eric A. Engels, Christopher H. Schmid, Norma Terrin, Ingram Olkin, and Joseph Lau. Heterogeneity and statistical significance in meta-analysis: an empirical study of 125 meta-analyses. Statistics in Medicine, 19(13): 1707–1728, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cronin:2000:CCS

[175] Kathleen A. Cronin and Eric J. Feuer. Cumulative cause-specific mortality for cancer patients in the presence of other causes: a crude analogue of relative survival. Statistics in Medicine, 19(13):1729–1740, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dyba:2000:CDA

[176] Tadeusz Dyba and Timo Hakulinen. Comparison of different approaches to incidence prediction based on simple interpolation techniques. Statistics in Medicine, 19(13):1741–1752, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geskus:2000:IPC

[177] Ronald B. Geskus. On the inclusion of prevalent cases in HIV/AIDS natural history studies through a marker-based estimate of time since seroconversion. Statistics in Medicine, 19(13):1753–1769, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2000:SPP

[178] Beiyao Zheng and Alan Agresti. Summarizing the predictive power of a generalized linear model. Statistics in Medicine, 19(13):1771–1781, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guggenmoos-Holzmann:2000:VSS

[179] Irene Guggenmoos-Holzmann and Hans C. van Houwelingen. The (in)validity of sensitivity and specificity. Statistics in Medicine, 19(13): 1783–1792, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Littell:2000:MCS

[180] Ramon C. Littell, Jane Pendergast, and Ranjini Natarajan. Modelling covariance structure in the analysis of repeated measures data. Statistics in Medicine, 19(13):1793–1819, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldstein:2000:LET

[181] Harvey Goldstein. Letter to the Editor: Tutorial in Biostatistics — Longitudinal data analysis (repeated measures) in clinical trials, by P.-S. Albert, Statistics in Medicine, 1999, 18, 1707–1732. Statistics in Medicine, 19(13):1821, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henderson:2000:BRA

[182] Robin Henderson. Book review: Applied Survival Analysis. Regression Modeling of Time to Event Data. David W. Hosmer and Stanley Lemeshow, Wiley, New York, 1999. No. of pages: xiii + 386. Price: £51.95. ISBN 0-471-15410-5. Statistics in Medicine, 19(13):1823–1824, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2000:BRS

[183] Christopher Cox. Book review: Statistics for the Environment 4. Statistical Aspects of Health and the Environment. Vic Barnett, Alfred Stein and K. Feridun Turkman (eds). Wiley, London, 1999. No. of pages: xv + 404. Price: £95. ISBN 0-4719-7645-8. Statistics in Medicine, 19(13): 1824–1825, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Traat:2000:BRT

[184] Imbi Traat. Book review: Theory of Sample Surveys. M. E. Thompson, Chapman and Hall, London, 1997. No. of pages: xiv + 305. Price: £39.95. ISBN 0-412-31780-X. Statistics in Medicine, 19(13):1825–1827, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sandberg:2000:BRM

[185] Sverre Sandberg. Book review: Modelling and Estimation of Measurement Error. Michele Neuilly, Lavoisier Publishing, Paris, 1999. No. of pages: x + 694. Price: £119.50. ISBN 1-898298-59-9. Statistics in Medicine, 19(13):1827–1828, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vach:2000:BRS

[186] Werner Vach. Book review: Statistical Analysis of Medical Data. New Developments. Brian S. Everitt and Graham Dunn (eds), Arnold, London, 1998. No. of pages: xi + 340. Price: £29.99. ISBN 0340-6775-9. Statistics in Medicine, 19(13):1828–1829, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steyerberg:2000:BRM

[187] Ewout Steyerberg. Book review: Models for Discrete Data. Daniel Zelterman, Oxford Science Publications, Clarendon Press, Oxford, U.K. 1999. No. of pages: vii + 233. Price: £35. ISBN 0-19-852436-6. Statistics in Medicine, 19(13):1829–1830, July 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2000:SME

[188] Patrick Royston. A strategy for modelling the effect of a continuous covariate in medicine and epidemiology. Statistics in Medicine, 19(14): 1831–1847, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:2000:ETE

[189] Nico Nagelkerke, Vaclav Fidler, Roos Bernsen, and Martien Borgdorff. Estimating treatment effects in randomized clinical trials in the presence of non-compliance. Statistics in Medicine, 19(14):1849–1864, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [416].

Gelfand:2000:MCR

[190] Alan E. Gelfand and Fei Wang. Modelling the cumulative risk for a falsepositive under repeated screening events. Statistics in Medicine, 19(14): 1865–1879, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Formann:2000:RAG

[191] Anton K. Formann. Rater agreement and the generalized Rudas–Clogg– Lindsay index of fit. Statistics in Medicine, 19(14):1881–1888, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Akhtar-Danesh:2000:UAT

[192] Noori Akhtar-Danesh and David R. Appleton. Using an antedependence test to analyse post-operative pain measurements. Statistics in Medicine, 19(14):1889–1899, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deBEdwardes:2000:GOR

[193] Michael D. deB. Edwardes and Marc Baltzan. The generalization of the odds ratio, risk ratio and risk difference to r × k tables. Statistics in Medicine, 19(14):1901–1914, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [283].

Hsiao:2000:CPT

[194] Chuhsing Kate Hsiao, Jung-Ying Tzeng, and Chih-Hao Wang. Comparing the performance of two indices for spatial model selection: application to two mortality data. Statistics in Medicine, 19(14):1915–1930, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tiensuwan:2000:ALL

[195] Montip Tiensuwan, Satinee Lertprapai, Jeeraphat Sirichaisinthop, and Aporn Lawmepol. Application of log-linear models to malaria patients in Thailand. Statistics in Medicine, 19(14):1931–1945, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hopper:2000:BRM

[196] John L. Hopper. Book review: Mathematical and statistical methods for genetic analysis. K. Lange, Springer, Berlin, 1997. No. of pages: 290. Price: £36.50. ISBN 0-387-94909-7. Statistics in Medicine, 19(14):1947– 1949, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tusch:2000:BRE

[197] Günter Tusch. Book review: Expertensysteme in der Medizin — eine Einführung mit Beispielen. Manfred Schnabel, G. Fischer, Stuttgart, 1996. No. of pages: xi + 243. Price: DM 68. ISBN 3-437-21056-4. Statistics in Medicine, 19(14):1949–1950, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:2000:BRM

[198] Peter Bauer. Book review: Multiple comparisons — theory and methods. Jason C. Hsu, Chapman and Hall, London, 1996. No. of pages: xiv + 277. Price: £35. ISBN 0-412-9828-1. Statistics in Medicine, 19(14):1951, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feveile:2000:BRA

[199] Helene Feveile. Book review: Analyzing repeated surveys. G. Firebaugh, Sage University paper series on quantitative applications in the social sciences, series no. 07-115, 1997. No of pages: vii + 72. Price: £8.99. ISBN 0-8039-7398-5. Statistics in Medicine, 19(14):1951–1952, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2000:BRT

[200] Richard J. Cook. Book review: The theory of dispersion models. Bent Jørgensen, Chapman and Hall, 1997. No. of pages: 237. Price: £39.95. ISBN 0-412-99718-8. Statistics in Medicine, 19(14):1952–1953, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Girard:2000:BRD

[201] Pascal Girard. Book review: Design and analysis of animal studies in pharmaceutical development. Shein-Chung Chow and Jen Pei Liu (eds), Marcel Dekker, New York, 1998. No. of pages: 472. Price: $175. ISBN 0-8247-0130-5. Statistics in Medicine, 19(14):1953–1954, July 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kinukawa:2000:ICI

[202] Naoko Kinukawa, Tsuyoshi Nakamura, Kouhei Akazawa, and Yoshiaki Nose. The impact of covariate imbalance on the size of the logrank test in randomized clinical trials. Statistics in Medicine, 19(15):1955–1967, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [757].

Marshall:2000:LDM

[203] Guillermo Marshall and Anna E. Barón. Linear discriminant models for unbalanced longitudinal data. Statistics in Medicine, 19(15):1969–1981, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davidov:2000:DCP

[204] Ori Davidov and Marvin Zelen. Designing cancer prevention trials: a stochastic model approach. Statistics in Medicine, 19(15):1983–1995, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:2000:CTS

[205] Colin B. Begg, Laura D. Cramer, E. S. Venkatraman, and Juan Rosai. Comparing tumour staging and grading systems: a case study and a review of the issues, using thymoma as a model. Statistics in Medicine, 19(15):1997–2014, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2000:STI

[206] Dongchu Sun, Robert K. Tsutakawa, Hoon Kim, and Zhuoqiong He. Spatio-temporal interaction with disease mapping. Statistics in Medicine, 19(15):2015–2035, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armstrong:2000:FSA

[207] Paul W. Armstrong. First steps in analysing NHS waiting times: avoiding the ‘stationary and closed population’ fallacy. Statistics in Medicine, 19 (15):2037–2051, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2000:UMN

[208] Patrick Royston. A useful monotonic non-linear model with applications in medicine and epidemiology. Statistics in Medicine, 19(15):2053–2066, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2000:ATT

[209] Feifang Hu and William F. Rosenberger. Analysis of time trends in adaptive designs with application to a neurophysiology experiment. Statistics in Medicine, 19(15):2067–2075, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vermeulen:2000:BRS

[210] Roel Vermeulen. Book review: Statistics and Experimental Design for Toxicologists. 3rd edition. Shayne C. Gad, CRC Press, U.S.A. 1999. No. of pages: 437. Price: $79.95. ISBN 0-8493-3132-8. Statistics in Medicine, 19 (15):2077, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Helland:2000:BRS

[211] Inge S. Helland. Book review: Statistical Regression with Measurement Error. Chi-Lun Cheng and John W. Van Ness, Kendall’s Library of Statistics 6, Arnold, London, 1999. No. of pages: xiv + 262. Price: £35.

ISBN: 0-340-61461-7. Statistics in Medicine, 19(15):2077–2078, August 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hung:2000:ECD

[212] H. M. James Hung. Evaluation of a combination drug with multiple doses in unbalanced factorial design clinical trials. Statistics in Medicine, 19 (16):2079–2087, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2000:AUT

[213] Seung-Ho Kang and James J. Chen. An approximate unconditional test of non-inferiority between two proportions. Statistics in Medicine, 19 (16):2089–2100, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [538].

Gao:2000:EDP

[214] Sujuan Gao, Siu L. Hui, Kathleen S. Hall, and Hugh C. Hendrie. Estimating disease prevalence from two-phase surveys with non-response at the second phase. Statistics in Medicine, 19(16):2101–2114, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reiser:2000:MED

[215] Benjamin Reiser. Measuring the effectiveness of diagnostic markers in the presence of measurement error through the use of ROC curves. Statistics in Medicine, 19(16):2115–2129, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parker:2000:EVI

[216] Robert A. Parker. Estimating the value of an internal biostatistical consulting service. Statistics in Medicine, 19(16):2131–2145, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hundborg:2000:FTF

[217] Heidi H. Hundborg, Malene Højbjerre, Ole Bjarne Christiansen, and Steffen L. Lauritzen. Familial tendency to foetal loss analysed with Bayesian graphical models by Gibbs sampling. Statistics in Medicine, 19(16):2147– 2168, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamura:2000:CTO

[218] Roy N. Tamura, Douglas E. Faries, and Jin Feng. Comparing time to onset of response in antidepressant clinical trials using the cure model and the Cramér–von Mises test. Statistics in Medicine, 19(16):2169– 2184, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauptmann:2000:UST

[219] Michael Hauptmann, Jay H. Lubin, Philip Rosenberg, Jürgen Wellmann, and Lothar Kreienbrock. The use of sliding time windows for the exploratory analysis of temporal effects of smoking histories on lung cancer risk. Statistics in Medicine, 19(16):2185–2194, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schinazi:2000:PCC

[220] Rinaldo B. Schinazi. The probability of a cancer cluster due to chance alone. Statistics in Medicine, 19(16):2195–2198, August 30, 2000. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Madsen:2000:BRN

[221] Birger Madsen. Book review: Neural Networks. H. Abdi, D. Valentin and B. Edelman, Sage University Paper Series on Quantitative Applications in the Social Sciences, Series no. 07-124, 1999. No. of pages: vi + 90. Price: £8.99. ISBN 0-7619-1440-4. Statistics in Medicine, 19(16):2199– 2200, August 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2000:P

[222] Andrew Lawson, Annibale Biggeri, Charmaine Dean, and Lance Waller. Preface. Statistics in Medicine, 19(17-18):2201–2202, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bithell:2000:CDM

[223] J. F. Bithell. A classification of disease mapping methods. Statistics in Medicine, 19(17-18):2203–2215, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Group:2000:DMM

[224] Disease Mapping Collaborative Group. Disease mapping models: an empirical evaluation. Statistics in Medicine, 19(17-18):2217–2241, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byers:2000:ICM

[225] Simon Byers and Julian Besag. Inference on a collapsed margin in disease mapping. Statistics in Medicine, 19(17-18):2243–2249, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickle:2000:EST

[226] Linda Williams Pickle. Exploring spatio-temporal patterns of mortality using mixed effects models. Statistics in Medicine, 19(17-18):2251–2263, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2000:CHM

[227] Li Zhu and Bradley P. Carlin. Comparing hierarchical models for spatiotemporally misaligned data using the deviance information criterion. Statistics in Medicine, 19(17-18):2265–2278, September 15–30, 2000. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eberly:2000:ICI

[228] Lynn E. Eberly and Bradley P. Carlin. Identifiability and convergence issues for Markov chain Monte Carlo fitting of spatial models. Statistics in Medicine, 19(17-18):2279–2294, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2000:TGE

[229] Wei Shen and Thomas A. Louis. Triple-goal estimates for disease mapping. Statistics in Medicine, 19(17-18):2295–2308, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelman:2000:MQA

[230] Andrew Gelman, Phillip N. Price, and Chia yu Lin. A method for quantifying artefacts in mapping methods illustrated by application to headbanging. Statistics in Medicine, 19(17-18):2309–2320, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leroux:2000:MSD

[231] Brian G. Leroux. Modelling spatial disease rates using maximum likelihood. Statistics in Medicine, 19(17-18):2321–2332, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bohning:2000:STM

[232] Dankmar Böhning, Ekkehart Dietz, and Peter Schlattmann. Space-time mixture modelling of public health data. Statistics in Medicine, 19(17-18):2333–2344, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ranta:2000:PSA

[233] J. Ranta and A. Penttinen. Probabilistic small area risk assessment using GIS-based data: a case study on Finnish childhood diabetes. Statistics in Medicine, 19(17-18):2345–2359, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2000:CMD

[234] Andrew B. Lawson. Cluster modelling of disease incidence via RJMCMC methods: a comparative evaluation. Statistics in Medicine, 19(17-18): 2361–2375, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stern:2000:PPM

[235] Hal S. Stern and Noel Cressie. Posterior predictive model checks for disease mapping models. Statistics in Medicine, 19(17-18):2377–2397, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Talbot:2000:ESF

[236] Thomas O. Talbot, Martin Kulldorff, Steven P. Forand, and Valerie B. Haley. Evaluation of spatial filters to create smoothed maps of health data. Statistics in Medicine, 19(17-18):2399–2408, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yasui:2000:EEV

[237] Yutaka Yasui, Hao Liu, Joan Benach, and Marcy Winget. An empirical evaluation of various priors in the empirical Bayes estimation of small area disease risks. Statistics in Medicine, 19(17-18):2409–2420, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2000:PBP

[238] Y. C. MacNab and C. B. Dean. Parametric bootstrap and penalized quasi-likelihood inference in conditional autoregressive models. Statistics in Medicine, 19(17-18):2421–2435, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Semenciw:2000:MID

[239] Robert M. Semenciw, Nhu D. Le, Loraine D. Marrett, Diane L. Robson, Donna Turner, and Stephen D. Walter. Methodological issues in the development of the Canadian cancer incidence atlas. Statistics in Medicine, 19(17-18):2437–2449, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2000:SCR

[240] Andrew B. Lawson and Fiona L. R. Williams. Spatial competing risk models in disease mapping. Statistics in Medicine, 19(17-18):2451–2467, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leyland:2000:MSM

[241] Alastair H. Leyland, Ian H. Langford, Jon Rasbash, and Harvey Goldstein. Multivariate spatial models for event data. Statistics in Medicine, 19(17-18):2469–2478, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2000:MCS

[242] Richard Thomas and Tony E. Smith. Multiregion contact systems for modelling STD epidemics. Statistics in Medicine, 19(17-18):2479–2491, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pascutto:2000:SIA

[243] C. Pascutto, J. C. Wakefield, N. G. Best, S. Richardson, L. Bernardinelli, A. Staines, and P. Elliott. Statistical issues in the analysis of disease mapping data. Statistics in Medicine, 19(17-18):2493–2519, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carr:2000:TNT

[244] Daniel B. Carr, John F. Wallin, and D. Andrew Carr. Two new templates for epidemiology applications: linked micromap plots and conditioned choropleth maps. Statistics in Medicine, 19(17-18):2521–2538, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biggeri:2000:NPM

[245] Annibale Biggeri, Marco Marchi, Corrado Lagazio, Marco Martuzzi, and Dankmar Böhning. Non-parametric maximum likelihood estimators for disease mapping. Statistics in Medicine, 19(17-18):2539–2554, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knorr-Held:2000:BMI

[246] Leonhard Knorr-Held. Bayesian modelling of inseparable space-time variation in disease risk. Statistics in Medicine, 19(17-18):2555–2567, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heisterkamp:2000:AHI

[247] S. H. Heisterkamp, G. Doornbos, and N. J. D. Nagelkerke. Assessing health impact of environmental pollution sources using space-time models. Statistics in Medicine, 19(17-18):2569–2578, September 15–30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giudici:2000:MCC

[248] Paolo Giudici, Leonhard Knorr-Held, and Günter Rasser. Modelling categorical covariates in Bayesian disease mapping by partition structures. Statistics in Medicine, 19(17-18):2579–2593, September 15–30, 2000. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bonetti:2000:GMA

[249] Marco Bonetti and Richard D. Gelber. A graphical method to assess treatment-covariate interactions using the Cox model on subsets of the data. Statistics in Medicine, 19(19):2595–2609, October 15, 2000. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2000:SES

[250] Wei Pan. Smooth estimation of the survival function for interval censored data. Statistics in Medicine, 19(19):2611–2624, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2000:CAB

[251] Nancy R. Cook, Matthew W. Gillman, Bernard A. Rosner, James O. Taylor, and Charles H. Hennekens. Combining annual blood pressure measurements in childhood to improve prediction of young adult blood pressure. Statistics in Medicine, 19(19):2625–2640, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gomez:2000:RHI

[252] Guadalupe Gómez, M. Luz Calle, José M. Egea, and Robert Muga. Risk of HIV infection as a function of the duration of intravenous drug use: a non-parametric Bayesian approach. Statistics in Medicine, 19(19):2641– 2656, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qian:2000:AML

[253] W. Qian, M. K. B. Parmar, R. J. Sambrook, P. M. Fayers, D. J. Girling, and R. J. Stephens. Analysis of messy longitudinal data from a randomized clinical trial. Statistics in Medicine, 19(19):2657–2674, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Omar:2000:ACR

[254] Rumana Z. Omar and Simon G. Thompson. Analysis of a cluster randomized trial with binary outcome data using a multi-level model. Statistics in Medicine, 19(19):2675–2688, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sato:2000:SSC

[255] Tosiya Sato. Sample size calculations with compliance information. Statistics in Medicine, 19(19):2689–2697, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohen:2000:TWT

[256] Arthur Cohen, Harold B. Sackrowitz, and Marian Sackrowitz. Testing whether treatment is ‘better’ than control with ordered categorical data: an evaluation of new methodology. Statistics in Medicine, 19(19):2699– 2712, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Postelnicu:2000:BRM

[257] Tiberiu Postelnicu. Book review: Multivariate Taxometric Procedures. Niels G. Waller and Paul E. Meehl, Sage Publications, U.S.A., 1998. No. of pages: 150. ISBN 0-7619-0257-0. Statistics in Medicine, 19(19):2713– 2714, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gibberd:2000:BRG

[258] Robert Gibberd. Book review: Globocan 1: Cancer Incidence and Mortality Worldwide. J. Ferlay, D. M. Parkin and P. Pisani, IARC Press, Lyon, 1999. Price: $90. Statistics in Medicine, 19(19):2714–2715, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tomba:2000:BRD

[259] Gianpaolo Scalia Tomba. Book reviews: Deciphering Global Epidemics: Analytical Approaches to the Disease Records of World Cities, 1888– 1912. Andrew Cliff, Peter Haggett and Matthew Smallman-Raynor, Cambridge Studies in Historical Geography 26, Cambridge University Press, UK, 1998. No. of pages: xxiii + 469. Hardback: Price: £74.95. ISBN 0-521-47266-0. Paperback: Price: £18.95. ISBN 0-521-47860-X. Statistics in Medicine, 19(19):2715–2716, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rudemo:2000:BRS

[260] Mats Rudemo. Book review: Statistical Shape Analysis. I. L. Dryden and K. V. Mardia, Wiley, Chichester 1998. No. of pages: xvii + 347. Price: £60.00. ISBN 0-471-95816-6. Statistics in Medicine, 19(19):2716– 2717, October 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2000:P

[261] Shein-Chung Chow and Jen pei Liu. Preface. Statistics in Medicine, 19(20):2719, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2000:PAE

[262] A. Lawrence Gould. A practical approach for evaluating population and individual bioequivalence. Statistics in Medicine, 19(20):2721–2740, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2000:BPI

[263] Jun Shao, Shein-Chung Chow, and Bill Wang. The bootstrap procedure in individual bioequivalence. Statistics in Medicine, 19(20):2741–2754, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wellek:2000:RDC

[264] Stefan Wellek. On a reasonable disaggregate criterion of population bioequivalence admitting of resampling-free testing procedures. Statistics in Medicine, 19(20):2755–2767, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauschke:2000:UDG

[265] Dieter Hauschke and Volker W. Steinijans. The U.S. draft guidance regarding population and individual bioequivalence approaches: comments by a research-based pharmaceutical company. Statistics in Medicine, 19 (20):2769–2774, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimanani:2000:NME

[266] Ebi K. Kimanani, Jean Lavigne, and Diane Potvin. Numerical methods for the evaluation of individual bioequivalence criteria. Statistics in Medicine, 19(20):2775–2795, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimanani:2000:DIB

[267] Ebi K. Kimanani. Definition of individual bioequivalence: occasion-to-occasion versus mean switchability. Statistics in Medicine, 19(20):2797– 2810, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zariffa:2000:CSP

[268] Névine M.-D. Zariffa, S. D. Patterson, D. Boyle, and M. Hyneck. Case studies, practical issues and observations on population and individual bioequivalence. Statistics in Medicine, 19(20):2811–2820, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2000:IBC

[269] Mei-Ling Chen, Rabindra Patnaik, Walter W. Hauck, Donald J. Schuirmann, Terry Hyslop, and Roger Williams. An individual bioequivalence criterion: regulatory considerations. Statistics in Medicine, 19(20):2821–2842, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munk:2000:CBA

[270] Axel Munk. Connections between average and individual bioequivalence. Statistics in Medicine, 19(20):2843–2854, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauger:2000:AIA

[271] David T. Mauger and Vernon M. Chinchilli. An alternative index for assessing profile similarity in bioequivalence trials. Statistics in Medicine, 19(20):2855–2866, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Endrenyi:2000:PEV

[272] Laszlo Endrenyi, Nathan Taback, and Laszlo Tothfalusi. Properties of the estimated variance component for subject-by-formulation interaction in studies of individual bioequivalence. Statistics in Medicine, 19 (20):2867–2878, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsuan:2000:SSC

[273] Francis C. Hsuan. Some statistical considerations on the FDA draft guidance for individual bioequivalence. Statistics in Medicine, 19(20):2879– 2884, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hyslop:2000:SSC

[274] Terry Hyslop, Francis Hsuan, and Daniel J. Holder. A small sample confidence interval approach to assess individual bioequivalence. Statistics in Medicine, 19(20):2885–2897, October 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bolland:2000:FAS

[275] Kim Bolland and John Whitehead. Formal approaches to safety monitoring of clinical trials in life-threatening conditions. Statistics in Medicine, 19(21):2899–2917, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2000:THR

[276] Kung-Jong Lui and Colleen Kelly. Tests for homogeneity of the risk ratio in a series of 2 × 2 tables. Statistics in Medicine, 19(21):2919– 2932, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2000:CIR

[277] Kung-Jong Lui, Joni A. Mayer, and Laura Eckhardt. Confidence intervals for the risk ratio under cluster sampling based on the beta-binomial model. Statistics in Medicine, 19(21):2933–2942, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2000:GFS

[278] P. Royston and E. M. Wright. Goodness-of-fit statistics for age-specific reference intervals. Statistics in Medicine, 19(21):2943–2962, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [960].

Cai:2000:PTC

[279] Jianwen Cai and Yu Shen. Permutation tests for comparing marginal survival functions with clustered failure time data. Statistics in Medicine, 19(21):2963–2973, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2000:MLE

[280] Kishan G. Mehrotra, Pandurang M. Kulkarni, Ram C. Tripathi, and Joel E. Michalek. Maximum likelihood estimation for longitudinal data with truncated observations. Statistics in Medicine, 19(21):2975–2988, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2000:UAT

[281] John R. Cook and Joseph F. Heyse. Use of an angular transformation for ratio estimation in cost-effectiveness analysis. Statistics in Medicine, 19(21):2989–3003, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klein:2000:ECL

[282] John P. Klein, Richard M. Szydlo, Charles Craddock, and John M. Goldman. Estimation of current leukaemia-free survival following donor lymphocyte infusion therapy for patients with leukaemia who relapse after allografting: application of a multistate model. Statistics in Medicine, 19(21):3005–3016, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deBEdwardes:2000:LEM

[283] Michael D. deB. Edwardes. Letter to the Editor: M. D. deB. Edwardes, ‘The generalization of the odds ratio, risk ratio and risk difference to r × k tables’. Statistics in Medicine 2000; 19 (14): 1901–1914. Statistics in Medicine, 19(21):3017, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [193].

Tango:2000:BRS

[284] Toshiro Tango. Book review: 100 Statistical Tests. Gopal K. Kanji, Sage Publications, London, 1999. No. of pages: 215. Price: £49.50. ISBN 0-7619-6151-8. Statistics in Medicine, 19(21):3018, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oldford:2000:BRD

[285] R. W. Oldford. Book review: Data Structures for Computational Statistics. Sigbert Klinke, Physica-Verlag, Springer, Heidelberg, 1997. No. of pages: viii + 284. Price: $60.40. ISBN 3-7908-0982-9. Statistics in Medicine, 19(21):3019, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2000:BRR

[286] Chen-Hsin Chen. Book review: Research Methods — Guidance for Post-Graduates. Tony Greenfield (ed.), Arnold, 1996. No. of pages: xii + 306. Price: £14.99. ISBN 0-340-64629-2. Statistics in Medicine, 19(21):3020, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roehmel:2000:BRD

[287] Joachim Roehmel. Book review: Design and Analysis of Clinical Trials. Shein-Chung Chow and Jen-Pei Liu, Wiley, Chichester, 1998. No. of pages: xi + 649. Price: £80.95. ISBN 0-471-13404-X. Statistics in Medicine, 19(21):3021, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dirksen:2000:BRR

[288] Asger Dirksen. Book review: Revealing Statistical Principles. J. K. Lindsey, Arnold, London, 1999. No. of pages: 220. Price: £12.99. ISBN 0-340-74120-1. Statistics in Medicine, 19(21):3022, November 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernardo:2000:GSD

[289] M. V. Patricia Bernardo and Joseph G. Ibrahim. Group sequential designs for cure rate models with early stopping in favour of the null hypothesis. Statistics in Medicine, 19(22):3023–3035, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Corcoran:2000:PCT

[290] Chris Corcoran, Cyrus Mehta, and Pralay Senchaudhuri. Power comparisons for tests of trend in dose-response studies. Statistics in Medicine, 19(22):3037–3050, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2000:EST

[291] Nigel Stallard and Susan Todd. Exact sequential tests for single samples of discrete responses using spending functions. Statistics in Medicine, 19(22):3051–3064, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2000:CCR

[292] Peter B. Gilbert. Comparison of competing risks failure time methods and time-independent methods for assessing strain variations in vaccine protection. Statistics in Medicine, 19(22):3065–3086, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pinna:2000:RSM

[293] Gian Domenico Pinna, Cristina Opasich, Antonio Mazza, Antonio Tangenti, Roberto Maestri, and Maurizio Sanarico. Reproducibility of the six-minute walking test in chronic heart failure patients. Statistics in Medicine, 19(22):3087–3094, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:2000:EBA

[294] Ying-Hen Hsieh, Cathy W. S. Chen, and Shen-Ming Lee. Empirical Bayes approach to estimating the number of HIV-infected individuals in hidden and elusive populations. Statistics in Medicine, 19(22):3095– 3108, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:2000:SSP

[295] T. J. Cole. Sympercents: symmetric percentage differences on the 100 log_{e} scale simplify the presentation of log transformed data. Statistics in Medicine, 19(22):3109–3125, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [841].

Chinn:2000:SMC

[296] Susan Chinn. A simple method for converting an odds ratio to effect size for use in meta-analysis. Statistics in Medicine, 19(22):3127–3131, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julious:2000:LER

[297] Steven A. Julious. Letter to the Editor: Repeated measures in clinical trials: analysis using mean summary statistics and its implications for design by L. Frison and S. J. Pocock, Statistics in Medicine 1992; 12: 1685–1704. Statistics in Medicine, 19(22):3133–3135, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [5].

Skovgaard:2000:BRA

[298] L. T. Skovgaard. Book review: Applied regression analysis. 3rd edn. N. R. Draper and H. Smith, Wiley, New York, 1998. No. of pages: xvii + 706. Price: £45. ISBN 0-471-17082-8. Statistics in Medicine, 19(22): 3136–3139, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lumley:2000:BRK

[299] Thomas Lumley. Book review: Kendall’s advanced theory of statistics. Volume 2A: classical inference and the linear model. Alan Stuart, Keith Ord and Steven Arnold, Arnold, London, 1998, No. of pages: xiv + 885. Price: £85.00. ISBN 0-340-66230-1. Statistics in Medicine, 19(22):3139– 3140, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:2000:BRS

[300] Mark Little. Book review: A survey of models for tumor-immune system dynamics. J. A. Adam and N. Bellomo (eds), Birkhäuser, Boston, 1997. No. of pages: x + 344. Price: 138DM. ISBN 0-8176-3901-2. Statistics in Medicine, 19(22):3140–3141, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dalgaard:2000:BRD

[301] Peter Dalgaard. Book review: Dynamic modeling in the health sciences. James L. Hargrove, Springer-Verlag, New York, 1998. No. of pages: 298 (includes CD-ROM). Price: £37.50. ISBN 0-387-94996-8. Statistics in Medicine, 19(22):3142–3143, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:2000:BRB

[302] T. A. Louis. Book review: Bayesian statistics 6. J. M. Bernardo, J. O. Berger, A. P. Dawid and A. F. M. Smith (eds), Oxford Science Publications, Oxford, 1999. No. of pages: vii + 867. Price: £00.00. ISBN 0-19-850485-3. Statistics in Medicine, 19(22):3143–3144, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:2000:BRD

[303] Stephanie Green. Book review: Dictionary for clinical trials. Simon Day. Wiley, Chichester, U.K., 19. No. of pages: 250. Price: £19.99. ISBN 0-471-98611-9. Statistics in Medicine, 19(22):3144–3145, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:2000:P

[304] Stuart Pocock and Diana Elbourne. Preface. Statistics in Medicine, 19(23):3147, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horton:2000:CSF

[305] Richard Horton. Common sense and figures: the rhetoric of validity in medicine (Bradford Hill Memorial Lecture 1999). Statistics in Medicine, 19(23):3149–3164, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:2000:RSL

[306] Peter Armitage. Reflections on statistics at the London school of hygiene and tropical medicine 30 years ago. Statistics in Medicine, 19(23):3165– 3170, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:2000:ADI

[307] Michael D. Hughes. Analysis and design issues for studies using censored biomarker measurements with an example of viral load measurements in HIV clinical trials. Statistics in Medicine, 19(23):3171–3191, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robinson:2000:EPI

[308] Noah Jamie Robinson. Epidemiology in the pharmaceutical industry: a perspective for infectious diseases. Statistics in Medicine, 19(23):3193– 3198, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:2000:PSF

[309] Stephen J. W. Evans. Pharmacovigilance: a science or fielding emergencies? Statistics in Medicine, 19(23):3199–3209, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rockhold:2000:SUS

[310] Frank W. Rockhold. Strategic use of statistical thinking in drug development. Statistics in Medicine, 19(23):3211–3217, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barber:2000:ACD

[311] Julie A. Barber and Simon G. Thompson. Analysis of cost data in randomized trials: an application of the non-parametric bootstrap. Statistics in Medicine, 19(23):3219–3236, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ribaudo:2000:JAQ

[312] Heather J. Ribaudo, Simon G. Thompson, and Timothy G. Allen-Mersh. A joint analysis of quality of life and survival using a random effect selection model. Statistics in Medicine, 19(23):3237–3250, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sharp:2000:ARB

[313] Stephen J. Sharp and Simon G. Thompson. Analysing the relationship between treatment effect and underlying risk in meta-analysis: comparison and development of approaches. Statistics in Medicine, 19(23):3251– 3274, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:2000:SMJ

[314] Douglas G. Altman. Statistics in medical journals: some recent trends. Statistics in Medicine, 19(23):3275–3289, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [537].

Ashby:2000:EBM

[315] Deborah Ashby and Adrian F. M. Smith. Evidence-based medicine as Bayesian decision-making. Statistics in Medicine, 19(23):3291–3305, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2000:P

[316] Ralph D’Agostino, David Machin, and Chris Palmer. Preface. Statistics in Medicine, 19(24):3307–3308, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vaida:2000:PHM

[317] Florin Vaida and Ronghui Xu. Proportional hazards model with random effects. Statistics in Medicine, 19(24):3309–3324, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hahn:2000:APB

[318] Seokyung Hahn, Paula R. Williamson, Jane L. Hutton, Paul Garner, and E. Victor Flynn. Assessing the potential for bias in meta-analysis due to selective reporting of subgroup analyses within studies. Statistics in Medicine, 19(24):3325–3336, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horvath:2000:FDR

[319] Stefan Horvath and Max P. Baur. Future directions of research in statistical genetics. Statistics in Medicine, 19(24):3337–3343, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ziegler:2000:FAL

[320] Andreas Ziegler, Christian Kastner, Daniel Brunner, and Maria Blettner. Familial associations of lipid profiles: a generalized estimating equations approach. Statistics in Medicine, 19(24):3345–3357, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prevost:2000:HMG

[321] Teresa C. Prevost, Keith R. Abrams, and David R. Jones. Hierarchical models in generalized synthesis of evidence: an example based on studies of breast cancer screening. Statistics in Medicine, 19(24):3359–3376, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Santos:2000:CRR

[322] Dirley M. Dos Santos and Damon M. Berridge. A continuation ratio random effects model for repeated ordinal responses. Statistics in Medicine, 19(24):3377–3388, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2000:SPI

[323] Julia Brown, Damien McElvenny, Jane Nixon, Jackie Bainbridge, and Su Mason. Some practical issues in the design, monitoring and analysis of a sequential randomized trial in pressure sore prevention. Statistics in Medicine, 19(24):3389–3400, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanHouwelingen:2000:VCR

[324] Hans C. van Houwelingen. Validation, calibration, revision and combination of prognostic survival models. Statistics in Medicine, 19(24): 3401–3415, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2000:MMF

[325] Rebecca M. Turner, Rumana Z. Omar, Min Yang, Harvey Goldstein, and Simon G. Thompson. A multilevel model framework for meta-analysis of clinical trials with binary outcomes. Statistics in Medicine, 19(24): 3417–3432, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zwinderman:2000:DSH

[326] Aeilko H. Zwinderman, Alexandre E. Voskuyl, David D. Schelhaas, Sjoerd G. van Duinen, Johanna M. van der Bas, and Johanna M. H. Hazes. Diagnostic strategies for the histological examination of muscle biopsy specimens for the assessment of vasculitis in rheumatoid arthritis. Statistics in Medicine, 19(24):3433–3447, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

derTweel:2000:SAM

[327] Ingeborg Van der Tweel and Paulus A. H. Van Noord. Sequential analysis of matched dichotomous data from prospective case-control studies. Statistics in Medicine, 19(24):3449–3464, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Escolano:2000:MSM

[328] Sylvie Escolano, Jean-Louis Golmard, Anne-Marie Korinek, and Alain Mallet. A multi-state model for evolution of intensive care unit patients: prediction of nosocomial infections and deaths. Statistics in Medicine, 19 (24):3465–3482, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Svensson:2000:CBR

[329] Elisabeth Svensson. Concordance between ratings using different scales for the same variable. Statistics in Medicine, 19(24):3483–3496, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arends:2000:BRP

[330] Lidia R. Arends, Arno W. Hoes, Jacobus Lubsen, Diederik E. Grobbee, and Theo Stijnen. Baseline risk as predictor of treatment benefit: three clinical meta-re-analyses. Statistics in Medicine, 19(24):3497–3518, December 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hunsberger:2001:TTI

[331] Sally Hunsberger and Dean A. Follmann. Testing for treatment and interaction effects in semi-parametric analysis of covariance. Statistics in Medicine, 20(1):1–19, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carey:2001:ALO

[332] Vincent J. Carey and Bernard A. Rosner. Analysis of longitudinally observed irregularly timed multivariate outcomes: regression with focus on cross-component correlation. Statistics in Medicine, 20(1):21–31, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lynn:2001:MLI

[333] Henry S. Lynn. Maximum likelihood inference for left-censored HIV RNA data. Statistics in Medicine, 20(1):33–45, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rostgaard:2001:APC

[334] Klaus Rostgaard, Michael Væth, Helle Holst, Mette Madsen, and Elsebeth Lynge. Age-period-cohort modelling of breast cancer incidence in the Nordic countries. Statistics in Medicine, 20(1):47–61, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2001:SAR

[335] Jianguo Sun and Peter X.-K. Song. Statistical analysis of repeated measurements with informative censoring times. Statistics in Medicine, 20 (1):63–73, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kowalski:2001:DEP

[336] Kenneth G. Kowalski and Matthew M. Hutmacher. Design evaluation for a population pharmacokinetic study using clinical trial simulations: a case study. Statistics in Medicine, 20(1):75–91, January 15, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2001:ADC

[337] Margaret C. Wu, Paul S. Albert, and Bechien U. Wu. Adjusting for drop-out in clinical trials with repeated measures: design and analysis issues. Statistics in Medicine, 20(1):93–108, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perez-Ocon:2001:PMP

[338] Rafael Pérez-Ocón, Juan Eloy Ruiz-Castro, and M. Luz Gámiz-Pérez. A piecewise Markov process for analysing survival from breast cancer in different risk groups. Statistics in Medicine, 20(1):109–122, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhu:2001:BOD

[339] Wei Zhu and Weng Kee Wong. Bayesian optimal designs for estimating a set of symmetrical quantiles. Statistics in Medicine, 20(1):123–137, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelman:2001:ERC

[340] Donna Spiegelman, Raymond J. Carroll, and Victor Kipnis. Efficient regression calibration for logistic regression in main study/internal validation study designs with an imperfect reference instrument. Statistics in Medicine, 20(1):139–160, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2001:BRE

[341] Per Kragh Andersen. Book review: Epidemiology. Study Design and Data Analysis. Mark Woodward. Chapman and Hall/CRC, London, 1999. No. of pages: 699. Price: £79.95. ISBN 0-8493-0398. Statistics in Medicine, 20(1):161–162, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hill:2001:BRC

[342] Catherine Hill. Book review: Clinical Trials in Oncology. Stephanie Green, Jacqueline Benedetti and John Crowley, Chapman and Hall, London, U.K., 1997. No. of pages: 203. Price: £48.95. ISBN 0-412-99631-6. Statistics in Medicine, 20(1):162–163, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eichler:2001:BRM

[343] Michael Eichler. Book review: Modelldiagnose in Der Bayesschen Inferenz. Reinhard Vontheim, Lang, Frankfurt am Main, 1999. No. of pages:

149. Price: DM 54. ISBN 3-631-34610-7. Statistics in Medicine, 20(1): 163–164, January 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2001:MTD

[344] John Whitehead, Anne Whitehead, Susan Todd, Kim Bolland, and M. Roshini Sooriyarachchi. Mid-trial design reviews for sequential clinical trials. Statistics in Medicine, 20(2):165–176, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marschner:2001:IMC

[345] Ian C. Marschner and Simone L. Becker. Interim monitoring of clinical trials based on long-term binary endpoints. Statistics in Medicine, 20(2): 177–192, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Potthoff:2001:DTC

[346] Richard F. Potthoff, Bercedis L. Peterson, and Stephen L. George. Detecting treatment-by-centre interaction in multi-centre clinical trials. Statistics in Medicine, 20(2):193–213, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stine:2001:NPE

[347] Robert A. Stine and Joseph F. Heyse. Non-parametric estimates of overlap. Statistics in Medicine, 20(2):215–236, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molinari:2001:RST

[348] Nicolas Molinari, Jean-Pierre Daurès, and Jean-François Durand. Regression splines for threshold selection in survival data analysis. Statistics in Medicine, 20(2):237–247, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartfay:2001:SIT

[349] E. Bartfay and A. Donner. Statistical inferences for a twin correlation with multinomial outcomes. Statistics in Medicine, 20(2):249–262, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2001:IRS

[350] Peter B. Gilbert. Interpretability and robustness of sieve analysis models for assessing HIV strain variations in vaccine efficacy. Statistics in Medicine, 20(2):263–279, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2001:SSI

[351] Yunchan Chi. Simulation study and implementation of the tests based on weighted Turnbull’s estimators for interval-censored data. Statistics in Medicine, 20(2):281–294, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailey:2001:ORA

[352] Norman T. J. Bailey. An operational research approach to estimating the age distribution of residence time in an HIV/AIDS core group. Statistics in Medicine, 20(2):295–303, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hunsberger:2001:ISM

[353] Sally Hunsberger, David Murray, C. Ed Davis, and Richard R. Fabsitz. Imputation strategies for missing data in a school-based multi-centre study: the pathways study. Statistics in Medicine, 20(2):305–316, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tamhane:2001:LEM

[354] A. C. Tamhane and L. A. Hothorn. Letter to the Editor: A multiple comparison procedure for three- and four-armed controlled clinical trials by M. A. Proschan, Statistics in Medicine 1999; 18:787–798. Statistics in Medicine, 20(2):317–318, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [16].

Bauer:2001:LEM

[355] Peter Bauer. Letter to the Editor: A multiple comparison procedure for three- and four-armed controlled clinical trials by M. A. Proschan, Statistics in Medicine 1999; 18:787–798. Statistics in Medicine, 20(2):

319–320, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [16].

Proschan:2001:AR

[356] Michael A. Proschan. Author’s reply. Statistics in Medicine, 20(2):320– 321, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mortensen:2001:BRD

[357] Erik Lykke Mortensen. Book review: A Dictionary of Statistics for Psychologists. Brian S. Everitt and Til Wykes, Arnold, London, 1999. No. of pages: vi + 188. Price: £40.00. ISBN 0-340-71997-4. Statistics in Medicine, 20(2):322–323, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Endahl:2001:BRA

[358] Lars A. Endahl. Book review: Applied Mixed Models in Medicine. Helen Brown and Robin Prescott, Wiley, Chichester, 1999. No. of pages: 408. Price: £45. ISBN 0-471-96554-5. Statistics in Medicine, 20(2):323–324, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Metcalfe:2001:BRB

[359] Chris Metcalfe. Book review: Biostatistics: a Foundation for Analysis in the Health Sciences. 7th edn. Wayne W. Daniel, Wiley, 1999. No. of. pages: xiv + 755 + appendices. Price: £28.95. ISBN 0-471-16386-4. Statistics in Medicine, 20(2):324–326, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bock:2001:BRC

[360] Hans-Hermann Bock. Book review: Construction and Assessment of Classification Rules. David J. Hand, Wiley, Chichester, 1997. No. of pages: xii + 214. Price: £34.95. ISBN 0-471-96583-9. Statistics in Medicine, 20(2):326–327, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Christensen:2001:BRL

[361] Karl Bang Christensen. Book review: Latent Variable Models and Factor Analysis. Kendall’s Library of Statistics 7, 2nd Edn. David Bartholomew and Martin Knott, Arnold, London, 1999. No. of pages: 224. Price: £35.00. ISBN 0-340-69243-X. Statistics in Medicine, 20(2):327–328, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dietz:2001:BRS

[362] Klaus Dietz. Book review: Statistical Aspects of BSE and vCJD: Models for Epidemics. Christl A. Donnelly and Neil M. Ferguson, Chapman and Hall/CRC, Boca Raton, 1999. No. of pages: v + 229. Price: £39.00. ISBN 0-8493-0386-9. Statistics in Medicine, 20(2):328, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:2001:P

[363] Michael J. Campbell, Allan Donner, and Diana R. Elbourne. Preface. Statistics in Medicine, 20(3):329–330, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moore:2001:DFP

[364] Helen Moore, Carolyn D. Summerbell, Andy Vail, Darren C. Greenwood, and Ashley J. Adamson. The design features and practicalities of conducting a pragmatic cluster randomized trial of obesity management in primary care. Statistics in Medicine, 20(3):331–340, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yudkin:2001:PTP

[365] P. L. Yudkin and M. Moher. Putting theory into practice: a cluster randomized trial with a small number of clusters. Statistics in Medicine, 20(3):341–349, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raab:2001:BCR

[366] Gillian M. Raab and Izzy Butcher. Balance in cluster randomized trials. Statistics in Medicine, 20(3):351–365, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eldridge:2001:SSC

[367] Sandra Eldridge, Colin Cryer, Gene Feder, and Martin Underwood. Sample size calculations for intervention trials in primary care randomizing by primary care group: an empirical illustration from one proposed intervention trial. Statistics in Medicine, 20(3):367–376, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kerry:2001:UCS

[368] Sally M. Kerry and J. Martin Bland. Unequal cluster sizes for trials in English and Welsh general practice: implications for sample size calculations. Statistics in Medicine, 20(3):377–390, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:2001:CTI

[369] Marion K. Campbell, Jill Mollison, and Jeremy M. Grimshaw. Cluster trials in implementation research: estimation of intracluster correlation coefficients and sample size. Statistics in Medicine, 20(3):391–399, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piaggio:2001:MCD

[370] Gilda Piaggio, Guillermo Carroli, José Villar, Alain Pinol, Leiv Bakketeig, Pisake Lumbiganon, Per Bergsjø, Yagob Al-Mazrou, Hassan Ba’aqeel, José Miguel Belizán, Ubaldo Farnot, and Heinz Berendes. Methodological considerations on the design and analysis of an equivalence stratified cluster randomization trial. Statistics in Medicine, 20 (3):401–416, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ukoumunne:2001:ACR

[371] Obioha C. Ukoumunne and Simon G. Thompson. Analysis of cluster randomized trials with repeated cross-sectional binary measurements. Statistics in Medicine, 20(3):417–433, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelhalter:2001:BMC

[372] David J. Spiegelhalter. Bayesian methods for cluster randomized trials with continuous responses. Statistics in Medicine, 20(3):435–452, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2001:BMA

[373] Rebecca M. Turner, Rumana Z. Omar, and Simon G. Thompson. Bayesian methods of analysis for cluster randomized trials with binary outcome data. Statistics in Medicine, 20(3):453–472, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutton:2001:DEP

[374] J. L. Hutton. Are distinctive ethical principles required for cluster randomized controlled trials? Statistics in Medicine, 20(3):473–488, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elbourne:2001:ECS

[375] Diana R. Elbourne and Marion K. Campbell. Extending the CON-SORT statement to cluster randomized trials: for discussion. Statistics in Medicine, 20(3):489–496, February 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shun:2001:TES

[376] Zhenming Shun, William Yuan, William E. Brady, and Huang Hsu. Type I error in sample size re-estimations based on observed treatment difference. Statistics in Medicine, 20(4):497–513, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2001:SSR

[377] Weichung Joseph Shih. Sample size re-estimation — journey for a decade. Statistics in Medicine, 20(4):515–518, February 28, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shun:2001:R

[378] Zhenming Shun, William Yuan, William E. Brady, and Huang Hsu. Rejoinder. Statistics in Medicine, 20(4):519–520, February 28, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bouckaert:2001:SOR

[379] A. Bouckaert and M. Mouchart. Sure outcomes of random events: a model for clinical trials. Statistics in Medicine, 20(4):521–543, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marcus:2001:SAS

[380] Sue M. Marcus. A sensitivity analysis for subverting randomization in controlled trials. Statistics in Medicine, 20(4):545–555, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernardo:2001:SSC

[381] M. V. Patricia Bernardo and David P. Harrington. Sample size calculations for the two-sample problem using the multiplicative intensity model. Statistics in Medicine, 20(4):557–579, February 28, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chaloner:2001:QDP

[382] Kathryn Chaloner and Frank S. Rhame. Quantifying and documenting prior beliefs in clinical trials. Statistics in Medicine, 20(4):581–600, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pulkstenis:2001:MEM

[383] Erik Pulkstenis, Thomas R. Ten Have, and J. Richard Landis. A mixed effects model for the analysis of ordinal longitudinal pain data subject to informative drop-out. Statistics in Medicine, 20(4):601–622, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reboussin:2001:LCM

[384] Beth A. Reboussin and James C. Anthony. Latent class marginal regression models for modelling youthful drug involvement and its suspected influences. Statistics in Medicine, 20(4):623–639, February 28, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Macaskill:2001:CMD

[385] Petra Macaskill, Stephen D. Walter, and Les Irwig. A comparison of methods to detect publication bias in meta-analysis. Statistics in Medicine, 20(4):641–654, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [803].

Kim:2001:CKH

[386] Hyune-Ju Kim. Correction: Kim H-J, Fay MP, Feuer EJ, Midthune DN, ‘Permutation tests for joinpoint regression with applications to cancer rates’. Statistics in Medicine 2000 19:335–351. Statistics in Medicine, 20 (4):655, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [48].

Hadgu:2001:LEU

[387] Alula Hadgu and William Miller. Letter to the Editor: Using a combination of reference tests to assess the accuracy of a diagnostic test by A. Alonzo and M. Pepe, Statistics in Medicine 1999; 18: 2987–3003. Statistics in Medicine, 20(4):656–658, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [18].

Pepe:2001:AR

[388] Margaret Sullivan Pepe and Todd A. Alonzo. Authors’ reply. Statistics in Medicine, 20(4):658–660, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2001:CRAa

[389] Bee-Choo Tai, David Machin, Ian White, and Val Gebski. Competing risks analysis of patients with osteosarcoma: a comparison of four different approaches. Statistics in Medicine, 20(5):661–684, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tilling:2001:MGC

[390] Kate Tilling, Jonathan A. C. Sterne, and Charles D. A. Wolfe. Multilevel growth curve models with covariate effects: application to recovery after stroke. Statistics in Medicine, 20(5):685–704, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weinberg:2001:ECR

[391] Janice M. Weinberg and Stephen W. Lagakos. Efficiency comparisons of rank and permutation tests based on summary statistics computed from repeated measures data. Statistics in Medicine, 20(5):705–731, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1337].

OHagan:2001:BCE

[392] Anthony O’Hagan, John W. Stevens, and Jacques Montmartin. Bayesian cost-effectiveness analysis from clinical trial data. Statistics in Medicine, 20(5):733–753, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Erkanli:2001:BAL

[393] Alaattin Erkanli, Refik Soyer, and Adrian Angold. Bayesian analyses of longitudinal binary data using Markov regression models of unknown order. Statistics in Medicine, 20(5):755–770, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Veierod:2001:EMB

[394] Marit B. Veierød and Petter Laake. Exposure misclassification: bias in category specific Poisson regression coefficients. Statistics in Medicine, 20(5):771–784, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2001:TCT

[395] Sin-Ho Jung and Seung-Ho Kang. Tests for 2 × K contingency tables with clustered ordered categorical data. Statistics in Medicine, 20(5): 785–794, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geskus:2001:MEA

[396] Ronald B. Geskus. Methods for estimating the AIDS incubation time distribution when date of seroconversion is censored. Statistics in Medicine, 20(5):795–812, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taves:2001:LEO

[397] Donald R. Taves. Letter to the Editor: Optimum biased-coin designs for sequential treatment allocation with covariate information by A. C. Atkinson, Statistics in Medicine 1999; 18:1741–1752. Statistics in Medicine, 20(5):813–816, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [17].

Atkinson:2001:AR

[398] Anthony C. Atkinson. Authors’ reply. Statistics in Medicine, 20(5): 816–818, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foldspang:2001:BRR

[399] Anders Foldspang. Book review: Research Methods in Health. Investigating Health and Health Services. Ann Bowling, Open University Press, Milton Keynes, 1997. No. of pages: xv + 430. Price: £16.99. ISBN 0-335-19885-6. Statistics in Medicine, 20(5):819, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skovgaard:2001:BRA

[400] L. T. Skovgaard. Book review: Applied Regression Including Computing and Graphics. R. D. Cook and S. Weisberg, Wiley, New York, 1999. No. of pages: xxvi + 593. Price: £58.50. ISBN 0-471-31711-X. Statistics in Medicine, 20(5):820–821, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wall:2001:BRC

[401] Stig Wall. Book review: Cancer Epidemiology: Principles and Methods. Isabel dos Santos Silva. IARC Press, Lyon, France, 1999. No. of pages: ix + 442. Price: £40.00 ISBN 92-832-0405-0. Statistics in Medicine, 20(5):

821–822, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Denne:2001:BRS

[402] Jon Denne. Book review: Statistical Analysis of Categorical Data. Chris Lloyd, Wiley, 1999. No. of pages: xii + 468. Price: £70.00. ISBN 0-471-29008-4. Statistics in Medicine, 20(5):822–824, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brockwell:2001:CSM

[403] Sarah E. Brockwell and Ian R. Gordon. A comparison of statistical methods for meta-analysis. Statistics in Medicine, 20(6):825–840, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halpern:2001:SSC

[404] Jerry Halpern, Byron Wm. Brown, Jr., and John Hornberger. The sample size for a clinical trial: a Bayesian-decision theoretic approach. Statistics in Medicine, 20(6):841–858, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

AHern:2001:SST

[405] R. P. A’Hern. Sample size tables for exact single-stage phase II designs. Statistics in Medicine, 20(6):859–866, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Legedza:2001:HPC

[406] Anna T. R. Legedza and Joseph G. Ibrahim. Heterogeneity in phase I clinical trials: prior elicitation and computation using the continual reassessment method. Statistics in Medicine, 20(6):867–882, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mazumdar:2001:SMA

[407] Madhu Mazumdar, Melissa Fazzari, and Katherine S. Panageas. A standardization method to adjust for the effect of patient selection in phase II clinical trials. Statistics in Medicine, 20(6):883–892, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2001:ENN

[408] S. D. Walter and L. Irwig. Estimating the number needed to treat (NNT) index when the data are subject to error. Statistics in Medicine, 20(6):

893–906, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:2001:SCS

[409] Robert G. Newcombe. Simultaneous comparison of sensitivity and specificity of two tests in the paired design: a straightforward graphical approach. Statistics in Medicine, 20(6):907–915, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Downer:2001:TED

[410] Robert G. Downer. Testing for elevated disease rates using smoothed estimates. Statistics in Medicine, 20(6):917–933, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sasieni:2001:EPW

[411] Peter Sasieni. Estimating prevalence when the true disease status is incompletely ascertained. Statistics in Medicine, 20(6):935–949, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Renard:2001:RMM

[412] Didier Renard, Liesbeth Bruckers, Geert Molenberghs, Akke Vellinga, and Pierre Van Damme. Repeated-measures models to evaluate a hepatitis B vaccination programme. Statistics in Medicine, 20(6):951–963, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunson:2001:AHF

[413] D. B. Dunson, C. R. Weinberg, D. D. Baird, J. S. Kesner, and A. J. Wilcox. Assessing human fertility using several markers of ovulation. Statistics in Medicine, 20(6):965–978, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2001:LEM

[414] Tuhao Chen. Letter to the Editor: Methods of Adjustment for Estimating the Attributable Risk in Case-Control Studies: a Review, by Jacques Benichou, Statistics in Medicine 1991; 10:1753–1773. Statistics in Medicine, 20(6):979–981, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3].

Benichou:2001:AR

[415] Jacques Benichou. Author’s reply. Statistics in Medicine, 20(6):981–982, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:2001:LEN

[416] Nico Nagelkerke. Letter to the Editor: Nagelkerke N, Fidler V, Bernsen R, Borgdorff M, ‘Estimating treatment effects in randomized clinical trials in the presence of non-compliance’, Statistics in Medicine 2000; 19:1849–1864. Statistics in Medicine, 20(6):982, March 30, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [189].

Ickstadt:2001:BRD

[417] Katja Ickstadt. Book review: Disease Mapping and Risk Assessment for Public Health. Andrew Lawson, Annibale Biggeri, Dankmar Böhning, Emmanuel Lesaffre, Jean-François Viel and Roberto Bertollini (eds), Wiley, Chichester, 1999. No. of pages: xix + 482. Price: £60. ISBN 0-471-98634-8. Statistics in Medicine, 20(6):983–984, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skovlund:2001:BRD

[418] Eva Skovlund. Book review: The Design and Analysis of Clinical Experiments. J. L. Fleiss, Wiley Classics Library, Chichester, U.K., 1999. No. of pages: xiv + 432. Price: £32.50. ISBN 0-471-34991-7. Statistics in Medicine, 20(6):984–985, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Plummer:2001:BRP

[419] Martyn Plummer. Book review: Programming with Data: a Guide to the S Language. John M. Chambers, Springer, Godalming, U.K., 1998. No. of pages:xv + 469. Price: £32.50. ISBN 0-387-98503-4. Statistics in Medicine, 20(6):985–986, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasser:2001:BRN

[420] Theo Gasser. Book review: The National Study of Health and Growth. Roberto J. Rona and Susan Chinn, Oxford University Press, 1999. No. of pages: viii + 133. Price: £22.50. ISBN 0-19-262919-0. Statistics in Medicine, 20(6):986–987, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bacelar-Nicolau:2001:BRM

[421] Helena Bacelar-Nicolau. Book review: Mathematical Classification and Clustering. Boris Mirkin, Kluwer Academic Publishers, 1996. No. of Pages: xv + 428. Price: £140. ISBN 0-7923-4159-7. Statistics in Medicine, 20(6):987–988, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schluchter:2001:ACP

[422] Mark D. Schluchter, Tom Greene, and Gerald J. Beck. Analysis of change in the presence of informative censoring: application to a longitudinal clinical trial of progressive renal disease. Statistics in Medicine, 20(7): 989–1007, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fitzmaurice:2001:APG

[423] Garrett M. Fitzmaurice, Nan M. Laird, and Lucy Shneyer. An alternative parameterization of the general linear mixture model for longitudinal data with non-ignorable drop-outs. Statistics in Medicine, 20(7):1009– 1021, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:2001:OTS

[424] Peter F. Thall and Su-Chun Cheng. Optimal two-stage designs for clinical trials based on safety and efficacy. Statistics in Medicine, 20(7):1023– 1032, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yafune:2001:MPS

[425] Akifumi Yafune and Makio Ishiguro. Modification of pharmacokinetic sampling schedules in clinical phase I trials: a use of Kullback–Leibler information. Statistics in Medicine, 20(7):1033–1049, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sargent:2001:FDM

[426] Daniel J. Sargent and Richard M. Goldberg. A flexible design for multiple armed screening trials. Statistics in Medicine, 20(7):1051–1060, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [714].

vanBuuren:2001:RIS

[427] S. van Buuren and M. Hopman-Rock. Revision of the ICIDH severity of disabilities scale by data linking and item response theory. Statistics in Medicine, 20(7):1061–1076, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2001:DCD

[428] Roger J. Marshall. Displaying clinical data relationships using scaled rectangle diagrams. Statistics in Medicine, 20(7):1077–1088, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Singer:2001:SPC

[429] Júlia Singer. A simple procedure to compute the sample size needed to compare two independent groups when the population variances are unequal. Statistics in Medicine, 20(7):1089–1095, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:2001:ECH

[430] B. A. Davis, C. R. Heathcote, and T. J. O’Neill. Estimating cohort health expectancies from cross-sectional surveys of disability. Statistics in Medicine, 20(7):1097–1111, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:2001:LDM

[431] Nico J. D. Nagelkerke, Martien W. Borgdorff, and Sang Jae Kim. Logistic discrimination of mixtures of M. tuberculosis and non-specific tuberculin reactions. Statistics in Medicine, 20(7):1113–1124, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2001:IED

[432] Meihui Guo, Mong-Na Lo Huang, Zhidong Bai, and Kai-Sheng Hsieh. Important ECG diagnosis-aiding indices of ventricular septal defect children with or without congestive heart failure. Statistics in Medicine, 20(7):1125–1141, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verbeke:2001:BRR

[433] Geert Verbeke. Book review: Regression Analysis of Count Data. A. C. Cameron and P. K. Trivedi (eds.), Cambridge University Press, Cambridge, 1998. No. of pages: xv + 404. Price: £ 17.95. ISBN 0-512-63567-5. Statistics in Medicine, 20(7):1143–1144, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Budtz-Jorgensen:2001:BRM

[434] Esben Budtz-Jørgensen. Book review: Model Selection and Inference. A Practical Information-Theoretic Approach. Kenneth P. Burnham and David R. Anderson, Springer-Verlag, Heidelberg, 1998. No. of pages: xix + 353. Price: DM 149.00. ISBN 0-387-98504-2. Statistics in Medicine, 20(7):1144–1146, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kochar:2001:BRT

[435] Subhash Kochar. Book review: Theory of Rank Tests. Jaroslav Hájek, Zbynek Sidák and Pranab K. Sen, Academic Press, Kent, U.K., 1999. No. of pages: vii + 437. Price: £ 39.95. ISBN 0-1264-2350-4. Statistics in Medicine, 20(7):1146–1147, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacobsen:2001:BRK

[436] Martin Jacobsen. Book review: Kendall’s Library of Statistics 8. Statistical Inference for Diffusion Type Processes. B. L. S. Prakasa Rao, Arnold, Euston, U.K., 1999. No. of pages: xvi + 349. Price: £ 55.00. ISBN 0-340-74149X. Statistics in Medicine, 20(7):1147–1148, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmiegelow:2001:BRI

[437] Kjeld Schmiegelow. Book reviews: International Incidence of Childhood Cancer, Vol. II. D. M. Parkin, E. Kramárová, G. J. Draper, E. Masuyer, J. Michaelis, J. Neglia, S. Qureshi and C. A. Stiller, IARC Scientific Publications No. 144, Oxford, U.K., 1999. No. of pages: xiii + 391. Price: £41.00. ISBN 92832-2144-3. Epidemiology of Childhood Cancer. Julian Little, IARC Scientific Publications No. 149, Oxford, U.K., 1999. No. of pages: xiv + 385. Price: £36.00. ISBN 92-832-2149-4. Statistics in Medicine, 20(7):1148–1149, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jockel:2001:BRA

[438] Karl-Heinz Jöckel. Book review: Analysis of Health Surveys. Edward L. Korn and Barry I. Graubard, Wiley, New York, 1999. No. of pages:

382. Price: £ 50.50. ISBN 0-471-13773-1. Statistics in Medicine, 20(7): 1149–1150, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holst:2001:BRS

[439] Helle Holst. Book review: Statistics for Environmental Biology and Toxicology. Walter W. Piegorsch and John Bailer, Chapman and Hall, London, 1997. No. of pages: 600. Price: £ 45.95. ISBN 0-412-04731-4. Statistics in Medicine, 20(7):1150–1151, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nilsson:2001:BRM

[440] Jens C. Nilsson and Bjørn A. Grønning. Book review: Multivariable Analysis. A Practical Guide for Clinicians. Mitchell H. Katz, Cambridge University Press, Cambridge, 1999. No. of. pages: xv + 192. Price: £ 17.95. ISBN 0-521-59693-9. Statistics in Medicine, 20(7):1151–1152, April 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meinert:2001:EGB

[441] Curtis L. Meinert and Adele Kaplan Gilpin. Estimation of gender bias in clinical trials. Statistics in Medicine, 20(8):1153–1164, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McMahon:2001:JTM

[442] Robert P. McMahon and Frank E. Harrell, Jr. Joint testing of mortality and a non-fatal outcome in clinical trials. Statistics in Medicine, 20(8): 1165–1172, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rochon:2001:MAR

[443] James Rochon and Brenda W. Gillespie. A methodology for analysing a repeated measures and survival outcome simultaneously. Statistics in Medicine, 20(8):1173–1184, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Correa:2001:PCR

[444] José A. Correa and François Bellavance. Power comparison of robust approximate and non-parametric tests for the analysis of cross-over trials. Statistics in Medicine, 20(8):1185–1196, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [813].

Wei:2001:PIA

[445] Lynn Wei and Weichung J. Shih. Partial imputation approach to analysis of repeated measurements with dependent drop-outs. Statistics in Medicine, 20(8):1197–1214, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:2001:CTP

[446] Peter A. Lachenbruch. Comparisons of two-part models with competitors. Statistics in Medicine, 20(8):1215–1234, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:2001:PSS

[447] Peter A. Lachenbruch. Power and sample size requirements for two-part models. Statistics in Medicine, 20(8):1235–1238, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2001:TPS

[448] Jun Shao and Shein-Chung Chow. Two-phase shelf-life estimation. Statistics in Medicine, 20(8):1239–1248, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2001:VES

[449] Jianguo Sun. Variance estimation of a survival function for intervalcensored survival data. Statistics in Medicine, 20(8):1249–1257, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBuuren:2001:WPS

[450] Stef van Buuren and Miranda Fredriks. Worm plot: a simple diagnostic device for modelling growth reference curves. Statistics in Medicine, 20 (8):1259–1277, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:2001:SCE

[451] Eugene M. Laska, Morris Meisner, Carole Siegel, and Joseph Wanderling. Statistical cost-effectiveness analysis of two treatments based on net health benefits. Statistics in Medicine, 20(8):1279–1302, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Helmers:2001:BRB

[452] Roelof Helmers. Book review: Bootstrap Methods. A Practitioner’s Guide. Michael R. Chernick, Wiley, Chichester, 1999. No. of pages: 280. Price: £48.50. ISBN 0471-34912-7. Statistics in Medicine, 20(8):1303– 1304, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kessing:2001:BRS

[453] Lars Vedel Kessing. Book review: Statistics in Psychiatry. Graham Dunn, Arnold, U.K., 1999. No. of pages: viii + 132. Price: £24.99. ISBN 034-067668. Statistics in Medicine, 20(8):1304, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Azzalini:2001:BRS

[454] Adelchi Azzalini. Book review: Statistical Inference: an Integrated Approach. Helio S. Migon and Dani Gamerman, Arnold, London, 1999. No. of pages: ix + 262. Price: £24.99. ISBN 0-340-740590. Statistics in Medicine, 20(8):1305, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bjorner:2001:BRL

[455] Jakob B. Bjorner. Book review: Latent Class Scaling Analysis. C. Mitchell Dayton, Sage University Papers Series on Quantitative Applications in the Social Sciences, 07-126, Sage, Thousand Oaks, 1998. No. of pages: vii + 95. Price: $13.95. ISBN 0-7619-1323-8. Statistics in Medicine, 20(8):1305–1306, April 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sieber:2001:P

[456] W. Karl Sieber. Preface. Statistics in Medicine, 20(9-10):1307–1308, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2001:LSC

[457] Anonymous. List of staff and committees. Statistics in Medicine, 20(9-10):1309–1314, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2001:PD

[458] Anonymous. Presenters and discussants. Statistics in Medicine, 20(9-10):1315–1320, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Speers:2001:WOR

[459] M. A. Speers. Welcome and opening remarks. Statistics in Medicine, 20 (9-10):1321–1322, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Braithwaite:2001:DFS

[460] William R. Braithwaite. Data field standards and the health insurance portability and accountability act. Statistics in Medicine, 20(9-10):1323– 1330, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carr:2001:DLM

[461] Daniel B. Carr. Designing linked micromap plots for states with many counties. Statistics in Medicine, 20(9-10):1331–1339, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohn:2001:PPH

[462] Victor Cohn. A perspective from the press: how to help reporters tell the truth (sometimes). Statistics in Medicine, 20(9-10):1341–1346, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fienberg:2001:SPC

[463] Stephen E. Fienberg. Statistical perspectives on confidentiality and data access in public health. Statistics in Medicine, 20(9-10):1347–1356, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Howard:2001:PMC

[464] Robert J. Howard. Perspective: media coverage of emerging and reemerging diseases behind the headlines. Statistics in Medicine, 20(9-10):1357–1361, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malec:2001:CBA

[465] Donald J. Malec. Contrasting Bayesian analysis of survey data and clinical trials. Statistics in Medicine, 20(9-10):1363–1371, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:2001:SMA

[466] David M. Murray. Statistical models appropriate for designs often used in group-randomized trials. Statistics in Medicine, 20(9-10):1373–1385, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Portier:2001:LTE

[467] Christopher J. Portier. Linking toxicology and epidemiology: the role of mechanistic modelling. Statistics in Medicine, 20(9-10):1387–1393, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:2001:ECE

[468] Donald B. Rubin. Estimating the causal effects of smoking. Statistics in Medicine, 20(9-10):1395–1414, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steinberg:2001:ECB

[469] Karen K. Steinberg. Ethical challenges at the beginning of the millennium. Statistics in Medicine, 20(9-10):1415–1419, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Banks:2001:PPE

[470] Steven M. Banks and John A. Pandiani. Probabilistic population estimation of the size and overlap of data sets based on date of birth. Statistics in Medicine, 20(9-10):1421–1430, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barker:2001:LRW

[471] Lawrence Barker and Cedric Brown. Logistic regression when binary predictor variables are highly correlated. Statistics in Medicine, 20(9-10):1431–1442, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burr:2001:MSM

[472] Tom L. Burr. Maximally selected measures of evidence of disease clusters. Statistics in Medicine, 20(9-10):1443–1460, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cannon:2001:WCG

[473] Michael J. Cannon, Lee Warner, J. Augusto Taddei, and David G. Kleinbaum. What can go wrong when you assume that correlated data are independent: an illustration from the evaluation of a childhood health intervention in Brazil. Statistics in Medicine, 20(9-10):1461–1467, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Indurkhya:2001:EOC

[474] Alka Indurkhya, Joseph C. Gardiner, and Zhehui Luo. The effect of outliers on confidence interval procedures for cost-effectiveness ratios. Statistics in Medicine, 20(9-10):1469–1477, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kresnow:2001:ATL

[475] Marcie jo Kresnow, Kenneth E. Powell, Kevin B. Webb, James A. Mercy, Lloyd B. Potter, Thomas A. Simon, Robin M. Ikeda, and Ralph Frankowski. Assigning time-linked exposure status to controls in unmatched case-control studies: alcohol use and nearly lethal suicide attempts. Statistics in Medicine, 20(9-10):1479–1485, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:2001:DAP

[476] Andrew C. Leon, Timothy I. Mueller, David A. Solomon, and Martin B. Keller. A dynamic adaptation of the propensity score adjustment for effectiveness analyses of ordinal doses of treatment. Statistics in Medicine, 20(9-10):1487–1498, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Merrill:2001:UDE

[477] Deane W. Merrill. Use of a density equalizing map projection in analysing childhood cancer in four California counties. Statistics in Medicine, 20 (9-10):1499–1513, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Price:2001:MSD

[478] Dionne L. Price and Amita K. Manatunga. Modelling survival data with a cured fraction using frailty models. Statistics in Medicine, 20(9-10):1515–1527, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sternberg:2001:GAE

[479] Maya R. Sternberg and Alula Hadgu. A GEE approach to estimating sensitivity and specificity and coverage properties of the confidence intervals. Statistics in Medicine, 20(9-10):1529–1539, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2001:MIP

[480] Xiao-Hua Zhou, George J. Eckert, and William M. Tierney. Multiple imputation in public health research. Statistics in Medicine, 20(9-10):1541– 1549, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2001:OP

[481] Anonymous. Oral presentations. Statistics in Medicine, 20(9-10):1551– 1553, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2001:PP

[482] Anonymous. Poster presentations. Statistics in Medicine, 20(9-10):1555– 1557, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sieber:2001:MSC

[483] W. Karl Sieber. Meeting summary and closing remarks. Statistics in Medicine, 20(9-10):1559–1561, May 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willan:2001:INB

[484] Andrew R. Willan and D. Y. Lin. Incremental net benefit in randomized clinical trials. Statistics in Medicine, 20(11):1563–1574, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldman:2001:OCS

[485] Anne I. Goldman and Peter J. Hannan. Optimal continuous sequential boundaries for monitoring toxicity in clinical trials: a restricted search algorithm. Statistics in Medicine, 20(11):1575–1589, June 15, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yau:2001:LTS

[486] Kelvin K. W. Yau and Angus S. K. Ng. Long-term survivor mixture model with random effects: application to a multi-centre clinical trial of carcinoma. Statistics in Medicine, 20(11):1591–1607, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chick:2001:BRR

[487] Stephen E. Chick, Daniel C. Barth-Jones, and James S. Koopman. Bias reduction for risk ratio and vaccine effect estimators. Statistics in Medicine, 20(11):1609–1624, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:2001:GFD

[488] J. K. Lindsey. A general family of distributions for longitudinal dependence with special reference to event histories. Statistics in Medicine, 20(11):1625–1638, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanderLinde:2001:ENP

[489] Angelika van der Linde and Gerhard Osius. Estimation of non-parametric multivariate risk functions in matched case-control studies with application to the assessment of interactions of risk factors in the study of cancer. Statistics in Medicine, 20(11):1639–1662, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cengiz:2001:MMG

[490] Mehmet A. Cengiz and David F. Percy. Mixed multivariate generalized linear models for assessing lower-limb arterial stenoses. Statistics in Medicine, 20(11):1663–1679, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seltman:2001:BMS

[491] Howard Seltman, Joel Greenhouse, and Larry Wasserman. Bayesian model selection: analysis of a survival model with a surviving fraction. Statistics in Medicine, 20(11):1681–1691, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutson:2001:BTC

[492] Alan D. Hutson. Bootstrap-type confidence intervals for quantiles of the survival distribution. Statistics in Medicine, 20(11):1693–1702, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2001:MTE

[493] Xiao-Hua Zhou, Chunming Li, Sujuan Gao, and William M. Tierney. Methods for testing equality of means of health care costs in a paired design study. Statistics in Medicine, 20(11):1703–1720, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muller:2001:BRV

[494] Hans-Georg Müller. Book review: Verfahrensbibliothek — Versuchsplanung und Auswertung. D. Rasch, G. Herrendörfer, J. Bock, N. Victor and V. Guiard, Oldenbourg Verlag, München 1996 (book plus diskette). No. of pages: x + 940. Price: DM 178. ISBN 3-486-23146-4. Statistics in Medicine, 20(11):1721–1722, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kubler:2001:BRD

[495] J. Kübler. Book review: Design and analysis of bioavailability and bioequivalence studies, 2nd edn. Shein-Chung Chow and Jen-Pei Liu, Marcel Dekker, New York, 2000. No. of pages: x + 519. Price: $195.00. ISBN 0-8247-7572-4. Statistics in Medicine, 20(11):1722–1723, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Traat:2001:BRS

[496] Imbi Traat. Book review: Sampling of populations. Methods and applications, 3rd edn. P. S. Levy and S. Lemeshow, Wiley, New York, 1999. No. of pages: xxxi + 525. Price: £70. ISBN 0-471-15575-6. Statistics in Medicine, 20(11):1723–1725, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cappelli:2001:BRR

[497] Carmela Cappelli and William Shannon. Book review: Recursive partitioning in the health sciences. H. Zhang and B. Singer, Springer Verlag, New York, 1999. No. of pages: 230. Price: £46.00. ISBN 0-387-98671-5. Statistics in Medicine, 20(11):1725–1726, June 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nixon:2001:RLP

[498] R. M. Nixon, S. W. Duffy, G. R. K. Fender, N. E. Day, and T. C. Prevost. Randomization at the level of primary care practice: use of pre-intervention data and random effects models. Statistics in Medicine, 20(12):1727–1738, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pavlic:2001:EPN

[499] Maja Pavlic, Richard J. Brand, and Steven R. Cummings. Estimating probability of non-response to treatment using mixture distributions. Statistics in Medicine, 20(12):1739–1753, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2001:MIM

[500] Hulin Wu and Lang Wu. A multiple imputation method for missing covariates in non-linear mixed-effects models with application to HIV dynamics. Statistics in Medicine, 20(12):1755–1769, June 30, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hartung:2001:TOT

[501] Joachim Hartung and Guido Knapp. On tests of the overall treatment effect in meta-analysis with normally distributed responses. Statistics in Medicine, 20(12):1771–1782, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peng:2001:TPC

[502] Yingwei Peng, Keith B. G. Dear, and K. C. Carriere. Testing for the presence of cured patients: a simulation study. Statistics in Medicine, 20(12):1783–1796, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2001:NIE

[503] Kung-Jong Lui. Notes on interval estimation of the attributable risk in cross-sectional sampling. Statistics in Medicine, 20(12):1797–1809, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malec:2001:CLC

[504] Donald Malec. A closer look at combining data among a small number of binomial experiments. Statistics in Medicine, 20(12):1811–1824, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bustami:2001:MBO

[505] Rami Bustami, Emmanuel Lesaffre, Geert Molenberghs, Ruth Loos, Marina Danckaerts, and Robert Vlietinck. Modelling bivariate ordinal responses smoothly with examples from ophthalmology and genetics. Statistics in Medicine, 20(12):1825–1842, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2001:CAN

[506] Zhengqing Li. Covariate adjustment for non-parametric tests for censored survival data. Statistics in Medicine, 20(12):1843–1853, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seaman:2001:PHM

[507] S. R. Seaman and S. M. Bird. Proportional hazards model for intervalcensored failure times and time-dependent covariates: application to hazard of HIV infection of injecting drug users in prison. Statistics in Medicine, 20(12):1855–1870, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:2001:MSM

[508] Niels Keiding, John P. Klein, and Mary M. Horowitz. Multi-state models and outcome prediction in bone marrow transplantation. Statistics in Medicine, 20(12):1871–1885, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becher:2001:BRM

[509] Heiko Becher. Book review: Medical statistics. A commonsense approach. M. J. Campbell and D. Machin, Wiley, Chichester, 1999. No. of pages: viii + 203. Price: £16.99. ISBN 0-471-98731-2. Statistics in Medicine, 20 (12):1887, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stephens:2001:BRS

[510] Matthew Stephens. Book review: Statistics in genetics. M. Elizabeth Halloran and Seymour Geisser (eds), Springer, Surrey, 1999. No. of pages:

260. Price: £53.50. ISBN 0-387-98828-9. Statistics in Medicine, 20(12): 1888, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gronning:2001:BRM

[511] Bjørn Grønning and Jens C. Nilsson. Book review: Multiple regression: a primer. Paul D. Allison, Sage, London, 1999. No. of pages: 220. Price: £11.99. ISBN 0-7619-8533-6. Statistics in Medicine, 20(12):1888–1889, June 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2001:UOR

[512] Mei Lu and Barbara C. Tilley. Use of odds ratio or relative risk to measure a treatment effect in clinical trials with multiple correlated binary outcomes: data from the NINDS t-PA stroke trial. Statistics in Medicine, 20(13):1891–1901, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2001:GST

[513] Sue-Jane Wang, H. M. James Hung, Yi Tsong, and Lu Cui. Group sequential test strategies for superiority and non-inferiority hypotheses in active controlled clinical trials. Statistics in Medicine, 20(13):1903–1912, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2001:IME

[514] Liji Shen. An improved method of evaluating drug effect in a multiple dose clinical trial. Statistics in Medicine, 20(13):1913–1929, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghahramani:2001:SMO

[515] M. Ghahramani, C. B. Dean, and J. J. Spinelli. Simultaneous modelling of operative mortality and long-term survival after coronary artery bypass surgery. Statistics in Medicine, 20(13):1931–1945, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2001:MAB

[516] Bei-Hung Chang, Christine Waternaux, and Stuart Lipsitz. Metaanalysis of binary data: which within study variance estimate to use? Statistics in Medicine, 20(13):1947–1956, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2001:RAG

[517] Minge Xie. Regression analysis of group testing samples. Statistics in Medicine, 20(13):1957–1969, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2001:SSC

[518] Sin-Ho Jung, Seung-Ho Kang, and Chul Ahn. Sample size calculations for clustered binary data. Statistics in Medicine, 20(13):1971–1982, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hernandez-Suarez:2001:NDN

[519] Carlos M. Hernández-Suárez. A note on the distribution of the number of vaccinated infected under non-random mixing conditions. Statistics in Medicine, 20(13):1983–1986, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hawkins:2001:SIR

[520] Douglas M. Hawkins, James A. Garrett, and Betty Stephenson. Some issues in resolution of diagnostic tests using an imperfect gold standard. Statistics in Medicine, 20(13):1987–2001, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williams:2001:EHI

[521] Brian Williams, Eleanor Gouws, David Wilkinson, and Salim Abdool Karim. Estimating HIV incidence rates from age prevalence data in epidemic situations. Statistics in Medicine, 20(13):2003–2016, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bellocco:2001:MAS

[522] Rino Bellocco and Marcello Pagano. Multinomial analysis of smoothed HIV back-calculation models incorporating uncertainty in the AIDS incidence. Statistics in Medicine, 20(13):2017–2033, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Militino:2001:UMM

[523] A. F. Militino, M. D. Ugarte, and C. B. Dean. The use of mixture models for identifying high risks in disease mapping. Statistics in Medicine, 20 (13):2035–2049, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2001:BRR

[524] Guobing Lu. Book review: Regression Graphics: Ideas for Studying Regressions Through Graphics. R. D. Cook, Wiley, New York, 1999. No. of pages. xviii + 349, Price: £48.50. ISBN 0-471-19365-8. Statistics in Medicine, 20(13):2051–2052, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nerman:2001:BRP

[525] Olle Nerman. Book review: Progress in Population Genetics and Human Evolution. The IMA Volumes in Mathematics and its Applications, vol 87. Peter Donnelly and Simon Tavaré, Springer, Heidelberg, Germany, 1996. No. of pages: xi + 329. Price: £40.50. ISBN 0-387-94803-1. Statistics in Medicine, 20(13):2052, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borgan:2001:BRM

[526] Ørnulf Borgan. Book review: Modeling Survival Data: Extending the Cox Model. Terry M. Therneau and Patricia M. Grambsch, Springer-Verlag, New York, 2000. No. of pages: xiii + 350. Price: $69.95. ISBN 0-387-98784-3. Statistics in Medicine, 20(13):2053–2054, July 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bagiella:2001:PAT

[527] Emilia Bagiella and Daniel F. Heitjan. Predicting analysis times in randomized clinical trials. Statistics in Medicine, 20(14):2055–2063, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2001:EOM

[528] Mimi Y. Kim and Judith D. Goldberg. The effects of outcome misclassification and measurement error on the design and analysis of therapeutic equivalence trials. Statistics in Medicine, 20(14):2065–2078, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Babb:2001:PSD

[529] James S. Babb and André Rogatko. Patient specific dosing in a cancer phase I clinical trial. Statistics in Medicine, 20(14):2079–2090, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cantor:2001:PSE

[530] Alan B. Cantor. Projecting the standard error of the Kaplan–Meier estimator. Statistics in Medicine, 20(14):2091–2097, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2001:RRE

[531] Devan V. Mehrotra and Arthur J. Roth. Relative risk estimation and inference using a generalized logrank statistic. Statistics in Medicine, 20(14):2099–2113, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stukel:2001:TSM

[532] Therese A. Stukel, Eugene Demidenko, James Dykes, and Margaret R. Karagas. Two-stage methods for the analysis of pooled data. Statistics in Medicine, 20(14):2115–2130, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

King:2001:GCC

[533] Tonya S. King and Vernon M. Chinchilli. A generalized concordance correlation coefficient for continuous and categorical data. Statistics in Medicine, 20(14):2131–2147, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2001:EAC

[534] Sin-Ho Jung, Chul Ahn, and Allan Donner. Evaluation of an adjusted chi-square statistic as applied to observational studies involving clustered binary data. Statistics in Medicine, 20(14):2149–2161, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2001:DSS

[535] Kelly H. Zou and Sharon-Lise T. Normand. On determination of sample size in hierarchical binomial models. Statistics in Medicine, 20(14):2163– 2182, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2001:DMR

[536] Andrew B. Lawson. Disease map reconstruction. Statistics in Medicine, 20(14):2183–2204, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelder:2001:LES

[537] John A. Nelder. Letter to the Editor: Statistics in medical journals: some recent trends by Douglas G. Altman, Statistics in Medicine 2000; 19: 3275–3289. Statistics in Medicine, 20(14):2205, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [314].

Andres:2001:LEA

[538] Antonio Martín Andrés and Inmaculada Herranz Tejedor. Letter to the Editor: An approximate unconditional test of non-inferiority between two proportions by Seung-Ho Kang and James J. Chen, Statistics in Medicine 2000; 19: 2089–2100. Statistics in Medicine, 20(14):2206–2207, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [213].

Kang:2001:AR

[539] Seung-Ho Kang and James J. Chen. Authors’ reply. Statistics in Medicine, 20(14):2207–2208, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sorensen:2001:BRC

[540] Thorkild. I. A. Sørensen. Book review: Critical Appraisal of Epidemiological Studies and Clinical Trials. 2nd edn. J. Mark Elwood, Oxford University Press, Oxford, 1998. No of pages: 448. Price: £29.50. ISBN 0-19-262745-7. Statistics in Medicine, 20(14):2209–2210, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrington:2001:BRG

[541] David Harrington. Book review: Group Sequential Methods with Applications to Clinical Trials. Christopher Jennison and Bruce W. Turnbull, CRC/Chapman and Hall, U.K., 2000. No. of pages: xviii + 390. Price: £39.00. ISBN 0-849-30316-8. Statistics in Medicine, 20(14):2210–2211, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailer:2001:BRP

[542] A. John Bailer. Book review: Probabilistic Techniques in Exposure Assessment. A Handbook for Dealing with Variability and Uncertainty in Models and Inputs. A. C. Cullen and H. C. Frey, Plenum Press, New York and London, 1999. No. of pages: ix + 335. Price: $99.50. ISBN 0-306-45956-6. Statistics in Medicine, 20(14):2211–2213, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Egeland:2001:BRS

[543] Thore Egeland. Book review: Statistical Science in the Courtroom. Joseph L. Gastwirth (ed.), Springer-Verlag, New York, 2000. No. of pages: xxii + 443. Price: $59.95. ISBN 0-387-98997-7. Statistics in Medicine, 20(14):2213–2214, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:2001:BRQ

[544] Jørgen Holm Petersen. Book review: Quality of Life. Assessment, Analysis and Interpretation. P. M. Fayers and D. Machin, Wiley, Chichester, 2000. No. of pages: xii + 404. Price: £60.00. ISBN 0-471-96861-7. Statistics in Medicine, 20(14):2214–2216, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryan:2001:BRH

[545] Louise Ryan. Book review: Human Variability in Response to Chemical Exposures. David Neumann and Carole Kimmel, ILSI Life Sciences Institute, CRC Press, 1998. No. of pages: 304. Price: $99.95. ISBN 0-8493-2805-5. Statistics in Medicine, 20(14):2216–2217, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Higgins:2001:MAC

[546] Julian P. T. Higgins, Anne Whitehead, Rebecca M. Turner, Rumana Z. Omar, and Simon G. Thompson. Meta-analysis of continuous outcome data from individual patients. Statistics in Medicine, 20(15):2219–2241, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2001:MAO

[547] Anne Whitehead, Rumana Z. Omar, Julian P. T. Higgins, Elly Savaluny, Rebecca M. Turner, and Simon G. Thompson. Meta-analysis of ordinal outcomes using individual patient data. Statistics in Medicine, 20 (15):2243–2260, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lunn:2001:CLM

[548] David J. Lunn, Jon Wakefield, and Amy Racine-Poon. Cumulative logit models for ordinal data: a case study involving allergic rhinitis severity scores. Statistics in Medicine, 20(15):2261–2285, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joffe:2001:AAC

[549] Marshall M. Joffe. Administrative and artificial censoring in censored regression models. Statistics in Medicine, 20(15):2287–2304, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mancuso:2001:ORD

[550] Jessica Y. Mancuso, Hongshik Ahn, and James J. Chen. Order-restricted dose-related trend tests. Statistics in Medicine, 20(15):2305–2318, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Assuncao:2001:DPL

[551] Renato M. Assunção, Ilka A. Reis, and Claudia Di Lorenzo Oliveira. Diffusion and prediction of leishmaniasis in a large metropolitan area in Brazil with a Bayesian space-time model. Statistics in Medicine, 20 (15):2319–2335, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2001:UOE

[552] Hsing-Yi Chang, Chirayath M. Suchindran, and Wen-Harn Pan. Using the overdispersed exponential family to estimate the distribution of usual daily intakes of people aged between 18 and 28 in Taiwan. Statistics in Medicine, 20(15):2337–2350, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Batista-Foguet:2001:USE

[553] J. M. Batista-Foguet, G. Coenders, and M. Artés Ferragud. Using structural equation models to evaluate the magnitude of measurement error in blood pressure. Statistics in Medicine, 20(15):2351–2368, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gmel:2001:IMV

[554] Gerhard Gmel. Imputation of missing values in the case of a multiple item instrument measuring alcohol consumption. Statistics in Medicine, 20(15):2369–2381, August 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:2001:GCI

[555] Michael E. Miller, Timothy M. Morgan, Mark A. Espeland, and Scott S. Emerson. Group comparisons involving missing data in clinical trials: a comparison of estimates and power (size) for some simple approaches. Statistics in Medicine, 20(16):2383–2397, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Storer:2001:EPC

[556] Barry E. Storer. An evaluation of phase I clinical trial designs in the continuous dose-response setting. Statistics in Medicine, 20(16):2399– 2408, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tanaka:2001:OET

[557] Mark M. Tanaka and Noah A. Rosenberg. Optimal estimation of transposition rates of insertion sequences for molecular epidemiology. Statistics in Medicine, 20(16):2409–2420, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2001:DAP

[558] Sander Greenland and Ronald Christensen. Data augmentation priors for Bayesian and semi-Bayes analyses of conditional-logistic and proportional-hazards regression. Statistics in Medicine, 20(16):2421– 2428, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arani:2001:PSA

[559] Ramin B. Arani, Seng-Jaw Soong, Heidi L. Weiss, Martin J. Wood, Paul A. Fiddian, John W. Gnann, and Richard Whitley. Phase specific analysis of herpes zoster associated pain data: a new statistical approach. Statistics in Medicine, 20(16):2429–2439, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1925].

Park:2001:ECM

[560] Taesung Park, Jin-Kyung Park, and Charles S. Davis. Effects of covariance model assumptions on hypothesis tests for repeated measurements: analysis of ovarian hormone data and pituitary–pteryomaxillary distance data. Statistics in Medicine, 20(16):2441–2453, August 30, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2001:PMN

[561] Yin Bun Cheung, Paul S. F. Yip, and Johan P. E. Karlberg. Parametric modelling of neonatal mortality in relation to size at birth. Statistics in Medicine, 20(16):2455–2466, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2001:ODL

[562] Jixian Wang. Optimal design for linear interpolation of curves. Statistics in Medicine, 20(16):2467–2477, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altaye:2001:GGF

[563] Mekibib Altaye, Allan Donner, and Michael Eliasziw. A general goodness-of-fit approach for inference procedures concerning the kappa statistic. Statistics in Medicine, 20(16):2479–2488, August 30, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2001:IFB

[564] Elizabeth R. Brown, Samantha MaWhinney, Richard H. Jones, Karen Kafadar, and Benjamin Young. Improving the fit of bivariate smoothing splines when estimating longitudinal immunological and virological markers in HIV patients with individual antiretroviral treatment strategies. Statistics in Medicine, 20(16):2489–2504, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeLong:2001:UOD

[565] Elizabeth R. DeLong, Charlotte L. Nelson, John B. Wong, David B. Pryor, Eric D. Peterson, Kerry L. Lee, Daniel B. Mark, Robert M. Califf, and Stephen G. Pauker. Using observational data to estimate prognosis: an example using a coronary artery disease registry. Statistics in Medicine, 20(16):2505–2532, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henderson:2001:BRA

[566] Rob Henderson. Book review: Analysis of Multivariate Survival Data. Philip Hougaard, Springer, New York, 2000. No. of pages: xvii + 542. Price: $84.95. ISBN 0-387-98873-4. Statistics in Medicine, 20(16):2533– 2534, August 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:2001:P

[567] Byron Jones. Preface. Statistics in Medicine, 20(17-18):2535, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:2001:TPM

[568] Peter Armitage. Theory and practice in medical statistics. Statistics in Medicine, 20(17-18):2537–2548, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:2001:IIG

[569] J. Lewis, W. Louv, F. Rockhold, and T. Sato. The impact of the international guideline entitled Statistical Principles for Clinical Trials (ICH E9). Statistics in Medicine, 20(17-18):2549–2560, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rohmel:2001:SCF

[570] Joachim Röhmel. Statistical considerations of FDA and CPMP rules for the investigation of new anti-bacterial products. Statistics in Medicine, 20(17-18):2561–2571, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2001:IDM

[571] Susan S. Ellenberg. Independent data monitoring committees: rationale, operations and controversies. Statistics in Medicine, 20(17-18): 2573–2583, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tangen:2001:NPA

[572] Catherine M. Tangen and Gary G. Koch. Non-parametric analysis of covariance for confirmatory randomized clinical trials to evaluate dose-response relationships. Statistics in Medicine, 20(17-18):2585–2607, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

LaVange:2001:ASS

[573] L. M. LaVange, G. G. Koch, and T. A. Schwartz. Applying sample survey methods to clinical trials data. Statistics in Medicine, 20(17-18): 2609–2623, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2001:SSR

[574] A. Lawrence Gould. Sample size re-estimation: recent developments and practical considerations. Statistics in Medicine, 20(17-18):2625–2643, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Denne:2001:SSR

[575] Jonathan S. Denne. Sample size recalculation using conditional power. Statistics in Medicine, 20(17-18):2645–2660, September 15–30, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ishizuka:2001:CRM

[576] Naoki Ishizuka and Yasuo Ohashi. The continual reassessment method and its applications: a Bayesian methodology for phase I cancer clinical trials. Statistics in Medicine, 20(17-18):2661–2681, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marks:2001:PSC

[577] Ronald G. Marks, Michael Conlon, and Stephen J. Ruberg. Paradigm shifts in clinical trials enabled by information technology. Statistics in Medicine, 20(17-18):2683–2696, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Enas:2001:EVD

[578] Gregory G. Enas and John S. Andersen. Enhancing the value delivered by the statistician throughout drug discovery and development: putting statistical science into regulated pharmaceutical innovation. Statistics in Medicine, 20(17-18):2697–2708, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agresti:2001:EIC

[579] Alan Agresti. Exact inference for categorical data: recent advances and continuing controversies. Statistics in Medicine, 20(17-18):2709–2722, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Corcoran:2001:CTE

[580] Chris Corcoran, Cyrus Mehta, Nitin Patel, and Pralay Senchaudhuri. Computational tools for exact conditional logistic regression. Statistics in Medicine, 20(17-18):2723–2739, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Landrum:2001:MIS

[581] Mary Beth Landrum and Mark P. Becker. A multiple imputation strategy for incomplete longitudinal data. Statistics in Medicine, 20(17-18): 2741–2760, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sato:2001:MAR

[582] Tosiya Sato. A method for the analysis of repeated binary outcomes in randomized clinical trials with non-compliance. Statistics in Medicine, 20(17-18):2761–2774, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2911].

Lindsey:2001:SSI

[583] J. K. Lindsey, B. Jones, and P. Jarvis. Some statistical issues in modelling pharmacokinetic data. Statistics in Medicine, 20(17-18):2775–2783, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2001:SIB

[584] Stephen Senn. Statistical issues in bioequivalance. Statistics in Medicine, 20(17-18):2785–2799, September 15–30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Denne:2001:MCT

[585] Jonathan S. Denne and Gary G. Koch. Monitoring a clinical trial with multiple hypotheses concerning the treatment effect on a single primary endpoint. Statistics in Medicine, 20(19):2801–2812, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2001:D

[586] John Whitehead. Discussion. Statistics in Medicine, 20(19):2813–2814, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Denne:2001:AR

[587] Jonathan S. Denne and Gary G. Koch. Authors’ reply. Statistics in Medicine, 20(19):2814, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liquet:2001:CVA

[588] Benoit Liquet and Daniel Commenges. Correction of the P -value after multiple coding of an explanatory variable in logistic regression. Statistics in Medicine, 20(19):2815–2826, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zohar:2001:CRM

[589] Sarah Zohar and Sylvie Chevret. The continual reassessment method: comparison of Bayesian stopping rules for dose-ranging studies. Statistics in Medicine, 20(19):2827–2843, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Obuchowski:2001:CEM

[590] Nancy A. Obuchowski. Can electronic medical images replace hard-copy film? Defining and testing the equivalence of diagnostic tests. Statistics in Medicine, 20(19):2845–2863, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rutter:2001:HRA

[591] Carolyn M. Rutter and Constantine A. Gatsonis. A hierarchical regression approach to meta-analysis of diagnostic test accuracy evaluations. Statistics in Medicine, 20(19):2865–2884, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:2001:EDB

[592] Robert G. Newcombe. Estimating the difference between differences: measurement of additive scale interaction for proportions. Statistics in Medicine, 20(19):2885–2893, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2001:EEK

[593] John R. Thompson. Estimating equations for kappa statistics. Statistics in Medicine, 20(19):2895–2906, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yau:2001:ZIP

[594] Kelvin K. W. Yau and Andy H. Lee. Zero-inflated Poisson regression with random effects to evaluate an occupational injury prevention programme. Statistics in Medicine, 20(19):2907–2920, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:2001:CCE

[595] Robert H. Lyles, Dongjie Fan, and Rutt Chuachoowong. Correlation coefficient estimation involving a left censored laboratory assay variable. Statistics in Medicine, 20(19):2921–2933, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2001:GMC

[596] Wei Pan, John E. Connett, Giovanni C. Porzio, and Sanford Weisberg. Graphical model checking with correlated response data. Statistics in Medicine, 20(19):2935–2949, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dubin:2001:EHG

[597] Joel A. Dubin, Hans-Georg Müller, and Jane-Ling Wang. Event history graphs for censored survival data. Statistics in Medicine, 20(19):2951– 2964, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraggi:2001:UNN

[598] David Faraggi, Michael LeBlanc, and John Crowley. Understanding neural networks using regression trees: an application to multiple myeloma survival data. Statistics in Medicine, 20(19):2965–2976, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1016].

Gangnon:2001:WAL

[599] Ronald E. Gangnon and Murray K. Clayton. A weighted average likelihood ratio test for spatial clustering of disease. Statistics in Medicine, 20 (19):2977–2987, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Engholm:2001:BRA

[600] Gerda Engholm. Book review: Analyzing Census Microdata. Angela Dale, Ed Fieldhouse and Clare Holdsworth, Arnold, London, 2000. No. of pages: xiii + 241. Price: £24.99. ISBN 0-340-69228-6. Statistics in Medicine, 20(19):2989–2990, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bithell:2001:BRC

[601] John F. Bithell. Book review: Computer-Assisted Analysis of Mixtures and Applications. Meta-Analysis, Disease Mapping and Others. D. Dankmar Böhning, Chapman and Hall/CRC Press, Hertfordshire, 1998. No. of pages: 224. Price: £69.95. ISBN 0-8493-0385-0. Statistics in Medicine, 20(19):2990–2991, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

George:2001:BRI

[602] Stephen L. George. Book review: An Introduction to Randomized Controlled Clinical Trials. J. N. S. Matthews, Arnold, London, U.K., 2000. Number of pages: xiv + 189. Price: £19.99. ISBN 0-340-761431. Statistics in Medicine, 20(19):2992–2993, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dwivedi:2001:BRC

[603] S. N. Dwivedi. Book review: Cancer Survival in Developing Countries. R. Sankarnarayanan, R. L. Black and D. M. Parkin (eds), International Agency for Research on Cancer, Lyon, France (Scientific Publications No. 145), 1998. Number of pages: x + 173. Price: $69.0. ISBN 92-832-2145-1. Statistics in Medicine, 20(19):2993–2994, October 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2001:RCT

[604] Ian R. White, Christina Bamias, Pollyanna Hardy, Stuart Pocock, and Jacquie Warner. Randomized clinical trials with added rescue medication: some approaches to their analysis and interpretation. Statistics in Medicine, 20(20):2995–3008, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Honkanen:2001:TSC

[605] Visa E. A. Honkanen, Andrew F. Siegel, John Paul Szalai, Vance Berger, Brian M. Feldman, and Jeffrey N. Siegel. A three-stage clinical trial design for rare disorders. Statistics in Medicine, 20(20):3009–3021, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molenberghs:2001:ESE

[606] Geert Molenberghs, Helena Geys, and Marc Buyse. Evaluation of surrogate endpoints in randomized experiments with mixed discrete and continuous outcomes. Statistics in Medicine, 20(20):3023–3038, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vines:2001:WSE

[607] S. K. Vines and C. P. Farrington. Within-subject exposure dependency in case-crossover studies. Statistics in Medicine, 20(20):3039–3049, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nobre:2001:DLM

[608] Flávio Fonseca Nobre, Ana Beatriz Soares Monteiro, Paulo Roberto Telles, and G. David Williamson. Dynamic linear model and SARIMA: a comparison of their forecasting performance in epidemiology. Statistics in Medicine, 20(20):3051–3069, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoover:2001:AST

[609] Donald R. Hoover and William C. Blackwelder. Allocation of subjects to test null relative risks smaller than one. Statistics in Medicine, 20 (20):3071–3082, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henderson:2001:APP

[610] Robin Henderson, Margaret Jones, and Janez Stare. Accuracy of point predictions in survival analysis. Statistics in Medicine, 20(20):3083–3096, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fong:2001:MVL

[611] D. Y. T. Fong, S. W. S. Wong, and J. Karlberg. Modelling variation of lower leg length growth in early life. Statistics in Medicine, 20(20):3097– 3108, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gossl:2001:BAL

[612] Christoff Gössl and Helmut Küchenhoff. Bayesian analysis of logistic regression with an unknown change point and covariate measurement error. Statistics in Medicine, 20(20):3109–3121, October 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chao:2001:ACR

[613] Anne Chao, P. K. Tsay, Sheng-Hsiang Lin, Wen-Yi Shau, and Day-Yu Chao. The applications of capture-recapture models to epidemiological data. Statistics in Medicine, 20(20):3123–3157, October 30, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quan:2001:AEM

[614] Hui Quan, Jim Bolognese, and Weiying Yuan. Assessment of equivalence on multiple endpoints. Statistics in Medicine, 20(21):3159–3173, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2001:MAP

[615] Zhengqing Li, Michael P. Meredith, and Mohammad S. Hoseyni. A method to assess the proportion of treatment effect explained by a surrogate endpoint. Statistics in Medicine, 20(21):3175–3188, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:2001:SEM

[616] Nicholas T. Longford. Synthetic estimators with moderating influence: the carry-over in cross-over trials revisited. Statistics in Medicine, 20 (21):3189–3203, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giraudeau:2001:PRS

[617] B. Giraudeau and J. Y. Mary. Planning a reproducibility study: how many subjects and how many replicates per subject for an expected width of the 95 per cent confidence interval of the intraclass correlation coefficient. Statistics in Medicine, 20(21):3205–3214, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Viallefont:2001:VSB

[618] Valerie Viallefont, Adrian E. Raftery, and Sylvia Richardson. Variable selection and Bayesian model averaging in case-control studies. Statistics in Medicine, 20(21):3215–3230, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lumley:2001:CCQ

[619] T. Lumley, R. J. Simes, V. Gebski, and H. M. Hudson. Combining components of quality of life to increase precision and evaluate trade-offs.

Statistics in Medicine, 20(21):3231–3249, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willan:2001:CEA

[620] Andrew R. Willan, Bernie J. O’Brien, and Rina A. Leyva. Costeffectiveness analysis when the WTA is greater than the WTP. Statistics in Medicine, 20(21):3251–3259, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Obuchowski:2001:APA

[621] Nancy A. Obuchowski, Marilyn J. Goske, and Kimberly E. Applegate. Assessing physicians’ accuracy in diagnosing paediatric patients with acute abdominal pain: measuring accuracy for multiple diseases. Statistics in Medicine, 20(21):3261–3278, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dominik:2001:SME

[622] Rosalie Dominik, Haibo Zhou, and Jianwen Cai. A statistical model for the evaluation of barrier contraceptive efficacy. Statistics in Medicine, 20(21):3279–3294, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmid:2001:MDR

[623] Christopher H. Schmid. Marginal and dynamic regression models for longitudinal data. Statistics in Medicine, 20(21):3295–3311, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:2001:SPR

[624] Jianguo Sun. A semi-parametric regression analysis of CD4 cell counts. Statistics in Medicine, 20(21):3313–3322, November 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Husing:2001:SDE

[625] Johannes Hüsing, Wolfgang Sauerwein, Katalin Hideghéty, and Karl-Heinz Jöckel. A scheme for a dose-escalation study when the event is lagged. Statistics in Medicine, 20(22):3323–3334, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leuraud:2001:CSM

[626] Klervi Leuraud and Jacques Benichou. A comparison of several methods to test for the existence of a monotonic dose-response relationship in clinical and epidemiological studies. Statistics in Medicine, 20(22):3335– 3351, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:2001:CGL

[627] Brent A. Evans, Ziding Feng, and Arthur V. Peterson. A comparison of generalized linear mixed model procedures with estimating equations for variance and covariance parameter estimation in longitudinal studies and group randomized trials. Statistics in Medicine, 20(22):3353–3373, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [958].

Taylor:2001:EPB

[628] Jeremy M. G. Taylor, Robert E. Weiss, Wenzhi Li, Chiu-Hsieh Hsu, and Rafal Suwinski. Estimation for paired binomial data with application to radiation therapy. Statistics in Medicine, 20(22):3375–3390, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gurrin:2001:SMF

[629] Lyle C. Gurrin, Kevin V. Blake, Sharon F. Evans, and John P. Newnham. Statistical measures of foetal growth using linear mixed models applied to the foetal origins hypothesis. Statistics in Medicine, 20(22):3391– 3409, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2001:ARE

[630] Mingxiu Hu and John M. Lachin. Application of robust estimating equations to the analysis of quantitative longitudinal data. Statistics in Medicine, 20(22):3411–3428, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diaz:2001:TAS

[631] M. E. Díaz, G. Ayala, S. Quesada, and L. Martínez-Costa. Testing abnormality in the spatial arrangement of cells in the corneal endothelium using spatial point processes. Statistics in Medicine, 20(22):3429–3439, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2001:CME

[632] Ian White, Chris Frost, and Shoji Tokunaga. Correcting for measurement error in binary and continuous variables using replicates. Statistics in Medicine, 20(22):3441–3457, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xue:2001:ACB

[633] Xiaonan Xue. Analysis of childhood brain tumour data in new York city using frailty models. Statistics in Medicine, 20(22):3459–3473, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heller:2001:APR

[634] Glenn Heller. An adjustment for a post-randomization variable in the comparison of two treatments for survival. Statistics in Medicine, 20 (22):3475–3485, November 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piaggio:2001:P

[635] Gilda Piaggio, Catherine d’Arcangues, and David Machin. Preface. Statistics in Medicine, 20(23):3487–3488, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Look:2001:I

[636] Paul F. A. Van Look. Introduction. Statistics in Medicine, 20(23):3489– 3490, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Platt:2001:DEE

[637] Robert W. Platt, Michal Abrahamowicz, Michael S. Kramer, K. S. Joseph, Les Mery, Béatrice Blondel, Gérard Bréart, and S. W. Wen. Detecting and eliminating erroneous gestational ages: a normal mixture model. Statistics in Medicine, 20(23):3491–3503, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hakkert:2001:CEM

[638] Ralph Hakkert. Country estimates of maternal mortality: an alternative model. Statistics in Medicine, 20(23):3505–3524, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gaillard:2001:CSI

[639] Philippe Gaillard, Ellen Piwoz, and Tim M. M. Farley. Collection of standardized information on infant feeding in the context of motherto-child transmission of HIV. Statistics in Medicine, 20(23):3525–3537, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alioum:2001:EEI

[640] Ahmadou Alioum, François Dabis, Laurence Dequae-Merchadou, Geert Haverkamp, Michael Hudgens, James Hughes, John Karon, Valeriane Leroy, Marie-Louise Newell, Barbra Richardson, and Gerrit-Jan Weverling. Estimating the efficacy of interventions to prevent mother-to-child transmission of HIV in breast-feeding populations: development of a consensus methodology. Statistics in Medicine, 20(23):3539–3556, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skegg:2001:ESM

[641] David C. G. Skegg. Evaluating the safety of medicines, with particular reference to contraception. Statistics in Medicine, 20(23):3557–3569, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piaggio:2001:UEA

[642] Gilda Piaggio and Alain P. Y. Pinol. Use of the equivalence approach in reproductive health clinical trials. Statistics in Medicine, 20(23):3571– 3578, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2001:CRAb

[643] Bee-Choo Tai, Alexander Peregoudov, and David Machin. A competing risk approach to the analysis of trials of alternative intra-uterine devices (IUDs) for fertility regulation. Statistics in Medicine, 20(23):3589–3600, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farley:2001:CAA

[644] Timothy M. M. Farley, Mohamed M. Ali, and Emma Slaymaker. Competing approaches to analysis of failure times with competing risks. Statistics in Medicine, 20(23):3601–3610, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ali:2001:AFT

[645] Mohamed M. Ali, Abdel G. Babiker, and John G. Cleland. Analysis of failure time hierarchical data in the presence of competing risks with application to oral contraceptive pill use in Egypt. Statistics in Medicine, 20(23):3611–3624, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petitti:2001:AHM

[646] Diana B. Petitti. Approaches to heterogeneity in meta-analysis. Statistics in Medicine, 20(23):3625–3633, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Villar:2001:MAS

[647] José Villar, Maria Eugenia Mackey, Guillermo Carroli, and Allan Donner. Meta-analyses in systematic reviews of randomized controlled trials in perinatal medicine: comparison of fixed and random effects models. Statistics in Medicine, 20(23):3635–3647, December 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valsecchi:2001:P

[648] Maria Grazia Valsecchi. Preface. Statistics in Medicine, 20(24):3649, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rossi:2001:BFM

[649] Carla Rossi. Bruno de Finetti: the mathematician, the statistician, the economist, the forerunner. Statistics in Medicine, 20(24):3651–3666, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sypsa:2001:CST

[650] V. Sypsa, G. Touloumi, M. Kenward, A. Karafoulidou, and A. Hatzakis. Comparison of smoothing techniques for CD4 data in a Markov model with states defined by CD4: an example on the estimation of the HIV incubation time distribution. Statistics in Medicine, 20(24):3667–3676, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boracchi:2001:MCS

[651] Patrizia Boracchi, Elia Biganzoli, and Ettore Marubini. Modelling causespecific hazards with radial basis function artificial neural networks: application to 2233 breast cancer patients. Statistics in Medicine, 20(24): 3677–3694, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hall:2001:EBM

[652] Charles B. Hall, Jun Ying, Lynn Kuo, Martin Sliwinski, Herman Buschke, Mindy Katz, and Richard B. Lipton. Estimation of bivariate measurements having different change points, with application to cognitive ageing. Statistics in Medicine, 20(24):3695–3714, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Touloumi:2001:IMD

[653] G. Touloumi, A. G. Babiker, S. J. Pocock, and J. H. Darbyshire. Impact of missing data due to drop-outs on estimators for rates of change in longitudinal studies: a simulation study. Statistics in Medicine, 20(24): 3715–3728, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klar:2001:CFC

[654] Neil Klar and Allan Donner. Current and future challenges in the design and analysis of cluster randomization trials. Statistics in Medicine, 20 (24):3729–3740, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schafer:2001:MSS

[655] Helmut Schäfer and Hans-Helge Müller. Modification of the sample size and the schedule of interim analyses in survival trials based on data inspections. Statistics in Medicine, 20(24):3741–3751, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1374].

Loeys:2001:ACC

[656] T. Loeys, S. Vansteelandt, and E. Goetghebeur. Accounting for correlation and compliance in cluster randomized trials. Statistics in Medicine, 20(24):3753–3767, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marinacci:2001:ARE

[657] Chiara Marinacci, Patrizia Schifano, Piero Borgia, and Carlo A. Perucci. Application of random effect ordinal regression model for outcome evaluation of two randomized controlled trials. Statistics in Medicine, 20(24): 3769–3776, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vail:2001:PAB

[658] A. Vail, J. Hornbuckle, D. J. Spiegelhalter, and J. G. Thornton. Prospective application of Bayesian monitoring and analysis in an ‘open’ randomized clinical trial. Statistics in Medicine, 20(24):3777–3787, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2001:AAB

[659] Paul C. Lambert, Keith R. Abrams, David R. Jones, Aidan W. F. Halligan, and Andrew Shennan. Analysis of ambulatory blood pressure monitor data using a hierarchical model incorporating restricted cubic splines and heterogeneous within-subject variances. Statistics in Medicine, 20 (24):3789–3805, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahe:2001:ARF

[660] Cédric Mahé and Sylvie Chevret. Analysis of recurrent failure times data: should the baseline hazard be stratified. Statistics in Medicine, 20(24): 3807–3815, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deBruijne:2001:UCR

[661] Mattheus H. J. de Bruijne, Saskia le Cessie, Hanneke C. Kluin-Nelemans, and Hans C. van Houwelingen. On the use of Cox regression in the presence of an irregularly observed time-dependent covariate. Statistics in Medicine, 20(24):3817–3829, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Svensson:2001:CSG

[662] Elisabeth Svensson. Construction of a single global scale for multi-item assessments of the same variable. Statistics in Medicine, 20(24):3831– 3846, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steyerberg:2001:RMH

[663] Ewout W. Steyerberg, Yvonne Vergouwe, H. Jan Keizer, J. Dik F. Habbema, and for the ReHiT study group. Residual mass histology in testicular cancer: development and validation of a clinical prediction rule. Statistics in Medicine, 20(24):3847–3859, December 30, 2001. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2001:CMA

[664] Tim Friede and Meinhard Kieser. A comparison of methods for adaptive sample size adjustment. Statistics in Medicine, 20(24):3861–3873, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hartung:2001:RMM

[665] Joachim Hartung and Guido Knapp. A refined method for the metaanalysis of controlled clinical trials with binary outcome. Statistics in Medicine, 20(24):3875–3889, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eshima:2001:AIS

[666] Nobuoki Eshima, Minoru Tabata, Hiroshi Kikuchi, Shigeru Karukaya, and Toshio Taguchi. Analysis of the infection system of human Tcell leukaemia virus type I based on a mathematical epidemic model. Statistics in Medicine, 20(24):3891–3900, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steen:2001:SAL

[667] Kristel Van Steen, Desmond Curran, and Geert Molenberghs. Sensitivity analysis of longitudinal binary quality of life data with drop-out: an example using the EORTC QLQ-C30. Statistics in Medicine, 20(24): 3901–3920, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tobi:2001:TCC

[668] Hilde Tobi, D. Joop Kuik, P. Dick Bezemer, and Paul Ket. Towards a curriculum for the consultant biostatistician: identification of central disciplines. Statistics in Medicine, 20(24):3921–3929, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Unnebrink:2001:ITM

[669] Kristina Unnebrink and Jürgen Windeler. Intention-to-treat: methods for dealing with missing values in clinical trials of progressively deteriorating diseases. Statistics in Medicine, 20(24):3931–3946, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2001:NNT

[670] S. D. Walter. Number needed to treat (NNT): estimation of a measure of clinical benefit. Statistics in Medicine, 20(24):3947–3962, December 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2002:F

[671] Ralph B. D’Agostino and David Machin. Foreword. Statistics in Medicine, 21(1):1–2, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Satten:2002:MEM

[672] Glen A. Satten and Somnath Datta. Marginal estimation for multistage models: waiting time distributions and competing risks analyses. Statistics in Medicine, 21(1):3–19, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galai:2002:EIH

[673] Noya Galai, Elisheva Simchen, Dalit Braun, Micha Mandel, and Yana Zitser-Gurevich. Evaluating inter-hospital variability in mortality rates over time, allowing for time-varying effects. Statistics in Medicine, 21(1): 21–33, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gauderman:2002:SSR

[674] W. James Gauderman. Sample size requirements for matched casecontrol studies of gene-environment interaction. Statistics in Medicine, 21(1):35–50, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leung:2002:ECR

[675] Denis Heng-Yan Leung and You-Gan Wang. An extension of the continual reassessment method using decision theory. Statistics in Medicine, 21(1):51–63, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freidlin:2002:TPS

[676] Boris Freidlin and Edward L. Korn. A testing procedure for survival data with few responders. Statistics in Medicine, 21(1):65–78, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pulkstenis:2002:TGF

[677] Erik Pulkstenis and Timothy J. Robinson. Two goodness-of-fit tests for logistic regression models with continuous covariates. Statistics in Medicine, 21(1):79–93, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hooper:2002:MFG

[678] Peter M. Hooper, Damon C. Mayes, and Nestor N. Demianczuk. A model for foetal growth and diagnosis of intrauterine growth restriction. Statistics in Medicine, 21(1):95–112, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2002:HMM

[679] Christopher H. Jackson and Linda D. Sharples. Hidden Markov models for the onset and progression of bronchiolitis obliterans syndrome in lung transplant recipients. Statistics in Medicine, 21(1):113–128, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2002:IRM

[680] Chen Wang, Jeffrey Douglas, and Susan Anderson. Item response models for joint analysis of quality of life and survival. Statistics in Medicine, 21(1):129–142, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2002:CMN

[681] Taesung Park and Young Jack Lee. Covariance models for nested repeated measures data: analysis of ovarian steroid secretion data. Statistics in Medicine, 21(1):143–164, January 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2002:IEA

[682] Tim Friede and Meinhard Kieser. On the inappropriateness of an EM algorithm based procedure for blinded sample size re-estimation. Statistics in Medicine, 21(2):165–176, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1589].

Murray:2002:GSM

[683] Susan Murray. Group sequential monitoring of years of life saved with paired censored survival data. Statistics in Medicine, 21(2):177–189, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2002:AUD

[684] Youngjo Lee and John A. Nelder. Analysis of ulcer data using hierarchical generalized linear models. Statistics in Medicine, 21(2):191–202, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

West:2002:CIM

[685] Colin P. West and Jeffrey D. Dawson. Complete imputation of missing repeated categorical data: one-sample applications. Statistics in Medicine, 21(2):203–217, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pauler:2002:NLH

[686] Donna K. Pauler and Nan M. Laird. Non-linear hierarchical models for monitoring compliance. Statistics in Medicine, 21(2):219–229, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2002:TEN

[687] Jen pei Liu, Huey miin Hsueh, Eric Hsieh, and James J. Chen. Tests for equivalence or non-inferiority for paired binary data. Statistics in Medicine, 21(2):231–245, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2002:SUH

[688] Wen-Chung Lee. Standardization using the harmonically weighted ratios: internal and external comparisons. Statistics in Medicine, 21(2):247–261, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1056].

Betensky:2002:LLP

[689] Rebecca A. Betensky, Jane C. Lindsey, Louise M. Ryan, and M. P. Wand. A local likelihood proportional hazards model for interval censored data. Statistics in Medicine, 21(2):263–275, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2002:BAA

[690] Soomin Park, Mari Palta, Jun Shao, and Lei Shen. Bias adjustment in analysing longitudinal data with informative missingness. Statistics in Medicine, 21(2):277–291, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sposto:2002:CMA

[691] Richard Sposto. Cure model analysis in cancer: an application to data from the Children’s cancer group. Statistics in Medicine, 21(2):293–312, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keles:2002:RBT

[692] Sündüz Keles and Mark R. Segal. Residual-based tree-structured survival analysis. Statistics in Medicine, 21(2):313–326, January 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carriere:2002:P

[693] Keumhee Chough Carrière, Allan Donner, and Stephen Walter. Preface. Statistics in Medicine, 21(3):327–329, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Normand:2002:SSC

[694] Sharon-Lise T. Normand and Kelly H. Zou. Sample size considerations in observational health care quality studies. Statistics in Medicine, 21 (3):331–345, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2002:STM

[695] Ying C. MacNab and C. B. Dean. Spatio-temporal modelling of rates for the construction of disease maps. Statistics in Medicine, 21(3):347– 358, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2002:SMR

[696] Andrew B. Lawson and Allan Clark. Spatial mixture relative risk models applied to disease mapping. Statistics in Medicine, 21(3):359–370, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1054].

Berlin:2002:IPV

[697] Jesse A. Berlin, Jill Santanna, Christopher H. Schmid, Lynda A. Szczech, and Harold I. Feldman. Individual patient- versus group-level data metaregressions for the investigation of treatment effect modifiers: ecological bias rears its ugly head. Statistics in Medicine, 21(3):371–387, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2002:RMT

[698] Sander Greenland. A review of multilevel theory for ecologic analyses. Statistics in Medicine, 21(3):389–395, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flynn:2002:RSC

[699] Terry N. Flynn, Elise Whitley, and Tim J. Peters. Recruitment strategies in a cluster randomized trial — cost implications. Statistics in Medicine, 21(3):397–405, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheng:2002:ICE

[700] Xiaoming Sheng and K. C. Carrière. An improved CML estimation procedure for the Rasch model with item response data. Statistics in Medicine, 21(3):407–416, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCulloch:2002:DSS

[701] C. E. McCulloch, H. Lin, E. H. Slate, and B. W. Turnbull. Discovering subpopulation structure with latent class mixed models. Statistics in Medicine, 21(3):417–429, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bull:2002:RMA

[702] Shelley B. Bull, Celia M. T. Greenwood, Lucia Mirea, and Kenneth Morgan. Regression models for allele sharing: analysis of accumulating data in affected sib pair studies. Statistics in Medicine, 21(3):431–444, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2002:NPM

[703] Mei-Cheng Wang and Chin-Tsang Chiang. Non-parametric methods for recurrent event data with informative and non-informative censorings. Statistics in Medicine, 21(3):445–456, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hussein:2002:MCC

[704] A. Hussein and K. C. Carrière. A measure of continuity of care based on the multiplicative intensity model. Statistics in Medicine, 21(3):457– 465, February 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2002:EGS

[705] Nigel Stallard and William F. Rosenberger. Exact group-sequential designs for clinical trials with randomized play-the-winner allocation. Statistics in Medicine, 21(4):467–480, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longini:2002:MBE

[706] Ira M. Longini, Jr., M. Elizabeth Halloran, and Azhar Nizam. Modelbased estimation of vaccine effects from community vaccine trials. Statistics in Medicine, 21(4):481–495, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2002:STD

[707] Toshiro Tango. Score tests for detecting excess risks around putative sources. Statistics in Medicine, 21(4):497–514, February 28, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2002:SER

[708] Richard J. Cook and Wei Wei. Selection effects in randomized trials with count data. Statistics in Medicine, 21(4):515–531, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bortot:2002:PTA

[709] Paola Bortot, Karl Thomaseth, and Alberto Salvan. Population toxicokinetic analysis of 2,3,7,8-tetrachlorodibenzo-p-dioxin using Bayesian techniques. Statistics in Medicine, 21(4):533–547, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biswas:2002:BAB

[710] Atanu Biswas and Kalyan Das. A Bayesian analysis of bivariate ordinal data: Wisconsin epidemiologic study of diabetic retinopathy revisited. Statistics in Medicine, 21(4):549–559, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2002:PPP

[711] Guillermo Marshall, Bradley Warner, Samantha MaWhinney, and Karl Hammermeister. Prospective prediction in the presence of missing data. Statistics in Medicine, 21(4):561–570, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sithole:2002:RMM

[712] Jabulani S. Sithole and Peter W. Jones. Repeated measures models for prescribing change. Statistics in Medicine, 21(4):571–587, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanHouwelingen:2002:AMM

[713] Hans C. van Houwelingen, Lidia R. Arends, and Theo Stijnen. Advanced methods in meta-analysis: multivariate approach and metaregression. Statistics in Medicine, 21(4):589–624, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2002:LEF

[714] Kung-Jong Lui. Letter to the Editor: A flexible design for multiple armed screening trials by Daniel J. Sargent and Richard M. Goldberg, Statistics in Medicine 2001; 20: 1051–1060. Statistics in Medicine, 21 (4):625–627, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [426].

Sargent:2002:AR

[715] Daniel J. Sargent. Author’s reply. Statistics in Medicine, 21(4):628, February 28, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2002:IBT

[716] Shein-Chung Chow, Jun Shao, and Hansheng Wang. Individual bioequivalence testing under 2 × 3 designs. Statistics in Medicine, 21(5): 629–648, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2002:EEC

[717] Jeffrey M. Albert. Estimating efficacy in clinical trials with clustered binary responses. Statistics in Medicine, 21(5):649–661, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lecoutre:2002:AMC

[718] Bruno Lecoutre, Bienvenu Mabika, and Gérard Derzko. Assessment and monitoring in clinical trials when survival curves have distinct shapes: a Bayesian approach with Weibull modelling. Statistics in Medicine, 21(5): 663–674, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsuyama:2002:CNC

[719] Yutaka Matsuyama. Correcting for non-compliance of repeated binary outcomes in randomized clinical trials: randomized analysis approach. Statistics in Medicine, 21(5):675–687, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter

Nam:2002:ARM

[720] Jun mo Nam and William C. Blackwelder. Analysis of the ratio of marginal probabilities in a matched-pair setting. Statistics in Medicine, 21(5):689–699, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2002:NPM

[721] Dong D. Zhang, Xia-Hua Zhou, Daniel H. Freeman, Jr., and Jean L. Freeman. A non-parametric method for the comparison of partial areas under ROC curves and its application to large health care data sets. Statistics in Medicine, 21(5):701–715, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valenta:2002:ESF

[722] Zdenek Valenta and Lisa Weissfeld. Estimation of the survival function for Gray’s piecewise-constant time-varying coefficients model. Statistics in Medicine, 21(5):717–727, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2002:TMF

[723] Jin Wang, Jun Shao, and Mari Palta. Testing model fit in longitudinal data analysis against alternatives with omitted covariates. Statistics in Medicine, 21(5):729–741, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lim:2002:MLE

[724] Heejeong Lim, Jianguo Sun, and David E. Matthews. Maximum likelihood estimation of a survival function with a change point for truncated and interval-censored data. Statistics in Medicine, 21(5):743–752, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2002:ISH

[725] Hulin Wu and Lang Wu. Identification of significant host factors for HIV dynamics modelled by non-linear mixed-effects models. Statistics in Medicine, 21(5):753–771, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ekholm:2002:CRA

[726] Anders Ekholm, Jukka Jokinen, and Terhi Kilpi. Combining regression and association modelling for longitudinal data on bacterial carriage. Statistics in Medicine, 21(5):773–791, March 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moses:2002:CRL

[727] Lincoln E. Moses, Frederick Mosteller, and John H. Buehler. Comparing results of large clinical trials to those of meta-analyses. Statistics in Medicine, 21(6):793–800, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deutsch:2002:SSA

[728] Reena Deutsch. A seminar series in applied biostatistics for clinical research fellows, faculty and staff. Statistics in Medicine, 21(6):801–810, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cowles:2002:BEP

[729] Mary Kathryn Cowles. Bayesian estimation of the proportion of treatment effect captured by a surrogate marker. Statistics in Medicine, 21(6): 811–834, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonzo:2002:SSC

[730] Todd A. Alonzo, Margaret S. Pepe, and Chaya S. Moskowitz. Sample size calculations for comparative studies of medical tests for detecting presence of disease. Statistics in Medicine, 21(6):835–852, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:2002:AMC

[731] Geoffrey Berry, Catherine L. Smith, Petra Macaskill, and Les Irwig. Analytic methods for comparing two dichotomous screening or diagnostic tests applied to two populations of differing disease prevalence when individuals negative on both tests are unverified. Statistics in Medicine, 21(6):853–862, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2002:SFR

[732] Li Hsu, Ross L. Prentice, and Janet L. Stanford. Some further results on incorporating risk factor information in assessing the dependence between paired failure times arising from case-control family studies: an application to prostate cancer. Statistics in Medicine, 21(6):863–876, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dragalin:2002:SAE

[733] Vladimir Dragalin, Valerii Fedorov, and Brigitte Cheuvart. Statistical approaches to establishing vaccine safety. Statistics in Medicine, 21(6): 877–893, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsodikov:2002:SPM

[734] A. Tsodikov. Semi-parametric models of long- and short-term survival: an application to the analysis of breast cancer survival in Utah by age and stage. Statistics in Medicine, 21(6):895–920, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McClelland:2002:RBV

[735] R. L. McClelland and R. A. Kronmal. Regression-based variable clustering for data reduction. Statistics in Medicine, 21(6):921–941, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2002:QEM

[736] Xun Chen. A quasi-exact method for the confidence intervals of the difference of two independent binomial proportions in small sample cases. Statistics in Medicine, 21(6):943–956, March 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moore:2002:F

[737] Laurence Moore, Tim J. Peters, and Jonathan A. C. Sterne. Foreword. Statistics in Medicine, 21(7):957, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:2002:ITY

[738] Douglas Altman. Invited talk: 20 years of Burwalls. Statistics in Medicine, 21(7):959–967, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morris:2002:DEO

[739] Richard W. Morris. Does EBM offer the best opportunity yet for teaching medical statistics? Statistics in Medicine, 21(7):969–977, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armstrong:2002:DDE

[740] Gillian K. Armstrong and Stephen J. Walters. Discussion: Does EBM offer the best opportunity yet for teaching medical statistics? Statistics in Medicine, 21(7):979–981, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sterne:2002:D

[741] J. A. C. Sterne. Discussion. Statistics in Medicine, 21(7):983–984, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sterne:2002:THT

[742] Jonathan A. C. Sterne. Teaching hypothesis tests — time for significant change? Statistics in Medicine, 21(7):985–994, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:2002:DTH

[743] Christopher R. Palmer. Discussion: Teaching hypothesis tests: time for significant change? Statistics in Medicine, 21(7):995–999, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peters:2002:D

[744] T. J. Peters. Discussion. Statistics in Medicine, 21(7):1001, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Astin:2002:MSP

[745] Jeremy Astin, Tom Jenkins, and Laurence Moore. Medical students’ perspective on the teaching of medical statistics in the undergraduate medical curriculum. Statistics in Medicine, 21(7):1003–1006, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moore:2002:D

[746] Laurence Moore. Discussion. Statistics in Medicine, 21(7):1007, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bruce:2002:TES

[747] Jane C. Bruce, Stephen T. Bond, and Michael E. Jones. Teaching epidemiology and statistics by distance learning. Statistics in Medicine, 21(7):1009–1020, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilmour:2002:DTE

[748] Harper Gilmour. Discussion: Teaching epidemiology and statistics by distance learning. Statistics in Medicine, 21(7):1021–1022, April 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Michiels:2002:SMP

[749] Bart Michiels, Geert Molenberghs, Luc Bijnens, Tony Vangeneugden, and Herbert Thijs. Selection models and pattern-mixture models to analyse longitudinal quality of life data subject to drop-out. Statistics in Medicine, 21(8):1023–1041, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carpenter:2002:CMD

[750] James Carpenter, Stuart Pocock, and Carl Johan Lamm. Coping with missing data in clinical trials: a model-based approach applied to asthma trials. Statistics in Medicine, 21(8):1043–1066, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bronskill:2002:LPH

[751] Susan E. Bronskill, Sharon-Lise T. Normand, Mary Beth Landrum, and Robert A. Rosenheck. Longitudinal profiles of health care providers. Statistics in Medicine, 21(8):1067–1088, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulathinal:2002:EEV

[752] S. B. Kulathinal, Kari Kuulasmaa, and Dario Gasbarra. Estimation of an errors-in-variables regression model when the variances of the measurement errors vary between the observations. Statistics in Medicine, 21(8):1089–1101, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ratcliffe:2002:FDAa

[753] Sarah J. Ratcliffe, Leo R. Leader, and Gillian Z. Heller. Functional data analysis with application to periodically stimulated foetal heart rate data. i: Functional regression. Statistics in Medicine, 21(8):1103–1114, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ratcliffe:2002:FDAb

[754] Sarah J. Ratcliffe, Gillian Z. Heller, and Leo R. Leader. Functional data analysis with application to periodically stimulated foetal heart rate data. II: Functional logistic regression. Statistics in Medicine, 21(8):1115–1127, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choudhury:2002:NPC

[755] Jahar B. Choudhury. Non-parametric confidence interval estimation for competing risks analysis: application to contraceptive data. Statistics in Medicine, 21(8):1129–1144, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ciampi:2002:PTS

[756] Antonio Ciampi, André Couturier, and Shaolin Li. Prediction trees with soft nodes for binary outcomes. Statistics in Medicine, 21(8):1145–1165, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mazumdar:2002:LEI

[757] Madhu Mazumdar. Letter to the Editor: The impact of covariate imbalance on the size of the logrank test in randomized clinical trials N. Kinukawa, T. Nakamura, K. Akazawa and Y. Nose, Statistics in Medicine 2000; 19: 1955–1967. Statistics in Medicine, 21(8):1167, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [202].

Kinukawa:2002:AR

[758] Naoko Kinukawa, Tsuyoshi Nakamura, Kouhei Akazawa, and Yoshiaki Nose. Authors’ reply. Statistics in Medicine, 21(8):1167–1171, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2002:LEH

[759] John Whitehead. Letter to the Editor: Heterogeneity in Phase I clinical trials: prior elicitation and computation using the continual reassessment method, by A. Legedza and J. G. Ibrahim, Statistics in Medicine 2001; 20: 867–882. Statistics in Medicine, 21(8):1172, April 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [406].

Loeys:2002:BIS

[760] T. Loeys and E. Goetghebeur. Baseline information in structural failure time estimators for the effect of observed treatment compliance. Statistics in Medicine, 21(9):1173–1188, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nandram:2002:BAP

[761] Balgobin Nandram and Jai Won Choi. A Bayesian analysis of a proportion under non-ignorable non-response. Statistics in Medicine, 21(9): 1189–1212, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sorribas:2002:NPM

[762] Albert Sorribas, Jaume March, and Javier Trujillano. A new parametric method based on S-distributions for computing receiver operating characteristic curves for continuous diagnostic tests. Statistics in Medicine, 21(9):1213–1235, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2002:PSR

[763] S. D. Walter. Properties of the summary receiver operating characteristic (SROC) curve for diagnostic test data. Statistics in Medicine, 21(9): 1237–1256, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phillips:2002:ECP

[764] Norman Phillips, Andrew Coldman, and Mary L. McBride. Estimating cancer prevalence using mixture models for cancer survival. Statistics in Medicine, 21(9):1257–1270, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schluchter:2002:JMR

[765] Mark D. Schluchter, Michael W. Konstan, and Pamela B. Davis. Jointly modelling the relationship between survival and pulmonary function in cystic fibrosis patients. Statistics in Medicine, 21(9):1271–1287, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Garrett:2002:MEP

[766] Elizabeth S. Garrett, William W. Eaton, and Scott Zeger. Methods for evaluating the performance of diagnostic tests in the absence of a gold standard: a latent class model approach. Statistics in Medicine, 21(9): 1289–1307, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ciampi:2002:NAT

[767] Antonio Ciampi and Fulin Zhang. A new approach to training backpropagation artificial neural networks: empirical evaluation on ten data sets from clinical studies. Statistics in Medicine, 21(9):1309–1330, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bonett:2002:SSR

[768] Douglas G. Bonett. Sample size requirements for estimating intraclass correlations with desired precision. Statistics in Medicine, 21(9):1331– 1335, May 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lake:2002:SSR

[769] Stephen Lake, Erin Kammann, Neil Klar, and Rebecca Betensky. Sample size re-estimation in cluster randomization trials. Statistics in Medicine, 21(10):1337–1350, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoover:2002:CTB

[770] Donald R. Hoover. Clinical trials of behavioural interventions with heterogeneous teaching subgroup effects. Statistics in Medicine, 21(10): 1351–1364, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McMahon:2002:SCV

[771] Alex D. McMahon. Study control, violators, inclusion criteria and defining explanatory and pragmatic trials. Statistics in Medicine, 21(10): 1365–1376, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2002:NAA

[772] Xun Chen, Peng-Liang Zhao, and Ji Zhang. A note on ANOVA assumptions and robust analysis for a cross-over study. Statistics in Medicine, 21(10):1377–1386, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molenberghs:2002:PSO

[773] Geert Molenberghs, Paige L. Williams, and Stuart R. Lipsitz. Prediction of survival and opportunistic infections in HIV-infected patients: a comparison of imputation methods of incomplete CD4 counts. Statistics in Medicine, 21(10):1387–1408, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

King:2002:ERR

[774] Gary King and Langche Zeng. Estimating risk and rate levels, ratios and differences in case-control studies. Statistics in Medicine, 21(10): 1409–1427, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2002:SSA

[775] Wei Pan and Melanie M. Wall. Small-sample adjustments in using the sandwich variance estimator in generalized estimating equations. Statistics in Medicine, 21(10):1429–1441, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:2002:IAC

[776] Douglas J. Taylor, Lawrence L. Kupper, and Keith E. Muller. Improved approximate confidence intervals for the mean of a log-normal random variable. Statistics in Medicine, 21(10):1443–1459, May 30, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2002:ZIM

[777] Yin Bin Cheung. Zero-inflated models for regression analysis of count data: a study of growth and development. Statistics in Medicine, 21(10): 1461–1469, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gautam:2002:AMC

[778] Shiva Gautam. Analysis of mixed categorical data in 2 × K contingency tables. Statistics in Medicine, 21(10):1471–1484, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gomatam:2002:ECR

[779] Shanti Gomatam, Randy Carter, Mario Ariet, and Glenn Mitchell. An empirical comparison of record linkage procedures. Statistics in Medicine, 21(10):1485–1496, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Levy:2002:BRT

[780] Paul S. Levy. Book review: The Theory of the Design of Experiments. D. R. Cox and N. Reid, Chapman and Hall/CRC, Boca Raton, 2000. No. of pages: 323. Price: £39.99. ISBN 1-5-488-195-X. Statistics in Medicine, 21(10):1497–1498, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moreau:2002:BRB

[781] Thierry Moreau. Book review: Biostatistical Methods; The Assessment of Relative Risks. J. M. Lachin. Wiley Series in Probability and Statistics, No. of pages: 544. Price: £58.50. ISBN 0471-36996-9. Statistics in Medicine, 21(10):1498–1499, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiu:2002:BRM

[782] Yen-Feng Chiu. Book review: Multiple Comparisons and Multiple Tests. Using the SAS System. Peter H. Westfall, Randall D.Tobias, Dror Rom, Russell D. Wolfinger and Yosef Hochberg, SAS Institute, Cary, U.S.A. 2000. No. of pages: xiv + 397. Price: DKK 412.00. ISBN 1-58025-397-0. Statistics in Medicine, 21(10):1499–1500, May 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deeks:2002:P

[783] Jonathan J. Deeks, Jesse A. Berlin, and Peter M. Rothwell. Preface. Statistics in Medicine, 21(11):1501–1502, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glasziou:2002:ICH

[784] P. P. Glasziou and S. L. Sanders. Investigating causes of heterogeneity in systematic reviews. Statistics in Medicine, 21(11):1503–1511, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sterne:2002:SMA

[785] Jonathan A. C. Sterne, Peter Jüni, Kenneth F. Schulz, Douglas G. Altman, Christopher Bartlett, and Matthias Egger. Statistical methods for assessing the influence of study characteristics on treatment effects in ‘meta-epidemiological’ research. Statistics in Medicine, 21(11):1513– 1524, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lijmer:2002:ESH

[786] Jeroen G. Lijmer, Patrick M. M. Bossuyt, and Siem H. Heisterkamp. Exploring sources of heterogeneity in systematic reviews of diagnostic tests. Statistics in Medicine, 21(11):1525–1537, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Higgins:2002:QHM

[787] Julian P. T. Higgins and Simon G. Thompson. Quantifying heterogeneity in a meta-analysis. Statistics in Medicine, 21(11):1539–1558, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2002:HSM

[788] Simon G. Thompson and Julian P. T. Higgins. How should metaregression analyses be undertaken and interpreted? Statistics in Medicine, 21(11):1559–1573, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deeks:2002:ISS

[789] Jonathan J. Deeks. Issues in the selection of a summary statistic for metaanalysis of clinical trials with binary outcomes. Statistics in Medicine, 21(11):1575–1600, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Warn:2002:BRE

[790] D. E. Warn, S. G. Thompson, and D. J. Spiegelhalter. Bayesian random effects meta-analysis of trials with binary outcomes: methods for the absolute risk difference and relative risk scales. Statistics in Medicine, 21(11):1601–1623, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1799].

Hopewell:2002:CHV

[791] S. Hopewell, M. Clarke, A. Lusher, C. Lefebvre, and M. Westby. A comparison of handsearching versus MEDLINE searching to identify reports of randomized controlled trials. Statistics in Medicine, 21(11):1625–1634, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:2002:IRC

[792] Phil Edwards, Mike Clarke, Carolyn DiGuiseppi, Sarah Pratap, Ian Roberts, and Reinhard Wentz. Identification of randomized controlled trials in systematic reviews: accuracy and reliability of screening records. Statistics in Medicine, 21(11):1635–1640, June 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dawson:2002:UIW

[793] Ree Dawson and Philip W. Lavori. Using inverse weighting and predictive inference to estimate the effects of time-varying treatments on the discrete-time hazard. Statistics in Medicine, 21(12):1641–1661, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:2002:CUI

[794] James M. Robins. Commentary on ‘using inverse weighting and predictive inference to estimate the effects of time-varying treatments on the discrete-time hazard’. Statistics in Medicine, 21(12):1663–1680, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dawson:2002:RUI

[795] Ree Dawson and Philip W. Lavori. Rejoinder for ‘using inverse weighting and predictive inference to estimate the effects of time-varying treatments on the discrete-time hazard’. Statistics in Medicine, 21(12):1681–1687, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hernan:2002:ECE

[796] Miguel A. Hernán, Babette A. Brumback, and James M. Robins. Estimating the causal effect of zidovudine on CD4 count with a marginal structural model for repeated measures. Statistics in Medicine, 21(12): 1689–1709, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steinberg:2002:ESR

[797] Seth M. Steinberg and David J. Venzon. Early selection in a randomized phase II clinical trial. Statistics in Medicine, 21(12):1711–1726, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2002:RPC

[798] Jun Shao and Shein-Chung Chow. Reproducibility probability in clinical trials. Statistics in Medicine, 21(12):1727–1742, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perkins:2002:ARA

[799] Susan M. Perkins and Mark P. Becker. Assessing rater agreement using marginal association models. Statistics in Medicine, 21(12):1743–1760, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kirchner:2002:SEI

[800] H. Lester Kirchner and Jon H. Lemke. Simultaneous estimation of intrarater and interrater agreement for multiple raters under order restrictions for a binary trait. Statistics in Medicine, 21(12):1761–1772, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gunnarsson:2002:PVM

[801] Ronny K. Gunnarsson and Jan Lanke. The predictive value of microbiologic diagnostic tests if asymptomatic carriers are present. Statistics in Medicine, 21(12):1773–1785, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1053].

Liu:2002:SSP

[802] Aiyi Liu, W. J. Shih, and E. Gehan. Sample size and power determination for clustered repeated measurements. Statistics in Medicine, 21(12): 1787–1801, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:2002:LEC

[803] Colin B. Begg. Letter to the Editor: A comparison of methods to detect publication bias in meta-analysis by P. Macaskill, S. D. Walter and L. Irwig, Statistics in Medicine, 2001; 20:641–654. Statistics in Medicine, 21(12):1803, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [385].

Macaskill:2002:AR

[804] Petra Macaskill, Les Irwig, and Stephen D. Walter. Authors’ reply. Statistics in Medicine, 21(12):1804, June 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Potter:2002:ADF

[805] Douglas M. Potter. Adaptive dose finding for phase I clinical trials of drugs used for chemotherapy of cancer. Statistics in Medicine, 21(13): 1805–1823, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quiroz:2002:ACI

[806] Jorge Quiroz, Naitee Ting, Greg C. G. Wei, and Richard K. Burdick. Alternative confidence intervals for the assessment of bioequivalence in four-period cross-over designs. Statistics in Medicine, 21(13):1825–1847, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2002:ROC

[807] Ying Lu, Daniel N. Heller, and Shoujun Zhao. Receiver operating characteristic (ROC) analysis for diagnostic examinations with uninterpretable cases. Statistics in Medicine, 21(13):1849–1865, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fijal:2002:AMS

[808] Bonnie A. Fijal, Ramana M. Idury, and John S. Witte. Analysis of mutational spectra: locating hotspots and clusters of mutations using recursive segmentation. Statistics in Medicine, 21(13):1867–1885, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Winawer:2002:CDF

[809] Melodie Winawer, Ruth Ottman, and Daniel Rabinowitz. Concordance of disease form in kindreds ascertained through affected individuals. Statistics in Medicine, 21(13):1887–1897, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aguirre-Hernandez:2002:PTG

[810] R. Aguirre-Hernández and V. T. Farewell. A Pearson-type goodness-of-fit test for stationary and time-continuous Markov regression models. Statistics in Medicine, 21(13):1899–1911, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hawkins:2002:DCP

[811] Douglas M. Hawkins. Diagnostics for conformity of paired quantitative measurements. Statistics in Medicine, 21(13):1913–1935, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lo:2002:PTD

[812] Yungtai Lo, Steven Matthysse, Donald B. Rubin, and Philip S. Holzman. Permutation tests for detecting and estimating mixtures in task performance within groups. Statistics in Medicine, 21(13):1937–1953, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2002:LEP

[813] Xun Chen. Letter to the Editor: Power comparison of robust approximate and non-parametric tests for the analysis of cross-over trials by J. A. Correa and F. Bellavance, Statistics in Medicine 2001; 20:1185– 1196. Statistics in Medicine, 21(13):1955–1956, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [444].

Correa:2002:AR

[814] Jose A. Correa and François Bellavance. Author’s reply. Statistics in Medicine, 21(13):1956–1958, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schoder:2002:LET

[815] Volker Schoder. Letter to the Editor: Test statistics and sample size formulae for comparative binomial trials with null hypothesis of non-zero risk difference or non-unity relative risk by C. P. Farrington and G. Manning, Statistics in Medicine 1990; 9:1447–1454. Statistics in Medicine, 21(13):1958–1960, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2].

Farrington:2002:AR

[816] C. P. Farrington. Author’s reply. Statistics in Medicine, 21(13):1960, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2002:BRD

[817] Per Kragh Andersen. Book review: 1. Design and Analysis of Cluster Randomization Trials in Health Research. Allan Donner and Neil Klar, Arnold, London, 2000. No. of pages: x + 178. Price: £35. ISBN 0-340-691530. Statistics in Medicine, 21(13):1961–1962, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knudsen:2002:BRB

[818] Lisbeth B. Knudsen. Book review: 2. Birth Counts. Statistics of Pregnancy and Childbirth. 2nd edn. Volumes 1 and 2. Alison Macfarlane and Miranda Mugford, The Stationery Office, London, 2000. No. of pages: xxv + 447 and xxx + 673. Price: £29.95 per volume or £45.00 for both. ISBN 0-11-621049-4 (volume 1) and 0-11-620917-8 (volume 2). Statistics in Medicine, 21(13):1962–1963, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2002:BRA

[819] Per Kragh Andersen. Book review: 3. Applied Logistic Regression. 2nd edn. David W. Hosmer and Stanley Lemeshow. Wiley, New York, 2000. No. of pages: xii + 373. Price: £60.95. ISBN 0-471-35632-8. Statistics in Medicine, 21(13):1963–1964, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skrondal:2002:BRM

[820] Anders Skrondal. Book review: 4. Multilevel Modelling of Health Statistics. A. H. Leyland and H. Goldstein (eds), Wiley, Chichester, 2001. No. of pages: xvii + 217. Price $95.00. ISBN 0-471-99890-7. Statistics in Medicine, 21(13):1964–1966, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2002:BRM

[821] Roger Marshall. Book review: 5. Multidimensional Scaling. 2nd edn. Trevor F. Cox and Michael A. A. Cox, Chapman and Hall/CRC, Boca Raton, London, New York, Washington DC, 2000. No. of pages: xiv + 309. Price: $79.95. ISBN 1-58488-094-5. Statistics in Medicine, 21(13): 1966–1967, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lakatos:2002:DCG

[822] Edward Lakatos. Designing complex group sequential survival trials. Statistics in Medicine, 21(14):1969–1989, July 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2002:BTS

[823] Say-Beng Tan and David Machin. Bayesian two-stage designs for phase II clinical trials. Statistics in Medicine, 21(14):1991–2012, July 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2160].

Williamson:2002:MAM

[824] P. R. Williamson, G. A. Lancaster, J. V. Craig, and R. L. Smyth. Metaanalysis of method comparison studies. Statistics in Medicine, 21(14): 2013–2025, July 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanderHeijden:2002:AAD

[825] Peter G. M. van der Heijden, Eugène P. van Puijenbroek, Stef van Buuren, and Jacques W. van der Hofstede. On the assessment of adverse drug reactions from spontaneous reporting systems: the influence of under-reporting on odds ratios. Statistics in Medicine, 21(14):2027–2044, July 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2002:CBS

[826] Xiao-Hua Zhou and Jaroslaw Harezlak. Comparison of bandwidth selection methods for kernel smoothing of ROC curves. Statistics in Medicine, 21(14):2045–2055, July 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Assuncao:2002:BSV

[827] Renato M. Assunção, Joseph E. Potter, and Suzana M. Cavenaghi. A Bayesian space varying parameter model applied to estimating fertility schedules. Statistics in Medicine, 21(14):2057–2075, July 30, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wainwright:2002:TVE

[828] Nicholas W. J. Wainwright and Paul G. Surtees. Time-varying exposure and the impact of stressful life events on onset of affective disorder. Statistics in Medicine, 21(14):2077–2091, July 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fitzgerald:2002:MHV

[829] Anthony P. Fitzgerald, Victor G. DeGruttola, and Florin Vaida. Modelling HIV viral rebound using non-linear mixed effects models. Statistics in Medicine, 21(14):2093–2108, July 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:2002:KCM

[830] Helena Chmura Kraemer, Vyjeyanthi S. Periyakoil, and Art Noda. Kappa coefficients in medical research. Statistics in Medicine, 21(14): 2109–2129, July 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curtin:2002:MACa

[831] François Curtin, Douglas G. Altman, and Diana Elbourne. Meta-analysis combining parallel and cross-over clinical trials. i: Continuous outcomes. Statistics in Medicine, 21(15):2131–2144, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curtin:2002:MACb

[832] François Curtin, Diana Elbourne, and Douglas G. Altman. Meta-analysis combining parallel and cross-over clinical trials. II: Binary outcomes. Statistics in Medicine, 21(15):2145–2159, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curtin:2002:MACc

[833] François Curtin, Diana Elbourne, and Douglas G. Altman. Meta-analysis combining parallel and cross-over clinical trials. III: The issue of carryover. Statistics in Medicine, 21(15):2161–2173, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2002:FPP

[834] Patrick Royston and Mahesh K. B. Parmar. Flexible parametric proportional-hazards and proportional-odds models for censored survival data, with application to prognostic modelling and estimation of treatment effects. Statistics in Medicine, 21(15):2175–2197, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Putter:2002:BAP

[835] H. Putter, S. H. Heisterkamp, J. M. A. Lange, and F. de Wolf. A Bayesian approach to parameter estimation in HIV dynamical models. Statistics in Medicine, 21(15):2199–2214, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2002:RCN

[836] Yangxin Huang. Robustness of choice of number of doses for maximum likelihood estimation of the ED_{50} in bioassay. Statistics in Medicine, 21 (15):2215–2223, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schumann:2002:PFS

[837] Agnes Schumann, Niels Wessel, Alexander Schirdewan, Karl Josef Osterziel, and Andreas Voss. Potential of feature selection methods in heart rate variability analysis for the classification of different cardiovascular diseases. Statistics in Medicine, 21(15):2225–2242, August 15, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2002:IMF

[838] Bee-Choo Tai, Ian R. White, Val Gebski, and David Machin. On the issue of ‘multiple’ first failures in competing risks analysis. Statistics in Medicine, 21(15):2243–2255, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sasahara:2002:TDM

[839] Hisako Sasahara, Hideo Sasahara, and Makoto Kawamura. A threedimensional model for the tooth-loss patterns by multi-plane regression analysis. Statistics in Medicine, 21(15):2257–2269, August 15, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoogenveen:2002:DPE

[840] Rudolf Hoogenveen, Gert Westert, Marcel Dijkgraaf, François Schellevis, and Dinny de Bakker. Disease prevalence estimations based on contact registrations in general practice. Statistics in Medicine, 21(15):2271– 2285, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:2002:SSP

[841] Tim J. Cole and Yuri V. Kryakin. Sympercents: symmetric percentage differences on the 100 log_{e} scale simplify the presentation of log transformed data by T. J. Cole, Statistics in Medicine 2000; 19: 3109–3125. Statistics in Medicine, 21(15):2287–2290, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [295].

Tutz:2002:BRB

[842] Gerhard Tutz. Book review: A Basic Course in Statistics, 4th edn, G. M. Clarke and D. Cooke, Arnold, 1997. No. of pages:xxiii + 672.Price: £ 19.99. ISBN 0-340-71995-8. Statistics in Medicine, 21(15):2291–2292, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Didelez:2002:BRC

[843] Vanessa Didelez. Book review: Causality: Models, Reasoning and Inference. Judea Pearl, Cambridge University Press, New York, 2000. No. of pages: 384. Price: $39.95. ISBN 0-521-77362-8. Statistics in Medicine, 21(15):2292–2293, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoffmann:2002:BRH

[844] Raymond Hoffmann. Book review: A Handbook of Statistical Analyses Using Stata, 2nd edn, Sophia Rabe-Hesketh and Brian Everitt, Chapman & Hall/CRC, London, U.K., 1998. No. of pages: xi + 215. Price: £19.95. ISBN 0-849-30387-7. Statistics in Medicine, 21(15):2293–2294, August 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leonard:2002:BFE

[845] Thomas Leonard and John C. Duffy. A Bayesian fixed effects analysis of the Mantel–Haenszel model applied to meta-analysis. Statistics in Medicine, 21(16):2295–2312, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lumley:2002:NMA

[846] Thomas Lumley. Network meta-analysis for indirect treatment comparisons. Statistics in Medicine, 21(16):2313–2324, August 30, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2002:FEV

[847] Y. Lee. Fixed-effect versus random-effect models for evaluating therapeutic preferences. Statistics in Medicine, 21(16):2325–2330, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grambsch:2002:PPD

[848] Patricia Grambsch, William H. Meller, and P. Victor Grambsch. Periodograms and pulse detection methods for pulsatile hormone data. Statistics in Medicine, 21(16):2331–2344, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hatteville:2002:PLT

[849] Laurence Hatteville, Cédric Mahe, and Catherine Hill. Prediction of the long-term survival in breast cancer patients according to the present oncological status. Statistics in Medicine, 21(16):2345–2354, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eichner:2002:ETD

[850] M. Eichner, H. H. Diebner, C. Schubert, H. W. Kreth, and K. Dietz. Estimation of the time-dependent vaccine efficacy from a measles epidemic. Statistics in Medicine, 21(16):2355–2368, August 30, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2002:MLE

[851] Haiqun Lin, Charles E. McCulloch, and Susan T. Mayne. Maximum likelihood estimation in the joint analysis of time-to-event and multiple longitudinal variables. Statistics in Medicine, 21(16):2369–2382, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:2002:ESO

[852] Stuart G. Baker and Melvyn S. Tockman. Evaluating serial observations of precancerous lesions for further study as a trigger for early intervention. Statistics in Medicine, 21(16):2383–2390, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacqmin-Gadda:2002:PLA

[853] Hélène Jacqmin-Gadda, Pierre Joly, Daniel Commenges, Christine Binquet, and Geneviève Chêne. Penalized likelihood approach to estimate a smooth mean curve on longitudinal data. Statistics in Medicine, 21 (16):2391–2402, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mukherjee:2002:ATD

[854] Ajit Mukherjee, Shiv Kumar, Rakesh Mittal, and N. C. Saxena. An application of the technique of density estimation in geographical epidemiology of goitre in the Muzaffarpur district of Bihar state of India. Statistics in Medicine, 21(16):2403–2407, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinze:2002:SPS

[855] Georg Heinze and Michael Schemper. A solution to the problem of separation in logistic regression. Statistics in Medicine, 21(16):2409–2419, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parzen:2002:WEE

[856] Michael Parzen, Stuart R. Lipsitz, Joseph G. Ibrahim, and Steven Lipshultz. A weighted estimating equation for linear regression with missing covariate data. Statistics in Medicine, 21(16):2421–2436, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2002:LEC

[857] Stephen Senn. Letter to the Editor: A comment on replication, p-values and evidence, S. N. Goodman, Statistics in Medicine 1992; 11:875–879. Statistics in Medicine, 21(16):2437–2444, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [4].

Goodman:2002:AR

[858] Steven N. Goodman. Author’s reply. Statistics in Medicine, 21(16):2445– 2447, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Branson:2002:ETE

[859] Michael Branson and John Whitehead. Estimating a treatment effect in survival studies in which patients switch treatment. Statistics in Medicine, 21(17):2449–2463, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwarzer:2002:ITE

[860] Guido Schwarzer, Gerd Antes, and Martin Schumacher. Inflation of type I error rate in two statistical tests for the detection of publication bias in meta-analyses with binary outcomes. Statistics in Medicine, 21(17): 2465–2477, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2002:MTS

[861] Jonathan Shuster, Michael Link, Bruce Camitta, Jeanette Pullen, and Fredrick Behm. Minimax two-stage-designs with applications to tissue banking case-control studies. Statistics in Medicine, 21(17):2479–2493, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williams:2002:BRE

[862] K. L. Williams and A. B. J. Nix. Bias in risk estimation: application to Down’s syndrome screening. Statistics in Medicine, 21(17):2495–2509, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giard:2002:MMG

[863] Nicole Giard, Paul Lichtenstein, and Anatoli I. Yashin. A multistate model for the genetic analysis of the ageing process. Statistics in Medicine, 21(17):2511–2526, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Macaskill:2002:AGD

[864] Petra Macaskill, Stephen D. Walter, Les Irwig, and Eduardo L. Franco. Assessing the gain in diagnostic performance when combining two diagnostic tests. Statistics in Medicine, 21(17):2527–2546, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:2002:MSA

[865] Geoffrey Jones and David M. Rocke. Multivariate survival analysis with doubly-censored data: application to the assessment of accutane treatment for fibrodysplasia ossificans progressiva. Statistics in Medicine, 21 (17):2547–2562, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blume:2002:LMM

[866] Jeffrey D. Blume. Likelihood methods for measuring statistical evidence. Statistics in Medicine, 21(17):2563–2599, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2002:LER

[867] Chii-Dean Lin and Kung-Jong Lui. Letter to the Editor: Randomized consent designs for clinical trials: an update by M. Zelen, Statistics in Medicine 1990; 9: 645–656. Statistics in Medicine, 21(17):2601–2603, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1].

Zelen:2002:AR

[868] Marvin Zelen. Authors’ reply. Statistics in Medicine, 21(17):2603–2605, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:2002:BRE

[869] Jørgen Holm Petersen. Book review: Environmental Statistics with S-Plus. Steven P. Millard and Nagaraj K. Neerchal. CRC Press LLC, London, 2001. No. of pages: 830. Price: $79.95. ISBN 0-0893-7168-6. Statistics in Medicine, 21(17):2607–2608, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1169].

Rogind:2002:BRS

[870] Henrik Rogind. Book review: Statistics for People Who (Think They) Hate Statistics. Neil J. Salkind, Sage Publications, London, 2000. No. of pages: xxi + 385. Price: £44.00. ISBN 0-7619-1621-0. Statistics in Medicine, 21(17):2608–2609, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2002:BRM

[871] Paula Williamson. Book review: Meta-Analysis in Medicine and Health Policy. Dalene K. Stangl and Donald A. Berry (eds), Marcel Dekker, New York, 2000. No. of pages: 398. Price: $155.00. ISBN 0-8247-9030-8. Statistics in Medicine, 21(17):2610–2611, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pinna:2002:BRF

[872] Gian Domenico Pinna. Book review: Fundamentals of Modern Statistical Methods. Rand R. Wilcox, Springer-Verlag, New York, 2001. No. of pages: xiii + 258. Price: $49.95. ISBN 0-387-95157-1. Statistics in Medicine, 21 (17):2611–2612, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2002:SIA

[873] Paula R. Williamson, Helen E. Clough, Jane L. Hutton, Anthony G. Marson, and David W. Chadwick. Statistical issues in the assessment of the evidence for an interaction between factors in epilepsy trials. Statistics in Medicine, 21(18):2613–2622, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Retout:2002:FIM

[874] Sylvie Retout, France Mentré, and René Bruno. Fisher information matrix for non-linear mixed-effects models: evaluation and application for optimal design of enoxaparin population pharmacokinetics. Statistics in Medicine, 21(18):2623–2639, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baujat:2002:GME

[875] Bertrand Baujat, Cédric Mahé, Jean-Pierre Pignon, and Catherine Hill. A graphical method for exploring heterogeneity in meta-analyses: application to a meta-analysis of 65 trials. Statistics in Medicine, 21(18): 2641–2652, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Black:2002:EDP

[876] Michael A. Black and Bruce A. Craig. Estimating disease prevalence in the absence of a gold standard. Statistics in Medicine, 21(18):2653– 2669, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:2002:SCE

[877] Ron Brookmeyer and Frank C. Curriero. Survival curve estimation with partial non-random exposure information. Statistics in Medicine, 21(18): 2671–2683, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gather:2002:GMM

[878] Ursula Gather, Michael Imhoff, and Roland Fried. Graphical models for multivariate time series from intensive care monitoring. Statistics in Medicine, 21(18):2685–2701, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mugglin:2002:HSM

[879] Andrew S. Mugglin, Noel Cressie, and Islay Gemmell. Hierarchical statistical modelling of influenza epidemic dynamics in space and time. Statistics in Medicine, 21(18):2703–2721, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hosmer:2002:GFP

[880] David W. Hosmer and Nils Lid Hjort. Goodness-of-fit processes for logistic regression: simulation results. Statistics in Medicine, 21(18):2723– 2738, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ambrosius:2002:ISC

[881] Walter T. Ambrosius and Amita K. Manatunga. Intensive short courses in biostatistics for fellows and physicians. Statistics in Medicine, 21(18): 2739–2756, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberger:2002:CDP

[882] William F. Rosenberger and Linda M. Haines. Competing designs for phase I clinical trials: a review. Statistics in Medicine, 21(18):2757– 2770, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2002:C

[883] Andrew Lawson. Correction. Statistics in Medicine, 21(18):2771, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krickeberg:2002:BRD

[884] Klaus Krickeberg. Book review: Design and implementation of health information systems. Theo Lippeveld, Rainer Sauerborn and Claude Bodart (eds), World Health Organization, Geneva, 2000. No. of pages: xi + 270. Price: Sw. fr. 70. ISBN 92-4-156199-8. Statistics in Medicine, 21 (18):2772–2773, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Britton:2002:BRE

[885] Tom Britton. Book review: Epidemic modelling: an introduction. Daryl J. Daley and Joe Gani, Cambridge University Press, U.K., 1999. No. of pages: xii + 213. Price: £30.00. ISBN 0-521-64079-2. Statistics in Medicine, 21(18):2773–2774, September 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knatterud:2002:P

[886] Genell Knatterud, Marcus Flather, Michael Terrin, and Stuart Pocock. Preface. Statistics in Medicine, 21(19):2775–2776, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeMets:2002:CTN

[887] David L. DeMets. Clinical trials in the new millennium. Statistics in Medicine, 21(19):2779–2787, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:2002:CLT

[888] Janet Wittes. On changing a long-term clinical trial midstream. Statistics in Medicine, 21(19):2789–2795, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Warlow:2002:AID

[889] Charles Warlow. Advanced issues in the design and conduct of randomized clinical trials: the bigger the better? Statistics in Medicine, 21 (19):2797–2805, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hampton:2002:SIE

[890] J. R. Hampton. Size isn’t everything. Statistics in Medicine, 21(19):2807– 2814, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laupacis:2002:CCH

[891] Andreas Laupacis. The Cochrane collaboration — how is it progressing? Statistics in Medicine, 21(19):2815–2822, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilhelmsen:2002:RDS

[892] Lars Wilhelmsen. Role of the data and safety monitoring committee (DSMC). Statistics in Medicine, 21(19):2823–2829, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ONeill:2002:RPD

[893] Robert T. O’Neill. Regulatory perspectives on data monitoring. Statistics in Medicine, 21(19):2831–2842, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleming:2002:MCT

[894] Thomas R. Fleming, Susan Ellenberg, and David L. DeMets. Monitoring clinical trials: issues and controversies regarding confidentiality. Statistics in Medicine, 21(19):2843–2851, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brody:2002:EIC

[895] Baruch A. Brody. Ethical issues in clinical trials in developing countries. Statistics in Medicine, 21(19):2853–2858, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yusuf:2002:CRT

[896] Salim Yusuf. Clinical research and trials in developing countries. Statistics in Medicine, 21(19):2859–2867, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2002:PRD

[897] Richard Smith. Publishing research from developing countries. Statistics in Medicine, 21(19):2869–2877, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hlatky:2002:EEL

[898] Mark A. Hlatky, Derek B. Boothroyd, and Iain M. Johnstone. Economic evaluation in long-term clinical trials. Statistics in Medicine, 21 (19):2879–2888, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Califf:2002:ECE

[899] Robert M. Califf. Evaluating the costs and effectiveness of cardiovascular therapies: who cares about economic analyses? Statistics in Medicine, 21 (19):2889–2897, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ford:2002:RCE

[900] Ian Ford and John Norrie. The role of covariates in estimating treatment effects and risk in long-term clinical trials. Statistics in Medicine, 21 (19):2899–2908, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simon:2002:BSA

[901] Richard Simon. Bayesian subset analysis: application to studying treatment-by-gender interactions. Statistics in Medicine, 21(19):2909– 2916, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:2002:SAC

[902] Stuart J. Pocock, Susan E. Assmann, Laura E. Enos, and Linda E. Kasten. Subgroup analysis, covariate adjustment and baseline comparisons in clinical trial reporting: current practiceand problems. Statistics in Medicine, 21(19):2917–2930, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:2002:ERE

[903] John A. Lewis. The European regulatory experience. Statistics in Medicine, 21(19):2931–2938, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Temple:2002:PDR

[904] Robert Temple. Policy developments in regulatory approval. Statistics in Medicine, 21(19):2939–2948, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rockhold:2002:IPI

[905] Frank W. Rockhold. Industry perspectives on ICH guidelines. Statistics in Medicine, 21(19):2949–2957, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lubsen:2002:CEC

[906] Jacobus Lubsen and Bridget-Anne Kirwan. Combined endpoints: can we use them? Statistics in Medicine, 21(19):2959–2970, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:2002:IMA

[907] Allan Donner and Neil Klar. Issues in the meta-analysis of cluster randomized trials. Statistics in Medicine, 21(19):2971–2980, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2002:IDB

[908] Bruce Thompson, Rolfe G. Jacob, Margaret Frederick, Nancy V. Geller, and Sally Hunsberger. Innovative designs in behavioural trials. Statistics in Medicine, 21(19):2981–2989, October 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armstrong:2002:EFN

[909] Paul W. Armstrong. The ebb and flow of the NHS waiting list: how do recruitment and admission affect event-based measures of the length of ‘time-to-admission’? Statistics in Medicine, 21(20):2991–3009, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1424].

Manda:2002:BOM

[910] Samuel O. M. Manda. A Bayesian ordinal model for heterogeneity in a multi-centre myocardial infarction clinical trial. Statistics in Medicine, 21 (20):3011–3022, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:2002:DMC

[911] J. K. Lindsey, J. W. Stevens, B. Jones, and D. Jackson. Dynamic modelling of a challenge-escalation cross-over study of treatment of capsaicininduced coughing. Statistics in Medicine, 21(20):3023–3033, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Preisser:2002:PWE

[912] John S. Preisser, Kurt K. Lohman, and Paul J. Rathouz. Performance of weighted estimating equations for longitudinal binary data with dropouts missing at random. Statistics in Medicine, 21(20):3035–3054, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heidenreich:2002:MML

[913] W. F. Heidenreich, J. Wellmann, P. Jacob, and H. E. Wichmann. Mechanistic modelling in large case-control studies of lung cancer risk from smoking. Statistics in Medicine, 21(20):3055–3070, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hansson:2002:RNE

[914] Sven Ove Hansson. Replacing the no-effect level (NOEL) with bounded effect levels (OBEL and LEBEL). Statistics in Medicine, 21(20):3071– 3078, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:2002:IEP

[915] C. P. Farrington. Interval estimation for Poisson capture-recapture models in epidemiology. Statistics in Medicine, 21(20):3079–3092, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraggi:2002:EAU

[916] David Faraggi and Benjamin Reiser. Estimation of the area under the ROC curve. Statistics in Medicine, 21(20):3093–3106, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2002:IEG

[917] Kung-Jong Lui. Interval estimation of generalized odds ratio in data with repeated measurements. Statistics in Medicine, 21(20):3107–3117, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gannoun:2002:RCB

[918] Ali Gannoun, Stéphane Girard, Christiane Guinot, and Jérôme Saracco. Reference curves based on non-parametric quantile regression. Statistics in Medicine, 21(20):3119–3135, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2002:PAC

[919] Hansheng Wang and Shein-Chung Chow. A practical approach for comparing means of two groups without equal variance assumption. Statistics in Medicine, 21(20):3137–3151, October 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sidik:2002:SCI

[920] Kurex Sidik and Jeffrey N. Jonkman. A simple confidence interval for meta-analysis. Statistics in Medicine, 21(21):3153–3159, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1168].

Jo:2002:MMS

[921] Booil Jo. Model misspecification sensitivity analysis in estimating causal effects of interventions with non-compliance. Statistics in Medicine, 21 (21):3161–3181, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinavaara:2002:CSS

[922] S. Heinävaara, L. Teppo, and T. Hakulinen. Cancer-specific survival of patients with multiple cancers: an application to patients with multiple breast cancers. Statistics in Medicine, 21(21):3183–3195, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2002:CBM

[923] Philippe Lambert and François Vandenhende. A copula-based model for multivariate non-normal longitudinal data: analysis of a dose titration safety study on a new antidepressant. Statistics in Medicine, 21(21): 3197–3217, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OQuigley:2002:PHM

[924] John O’Quigley and Janez Stare. Proportional hazards models with frailties and random effects. Statistics in Medicine, 21(21):3219–3233, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2002:SPA

[925] Chin-Shang Li and Jeremy M. G. Taylor. A semi-parametric accelerated failure time cure model. Statistics in Medicine, 21(21):3235–3247, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byrnes:2002:MMS

[926] G. B. Byrnes. A Markov model for sample size calculation and inference in vaccine cost-effectiveness studies. Statistics in Medicine, 21(21):3249– 3260, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heid:2002:PME

[927] I. M. Heid, H. Küchenhoff, J. Wellmann, M. Gerken, L. Kreienbrock, and H. E. Wichmann. On the potential of measurement error to induce differential bias on odds ratio estimates: an example from radon epidemiology. Statistics in Medicine, 21(21):3261–3278, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frankel:2002:PMA

[928] Paul Frankel and Jeffrey Longmate. Parametric models for accelerated and long-term survival: a comment on proportional hazards. Statistics in Medicine, 21(21):3279–3289, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldstein:2002:MMM

[929] Harvey Goldstein, William Browne, and Jon Rasbash. Multilevel modelling of medical data. Statistics in Medicine, 21(21):3291–3315, November 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pintilie:2002:DCR

[930] Melania Pintilie. Dealing with competing risks: testing covariates and calculating sample size. Statistics in Medicine, 21(22):3317–3324, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2002:DDR

[931] Frank Bretz and Ludwig A. Hothorn. Detecting dose-response using contrasts: asymptotic power and sample size determination for binomial data. Statistics in Medicine, 21(22):3325–3335, November 30, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2002:ADM

[932] Paula R. Williamson, Catrin Tudur Smith, Jane L. Hutton, and Anthony G. Marson. Aggregate data meta-analysis with time-to-event outcomes. Statistics in Medicine, 21(22):3337–3351, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:2002:RAB

[933] Dean A. Follmann. Regression analysis based on pairwise ordering of patients’ clinical histories. Statistics in Medicine, 21(22):3353–3367, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:2002:ACD

[934] Hsin-Ju Hsieh, Tony Hsiu-Hsi Chen, and Shu-Hui Chang. Assessing chronic disease progression using non-homogeneous exponential regression Markov models: an illustration using a selective breast cancer screening in Taiwan. Statistics in Medicine, 21(22):3369–3382, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dewanji:2002:CSP

[935] Anup Dewanji and Suresh H. Moolgavkar. Choice of stratification in Poisson process analysis of recurrent event data with environmental covariates. Statistics in Medicine, 21(22):3383–3393, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abdolell:2002:BPC

[936] M. Abdolell, M. LeBlanc, D. Stephens, and R. V. Harrison. Binary partitioning for continuous longitudinal data: categorizing a prognostic variable. Statistics in Medicine, 21(22):3395–3409, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chan:2002:USM

[937] Ivan S. F. Chan, Shu Li, Holly Matthews, Christina Chan, Rupert Vessey, Jerald Sadoff, and Joseph Heyse. Use of statistical models for evaluating antibody response as a correlate of protection against varicella. Statistics in Medicine, 21(22):3411–3430, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rousson:2002:AII

[938] Valentin Rousson, Theo Gasser, and Burkhardt Seifert. Assessing intrarater, interrater and test-retest reliability of continuous measurements. Statistics in Medicine, 21(22):3431–3446, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pfeiffer:2002:TAM

[939] Ruth M. Pfeiffer, Efstathia Bura, Amelia Smith, and Joni L. Rutter. Two approaches to mutation detection based on functional data. Statistics in Medicine, 21(22):3447–3464, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2002:BPC

[940] Aiyi Liu, Ying Zhang, Edmund Gehan, and Robert Clarke. Block principal component analysis with application to gene microarray data classification. Statistics in Medicine, 21(22):3465–3474, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bebchuk:2002:LLA

[941] Judith D. Bebchuk and Rebecca A. Betensky. Local likelihood analysis of the latency distribution with interval censored intermediate events. Statistics in Medicine, 21(22):3475–3491, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Orbe:2002:CPH

[942] Jesus Orbe, Eva Ferreira, and Vicente Núñez-Antón. Comparing proportional hazards and accelerated failure time models for survival analysis. Statistics in Medicine, 21(22):3493–3510, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verdecchia:2002:EPC

[943] Arduino Verdecchia, Giovanni De Angelis, and Riccardo Capocaccia. Estimation and projections of cancer prevalence from cancer registry data. Statistics in Medicine, 21(22):3511–3526, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OGorman:2002:ATS

[944] Thomas W. O’Gorman. An adaptive test of significance for a subset of regression coefficients. Statistics in Medicine, 21(22):3527–3542, November 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2002:PSS

[945] Mei-Ling Ting Lee and G. A. Whitmore. Power and sample size for DNA microarray studies. Statistics in Medicine, 21(23):3543–3570, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Truong:2002:ODR

[946] Young K. Truong, Rona S. Scott, and Jean-Michel H. Vos. The origin of DNA replication and Fieller’s problem. Statistics in Medicine, 21(23): 3571–3582, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fung:2002:EMS

[947] Wing K. Fung and Yue-Qing Hu. Evaluating mixed stains with contributors of different ethnic groups under the NRC-II recommendation 4.1. Statistics in Medicine, 21(23):3583–3593, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2002:AMS

[948] Elisabeth Wreford Andersen and Per Kragh Andersen. Adjustment for misclassification in studies of familial aggregation of disease using routine register data. Statistics in Medicine, 21(23):3595–3607, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McNamee:2002:ODT

[949] Roseanne McNamee. Optimal designs of two-stage studies for estimation of sensitivity, specificity and positive predictive value. Statistics in Medicine, 21(23):3609–3625, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookhart:2002:SEP

[950] M. Alan Brookhart, Alan E. Hubbard, Mark J. van der Laan, John M. Colford, Jr., and Joseph N. S. Eisenberg. Statistical estimation of parameters in a disease transmission model: analysis of a cryptosporidium outbreak. Statistics in Medicine, 21(23):3627–3638, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2002:HPM

[951] K. Wang, Kelvin K. W. Yau, and Andy H. Lee. A hierarchical Poisson mixture regression model to analyse maternity length of hospital stay. Statistics in Medicine, 21(23):3639–3654, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2002:SLT

[952] Hulin Wu and Jin-Ting Zhang. The study of long-term HIV dynamics using semi-parametric non-linear mixed-effects models. Statistics in Medicine, 21(23):3655–3675, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goedhart:2002:SBM

[953] Dick M. Goedhart and Pieter Zanen. Selecting the best method to evaluate bronchodilation when analysing bronchodilator studies. Statistics in Medicine, 21(23):3677–3685, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keene:2002:AHR

[954] Oliver N. Keene. Alternatives to the hazard ratio in summarizing efficacy in time-to-event studies: an example from influenza trials. Statistics in Medicine, 21(23):3687–3700, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liestol:2002:UCC

[955] Knut Liestøl and Per Kragh Andersen. Updating of covariates and choice of time origin in survival analysis: problems with vaguely defined disease states. Statistics in Medicine, 21(23):3701–3714, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2002:AMI

[956] Mimi Y. Kim and Xiaonan Xue. The analysis of multivariate intervalcensored survival data. Statistics in Medicine, 21(23):3715–3726, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yau:2002:ACI

[957] C. Lillian Yau and Aparna V. Huzurbazar. Analysis of censored and incomplete survival data using flowgraph models. Statistics in Medicine, 21(23):3727–3743, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2002:LEC

[958] Devan V. Mehrotra. Letter to the Editor: A comparison of generalized linear mixed model procedures with estimating equations for variance and covariance parameter estimation in longitudinal studies and group randomized trials by B. A. Evans, Z. Feng and A. V. Peterson, Statistics in Medicine 2001; 20: 3353–3373. Statistics in Medicine, 21(23):3745– 3747, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [627].

Evans:2002:AR

[959] Brent A. Evans, Ziding Feng, and Arthur V. Peterson. Authors’ reply. Statistics in Medicine, 21(23):3747–3748, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Metcalfe:2002:LEG

[960] Chris Metcalfe. Letter to the Editor: Goodness-of-fit statistics for agespecific reference intervals by P. Royston and E. M. Wright, Statistics in Medicine 2000; 19: 2943–2962. Statistics in Medicine, 21(23):3749– 3750, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [278].

Gomez:2002:BRB

[961] Guadalupe Gomez. Book review: 1. Bayes and Empirical Bayes Methods for Data Analysis. 2nd edn. Bradley P. Carlin and Thomas A. Louis. Chapman and Hall/CRC, Hertfordshire, U.K., 2000. No. of pages: xvii + 419. Price: 34.99. ISBN 1-58488-170-4. Statistics in Medicine, 21(23): 3751–3752, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McIntosh:2002:BRQ

[962] Martin W. McIntosh. Book review: Quantitative Methods for the Evaluation of Cancer Screening. Stephen Duffy, Jaques Esteve and Catherine Hill. Arnold, London, U.K., 1999. No. of pages: 320. Price: £40.00. ISBN 0-340-74125-2. Statistics in Medicine, 21(23):3752–3754, December 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmgren:2002:P

[963] Juni Palmgren. Preface. Statistics in Medicine, 21(24):3755, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ukoumunne:2002:CCI

[964] Obioha C. Ukoumunne. A comparison of confidence interval methods for the intraclass correlation coefficient in cluster randomized trials. Statistics in Medicine, 21(24):3757–3774, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1494].

Bogaerts:2002:MTE

[965] K. Bogaerts, R. Leroy, E. Lesaffre, and D. Declerck. Modelling tooth emergence data based on multivariate interval-censored data. Statistics in Medicine, 21(24):3775–3787, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuss:2002:GGF

[966] Oliver Kuss. Global goodness-of-fit tests in logistic regression with sparse data. Statistics in Medicine, 21(24):3789–3801, December 30, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ambler:2002:SPM

[967] Gareth Ambler, Anthony R. Brady, and Patrick Royston. Simplifying a prognostic model: a simulation study based on clinical data. Statistics in Medicine, 21(24):3803–3822, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hollis:2002:GSA

[968] Sally Hollis. A graphical sensitivity analysis for clinical trials with nonignorable missing binary outcome. Statistics in Medicine, 21(24):3823– 3834, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobbins:2002:CTC

[969] Timothy A. Dobbins and Judy M. Simpson. Comparison of tests for categorical data from a stratified cluster randomized trial. Statistics in Medicine, 21(24):3835–3846, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galimberti:2002:WKM

[970] Stefania Galimberti, Peter Sasieni, and Maria Grazia Valsecchi. A weighted Kaplan–Meier estimator for matched data with application to the comparison of chemotherapy and bone-marrow transplant in leukaemia. Statistics in Medicine, 21(24):3847–3864, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steen:2002:MPF

[971] Kristel Van Steen, Desmond Curran, Jocelyn Kramer, Geert Molenberghs, Ann Van Vreckem, Andrew Bottomley, and Richard Sylvester. Multicollinearity in prognostic factor analyses using the EORTC QLQ-C30: identification and impact on model selection. Statistics in Medicine, 21(24):3865–3884, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perez:2002:UMH

[972] Adriana Pérez, Rodolfo J. Dennis, Jacky F. A. Gil, Martín A. Rondón, and Adriana López. Use of the mean, hot deck and multiple imputation techniques to predict outcome in intensive care unit patients in Colombia. Statistics in Medicine, 21(24):3885–3896, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pauler:2002:PTP

[973] Donna K. Pauler and Dianne M. Finkelstein. Predicting time to prostate cancer recurrence based on joint models for non-linear longitudinal biomarkers and event time outcomes. Statistics in Medicine, 21(24): 3897–3911, December 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phillips:2003:IEG

[974] Alan Phillips and Vincent Haudiquet. ICH e9 guideline ‘statistical principles for clinical trials’: a case study. Statistics in Medicine, 22(1): 1–11, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2003:IEG

[975] David J. Brown. ICH E9 guideline ‘Statistical principles for clinical trials’: a case study Response to A. Phillips and V. Haudiquet. Statistics in Medicine, 22(1):13–17, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Law:2003:SET

[976] Matthew G. Law and Sean Emery. Selective exclusion of treatment arms in multi-arm randomized clinical trials. Statistics in Medicine, 22(1): 19–30, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McNally:2003:TIP

[977] Richard J. McNally, Hari Iyer, and Thomas Mathew. Tests for individual and population bioequivalence based on generalized p-values. Statistics in Medicine, 22(1):31–53, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2003:VBT

[978] Shein-Chung Chow, Jun Shao, and Hansheng Wang. In vitro bioequivalence testing. Statistics in Medicine, 22(1):55–68, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2003:IDP

[979] Anastasia Ivanova, Aliakbar Montazer-Haghighi, Sri Gopal Mohanty, and Stephen D. Durham. Improved up-and-down designs for phase I trials. Statistics in Medicine, 22(1):69–82, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kafadar:2003:ADC

[980] Karen Kafadar and Philip C. Prorok. Alternative definitions of comparable case groups and estimates of lead time and benefit time in randomized cancer screening trials. Statistics in Medicine, 22(1):83–111, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:2003:LDB

[981] Avis J. Thomas and Bradley P. Carlin. Late detection of breast and colorectal cancer in Minnesota counties: an application of spatial smoothing and clustering. Statistics in Medicine, 22(1):113–127, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2003:CCS

[982] Harland Austin and W. Dana Flanders. Case-control studies of genotypic relative risks using children of cases as controls. Statistics in Medicine, 22(1):129–145, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2003:TGM

[983] Y. Huang, P. J. G. Lisboa, and W. El-Deredy. Tumour grading from magnetic resonance spectroscopy: a comparison of feature extraction with variable selection. Statistics in Medicine, 22(1):147–164, January 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agostino:2003:NIT

[984] Ralph B. D’ Agostino, Sr. Non-inferiority trials: advances in concepts and methodology. Statistics in Medicine, 22(2):165–167, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2003:NIT

[985] Ralph B. D’Agostino, Sr., Joseph M. Massaro, and Lisa M. Sullivan. Non-inferiority trials: design concepts and issues — the encounters of academic consultants in statistics. Statistics in Medicine, 22(2):169–186, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laster:2003:NIT

[986] Larry L. Laster and Mary F. Johnson. Non-inferiority trials: the ‘at least as good as’ criterion. Statistics in Medicine, 22(2):187–200, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phillips:2003:NTN

[987] Kem F. Phillips. A new test of non-inferiority for anti-infective trials. Statistics in Medicine, 22(2):201–212, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hung:2003:SFI

[988] H. M. James Hung, Sue-Jane Wang, Yi Tsong, John Lawrence, and Robert T. O’Neil. Some fundamental issues with non-inferiority testing in active controlled trials. Statistics in Medicine, 22(2):213–225, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2003:TMN

[989] Sue-Jane Wang and H. M. James Hung. TACT method for non-inferiority testing in active controlled trials. Statistics in Medicine, 22(2):227–238, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rothmann:2003:DAN

[990] Mark Rothmann, Ning Li, Gang Chen, George Y. H. Chi, Robert Temple, and Hsiao-Hui Tsou. Design and analysis of non-inferiority mortality trials in oncology. Statistics in Medicine, 22(2):239–264, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sidik:2003:EUT

[991] Kurex Sidik. Exact unconditional tests for testing non-inferiority in matched-pairs design. Statistics in Medicine, 22(2):265–278, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Durkalski:2003:ACMa

[992] Valerie L. Durkalski, Yuko Y. Palesch, Stuart R. Lipsitz, and Philip F. Rust. Analysis of clustered matched-pair data for a non-inferiority study design. Statistics in Medicine, 22(2):279–290, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rashid:2003:RBT

[993] M. Mushfiqur Rashid. Rank-based tests for non-inferiority and equivalence hypotheses in multi-centre clinical trials using mixed models. Statistics in Medicine, 22(2):291–311, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OMalley:2003:AMM

[994] A. James O’Malley, Sharon lise T. Normand, and Richard E. Kuntz. Application of models for multivariate mixed outcomes to medical device trials: coronary artery stenting. Statistics in Medicine, 22(2):313–336, January 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2003:F

[995] R. B. D’Agostino, M. J. Campbell, and J. B. Greenhouse. Foreword. Statistics in Medicine, 22(3):337, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jeong:2003:PML

[996] Jong-Hyeon Jeong, Sin-Ho Jung, and Samuel Wieand. A parametric model for long-term follow-up data from phase III breast cancer clinical trials. Statistics in Medicine, 22(3):339–352, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willan:2003:INB

[997] Andrew R. Willan, Eric Bingshu Chen, Richard J. Cook, and D. Y. Lin. Incremental net benefit in randomized clinical trials with qualityadjusted survival. Statistics in Medicine, 22(3):353–362, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ambler:2003:NLM

[998] Gareth Ambler, Patrick Royston, and Jenny Head. Non-linear models for the relation between cardiovascular risk factors and intake of wine, beer and spirits. Statistics in Medicine, 22(3):363–383, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McKinnon:2003:VLD

[999] Elizabeth J. McKinnon, Ian R. James, Mina John, and Simon A. Mallal. Viral load detectability profiles for HIV infection. Statistics in Medicine, 22(3):385–396, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2003:MVJ

[1000] Xuming He, K. W. Ng, and Jian Shi. Marginal versus joint Box–Cox transformation with applications to percentile curve construction for IgG subclasses and blood pressures. Statistics in Medicine, 22(3):397–408, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gomez:2003:ILR

[1001] Guadalupe Gómez, Anna Espinal, and Stephen W. Lagakos. Inference for a linear regression model with an interval-censored covariate. Statistics in Medicine, 22(3):409–425, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wolfe:2003:TIO

[1002] Rory Wolfe, John B. Carlin, and George C. Patton. Transitions in an imperfectly observed binary variable: depressive symptomatology in adolescents. Statistics in Medicine, 22(3):427–440, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bureau:2003:ACT

[1003] Alexandre Bureau, Stephen Shiboski, and James P. Hughes. Applications of continuous time hidden Markov models to the study of misclassified disease outcomes. Statistics in Medicine, 22(3):441–462, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hudgens:2003:ECP

[1004] Michael G. Hudgens. Estimating cumulative probabilities from incomplete longitudinal binary responses with application to HIV vaccine trials. Statistics in Medicine, 22(3):463–479, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Satagopan:2003:SPG

[1005] Jaya M. Satagopan and Katherine S. Panageas. A statistical perspective on gene expression data analysis. Statistics in Medicine, 22(3):481–499, February 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiessens:2003:CGS

[1006] Bart Spiessens, Emmanuel Lesaffre, and Geert Verbeke. A comparison of group sequential methods for binary longitudinal data. Statistics in Medicine, 22(4):501–515, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:2003:MDT

[1007] Robert H. Lyles and Andrew S. Allen. Missing data in the 2 × 2 table: patterns and likelihood-based analysis for cross-sectional studies with supplemental sampling. Statistics in Medicine, 22(4):517–534, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stylianou:2003:EPT

[1008] Mario Stylianou, Michael Proschan, and Nancy Flournoy. Estimating the probability of toxicity at the target dose following an up-and-down design. Statistics in Medicine, 22(4):535–543, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demissie:2003:BDM

[1009] Serkalem Demissie, Michael P. LaValley, Nicholas J. Horton, Robert J. Glynn, and L. Adrienne Cupples. Bias due to missing exposure data using complete-case analysis in the proportional hazards regression model. Statistics in Medicine, 22(4):545–557, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mazumdar:2003:MCP

[1010] Madhu Mazumdar, Alex Smith, and Jennifer Bacik. Methods for categorizing a prognostic variable in a multivariable setting. Statistics in Medicine, 22(4):559–571, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2003:CHH

[1011] Peter B. Gilbert, Ian W. McKeague, Geoffrey Eisen, Christopher Mullins, Aissatou Guéye-Ndiaye, Souleymane Mboup, and Phyllis J. Kanki. Comparison of HIV-1 and HIV-2 infectivity from a prospective cohort study in Senegal. Statistics in Medicine, 22(4):573–593, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kowalski:2003:GCA

[1012] J. Kowalski, X. M. Tu, G. Jia, M. Perlis, E. Frank, P. Crits-Christoph, and D. J. Kupfer. Generalized covariance-adjusted canonical correlation analysis with application to psychiatry. Statistics in Medicine, 22(4): 595–610, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reiczigel:2003:CIB

[1013] Jenö Reiczigel. Confidence intervals for the binomial parameter: some new considerations. Statistics in Medicine, 22(4):611–621, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holmes:2003:GMR

[1014] C. C. Holmes and N. A. Heard. Generalized monotonic regression using random change points. Statistics in Medicine, 22(4):623–638, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2003:SMF

[1015] P. Royston and W. Sauerbrei. Stability of multivariable fractional polynomial models with selection of variables and transformations: a bootstrap investigation. Statistics in Medicine, 22(4):639–659, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2003:LEU

[1016] R. J. Marshall. Letter to the Editor: Understanding neural networks using regression trees: an application to multiple myeloma survival data by D. Faraggi, M. LeBlanc and J. Crowley, Statistics in Medicine 2001; 20:2965–2976. Statistics in Medicine, 22(4):661, February 28, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [598].

Faraggi:2003:AR

[1017] David Faraggi, Michael Leblanc, and John Crowley. Authors’ reply. Statistics in Medicine, 22(4):661–662, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:2003:BRM

[1018] Jørgen Hilden. Book review: Modeling In Medical Decision Making. A Bayesian Approach.Giovanni Parmigiani, Wiley, Chichester, 2002. No. of pages: xi + 267. Price: £45.00. ISBN 0-471-98608-9. Statistics in Medicine, 22(4):663–664, February 28, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2003:CTE

[1019] Weichung J. Shih, Peter Ouyang, Hui Quan, Yong Lin, Bart Michiels, and Luc Bijnens. Controlling type I error rate for fast track drug development programmes. Statistics in Medicine, 22(5):665–675, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2003:SCT

[1020] John Whitehead and Tatsuru Matsushita. Stopping clinical trials because of treatment ineffectiveness: a comparison of a futility design with a method of stochastic curtailment. Statistics in Medicine, 22(5):677–687, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2003:SDP

[1021] Nigel Stallard and Susan Todd. Sequential designs for phase III clinical trials incorporating treatment selection. Statistics in Medicine, 22(5): 689–703, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weir:2003:CSA

[1022] Christopher J. Weir and Kennedy R. Lees. Comparison of stratification and adaptive methods for treatment allocation in an acute stroke clinical trial. Statistics in Medicine, 22(5):705–726, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1563].

Mazumdar:2003:GSD

[1023] Madhu Mazumdar and Aiyi Liu. Group sequential design for comparative diagnostic accuracy studies. Statistics in Medicine, 22(5):727–739, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Garrett:2003:TEF

[1024] Andrew D. Garrett. Therapeutic equivalence: fallacies and falsification. Statistics in Medicine, 22(5):741–762, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:2003:HBA

[1025] Peter F. Thall, J. Kyle Wathen, B. Nebiyou Bekele, Richard E. Champlin, Laurence H. Baker, and Robert S. Benjamin. Hierarchical Bayesian approaches to phase II trials in diseases with multiple subtypes. Statistics in Medicine, 22(5):763–780, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2003:ATC

[1026] Ian R. White, James Carpenter, Stuart J. Pocock, and Robert A. Henderson. Adjusting treatment comparisons to account for non-randomized interventions: an example from an angina trial. Statistics in Medicine, 22(5):781–793, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pauler:2003:PMM

[1027] Donna K. Pauler, Sheryl McCoy, and Carol Moinpour. Pattern mixture models for longitudinal quality of life studies in advanced stage disease. Statistics in Medicine, 22(5):795–809, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsuyama:2003:SAE

[1028] Yutaka Matsuyama. Sensitivity analysis for the estimation of rates of change with non-ignorable drop-out: an application to a randomized clinical trial of the vitamin D_{3}. Statistics in Medicine, 22(5):811–827, March 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Auclair:2003:P

[1029] Jean Auclair and Peter Bauer. Preface. Statistics in Medicine, 22(6):829, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ane:2003:PPP

[1030] Cécile Ané and Didier Concordet. Population pharmacokinetics/ pharmacodynamics relationships of an anticancer drug. Statistics in Medicine, 22(6):833–846, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2003:IES

[1031] Frank Bretz, Ludwig A. Hothorn, and Jason C. Hsu. Identifying effective and/or safe doses by stepwise confidence intervals for ratios. Statistics in Medicine, 22(6):847–858, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Letierce:2003:TTC

[1032] Alexia Letierce, Pascale Tubert-Bitter, Andrew Kramar, and Jean Maccario. Two-treatment comparison based on joint toxicity and efficacy ordered alternatives in cancer trials. Statistics in Medicine, 22(6):859–868, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dilleen:2003:NPE

[1033] Maria Dilleen, Günter Heimann, and Ian Hirsch. Non-parametric estimators of a monotonic dose-response curve and bootstrap confidence intervals. Statistics in Medicine, 22(6):869–882, March 30, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pigeot:2003:ANI

[1034] Iris Pigeot, Juliane Schäfer, Joachim Röhmel, and Dieter Hauschke. Assessing non-inferiority of a new treatment in a three-arm clinical trial including a placebo. Statistics in Medicine, 22(6):883–899, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carrasco:2003:AIB

[1035] Josep-Lluís Carrasco and Lluís Jover. Assessing individual bioequivalence using the structural equation model. Statistics in Medicine, 22(6):901– 912, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dragalin:2003:KLD

[1036] Vladimir Dragalin, Valerii Fedorov, Scott Patterson, and Byron Jones. Kullback–Leibler divergence for evaluating bioequivalence. Statistics in Medicine, 22(6):913–930, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Broet:2003:STE

[1037] P. Broët, P. Tubert-Bitter, Y. De Rycke, and T. Moreau. A score test for establishing non-inferiority with respect to short-term survival in two-sample comparisons with identical proportions of long-term survivors. Statistics in Medicine, 22(6):931–940, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Durrleman:2003:UPP

[1038] Sylvain Durrleman and Philip Chaikin. The use of putative placebo in active control trials: two applications in a regulatory setting. Statistics in Medicine, 22(6):941–952, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Posch:2003:IDF

[1039] Martin Posch, Peter Bauer, and Werner Brannath. Issues in designing flexible trials. Statistics in Medicine, 22(6):953–969, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jennison:2003:MCS

[1040] Christopher Jennison and Bruce W. Turnbull. Mid-course sample size modification in clinical trials based on the observed treatment effect. Statistics in Medicine, 22(6):971–993, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2003:BSS

[1041] Tim Friede and Meinhard Kieser. Blinded sample size reassessment in non-inferiority and equivalence trials. Statistics in Medicine, 22(6):995– 1007, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hahn:2003:IMC

[1042] Seokyung Hahn and Anne Whitehead. An illustration of the modelling of cost and efficacy data from a clinical trial. Statistics in Medicine, 22 (6):1009–1024, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petit:2003:BAP

[1043] Claude Petit and Jean Maccario. A Bayesian analysis of pharmacoeconomic data from a clinical trial on schizophrenia. Statistics in Medicine, 22(6):1025–1039, March 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chambless:2003:AAC

[1044] Lloyd E. Chambless and Vicki Davis. Analysis of associations with change in a multivariate outcome variable when baseline is subject to measurement error. Statistics in Medicine, 22(7):1041–1067, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tosteson:2003:PSS

[1045] Tor D. Tosteson, Jeffrey S. Buzas, Eugene Demidenko, and Margaret Karagas. Power and sample size calculations for generalized regression models with covariate measurement error. Statistics in Medicine, 22(7): 1069–1082, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2003:URM

[1046] Ian R. White, Ilona Koupilova, and James Carpenter. The use of regression models for medians when observed outcomes may be modified by interventions. Statistics in Medicine, 22(7):1083–1096, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ng:2003:EBS

[1047] S. K. Ng and G. J. McLachlan. An EM-based semi-parametric mixture model approach to the regression analysis of competing-risks data. Statistics in Medicine, 22(7):1097–1111, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zahl:2003:RAM

[1048] Per-Henrik Zahl. Regression analysis with multiplicative and timevarying additive regression coefficients with examples from breast and colon cancer. Statistics in Medicine, 22(7):1113–1127, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2003:TNB

[1049] Andy H. Lee, K. Wang, Kelvin K. W. Yau, and Peter J. Somerford. Truncated negative binomial mixed regression modelling of ischaemic stroke hospitalizations. Statistics in Medicine, 22(7):1129–1139, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pastor-Barriuso:2003:TMC

[1050] Roberto Pastor-Barriuso, Eliseo Guallar, and Josef Coresh. Transition models for change-point estimation in logistic regression. Statistics in Medicine, 22(7):1141–1162, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2003:DCM

[1051] Ursula Berger, Juliane Schäfer, and Kurt Ulm. Dynamic Cox modelling based on fractional polynomials: time-variations in gastric cancer prognosis. Statistics in Medicine, 22(7):1163–1180, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2003:RAI

[1052] D. Y. Lin. Regression analysis of incomplete medical cost data. Statistics in Medicine, 22(7):1181–1200, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taube:2003:LEP

[1053] A. Taube. Letter to the Editor: The predictive value of microbiologic diagnostic tests if asymptomatic carriers are present. Ronny K. Gunnarsson and Jan Lanke, Statistics in Medicine 2002; 21:1773–1785. Statistics in Medicine, 22(7):1201–1202, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [801].

Law:2003:LES

[1054] Dionne C. G. Law. Letter to the Editor: Spatial mixture relative risk models applied to disease mapping. A. B. Lawson and A. Clark, Statistics in Medicine 2002; 21:359–370. Statistics in Medicine, 22(7):1202–1203, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [696].

Lawson:2003:AR

[1055] Andrew Lawson and Allan Clark. Authors’ reply. Statistics in Medicine, 22(7):1203, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2003:LES

[1056] Wen-Chung Lee. Letter to the Editor: Standardization using the harmonically weighted ratios: internal and external comparisons. Wen-Chung Lee, Statistics in Medicine 2002; 21: 247–261. Statistics in Medicine, 22(7):1204, April 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [688].

White:2003:CTC

[1057] Ian R. White and Simon G. Thompson. Choice of test for comparing two groups, with particular application to skewed outcomes. Statistics in Medicine, 22(8):1205–1215, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2003:TEN

[1058] Nian-Sheng Tang, Man-Lai Tang, and Ivan Siu Fung Chan. On tests of equivalence via non-unity relative risk for matched-pair design. Statistics in Medicine, 22(8):1217–1233, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Preisser:2003:IPA

[1059] John S. Preisser, Mary L. Young, Daniel J. Zaccaro, and Mark Wolfson. An integrated population-averaged approach to the design, analysis and sample size determination of cluster-unit trials. Statistics in Medicine, 22(8):1235–1254, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Have:2003:CLM

[1060] Thomas R. Ten Have, Marshall Joffe, and Mark Cary. Causal logistic models for non-compliance under randomized treatment with univariate binary response. Statistics in Medicine, 22(8):1255–1283, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joffe:2003:WIV

[1061] Marshall M. Joffe and Colleen Brensinger. Weighting in instrumental variables and G-estimation. Statistics in Medicine, 22(8):1285–1303, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2003:SSE

[1062] Sin-Ho Jung and Chul Ahn. Sample size estimation for GEE method for comparing slopes in repeated measurements data. Statistics in Medicine, 22(8):1305–1315, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sullivan:2003:RPA

[1063] L. M. Sullivan and R. B. D’Agostino, Sr. Robustness and power of analysis of covariance applied to ordinal scaled data as arising in randomized controlled trials. Statistics in Medicine, 22(8):1317–1334, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arends:2003:CMO

[1064] Lidia R. Arends, Zoltán Vokó, and Theo Stijnen. Combining multiple outcome measures in a meta-analysis: an application. Statistics in Medicine, 22(8):1335–1353, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2003:BRP

[1065] Nigel Stallard. Book review: 1.Principles and Practice of Clinical Research. John L. Gallin (ed.), Academic Press, San Diego, 2001. No. of pages: xv + 490. Price: £93.95. ISBN 0-12-274065-3. Statistics in Medicine, 22(8):1355–1356, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uesaka:2003:BRR

[1066] Hiroyuki Uesaka. Book review: 2. Randomised Controlled Clinical Trials. 2nd edn. Christopher J. Bulpitt, Kluwer Academic Publishers, Boston, 1996. No. of pages: ix + 437. Price: £120.75. ISBN 0-79234-257-7. Statistics in Medicine, 22(8):1356–1357, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guilbaud:2003:BRS

[1067] Olivier Guilbaud. Book review: 3. Statistical Aspects of the Design and Analysis of Clinical Trials. Brian S. Everitt and Andrew Pickles, Imperial College Press, London, 1999. No. of pages: xiv + 336. Price: £53.00. ISBN 1-86094-153-2. Statistics in Medicine, 22(8):1357–1358, April 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2003:P

[1068] Lillian S. Lin. Preface. Statistics in Medicine, 22(9):1359–1360, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Livengood:2003:OR

[1069] John R. Livengood. Opening remarks. Statistics in Medicine, 22(9): 1363–1364, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friedman:2003:MAR

[1070] Jerome H. Friedman and Jacqueline J. Meulman. Multiple additive regression trees with application in epidemiology. Statistics in Medicine, 22(9):1365–1381, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wegman:2003:VDM

[1071] Edward J. Wegman. Visual data mining. Statistics in Medicine, 22(9): 1383–1397, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2003:DRG

[1072] R. Dennis Cook. Dimension reduction and graphical exploration in regression including survival analysis. Statistics in Medicine, 22(9):1399– 1413, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gotway:2003:SPC

[1073] Carol A. Gotway and Russell D. Wolfinger. Spatial prediction of counts and rates. Statistics in Medicine, 22(9):1415–1432, May 15, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hedeker:2003:MEM

[1074] Donald Hedeker. A mixed-effects multinomial logistic regression model. Statistics in Medicine, 22(9):1433–1446, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laird:2003:RMA

[1075] Nan M. Laird and Karen T. Cuenco. Regression methods for assessing familial aggregation of disease. Statistics in Medicine, 22(9):1447–1455, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chakraborty:2003:ECU

[1076] Hrishikesh Chakraborty, Ronald W. Helms, Pranab K. Sen, and Myron S. Cohen. Estimating correlation by using a general linear mixed model: evaluation of the relationship between the concentration of HIV-1 RNA in blood and semen. Statistics in Medicine, 22(9):1457–1464, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harezlak:2003:IDS

[1077] Jaroslaw Harezlak, Sujuan Gao, and Siu L. Hui. An illness-death stochastic model in the analysis of longitudinal dementia data. Statistics in Medicine, 22(9):1465–1475, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borkowf:2003:ULR

[1078] Craig B. Borkowf, Paul S. Albert, and Christian C. Abnet. Using lowess to remove systematic trends over time in predictor variables prior to logistic regression with quantile categories. Statistics in Medicine, 22(9): 1477–1493, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burr:2003:IES

[1079] Tom L. Burr, James R. Gattiker, and Philip J. Gerrish. An investigation of error sources and their impact in estimating the time to the most recent ancestor of spatially and temporally distributed HIV sequences. Statistics in Medicine, 22(9):1495–1516, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Varbanov:2003:BAQ

[1080] Alex R. Varbanov and Charles H. Taylor. Bayesian analysis of quantitative antimicrobial assays. Statistics in Medicine, 22(9):1517–1526, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kalsbeek:2003:SMG

[1081] William D. Kalsbeek. Sampling minority groups in health surveys. Statistics in Medicine, 22(9):1527–1549, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anderson:2003:DCM

[1082] Robert N. Anderson and Harry M. Rosenberg. Disease classification: measuring the effect of the Tenth Revision of the International Classification of Diseases on cause-of-death data in the United States. Statistics in Medicine, 22(9):1551–1570, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schenker:2003:SRR

[1083] Nathaniel Schenker and Jennifer D. Parker. From single-race reporting to multiple-race reporting: using imputation methods to bridge the transition. Statistics in Medicine, 22(9):1571–1587, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:2003:WRN

[1084] Roderick J. Little and Sonya Vartivarian. On weighting the rates in nonresponse weights. Statistics in Medicine, 22(9):1589–1599, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Looney:2003:MCT

[1085] Stephen W. Looney and Peter W. Jones. A method for comparing two normal means using combined samples of correlated and uncorrelated data. Statistics in Medicine, 22(9):1601–1610, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1362].

Frankel:2003:ANT

[1086] Martin R. Frankel, K. P. Srinath, David C. Hoaglin, Michael P. Battaglia, Philip J. Smith, Robert A. Wright, and Meena Khare. Adjustments for non-telephone bias in random-digit-dialling surveys. Statistics in Medicine, 22(9):1611–1626, May 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daniels:2003:MRE

[1087] Michael J. Daniels and Yan D. Zhao. Modelling the random effects covariance matrix in longitudinal data. Statistics in Medicine, 22(10): 1631–1647, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2003:ACV

[1088] E. C. Marshall and D. J. Spiegelhalter. Approximate cross-validatory predictive checks in disease mapping models. Statistics in Medicine, 22 (10):1649–1660, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Myles:2003:BEB

[1089] Jonathan P. Myles, Richard M. Nixon, Stephen W. Duffy, Laszlo Tabar, Caroline Boggis, Gareth Evans, Andrew Shenton, and Anthony Howell.

Bayesian evaluation of breast cancer screening using data from two studies. Statistics in Medicine, 22(10):1661–1674, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kousignian:2003:MMC

[1090] I. Kousignian, B. Autran, C. Chouquet, V. Calvez, E. Gomard, C. Katlama, Y. Rivière, and D. Costagliola. Markov modelling of changes in HIV-specific cytotoxic T-lymphocyte responses with time in untreated HIV-1 infected patients. Statistics in Medicine, 22(10):1675–1690, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deGunst:2003:EHT

[1091] Mathisca C. M. de Gunst, Anup Dewanji, and E. Georg Luebeck. Exploring heterogeneity in tumour data using Markov chain Monte Carlo. Statistics in Medicine, 22(10):1691–1707, May 30, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2003:EIR

[1092] Tom Smith and Penelope Vounatsou. Estimation of infection and recovery rates for highly polymorphic parasites when detectability is imperfect, using hidden Markov models. Statistics in Medicine, 22(10): 1709–1724, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanson:2003:MMB

[1093] Timothy Hanson, Edward J. Bedrick, Wesley O. Johnson, and Mark C. Thurmond. A mixture model for bovine abortion and foetal survival. Statistics in Medicine, 22(10):1725–1739, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burr:2003:MAS

[1094] Deborah Burr, Hani Doss, Glen E. Cooke, and Pascal J. Goldschmidt-Clermont. A meta-analysis of studies on the association of the platelet PlA polymorphism of glycoprotein IIIa and risk of coronary heart disease. Statistics in Medicine, 22(10):1741–1760, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2003:HBS

[1095] Ying C. MacNab. Hierarchical Bayesian spatial modelling of small-area rates of non-rare disease. Statistics in Medicine, 22(10):1761–1773, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pekoz:2003:CID

[1096] Erol A. Peköz, Michael Shwartz, Lisa I. Iezzoni, Arlene S. Ash, Michael A. Posner, and Joseph D. Restuccia. Comparing the importance of disease rate versus practice style variations in explaining differences in small area hospitalization rates for two respiratory conditions. Statistics in Medicine, 22(10):1775–1786, May 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galbraith:2003:IAC

[1097] Sally Galbraith and Ian C. Marschner. Interim analysis of continuous long-term endpoints in clinical trials with longitudinal outcomes. Statistics in Medicine, 22(11):1787–1805, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strickland:2003:EPS

[1098] Pamela A. Ohman Strickland and Shou-En Lu. Estimates, power and sample size calculations for two-sample ordinal outcomes under beforeafter study designs. Statistics in Medicine, 22(11):1807–1818, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Govindarajulu:2003:RSS

[1099] Z. Govindarajulu. Robustness of sample size re-estimation procedure in clinical trials (arbitrary populations). Statistics in Medicine, 22(11): 1819–1828, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Natarajan:2003:IEP

[1100] Loki Natarajan and John O’Quigley. Interval estimates of the probability of toxicity at the maximum tolerated dose for small samples. Statistics in Medicine, 22(11):1829–1836, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:2003:ACP

[1101] Michael P. Fay, Ruth Pfeiffer, Kathleen A. Cronin, Chenxiong Le, and Eric J. Feuer. Age-conditional probabilities of developing cancer. Statistics in Medicine, 22(11):1837–1848, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2003:LBC

[1102] Jianrong Wu, A. C. M. Wong, and Guoyong Jiang. Likelihood-based confidence intervals for a log-normal mean. Statistics in Medicine, 22 (11):1849–1860, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cappelleri:2003:MLS

[1103] Joseph C. Cappelleri and Naitee Ting. A modified large-sample approach to approximate interval estimation for a particular intraclass correlation coefficient. Statistics in Medicine, 22(11):1861–1877, June 15, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mertens:2003:MPR

[1104] Bart J. A. Mertens. Microarrays, pattern recognition and exploratory data analysis. Statistics in Medicine, 22(11):1879–1899, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shah:2003:MLA

[1105] Amrik Shah and Gerald Hajian. A maximum likelihood approach for estimating the QT correction factor using mixed effects model. Statistics in Medicine, 22(11):1901–1909, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heo:2003:HLM

[1106] Moonseong Heo, Myles S. Faith, John W. Mott, Bernard S. Gorman, David T. Redden, and David B. Allison. Hierarchical linear models for the development of growth curves: an example with body mass index in overweight/obese adults. Statistics in Medicine, 22(11):1911–1942, June 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hasselblad:2003:PCL

[1107] Vic Hasselblad and Andrew S. Allen. Power calculations for large multiarm placebo-controlled studies of dichotomous outcomes. Statistics in Medicine, 22(12):1943–1954, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2003:AAC

[1108] John R. Cook, Michael Drummond, Henry Glick, and Joseph F. Heyse. Assessing the appropriateness of combining economic data from multinational clinical trials. Statistics in Medicine, 22(12):1955–1976, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraggi:2003:CIP

[1109] D. Faraggi, P. Izikson, and B. Reiser. Confidence intervals for the 50 per cent response dose. Statistics in Medicine, 22(12):1977–1988, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murthy:2003:IGT

[1110] B. N. Murthy. Implications of genetic traits on vaccine efficacy. Statistics in Medicine, 22(12):1989–1998, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:2003:CHA

[1111] Dean A. Follmann. Comparing HLA antigen frequencies between two groups of patients. Statistics in Medicine, 22(12):1999–2013, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Franco:2003:PMM

[1112] Glaura C. Franco, Paulo S. Lucio, Flávia C. Parra, and Sérgio D. J. Pena. A probability model for the meiosis I non-disjunction fraction in numerical chromosomal anomalies. Statistics in Medicine, 22(12):2015– 2024, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lam:2003:RME

[1113] K. F. Lam and David Ip. REML and ML estimation for clustered grouped survival data. Statistics in Medicine, 22(12):2025–2034, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosychuk:2003:BCT

[1114] Rhonda J. Rosychuk and Mary E. Thompson. Bias correction of twostate latent Markov process parameter estimates under misclassification. Statistics in Medicine, 22(12):2035–2055, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shelton:2003:CLB

[1115] Brent J. Shelton, Gregg H. Gilbert, Zhenmei Lu, Patrick Bradshaw, L. Scott Chavers, and George Howard. Comparing longitudinal binary outcomes in an observational oral health study. Statistics in Medicine, 22(12):2057–2070, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graham:2003:CLR

[1116] P. L. Graham, K. Mengersen, and A. P. Morton. Confidence limits for the ratio of two rates based on likelihood scores: non-iterative method. Statistics in Medicine, 22(12):2071–2083, June 30, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [1117, 1545, 1669].

Anonymous:2003:LEC

[1117] Anonymous. Letter to the Editor: Confidence limits for the ratio of two rates based on likelihood scores: non-iterative method. P. L. Graham, K. Mengersen and A. P. Morton, Statistics in Medicine, 2003; 22:2071– 2083. Statistics in Medicine, 22(12):2085–2086, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1116, 1545, 1669].

Louis:2003:BRS

[1118] Tthomas A. Louis. Book review: 1. Statistical Methods in Medical Research. 4th edition. P. Armitage, G. Berry, and J. N. S. Matthews, Blackwell Science, Oxford, 2001. No. of pages: 817. Price: £55.00. ISBN 0-632-052570. Statistics in Medicine, 22(12):2087–2088, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2003:BRP

[1119] Vern Farewell. Book review: 2. Principles of Medical Statistics. Alvan R. Feinstein, Chapman Hall/CRC, London, 2001. No. of pages: 679. Price: £53.99. ISBN 1-5848-8216-6. Statistics in Medicine, 22(12):2088–2090, June 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2003:EDS

[1120] Ming Tan, Hong-Bin Fang, Guo-Liang Tian, and Peter J. Houghton. Experimental design and sample size determination for testing synergism in drug combination studies based on uniform measures. Statistics in Medicine, 22(13):2091–2100, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andre:2003:OVS

[1121] Carmen D. S. André, Subhash C. Narula, Silvia N. Elian, and Rodrigo A. Tavares. An overview of the variables selection methods for the minimum sum of absolute errors regression. Statistics in Medicine, 22(13):2101– 2111, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Terrin:2003:APB

[1122] Norma Terrin, Christopher H. Schmid, Joseph Lau, and Ingram Olkin. Adjusting for publication bias in the presence of heterogeneity. Statistics in Medicine, 22(13):2113–2126, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1649].

Le:2003:SSM

[1123] Chap T. Le, Ping Liu, Bruce R. Lindgren, Kathleen A. Daly, and G. Scott Giebink. Some statistical methods for investigating the date of birth as a disease indicator. Statistics in Medicine, 22(13):2127–2135, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Slasor:2003:JME

[1124] Peter Slasor and Nan Laird. Joint models for efficient estimation in proportional hazards regression models. Statistics in Medicine, 22(13): 2137–2148, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nakashima:2003:ESP

[1125] Eiji Nakashima, Shuichi Tsuji, Haruhito Fukuoka, Megu Ohtaki, and Katsuhide Ito. Estimating the stenosis probabilities in arteriosclerosis obliterans using generalized estimating equations. Statistics in Medicine, 22(13):2149–2160, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnhart:2003:EMD

[1126] Huiman X. Barnhart and Andrzej S. Kosinski. Evaluating medical diagnostic tests at the subunit level in the presence of verification bias. Statistics in Medicine, 22(13):2161–2176, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:2003:ASB

[1127] Niels G. Becker, James J. C. Lewis, Zhengfeng Li, and Ann McDonald. Age-specific back-projection of HIV diagnosis data. Statistics in Medicine, 22(13):2177–2190, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:2003:EFA

[1128] Douglas J. Taylor, Mark A. Weaver, and Ronald E. Roddy. Evaluating factors associated with STD infection in a study with interval-censored event times and an unknown proportion of participants not at risk for disease. Statistics in Medicine, 22(13):2191–2204, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2003:EMS

[1129] James X. Song and Chul W. Ahn. An evaluation of methods for the stratified analysis of clustered binary data in community intervention trials. Statistics in Medicine, 22(13):2205–2216, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ziegler:2003:APO

[1130] Andreas Ziegler, Christian Kastner, and Jenny Chang-Claude. Analysis of pregnancy and other factors on detection of human papilloma virus (HPV) infection using weighted estimating equations for follow-up data. Statistics in Medicine, 22(13):2217–2233, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:2003:C

[1131] Jun-Mo Nam. Corrections. Statistics in Medicine, 22(13):2235, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoover:2003:C

[1132] Donald R. Hoover. Corrections. Statistics in Medicine, 22(13):2235, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koster:2003:BRC

[1133] Jan Koster. Book review: 1. Causation, Prediction, and Search. 2nd edn. Peter Spirtes, Clark Glymour and Richard Scheines, MIT Press, Cambridge, MA, 2000. No. of pages: 543. ISBN 0-262-19440-6. Statistics in Medicine, 22(13):2236–2237, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:2003:BRS

[1134] A. W. F. Edwards. Book review: 2. The Subjectivity of Scientists and the Bayesian Approach. S. J. Press and J. M. Tanur, Wiley, New York, 2001. No. of pages: x + 274. ISBN 0-471-39685-0. Statistics in Medicine, 22(13):2237–2238, July 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2003:NDM

[1135] Patrick Royston, Mahesh K. B. Parmar, and Wendi Qian. Novel designs for multi-arm clinical trials with survival outcomes with an application in ovarian cancer. Statistics in Medicine, 22(14):2239–2256, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2003:TEM

[1136] Margaret Shih, Chris Gennings, Vernon M. Chinchilli, and Walter H. Carter, Jr. Titrating and evaluating multi-drug regimens within subjects. Statistics in Medicine, 22(14):2257–2279, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hudgens:2003:AVL

[1137] Michael G. Hudgens, Antje Hoering, and Steven G. Self. On the analysis of viral load endpoints in HIV vaccine trials. Statistics in Medicine, 22 (14):2281–2298, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:2003:PAE

[1138] Michael Schemper. Predictive accuracy and explained variation. Statistics in Medicine, 22(14):2299–2308, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:2003:MMA

[1139] In-Sun Nam, Kerrie Mengersen, and Paul Garthwaite. Multivariate meta-analysis. Statistics in Medicine, 22(14):2309–2333, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rachet:2003:EDL

[1140] B. Rachet, M. Abrahamowicz, A. J. Sasco, and J. Siemiatycki. Estimating the distribution of lag in the effect of short-term exposures and interventions: adaptation of a non-parametric regression spline model. Statistics in Medicine, 22(14):2335–2363, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jordan:2003:DTD

[1141] K. Jordan, P. W. Jones, and K. Dziedzic. Describing three-dimensional cervical spine movement in a diseased and a non-diseased group using multilevel modelling. Statistics in Medicine, 22(14):2365–2380, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulldorff:2003:BRS

[1142] Martin Kulldorff. Book review: 1. Spatial Epidemiology: Methods and Applications. P. Elliott, J. C. Wakefield, N. G. Best and D. J. Briggs (eds), Oxford University Press, Oxford, 2000. No. of pages: xvii + 475. ISBN 0-19-262941-7. Statistics in Medicine, 22(14):2381–2382, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osler:2003:BRE

[1143] Merete Osler. Book review: 2. Epidemiologic Analysis. A Case-Oriented Approach. Steve Selvin, Oxford University Press, 2001. No. of pages: xiv + 323, ISBN 0-19-514489-9. Statistics in Medicine, 22(14):2383, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ickstadt:2003:BRS

[1144] Katja Ickstadt. Book review: Statistical Methods in Spatial Epidemiology. Andrew B. Lawson, Wiley, Chichester, 2001. No. of pages: x + 277. ISBN 0-471-97572-9. Statistics in Medicine, 22(14):2383–2384, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peddada:2003:BRB

[1145] Shyamal D. Peddada. Book review: A Biologist’s Guide to Analysis of DNA Micro-Array Data. Steen Knudsen, Wiley, New York, 2002. No. of pages: xiii + 125. ISBN 0-471-22490-1. Statistics in Medicine, 22 (14):2385, July 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2003:GSC

[1146] Alexei Dmitrienko, Walter W. Offen, and Peter H. Westfall. Gatekeeping strategies for clinical trials that do not require all primary effects to be significant. Statistics in Medicine, 22(15):2387–2400, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2003:MAE

[1147] Jian-Lun Xu and Philip C. Prorok. Modelling and analysing exchangeable binary data with random cluster sizes. Statistics in Medicine, 22 (15):2401–2416, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Durkalski:2003:ACMb

[1148] Valerie L. Durkalski, Yuko Y. Palesch, Stuart R. Lipsitz, and Philip F. Rust. Analysis of clustered matched-pair data. Statistics in Medicine, 22 (15):2417–2428, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2003:LOC

[1149] Jun Shao and Bob Zhong. Last observation carry-forward and last observation analysis. Statistics in Medicine, 22(15):2429–2441, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2003:IEA

[1150] Kung-Jong Lui. Interval estimation of the attributable risk for multiple exposure levels in case-control studies with confounders. Statistics in Medicine, 22(15):2443–2457, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duffy:2003:BRM

[1151] Stephen W. Duffy, Thomas E. Rohan, Rita Kandel, Teresa C. Prevost, Kenneth Rice, and Jonathan P. Myles. Book review: Misclassification in a matched case-control study with variable matching ratio: application to a study of c-erbB-2 overexpression and breast cancer. Statistics in Medicine, 22(15):2459–2468, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coffey:2003:PIP

[1152] Christopher S. Coffey and Keith E. Muller. Properties of internal pilots with the univariate approach to repeated measures. Statistics in Medicine, 22(15):2469–2485, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2003:IAB

[1153] Philip J. Smith, David C. Hoaglin, Michael P. Battaglia, and Lawrence Barker. Implementation and applications of bootstrap methods for the national immunization survey. Statistics in Medicine, 22(15):2487–2502, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nakas:2003:ADM

[1154] Christos Nakas, Constantin T. Yiannoutsos, Ronald J. Bosch, and Chronis Moyssiadis. Assessment of diagnostic markers by goodness-of-fit tests. Statistics in Medicine, 22(15):2503–2513, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraggi:2003:RCA

[1155] David Faraggi, Benjamin Reiser, and Enrique F. Schisterman. ROC curve analysis for biomarkers based on pooled assessments. Statistics in Medicine, 22(15):2515–2527, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lynch:2003:BRU

[1156] Andrew G. Lynch. Book reviews: Using SPSS for Windows. Kristin E. Voelkl and Susan B. Gerber, Springer, New York, 1999. No. of pages: xvi + 228. ISBN 0-387-98563-8. Doing Statistics with SPSS. Alistair W. Kerr, Howard K. Hall and Stephen A. Kozub, Sage Publications, London, 2002. No. of pages: vii + 238. ISBN 0-7169-7385-0. Data Analysis using SPSS for Windows Versions 8-10: a Beginner’s Guide New Edition. Jeremy J. Foster, Sage Publications, London, 2002. No. of pages: xvii + 252. ISBN 0-7169-6927-0. Statistics in Medicine, 22(15):2529–2531, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2003:BRR

[1157] Per Kragh Andersen. Book review: Regression Modeling Strategies with Applications to Linear Models, Logistic Regression and Survival Analysis. Frank E. Harrell, Jun, Springer-Verlag, New York, 2001. No. of pages:

568. ISBN 0-387-95232-2. Statistics in Medicine, 22(15):2531–2532, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burger:2003:BRR

[1158] Hans Ulrich Burger. Book review: Risk Stratification, A Practical Guide for Clinicians. Charles C. Miller III, Michael J. Reardon and Hazim J. Safi, Cambridge University Press, Cambridge, 2001. No. of pages: xiv + 174. ISBN 0-521-66945-6. Statistics in Medicine, 22(15):2532–2533, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2003:BRC

[1159] Kyungmann Kim. Book review: Clinical Trials. Lelia Duley and Barbara Farrell (eds), BMJ Books, London, 2002. No. of pages: x + 133. Price: ISBN 0-7279-1599-1. Statistics in Medicine, 22(15):2533–2534, August 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glueck:2003:APB

[1160] Deborah H. Glueck and Keith E. Muller. Adjusting power for a baseline covariate in linear models. Statistics in Medicine, 22(16):2535–2551, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demirtas:2003:PRC

[1161] Hakan Demirtas and Joseph L. Schafer. On the performance of randomcoefficient pattern-mixture models for non-ignorable drop-out. Statistics in Medicine, 22(16):2553–2575, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2003:ETR

[1162] C. Y. Wang and Yijian Huang. Error in timing in regression with observed longitudinal measurements. Statistics in Medicine, 22(16):2577– 2590, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:2003:SIL

[1163] Melissa D. Begg and Michael K. Parides. Separation of individual-level and cluster-level covariate effects in regression analysis of correlated data. Statistics in Medicine, 22(16):2591–2602, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:2003:TBG

[1164] Jørgen Holm Petersen. Two bivariate geometrically defined reference regions with applications to male reproductive hormones and human growth. Statistics in Medicine, 22(16):2603–2618, August 30, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moon:2003:EPT

[1165] Hojin Moon, Hongshik Ahn, Ralph L. Kodell, and J. Jack Lee. Estimation of k for the poly-k test with application to animal carcinogenicity studies. Statistics in Medicine, 22(16):2619–2636, August 30, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foulkes:2003:CCA

[1166] A. S. Foulkes and V. DeGruttola. Characterizing classes of antiretroviral drugs by genotype. Statistics in Medicine, 22(16):2637–2655, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2003:UMM

[1167] Gang Zheng. Use of max and min scores for trend tests for association when the genetic model is unknown. Statistics in Medicine, 22(16):2657– 2666, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2003:LES

[1168] Anonymous. Letter to the Editor: A simple confidence interval for meta-analysis. K. Sidik and J. N. Jonkman, Statistics in Medicine 2002; 21:3153–3159. Statistics in Medicine, 22(16):2667–2668, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [920].

Millard:2003:LEB

[1169] Steven P. Millard and Nagaraj K. Neerchal. Letter to the Editor: Book Review of Environmental Statistics with S-Plus. J. H. Petersen, Statistics in Medicine 2002; 21:2607–2608. Statistics in Medicine, 22(16):2668– 2670, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [869].

Petersen:2003:AR

[1170] Jørgen Holm Petersen. Author’s reply. Statistics in Medicine, 22(16): 2670–2672, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nixon:2003:BAB

[1171] R. M. Nixon and S. G. Thompson. Baseline adjustments for binary data in repeated cross-sectional cluster randomized trials. Statistics in Medicine, 22(17):2673–2692, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knapp:2003:ITR

[1172] Guido Knapp and Joachim Hartung. Improved tests for a random effects meta-regression with a single covariate. Statistics in Medicine, 22(17): 2693–2710, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kosinski:2003:GSA

[1173] Andrzej S. Kosinski and Huiman X. Barnhart. A global sensitivity analysis of performance of a medical diagnostic test when verification bias is present. Statistics in Medicine, 22(17):2711–2721, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quan:2003:ESD

[1174] Hui Quan and Ji Zhang. Estimate of standard deviation for a logtransformed variable using arithmetic means and standard deviations. Statistics in Medicine, 22(17):2723–2736, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1561].

Newcombe:2003:CIM

[1175] Robert G. Newcombe. Confidence intervals for the mean of a variable taking the values 0,1 and 2. Statistics in Medicine, 22(17):2737–2750, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moller:2003:PCI

[1176] Bjørn Møller, Harald Fekjær, Timo Hakulinen, Helgi Sigvaldason, Hans H. Storm, Mats Talbäck, and Tor Haldorsen. Prediction of cancer incidence in the Nordic countries: empirical comparison of different approaches. Statistics in Medicine, 22(17):2751–2766, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giorgi:2003:RSR

[1177] Roch Giorgi, Michal Abrahamowicz, Catherine Quantin, Philippe Bolard, Jacques Esteve, Joanny Gouvernet, and Jean Faivre. A relative survival regression model using B-spline functions to model nonproportional hazards. Statistics in Medicine, 22(17):2767–2784, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DiRienzo:2003:NPM

[1178] A. Gregory DiRienzo, Victor DeGruttola, Brendan Larder, and Kurt Hertogs. Non-parametric methods to predict HIV drug susceptibility phenotype from genotype. Statistics in Medicine, 22(17):2785–2798, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2003:CSR

[1179] Peter C. Austin, William A. Ghali, and Jack V. Tu. A comparison of several regression models for analysing cost of CABG surgery. Statistics in Medicine, 22(17):2799–2815, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gtzsche:2003:BRS

[1180] Peter C. Gtzsche. Book review: Statistical questions in evidence-based medicine. Martin Bland and Janet Peacock, Oxford University Press, Oxford, 2000. No. of pages: 251. Price: £19.95. ISBN 0-19-262992-1. Statistics in Medicine, 22(17):2817–2818, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ihaka:2003:BRC

[1181] Ross Ihaka. Book review: Computational visualisation graphics, abstraction and interactivity. Thomas Strothotte (ed.), Springer, Heidelberg, 1999. No. of pages: xxiii + 459. Price: £37.50. ISBN 3-540-63737-0. Statistics in Medicine, 22(17):2818–2819, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harbo:2003:BRA

[1182] Ingrid Sofie Harbo. Book review: Applied statistics in the pharmaceutical industry: With case studies using S-Plus. Steven P. Millard and Andreas Krause (eds), Springer Verlag, Heidelberg, Germany, 2001. No. of pages: xviii + 513. Price: £62.50. ISBN 0-387-98814-9. Statistics in Medicine, 22(17):2819–2820, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2003:CRM

[1183] Xun Chen and Lynn Wei. A comparison of recent methods for the analysis of small-sample cross-over studies. Statistics in Medicine, 22(18): 2821–2833, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2003:DSS

[1184] Tailiang Xie and Joel Waksman. Design and sample size estimation in clinical trials with clustered survival times as the primary endpoint. Statistics in Medicine, 22(18):2835–2846, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsuan:2003:AIB

[1185] Francis C. Hsuan and Russell Reeve. Assessing individual bioequivalence with high-order cross-over designs: a unified procedure. Statistics in Medicine, 22(18):2847–2860, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ismail:2003:OMR

[1186] Noor Azina Ismail, A. N. Pettitt, and R. A. Webster. ‘Online’ monitoring and retrospective analysis of hospital outcomes based on a scan statistic. Statistics in Medicine, 22(18):2861–2876, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spix:2003:LTO

[1187] Claudia Spix, Jörg Michaelis, Frank Berthold, Rudolf Erttmann, Johannes Sander, and Freimut H. Schilling. Lead-time and overdiagnosis estimation in neuroblastoma screening. Statistics in Medicine, 22(18): 2877–2892, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donaldson:2003:GLC

[1188] Gary W. Donaldson. General linear contrasts on latent variable means: structural equation hypothesis tests for multivariate clinical trials. Statistics in Medicine, 22(18):2893–2917, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Toledano:2003:TMA

[1189] Alicia Y. Toledano. Three methods for analysing correlated ROC curves: a comparison in real data sets from multi-reader, multi-case studies with a factorial design. Statistics in Medicine, 22(18):2919–2933, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Le:2003:QCI

[1190] Chap T. Le, Patricia M. Grambsch, and G. Scott Giebink. Quality control and the identification of vaccine responders using ELISA-derived antibody data. Statistics in Medicine, 22(18):2935–2942, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crook:2003:MSE

[1191] Angela M. Crook, Leonhard Knorr-Held, and Harry Hemingway. Measuring spatial effects in time to event data: a case study using months from angiography to coronary artery bypass graft (CABG). Statistics in Medicine, 22(18):2943–2961, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Posch:2003:BRS

[1192] Martin Posch. Book review: Statistical Methods for Clinical Trials. M. X. Norleans, Marcel Dekker, New York, 2001. No. of pages: 257. ISBN 0-8247-0467-3. Statistics in Medicine, 22(18):2963–2964, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nielsen:2003:BRS

[1193] Tine Nielsen. Book review: SPSS in Practice. An Illustrated Guide. 2nd edn. Pasant K. Puri, Arnold Publishers, London, 2002. No. of pages: x + 175. ISBN 0-340-76112-1. Statistics in Medicine, 22(18):2964–2965, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanseman:2003:BRE

[1194] Dennis Hanseman. Book review: Econometric Analysis of Health Data. Andrew Jones and Owen O’Donnell (eds), Wiley, Chichester, U.K., 2002. No. of pages: xi + 233. ISBN 0-470-84145-1. Statistics in Medicine, 22 (18):2965–2966, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:2003:BRS

[1195] Jørgen Hilden. Book review: Statistical Consulting. Javier Cabrera and Andrew McDougall, Springer-Verlag, New York, 2002. No. of pages: xii + 390. ISBN 0-387-98863-7. Statistics in Medicine, 22(18):2967, September 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simeone:2003:ESI

[1196] Ronald Simeone, Lynn Holland, and Roman Viveros-Aguilero. Estimating the size of an illicit-drug-using population. Statistics in Medicine, 22 (19):2969–2993, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ades:2003:CEM

[1197] A. E. Ades. A chain of evidence with mixed comparisons: models for multi-parameter synthesis and consistency of evidence. Statistics in Medicine, 22(19):2995–3016, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2003:MPW

[1198] Vance W. Berger, Anastasia Ivanova, and Maria Deloria Knoll. Minimizing predictability while retaining balance through the use of less restrictive randomization procedures. Statistics in Medicine, 22(19):3017–3028, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2003:NID

[1199] Ying Lu, Hua Jin, and Harry K. Genant. On the non-inferiority of a diagnostic test based on paired observations. Statistics in Medicine, 22 (19):3029–3044, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yafune:2003:SBC

[1200] Akifumi Yafune and Makio Ishiguro. A simulation-based confidence band for drug concentrations in blood: an application to a clinical phase I trial. Statistics in Medicine, 22(19):3045–3053, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muggeo:2003:ERM

[1201] Vito M. R. Muggeo. Estimating regression models with unknown breakpoints. Statistics in Medicine, 22(19):3055–3071, October 15, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2003:BAS

[1202] Sik-Yum Lee and Xin-Yuan Song. Bayesian analysis of structural equation models with dichotomous variables. Statistics in Medicine, 22(19): 3073–3088, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Budtz-Jorgensen:2003:CEM

[1203] Esben Budtz-Jørgensen, Niels Keiding, Philippe Grandjean, Pal Weihe, and Roberta F. White. Consequences of exposure measurement error for confounder identification in environmental epidemiology. Statistics in Medicine, 22(19):3089–3100, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2003:CTI

[1204] R. Chen and P. Froom. The CUSCORE test and the q-interval in cluster analyses of colon cancer and of lymphoma among asbestos workers. Statistics in Medicine, 22(19):3101–3109, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:2003:BRT

[1205] Jørgen Hilden. Book review: Testing Statistical Hypotheses of Equivalence. Stefan Wellek, Chapman and Hall/CRC, Boca Raton, 2003. No. of pages: xvi + 284. ISBN 1-58488-160-7. Statistics in Medicine, 22 (19):3111–3112, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwarzer:2003:BRM

[1206] Guido Schwarzer. Book review: Methods for Meta-analysis in Medical Research. Alex J. Sutton, Keith R. Abrams, David R. Jones, Trevor A. Sheldon and Fujian Song, Wiley, Chichester, U.K., 2000. No. of pages: xvii + 317. ISBN 0-471-49066-0. Statistics in Medicine, 22(19):3112– 3114, October 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Branson:2003:STB

[1207] Michael Branson and John Whitehead. A score test for binary data with patient non-compliance. Statistics in Medicine, 22(20):3115–3132, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sankoh:2003:EES

[1208] Abdul J. Sankoh, Ralph B. D’Agostino, Sr., and Mohammad F. Huque. Efficacy endpoint selection and multiplicity adjustment methods in clinical trials with inherent multiple endpoint issues. Statistics in Medicine, 22(20):3133–3150, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ancukiewicz:2003:MRB

[1209] Marek Ancukiewicz, Dianne M. Finkelstein, and David A. Schoenfeld. Modelling the relationship between continuous covariates and clinical events using isotonic regression. Statistics in Medicine, 22(20):3151– 3159, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Touloumi:2003:CTM

[1210] G. Touloumi, A. G. Babiker, M. G. Kenward, S. J. Pocock, and J. H. Darbyshire. A comparison of two methods for the estimation of precision with incomplete longitudinal data, jointly modelled with a time-to-event outcome. Statistics in Medicine, 22(20):3161–3175, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rice:2003:FLA

[1211] Kenneth Rice. Full-likelihood approaches to misclassification of a binary exposure in matched case-control studies. Statistics in Medicine, 22 (20):3177–3194, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mikulich:2003:CLN

[1212] Susan K. Mikulich, Gary O. Zerbe, Richard H. Jones, and Thomas J. Crowley. Comparing linear and nonlinear mixed model approaches to cosinor analysis. Statistics in Medicine, 22(20):3195–3211, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gangnon:2003:HMS

[1213] Ronald E. Gangnon and Murray K. Clayton. A hierarchical model for spatially clustered disease rates. Statistics in Medicine, 22(20):3213– 3228, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keen:2003:RAS

[1214] K. J. Keen and Robert C. Elston. Robust asymptotic sampling theory for correlations in pedigrees. Statistics in Medicine, 22(20):3229–3247, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gagnon:2003:ASD

[1215] David R. Gagnon, Mark E. Glickman, Richard H. Myers, and L. Adrienne Cupples. The analysis of survival data with a non-susceptible fraction and dual censoring mechanisms. Statistics in Medicine, 22(20):3249– 3262, October 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gehan:2003:P

[1216] Edmund A. Gehan. Preface. Statistics in Medicine, 22(21):3263–3264, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2003:PBS

[1217] Anonymous. Perspectives on the biostatistical sciences: a symposium in memory of Samuel W. Greenhouse. Statistics in Medicine, 22(21):3265– 3266, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2003:TSW

[1218] John M. Lachin. A tribute to Samuel W. Greenhouse. Statistics in Medicine, 22(21):3267–3276, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keller:2003:SGY

[1219] Jay Keller and Cynthia Z. F. Clark. Sam Greenhouse’s years at the Census Bureau and the UNRRA. Statistics in Medicine, 22(21):3277– 3279, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Butler:2003:SGS

[1220] Robert N. Butler. Sam Greenhouse’s scholarly contributions to human aging studies. Statistics in Medicine, 22(21):3281–3284, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ONeill:2003:SGH

[1221] Robert T. O’Neill. Sam Greenhouse: his contributions as a consultant to the Food and Drug Administration. Statistics in Medicine, 22(21): 3285–3289, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katz:2003:CSG

[1222] Martin M. Katz and Nancy Berman. The contributions of Sam Greenhouse to research on psychiatric diagnosis and psychopharmacology. Statistics in Medicine, 22(21):3291–3299, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:2003:SSG

[1223] Mitchell H. Gail. Some of Sam Greenhouse’s contributions to statistical methods. Statistics in Medicine, 22(21):3301–3308, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:2003:I

[1224] Joel B. Greenhouse. Introduction. Statistics in Medicine, 22(21):3309– 3310, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:2003:PES

[1225] Samuel W. Greenhouse. On psychiatry, epidemiology, and statistics: a view from the 1950s and 1960s. Statistics in Medicine, 22(21):3311–3322, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:2003:GFB

[1226] Samuel W. Greenhouse. The growth and future of biostatistics: (a view from the 1980s). Statistics in Medicine, 22(21):3323–3335, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geisser:2003:APD

[1227] Seymour Geisser. The analysis of profile data-revisited. Statistics in Medicine, 22(21):3337–3346, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lan:2003:RGS

[1228] K. K. Gordon Lan, John M. Lachin, and Oliver Bautista. Over-ruling a group sequential boundary — a stopping rule versus a guideline. Statistics in Medicine, 22(21):3347–3355, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2003:GSL

[1229] John M. Lachin, Samuel W. Greenhouse, and Oliver M. Bautista. Group sequential large sample T^{2}-like χ^{2} tests for multivariate observations. Statistics in Medicine, 22(21):3357–3368, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberg:2003:QER

[1230] Philip S. Rosenberg, Hormuzd Katki, Christine A. Swanson, Linda M. Brown, Sholom Wacholder, and Robert N. Hoover. Quantifying epidemiologic risk factors using non-parametric regression: model selection remains the greatest challenge. Statistics in Medicine, 22(21):3369–3381, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2003:URC

[1231] Binbing Yu and Joseph L. Gastwirth. The use of the ‘reverse Cornfield inequality’ to assess the sensitivity of a non-significant association to an omitted variable. Statistics in Medicine, 22(21):3383–3401, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:2003:EPS

[1232] Eugene M. Laska, Morris Meisner, Joseph Wanderling, and Carole Siegel. Estimating population size and duplication rates when records cannot be linked. Statistics in Medicine, 22(21):3403–3417, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seigel:2003:CTE

[1233] Daniel Seigel. Clinical trials, epidemiology, and public confidence. Statistics in Medicine, 22(21):3419–3425, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zelen:2003:TBS

[1234] M. Zelen. The training of biostatistical scientists. Statistics in Medicine, 22(21):3427–3430, November 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2003:MMC

[1235] Thomas D. Cook. Methods for mid-course corrections in clinical trials with survival outcomes. Statistics in Medicine, 22(22):3431–3447, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2003:PTE

[1236] Cong Chen, Hongwei Wang, and Steven M. Snapinn. Proportion of treatment effect (PTE) explained by a surrogate marker. Statistics in Medicine, 22(22):3449–3459, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Polansky:2003:SBT

[1237] Alan M. Polansky. Selecting the best treatment in designed experiments. Statistics in Medicine, 22(22):3461–3471, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parker:2003:RMW

[1238] Corette B. Parker and Elizabeth R. DeLong. ROC methodology within a monitoring framework. Statistics in Medicine, 22(22):3473–3488, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2003:RBI

[1239] Kai Li and Dale J. Poirier. The roles of birth inputs and outputs in predicting health, behaviour and test scores in early childhood. Statistics in Medicine, 22(22):3489–3514, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kattan:2003:CRN

[1240] Michael W. Kattan, Glenn Heller, and Murray F. Brennan. A competingrisks nomogram for sarcoma-specific death following local recurrence. Statistics in Medicine, 22(22):3515–3525, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:2003:MEV

[1241] Ronghui Xu. Measuring explained variation in linear mixed effects models. Statistics in Medicine, 22(22):3527–3541, November 30, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carvalho:2003:MDT

[1242] Marilia Sá Carvalho and Leonhard Knorr-Held. Modelling discrete time survival data with random slopes: evaluating haemodialysis centres. Statistics in Medicine, 22(22):3543–3555, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ogden:2003:EKP

[1243] R. Todd Ogden. Estimation of kinetic parameters in graphical analysis of PET imaging data. Statistics in Medicine, 22(22):3557–3568, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2003:BRA

[1244] Weijing Wang. Book review: 1. Analysis of Failure and Survival Data. Peter J. Smith, CRC, Boca Raton, U.S.A., 2002. No. of pages: 254. ISBN 1-58488-075-9. Statistics in Medicine, 22(22):3569–3570, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gronbaek:2003:BRM

[1245] Morten Grønbæk. Book review: Making Sense of Data. A Self-Instruction Manual on the Interpretation of Epidemiological Data. 3rd Edition. J. H. Abramson and Z. H. Abramson, Oxford University Press, Oxford, 2001. No. of pages: xvi + 367. ISBN 0-19-514525-9. Statistics in Medicine, 22(22):3570, November 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kieser:2003:SPB

[1246] Meinhard Kieser and Tim Friede. Simple procedures for blinded sample size adjustment that do not affect the type I error rate. Statistics in Medicine, 22(23):3571–3581, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:2003:NNP

[1247] Emmanuel Lesaffre and Stephen Senn. A note on non-parametric AN-COVA for covariate adjustment in randomized clinical trials. Statistics in Medicine, 22(23):3583–3596, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1696].

Nardi:2003:CCP

[1248] Alessandra Nardi and Michael Schemper. Comparing Cox and parametric models in clinical studies. Statistics in Medicine, 22(23):3597–3610, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2003:UIP

[1249] Debbie M. Cheng and Stephen W. Lagakos. Use of incomplete posttreatment data in the analysis of viral eradication studies. Statistics in Medicine, 22(23):3611–3628, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weusten:2003:SSQ

[1250] Jos J. A. M. Weusten, Pieter A. W. M. Wouters, and Martien C. A. van Zuijlen. Sensitivity studies for quantitative assays: use of censored data analysis. Statistics in Medicine, 22(23):3629–3640, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosen:2003:AGC

[1251] Ori Rosen and Ayala Cohen. Analysis of growth curves via mixtures. Statistics in Medicine, 22(23):3641–3654, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2003:MRL

[1252] Xuming He, Bo Fu, and Wing K. Fung. Median regression for longitudinal data. Statistics in Medicine, 22(23):3655–3669, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wall:2003:CMR

[1253] Melanie M. Wall and Ruifeng Li. Comparison of multiple regression to two latent variable techniques for estimation and prediction. Statistics in Medicine, 22(23):3671–3685, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelhalter:2003:BAM

[1254] David J. Spiegelhalter and Nicola G. Best. Bayesian approaches to multiple sources of evidence and uncertainty in complex cost-effectiveness modelling. Statistics in Medicine, 22(23):3687–3709, December 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2003:P

[1255] Stephen Senn, Harbajan Chadha-Boreham, and Catherine Quantin. Preface. Statistics in Medicine, 22(24):3711, December 30, 2003. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:2003:EIS

[1256] Nico J. D. Nagelkerke, Hendriek C. Boshuizen, Hester E. de Melker, Joop F. P. Schellekens, Marcel F. Peeters, and Marina Conyn van Spaendonck. Estimating the incidence of subclinical infections with legionella pneumonia using data augmentation: analysis of an outbreak in the Netherlands. Statistics in Medicine, 22(24):3713–3724, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friede:2003:IEO

[1257] Tim Friede and Robin Henderson. Intervention effects in observational survival studies with an application in total hip replacements. Statistics in Medicine, 22(24):3725–3737, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davidov:2003:AIB

[1258] Ori Davidov, Havi Murad, Ayala Lusky, Eric S. Shinwell, Brian Reichman, and Laurence Freedman. The analysis of the influence of birth order and other factors in multiple birth data. Statistics in Medicine, 22(24): 3739–3753, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saint-Pierre:2003:AAC

[1259] P. Saint-Pierre, C. Combescure, Jp. Daurès, and P. Godard. The analysis of asthma control under a Markov assumption with use of covariates. Statistics in Medicine, 22(24):3755–3770, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Panhard:2003:CGS

[1260] Xavière Panhard, Sébastien Dominique, Marie-Pierre Gaub, Vincent Ravery, Bernard Grandchamp, and France Mentré. Construction of a global score quantifying allelic imbalance among biallelic SIDP markers in bladder cancer. Statistics in Medicine, 22(24):3771–3779, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leffondre:2003:ECM

[1261] Karen Leffondré, Michal Abrahamowicz, and Jack Siemiatycki. Evaluation of Cox’s model and logistic regression for matched case-control data with time-dependent covariates: a simulation study. Statistics in Medicine, 22(24):3781–3794, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:2003:CCC

[1262] Liselotte Petersen, Thorkild I. A. Sørensen, and Per Kragh Andersen. Comparison of case-cohort estimators based on data on premature death of adult adoptees. Statistics in Medicine, 22(24):3795–3803, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ukoumunne:2003:NPB

[1263] Obioha C. Ukoumunne, Anthony C. Davison, Martin C. Gulliford, and Susan Chinn. Non-parametric bootstrap confidence intervals for the intraclass correlation coefficient. Statistics in Medicine, 22(24):3805–3821, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:2003:PMC

[1264] Peter A. Lachenbruch. Proper metrics for clinical trials: transformations and other procedures to remove non-normality effects. Statistics in Medicine, 22(24):3823–3842, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morgan:2003:SSR

[1265] Caroline C. Morgan. Sample size re-estimation in group-sequential response-adaptive clinical trials. Statistics in Medicine, 22(24):3843– 3857, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lydersen:2003:PCT

[1266] Stian Lydersen and Petter Laake. Power comparison of two-sided exact tests for association in 2 × 2 contingency tables using standard, mid p and randomized test versions. Statistics in Medicine, 22(24):3859–3871, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wienke:2003:HCS

[1267] Andreas Wienke, Niels V. Holm, Kaare Christensen, Axel Skytthe, James W. Vaupel, and Anatoli I. Yashin. The heritability of causespecific mortality: a correlated gamma-frailty model applied to mortality due to respiratory diseases in Danish twins born 1870–1930. Statistics in Medicine, 22(24):3873–3887, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ballone:2003:PAD

[1268] Enzo Ballone, Vittorio Colagrande, Marta Di Nicola, Rocco Di Mascio, Claudio Di Mascio, and Fabio Capani. Probabilistic approach to developing nephropathy in diabetic patients with retinopathy. Statistics in Medicine, 22(24):3889–3897, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kendziorski:2003:PEB

[1269] C. M. Kendziorski, M. A. Newton, H. Lan, and M. N. Gould. On parametric empirical Bayes methods for comparing multiple groups using replicated gene expression profiles. Statistics in Medicine, 22(24):3899– 3914, December 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2004:SMM

[1270] Ralph B. D’Agostino, Sr., Michael J. Campbell, Joel B. Greenhouse, and John N. S. Matthews. Statistics in Medicine moves into online submission. Statistics in Medicine, 23(1):1, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shi:2004:MAT

[1271] Jian Qing Shi and J. B. Copas. Meta-analysis for trend estimation. Statistics in Medicine, 23(1):3–19, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonzo:2004:SSE

[1272] Todd A. Alonzo, Thomas M. Braun, and Chaya S. Moskowitz. Small sample estimation of relative accuracy for binary screening tests. Statistics in Medicine, 23(1):21–34, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2004:EHT

[1273] Haitao Chu, Marie-Pierre Préziosi, and M. Elizabeth Halloran. Estimating heterogeneous transmission with multiple infectives using MCMC methods. Statistics in Medicine, 23(1):35–49, January 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dickman:2004:RMR

[1274] Paul W. Dickman, Andy Sloggett, Michael Hills, and Timo Hakulinen. Regression models for relative survival. Statistics in Medicine, 23(1): 51–64, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agresti:2004:ENE

[1275] Alan Agresti and Yongyi Min. Effects and non-effects of paired identical observations in comparing proportions with binary matched-pairs data. Statistics in Medicine, 23(1):65–75, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1450].

Hothorn:2004:BST

[1276] Torsten Hothorn, Berthold Lausen, Axel Benner, and Martin Radespiel-Tröger. Bagging survival trees. Statistics in Medicine, 23(1):77–91, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Plummer:2004:IEF

[1277] Martyn Plummer. Improved estimates of floating absolute risk. Statistics in Medicine, 23(1):93–104, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Magni:2004:AIV

[1278] Paolo Magni and Riccardo Bellazzi. Analysing Italian voluntary abortion data using a Bayesian approach to the time series decomposition. Statistics in Medicine, 23(1):105–123, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goehring:2004:SBQ

[1279] Catherine Goehring, Arnaud Perrier, and Alfredo Morabia. Spectrum bias: a quantitative and graphical analysis of the variability of medical diagnostic test performance. Statistics in Medicine, 23(1):125–135, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1728].

Islam:2004:MMM

[1280] M. Ataharul Islam, Rafiqul Islam Chowdhury, Nitai Chakraborty, and Wasimul Bari. A multistage model for maternal morbidity during antenatal, delivery and postpartum periods. Statistics in Medicine, 23(1): 137–158, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sidik:2004:AR

[1281] Kurex Sidik and Jeffrey N. Jonkman. Authors’ reply. Statistics in Medicine, 23(1):159–162, January 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kryscio:2004:F

[1282] Richard J. Kryscio and Frederick A. Schmitt. Foreword. Statistics in Medicine, 23(2):163–167, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2004:SMA

[1283] Anonymous. Statistical methodology in Alzheimer’s disease research II. Statistics in Medicine, 23(2):168, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edland:2004:GAS

[1284] Steven D. Edland, Susan Slager, and Matthew Farrer. Genetic association studies in Alzheimer’s disease research: challenges and opportunities. Statistics in Medicine, 23(2):169–178, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lange:2004:FBA

[1285] Christoph Lange, Deborah Blacker, and Nan M. Laird. Family-based association tests for survival and times-to-onset analysis. Statistics in Medicine, 23(2):179–189, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Launer:2004:PBS

[1286] L. J. Launer and D. B. Brock. Population-based studies of AD: message and methods: an epidemiologic view. Statistics in Medicine, 23(2):191– 197, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Commenges:2004:IMA

[1287] D. Commenges, P. Joly, L. Letenneur, and Jf. Dartigues. Incidence and mortality of Alzheimer’s disease or dementia using an illness-death model. Statistics in Medicine, 23(2):199–210, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2004:SRE

[1288] Sujuan Gao. A shared random effect parameter approach for longitudinal dementia data with non-ignorable missing data. Statistics in Medicine, 23(2):211–219, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2004:ANI

[1289] Xiao-Hua Zhou and Pete Castelluccio. Adjusting for non-ignorable verification bias in clinical studies for Alzheimer’s disease. Statistics in Medicine, 23(2):221–230, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beckett:2004:MLM

[1290] L. A. Beckett, D. J. Tancredi, and R. S. Wilson. Multivariate longitudinal models for complex change processes. Statistics in Medicine, 23(2):231– 239, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crane:2004:TBC

[1291] Paul K. Crane, Gerald van Belle, and Eric B. Larson. Test bias in a cognitive test: differential item functioning in the CASI. Statistics in Medicine, 23(2):241–256, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:2004:ROR

[1292] Helena Chmura Kraemer. Reconsidering the odds ratio as a measure of 2 × 2 association in a population. Statistics in Medicine, 23(2):257– 270, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2004:SPS

[1293] Lang Li, Jeffrey Huang, Sharon Sun, Jianzhao Shen, Frederick W. Unverzagt, Sujuan Gao, Hugh H. Hendrie, Kathleen Hall, and Siu L. Hui. Selecting pre-screening items for early intervention trials of dementia — a case study. Statistics in Medicine, 23(2):271–283, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kryscio:2004:DLP

[1294] Richard J. Kryscio, Marta S. Mendiondo, Frederick A. Schmitt, and William R. Markesbery. Designing a large prevention trial: statistical issues. Statistics in Medicine, 23(2):285–296, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2004:APR

[1295] Susan S. Ellenberg. Analytical, practical and regulatory issues in prevention studies. Statistics in Medicine, 23(2):297–303, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mani:2004:EDM

[1296] Ranjit B. Mani. The evaluation of disease modifying therapies in Alzheimer’s disease: a regulatory viewpoint. Statistics in Medicine, 23 (2):305–314, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Higdon:2004:CCM

[1297] Roger Higdon, Norman L. Foster, Robert A. Koeppe, Charles S. De-Carli, William J. Jagust, Christopher M. Clark, Nancy R. Barbas, Steven E. Arnold, R. Scott Turner, Judith L. Heidebrink, and Satoshi Minoshima. A comparison of classification methods for differentiating fronto-temporal dementia from Alzheimer’s disease using FDG-PET imaging. Statistics in Medicine, 23(2):315–326, January 30, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lange:2004:TMM

[1298] Nicholas Lange, Stephen Lake, Reisa Sperling, John Brown, Carol Routledge, Marilyn Albert, and Stephan Heckers. Two macroscopic and microscopic brain imaging studies of human hippocampus in early Alzheimer’s disease and schizophrenia research. Statistics in Medicine, 23(2):327–350, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Massaro:2004:MAD

[1299] Joseph M. Massaro, Ralph B. D’Agostino, Sr., Lisa M. Sullivan, Alexa Beiser, Charles DeCarli, Rhoda Au, Merrill F. Elias, and Philip A. Wolf. Managing and analysing data from a large-scale study on Framingham Offspring relating brain structure to cognitive function. Statistics in Medicine, 23(2):351–367, January 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glidden:2004:MCS

[1300] David V. Glidden and Eric Vittinghoff. Modelling clustered survival data from multicentre clinical trials. Statistics in Medicine, 23(3):369– 388, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2004:BMA

[1301] Simon G. Thompson, David E. Warn, and Rebecca M. Turner. Bayesian methods for analysis of binary outcome data in cluster randomized trials on the absolute risk scale. Statistics in Medicine, 23(3):389–410, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2004:ELR

[1302] Peter C. Austin and Jeffrey S. Hoch. Estimating linear regression models in the presence of a censored independent variable. Statistics in Medicine, 23(3):411–429, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Straatman:2004:ELE

[1303] Huub Straatman, André L. M. Verbeek, Petronella G. M. Peer, and George Borm. Estimating life expectancy and related probabilities in screen-detected breast cancer patients with restricted follow-up information. Statistics in Medicine, 23(3):431–448, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pawitan:2004:EGE

[1304] Y. Pawitan, M. Reilly, E. Nilsson, S. Cnattingius, and P. Lichtenstein. Estimation of genetic and environmental factors for binary traits using family data. Statistics in Medicine, 23(3):449–465, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1711].

Cui:2004:ERD

[1305] Jisheng S. Cui, Paul S. F. Yip, and P. H. Chau. Estimation of reporting delay and suicide incidence in Hong Kong. Statistics in Medicine, 23 (3):467–476, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:2004:ODM

[1306] J. N. S. Matthews and G. C. Allcock. Optimal designs for Michaelis– Menten kinetic studies. Statistics in Medicine, 23(3):477–491, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2004:LND

[1307] Byron J. Gajewski, Jack D. Sedwick, and Patrick J. Antonelli. A lognormal distribution model of the effect of bacteria and ear fenestration on hearing loss: a Bayesian approach. Statistics in Medicine, 23(3):493– 508, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carey:2004:GVA

[1308] V. J. Carey, F. H. Yong, L. M. Frenkel, and R. M. McKinney. Growth velocity assessment in paediatric AIDS: smoothing, penalized quantile regression and the definition of growth failure. Statistics in Medicine, 23(3):509–526, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knapp:2004:BRM

[1309] Guido Knapp. Book review: Meta-Analysis of Controlled Clinical Trials. Anne Whitehead, Wiley, Chichester, 2002. No. of pages: xiv + 337. ISBN 0-471-98370-5. Statistics in Medicine, 23(3):527–528, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Houwelingen:2004:BRE

[1310] Hans C. Van Houwelingen. Book review: The Elements of Statistical Learning, Data Mining, Inference, and Prediction. Trevor Hastie, Robert Tibshirani and Jerome Friedman, Springer, New York, 2001. No. of pages: xvi + 533. ISBN 0-387-95284-5. Statistics in Medicine, 23(3): 528–529, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julious:2004:BRM

[1311] Steven A. Julious. Book review: A Manager’s Guide to the Design and Conduct of Clinical Trials. Phillip I. Good (ed.), Wiley, Chichester, U.K., 2002. No. of pages: xiv + 228. ISBN: 0-471-22615-7. Statistics in Medicine, 23(3):529–530, February 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vrijens:2004:EMV

[1312] Bernard Vrijens and Els Goetghebeur. Electronic monitoring of variation in drug intakes can reduce bias and improve precision in pharmacokinetic/pharmacodynamic population studies. Statistics in Medicine, 23(4):531–544, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2004:TNI

[1313] Kung-Jong Lui and Xiao-Hua Zhou. Testing non-inferiority (and equivalence) between two diagnostic procedures in paired-sample ordinal data.

Statistics in Medicine, 23(4):545–559, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2004:ATS

[1314] Sin-Ho Jung, Taiyeong Lee, KyungMann Kim, and Stephen L. George. Admissible two-stage designs for phase II cancer clinical trials. Statistics in Medicine, 23(4):561–569, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walker:2004:PRB

[1315] A. Sarah Walker, Ian R. White, and Abdel G. Babiker. Parametric randomization-based methods for correcting for treatment changes in the assessment of the causal effect of treatment. Statistics in Medicine, 23(4):571–590, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2004:TSS

[1316] Chu-Chih Chen, Kuen-Yuh Wu, and N. Balakrishnan. Two-sample scale tests for comparison of metabolic rates for styrene in previously exposed and unexposed groups. Statistics in Medicine, 23(4):591–599, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Angers:2004:BRB

[1317] Jean-François Angers and Atanu Biswas. Book review: A Bayesian analysis of the 4-year follow-up data of the Wisconsin epidemiologic study of diabetic retinopathy. Statistics in Medicine, 23(4):601–615, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moger:2004:ATC

[1318] Tron A. Moger, Odd O. Aalen, Ketil Heimdal, and Håkon K. Gjessing. Analysis of testicular cancer data using a frailty modelwith familial dependence. Statistics in Medicine, 23(4):617–632, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2004:SMN

[1319] Tony Hsiu-Hsi Chen, Ming-Fang Yen, Ming-Neng Shiu, Tao-Hsin Tung, and Hui-Min Wu. Stochastic model for non-standard case-cohort design. Statistics in Medicine, 23(4):633–647, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hardt:2004:ALA

[1320] J. Hardt, F. Petrak, D. Filipas, and U. T. Egle. Adaptation to life after surgical removal of the bladder — an application of graphical Markov models for analysing longitudinal data. Statistics in Medicine, 23(4): 649–666, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2004:BLV

[1321] Z. G. Hu, C. M. Wong, T. Q. Thach, T. H. Lam, and A. J. Hedley. Binary latent variable modelling and its applicationin the study of air pollution in Hong Kong. Statistics in Medicine, 23(4):667–684, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2004:CCL

[1322] Man-Lai Tang and Hon Keung Tony Ng. Comment on: Confidence limits for the ratio of two rates based on likelihood scores: non-iterative method. Statistics in Medicine, 23(4):685–691, February 28, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graham:2004:ARa

[1323] P. L. Graham and K. Mengersen. Authors’ reply. Statistics in Medicine, 23(4):691–693, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siev:2004:LE

[1324] David Siev. Letter to the Editor. Statistics in Medicine, 23(4):693, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graham:2004:ARb

[1325] Petra Graham. Author’s reply. Statistics in Medicine, 23(4):694, February 28, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:2004:E

[1326] Mike Campbell. Editorial. Statistics in Medicine, 23(5):695, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grouin:2004:ABC

[1327] Jean-Marie Grouin, Simon Day, and John Lewis. Adjustment for baseline covariates: an introductory note. Statistics in Medicine, 23(5):697–699, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2004:CPM

[1328] Anonymous. Committee for proprietary medicinal products (CPMP) points to consider on adjustment for baseline covariates. Statistics in Medicine, 23(5):701–709, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2004:TQI

[1329] Xin Yan. Test for qualitative interaction in equivalence trials when the number of centres is large. Statistics in Medicine, 23(5):711–722, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2004:NMP

[1330] Patrick Royston and Willi Sauerbrei. A new measure of prognostic separation in survival data. Statistics in Medicine, 23(5):723–748, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2004:SAU

[1331] Babette A. Brumback, Miguel A. Hernán, Sebastien J. P. A. Haneuse, and James M. Robins. Sensitivity analyses for unmeasured confounding assuming a marginal structural model for repeated measures. Statistics in Medicine, 23(5):749–767, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gray:2004:MHR

[1332] Linsay Gray, Mario Cortina-Borja, and Marie-Louise Newell. Modelling HIV-RNA viral load in vertically infected children. Statistics in Medicine, 23(5):769–781, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knafl:2004:EMD

[1333] George J. Knafl, Kristopher P. Fennie, Carol Bova, Kevin Dieckhaus, and Ann B. Williams. Electronic monitoring device event modelling on an individual-subject basis using adaptive Poisson regression. Statistics in Medicine, 23(5):783–801, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troiani:2004:CBC

[1334] John S. Troiani and Bradley P. Carlin. Comparison of Bayesian, classical, and heuristic approaches in identifying acute disease events in lung transplant recipients. Statistics in Medicine, 23(5):803–824, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ripley:2004:NLS

[1335] Ruth M. Ripley, Adrian L. Harris, and Lionel Tarassenko. Non-linear survival analysis using neural networks. Statistics in Medicine, 23(5): 825–842, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parmigiani:2004:UVD

[1336] Giovanni Parmigiani. Uncertainty and the value of diagnostic information, with application to axillary lymph node dissection in breast cancer. Statistics in Medicine, 23(5):843–855, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Good:2004:LEE

[1337] Phillip I. Good. Letter to the Editor: Efficiency comparisons of rank and permutation tests by Statistics in Medicine 2001; 20:705–731. Statistics in Medicine, 23(5):857, March 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [391].

Yan:2004:EEA

[1338] Jun Yan and Jason Fine. Estimating equations for association structures. Statistics in Medicine, 23(6):859–874, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Franke:2004:GEE

[1339] D. Franke, C. Kastner, and A. Ziegler. Generalized estimating equations for association structures: familial correlations of lipid profiles. Statistics in Medicine, 23(6):875–877, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2004:RFK

[1340] Jun Yan and Jason Fine. Rejoinder to Franke, Kastner and Ziegler. Statistics in Medicine, 23(6):879–880, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2004:EBP

[1341] Sin-Ho Jung and Kyung Mann Kim. On the estimation of the binomial probability in multistage clinical trials. Statistics in Medicine, 23(6): 881–896, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewsey:2004:CCS

[1342] J. D. Lewsey. Comparing completely and stratified randomized designs in cluster randomized trials when the stratifying factor is cluster size: a simulation study. Statistics in Medicine, 23(6):897–905, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2004:CVP

[1343] Patrick Royston, Mahesh K. B. Parmar, and Richard Sylvester. Construction and validation of a prognostic model across several studies, with an application in superficial bladder cancer. Statistics in Medicine, 23(6):907–926, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ranta:2004:PMD

[1344] J. Ranta, P. H. Mäkelä, and E. Arjas. Predicting meningococcal disease outbreaks in structured populations. Statistics in Medicine, 23(6):927– 945, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tibaldi:2004:PLE

[1345] Fabián Tibaldi, Geert Molenberghs, Tomasz Burzykowski, and Helena Geys. Pseudo-likelihood estimation for a marginal multivariate survival model. Statistics in Medicine, 23(6):947–963, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowman:2004:MCI

[1346] F. DuBois Bowman and Lance A. Waller. Modelling of cardiac imaging data with spatial correlation. Statistics in Medicine, 23(6):965–985, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:2004:EMG

[1347] M. Y. Wong, N. E. Day, J. A. Luan, and N. J. Wareham. Estimation of magnitude in gene-environment interactions in the presence of measurement error. Statistics in Medicine, 23(6):987–998, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pulkstenis:2004:GFT

[1348] Erik Pulkstenis and Timothy J. Robinson. Goodness-of-fit tests for ordinal response regression models. Statistics in Medicine, 23(6):999–1014, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cleophas:2004:RDC

[1349] Ton J. Cleophas. Research data closer to expectation than compatible with random sampling. Statistics in Medicine, 23(6):1015–1017, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2004:C

[1350] Philip J. Smith. Correction. Statistics in Medicine, 23(6):1018, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bickeboller:2004:BRH

[1351] Heike Bickeböller. Book review: Handbook of Statistical Genetics D. J. Balding, M. Bishop, C. Cannings (eds), Wiley, Chichester, UK, 2001. No. of pages: xix + 863. ISBN: 0-471-86094-8. Statistics in Medicine, 23(6):1019–1020, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heller:2004:BRS

[1352] Gillian Z. Heller. Book review: Survival Analysis: Techniques For Truncated And Censored Data, 2nd Edition. John P. Klein and Melvin L. Moeschberger, Springer-Verlag, New York, 2003. Hardcover. No. of pages: xv + 536. ISBN 0-387-95399-X. Statistics in Medicine, 23(6): 1020–1021, March 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2004:ISS

[1353] Y. H. Joshua Chen, David L. DeMets, and K. K. Gordon Lan. Increasing the sample size when the unblinded interim result is promising. Statistics in Medicine, 23(7):1023–1038, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Minini:2004:SAL

[1354] Pascal Minini and Michel Chavance. Sensitivity analysis of longitudinal normal data with drop-outs. Statistics in Medicine, 23(7):1039–1054, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abraham:2004:RBD

[1355] Christophe Abraham and Jean-Pierre Daurès. Robust Bayesian decision theory applied to optimal dosage. Statistics in Medicine, 23(7):1055– 1073, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moodie:2004:NPP

[1356] Patricia F. Moodie, Norma A. Nelson, and Gary G. Koch. A nonparametric procedure for evaluating treatment effect in the meta-analysis of survival data. Statistics in Medicine, 23(7):1075–1093, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McInturff:2004:MRW

[1357] Pat McInturff, Wesley O. Johnson, David Cowling, and Ian A. Gardner. Modelling risk when binary outcomes are subject to error. Statistics in Medicine, 23(7):1095–1109, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tyrer:2004:BCP

[1358] Jonathan Tyrer, Stephen W. Duffy, and Jack Cuzick. A breast cancer prediction model incorporating familial and personal risk factors. Statistics in Medicine, 23(7):1111–1130, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1591].

Goodall:2004:ICS

[1359] Ruth L. Goodall, David T. Dunn, and Abdel G. Babiker. Intervalcensored survival time data: confidence intervals for the non-parametric survivor function. Statistics in Medicine, 23(7):1131–1145, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neumann:2004:RSM

[1360] Anke Neumann, Josiane Holstein, Gilles Chatellier, and Eric Lepage. A regression shrinkage method tailored to qualitative regressors and clustered data. Statistics in Medicine, 23(7):1147–1157, April 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2004:ITE

[1361] Peter C. Austin and Lawrence J. Brunner. Inflation of the type I error rate when a continuous confounding variable is categorized in logistic regression analyses. Statistics in Medicine, 23(7):1159–1178, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:2004:LEM

[1362] Devan V. Mehrotra. Letter to the Editor: A method for comparing two normal means using combined samples of correlated and uncorrelated data by S. W. Looney and P. W. Jones, Statistics in Medicine 2003; 22:1601–1610. Statistics in Medicine, 23(7):1179–1180, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1085].

Nielsen:2004:BRS

[1363] SØren Feodor Nielsen. Book review: 1. Statistical Analysis with Missing Data (2nd edn). Roderick J. Little and Donald B. Rubin, John Wiley and Sons, New York, 2002. No. of pages: xv + 381. ISBN: 0-471-18386-5. Statistics in Medicine, 23(7):1181, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burr:2004:BRS

[1364] Tom Burr. Book review: Spatial Cluster Modelling. Andrew B. Lawson and David G. T. Denison (eds), Chapman and Hall, CRC, Boca Raton, 2002. No. of pages: xiv + 287. ISBN: 1-58488-266-2. Statistics in Medicine, 23(7):1182–1184, April 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2004:ACD

[1365] Shein-Chung Chow and Jun Shao. Analysis of clinical data with breached blindness. Statistics in Medicine, 23(8):1185–1193, April 30, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [1645, 2002].

Turner:2004:AII

[1366] Rebecca M. Turner, A. Toby Prevost, and Simon G. Thompson. Allowing for imprecision of the intracluster correlation coefficient in the design of cluster randomized trials. Statistics in Medicine, 23(8):1195–1214, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shaffer:2004:BIR

[1367] Michele L. Shaffer and Vernon M. Chinchilli. Bayesian inference for randomized clinical trials with treatment failures. Statistics in Medicine, 23(8):1215–1228, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoover:2004:SAS

[1368] Donald R. Hoover. Subject allocation and study curtailment for fixed event comparative Poisson trials. Statistics in Medicine, 23(8):1229– 1245, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ismail:2004:SDH

[1369] Noor Azina Ismail and A. N. Pettitt. Smoothing a discrete hazard function for the number of patients colonized with methicillin-resistant Staphylococcus aureus in an intensive care unit. Statistics in Medicine, 23(8):1247–1258, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2004:TVM

[1370] Kelly H. Zou, William M. Wells III, Ron Kikinis, and Simon K. Warfield. Three validation metrics for automated probabilistic image segmentation of brain tumours. Statistics in Medicine, 23(8):1259–1282, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2004:EIE

[1371] Kung-Jong Lui. Eight interval estimators of a common rate ratio under stratified Poisson sampling. Statistics in Medicine, 23(8):1283–1296, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heitjan:2004:BEC

[1372] Daniel F. Heitjan, Clara Yuri Kim, and Huiling Li. Bayesian estimation of cost-effectiveness from censored data. Statistics in Medicine, 23(8): 1297–1309, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nixon:2004:PMC

[1373] R. M. Nixon and S. G. Thompson. Parametric modelling of cost data in medical studies. Statistics in Medicine, 23(8):1311–1331, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:2004:LEM

[1374] P. Bauer and M. Posch. Letter to the Editor: Modification of the sample size and the schedule of interim analyses in survival trials based on data inspections by H. Schäfer and H.-H. Müller, Statistics in Medicine 2001; 20: 3741–3751. Statistics in Medicine, 23(8):1333–1334, April 30, 2004.

CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [655].

Muller:2004:AR

[1375] H.-H. Müller and H. Schäfer. Authors’ reply. Statistics in Medicine, 23(8):1334–1335, April 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reuter:2004:ULM

[1376] Martin Reuter, Juergen Hennig, Petra Netter, Markus Buehner, and Michael Hueppe. Using latent mixed Markov models for the choice of the best pharmacological treatment. Statistics in Medicine, 23(9):1337–1349, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sweeting:2004:WAN

[1377] Michael J. Sweeting, Alexander J. Sutton, and Paul C. Lambert. What to add to nothing? Use and avoidance of continuity corrections in metaanalysis of sparse data. Statistics in Medicine, 23(9):1351–1375, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martus:2004:GAM

[1378] Peter Martus, Andrea Stroux, Anselm M. Jünemann, Matthias Korth, Jost B. Jonas, Folkert K. Horn, and Andreas Ziegler. GEE approaches to marginal regression models for medical diagnostic tests. Statistics in Medicine, 23(9):1377–1398, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nadziejko:2004:FLM

[1379] Christine Nadziejko, Lung Chi Chen, Arthur Nádas, and Jing-Shiang Hwang. The ‘fishing license’ method for analysing the time course of effects in repeated measurements. Statistics in Medicine, 23(9):1399– 1411, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schafer:2004:CGS

[1380] Helmut Schäfer and Hans-Helge Müller. Construction of group sequential designs in clinical trials on the basis of detectable treatment differences. Statistics in Medicine, 23(9):1413–1424, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moons:2004:TBL

[1381] E. Moons, M. Aerts, and G. Wets. A tree based lack-of-fit test for multiple logistic regression. Statistics in Medicine, 23(9):1425–1438, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2004:TSB

[1382] Nancy R. Cook, Robert Y. L. Zee, and Paul M. Ridker. Tree and spline based association analysis of gene-gene interaction models for ischemic stroke. Statistics in Medicine, 23(9):1439–1453, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hogan:2004:HDL

[1383] Joseph W. Hogan, Jason Roy, and Christina Korkontzelou. Handling drop-out in longitudinal studies. Statistics in Medicine, 23(9):1455–1497, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2004:BRS

[1384] Toshiro Tango. Book review: Scan Statistics. Joseph Glaz, Joseph Naus and Sylvan Wallenstein, Springer, New York, 2001. No. of pages: xv + 370. ISBN 0-387-098819-X. Statistics in Medicine, 23(9):1499, May 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2004:SDM

[1385] Ralph B. D’Agostino, Sr. The statistician and the data monitoring committee. Statistics in Medicine, 23(10):1501–1502, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2004:SSR

[1386] Susan S. Ellenberg and Stephen L. George. Should statisticians reporting to data monitoring committees be independent of the trial sponsor and leadership? Statistics in Medicine, 23(10):1503–1505, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bryant:2004:WAR

[1387] John Bryant. What is the appropriate role of the trial statistician in preparing and presenting interim findings to an independent Data Monitoring Committee in the U.S. Cancer Cooperative Group setting? Statistics in Medicine, 23(10):1507–1511, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeMets:2004:ISD

[1388] David L. DeMets and Thomas R. Fleming. The independent statistician for data monitoring committees. Statistics in Medicine, 23(10):1513– 1517, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2004:CID

[1389] John M. Lachin. Conflicts of interest in data monitoring of industry versus publicly financed clinical trials. Statistics in Medicine, 23(10): 1519–1521, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:2004:PSP

[1390] Janet Wittes. Playing safe and preserving integrity: making the FDA model work. Statistics in Medicine, 23(10):1523–1525, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siegel:2004:ISW

[1391] Jay P. Siegel, Robert T. O’Neill, Robert Temple, Gregory Campbell, and Mary A. Foulkes. Independence of the statistician who analyses unblinded data. Statistics in Medicine, 23(10):1527–1529, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snapinn:2004:RUS

[1392] Steven Snapinn, Thomas Cook, Deborah Shapiro, and Duane Snavely. The role of the unblinded sponsor statistician. Statistics in Medicine, 23(10):1531–1533, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:2004:MTN

[1393] Stuart J. Pocock. A major trial needs three statisticians: why, how and who? Statistics in Medicine, 23(10):1535–1539, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2004:TTE

[1394] Sudhir. R. Paul, Xing Jiang, Shesh Nath Rai, and Uditha Balasooriya. Test of treatment effect in pre-drug and post-drug count data with zeroinflation. Statistics in Medicine, 23(10):1541–1554, May 30, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moskowitz:2004:QCA

[1395] Chaya S. Moskowitz and Margaret S. Pepe. Quantifying and comparing the accuracy of binary biomarkers when predicting a failure time outcome. Statistics in Medicine, 23(10):1555–1570, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2004:DPC

[1396] Vance W. Berger. Does the Prentice criterion validate surrogate endpoints? Statistics in Medicine, 23(10):1571–1578, May 30, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittkowski:2004:CSO

[1397] Knut M. Wittkowski, Edmund Lee, Rachel Nussbaum, Francesca N. Chamian, and James G. Krueger. Combining several ordinal measures in clinical studies. Statistics in Medicine, 23(10):1579–1592, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grigg:2004:RAS

[1398] O. A. Grigg and V. T. Farewell. A risk-adjusted sets method for monitoring adverse medical outcomes. Statistics in Medicine, 23(10):1593–1602, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2004:MAR

[1399] Sarah A. White and Nynke R. van den Broek. Methods for assessing reliability and validity for a measurement tool: a case study and critique using the WHO haemoglobin colour scale. Statistics in Medicine, 23(10): 1603–1619, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2004:GLR

[1400] Qiang Zhao and Jianguo Sun. Generalized log-rank test for mixed interval-censored failure time data. Statistics in Medicine, 23(10):1621– 1629, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sullivan:2004:PMD

[1401] Lisa M. Sullivan, Joseph M. Massaro, and Ralph B. D’Agostino, Sr. Presentation of multivariate data for clinical use: the Framingham Study risk score functions. Statistics in Medicine, 23(10):1631–1660, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elbourne:2004:BRD

[1402] Diana Elbourne. Book review: Data Monitoring Committees In Clinical Trials: a Practical Perspective. Susan S. Ellenberg, Thomas R. Fleming and David L. DeMets, Wiley, Chichester, U.K., 2002. No. of pages: xiii + 191. Price: £45.00. ISBN 0-471-48986-7. Statistics in Medicine, 23(10): 1661–1662, May 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Higgins:2004:CRS

[1403] Julian P. T. Higgins and Simon G. Thompson. Controlling the risk of spurious findings from meta-regression. Statistics in Medicine, 23(11): 1663–1682, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deBoo:2004:MSS

[1404] Theo M. de Boo and Gerhard A. Zielhuis. Minimization of sample size when comparing two small probabilities in a non-inferiority safety trial. Statistics in Medicine, 23(11):1683–1699, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hollander:2004:CIE

[1405] Norbert Holländer, Willi Sauerbrei, and Martin Schumacher. Confidence intervals for the effect of a prognostic factor after selection of an ‘optimal’ cutpoint. Statistics in Medicine, 23(11):1701–1713, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2004:SIL

[1406] Lang Wu. Simultaneous inference for longitudinal data with detection limits and covariates measured with errors, with application to AIDS studies. Statistics in Medicine, 23(11):1715–1731, June 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2004:CFE

[1407] Binbing Yu, Ram C. Tiwari, Kathleen A. Cronin, and Eric J. Feuer. Cure fraction estimation from the mixture cure models for grouped survival data. Statistics in Medicine, 23(11):1733–1747, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2004:CGF

[1408] H. Pan and T. J. Cole. A comparison of goodness of fit tests for agerelated reference ranges. Statistics in Medicine, 23(11):1749–1765, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pawitan:2004:GEP

[1409] Y. Pawitan, J. Bjöhle, S. Wedren, K. Humphreys, L. Skoog, F. Huang, L. Amler, P. Shaw, P. Hall, and J. Bergh. Gene expression profiling for prognosis using Cox regression. Statistics in Medicine, 23(11):1767–1780, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vaeth:2004:SAP

[1410] Michael Væth and Eva Skovlund. A simple approach to power and sample size calculations in logistic regression and Cox regression models. Statistics in Medicine, 23(11):1781–1792, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halabi:2004:SSD

[1411] Susan Halabi and Bahadur Singh. Sample size determination for comparing several survival curves with unequal allocations. Statistics in Medicine, 23(11):1793–1815, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freels:2004:ESS

[1412] Sally Freels. Extracting summary statistics to perform meta-analysis of the published literature for survival endpoints. Statistics in Medicine, 23 (11):1817, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freidlin:2004:TPS

[1413] Boris Freidlin and Edward L. Korn. A testing procedure for survival data with few responders. Statistics in Medicine, 23(11):1818–1823, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henderson:2004:C

[1414] R. Henderson and H. Prince. Correction. Statistics in Medicine, 23 (11):1823, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobson:2004:BRO

[1415] Annette Dobson. Book review: 1. The Oxford Dictionary of Statistical Terms. Yadolah Dodge (ed.), Oxford University Press, Oxford, 2003,Hardcover. No. of pages: 506. Price: GBP 25.00. ISBN 0-19-850994-4. Statistics in Medicine, 23(11):1824–1825, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whittaker:2004:BRG

[1416] Joe Whittaker. Book review: 2. Graphical Models: Methods for Data Analysis and Mining. Christian Borgelt and Rudolf Kruse, Wiley, Chichester, 2002. No. of pages: 368. Price: $79.95. ISBN: 0-470-84337-3. Statistics in Medicine, 23(11):1825, June 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2004:APE

[1417] Weichung Joseph Shih, Pamela A. Ohman-Strickland, and Yong Lin. Analysis of pilot and early phase studies with small sample sizes. Statistics in Medicine, 23(12):1827–1842, June 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moser:2004:ORC

[1418] Barry Kurt Moser and Laura P. Coombs. Odds ratios for a continuous outcome variable without dichotomizing. Statistics in Medicine, 23(12): 1843–1860, June 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2004:NPA

[1419] Anastasia Ivanova and Kai Wang. A non-parametric approach to the design and analysis of two-dimensional dose-finding trials. Statistics in Medicine, 23(12):1861–1870, June 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:2004:SAM

[1420] Taesung Park, Moran Ki, and Sung-Gon Yi. Statistical analysis of MMR vaccine adverse events on aseptic meningitis using the case cross-over design. Statistics in Medicine, 23(12):1871–1883, June 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schaubel:2004:NPE

[1421] Douglas E. Schaubel and Jianwen Cai. Non-parametric estimation of gap time survival functions for ordered multivariate failure time data. Statistics in Medicine, 23(12):1885–1900, June 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2004:BAR

[1422] Yan Liu, Wesley O. Johnson, Ellen B. Gold, and Bill L. Lasley. Bayesian analysis of risk factors for anovulation. Statistics in Medicine, 23(12): 1901–1919, June 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julious:2004:SSC

[1423] Steven A. Julious. Sample sizes for clinical trials with normal data. Statistics in Medicine, 23(12):1921–1986, June 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armstrong:2004:LEA

[1424] P. W. Armstrong. Letter to the Editor: Armstrong, P. W. The ebb and flow of the NHS waiting list: how do recruitment and admission affect event-based measures of the length of ‘time-to-admission’? Statistics in Medicine 2002; 21:2991–3009. Statistics in Medicine, 23(12):1987, June 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [909].

Brunner:2004:BRA

[1425] Edgar Brunner. Book review: 1. Applied Nonparametric Statistical Methods. P. Sprent and N. C. Smeeton, Chapman and Hall/CRC, London, England, 2001. No. of pages: ix + 461. Price: £29.99. ISBN: 1-584-88145-3. Statistics in Medicine, 23(12):1988–1989, June 30, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yamaguchi:2004:ADPa

[1426] Takuhiro Yamaguchi and Yasuo Ohashi. Adjusting for differential proportions of second-line treatment in cancer clinical trials. Part I: Structural nested models and marginal structural models to test and estimate treatment arm effects. Statistics in Medicine, 23(13):1991–2003, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yamaguchi:2004:ADPb

[1427] Takuhiro Yamaguchi and Yasuo Ohashi. Adjusting for differential proportions of second-line treatment in cancer clinical trials. Part II: an application in a clinical trial of unresectable non-small-cell lung cancer. Statistics in Medicine, 23(13):2005–2022, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eberly:2004:CER

[1428] Lynn E. Eberly. Consequences of event rate heterogeneity across nonrandomized study sub-groups. Statistics in Medicine, 23(13):2023–2036, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wall:2004:ASR

[1429] Melanie M. Wall. Adjusting SIDS rates by seasonality in births in Minnesota. Statistics in Medicine, 23(13):2037–2048, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Du:2004:UBT

[1430] Joanna X. Du, Terry Watkins, Luis E. Bravo, Elizabeth T. H. Fontham, M. Constanza Camargo, Pelayo Correa, and Robertino Mera. ^{13}C-urea breath test for Helicobacter pylori in young children: cut-off point determination by finite mixture model. Statistics in Medicine, 23(13):2049– 2060, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Terpstra:2004:CBR

[1431] Jeff T. Terpstra and Lindsey A. Liudahl. Concomitant-based rank set sampling proportion estimates. Statistics in Medicine, 23(13):2061–2070, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goodman:2004:EEL

[1432] Allen C. Goodman, Yingwei Peng, Janet R. Hankin, David E. Kalist, and Stephen J. Spurr. Estimating episode lengths when some observations are probably censored. Statistics in Medicine, 23(13):2071–2087, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2004:TIC

[1433] Xuelin Huang, Robert A. Wolfe, and Chengcheng Hu. A test for informative censoring in clustered survival data. Statistics in Medicine, 23(13): 2089–2107, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2004:OMD

[1434] Michael J. Pencina and Ralph B. D’Agostino. Overall C as a measure of discrimination in survival analysis: model specific population value and confidence interval estimation. Statistics in Medicine, 23(13):2109–2123, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2004:NAI

[1435] Lili Tian and Joseph C. Cappelleri. A new approach for interval estimation and hypothesis testing of a certain intraclass correlation coefficient: the generalized variable method. Statistics in Medicine, 23(13): 2125–2135, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mugno:2004:AUD

[1436] Raymond Mugno, Wei Zhus, and William F. Rosenberger. Adaptive urn designs for estimating several percentiles of a dose-response curve. Statistics in Medicine, 23(13):2137–2150, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:2004:BRI

[1437] Kerrie Nelson. Book review: 1. An Introduction to Generalized Linear Models. 2nd edn. Annette J. Dobson, Chapman and Hall/CRC, London, 2002. No. of pages: ix + 225. Price: $49.95. ISBN 1-58488-165-8. Statistics in Medicine, 23(13):2151, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gebski:2004:BRS

[1438] Val Gebski. Book review: 2. Statistics in Clinical Oncology. John Crowley (ed.), Marcel Dekker, New York, 2001. Price: $175.00. ISBN: 0-8247-9025-1. Statistics in Medicine, 23(13):2151–2153, July 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Poon:2004:LND

[1439] Wai-Yin Poon. A latent normal distribution model for analysing ordinal responses with applications in meta-analysis. Statistics in Medicine, 23 (14):2155–2172, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tibaldi:2004:MAB

[1440] Fabián Tibaldi, Franz Torres Barbosa, and Geert Molenberghs. Modelling associations between time-to-event responses in pilot cancer clinical trials using a Plackett–Dale model. Statistics in Medicine, 23(14): 2173–2186, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kallen:2004:MER

[1441] Anders Källén. Multivariate estimation of the relative dose potency. Statistics in Medicine, 23(14):2187–2193, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rogerson:2004:MCS

[1442] Peter A. Rogerson and Ikuho Yamada. Monitoring change in spatial patterns of disease: comparing univariate and multivariate cumulative sum approaches. Statistics in Medicine, 23(14):2195–2214, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dufouil:2004:ALS

[1443] Carole Dufouil, Carol Brayne, and David Clayton. Analysis of longitudinal studies with death and drop-out: a case study. Statistics in Medicine, 23(14):2215–2226, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:2004:RBP

[1444] Richard B. Evans and Keith Erlandson. Robust Bayesian prediction of subject disease status and population prevalence using several similar diagnostic tests. Statistics in Medicine, 23(14):2227–2236, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2004:DTA

[1445] Chun-Lung Su, Ian A. Gardner, and Wesley O. Johnson. Diagnostic test accuracy and prevalence inferences based on joint and sequential testing with finite population sampling. Statistics in Medicine, 23(14): 2237–2255, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:2004:ATR

[1446] Nils P. Johnson. Advantages to transforming the receiver operating characteristic (ROC) curve into likelihood ratio co-ordinates. Statistics in Medicine, 23(14):2257–2266, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1682].

Zwane:2004:MRS

[1447] Eugene N. Zwane, Karin van der Pal-de Bruin, and Peter G. M. van der Heijden. The multiple-record systems estimator when registrations refer to different but overlapping populations. Statistics in Medicine, 23(14): 2267–2281, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gonen:2004:SSP

[1448] Mithat Gönen. Sample size and power for McNemar’s test with clustered data. Statistics in Medicine, 23(14):2283–2294, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yee:2004:QRV

[1449] Thomas W. Yee. Quantile regression via vector generalized additive models. Statistics in Medicine, 23(14):2295–2315, July 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittkowski:2004:LEE

[1450] Knut M. Wittkowski. Letter to the Editor: Effects and non-effects of paired identical observations in comparing proportions with binary matched-pairs data. By A. Agresti and Y. Min. Statistics in Medicine 2004; 23:65–75. Statistics in Medicine, 23(14):2317, July 30, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1275].

Levy:2004:CAC

[1451] Douglas E. Levy, A. James O’Malley, and Sharon-Lise T. Normand. Covariate adjustment in clinical trials with non-ignorable missing data and non-compliance. Statistics in Medicine, 23(15):2319–2339, August 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klar:2004:MMC

[1452] Neil Klar and Gerarda Darlington. Methods for modelling change in cluster randomization trials. Statistics in Medicine, 23(15):2341–2357, August 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cesana:2004:SST

[1453] Bruno Mario Cesana. Sample size for testing and estimating the difference between two paired and unpaired proportions: a ‘two-step’ procedure combining power and the probability of obtaining a precise estimate. Statistics in Medicine, 23(15):2359–2373, August 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

May:2004:DVP

[1454] Margaret May, Patrick Royston, Matthias Egger, Amy C. Justice, Jonathan Ac Sterne, and For the Art Cohort Collaboration. Development and validation of a prognostic model for survival time data: application to prognosis of HIV positive patients treated with antiretroviral therapy. Statistics in Medicine, 23(15):2375–2398, August 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sartori:2004:LIA

[1455] Nicola Sartori, Karl Thomaseth, and Alberto Salvan. Local influence analysis when interfacing toxicokinetic and proportional hazard models. Statistics in Medicine, 23(15):2399–2412, August 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:2004:IGT

[1456] Barry R. Davis, Charles E. Ford, Eric Boerwinkle, Donna Arnett, John Eckfeldt, and Henry Black. Imputing gene-treatment interactions when the genotype distribution is unknown using case-only and putative placebo analyses — a new method for the genetics of hypertension associated treatment (GenHAT) study. Statistics in Medicine, 23(15):2413– 2427, August 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitaker:2004:EID

[1457] H. J. Whitaker and C. P. Farrington. Estimation of infectious disease parameters from serological survey data: the impact of regular epidemics. Statistics in Medicine, 23(15):2429–2443, August 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tutz:2004:FMD

[1458] G. Tutz and H. Binder. Flexible modelling of discrete failure time including time-varying smooth effects. Statistics in Medicine, 23(15):2445– 2461, August 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Price:2004:MRS

[1459] Dionne L. Price and Amita K. Manatunga. Modelling relative survival using transformation methods. Statistics in Medicine, 23(15):2463–2477, August 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2004:BRG

[1460] Per Kragh Andersen. Book review: 1. Generalized estimating equations. James W. Hardin and Joseph M. Hilbe, Chapman and Hall/CRC, Boca Raton, 2003. No. of pages: xiii + 222 pp. Price: $79.95. ISBN 1-58488-307-3. Statistics in Medicine, 23(15):2479–2480, August 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Garrett-Mayer:2004:BRS

[1461] Elizabeth Garrett-Mayer. Book review: 2. Scaling procedures: Issues and applications. Richard G. Netemeyer, William O. Bearden and Subhash Sharma, Sage Publications, London, 2003. No. of pages: 206. Price: GBP 50.00 (cloth), GBP 25.00 (paper). ISBN 0-7619-2026-9. Statistics in Medicine, 23(15):2480–2481, August 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2004:QEF

[1462] Ranjan K. Paul, William F. Rosenberger, and Nancy Flournoy. Quantile estimation following non-parametric phase I clinical trials with ordinal response. Statistics in Medicine, 23(16):2483–2495, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muller:2004:GSP

[1463] Hans-Helge Müller and Helmut Schäfer. A general statistical principle for changing a design any time during the course of a trial. Statistics in Medicine, 23(16):2497–2508, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2004:NAM

[1464] Patrick Royston and Willi Sauerbrei. A new approach to modelling interactions between treatment and continuous covariates in clinical trials by using fractional polynomials. Statistics in Medicine, 23(16):2509–2525, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2004:MEM

[1465] Liang Li, Mari Palta, and Jun Shao. A measurement error model with a Poisson distributed surrogate. Statistics in Medicine, 23(16):2527–2536, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2004:SSS

[1466] Jialiang Li and Jason Fine. On sample size for sensitivity and specificity in prospective diagnostic accuracy studies. Statistics in Medicine, 23 (16):2537–2550, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2004:SSC

[1467] Dongwoo Kang, Janice B. Schwartz, and Davide Verotta. A sample size computation method for non-linear mixed effects models with applications to pharmacokinetics models. Statistics in Medicine, 23(16):2551– 2566, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steyerberg:2004:VUP

[1468] Ewout W. Steyerberg, Gerard J. J. M. Borsboom, Hans C. van Houwelingen, Marinus J. C. Eijkemans, and J. Dik F. Habbema. Validation and updating of predictive logistic regression models: a study on sample size and shrinkage. Statistics in Medicine, 23(16):2567–2586, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Redden:2004:SST

[1469] David T. Redden, José R. Fernández, and David B. Allison. A simple significance test for quantile regression. Statistics in Medicine, 23 (16):2587–2597, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rasmussen:2004:BRH

[1470] Søren Rasmussen. Book review: Hierarchical modelling of small area and hospital variation in short-term prognosis after acute myocardial infarction A longitudinal study of 35- to 74-year-old men in Denmark between 1978 and 1997. Statistics in Medicine, 23(16):2599–2621, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reilly:2004:BRS

[1471] Marie Reilly. Book review: 1. Statistical methods for the analysis of biomedical data. (2nd edn). Robert F. Woolson and William R. Clarke, Wiley, New York, 2002. No. of pages xxx. Price: £72.50. ISBN 0-471-39405-X. Statistics in Medicine, 23(16):2623–2624, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2004:C

[1472] Anonymous. Correction. Statistics in Medicine, 23(16):2624, August 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2004:QAC

[1473] Lang Li, Mehul Desai, Zeruesenay Desta, and David Flockhart. QT analysis: a complex answer to a ‘simple’ problem. Statistics in Medicine, 23(17):2625–2643, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

John:2004:CAC

[1474] J. A. John, K. G. Russell, and D. Whitaker. CrossOver: an algorithm for the construction of efficient cross-over designs. Statistics in Medicine, 23(17):2645–2658, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

George:2004:SAE

[1475] Stephen L. George, Boris Freidlin, and Edward L. Korn. Strength of accumulating evidence and data monitoring committee decision making. Statistics in Medicine, 23(17):2659–2672, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kurland:2004:MTM

[1476] Brenda F. Kurland and Patrick J. Heagerty. Marginalized transition models for longitudinal binary data with ignorable and non-ignorable drop-out. Statistics in Medicine, 23(17):2673–2695, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2004:EVE

[1477] Haitao Chu and M. Elizabeth Halloran. Estimating vaccine efficacy using auxiliary outcome data and a small validation sample. Statistics in Medicine, 23(17):2697–2711, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2004:SAM

[1478] Bin Huang, Siva Sivaganesan, Paul Succop, and Elizabeth Goodman. Statistical assessment of mediational effects for logistic mediational models. Statistics in Medicine, 23(17):2713–2728, September 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ng:2004:MDI

[1479] S. K. Ng, G. J. McLachlan, Kelvin K. W. Yau, and Andy H. Lee. Modelling the distribution of ischaemic stroke-specific survival time using an EM-based mixture approach with random effects adjustment. Statistics in Medicine, 23(17):2729–2744, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mann:2004:SWB

[1480] Vera Mann, Bianca L. De Stavola, and David A. Leon. Separating within and between effects in family studies: an application to the study of blood pressure in children. Statistics in Medicine, 23(17):2745–2756, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2004:SST

[1481] Andy H. Lee, Liming Xiang, and Wing K. Fung. Sensitivity of score tests for zero-inflation in count data. Statistics in Medicine, 23(17):2757–2769, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carroll:2004:DAN

[1482] Kevin Carroll, Bob Milsted, and John A. Lewis. Design and analysis of non-inferiority mortality trials in oncology. Statistics in Medicine, 23 (17):2771–2774, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rothmann:2004:AR

[1483] Mark Rothmann. Author’s reply. Statistics in Medicine, 23(17):2774– 2778, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Govindarajulu:2004:C

[1484] Z. Govindarajulu. Correction. Statistics in Medicine, 23(17):2779, September 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2004:TSA

[1485] Weichung J. Shih, Hui Quan, and Gang Li. Two-stage adaptive strategy for superiority and non-inferiority hypotheses in active controlled clinical trials. Statistics in Medicine, 23(18):2781–2798, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:2004:PAL

[1486] X. M. Tu, J. Kowalski, J. Zhang, K. G. Lynch, and P. Crits-Christoph. Power analyses for longitudinal trials and other clustered designs. Statistics in Medicine, 23(18):2799–2815, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2004:ICI

[1487] Myron N. Chang. Improved confidence intervals for a binomial parameter following a group sequential phase II clinical trial. Statistics in Medicine, 23(18):2817–2826, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2004:IIH

[1488] Juwon Song and Thomas R. Belin. Imputation for incomplete highdimensional multivariate normal data using a common factor model. Statistics in Medicine, 23(18):2827–2843, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ambrosius:2004:CCL

[1489] Walter T. Ambrosius and Siu L. Hui. Cross calibration in longitudinal studies. Statistics in Medicine, 23(18):2845–2861, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2004:CRA

[1490] Eric C. Yan and Ling Chen. A cost-related approach for evaluating drug development programs. Statistics in Medicine, 23(18):2863–2873, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Legler:2004:LVM

[1491] Julie M. Legler, William W. Davis, Arnold L. Potosky, and Richard M. Hoffman. Latent variable modelling of recovery trajectories: sexual function following radical prostatectomy. Statistics in Medicine, 23(18):2875– 2893, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Das:2004:ACC

[1492] Kalyan Das and Asis Kumar Chattopadhyay. An analysis of clustered categorical data-application in dental health. Statistics in Medicine, 23 (18):2895–2910, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horton:2004:RAM

[1493] Nicholas J. Horton and Garrett M. Fitzmaurice. Regression analysis of multiple source and multiple informant data from complex survey samples. Statistics in Medicine, 23(18):2911–2933, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ukoumunne:2004:CCU

[1494] Obioha C. Ukoumunne. Correction. O. C. Ukoumunne, A comparison of confidence interval methods for the intraclass correlation coefficient in cluster randomized trials Statistics in Medicine 2002; 21: 3757– 3774. Statistics in Medicine, 23(18):2935, September 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [964].

Lunceford:2004:SWP

[1495] Jared K. Lunceford and Marie Davidian. Stratification and weighting via the propensity score in estimation of causal treatment effects: a comparative study. Statistics in Medicine, 23(19):2937–2960, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:2004:EVE

[1496] Xiaohong M. Davis and Michael Haber. Estimating vaccine efficacy from household data observed over time. Statistics in Medicine, 36(14):2961– 2974, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2004:MCC

[1497] Siu Hung Cheung, Koon Shing Kwong, Wai Sum Chan, and Shun Piu Leung. Multiple comparisons with a control in families with both onesided and two-sided hypotheses. Statistics in Medicine, 36(14):2975– 2988, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bakker:2004:ICS

[1498] Bart Bakker, Tom Heskes, Jan Neijt, and Bert Kappen. Improving Cox survival analysis with a neural-Bayesian approach. Statistics in Medicine, 36(14):2989–3012, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reilly:2004:MID

[1499] Marie Reilly, Agus Salim, Emer Lawlor, Owen Smith, Ian Temperley, and Yudi Pawitan. Modelling infectious disease transmission with complex exposure pattern and sparse outcome data. Statistics in Medicine, 36 (14):3013–3032, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2004:SMH

[1500] Zuoheng Wang and Rongling Wu. A statistical model for high-resolution mapping of quantitative trait loci determining HIV dynamics. Statistics in Medicine, 36(14):3033–3051, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rigby:2004:SCC

[1501] Robert A. Rigby and D. Mikis Stasinopoulos. Smooth centile curves for skew and kurtotic data modelled using the Box–Cox power exponential distribution. Statistics in Medicine, 36(14):3053–3076, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mishra:2004:MIB

[1502] Gita D. Mishra and Annette J. Dobson. Multiple imputation for body mass index: lessons from the Australian longitudinal study on Women’s health. Statistics in Medicine, 36(14):3077–3087, October 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Welsh:2004:BRH

[1503] Alan H. Welsh. Book review: 1. Handbook of parametric and nonparametric statistical procedures (3rd edn). David J. Sheskin, Chapman and Hall/CRC, Boca Raton, 2003. No of pages 1232. Price: $139-95. ISBN: 1-58488-440-1. Statistics in Medicine, 36(14):3089–3090, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stoyan:2004:BRS

[1504] Dietrich Stoyan. Book review: 2. Statistical inference and simulation for spatial point processes. Jesper Møller and Rasmus Waagepetersen, Chapman and Hall/CRC, Boca Raton, 2003. No. of pages: 320. Price: £69.95. ISBN: 1-58488-265-4. Statistics in Medicine, 36(14):3090–3091, October 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fieuws:2004:JMM

[1505] Steffen Fieuws and Geert Verbeke. Joint modelling of multivariate longitudinal profiles: pitfalls of the random-effects approach. Statistics in Medicine, 23(20):3093–3104, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2004:CDI

[1506] G. Lu and A. E. Ades. Combination of direct and indirect evidence in mixed treatment comparisons. Statistics in Medicine, 23(20):3105–3124, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raab:2004:DFI

[1507] G. M. Raab, J. A. Davies, and A. B. Salter. Designing follow-up intervals. Statistics in Medicine, 23(20):3125–3137, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bottle:2004:CPT

[1508] Alex Bottle and Jon Wakefield. Controlling for provider of treatment in the modelling of respiratory disease risk near cokeworks. Statistics in Medicine, 23(20):3139–3158, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Langohr:2004:PSM

[1509] Klaus Langohr, Guadalupe Gómez, and Robert Muga. A parametric survival model with an interval-censored covariate. Statistics in Medicine, 23 (20):3159–3175, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2004:PAF

[1510] Philippe Lambert, Dave Collett, Alan Kimber, and Rachel Johnson. Parametric accelerated failure time models with random effects and an application to kidney transplant survival. Statistics in Medicine, 23 (20):3177–3192, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2004:ITY

[1511] Nigel Stallard and Lee Tombs. Including a time-of-year effect in the analysis of a matched case-control study. Statistics in Medicine, 23(20):3193– 3207, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1784].

Shults:2004:ADM

[1512] Justine Shults, Melicia C. Whitt, and Shiriki Kumanyika. Analysis of data with multiple sources of correlation in the framework of generalized estimating equations. Statistics in Medicine, 23(20):3209–3226, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lie:2004:DIS

[1513] Stein Atle Lie, Lars B. Engesæter, Leif I. Havelin, Håkon K. Gjessing, and Stein E. Vollset. Dependency issues in survival analyses of 55 782 primary hip replacements from 47 355 patients. Statistics in Medicine, 23 (20):3227–3240, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carpenter:2004:LOC

[1514] James Carpenter, Mike Kenward, Stephen Evans, and Ian White. Last observation carry-forward and last observation analysis. Statistics in Medicine, 23(20):3241–3242, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2004:AR

[1515] Jun Shao and Bob Zhong. Authors’ reply. Statistics in Medicine, 23 (20):3242–3244, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sabatti:2004:BRA

[1516] Chiara Sabatti. Book review: 1. The analysis of gene expression data: Methods and software. Giovanni Parmigiani, Elizabeth S. Garrett, Rafael A. Irizarry and Scott L. Zeger (eds), Springer, New York, 2003. No. of pages: 455. Price: EUR 89.95; £63.00; US $89.95. ISBN: 0-387-95577-1. Statistics in Medicine, 23(20):3245–3246, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spizzichino:2004:BRB

[1517] Fabio Spizzichino. Book review: 2. Bayesian survival analysis. Joseph G. Ibrahim, Ming-Hui Chen and Debajyoti Sinha, Springer, Heidelberg, Germany, 2001. No. of pages: xiv + 479. Price: DM 171.09. ISBN: 0-387-95277-2. Statistics in Medicine, 23(20):3246–3247, October 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dawson:2004:PFD

[1518] Ree Dawson and Philip W. Lavori. Placebo-free designs for evaluating new mental health treatments: the use of adaptive treatment strategies. Statistics in Medicine, 23(21):3249–3262, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Latouche:2004:SSF

[1519] Aurélien Latouche, Raphaël Porcher, and Sylvie Chevret. Sample size formula for proportional hazards modelling of competing risks. Statistics in Medicine, 23(21):3263–3274, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frost:2004:ARD

[1520] Chris Frost, Michael G. Kenward, and Nick C. Fox. The analysis of repeated ‘direct’ measures of change illustrated with an application in longitudinal imaging. Statistics in Medicine, 23(21):3275–3286, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kieser:2004:PSS

[1521] Meinhard Kieser, Joachim Röhmel, and Tim Friede. Power and sample size determination when assessing the clinical relevance of trial results by ‘responder analyses’. Statistics in Medicine, 23(21):3287–3305, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glynn:2004:MIE

[1522] Robert J. Glynn and Bernard Rosner. Multiple imputation to estimate the association between eyes in disease progression with interval-censored data. Statistics in Medicine, 23(21):3307–3318, November 15, 2004. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schisterman:2004:AGR

[1523] Enrique F. Schisterman, David Faraggi, and Benjamin Reiser. Adjusting the generalized ROC curve for covariates. Statistics in Medicine, 23(21): 3319–3331, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaiser:2004:CRM

[1524] J. C. Kaiser and W. F. Heidenreich. Comparing regression methods for the two-stage clonal expansion model of carcinogenesis. Statistics in Medicine, 23(21):3333–3350, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:2004:ALC

[1525] Stephen R. Cole, Rui Li, Kathryn Anastos, Roger Detels, Mary Young, Joan S. Chmiel, and Alvaro Muñoz. Accounting for leadtime in cohort studies: evaluating when to initiate HIV therapies. Statistics in Medicine, 23(21):3351–3363, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2004:BCI

[1526] Hongyu Jiang and Xiao-Hua Zhou. Bootstrap confidence intervals for medical costs with censored observations. Statistics in Medicine, 23(21): 3365–3376, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Topp:2004:RAL

[1527] Rebekka Topp and Guadalupe Gómez. Residual analysis in linear regression models with an interval-censored covariate. Statistics in Medicine, 23(21):3377–3391, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rohmel:2004:MSS

[1528] Joachim Röhmel and Gernot Wassmer. Minimization of sample size when comparing two small probabilities in a non-inferiority safety trial. Statistics in Medicine, 23(21):3393–3394, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boo:2004:AR

[1529] Theo De Boo and Gerhard Zielhuis. Authors’ reply. Statistics in Medicine, 23(21):3395–3396, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:2004:BRS

[1530] Odd O. Aalen. Book review: 1. The statistical analysis of failure time data (2nd edn). J. D. Kalbfleisch and R. L. Prentice, Wiley-Interscience, Hoboken, New Jersey, 2002. No. of pages: 439. Price: £62.95. ISBN: 0-471-36357-X. Statistics in Medicine, 23(21):3397–3398, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Love:2004:BRC

[1531] Sharon Love. Book review: 2. Clinical trials in cancer: Principles and practice. David J. Girling, Mahesh K. B. Parmer, Sally P. Stenning, Richard J. Stephens and Lesley A. Stewart (eds), Oxford University Press, Oxford, 2003. No. of pages: 386. Price: £49.95. ISBN: 0-19-262959-X. Statistics in Medicine, 23(21):3398–3399, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmgren:2004:BRA

[1532] Juni Palmgren. Book review: 3. Analysis of longitudinal data (2nd edn). P. J. Diggle, P. Heagarty, K.-Y. Liang and S. L. Zeger, Oxford University Press, Oxford, 2002. No. of pages: xi + 379. Statistics in Medicine, 23 (21):3399–3401, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joseph:2004:BRB

[1533] Lawrence Joseph. Book review: 4. Bayesian data analysis (2nd edn). Andrew Gelman, John B. Carlin, Hal S. Stern and Donald B. Rubin (eds), Chapman and Hall/CRC, Boca Raton, 2003. No. of pages: xxv + 668. Price: $59.95. ISBN 1-58488-388-X. Statistics in Medicine, 23(21):

3401–3403, November 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2004:AL

[1534] Simon Thompson. The Armitage Lectures. Statistics in Medicine, 23 (22):3405–3407, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prentice:2004:CDP

[1535] Ross L. Prentice. Chronic disease prevention: public health potential and research needs. Statistics in Medicine, 23(22):3409–3420, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2004:UAU

[1536] Nancy R. Cook, Bernard A. Rosner, Wei Chen, Sathanur R. Srinivasan, and Gerald S. Berenson. Using the area under the curve to reduce measurement error in predicting young adult blood pressure from childhood measures. Statistics in Medicine, 23(22):3421–3435, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nakas:2004:OMC

[1537] Christos T. Nakas and Constantin T. Yiannoutsos. Ordered multipleclass ROC analysis with continuous measurements. Statistics in Medicine, 23(22):3437–3449, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2004:CBM

[1538] Duolao Wang, Wenyang Zhang, and Ameet Bakhai. Comparison of Bayesian model averaging and stepwise methods for model selection in logistic regression. Statistics in Medicine, 23(22):3451–3467, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cauchemez:2004:BMA

[1539] S. Cauchemez, F. Carrat, C. Viboud, A. J. Valleron, and P. Y. Boëlle. A Bayesian MCMC approach to study transmission of influenza: application to household longitudinal data. Statistics in Medicine, 23(22): 3469–3487, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2004:CMM

[1540] Jingjin Li and Mark D. Schluchter. Conditional mixed models adjusting for non-ignorable drop-out with administrative censoring in longitudinal studies. Statistics in Medicine, 23(22):3489–3503, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnett:2004:ETS

[1541] A. G. Barnett, A. J. Dobson, and WHO MONICA (monitoring trends and determinants in cardiovascular disease) Project. Estimating trends and seasonality in coronary heart disease. Statistics in Medicine, 23(22): 3505–3523, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2004:NPM

[1542] Quanxi Shao and Xian Zhou. A new parametric model for survival data with long-term survivors. Statistics in Medicine, 23(22):3525–3543, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rudas:2004:ROR

[1543] Tamás Rudas and Wicher Bergsma. Reconsidering the odds ratio as a measure of 2 × 2 association in a population. Statistics in Medicine, 23 (22):3545–3547, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:2004:RTR

[1544] Helena Chmura Kraemer. Reply to Tamas Rudas and Wicher Bergsma regarding ‘Reconsidering the odds ratio as a measure of 2 × 2 association in a population’. Statistics in Medicine, 23(22):3547–3548, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2004:LEC

[1545] Anonymous. Letter to the Editor: Confidence limits for the ratio of two rates based on likelihood scores: non-iterative method by P. L. Graham, K. Mengersen and A. P. Morton, Statistics in Medicine 2003; 22:2071– 2083. Statistics in Medicine, 23(22):3549, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1116, 1117, 1669].

Proschan:2004:BRM

[1546] Michael Proschan. Book review: 1. Multiple analyses in clinical trials: Fundamentals for investigators. Lemuel A. Moyé, Springer, New York, 2003. No. of pages: xxiii + 436. Price: $79.95, £69.00. EUR 89.95. ISBN: 0-387-00727. Statistics in Medicine, 23(22):3551–3552, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skrondal:2004:BRT

[1547] Anders Skrondal and Sophia Rabe-Hesketh. Book review: 2. Topics in modelling of clustered data. M. Aerts, H. Geys, G. Molenberghs and L. M. Ryan (eds), Chapman and Hall, London, 2002. Price: £48.99. ISBN: 1-58488-1852. Statistics in Medicine, 23(22):3552–3553, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gurrin:2004:BRM

[1548] Lyle C. Gurrin and John L. Hopper. Book review: 3. Mathematical and statistical methods for genetic analysis (2nd edn). Kenneth Lange, Springer, Berlin, 2002. No. of pages: 361. Price: £52.50. ISBN: 0-387-95389-2. Statistics in Medicine, 23(22):3554–3556, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wall:2004:BRI

[1549] Melanie M. Wall. Book review: 4. Introduction to applied statistics: a modelling approach (2nd edn). J. K. Lindsey, Oxford University Press, Oxford, 2004. No.of pages: xiv + 321. Price: £60.00 (Hardback), £27.99 (Paperback). ISBN 0-19-852894-9 (Hbk), 0-19-852895-7 (Pbk). Statistics in Medicine, 23(22):3556–3557, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edler:2004:BRB

[1550] Lutz Edler. Book review: 5. Biostatistical methods. Stephen W. Looney (ed.), Human Press, Totowa, New Jersey, U.S.A., 2002. No. of pages:

214. Price: $ 89.00. ISBN 0-89603-951-X. Statistics in Medicine, 23(22): 3557–3559, November 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kenna:2004:ETE

[1551] Leslie A. Kenna and Lewis B. Sheiner. Estimating treatment effect in the presence of non-compliance measured with error: precision and robustness of data analysis methods. Statistics in Medicine, 23(23):3561–3580, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2004:STD

[1552] Xiaohui Luo, Dennis D. Boos, and Roy N. Tamura. Score tests for dose effect in the presence of non-responders. Statistics in Medicine, 23(23): 3581–3591, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2004:SAS

[1553] Man-Lai Tang. On simultaneous assessment of sensitivity and specificity when combining two diagnostic tests. Statistics in Medicine, 23(23): 3593–3605, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bellamy:2004:ACI

[1554] Scarlett L. Bellamy, Yi Li, Louise M. Ryan, Stuart Lipsitz, Marina J. Canner, and Rosalind Wright. Analysis of clustered and interval censored data from a community-based study in asthma. Statistics in Medicine, 23(23):3607–3621, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wade:2004:LBM

[1555] A. M. Wade, A. T. Salt, R. V. Proffitt, S. J. Heavens, and P. M. Sonksen. Likelihood-based modelling of age-related normal ranges for ordinal measurements: changes in visual acuity through early childhood. Statistics in Medicine, 23(23):3623–3640, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gasser:2004:MNR

[1556] Theo Gasser and Valentin Rousson. Modelling neuromotor ratings with floor-effects. Statistics in Medicine, 23(23):3641–3653, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulinskaya:2004:WTT

[1557] E. Kulinskaya, M. B. Dollinger, E. Knight, and H. Gao. A welch-type test for homogeneity of contrasts under heteroscedasticity with application to meta-analysis. Statistics in Medicine, 23(23):3655–3670, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2004:BCT

[1558] Cong Han and Richard Kronmal. Box–Cox transformation of leftcensored data with application to the analysis of coronary artery calcification and pharmacokinetic data. Statistics in Medicine, 23(23):3671– 3679, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2004:ELR

[1559] Minhee Kang and Stephen W. Lagakos. Evaluation of log-rank tests for infrequent observations from a multi-state process, with application to HPV vaccine efficacy. Statistics in Medicine, 23(23):3681–3696, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraft:2004:CSG

[1560] Peter Kraft and Duncan C. Thomas. Case-sibling gene-association studies for diseases with variable age at onset. Statistics in Medicine, 23(23): 3697–3712, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2004:CES

[1561] Ling Chen and Yi Tsong. Comment on: Estimate of standard deviation for a log-transformed variable using arithmetic means and standard deviations by Quan Hui and Zhang Ji, Statistics in Medicine 2003; 22: 2723–2736. Statistics in Medicine, 23(23):3713–3716, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1174].

Quan:2004:AR

[1562] Hui Quan and Ji Zhang. Authors’ reply. Statistics in Medicine, 23(23): 3716–3718, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McEntegart:2004:LEC

[1563] Damian J. McEntegart. Letter to the Editor: Comparison of stratification and adaptive methods for treatment allocation in an acute stroke clinical trial by Christopher J. Weir and Kennedy R. Lees, Statistics in Medine 2003; 22:705–726. Statistics in Medicine, 23(23):3719–3720, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1022].

Weir:2004:AR

[1564] Christopher J. Weir and Kennedy R. Lees. Authors’ reply. Statistics in Medicine, 23(23):3720–3721, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:2004:BRA

[1565] Peter Armitage. Book review: 1. Annotated readings in the history of statistics. H. A. David and A. W. F. Edwards, Springer, New York, 2001. No. of pages: xv + 252. Price £53.50. ISBN 0-387-98844-0. Statistics in Medicine, 23(23):3722, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ecochard:2004:BRB

[1566] René Ecochard. Book review: 2. Biostatistical methods in epidemiology. Stephen C. Newman, Wiley, New York, 2001. No. of pages: 382. Price: £60.95. ISBN 0-471-36914-4. Statistics in Medicine, 23(23):3723, December 15, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:2004:P

[1567] Deborah Ashby. Preface. Statistics in Medicine, 23(24):3725, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2004:ISC

[1568] Anonymous. International Society for Clinical Biostatistics. Statistics in Medicine, 23(24):3727–3728, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2004:CCR

[1569] Stephen Senn. Controversies concerning randomization and additivity in clinical trials. Statistics in Medicine, 23(24):3729–3753, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanderTweel:2004:STG

[1570] Ingeborg van der Tweel and Maria Schipper. Sequential tests for geneenvironment interactions in matched case-control studies. Statistics in Medicine, 23(24):3755–3771, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Garcia:2004:ECD

[1571] Raquel Garcia, Marta Benet, Catalina Arnau, and Erik Cobo. Efficiency of the cross-over design: an empirical estimation. Statistics in Medicine, 23(24):3773–3780, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benedetti:2004:UGA

[1572] Andrea Benedetti and Michal Abrahamowicz. Using generalized additive models to reduce residual confounding. Statistics in Medicine, 23(24): 3781–3801, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thiebaut:2004:CTS

[1573] Anne C. M. Thiébaut and Jacques Bénichou. Choice of time-scale in Cox’s model analysis of epidemiologic cohort data: a simulation study. Statistics in Medicine, 23(24):3803–3820, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2004:APE

[1574] Paul C. Lambert, Paul R. Burton, Keith R. Abrams, and Adrian M. Brooke. The analysis of peak expiratory flow data using a three-level hierarchical model. Statistics in Medicine, 23(24):3821–3839, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Law:2004:ADP

[1575] Graham R. Law, Sara Rollinson, Richard Feltbower, James M. Allan, Gareth J. Morgan, and Eve Roman. Application of DNA pooling to large studies of disease. Statistics in Medicine, 23(24):3841–3850, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katsahian:2004:GVL

[1576] Sandrine Katsahian, Raphaël Porcher, Jean-Yves Mary, and Sylvie Chevret. The graft-versus-leukaemia effect after allogeneic bone-marrow transplantation: assessment through competing risks approaches. Statistics in Medicine, 23(24):3851–3863, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2004:A

[1577] Anonymous. Acknowledgement. Statistics in Medicine, 23(24):3865– 3868, December 30, 2004. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brittain:2005:CIT

[1578] Erica Brittain and Daphne Lin. A comparison of intent-to-treat and perprotocol results in antibiotic non-inferiority trials. Statistics in Medicine, 24(1):1–10, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McMahon:2005:MPT

[1579] Robert P. McMahon, Stephan Arndt, and Robert R. Conley. More powerful two-sample tests for differences in repeated measures of adverse effects in psychiatric trials when only some patients may be at risk. Statistics in Medicine, 24(1):11–21, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2005:MTP

[1580] Yunchan Chi. Multiple testing procedures based on weighted Kaplan– Meier statistics for right-censored survival data. Statistics in Medicine, 24(1):23–35, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2005:LCB

[1581] Aiyi Liu, Enrique F. Schisterman, and Yan Zhu. On linear combinations of biomarkers to improve diagnostic accuracy. Statistics in Medicine, 24(1):37–47, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsiao:2005:BTE

[1582] Chuhsing Kate Hsiao, Mei hsien Lee, and Robert E. Kass. Bayesian tests of extra-binomial variability. Statistics in Medicine, 24(1):49–64, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thiebaut:2005:JMB

[1583] Rodolphe Thiébaut, Hélène Jacqmin-Gadda, Abdel Babiker, Daniel Commenges, and The Cascade Collaboration. Joint modelling of bivariate longitudinal data with informative dropout and left-censoring, with application to the evolution of CD4+ cell count and HIV RNA viral load in response to treatment of HIV infection. Statistics in Medicine, 24(1):

65–82, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bacchetti:2005:UDM

[1584] Peter Bacchetti. Uncertainty due to model choice in variant Creutzfeldt– Jakob disease projections. Statistics in Medicine, 24(1):83–93, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kapitula:2005:DEN

[1585] Laura Ring Kapitula and Edward J. Bedrick. Diagnostics for the exponential normal growth curve model. Statistics in Medicine, 24(1):95–108, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2005:RMM

[1586] Ming Tan, Hong-Bin Fang, Guo-Liang Tian, and Peter J. Houghton. Repeated-measures models with constrained parameters for incomplete data in tumour xenograft experiments. Statistics in Medicine, 24(1): 109–119, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:2005:TGF

[1587] Nico Nagelkerke, Jeroen Smits, Saskia le Cessie, and Hans van Houwelingen. Testing goodness-of-fit of the logistic regression model in casecontrol studies using sample reweighting. Statistics in Medicine, 24(1): 121–130, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willan:2005:RMC

[1588] Andrew R. Willan, D. Y. Lin, and Andrea Manca. Regression methods for cost-effectiveness analysis with censored data. Statistics in Medicine, 24(1):131–145, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2005:LEI

[1589] A. Lawrence Gould and Weichung J. Shih. Letter to the Editor: On the inappropriateness of an EM algorithm based procedure for blinded sample size re-estimation by T. Friede and M. Kieser, Statistics in Medicine 2002; 21:165–176. Statistics in Medicine, 24(1):147–154, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [344, 3387, 682].

Friede:2005:AR

[1590] Tim Friede and Meinhard Kieser. Authors’ reply. Statistics in Medicine, 24(1):154–156, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2005:LET

[1591] Anonymous. Letter to the Editor: Tyrer J., Duffy, S. W. and Cuzick, J. A breast cancer prediction model incorporating familial and personal risk factors. Statistics in Medicine 2004; 23:1111–1130. Statistics in Medicine, 24(1):156, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1358].

Azzalini:2005:BRA

[1592] Adelchi Azzalini. Book review: 1. In all likelihood. Yudi Pawitan, Oxford University Press, Oxford, 2001. No. of pages: xiii + 515. Price: £40.00. ISBN: 0-19-850765-8. Statistics in Medicine, 24(1):157–158, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klein:2005:BRC

[1593] John P. Klein. Book review: 2. Classical competing risks. Martin J. Crowder, CRC Press, London, 2001. No. of pages: 200. Price: £49.99. ISBN 1-58488-175-5. Statistics in Medicine, 24(1):158–159, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altham:2005:BRC

[1594] Pat Altham. Book review: 3. Computer-aided multivariate analysis (4th edn). Abdelmonem Afifi, Virginia A. Clark and Susanne May, Chapman and Hall/CRC, Boca Raton, 2004. No. of pages: xvi + 489. Price: £39.99, $79.95. ISBN: 1-58488-308-1. Statistics in Medicine, 24(1):159–160, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lumley:2005:BRA

[1595] Thomas Lumley. Book review: 4. Applied longitudinal data analysis: Modelling change and event occurrence. Judith D. Singer and John B. Willett, Oxford University Press, New York, 2003. No. of pages: xx + 644. Price: £39.50. ISBN: 0-19-515296-4. Statistics in Medicine, 24(1): 160–161, January 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:2005:AST

[1596] Edward L. Korn, Paul S. Albert, and Lisa M. McShane. Assessing surrogates as trial endpoints using mixed models. Statistics in Medicine, 24(2):163–182, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:2005:CAS

[1597] Laurence Freedman. Commentary on Assessing surrogates as trial endpoints using mixed models, by E. L. Korn, P. S. Albert and L. M. Mc-Shane. Statistics in Medicine, 24(2):183–185, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:2005:RCD

[1598] Edward L. Korn, Paul S. Albert, and Lisa M. Mcshane. Rejoinder to commentary by Dr Freedman of ‘Assessing surrogates as trial endpoints using mixed models’. Statistics in Medicine, 24(2):187–190, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vansteelandt:2005:SSW

[1599] S. Vansteelandt and E. Goetghebeur. Sense and sensitivity when correcting for observed exposures in randomized clinical trials. Statistics in Medicine, 24(2):191–210, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2005:TCP

[1600] Y. H. Joshua Chen, A. Lawrence Gould, and Michael L. Nessly. Treatment comparisons for a partially categorical outcome applied to a biomarker with assay limit. Statistics in Medicine, 24(2):211–228, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arbogast:2005:MCT

[1601] Patrick G. Arbogast and D. Y. Lin. Model-checking techniques for stratified case-control studies. Statistics in Medicine, 24(2):229–247, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2005:BIP

[1602] Xiaoyin Wang, Chong Z. He, and Dongchu Sun. Bayesian inference on the patient population size given list mismatches. Statistics in Medicine, 24(2):249–267, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2005:MEM

[1603] Surupa Roy, T. Banerjee, and Tapabrata Maiti. Measurement error model for misclassified binary responses. Statistics in Medicine, 24(2): 269–283, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Janes:2005:OBC

[1604] Holly Janes, Lianne Sheppard, and Thomas Lumley. Overlap bias in the case-crossover design, with application to air pollution exposures. Statistics in Medicine, 24(2):285–300, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Messinger:2005:NPP

[1605] Shari Messinger and Susan Murray. Non-parametric paired two-sample tests for censored survival data incorporating longitudinal covariate information. Statistics in Medicine, 24(2):301–318, January 30, 2005. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kieser:2005:LE

[1606] M. Kieser. Letter to the Editor. Statistics in Medicine, 24(2):319–320, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cleophas:2005:AR

[1607] Tom. J. Cleophas. Author’s reply. Statistics in Medicine, 24(2):321– 323, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carlin:2005:BRB

[1608] John Carlin. Book review: 1. Bayesian Statistics 7: Proceedings of the Seventh Valencia International Meeting. J. M. Bernardo, M. J. Bayarri, J. O. Berger, A. P. Dawid, D. Heckerman, A. F. M. Smith and M. West, Oxford University Press, Oxford, 2003. No. of pages: xiii + 750. Price: £85.00. ISBN 0-19-852615-6. Statistics in Medicine, 24(2):325–326, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2005:BRM

[1609] Roger Marshall. Book review: 2. Measurement error and misclassification in statistics and epidemiology: Impacts and Bayesian adjustment. Paul Gustafson (ed.), Chapman and Hall/CRC, Boca Raton, 2003. No. of pages: xi + 188. Price: $79.95. ISBN: 1-58488-335-9. Statistics in Medicine, 24(2):326–327, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stare:2005:BRS

[1610] Janez Stare. Book review: 3. Survival analysis using S: Analysis of time to event data. Mara Tableman and Jong Sung Kim with a contribution from Stephen Portnoy, Chapman and Hall/CRC, Boca Raton, 2004. No. of pages: xv + 280. Price: $69.95. ISBN 1-58488-408-8. Statistics in Medicine, 24(2):327–328, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maitournam:2005:ETC

[1611] A. Maitournam and R. Simon. On the efficiency of targeted clinical trials. Statistics in Medicine, 24(3):329–339, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OBrien:2005:SPN

[1612] Peter C. O’Brien, David Zhang, and Kent R. Bailey. Semi-parametric and non-parametric methods for clinical trials with incomplete data. Statistics in Medicine, 24(3):341–358, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1851].

Bebchuk:2005:TTG

[1613] Judith D. Bebchuk and Rebecca A. Betensky. Tests for treatment group differences in the hazards for survival, before and after the occurrence of an intermediate event. Statistics in Medicine, 24(3):359–378, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:2005:BAM

[1614] Gordon J. Prescott and Paul H. Garthwaite. Bayesian analysis of misclassified binary data from a matched case-control study with a validation sub-study. Statistics in Medicine, 24(3):379–401, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonzo:2005:VBC

[1615] Todd A. Alonzo. Verification bias-corrected estimators of the relative true and false positive rates of two binary screening tests. Statistics in Medicine, 24(3):403–417, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2005:EIE

[1616] Hung-Mo Lin, Robert H. Lyles, John M. Williamson, and Allen R. Kunselman. Estimation of the intervention effect in a non-randomized study with pre- and post-mismeasured binary responses. Statistics in Medicine, 24(3):419–435, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandrekar:2005:SMD

[1617] Sumithra J. Mandrekar, Haikady N. Nagaraja, and Gary G. Berntson. Statistical modelling of the differences between successive R–R intervals. Statistics in Medicine, 24(3):437–451, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2005:UAI

[1618] Ruiguang Song, H. Irene Hall, and Robert Frey. Uncertainties associated with incidence estimates of HIV/AIDS diagnoses adjusted for reporting delay and risk redistribution. Statistics in Medicine, 24(3):453–464, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2005:ICI

[1619] Xiao-Hua Zhou and Gengsheng Qin. Improved confidence intervals for the sensitivity at a fixed level of specificity of a continuous-scale diagnostic test. Statistics in Medicine, 24(3):465–477, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OQuigley:2005:ERP

[1620] John O’Quigley, Ronghui Xu, and Janez Stare. Explained randomness in proportional hazards models. Statistics in Medicine, 24(3):479–489, February 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipman:2005:P

[1621] Harvey B. Lipman. Preface. Statistics in Medicine, 24(4):491–492, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morton:2005:SCI

[1622] Sally C. Morton. Statistical collaboration to impact policy decisions. Statistics in Medicine, 24(4):493–501, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stangl:2005:BGB

[1623] Dalene K. Stangl. Bridging the gap between statistical analysis and decision making in public health research. Statistics in Medicine, 24(4): 503–511, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fienberg:2005:SIC

[1624] Stephen E. Fienberg and Galit Shmueli. Statistical issues and challenges associated with rapid detection of bio-terrorist attacks. Statistics in Medicine, 24(4):513–529, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:2005:MIP

[1625] Ron Brookmeyer, Elizabeth Johnson, and Sarah Barry. Modelling the incubation period of anthrax. Statistics in Medicine, 24(4):531–542, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutwagner:2005:SMA

[1626] L. C. Hutwagner, W. W. Thompson, G. M. Seeman, and T. Treadwell. A simulation model for assessing aberration detection methods used in public health surveillance for systems with limited baselines. Statistics in Medicine, 24(4):543–550, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rolka:2005:USA

[1627] Henry Rolka, Dwayne Bracy, Channing Russell, David Fram, and Robert Ball. Using simulation to assess the sensitivity and specificity of a signal detection tool for multidimensional public health surveillance data. Statistics in Medicine, 24(4):551–562, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kadane:2005:BMH

[1628] Joseph B. Kadane. Bayesian methods for health-related decision making. Statistics in Medicine, 24(4):563–567, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kafadar:2005:CMM

[1629] Karen Kafadar and Philip C. Prorok. Computational methods in medical decision making: to screen or not to screen? Statistics in Medicine, 24 (4):569–581, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Casella:2005:IMA

[1630] George Casella and Elías Moreno. Intrinsic meta-analysis of contingency tables. Statistics in Medicine, 24(4):583–604, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barker:2005:RSI

[1631] Lawrence E. Barker, Philip J. Smith, Robert B. Gerzoff, Elizabeth T. Luman, Mary M. McCauley, and Tara W. Strine. Ranking states’ immunization coverage: an example from the national immunization survey. Statistics in Medicine, 24(4):605–613, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caudill:2005:ECP

[1632] Samuel P. Caudill, Daniel L. Van Dyke, Andrew T. L. Chen, John A. Reidy, Paul S. Ing, Stuart Schwartz, and Gail H. Vance. Evaluating current policy for detecting mosaicism in amniotic fluid cultures: implications for current cell counting practices. Statistics in Medicine, 24 (4):615–622, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borkowf:2005:EER

[1633] Craig B. Borkowf and Paul S. Albert. Efficient estimation of the risk of a disease by quantile-categories of a key predictor variable using generalized additive models. Statistics in Medicine, 24(4):623–645, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:2005:MEQ

[1634] Andrew C. Leon and Donald Hedeker. A mixed-effects quintile-stratified propensity adjustment for effectiveness analyses of ordered categorical doses. Statistics in Medicine, 24(4):647–658, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gonzalez:2005:STD

[1635] Joe Fred Gonzalez, Jr. and Lawrence H. Cox. Software for tabular data protection. Statistics in Medicine, 24(4):659–669, February 28, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2005:SMM

[1636] Bingshu E. Chen, Richard J. Cook, Jerald F. Lawless, and Min Zhan. Statistical methods for multivariate interval-censored recurrent events. Statistics in Medicine, 24(5):671–691, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Potter:2005:PTI

[1637] Douglas M. Potter. A permutation test for inference in logistic regression with small- and moderate-sized data sets. Statistics in Medicine, 24(5): 693–708, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1951].

Adebayo:2005:ACM

[1638] Samson B. Adebayo and Ludwig Fahrmeir. Analysing child mortality in Nigeria with geoadditive discrete-time survival models. Statistics in Medicine, 24(5):709–728, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agresti:2005:SIC

[1639] Alan Agresti and Yongyi Min. Simple improved confidence intervals for comparing matched proportions. Statistics in Medicine, 24(5):729–740, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2032].

Kang:2005:HVC

[1640] Weechang Kang, Moo-Song Lee, and Youngjo Lee. HGLM versus conditional estimators for the analysis of clustered binary data. Statistics in Medicine, 24(5):741–752, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bonetti:2005:IDD

[1641] Marco Bonetti and Marcello Pagano. The interpoint distance distribution as a descriptor of point patterns, with an application to spatial disease clustering. Statistics in Medicine, 24(5):753–773, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foulkes:2005:MMC

[1642] A. S. Foulkes, M. Reilly, L. Zhou, M. Wolfe, and D. J. Rader. Mixed modelling to characterize genotype-phenotype associations. Statistics in Medicine, 24(5):775–789, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2005:UQR

[1643] Peter C. Austin, Jack V. Tu, Paul A. Daly, and David A. Alter. The use of quantile regression in health care research: a case study examining gender differences in the timeliness of thrombolytic therapy. Statistics in Medicine, 24(5):791–816, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Singer:2005:LEP

[1644] Julia Singer. Letter to the Editor: A practical approach for comparing means of two groups without equal variance assumption H. Wang and S. Chow, Statistics in Medicine 2002; 21:3137–3151. Statistics in Medicine, 24(5):817–818, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [919, 19].

Ganju:2005:LEA

[1645] Jitendra Ganju. Letter to the Editor: Analysis of clinical data with breached blindness by Shein-Chung Chow and Jun Shao Statistics in Medicine 2004; 23:1185–1193. Statistics in Medicine, 24(5):819–821, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1365, 2002].

Shao:2005:AR

[1646] Jun Shao and Shein-Chung Chow. Authors’ reply. Statistics in Medicine, 24(5):821–822, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2005:BRS

[1647] Niels Trolle Andersen. Book review: 1. Statistical methods for the analysis of repeated measurements. Charles S. Davis, Springer-Verlag, New York, 2002. No of pages: xxiv + 415. Price: $79.95. ISBN 0-387-95370-1. Statistics in Medicine, 24(5):823–824, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rudnicka:2005:BRE

[1648] Alicja R. Rudnicka. Book review: 2. Essential medical statistics (2nd edn). Betty R. Kirkwood and Jonathan A. C. Sterne, Blackwell Science, Oxford, 2003. No. of pages: 512. Price: £22.95. ISBN 0-86542-871-9. Statistics in Medicine, 24(5):824, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2005:LEA

[1649] Anonymous. Letter to the Editor: Adjusting for publication bias in the presence of heterogeneity by N. Terrin, C. H. Schmid, J. Lau, Statistics in Medicine 2003; 22:2113–2126. Statistics in Medicine, 24(5):825–826, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1122].

Baker:2005:C

[1650] S. G. Baker. Correction. Statistics in Medicine, 24(5):826, March 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borkowf:2005:SHV

[1651] Craig B. Borkowf. A simple hybrid variance estimator for the Kaplan– Meier survival function. Statistics in Medicine, 24(6):827–851, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baksh:2005:DCS

[1652] M. Fazil Baksh, Susan Todd, John Whitehead, and María M. Lucini. Design considerations in the sequential analysis of matched case-control data. Statistics in Medicine, 24(6):853–867, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2005:IDE

[1653] Gang Zheng and Xin Tian. The impact of diagnostic error on testing genetic association in case-control studies. Statistics in Medicine, 24(6): 869–882, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liao:2005:CCC

[1654] Jason J. Z. Liao. Comparing the concentration curves directly in a pharmacokinetics, bioavailability/bioequivalence study. Statistics in Medicine, 24(6):883–891, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demidenko:2005:IAL

[1655] Eugene Demidenko and Therese A. Stukel. Influence analysis for linear mixed-effects models. Statistics in Medicine, 24(6):893–909, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heo:2005:CSM

[1656] Moonseong Heo and Andrew C. Leon. Comparison of statistical methods for analysis of clustered binary observations. Statistics in Medicine, 24(6):911–923, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wall:2005:GEM

[1657] Melanie M. Wall, Yu Dai, and Lynn E. Eberly. GEE estimation of a misspecified time-varying covariate: an example with the effect of alcoholism treatment on medical utilization. Statistics in Medicine, 24(6): 925–939, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coffey:2005:EIW

[1658] Christopher S. Coffey, Gary L. Gadbury, Kevin R. Fontaine, Chenxi Wang, Richard Weindruch, and David B. Allison. The effects of intentional weight loss as a latent variable problem. Statistics in Medicine, 24(6):941–954, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ng:2005:TET

[1659] Hon Keung Tony Ng and Man-Lai Tang. Testing the equality of two Poisson means using the rate ratio. Statistics in Medicine, 24(6):955–965, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martin-Rodriguez:2005:HFH

[1660] Gloria Martín-Rodríguez and Carlos Murillo-Fort. High-frequency health data and spline functions. Statistics in Medicine, 24(6):967–981, March 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hernandez-Suarez:2005:MHV

[1661] C. M. Hernández-Suárez. Measuring HIV vaccine efficacy. Statistics in Medicine, 24(7):983–992, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2005:APM

[1662] Ian R. White and Simon G. Thompson. Adjusting for partially missing baseline measurements in randomized trials. Statistics in Medicine, 24(7):993–1007, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2005:SAP

[1663] Aiyi Liu, Chengqing Wu, Kai F. Yu, and Edmund Gehan. Supplementary analysis of probabilities at the termination of a group sequential phase II trial. Statistics in Medicine, 24(7):1009–1027, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vandebosch:2005:SAF

[1664] An Vandebosch, Els Goetghebeur, and Lut Van Damme. Structural accelerated failure time models for the effects of observed exposures on repeated events in a clinical trial. Statistics in Medicine, 24(7):1029–1046, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nandram:2005:BNR

[1665] Balgobin Nandram, Ning Liu, Jai Won Choi, and Lawrence Cox. Bayesian non-response models for categorical data from small areas: an application to BMD and age. Statistics in Medicine, 24(7):1047–1074, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walters:2005:UBM

[1666] Stephen J. Walters and Michael J. Campbell. The use of bootstrap methods for estimating sample size and analysing health-related quality of life outcomes. Statistics in Medicine, 24(7):1075–1102, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rondeau:2005:TLM

[1667] Virginie Rondeau, Kiros Berhane, and Duncan C. Thomas. A three-level model for binary time-series data: the effects of air pollution on school absences in the Southern California Children’s Health Study. Statistics in Medicine, 24(7):1103–1115, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gorlova:2005:EGR

[1668] Olga Gorlova, Bo Peng, David Yankelevitz, Claudia Henschke, and Marek Kimmel. Estimating the growth rates of primary lung tumours from samples with missing measurements. Statistics in Medicine, 24(7): 1117–1134, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sato:2005:LEC

[1669] T. Shun Sato. Letter to the Editor: Confidence limits for the ratio of two rates based on likelihood scores: non-iterative methods by P. L. Graham, K. Mengersen and A. P. Morton, Statistics in Medicine 2003; 22:2071–2083. Statistics in Medicine, 24(7):1135, April 15, 2005. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1116, 1117, 1545].

Graham:2005:AR

[1670] Petra Graham. Author’s reply. Statistics in Medicine, 24(7):1135, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mar:2005:BRE

[1671] Chris Del Mar. Book review: 1. Essential evidence-based medicine. Dan Mayer, Cambridge University Press, Cambridge, 2004. No. of pages: x + 381 (+ CD). Price: £24.99, $34.99 (paperback); £55.00, $80.00 (hardback). ISBN 0-521-54027-5 (paperback + CD); 0-521-83261-6 (hardback + CD). Statistics in Medicine, 24(7):1137, April 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chandrasekaran:2005:SDT

[1672] V. Chandrasekaran, G. Gopal, and A. Thomas. Summarizing data through a piecewise linear growth curve model. Statistics in Medicine, 24(8):1139–1151, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Durban:2005:SFS

[1673] M. Durbán, J. Harezlak, M. P. Wand, and R. J. Carroll. Simple fitting of subject-specific curves for longitudinal data. Statistics in Medicine, 24(8):1153–1167, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cadarso-Suarez:2005:NPE

[1674] Carmen Cadarso-Suárez, Javier Roca-Pardiñas, Adolfo Figueiras, and Wenceslao González-Manteiga. Non-parametric estimation of the odds ratios for continuous exposures using generalized additive models with an unknown link function. Statistics in Medicine, 24(8):1169–1184, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelhalter:2005:FPC

[1675] David J. Spiegelhalter. Funnel plots for comparing institutional performance. Statistics in Medicine, 24(8):1185–1202, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gustafson:2005:UPI

[1676] Paul Gustafson. The utility of prior information and stratification for parameter estimation with two screening tests but no gold standard. Statistics in Medicine, 24(8):1203–1217, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilliam:2005:SDP

[1677] R. M. Hilliam. Statistical discrimination in the presence of selection effects. Statistics in Medicine, 24(8):1219–1232, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moon:2005:AAB

[1678] Hojin Moon, Hongshik Ahn, and Ralph L. Kodell. An age-adjusted bootstrap-based poly-k test. Statistics in Medicine, 24(8):1233–1244, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:2005:CGF

[1679] Scott Evans and Lingling Li. A comparison of goodness of fit tests for the logistic GEE model. Statistics in Medicine, 24(8):1245–1261, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Manda:2005:BIR

[1680] Samuel O. M. Manda and Renate Meyer. Bayesian inference for recurrent events data using time-dependent frailty. Statistics in Medicine, 24(8): 1263–1274, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2005:IEP

[1681] Kung-Jong Lui. Interval estimation of the proportion ratio under multiple matching. Statistics in Medicine, 24(8):1275–1285, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fosgate:2005:LEA

[1682] Geoffrey T. Fosgate. Letter to the Editor: Advantages to transforming the receiver operating characteristic (ROC) curve into likelihood ratio co-ordinates by Nils P. Johnson, Statistics in Medicine 2004; 23:2257– 2266. Statistics in Medicine, 24(8):1287–1288, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1446].

Johnson:2005:AR

[1683] Nills P. Johnson. Author’s reply. Statistics in Medicine, 24(8):1288, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vermunt:2005:BRA

[1684] Jeroen K. Vermunt. Book review: 1. Analyzing categorical data. Jeffrey S. Simonoff, Springer, New York, 2003. No. of pages: xv + 496. Price: EUR 84.95, £65.50, $84.95. ISBN: 0-387-00749-0 (hardcover). Statistics in Medicine, 24(8):1289–1290, April 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thakkinstian:2005:MMA

[1685] Ammarin Thakkinstian, Patrick McElduff, Catherine D’Este, David Duffy, and John Attia. A method for meta-analysis of molecular association studies. Statistics in Medicine, 24(9):1291–1306, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2005:IHI

[1686] Catrin Tudur Smith, Paula R. Williamson, and Anthony G. Marson. Investigating heterogeneity in an individual patient data meta-analysis of time to event outcomes. Statistics in Medicine, 24(9):1307–1319, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Janes:2005:ITP

[1687] Holly Janes, Margaret Pepe, Charles Kooperberg, and Polly Newcomb. Identifying target populations for screening or not screening using logic regression. Statistics in Medicine, 24(9):1321–1338, May 15, 2005. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsui:2005:SSC

[1688] Shigeyuki Matsui. Sample size calculations for comparative clinical trials with over-dispersed Poisson process data. Statistics in Medicine, 24(9): 1339–1356, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lunt:2005:POO

[1689] Mark Lunt. Prediction of ordinal outcomes when the association between predictors and outcome differs between outcome levels. Statistics in Medicine, 24(9):1357–1369, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnhart:2005:AII

[1690] Huiman X. Barnhart, Jingli Song, and Michael J. Haber. Assessing intra, inter and total agreement with replicated readings. Statistics in Medicine, 24(9):1371–1384, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2005:AEP

[1691] Xun Chen, Xiaohui Luo, and Tom Capizzi. The application of enhanced parallel gatekeeping strategies. Statistics in Medicine, 24(9):1385–1397, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2005:SAC

[1692] Zhigang Zhang, Jianguo Sun, and Liuquan Sun. Statistical analysis of current status data with informative observation times. Statistics in Medicine, 24(9):1399–1407, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2005:CCM

[1693] Xinhua Liu, Yunling Du, Jeanne Teresi, and Deborah S. Hasin. Concordance correlation in the measurements of time to event. Statistics in Medicine, 24(9):1409–1420, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2005:CIB

[1694] Peter T. Kim and Christine H. Lee. Concomitant information in bioassay and semi-parametric estimation. Statistics in Medicine, 24(9):1421–1433, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bascoul-Mollevi:2005:TPS

[1695] C. Bascoul-Mollevi, S. Gourgou-Bourgade, and A. Kramar. Two-part statistics with paired data. Statistics in Medicine, 24(9):1435–1448, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2005:LEN

[1696] Xun Chen. Letter to the Editor: A note on non-parametric ANCOVA for covariate adjustment in randomized clinical trials by Emmanuel Lesaffre and Stephen Senn, Statistics in Medicine 2003; 22:3583–3596. Statistics in Medicine, 24(9):1449–1452, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1247].

Lesaffre:2005:AR

[1697] Emmanuel Lesaffre and Stephen Senn. Authors’ reply. Statistics in Medicine, 24(9):1452, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thabane:2005:BRD

[1698] Lehana Thabane. Book review: 1. Dicing with death: chance, risk and health. Stephen Senn, Cambridge University Press, Cambridge, 2003. No of pages: xii + 251. Price: £ 14.99, $28.00 (paperback). ISBN: 0-521-54023-2. Statistics in Medicine, 24(9):1453, May 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murphy:2005:EDD

[1699] S. A. Murphy. An experimental design for the development of adaptive treatment strategies. Statistics in Medicine, 24(10):1455–1481, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coad:2005:UTT

[1700] D. Stephen Coad and Anastasia Ivanova. The use of the triangular test with response-adaptive treatment allocation. Statistics in Medicine, 24 (10):1483–1493, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kelly:2005:SDA

[1701] Patrick J. Kelly, Nigel Stallard, and John C. Whittaker. Statistical design and analysis of pharmacogenetic trials. Statistics in Medicine, 24(10): 1495–1508, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Panhard:2005:EST

[1702] Xavière Panhard and France Mentré. Evaluation by simulation of tests based on non-linear mixed-effects models in pharmacokinetic interaction and bioequivalence cross-over trials. Statistics in Medicine, 24(10):1509– 1524, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duchateau:2005:PNP

[1703] Luc Duchateau and Paul Janssen. Pairwise non-parametric noninferiority tests in 3 × 3 cross-over trials: should we adjust for period? Statistics in Medicine, 24(10):1525–1536, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2005:AOS

[1704] Anastasia Ivanova, Robert C. Barrier, Jr., and Vance W. Berger. Adjusting for observable selection bias in block randomized trials. Statistics in Medicine, 24(10):1537–1546, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2005:IIO

[1705] P. R. Williamson and C. Gamble. Identification and impact of outcome selection bias in meta-analysis. Statistics in Medicine, 24(10):1547–1561, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2005:UPS

[1706] Peter C. Austin, Muhammad M. Mamdani, Therese A. Stukel, Geoffrey M. Anderson, and Jack V. Tu. The use of the propensity score for estimating treatment effects: administrative versus clinical data. Statistics in Medicine, 24(10):1563–1578, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hillis:2005:CDB

[1707] Stephen L. Hillis, Nancy A. Obuchowski, Kevin M. Schartz, and Kevin S. Berbaum. A comparison of the Dorfman–Berbaum–Metz and Obuchowski–Rockette methods for receiver operating characteristic (ROC) data. Statistics in Medicine, 24(10):1579–1607, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:2005:LE

[1708] Jack Cuzick. Letter to the editor. Statistics in Medicine, 24(10):1609, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deBock:2005:LEB

[1709] Geertruida H. de Bock, Catharina E. Jacobi, Minka A. Jonker, Nico J. D. Nagelkerke, and Johannes C. van Houwelingen. Letter to the Editor: A breast cancer prediction model by J. Tyrer, S. W. Duffy and J. Cuzick, Statistics in Medicine 2004; 23 (7):1111–1130. Statistics in Medicine, 24(10):1610–1612, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:2005:AR

[1710] Jack Cuzick, Stephen Duffy, and Jonathan Tyrer. Authors’ reply. Statistics in Medicine, 24(10):1612, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scurrah:2005:LEE

[1711] Katrina Scurrah, Lyle Gurrin, Lyle Palmer, and Paul Burton. Letter to the Editor: Estimation of genetic and environmental factors for binary traits using family data by Y. Pawitan, M. Reilly, E. Nilsson, S. Cnattingius and P. Lichtenstein, Statistics in Medicine 2004; 23:449– 465. Statistics in Medicine, 24(10):1613–1617, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1304].

Pawitan:2005:AR

[1712] Y. Pawitan, M. Reilly, E. Nilsson, S. Cnattingius, and P. Lichtenstein. Authors’ reply. Statistics in Medicine, 24(10):1617–1618, May 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shun:2005:SCS

[1713] Zhenming Shun, Eric Chi, Sylvain Durrleman, and Lloyd Fisher. Statistical consideration of the strategy for demonstrating clinical evidence of effectiveness-one larger vs two smaller pivotal studies. Statistics in Medicine, 24(11):1619–1637, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1716, 1714, 1715].

Koch:2005:LES

[1714] Gary G. Koch. Letter to the Editor: Statistical consideration of the strategy for demonstrating clinical evidence of effectiveness-one larger vs two smaller pivotal studies by Z. Shun, E. Chi, S. Durrleman and L. Fisher, Statistics in Medicine 2005; 24:1619–1637. Statistics in Medicine, 24(11):1639–1646, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1713, 1716, 1715].

Huque:2005:LES

[1715] Mohammad F. Huque. Letter to the Editor: Statistical consideration of the strategy for demonstrating clinical evidence of effectiveness-one larger vs two smaller pivotal studies by Z. Shun, E. Chi, S. Durrleman and L. Fisher, Statistics in Medicine 2005; 24:1619–1637. Statistics in Medicine, 24(11):1646–1651, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1713, 1716, 1714].

Shun:2005:LES

[1716] Zhenming Shun, Eric Chi, Sylvain Durrleman, and Lloyd Fisher. Letter to the Editor: Statistical consideration of the strategy for demonstrating clinical evidence of effectiveness-one larger vs two smaller pivotal studies Statistics in Medicine 2005; 24:1619–1637. Statistics in Medicine, 24(11): 1652–1656, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1713, 1714, 1715].

Spiegelman:2005:CEB

[1717] D. Spiegelman, B. Zhao, and J. Kim. Correlated errors in biased surrogates: study designs and methods for measurement error correction. Statistics in Medicine, 24(11):1657–1682, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaufman:2005:IEC

[1718] Sol Kaufman, Jay S. Kaufman, Richard F. MacLehose, Sander Greenland, and Charles Poole. Improved estimation of controlled direct effects in the presence of unmeasured confounding of intermediate variables. Statistics in Medicine, 24(11):1683–1702, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freidlin:2005:TTE

[1719] Boris Freidlin and Edward L. Korn. Testing treatment effects in the presence of competing risks. Statistics in Medicine, 24(11):1703–1712, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2183].

Bender:2005:GST

[1720] Ralf Bender, Thomas Augustin, and Maria Blettner. Generating survival times to simulate Cox proportional hazards models. Statistics in Medicine, 24(11):1713–1723, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2052].

OBrien:2005:RMA

[1721] Liam M. O’Brien and Garrett M. Fitzmaurice. Regression models for the analysis of longitudinal Gaussian data from multiple sources. Statistics in Medicine, 24(11):1725–1744, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2005:IEH

[1722] Lili Tian. Interval estimation and hypothesis testing of intraclass correlation coefficients: the generalized variable approach. Statistics in Medicine, 24(11):1745–1753, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ou:2005:MAA

[1723] San-San Ou, James P. Hughes, and Barbra A. Richardson. Methods for assessing the accuracy of PCR-based tests: comparisons and extensions. Statistics in Medicine, 24(11):1755–1764, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2005:IED

[1724] Kung-Jong Lui. Interval estimation of the difference in proportions under m-to-one matching. Statistics in Medicine, 24(11):1765–1776, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neuhauser:2005:LEE

[1725] Markus Neuhäuser. Letter to the Editor: Efficiency comparisons of rank and permutation tests by Phillip I. Good Statistics in Medicine 2004; 23:857. Statistics in Medicine, 24(11):1777–1778, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pesarin:2005:CPP

[1726] Fortunato Pesarin. Comments on a problem posed by Markus Neuhäuser. Statistics in Medicine, 24(11):1778–1780, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Good:2005:AR

[1727] Phillip I. Good. Author’s reply. Statistics in Medicine, 24(11):1780–1781, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morabia:2005:LEC

[1728] Alfredo Morabia. Letter to the Editor: C. Goehring, A. Perrier and A. Morabia, ‘Spectrum bias: a quantitative and graphical analysis of the variability of medical diagnostic test performance’. Statistics in Medicine 2004; 23:(1) 125–135. Statistics in Medicine, 24(11):1782, June 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1279].

Shao:2005:SIC

[1729] Jun Shao, Mark Chang, and Shein-Chung Chow. Statistical inference for cancer trials with treatment switching. Statistics in Medicine, 24 (12):1783–1790, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willan:2005:VIO

[1730] Andrew R. Willan and Eleanor M. Pinto. The value of information and optimal clinical trial design. Statistics in Medicine, 24(12):1791–1806, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xing:2005:MEV

[1731] Biao Xing and Jitendra Ganju. A method to estimate the variance of an endpoint from an on-going blinded trial. Statistics in Medicine, 24(12): 1807–1814, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:2005:TFA

[1732] A. Lawrence Gould. Timing of futility analyses for ‘proof of concept’ trials. Statistics in Medicine, 24(12):1815–1835, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duncan:2005:BEN

[1733] Bradford W. Duncan and Ingram Olkin. Bias of estimates of the number needed to treat. Statistics in Medicine, 24(12):1837–1848, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Toti:2005:AML

[1734] Simona Toti, Annibale Biggeri, and Francesco Forastiere. Adult myeloid leukaemia and radon exposure: a Bayesian model for a case-control study with error in covariates. Statistics in Medicine, 24(12):1849–1864, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lam:2005:EPC

[1735] K. F. Lam, Daniel Y. T. Fong, and O. Y. Tang. Estimating the proportion of cured patients in a censored sample. Statistics in Medicine, 24(12): 1865–1879, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vogel:2005:EJM

[1736] C. Vogel, H. Brenner, A. Pfahlberg, and O. Gefeller. The effects of joint misclassification of exposure and disease on the attributable risk. Statistics in Medicine, 24(12):1881–1896, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cooil:2005:PPM

[1737] Bruce Cooil and Paolo Raggi. On the prediction and prevention of myocardial infarctions: models based on retrospective and doubly censored prospective data. Statistics in Medicine, 24(12):1897–1918, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dreassi:2005:STM

[1738] Emanuela Dreassi, Annibale Biggeri, and Dolores Catelan. Space-time models with time-dependent covariates for the analysis of the temporal lag between socioeconomic factors and lung cancer mortality. Statistics in Medicine, 24(12):1919–1932, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandel:2005:ESM

[1739] M. Mandel, N. Galai, and E. Simchen. Evaluating survival model performance: a graphical approach. Statistics in Medicine, 24(12):1933–1945, June 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:2005:CAA

[1740] Peter F. Thall and J. Kyle Wathen. Covariate-adjusted adaptive randomization in a sarcoma trial with multi-stage treatments. Statistics in Medicine, 24(13):1947–1964, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pinto:2005:CEA

[1741] Eleanor M. Pinto, Andrew R. Willan, and Bernie J. O’Brien. Costeffectiveness analysis for multinational clinical trials. Statistics in Medicine, 24(13):1965–1982, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:2005:UBD

[1742] James P. Hughes. Using baseline data to design a group randomized trial. Statistics in Medicine, 24(13):1983–1994, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coad:2005:SUD

[1743] D. Stephen Coad and Anastasia Ivanova. Sequential urn designs with elimination for comparing K ≥ 3 treatments. Statistics in Medicine, 24(13):1995–2009, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ONeill:2005:BMC

[1744] Philip D. O’Neill and Peter J. Marks. Bayesian model choice and infection route modelling in an outbreak of norovirus. Statistics in Medicine, 24(13):2011–2024, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2005:PAU

[1745] S. D. Walter. The partial area under the summary ROC curve. Statistics in Medicine, 24(13):2025–2040, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2817].

Cadwell:2005:MCR

[1746] Betsy L. Cadwell, Philip J. Smith, and Andrew L. Baughman. Methods for capture-recapture analysis when cases lack personal identifiers. Statistics in Medicine, 24(13):2041–2051, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2005:AEI

[1747] Haitao Chu, Thomas W. Kensler, and Alvaro Muñoz. Assessing the effect of interventions in the context of mixture distributions with detection limits. Statistics in Medicine, 24(13):2053–2067, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2005:GSM

[1748] Antai Wang and Edmund A. Gehan. Gene selection for microarray data analysis using principal component analysis. Statistics in Medicine, 24 (13):2069–2087, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gustafson:2005:ELR

[1749] Paul Gustafson, Azad M. R. Kazi, and Adrian R. Levy. Extending logistic regression to model diffuse interactions. Statistics in Medicine, 24(13): 2089–2104, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2005:BRS

[1750] Christopher Cox. Book review: 1. Statistical tools for non-linear regression. A practical guide with S-Plus and R examples (2nd edn). Sylvie Huet, Annie Bouvier, Marie-Anne Poursat and Emmanuel Jolivet, Springer, New York, 2004. No. of pages: xiv + 232. Price: $74.94, £57.50, EUR74.95 (+ VAT). ISBN 0-387-40081-8. Statistics in Medicine, 24(13):2105–2106, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collier:2005:BRH

[1751] Timothy Collier. Book review: 2. A handbook of statistical analyses using Stata (3rd edn). Sophia Rabe-Hesketh and Brian Everitt, Chapman and Hall/CRC, Boca Raton, 2004. No. of pages: xiii + 308. Price: $49.95. ISBN: 1-58488-404-5. Statistics in Medicine, 24(13):2106–2107, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:2005:BRB

[1752] Donald A. Berry. Book review: 3. Bayesian approaches to clinical trials and health-care evaluation. David J. Spiegelhalter, Keith R. Abrams and Jonathan P. Myles, Wiley, Chichester, 2004. No. of Pages: 406. Price: £45.00, EUR67.50, $85.00. ISBN: 0-471-49975-7. Statistics in Medicine, 24(13):2107–2109, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gustafson:2005:BRM

[1753] Paul Gustafson. Book review: 4. Multilevel statistical models (3rd edn). Harvey Goldstein, Arnold, London, 2003. No. of pages: xv + 253. Price: £45.00 (hardback). ISBN 0-340-80655-9 (hardback). Statistics in Medicine, 24(13):2109–2110, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2005:CIM

[1754] Lingqi Tang, Juwon Song, Thomas R. Belin, and Jürgen Unützer. A comparison of imputation methods in a longitudinal randomized clinical trial. Statistics in Medicine, 24(14):2111–2128, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2005:ILS

[1755] Guoguang Ma, Andrea B. Troxel, and Daniel F. Heitjan. An index of local sensitivity to nonignorable drop-out in longitudinal modelling. Statistics in Medicine, 24(14):2129–2150, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2134].

Quan:2005:MAM

[1756] Hui Quan, Xiaohui Luo, and Tom Capizzi. Multiplicity adjustment for multiple endpoints in clinical trials with multiple doses of an active treatment. Statistics in Medicine, 24(14):2151–2170, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hunsberger:2005:DET

[1757] Sally Hunsberger, Lawrence V. Rubinstein, Janet Dancey, and Edward L. Korn. Dose escalation trial designs based on a molecularly targeted endpoint. Statistics in Medicine, 24(14):2171–2181, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tighiouart:2005:FBM

[1758] Mourad Tighiouart, André Rogatko, and James S. Babb. Flexible Bayesian methods for cancer phase I clinical trials. dose escalation with overdose control. Statistics in Medicine, 24(14):2183–2196, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tchernis:2005:UDC

[1759] Rusty Tchernis, Marcela Horvitz-Lennon, and Sharon-Lise T. Normand. On the use of discrete choice models for causal inference. Statistics in Medicine, 24(14):2197–2212, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2005:ICC

[1760] Lili Tian. Inferences on the common coefficient of variation. Statistics in Medicine, 24(14):2213–2220, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luan:2005:DIA

[1761] Xianqun Luan, Wei Pan, Susan G. Gerberich, and Bradley P. Carlin. Does it always help to adjust for misclassification of a binary outcome in logistic regression? Statistics in Medicine, 24(14):2221–2234, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhattacharyya:2005:NTA

[1762] Mouchumi Bhattacharyya and John P. Klein. A note on testing in Aalen’s additive hazards regression models. Statistics in Medicine, 24(14):2235– 2240, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2005:MAG

[1763] John R. Thompson, Cosetta Minelli, Keith R. Abrams, Martin D. Tobin, and Richard D. Riley. Meta-analysis of genetic studies using Mendelian randomization — a multivariate approach. Statistics in Medicine, 24(14): 2241–2254, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaufman:2005:SBN

[1764] Cari G. Kaufman, Valérie Ventura, and Robert E. Kass. Spline-based non-parametric regression for periodic functions and its application to directional tuning of neurons. Statistics in Medicine, 24(14):2255–2265, July 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2005:FCT

[1765] Shuying S. Li, Jeannette Bigler, Johanna W. Lampe, John D. Potter, and Ziding Feng. FDR-controlling testing procedures and sample size determination for microarrays. Statistics in Medicine, 24(15):2267–2280, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kooperberg:2005:STS

[1766] Charles Kooperberg, Aaron Aragaki, Andrew D. Strand, and James M. Olson. Significance testing for small microarray experiments. Statistics in Medicine, 24(15):2281–2298, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flanders:2005:TTH

[1767] W. D. Flanders, M. J. Khoury, Q. H. Yang, and H. Austin. Tests of traithaplotype association when linkage phase is ambiguous, appropriate for matched case-control and cohort studies with competing risks. Statistics in Medicine, 24(15):2299–2316, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:2005:MMC

[1768] Bernard F. Cole, Marco Bonetti, Alan M. Zaslavsky, and Richard D. Gelber. A multistate Markov chain model for longitudinal, categorical quality-of-life data subject to non-ignorable missingness. Statistics in Medicine, 24(15):2317–2334, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:2005:AHM

[1769] Rachel MacKay Altman and A. John Petkau. Application of hidden Markov models to multiple sclerosis lesion count data. Statistics in Medicine, 24(15):2335–2344, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demirtas:2005:MIU

[1770] Hakan Demirtas. Multiple imputation under Bayesianly smoothed pattern-mixture models for non-ignorable drop-out. Statistics in Medicine, 24(15):2345–2363, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernatsky:2005:BMI

[1771] Sasha Bernatsky, Lawrence Joseph, Patrick Bélisle, Jean-François Boivin, Raghu Rajan, Andrew Moore, and Ann Clarke. Bayesian modelling of imperfect ascertainment methods in cancer studies. Statistics in Medicine, 24(15):2365–2379, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:2005:PBB

[1772] Kim E. Andersen and Malene Højbjerre. A population-based Bayesian approach to the minimal model of glucose and insulin homeostasis. Statistics in Medicine, 24(15):2381–2400, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2005:HVV

[1773] Paul C. Lambert, Alex J. Sutton, Paul R. Burton, Keith R. Abrams, and David R. Jones. How vague is vague? A simulation study of the impact of the use of vague prior distributions in MCMC using WinBUGS. Statistics in Medicine, 24(15):2401–2428, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holford:2005:BRE

[1774] Theodore R. Holford. Book review: 1. Environmental statistics: methods and applications. Vic Barnett, Wiley, Chichester, 2004. No. of pages: xi + 293. Price: $78.00, £55.00, EUR82.50. ISBN: 0-471-48971-90. Statistics in Medicine, 24(15):2429–2430, August 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2005:DAM

[1775] Jonathan J. Shuster. Diagnostics for assumptions in moderate to large simple clinical trials: do they really help? Statistics in Medicine, 24 (16):2431–2438, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2033].

Koyama:2005:FTS

[1776] Tatsuki Koyama, Allan R. Sampson, and Leon J. Gleser. A framework for two-stage adaptive procedures to simultaneously test non-inferiority and superiority. Statistics in Medicine, 24(16):2439–2456, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kittelson:2005:GSC

[1777] John M. Kittelson, Katrina Sharples, and Scott S. Emerson. Group sequential clinical trials for longitudinal data with analyses using summary statistics. Statistics in Medicine, 24(16):2457–2475, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Atkinson:2005:ABC

[1778] A. C. Atkinson and A. Biswas. Adaptive biased-coin designs for skewing the allocation proportion in clinical trials with normal responses. Statistics in Medicine, 24(16):2477–2492, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hamm:2005:AIT

[1779] Adam K. Hamm, W. Hans Carter, Jr., and Chris Gennings. Analysis of an interaction threshold in a mixture of drugs and/or chemicals. Statistics in Medicine, 24(16):2493–2507, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Logan:2005:PMC

[1780] Brent R. Logan, Hong Wang, and Mei-Jie Zhang. Pairwise multiple comparison adjustment in survival analysis. Statistics in Medicine, 24 (16):2509–2523, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuk:2005:ESI

[1781] Anthony Y. C. Kuk and Stefan Ma. The estimation of SARS incubation distribution from serial interval data using a convolution likelihood. Statistics in Medicine, 24(16):2525–2537, August 30, 2005. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Christensen:2005:STA

[1782] William F. Christensen and F. Zerrin Yetkin. Spatio-temporal analysis of auditory cortex activation as detected with silent event related fMRI. Statistics in Medicine, 24(16):2539–2556, August 30, 2005. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Solis-Trapala:2005:RAO

[1783] I. L. Solis-Trapala and V. T. Farewell. Regression analysis of overdispersed correlated count data with subject specific covariates. Statistics in Medicine, 24(16):2557–2575, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coen:2005:LES

[1784] Pietro Coen. Letter to the Editor: Seasonal recruitment in case-control studies by Stallard and Tombs, Statistics in Medicine 2004; 23:3193– 3207. Statistics in Medicine, 24(16):2577–2580, August 30, 2005. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1511].

Stallard:2005:AR

[1785] Nigel Stallard and Lee Tombs. Authors’ reply. Statistics in Medicine, 24 (16):2580, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beyersmann:2005:BRH

[1786] Jan Beyersmann. Book review: 1. Handbook of statistics 23: Advances in survival analysis. N. Balakrishnan and C. R. Rao (eds), Elsevier, Amsterdam, 2004. No. of pages: xxv + 795. Price: $175.00, £116.50, EUR 175.00 (see www.elsevier.com for availability). ISBN 0-444-50079-0. Statistics in Medicine, 24(16):2581, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). URL http://www.elsevier.

com.

Glasziou:2005:BRS

[1787] Paul Glasziou. Book review: 2. The statistical evaluation of medical tests for classification and prediction. Margaret Sullivan Pepe, Oxford University Press, Oxford, 2003. No. of pages: xvi + 302. Price: £65.00, $124.50 (Hardback); £39.50, $69.50 (Paperback). ISBN: 0-19-850984-7 (Hardback), 0-19-856582-8 (Paperback). Statistics in Medicine, 24(16): 2582, August 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2005:SST

[1788] Sin-Ho Jung and Chul W. Ahn. Sample size for a two-group comparison of repeated binary measurements using GEE. Statistics in Medicine, 24 (17):2583–2596, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

London:2005:OTS

[1789] Wendy B. London and Myron N. Chang. One- and two-stage designs for stratified phase II clinical trials. Statistics in Medicine, 24(17):2597– 2611, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gonen:2005:PDR

[1790] Mithat Gönen. Planning a dose-response study with subject-specific doses. Statistics in Medicine, 24(17):2613–2623, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schluter:2005:SPT

[1791] P. J. Schluter and R. S. Ware. Single patient (n-of-1) trials with binary treatment preference. Statistics in Medicine, 24(17):2625–2636, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alexander:2005:AIR

[1792] Neal Alexander and Paul Emerson. Analysis of incidence rates in cluster-randomized trials of interventions against recurrent infections, with an application to trachoma. Statistics in Medicine, 24(17):2637– 2647, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutson:2005:MDT

[1793] Alan D. Hutson. Medical diagnostic testing models for multifactorial disease classifications. Statistics in Medicine, 24(17):2649–2657, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graubard:2005:UPB

[1794] B. I. Graubard, R. Sowmya Rao, and Joseph L. Gastwirth. Using the Peters–Belson method to measure health care disparities from complex survey data. Statistics in Medicine, 24(17):2659–2668, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DellOmodarme:2005:PFD

[1795] M. Dell’Omodarme and M. C. Prati. The probability of failing in detecting an infectious disease at entry points into a country. Statistics in Medicine, 24(17):2669–2679, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kheiri:2005:CFM

[1796] Soleiman Kheiri, Mohammad Reza Meshkani, and Soghrat Faghihzadeh. A correlated frailty model for analysing risk factors in bilateral corneal graft rejection for keratoconus: a Bayesian approach. Statistics in Medicine, 24(17):2681–2693, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nakaya:2005:GWP

[1797] T. Nakaya, A. S. Fotheringham, C. Brunsdon, and M. Charlton. Geographically weighted Poisson regression for disease association mapping. Statistics in Medicine, 24(17):2695–2717, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vrijens:2005:MAB

[1798] Bernard Vrijens, Els Goetghebeur, Erik de Klerk, Richard Rode, Steve Mayer, and John Urquhart. Modelling the association between adherence and viral load in HIV-infected patients. Statistics in Medicine, 24(17): 2719–2731, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ORourke:2005:LEB

[1799] Keith O’Rourke and Douglas G. Altman. Letter to the Editor: Bayesian random effects meta-analysis of trials with binary outcomes: methods for the absolute risk difference and relative risk scales by D. E. Warn, S. G. Thompson and D. J. Spiegelhalter, Statistics in Medicine 2002; 21: 1601–1623. Statistics in Medicine, 24(17):2733–2742, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [790].

Spiegelhalter:2005:AR

[1800] David J. Spiegelhalter, David E. Warn, and Simon G. Thompson. Authors’ reply. Statistics in Medicine, 24(17):2743, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McPherson:2005:BRS

[1801] Klim McPherson. Book review: 1. Statistical methods for rates and proportions (3rd edn). Joseph L. Fleiss, Bruce Levin and Myunghee Cho Paik, Wiley, New Jersey, 2003. No. of pages: xxvii + 760. Price: $99.95 (hardcover). ISBN: 0-471-52629-0. Statistics in Medicine, 24(17):2744– 2745, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holford:2005:BRF

[1802] Nicholas H. G. Holford. Book review: 2. Fitting models to biological data using linear and non-linear regression: a practical guide to curve fitting. Harvey Motulsky and Arthur Christopoulos, Oxford University Press, Oxford, 2004. No. of pages: 352. Price: £19.99, $29.29 (paperback); £40.00, $65.00 (hardcover). ISBN: 0-19-517180-2 (paperback), 0-19-517179-9 (hardcover). Statistics in Medicine, 24(17):2745–2746, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carroll:2005:C

[1803] Raymond J. Carroll. Correction. Statistics in Medicine, 24(17):2746, September 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2005:RMF

[1804] John M. Lachin. A review of methods for futility stopping based on conditional power. Statistics in Medicine, 24(18):2747–2764, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2005:MCT

[1805] Wenlei Liu, Wei Zhao, Michele L. Shaffer, Nikolina Icitovic, and Gary A. Chase. Modelling clinical trials in heterogeneous samples. Statistics in Medicine, 24(18):2765–2775, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2005:RPS

[1806] Vance W. Berger. The reverse propensity score to detect selection bias and correct for baseline imbalances. Statistics in Medicine, 24(18):2777– 2787, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2005:ETE

[1807] Leena Choi, Francesca Dominici, Scott L. Zeger, and Peter Ouyang. Estimating treatment efficacy over time: a logistic regression model for binary longitudinal outcomes. Statistics in Medicine, 24(18):2789–2805, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Putter:2005:LTS

[1808] H. Putter, M. Sasako, H. H. Hartgrink, C. J. H. van de Velde, and J. C. van Houwelingen. Long-term survival with non-proportional hazards: results from the Dutch gastric cancer trial. Statistics in Medicine, 24 (18):2807–2821, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanHouwelingen:2005:IAS

[1809] Hans C. van Houwelingen, Cornelis J. H. van de Velde, and Theo Stijnen. Interim analysis on survival data: its potential bias and how to repair it. Statistics in Medicine, 24(18):2823–2835, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zaslavsky:2005:AEN

[1810] Boris G. Zaslavsky. Assessing equivalence or non-inferiority in screening tests with paired and independent binomial endpoints through confidence intervals for the odds ratio. Statistics in Medicine, 24(18):2837–2855, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fosgate:2005:MES

[1811] Geoffrey T. Fosgate. Modified exact sample size for a binomial proportion with special emphasis on diagnostic test parameter estimation. Statistics in Medicine, 24(18):2857–2866, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinzl:2005:CUP

[1812] Harald Heinzl, Thomas Waldhör, and Martina Mittlböck. Careful use of pseudo R-squared measures in epidemiological studies. Statistics in Medicine, 24(18):2867–2872, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bandos:2005:PTS

[1813] Andriy I. Bandos, Howard E. Rockette, and David Gur. A permutation test sensitive to differences in areas for comparing ROC curves from a paired design. Statistics in Medicine, 24(18):2873–2893, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chung:2005:LTA

[1814] Hwan Chung, YouSung Park, and Stephanie T. Lanza. Latent transition analysis with covariates: pubertal timing and substance use behaviours in adolescent females. Statistics in Medicine, 24(18):2895–2910, September 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tobin:2005:ATE

[1815] Martin D. Tobin, Nuala A. Sheehan, Katrina J. Scurrah, and Paul R. Burton. Adjusting for treatment effects in studies of quantitative traits: antihypertensive therapy and systolic blood pressure. Statistics in Medicine, 24(19):2911–2935, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tong:2005:ISM

[1816] Xingwei Tong, Zhongguo Zheng, and Zhi Geng. Identifiability, stratification and minimum variance estimation of causal effects. Statistics in Medicine, 24(19):2937–2952, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2005:EAR

[1817] Kung-Jong Lui. Estimation of attributable risk for case-control studies with multiple matching. Statistics in Medicine, 24(19):2953–2962, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stamey:2005:BSS

[1818] James D. Stamey, John W. Seaman, and Dean M. Young. Bayesian sample-size determination for inference on two binomial populations with no gold standard classifier. Statistics in Medicine, 24(19):2963–2976, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2005:SSC

[1819] Hae-Young Kim, John M. Williamson, and Cynthia M. Lyles. Samplesize calculations for studies with correlated ordinal outcomes. Statistics in Medicine, 24(19):2977–2987, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moltchanova:2005:HBB

[1820] E. Moltchanova, A. Penttinen, and M. Karvonen. A hierarchical Bayesian birth cohort analysis from incomplete registry data: evaluating the trends in the age of onset of insulin-dependent diabetes mellitus (T1DM). Statistics in Medicine, 24(19):2989–3004, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hollenbeak:2005:FFR

[1821] Christopher S. Hollenbeak. Functional form and risk adjustment of hospital costs: Bayesian analysis of a Box–Cox random coefficients model. Statistics in Medicine, 24(19):3005–3018, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

James:2005:RSS

[1822] Kenneth E. James, Roberta F. White, and Helena C. Kraemer. Repeated split sample validation to assess logistic regression and recursive partitioning: an application to the prediction of cognitive impairment. Statistics in Medicine, 24(19):3019–3035, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schaubel:2005:VEC

[1823] Douglas E. Schaubel. Variance estimation for clustered recurrent event data with a small number of clusters. Statistics in Medicine, 24(19):3037– 3051, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiang:2005:IDT

[1824] Liming Xiang, Kelvin K. W. Yau, Andy H. Lee, and Wing K. Fung. Influence diagnostics for two-component Poisson mixture regression models: applications in public health. Statistics in Medicine, 24(19):3053–3071, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teresi:2005:BRM

[1825] Jeanne A. Teresi. Book review: 0. Measurement theory and practice: the world through quantification. David J. Hand, Hodder Arnold, London (distributed in U.S.A. by Oxford University Press, New York), 2004. No of pages: x + 320. Price: £45.00, $60.00. ISBN: 0-340-67783-X. Statistics in Medicine, 24(19):3073–3074, October 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2005:REB

[1826] Ralph B. D’Agostino, Michael J. Campbell, and Joel B. Greenhouse. Restructuring the Editorial Board. Statistics in Medicine, 24(20):3075, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2005:MTP

[1827] Sin-Ho Jung, Kouros Owzar, and Stephen L. George. A multiple testing procedure to associate gene expression levels with survival. Statistics in Medicine, 24(20):3077–3088, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2005:AKM

[1828] Jun Xie and Chaofeng Liu. Adjusted Kaplan–Meier estimator and logrank test with inverse probability of treatment weighting for survival data. Statistics in Medicine, 24(20):3089–3110, October 30, 2005. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samuelsen:2005:SCT

[1829] S. O. Samuelsen, T. F. Wisløff, and A. Skrondal. A simple correction for ties when censoring times depend on covariates. Statistics in Medicine, 24 (20):3111–3121, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cong:2005:EAS

[1830] Xiuyu J. Cong, Yu Shen, and Anthony B. Miller. Estimation of agespecific sensitivity and sojourn time in breast cancer screening studies. Statistics in Medicine, 24(20):3123–3138, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stovring:2005:EPI

[1831] Henrik Støvring and Werner Vach. Estimation of prevalence and incidence based on occurrence of health-related events. Statistics in Medicine, 24(20):3139–3154, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2005:DPI

[1832] Ying Lu, Hua Jin, and Kathleen R. Lamborn. A design of phase II cancer trials using total and complete response endpoints. Statistics in Medicine, 24(20):3155–3170, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conigliani:2005:SPM

[1833] C. Conigliani and A. Tancredi. Semi-parametric modelling for costs of health care technologies. Statistics in Medicine, 24(20):3171–3184, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katki:2005:AUR

[1834] Hormuzd A. Katki, Eric A. Engels, and Philip S. Rosenberg. Assessing uncertainty in reference intervals via tolerance intervals: application to a mixed model describing HIV infection. Statistics in Medicine, 24 (20):3185–3198, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jennen-Steinmetz:2005:NAS

[1835] C. Jennen-Steinmetz and S. Wellek. A new approach to sample size calculation for reference interval studies. Statistics in Medicine, 24(20):3199– 3212, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2005:LER

[1836] Toshiro Tango. Linear equations with random variables. Statistics in Medicine, 24(20):3213–3222, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2186].

Tian:2005:IMZ

[1837] Lili Tian. Inferences on the mean of zero-inflated lognormal data: the generalized variable approach. Statistics in Medicine, 24(20):3223–3232, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julious:2005:INN

[1838] Steven A. Julious. Issues with number needed to treat. Statistics in Medicine, 24(20):3233–3235, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wade:2005:BRH

[1839] Angie Wade. Book review: 1. A handbook of statistical analyses using SPSS. Sabine Landau and Brian S. Everitt, Chapman and Hall/ CRC, Boca Raton, 2004. No. of pages: xii + 354. price: $44.95, £24.99. ISBN: 1-58488-369-3. Statistics in Medicine, 24(20):3236–3237, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nair:2005:BRS

[1840] Rama Chandran Nair. Book review: 2. Statistics for epidemiology. Nicholas P. Jewell, Chapman and Hall/CRC, Boca Raton, 2004. No. of pages: xiv + 333. Price: $69.95 (hardcover). ISBN: 1-58488-433-9. Statistics in Medicine, 24(20):3237–3238, October 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:2005:STE

[1841] James F. Troendle, Aiyi Liu, Chengqing Wu, and Kai F. Yu. Sequential testing for efficacy in clinical trials with non-transient effects. Statistics in Medicine, 24(21):3239–3250, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsiao:2005:INV

[1842] Chuhsing Kate Hsiao, Miao-Yu Tsai, and Ho-Min Chen. Inference of nested variance components in a longitudinal myopia intervention trial. Statistics in Medicine, 24(21):3251–3267, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sama:2005:IDM

[1843] Wilson Sama, Seth Owusu-Agyei, Ingrid Felger, Penelope Vounatsou, and Tom Smith. An immigration-death model to estimate the duration of malaria infection when detectability of the parasite is imperfect. Statistics in Medicine, 24(21):3269–3288, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2005:RAC

[1844] K. F. Cheng and W. J. Lin. Retrospective analysis of case-control studies when the population is in Hardy-Weinberg equilibrium. Statistics in Medicine, 24(21):3289–3310, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2005:CIC

[1845] Lili Tian. On confidence intervals of a common intraclass correlation coefficient. Statistics in Medicine, 24(21):3311–3318, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2005:RSS

[1846] Haiying Chen, Elizabeth A. Stasny, and Douglas A. Wolfe. Ranked set sampling for efficient estimation of a population proportion. Statistics in Medicine, 24(21):3319–3329, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raman:2005:MER

[1847] Rema Raman and Donald Hedeker. A mixed-effects regression model for three-level ordinal response data. Statistics in Medicine, 24(21):3331– 3345, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnhart:2005:AVL

[1848] Huiman X. Barnhart, Jingli Song, and Robert H. Lyles. Assay validation for left-censored data. Statistics in Medicine, 24(21):3347–3360, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gurrin:2005:TBS

[1849] Lyle C. Gurrin, Katrina J. Scurrah, and Martin L. Hazelton. Tutorial in biostatistics: spline smoothing with linear mixed models. Statistics in Medicine, 24(21):3361–3381, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julious:2005:LET

[1850] Steven A. Julious. Letter to the Editor: Two-sided confidence intervals for the single proportion: comparison of seven methods by Robert G. Newcombe, Statistics in Medicine 1998; 17:857–872. Statistics in Medicine, 24(21):3383–3384, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [9].

OBrien:2005:LES

[1851] Peter C. O’Brien. Letter to the Editor: Semi-parametric and nonparametric methods for clinical trials with incomplete data Statistics in Medicine 2005; 24:341–358. Statistics in Medicine, 24(21):3385, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1612].

Monaco:2005:BAM

[1852] Jane Monaco, Jianwen Cai, and James Grizzle. Bootstrap analysis of multivariate failure time data. Statistics in Medicine, 24(22):3387–3400, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clarke:2005:ACB

[1853] Paul S. Clarke. Analysing change based on two measures taken under different conditions. Statistics in Medicine, 24(22):3401–3415, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yafune:2005:EIC

[1854] Akifumi Yafune, Takashi Funatogawa, and Makio Ishiguro. Extended information criterion (EIC) approach for linear mixed effects models under restricted maximum likelihood (REML) estimation. Statistics in Medicine, 24(22):3417–3429, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2005:SIQ

[1855] V. T. Farewell, A. M. Herzberg, K. W. James, L. M. Ho, and G. M. Leung. SARS incubation and quarantine times: when is an exposed individual known to be disease free? Statistics in Medicine, 24(22): 3431–3445, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finkenstadt:2005:MAD

[1856] B. F. Finkenstädt, A. Morton, and D. A. Rand. Modelling antigenic drift in weekly flu incidence. Statistics in Medicine, 24(22):3447–3461, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:2005:BAP

[1857] G. J. Prescott and P. H. Garthwaite. A Bayesian approach to prospective binary outcome studies with misclassification in a binary risk factor. Statistics in Medicine, 24(22):3463–3477, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guo:2005:SSP

[1858] Xu Guo, Wei Pan, John E. Connett, Peter J. Hannan, and Simone A. French. Small-sample performance of the robust score test and its modifications in generalized estimating equations. Statistics in Medicine, 24 (22):3479–3495, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saha:2005:BCM

[1859] Krishna K. Saha and Sudhir R. Paul. Bias-corrected maximum likelihood estimator of the intraclass correlation parameter for binary data. Statistics in Medicine, 24(22):3497–3512, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2005:TBM

[1860] Tae Young Yang. A tree-based model for homogeneous groupings of multinomials. Statistics in Medicine, 24(22):3513–3522, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Behseta:2005:TET

[1861] Sam Behseta and Robert E. Kass. Testing equality of two functions using BARS. Statistics in Medicine, 24(22):3523–3534, November 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borm:2005:PCR

[1862] George F. Borm, René J. F. Melis, Steven Teerenstra, and Petronella G. Peer. Pseudo cluster randomization: a treatment allocation method to minimize contamination and selection bias. Statistics in Medicine, 24 (23):3535–3547, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hines:2005:AMA

[1863] R. J. O’Hara Hines and W. G. S. Hines. An appraisal of methods for the analysis of longitudinal categorical data with MAR drop-outs. Statistics in Medicine, 24(23):3549–3563, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2005:CIC

[1864] Man-Lai Tang, Nian-Sheng Tang, and Ivan S. F. Chan. Confidence interval construction for proportion difference in small-sample paired studies. Statistics in Medicine, 24(23):3565–3579, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:2005:MCB

[1865] James F. Troendle. Multiple comparisons between two groups on multiple Bernoulli outcomes while accounting for covariates. Statistics in Medicine, 24(23):3581–3591, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersson:2005:MSP

[1866] Mikael Andersson, Karl Ekdahl, Sigvard Mölstad, Kristina Persson, Hans Bertil Hansson, and Johan Giesecke. Modelling the spread of penicillin-resistant streptococcus pneumoniae in day-care and evaluation of intervention. Statistics in Medicine, 24(23):3593–3607, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2005:EPL

[1867] Jason Roy and Vincent Mor. The effect of provider-level ascertainment bias on profiling nursing homes. Statistics in Medicine, 24(23):3609– 3629, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Annis:2005:SPI

[1868] David H. Annis and Bruce A. Craig. Statistical properties and inference of the antimicrobial MIC test. Statistics in Medicine, 24(23):3631–3644, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chakraborty:2005:BNN

[1869] Sounak Chakraborty, Malay Ghosh, Tapabrata Maiti, and Ashutosh Tewari. Bayesian neural networks for bivariate binary data: an application to prostate cancer study. Statistics in Medicine, 24(23):3645–3662, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanDommelen:2005:IGC

[1870] P. van Dommelen, S. van Buuren, G. R. J. Zandwijken, and P. H. Verkerk. Individual growth curve models for assessing evidence-based referral criteria in growth monitoring. Statistics in Medicine, 24(23):3663–3674, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2005:MRC

[1871] Stephen Senn. Misunderstandings regarding clinical cross-over trials. Statistics in Medicine, 24(23):3675–3678, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McMichael:2005:BRH

[1872] Anthony J. McMichael. Book review: 1. A history of epidemiologic methods and concepts. Alfredo Morabia (editor), Birkhauser Verlag, Basel, 2004. Pages: xvii + 403. Price: EUR 98 (softcover). ISBN: 3-7643-6818-7. Statistics in Medicine, 24(23):3679–3680, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cruz-Orive:2005:BRS

[1873] Luis M. Cruz-Orive. Book review: 2. Stereology for statisticians. Adrian Baddeley and Eva B. Vedel Jensen, Chapman and Hall/CRC, Boca Raton, 2005. No. of pages: 395 (hardcover). Price: $79.95. ISBN: 1-58488-405-3. Statistics in Medicine, 24(23):3681–3682, December 15, 2005. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2005:C

[1874] Kung-Jong Lui. Correction. Statistics in Medicine, 24(23):3682, December 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stijnen:2005:P

[1875] Theo Stijnen and Hans Van Houwelingen. Preface. Statistics in Medicine, 24(24):3683, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2005:ISC

[1876] Anonymous. International Society for Clinical Biostatistics. Statistics in Medicine, 24(24):3685–3686, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:2005:ETC

[1877] M. S. Pepe. Evaluating technologies for classification and prediction in medicine. Statistics in Medicine, 24(24):3687–3696, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Posch:2005:TEF

[1878] Martin Posch, Franz Koenig, Michael Branson, Werner Brannath, Cornelia Dunger-Baldauf, and Peter Bauer. Testing and estimation in flexible group sequential designs with adaptive treatment selection. Statistics in Medicine, 24(24):3697–3714, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2005:MRP

[1879] Sarah Brown, Helen Thorpe, Kim Hawkins, and Julia Brown. Minimization-reducing predictability for multi-centre trials whilst retaining balance within centre. Statistics in Medicine, 24(24):3715–3727, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heritier:2005:DBR

[1880] Stephane Heritier, Val Gebski, and Avinesh Pillai. Dynamic balancing randomization in controlled clinical trials. Statistics in Medicine, 24(24): 3729–3741, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Winkens:2005:OTP

[1881] Bjorn Winkens, Hubert J. A. Schouten, Gerard J. P. van Breukelen, and Martijn P. F. Berger. Optimal time-points in clinical trials with linearly divergent treatment effects. Statistics in Medicine, 24(24):3743– 3756, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hendriks:2005:SSC

[1882] J. C. M. Hendriks, S. Teerenstra, J. P. E. Punt-Van der Zalm, A. M. M. Wetzels, J. R. Westphal, and G. F. Borm. Sample size calculations for a split-cluster, beta-binomial design in the assessment of toxicity. Statistics in Medicine, 24(24):3757–3772, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harkanen:2005:CSC

[1883] T. Härkänen, P. Knekt, E. Virtala, and O. Lindfors. A case study in comparing therapies involving informative drop-out, non-ignorable noncompliance and repeated measurements. Statistics in Medicine, 24(24): 3773–3787, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Legrand:2005:BAJ

[1884] Catherine Legrand, Vincent Ducrocq, Paul Janssen, Richard Sylvester, and Luc Duchateau. A Bayesian approach to jointly estimate centre and treatment by centre heterogeneity in a proportional hazards model. Statistics in Medicine, 24(24):3789–3804, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2005:EUE

[1885] Ian R. White, Stuart J. Pocock, and Duolao Wang. Eliciting and using expert opinions about influence of patient characteristics on treatment effects: a Bayesian analysis of the CHARM trials. Statistics in Medicine, 24(24):3805–3821, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abrams:2005:MAH

[1886] Keith R. Abrams, Clare L. Gillies, and Paul C. Lambert. Meta-analysis of heterogeneously reported trials assessing change from baseline. Statistics in Medicine, 24(24):3823–3844, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Minelli:2005:BIG

[1887] Cosetta Minelli, John R. Thompson, Keith R. Abrams, and Paul C. Lambert. Bayesian implementation of a genetic model-free approach to the meta-analysis of genetic association studies. Statistics in Medicine, 24(24):3845–3861, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giorgi:2005:ATD

[1888] Roch Giorgi and Joanny Gouvernet. Analysis of time-dependent covariates in a regressive relative survival model. Statistics in Medicine, 24 (24):3863–3870, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2005:AMC

[1889] Paul C. Lambert, Lucy K. Smith, David R. Jones, and Johannes L. Botha. Additive and multiplicative covariate regression models for relative survival incorporating fractional polynomials for time-dependent effects. Statistics in Medicine, 24(24):3871–3885, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teuff:2005:CCR

[1890] Gwenaël Le Teuff, Michal Abrahamowicz, Philippe Bolard, and Catherine Quantin. Comparison of Cox’s and relative survival models when estimating the effects of prognostic factors on disease-specific mortality: a simulation study under proportional excess hazards. Statistics in Medicine, 24(24):3887–3909, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stare:2005:GFR

[1891] Janez Stare, Maja Pohar, and Robin Henderson. Goodness of fit of relative survival models. Statistics in Medicine, 24(24):3911–3925, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Antolini:2005:TDD

[1892] Laura Antolini, Patrizia Boracchi, and Elia Biganzoli. A time-dependent discrimination index for survival data. Statistics in Medicine, 24(24): 3927–3944, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lorenzo:2005:AIC

[1893] Paola De Lorenzo, Robin Henderson, and Maria Grazia Valsecchi. Analysis of interval-censored longitudinal data with application to oncohaematology. Statistics in Medicine, 24(24):3945–3957, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gonzalez:2005:MIE

[1894] Juan R. González, Edsel A. Peña, and Elizabeth H. Slate. Modelling intervention effects after cancer relapses. Statistics in Medicine, 24(24): 3959–3975, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2005:BPH

[1895] Philippe Lambert and Paul H. C. Eilers. Bayesian proportional hazards model with time-varying regression coefficients: a penalized Poisson regression approach. Statistics in Medicine, 24(24):3977–3989, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sweeting:2005:BBC

[1896] Michael J. Sweeting, Daniela De Angelis, and Odd O. Aalen. Bayesian back-calculation using a multi-state model with application to HIV. Statistics in Medicine, 24(24):3991–4007, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reilly:2005:RUC

[1897] Marie Reilly, Anna Torrång, and Åsa Klint. Re-use of case-control data for analysis of new outcome variables. Statistics in Medicine, 24(24): 4009–4019, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carrasco:2005:CCC

[1898] Josep L. Carrasco and Lluis Jover. Concordance correlation coefficient applied to discrete data. Statistics in Medicine, 24(24):4021–4034, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hocine:2005:TIB

[1899] Mounia Hocine, Didier Guillemot, Pascale Tubert-Bitter, and Thierry Moreau. Testing independence between two Poisson-generated multinomial variables in case-series and cohort studies. Statistics in Medicine, 24 (24):4035–4044, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kivi:2005:AAN

[1900] Mårten Kivi, Anna L. V. Johansson, Agus Salim, Ylva Tindberg, and Marie Reilly. Accommodation of additional non-randomly sampled cases in a study of Helicobacter pylori infection in families. Statistics in Medicine, 24(24):4045–4054, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moerkerke:2005:PBM

[1901] Beatrijs Moerkerke, Els Goetghebeur, Kristel Van Steen, Simon Van Belle, and Veronique Cocquyt. Permutation based methods for comparing quality of life between observed treatments. Statistics in Medicine, 24 (24):4055–4066, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2005:A

[1902] Anonymous. Acknowledgement. Statistics in Medicine, 24(24):4067– 4072, December 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagostino:2006:E

[1903] Ralph B. D’agostino, Michael J. Campbell, and Joel B. Greenhouse. Editorial. Statistics in Medicine, 25(1):1–2, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Loke:2006:BDF

[1904] Yee-Chong Loke, Say-Beng Tan, YiYu Cai, and David Machin. A Bayesian dose finding design for dual endpoint phase I trials. Statistics in Medicine, 25(1):3–22, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:2006:RTP

[1905] Peter Bauer and Franz Koenig. The reassessment of trial perspectives from interim data — a critical view. Statistics in Medicine, 25(1):23–36, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2006:BDP

[1906] John Whitehead, Yinghui Zhou, John Stevens, Graham Blakey, Jim Price, and Joanna Leadbetter. Bayesian decision procedures for doseescalation based on evidence of undesirable events and therapeutic benefit. Statistics in Medicine, 25(1):37–53, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2006:CBA

[1907] Ying Kuen Cheung, Lurdes Y. T. Inoue, J. Kyle Wathen, and Peter F. Thall. Continuous Bayesian adaptive randomization based on event times with covariates. Statistics in Medicine, 25(1):55–70, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2439].

Salehi:2006:AML

[1908] M. Mohammad Salehi, Paul S. Levy, Mohammad A. Jamalzadeh, and K.-C. Chang. Adaptation of multiple logistic regression to a multiple inverse sampling design: application to the Isfahan Healthy Heart Program. Statistics in Medicine, 25(1):71–85, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gustafson:2006:CPB

[1909] Paul Gustafson and Sander Greenland. Curious phenomena in Bayesian adjustment for exposure misclassification. Statistics in Medicine, 25(1): 87–103, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2006:PMC

[1910] Kijoung Song and Robert C. Elston. A powerful method of combining measures of association and Hardy-Weinberg disequilibrium for finemapping in case-control studies. Statistics in Medicine, 25(1):105–126, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:2006:DCP

[1911] Patrick Royston, Douglas G. Altman, and Willi Sauerbrei. Dichotomizing continuous predictors in multiple regression: a bad idea. Statistics in Medicine, 25(1):127–141, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vonesh:2006:SPM

[1912] Edward F. Vonesh, Tom Greene, and Mark D. Schluchter. Shared parameter models for the joint analysis of longitudinal data and event times. Statistics in Medicine, 25(1):143–163, January 15, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Metcalfe:2006:IVE

[1913] Chris Metcalfe and Simon G. Thompson. The importance of varying the event generation process in simulation studies of statistical methods for recurrent events. Statistics in Medicine, 25(1):165–179, January 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2006:SMB

[1914] Ralph B. D’Agostino, Michael J. Campbell, and Joel B. Greenhouse. Surrogate markers: back to the future: Special papers for the 25th Anniversary of Statistics in Medicine. Statistics in Medicine, 25(2):181–182, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weir:2006:SEB

[1915] Christopher J. Weir and Rosalind J. Walley. Statistical evaluation of biomarkers as surrogate endpoints: a literature review. Statistics in Medicine, 25(2):183–203, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2327].

Alonso:2006:UAS

[1916] Ariel Alonso, Geert Molenberghs, Helena Geys, Marc Buyse, and Tony Vangeneugden. A unifying approach for surrogate marker validation based on Prentice’s criteria. Statistics in Medicine, 25(2):205–221, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2006:QIT

[1917] Yongming Qu and Michael Case. Quantifying the indirect treatment effect via surrogate markers. Statistics in Medicine, 25(2):223–231, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3384].

Barnes:2006:MIT

[1918] Sunni A. Barnes, Stacy R. Lindborg, and John W. Seaman, Jr. Multiple imputation techniques in small sample clinical trials. Statistics in Medicine, 25(2):233–245, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borghi:2006:CWH

[1919] E. Borghi, M. de Onis, C. Garza, J. Van den Broeck, E. A. Frongillo, L. Grummer-Strawn, S. Van Buuren, H. Pan, L. Molinari, R. Martorell, A. W. Onyango, J. C. Martines, and WHO Multicentre Growth Reference Study Group. Construction of the World Health Organization child growth standards: selection of methods for attained growth curves. Statistics in Medicine, 25(2):247–265, January 30, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paliwal:2006:EAS

[1920] Prashni Paliwal, Alan E. Gelfand, Linn Abraham, William Barlow, and Joann G. Elmore. Examining accuracy of screening mammography using an event order model. Statistics in Medicine, 25(2):267–283, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2006:BSP

[1921] Song Zhang, Dongchu Sun, Chong Z. He, and Mario Schootman. A Bayesian semi-parametric model for colorectal cancer incidences. Statistics in Medicine, 25(2):285–309, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Naus:2006:TSU

[1922] Joseph Naus and Sylvan Wallenstein. Temporal surveillance using scan statistics. Statistics in Medicine, 25(2):311–324, January 30, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2006:IRR

[1923] Byron J. Gajewski, Sarah Thompson, Nancy Dunton, Annette Becker, and Marcia Wrona. Inter-rater reliability of nursing home surveys: a Bayesian latent class approach. Statistics in Medicine, 25(2):325–344, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abellan:2006:ART

[1924] J. J. Abellán, C. Armero, D. Conesa, J. Pérez-Panadés, M. A. Martínez-Beneito, O. Zurriaga, M. J. García-Blasco, and H. Vanaclocha. Analysis of the renal transplant waiting list in the País Valencià (Spain). Statistics in Medicine, 25(2):345–358, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kay:2006:LEP

[1925] Richard Kay. Letter to the Editor: Phase specific analysis of herpes zoster associated pain data: a new statistical approach by R. B. Arani, S.-J. Soong, H. L. Weiss, M. J. Wood, P. A. Fiddian, J. W. Gnann and R. Whitley, Statistics in Medicine 2001; 20:2429–2439. Statistics in Medicine, 25(2):359–360, January 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [559].

Jeong:2006:RTC

[1926] Jong-Hyeon Jeong and Sin-Ho Jung. Rank tests for clustered survival data when dependent subunits are randomized. Statistics in Medicine, 25(3):361–373, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:2006:CPM

[1927] David M. Murray, Peter J. Hannan, Sherri P. Pals, Richard G. McCowen, William L. Baker, and Jonathan L. Blitstein. A comparison of permutation and mixed-model regression methods for the analysis of simulated data in the context of a group-randomized trial. Statistics in Medicine, 25(3):375–388, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhattacharya:2006:EPM

[1928] Jay Bhattacharya, Dana Goldman, and Daniel McCaffrey. Estimating probit models with self-selected treatments. Statistics in Medicine, 25 (3):389–413, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galbraith:2006:MDM

[1929] Sally Galbraith, Ian C. Marschner, and John Simes. Missing data methods for the assessment of surrogate outcomes and treatment mechanisms in clinical trial substudies. Statistics in Medicine, 25(3):415–431, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2006:EBD

[1930] John Whitehead, Yinghui Zhou, Adrian Mander, Sally Ritchie, Antony Sabin, and Andrew Wright. An evaluation of Bayesian designs for doseescalation studies in healthy volunteers. Statistics in Medicine, 25(3): 433–445, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:2006:MMC

[1931] May C. M. Wong, K. F. Lam, and Edward C. M. Lo. Multilevel modelling of clustered grouped survival data using Cox regression model: an application to ART dental restorations. Statistics in Medicine, 25 (3):447–457, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OMalley:2006:BMH

[1932] A. James O’Malley and Kelly H. Zou. Bayesian multivariate hierarchical transformation models for ROC analysis. Statistics in Medicine, 25(3): 459–479, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Obuchowski:2006:RTM

[1933] Nancy A. Obuchowski. An ROC-type measure of diagnostic accuracy when the gold standard is continuous-scale. Statistics in Medicine, 25 (3):481–493, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2006:SMA

[1934] Zuoheng Wang, Wei Hou, and Rongling Wu. A statistical model to analyse quantitative trait locus interactions for HIV dynamics from the virus and human genomes. Statistics in Medicine, 25(3):495–511, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:2006:MFR

[1935] Lloyd J. Edwards, Paul W. Stewart, James E. MacDougall, and Ronald W. Helms. A method for fitting regression splines with varying polynomial order in the linear mixed model. Statistics in Medicine, 25(3):513–527, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leuraud:2006:CSA

[1936] Klervi Leuraud and Jacques Benichou. A comparison of stratified and adjusted trend tests for binomial proportions. Statistics in Medicine, 25 (3):529–535, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rice:2006:BAM

[1937] Kenneth Rice. On Bayesian analysis of misclassified data from a matched case-control study with a validation sub-study by Gordon J. Prescott and Paul H. Garthwaite. Statistics in Medicine, 25(3):537–539, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:2006:AR

[1938] Gordon J. Prescott and Paul H. Garthwaite. Authors’ reply. Statistics in Medicine, 25(3):539–540, February 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2006:MWS

[1939] Ralph B. D’Agostino, Michael Campbell, and Joel Greenhouse. The Mann–Whitney statistic: continuous use and discovery: Special papers for the 25th Anniversary of Statistics in Medicine. Statistics in Medicine, 25(4):541–542, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:2006:CIEa

[1940] Robert G. Newcombe. Confidence intervals for an effect size measure based on the Mann–Whitney statistic. Part 1: general issues and tailarea-based methods. Statistics in Medicine, 25(4):543–557, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:2006:CIEb

[1941] Robert G. Newcombe. Confidence intervals for an effect size measure based on the Mann–Whitney statistic. Part 2: asymptotic methods and evaluation. Statistics in Medicine, 25(4):559–573, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2006:URC

[1942] Lyndia C. Brumback, Margaret S. Pepe, and Todd A. Alonzo. Using the ROC curve for gauging treatment effect in clinical trials. Statistics in Medicine, 25(4):575–590, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Acion:2006:PII

[1943] Laura Acion, John J. Peterson, Scott Temple, and Stephan Arndt. Probabilistic index: an intuitive non-parametric approach to measuring the size of treatment effects. Statistics in Medicine, 25(4):591–602, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cadarso-Suarez:2006:EMN

[1944] Carmen Cadarso-Suárez, Javier Roca-Pardiñas, and Adolfo Figueiras. Effect measures in non-parametric regression with interactions between continuous exposures. Statistics in Medicine, 25(4):603–621, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schisterman:2006:RAM

[1945] Enrique F. Schisterman, Benjamin Reiser, and David Faraggi. ROC analysis for markers with mass at zero. Statistics in Medicine, 25(4): 623–638, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2006:CIE

[1946] Jianrong Wu, Guoyong Jiang, and Wei Wei. Confidence intervals of effect size in randomized comparative parallel-group studies. Statistics in Medicine, 25(4):639–651, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ranyimbo:2006:EFN

[1947] A. O. Ranyimbo and L. Held. Estimation of the false negative fraction of a diagnostic kit through Bayesian regression model averaging. Statistics in Medicine, 25(4):653–667, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2006:CCD

[1948] Hsiao-Mei Wang, Michael P. Jones, and Barry E. Storer. Comparison of case-deletion diagnostic methods for Cox regression. Statistics in Medicine, 25(4):669–683, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2006:PTH

[1949] Kung-Jong Lui. Procedures for testing the homogeneity of relative difference in sparse data. Statistics in Medicine, 25(4):685–697, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2006:DUP

[1950] Young-Ku Choi, Wesley O. Johnson, and Mark C. Thurmond. Diagnosis using predictive probabilities without cut-offs. Statistics in Medicine, 25 (4):699–717, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2006:LEP

[1951] Anonymous. Letter to the Editor: A permutation test for inference in logistic regression with small- and moderate-sized data sets by D. M. Potter, Statistics in Medicine 2005; 24:693–708. Statistics in Medicine, 25 (4):719, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1637].

Willan:2006:C

[1952] Andrew R. Willan. Correction. Statistics in Medicine, 25(4):720, February 28, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2006:DDC

[1953] Andrew Lawson, Ronald Gangnon, and Dan Wartenberg. Developments in disease cluster detection. Statistics in Medicine, 25(5):721, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Assuncao:2006:FDA

[1954] R. Assunção, M. Costa, A. Tavares, and S. Ferreira. Fast detection of arbitrarily shaped disease clusters. Statistics in Medicine, 25(5):723–742, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duczmal:2006:WSS

[1955] Luiz Duczmal and David L. Buckeridge. A workflow spatial scan statistic. Statistics in Medicine, 25(5):743–754, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kleinman:2006:ESS

[1956] Ken Kleinman, Allyson Abrams, W. Katherine Yih, Richard Platt, and Martin Kulldorff. Evaluating spatial surveillance: detection of known outbreaks in real data. Statistics in Medicine, 25(5):755–769, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hossain:2006:CDD

[1957] Md. Monir Hossain and Andrew B. Lawson. Cluster detection diagnostics for small area health data: with reference to evaluation of local likelihood models. Statistics in Medicine, 25(5):771–786, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louie:2006:MDL

[1958] Mary M. Louie and Eric D. Kolaczyk. Multiscale detection of localized anomalous structure in aggregate disease incidence data. Statistics in Medicine, 25(5):787–810, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rogerson:2006:SMD

[1959] Peter A. Rogerson. Statistical methods for the detection of spatial clustering in case-control data. Statistics in Medicine, 25(5):811–823, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2006:LBT

[1960] Changhong Song and Martin Kulldorff. Likelihood based tests for spatial randomness. Statistics in Medicine, 25(5):825–839, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Takahashi:2006:EPC

[1961] Kunihiko Takahashi and Toshiro Tango. An extended power of cluster detection tests. Statistics in Medicine, 25(5):841–852, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waller:2006:GPS

[1962] Lance A. Waller, Elizabeth G. Hill, and Rose Anne Rudd. The geography of power: statistical performance of tests of clusters and clustering in heterogeneous populations. Statistics in Medicine, 25(5):853–865, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2006:CMS

[1963] Ping Yan and Murray K. Clayton. A cluster model for space-time disease counts. Statistics in Medicine, 25(5):867–881, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gangnon:2006:IPC

[1964] Ronald E. Gangnon. Impact of prior choice on local Bayes factors for cluster detection. Statistics in Medicine, 25(5):883–895, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:2006:DCD

[1965] Andrew B. Lawson. Disease cluster detection: a critique and a Bayesian proposal. Statistics in Medicine, 25(5):897–916, March 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jennison:2006:EGS

[1966] Christopher Jennison and Bruce W. Turnbull. Efficient group sequential designs when there are several effect sizes under consideration. Statistics in Medicine, 25(6):917–932, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:2006:GSS

[1967] Weichung Joe Shih. Group sequential, sample size re-estimation and two-stage adaptive designs in clinical trials: a comparison. Statistics in Medicine, 25(6):933–941, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelfand:2006:BRU

[1968] Lois A. Gelfand, Daniel R. Strunk, Xin M. Tu, Ronald E. S. Noble, and Robert J. DeRubeis. Bias resulting from the use of ‘assay sensitivity’ as an inclusion criterion for meta-analysis. Statistics in Medicine, 25(6): 943–955, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giraudeau:2006:MMS

[1969] Bruno Giraudeau. Model mis-specification and overestimation of the intraclass correlation coefficient in cluster randomized trials. Statistics in Medicine, 25(6):957–964, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barry:2006:ICR

[1970] Daniel Barry and Victoria Livingstone. The investigation and correction of recall bias for an ordinal response in a case-control study. Statistics in Medicine, 25(6):965–975, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holford:2006:AFA

[1971] Theodore R. Holford. Approaches to fitting age-period-cohort models with unequal intervals. Statistics in Medicine, 25(6):977–993, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zucker:2006:IFS

[1972] David M. Zucker and Song Yang. Inference for a family of survival models encompassing the proportional hazards and proportional odds models. Statistics in Medicine, 25(6):995–1014, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klein:2006:MCR

[1973] John P. Klein. Modelling competing risks in cancer studies. Statistics in Medicine, 25(6):1015–1034, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Isaman:2006:DSD

[1974] Deanna J. M. Isaman, William H. Herman, and Morton B. Brown. A discrete-state discrete-time model using indirect observation. Statistics in Medicine, 25(6):1035–1049, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2006:ARC

[1975] Wenqing He, Shelley B. Bull, Nalan Gokgoz, Irene Andrulis, and Jay Wunder. Application of reliability coefficients in cDNA microarray data analysis. Statistics in Medicine, 25(6):1051–1066, March 30, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riffenburgh:2006:SGC

[1976] R. H. Riffenburgh and K. M. Cummins. A simple and general changepoint identifier. Statistics in Medicine, 25(6):1067–1077, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:2006:EVE

[1977] Niels G. Becker, Tom Britton, and Philip D. O’Neill. Estimating vaccine effects from studies of outbreaks in household pairs. Statistics in Medicine, 25(6):1079–1093, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2006:C

[1978] Lingqi Tang. Correction. Statistics in Medicine, 25(6):1095, March 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agostino:2006:ENI

[1979] Ralph B. D’Agostino, Sr., Michael Campbell, and Joel Greenhouse. Editorial: Non-inferiority trials: continued advancements in concepts and methodology (special papers for the 25th Anniversary of Statistics in Medicine 25 (7)). Statistics in Medicine, 25(7):1097–1099, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2006:NIM

[1980] Shein-Chung Chow and Jun Shao. On non-inferiority margin and statistical tests in active control trials. Statistics in Medicine, 25(7):1101–1113, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laster:2006:NIT

[1981] Larry L. Laster, Mary F. Johnson, and Mitchell L. Kotler. Non-inferiority trials: the ‘at least as good as’ criterion with dichotomous data. Statistics in Medicine, 25(7):1115–1130, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dilba:2006:PSS

[1982] Gemechis Dilba, Frank Bretz, Ludwig A. Hothorn, and Volker Guiard. Power and sample size computations in simultaneous tests for noninferiority based on relative margins. Statistics in Medicine, 25(7):1131– 1147, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lai:2006:MHP

[1983] Tze Leung Lai, Mei-Chiung Shih, and Guangrui Zhu. Modified Haybittle–Peto group sequential designs for testing superiority and noninferiority hypotheses in clinical trials. Statistics in Medicine, 25(7): 1149–1167, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sanchez:2006:CAP

[1984] M. Matilde Sanchez and Xun Chen. Choosing the analysis population in non-inferiority studies: per protocol or intent-to-treat. Statistics in Medicine, 25(7):1169–1181, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheng:2006:ENC

[1985] Dan Sheng and Mimi Y. Kim. The effects of non-compliance on intentto-treat analysis of equivalence trials. Statistics in Medicine, 25(7):1183– 1199, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freitag:2006:NPA

[1986] G. Freitag, S. Lange, and A. Munk. Non-parametric assessment of noninferiority with censored data. Statistics in Medicine, 25(7):1201–1217, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2006:TEN

[1987] Jen-Pei Liu, Mi-Chia Ma, Chin yu Wu, and Jia-Yen Tai. Tests of equivalence and non-inferiority for diagnostic accuracy based on the paired areas under ROC curves. Statistics in Medicine, 25(7):1219–1238, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2006:CDT

[1988] Yuqing Yang and Zhezhen Jin. Combining dependent tests to compare the diagnostic accuracies — a non-parametric approach. Statistics in Medicine, 25(7):1239–1250, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiong:2006:MED

[1989] Chengjie Xiong, Gerald van Belle, J. Philip Miller, and John C. Morris. Measuring and estimating diagnostic accuracy when there are three ordinal diagnostic groups. Statistics in Medicine, 25(7):1251–1273, April 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2006:CIE

[1990] G. Frank Liu, Junyuan Wang, Kenneth Liu, and Duane B. Snavely. Confidence intervals for an exposure adjusted incidence rate difference with applications to clinical trials. Statistics in Medicine, 25(8):1275–1286, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louie:2006:MMD

[1991] Mary M. Louie and Eric D. Kolaczyk. A multiscale method for disease mapping in spatial epidemiology. Statistics in Medicine, 25(8):1287–1306, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gueorguieva:2006:JAR

[1992] R. V. Gueorguieva and G. Sanacora. Joint analysis of repeatedly observed continuous and ordinal measures of disease severity. Statistics in Medicine, 25(8):1307–1322, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2006:SAC

[1993] Yannan Jiang, Alastair J. Scott, and Chris J. Wild. Secondary analysis of case-control data. Statistics in Medicine, 25(8):1323–1339, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noh:2006:DFM

[1994] Maengseok Noh, Il Do Ha, and Youngjo Lee. Dispersion frailty models and HGLMs. Statistics in Medicine, 25(8):1341–1354, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

York:2006:MAR

[1995] Timothy P. York, Lindon J. Eaves, and Edwin J. C. G. van den Oord. Multivariate adaptive regression splines: a powerful method for detecting disease-risk relationship differences among subgroups. Statistics in Medicine, 25(8):1355–1367, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2006:QRM

[1996] Ying Wei, Anneli Pere, Roger Koenker, and Xuming He. Quantile regression methods for reference growth charts. Statistics in Medicine, 25(8): 1369–1382, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2006:PSL

[1997] Bo Hu, Mari Palta, and Jun Shao. Properties of R^{2} statistics for logistic regression. Statistics in Medicine, 25(8):1383–1395, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2006:RAL

[1998] Tsung I. Lin and Jack C. Lee. A robust approach to t linear mixed models applied to multiple sclerosis data. Statistics in Medicine, 25(8): 1397–1412, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:2006:SMA

[1999] Kerrie P. Nelson and Brian G. Leroux. Statistical models for autocorrelated count data. Statistics in Medicine, 25(8):1413–1430, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2006:LES

[2000] Anonymous. Letter to the Editor: Survival analysis for recurrent event data: an application to childhood infectious diseases by P. J. Kelly and L. L. Lim, Statistics in Medicine 2000; 19:13–33. Statistics in Medicine, 25(8):1431–1433, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [21].

Kelly:2006:AR

[2001] Patrick Kelly and Lynette Lim. Authors’ reply. Statistics in Medicine, 25(8):1433, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2006:LEA

[2002] Anonymous. Letter to the Editor: Analysis of clinical data with breached blindness by Shein-Chung Chow and Jun Shao, Statistics in Medicine 2004; 23:1185–1193. Statistics in Medicine, 25(8):1434–1437, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1365, 1645].

Shao:2006:AR

[2003] Jun Shao and Shein-Chung Chow. Authors’ reply. Statistics in Medicine, 25(8):1437–1438, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Esterman:2006:BRE

[2004] Adrian J. Esterman. Book review: 1. Epidemiology, study design and data analysis (2nd edn). Mark Woodward, Chapman and Hall/CRC, Boca Raton, 2005. No. of pages: xxii + 849. Price: $79.95. ISBN: 1-58488-415-0. Statistics in Medicine, 25(8):1439–1440, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Putt:2006:BRL

[2005] Mary E. Putt. Book review: 2. Linear models with R. Julian J. Faraway, Chapman and Hall/CRC, Boca Raton, 2005. No of pages: x + 229. Price: $69.95. ISBN: 1-58488-425-8. Statistics in Medicine, 25(8):1440–1441, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunne:2006:BRA

[2006] Adrian Dunne. Book review: 3. Applied statistics in toxicology and pharmacology. Katsumi Kobayashi and K. Sadasivan Pillai, Science Publishers Inc, Enfield, Hew Hampshire, 2003. No. of pages: 111. Price: $45.00. ISBN: 1-57808-304-4. Statistics in Medicine, 25(8):1441–1442, April 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2006:CII

[2007] Rebecca M. Turner, Rumana Z. Omar, and Simon G. Thompson. Constructing intervals for the intracluster correlation coefficient using Bayesian modelling, and application in cluster randomized trials. Statistics in Medicine, 25(9):1443–1456, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Formann:2006:MAL

[2008] Anton K. Formann. Mixture analysis of longitudinal binary data. Statistics in Medicine, 25(9):1457–1469, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

laCruz-Mesia:2006:NLR

[2009] Rolando De la Cruz-Mesía and Guillermo Marshall. Non-linear random effects models with continuous time autoregressive errors: a Bayesian approach. Statistics in Medicine, 25(9):1471–1484, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunning:2006:MIC

[2010] Andrew J. Dunning. A model for immunological correlates of protection. Statistics in Medicine, 25(9):1485–1497, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Houseman:2006:PQD

[2011] E. Andrés Houseman and Donald K. Milton. Partial questionnaire designs, questionnaire non-response, and attributable fraction: applications to adult onset asthma. Statistics in Medicine, 25(9):1499–1519, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:2006:AHT

[2012] Jun mo Nam. Assessment on homogeneity tests for kappa statistics under equal prevalence across studies in reliability. Statistics in Medicine, 25(9): 1521–1531, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2006:PMM

[2013] Huiling Li and Daniel F. Heitjan. A pattern-mixture model for the analysis of censored quality-of-life data. Statistics in Medicine, 25(9): 1533–1546, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saito:2006:ENS

[2014] Yuki Saito, Takashi Sozu, Chikuma Hamada, and Isao Yoshimura. Effective number of subjects and number of raters for inter-rater reliability studies. Statistics in Medicine, 25(9):1547–1560, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Indurkhya:2006:UPS

[2015] Alka Indurkhya, Nandita Mitra, and Deborah Schrag. Using propensity scores to estimate the cost-effectiveness of medical therapies. Statistics in Medicine, 25(9):1561–1576, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shkedy:2006:MAD

[2016] Z. Shkedy, M. Aerts, G. Molenberghs, Ph. Beutels, and P. Van Damme. Modelling age-dependent force of infection from prevalence data using fractional polynomials. Statistics in Medicine, 25(9):1577–1591, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2006:AMD

[2017] Fei Gao, Kee-Seng Chia, Ingela Krantz, Per Nordin, and David Machin. On the application of the von Mises distribution and angular regression methods to investigate the seasonality of disease onset. Statistics in Medicine, 25(9):1593–1618, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2006:ETE

[2018] Ian R. White. Estimating treatment effects in randomized trials with treatment switching. Statistics in Medicine, 25(9):1619–1622, May 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2006:INC

[2019] David Brown, Peter Volkers, and Simon Day. An introductory note to CHMP guidelines: choice of the non-inferiority margin and data monitoring committees. Statistics in Medicine, 25(10):1623–1627, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2245].

Anonymous:2006:CMPa

[2020] Anonymous. Committee for medicinal products for human use (CHMP) guideline on the choice of the non-inferiority margin. Statistics in Medicine, 25(10):1628–1638, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2006:CMPb

[2021] Anonymous. Committee for medicinal products for human use (CHMP) guideline on data monitoring committees. Statistics in Medicine, 25(10): 1639–1645, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schill:2006:MDP

[2022] W. Schill and P. Wild. Minmax designs for planning the second phase in a two-phase case-control study. Statistics in Medicine, 25(10):1646–1659, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiang:2006:STZ

[2023] Liming Xiang, Andy H. Lee, Kelvin K. W. Yau, and Geoffrey J. McLachlan. A score test for zero-inflation in correlated count data. Statistics in Medicine, 25(10):1660–1671, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olesen:2006:CSU

[2024] Anne Vingaard Olesen and Erik Thorlund Parner. Correcting for selection using frailty models. Statistics in Medicine, 25(10):1672–1684, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2006:MCG

[2025] Xin-Yuan Song and Sik-Yum Lee. Model comparison of generalized linear mixed models. Statistics in Medicine, 25(10):1685–1698, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fan:2006:DTD

[2026] Shenghua K. Fan and You-Gan Wang. Decision-theoretic designs for dose-finding clinical trials with multiple outcomes. Statistics in Medicine, 25(10):1699–1714, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2006:NCR

[2027] Anuradha Roy. A new classification rule for incomplete doubly multivariate data using mixed effects model with performance comparisons on the imputed data. Statistics in Medicine, 25(10):1715–1728, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2006:ITD

[2028] Yali Liu and Bruce A. Craig. Incorporating time-dependent covariates in survival analysis using the LVAR method. Statistics in Medicine, 25(10): 1729–1740, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldfarb:2006:ESM

[2029] Bernard Goldfarb and Catherine M. Pardoux. Exploring series of multivariate censored temporal data through fuzzy coding and correspondence analysis. Statistics in Medicine, 25(10):1741–1750, May 30, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Asgharian:2006:CSI

[2030] Masoud Asgharian, David B. Wolfson, and Xun Zhang. Checking stationarity of the incidence rate using prevalent cohort survival data. Statistics in Medicine, 25(10):1751–1767, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitaker:2006:TBS

[2031] Heather J. Whitaker, C. Paddy Farrington, Bart Spiessens, and Patrick Musonda. Tutorial in biostatistics: the self-controlled case series method. Statistics in Medicine, 25(10):1768–1797, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Singer:2006:LES

[2032] Julia Singer. Letter to the Editor: Simple improved confidence intervals for comparing matched proportions by Alan Agresti and Yongyi Min, Statistics in Medicine 2005; 24:729–740. Statistics in Medicine, 25(10): 1798–1799, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1639].

Ganju:2006:LED

[2033] Jitendra Ganju. Letter to the Editor: Diagnostics for assumptions in moderate to large simple clinical trials: do they really help? by Jonathan J. Shuster, Statistics in Medicine 2005; 24:2431–2438. Statistics in Medicine, 25(10):1799–1800, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1775].

Shrout:2006:BRM

[2034] Patrick E. Shrout and Petra Macaskill. Book review: 1. Measures of interobserver agreement. Mohamed M. Shoukri, Chapman and Hall/ CRC, Boca Raton, 2004. Pages: 152. Price: $79.95. ISBN: 1-58488-321-9. Statistics in Medicine, 25(10):1801–1802, May 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sebastiani:2006:BDMa

[2035] Paola Sebastiani, Kenneth D. Mandl, Peter Szolovits, Isaac S. Kohane, and Marco F. Ramoni. A Bayesian dynamic model for influenza surveillance. Statistics in Medicine, 25(11):1803–1816, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stoto:2006:BDM

[2036] Michael A. Stoto. A Bayesian dynamic model for influenza surveillance by sebastiani et al., Statistics in Medicine (this issue). Statistics in Medicine, 25(11):1817–1818, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2006:BDM

[2037] V. T. Farewell and D. J. Spiegelhalter. A Bayesian dynamic model for influenza surveillance. Statistics in Medicine, 25(11):1819–1820, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fienberg:2006:BDM

[2038] Stephen E. Fienberg and Galit Shmueli. A Bayesian dynamics model for influenza surveillance. Statistics in Medicine, 25(11):1821–1822, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sebastiani:2006:BDMb

[2039] Paola Sebastiani, Kenneth D. Mandl, Peter Szolovits, Isaac S. Kohane, and Marco F. Ramoni. A Bayesian dynamic model for influenza surveillance. Statistics in Medicine, 25(11):1823–1825, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chan:2006:MSD

[2040] Jennifer S. K. Chan, Philip L. H. Yu, Yeh Lam, and Alvin P. K. Ho. Modelling SARS data using threshold geometric process. Statistics in Medicine, 25(11):1826–1839, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chowell:2006:MTD

[2041] G. Chowell, E. Shim, F. Brauer, P. Diaz-Dueñas, J. M. Hyman, and C. Castillo-Chavez. Modelling the transmission dynamics of acute haemorrhagic conjunctivitis: application to the 2003 outbreak in Mexico. Statistics in Medicine, 25(11):1840–1857, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:2006:BHA

[2042] Timothy D. Johnson and Valen E. Johnson. A Bayesian hierarchical approach to multirater correlated ROC analysis. Statistics in Medicine, 25(11):1858–1871, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2006:MTB

[2043] Debashis Ghosh. Modelling tumour biology-progression relationships in screening trials. Statistics in Medicine, 25(11):1872–1884, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karlis:2006:BAD

[2044] D. Karlis and I. Ntzoufras. Bayesian analysis of the differences of count data. Statistics in Medicine, 25(11):1885–1905, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosychuk:2006:CVE

[2045] R. J. Rosychuk, X. Sheng, and J. L. Stuber. Comparison of variance estimation approaches in a two-state Markov model for longitudinal data with misclassification. Statistics in Medicine, 25(11):1906–1921, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2006:DLE

[2046] Charles C. Lin and Norman J. Johnson. Decomposition of life expectancy and expected life-years lost by disease. Statistics in Medicine, 25(11): 1922–1936, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Resche-Rigon:2006:LIS

[2047] Matthieu Resche-Rigon and Sylvie Chevret. Local influence for the subdistribution of a competing risk. Statistics in Medicine, 25(11):1937– 1947, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2006:MRE

[2048] Hong-Bin Fang, Guo-Liang Tian, Xiaoping Xiong, and Ming Tan. A multivariate random-effects model with restricted parameters: application to assessing radiation therapy for brain tumours. Statistics in Medicine, 25 (11):1948–1959, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demiris:2006:BES

[2049] N. Demiris and L. D. Sharples. Bayesian evidence synthesis to extrapolate survival estimates in cost-effectiveness studies. Statistics in Medicine, 25(11):1960–1975, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2006:LEE

[2050] Byron J. Gajewski, Greg Petroski, Sarah Thompson, Nancy Dunton, Marcia Wrona, Annette Becker, and Valorie Coffland. Letter to the Editor: The effect of provider-level ascertainment bias on profiling nursing homes by J. Roy and V. Mor, Statistics in Medicine 2005; 24 (23):3609– 3629. Statistics in Medicine, 25(11):1976–1977, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:2006:AR

[2051] Jason Roy and Vincent Mor. Authors’ reply. Statistics in Medicine, 25(11):1977–1978, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bender:2006:LEG

[2052] Ralf Bender, Thomas Augustin, and Maria Blettner. Letter to the Editor: Generating survival times to simulate Cox proportional hazards models by Ralf Bender, Thomas Augustin and Maria Blettner, Statistics in Medicine 2005; 24:1713–1723. Statistics in Medicine, 25(11): 1978–1979, June 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1720].

Small:2006:REL

[2053] Dylan S. Small, Thomas R. Ten Have, Marshall M. Joffe, and Jing Cheng. Random effects logistic models for analysing efficacy of a longitudinal randomized treatment with non-adherence. Statistics in Medicine, 25 (12):1981–2007, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2006:IWS

[2054] Lili Tian. Inferences on the within-subject coefficient of variation. Statistics in Medicine, 25(12):2008–2017, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2006:BID

[2055] Anastasia Ivanova and Kai Wang. Bivariate isotonic design for dosefinding with ordered groups. Statistics in Medicine, 25(12):2018–2026, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2006:MBA

[2056] Weili He, Jun Liu, Bruce Binkowitz, and Hui Quan. A model-based approach in the estimation of the maximum tolerated dose in phase I cancer clinical trials. Statistics in Medicine, 25(12):2027–2042, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:2006:EDS

[2057] Lynn Wei. An efficient design for a study comparing two drugs, their combination and placebo. Statistics in Medicine, 25(12):2043–2058, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2006:SIA

[2058] Debashis Ghosh. Semiparametric inferences for association with semicompeting risks data. Statistics in Medicine, 25(12):2059–2070, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Braun:2006:GTC

[2059] Thomas M. Braun. Generalizing the TITE–CRM to adapt for earlyand late-onset toxicities. Statistics in Medicine, 25(12):2071–2083, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2006:CPS

[2060] Peter C. Austin and Muhammad M. Mamdani. A comparison of propensity score methods: a case-study estimating the effectiveness of post-AMI statin use. Statistics in Medicine, 25(12):2084–2106, June 30, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2006:AIL

[2061] Xiaoming Li, Devan V. Mehrotra, and John Barnard. Analysis of incomplete longitudinal binary data using multiple imputation. Statistics in Medicine, 25(12):2107–2124, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2006:IEM

[2062] Jianrong Wu, A. C. M. Wong, and Wei Wei. Interval estimation of the mean response in a log-regression model. Statistics in Medicine, 25(12): 2125–2135, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2006:IEI

[2063] Christopher Jackson, Nicky Best, and Sylvia Richardson. Improving ecological inference using individual-level data. Statistics in Medicine, 25(12):2136–2159, June 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2006:MPP

[2064] Vern Farewell, Tony Johnson, and Peter Armitage. ‘A memorandum on the Present Position and Prospects of Medical Statistics and Epidemiology’ by Major Greenwood. Statistics in Medicine, 25(13):2161–2177, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2006:BPA

[2065] Alexei Dmitrienko and Ming-Dauh Wang. Bayesian predictive approach to interim monitoring in clinical trials. Statistics in Medicine, 25(13): 2178–2195, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sooriyarachchi:2006:SAR

[2066] Marina Roshini Sooriyarachchi, John Whitehead, Anne Whitehead, and Kim Bolland. The sequential analysis of repeated binary responses: a score test for the case of three time points. Statistics in Medicine, 25(13): 2196–2214, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2006:CPV

[2067] Wenquan Wang, Charles S. Davis, and Seng-Jaw Soong. Comparison of predictive values of two diagnostic tests from the same sample of subjects using weighted least squares. Statistics in Medicine, 25(13):2215–2229, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hill:2006:IET

[2068] Jennifer Hill and Jerome P. Reiter. Interval estimation for treatment effects using propensity score matching. Statistics in Medicine, 25(13): 2230–2256, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2006:CCM

[2069] Liansheng Wang and Abba M. Krieger. Causal conclusions are most sensitive to unobserved binary covariates. Statistics in Medicine, 25(13): 2257–2271, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacobusse:2006:ISD

[2070] Gert Jacobusse, Stef van Buuren, and Paul H. Verkerk. An interval scale for development of children aged 0–2 years. Statistics in Medicine, 25(13):2272–2283, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:2006:LAN

[2071] James F. Troendle and Kai F. Yu. Likelihood approaches to the nonparametric two-sample problem for right-censored data. Statistics in Medicine, 25(13):2284–2298, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pietrantonj:2006:FFT

[2072] Carlo Di Pietrantonj. Four-fold table cell frequencies imputation in meta analysis. Statistics in Medicine, 25(13):2299–2322, July 15, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baharith:2006:CAS

[2073] Lamya A. Baharith, Abeer Al-Khouli, and Gillian M. Raab. Cytotoxic assays for screening anticancer agents. Statistics in Medicine, 25(13): 2323–2339, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:2006:ETC

[2074] M. S. Pepe. Evaluating technologies for classification and prediction in medicine. Statistics in Medicine, 25(13):2340, July 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:2006:SAL

[2075] Simon G. Thompson. The Second Armitage Lecture. Statistics in Medicine, 25(14):2341–2342, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:2006:EHA

[2076] Niels Keiding. Event history analysis and the cross-section. Statistics in Medicine, 25(14):2343–2364, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2006:ADF

[2077] Wei Zhang, Daniel J. Sargent, and Sumithra Mandrekar. An adaptive dose-finding design incorporating both toxicity and efficacy. Statistics in Medicine, 25(14):2365–2383, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mallinckrodt:2006:COA

[2078] Craig H. Mallinckrodt, Michael J. Detke, Christopher J. Kaiser, John G. Watkin, Geert Molenberghs, and Raymond J. Carroll. Comparing onset of antidepressant action using a repeated measures approach and a traditional assessment schedule. Statistics in Medicine, 25(14):2384–2397, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gallop:2006:MEM

[2079] Robert J. Gallop, Thomas R. Ten Have, and Paul Crits-Christoph. A mixed effects Markov model for repeated binary outcomes with nonignorable dropout. Statistics in Medicine, 25(14):2398–2426, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2006:CMR

[2080] Changyu Shen and Lisa Weissfeld. A copula model for repeated measurements with non-ignorable non-monotone missing outcome. Statistics in Medicine, 25(14):2427–2440, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:2006:EHH

[2081] Kam-Fai Wong, Wei-Yann Tsai, and Louise Kuhn. Estimating HIV hazard rates from cross-sectional HIV prevalence data. Statistics in Medicine, 25(14):2441–2449, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chao:2006:SCM

[2082] Edward C. Chao. Structured correlation in models for clustered data. Statistics in Medicine, 25(14):2450–2468, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganguli:2006:AMG

[2083] B. Ganguli and M. P. Wand. Additive models for geo-referenced failure time data. Statistics in Medicine, 25(14):2469–2482, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roca-Pardinas:2006:BBM

[2084] Javier Roca-Pardiñas, Carmen Cadarso-Suárez, Verónica Nácher, and Carlos Acuña. Bootstrap-based methods for testing factor-by-curve interactions in generalized additive models: assessing prefrontal cortex neural activity related to decision-making. Statistics in Medicine, 25(14): 2483–2501, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hens:2006:MSI

[2085] N. Hens, M. Aerts, and G. Molenberghs. Model selection for incomplete and design-based samples. Statistics in Medicine, 25(14):2502–2520, July 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barthel:2006:ESS

[2086] F. M.-S. Barthel, A. Babiker, P. Royston, and M. K. B. Parmar. Evaluation of sample size and power for multi-arm survival trials allowing for non-uniform accrual, non-proportional hazards, loss to follow-up and cross-over. Statistics in Medicine, 25(15):2521–2542, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heller:2006:PCP

[2087] Glenn Heller. Power calculations for preclinical studies using a K-sample rank test and the Lehmann alternative hypothesis. Statistics in Medicine, 25(15):2543–2553, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2006:BSS

[2088] Byron J. Gajewski and Matthew S. Mayo. Bayesian sample size calculations in phase II clinical trials using a mixture of informative priors. Statistics in Medicine, 25(15):2554–2566, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Putt:2006:PDC

[2089] Mary E. Putt. Power to detect clinically relevant carry-over in a series of cross-over studies. Statistics in Medicine, 25(15):2567–2586, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:2006:PAC

[2090] X. M. Tu, J. Kowalski, P. Crits-Christoph, and R. Gallop. Power analyses for correlations from clustered study designs. Statistics in Medicine, 25 (15):2587–2606, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moerbeek:2006:PMC

[2091] Mirjam Moerbeek. Power and money in cluster randomized trials: when is it worth measuring a covariate? Statistics in Medicine, 25(15):2607– 2617, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Musonda:2006:SSS

[2092] Patrick Musonda, C. Paddy Farrington, and Heather J. Whitaker. Sample sizes for self-controlled case series studies. Statistics in Medicine, 25 (15):2618–2631, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bukszar:2006:AEP

[2093] József Bukszár and Edwin J. C. G. van den Oord. Accurate and efficient power calculations for 2 × m tables in unmatched case-control designs. Statistics in Medicine, 25(15):2632–2646, August 15, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2006:SSS

[2094] Haitao Chu, Lei Nie, and Stephen R. Cole. Sample size and statistical power assessing the effect of interventions in the context of mixture distributions with detection limits. Statistics in Medicine, 25(15):2647– 2657, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Branscum:2006:SSC

[2095] Adam J. Branscum, Wesley O. Johnson, and Ian A. Gardner. Sample size calculations for disease freedom and prevalence estimation surveys. Statistics in Medicine, 25(15):2658–2674, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2006:SSE

[2096] Yan D. Zhao. Sample size estimation for the van Elteren test — a stratified Wilcoxon–Mann–Whitney test. Statistics in Medicine, 25(15):2675– 2687, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2006:PST

[2097] Dan Jackson. The power of the standard test for the presence of heterogeneity in meta-analysis. Statistics in Medicine, 25(15):2688–2699, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sweeting:2006:C

[2098] Michael Sweeting, Alexander Sutton, and Paul Lambert. Correction. Statistics in Medicine, 25(15):2700, August 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dragalin:2006:DMC

[2099] Vladimir Dragalin and Valerii Fedorov. Design of multi-centre trials with binary response. Statistics in Medicine, 25(16):2701–2719, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Localio:2006:LRC

[2100] A. Russell Localio, Jesse A. Berlin, and Thomas R. Ten Have. Longitudinal and repeated cross-sectional cluster-randomization designs using mixed effects regression for binary outcomes: bias and coverage of frequentist and Bayesian methods. Statistics in Medicine, 25(16):2720– 2736, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2006:IAR

[2101] Xiao-Hua Zhou and Sierra M. Li. ITT analysis of randomized encouragement design studies with missing data. Statistics in Medicine, 25 (16):2737–2761, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marcus:2006:RBB

[2102] Sue M. Marcus, Jack M. Gorman, Xin Tu, Robert D. Gibbons, David H. Barlow, Scott W. Woods, and M. Katharine Shear. Rater bias in a blinded randomized placebo-controlled psychiatry trial. Statistics in Medicine, 25(16):2762–2770, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2006:SIC

[2103] Man-Lai Tang, Nian-Sheng Tang, and Bernard Rosner. Statistical inference for correlated data in ophthalmologic studies. Statistics in Medicine, 25(16):2771–2783, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parzen:2006:PLM

[2104] Michael Parzen, Stuart R. Lipsitz, Garrett M. Fitzmaurice, Joseph G. Ibrahim, and Andrea Troxel. Pseudo-likelihood methods for longitudinal binary data with non-ignorable missing responses and covariates. Statistics in Medicine, 25(16):2784–2796, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Balakrishnan:2006:GGF

[2105] N. Balakrishnan and Yingwei Peng. Generalized gamma frailty model. Statistics in Medicine, 25(16):2797–2816, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2006:NLR

[2106] Guillermo Marshall, Rolando De la Cruz-Mesía, Anna E. Barón, James H. Rutledge, and Gary O. Zerbe. Non-linear random effects model for multivariate responses with missing data. Statistics in Medicine, 25 (16):2817–2830, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perperoglou:2006:RRH

[2107] Aris Perperoglou, Saskia le Cessie, and Hans C. van Houwelingen. Reduced-rank hazard regression for modelling non-proportional hazards. Statistics in Medicine, 25(16):2831–2845, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsodikov:2006:PMP

[2108] A. Tsodikov, A. Szabo, and J. Wegelin. A population model of prostate cancer incidence. Statistics in Medicine, 25(16):2846–2866, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reda:2006:BRS

[2109] Domenic J. Reda and Petra Macaskill. Book review: 1. Statistical aspects of the design and analysis of clinical trials (revised edn). Brian S. Everitt and Andrew Pickles, Imperial College Press, London, 2004. No. of pages: xiii + 323. Price: $78.00, £48.00. ISBN: 1-86094-441-8. Statistics in Medicine, 25(16):2867–2868, August 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

London:2006:RRC

[2110] Alex John London. Reasonable risks in clinical research: a critique and a proposal for the integrative approach. Statistics in Medicine, 25(17): 2869–2885, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Basu:2006:ETB

[2111] Anirban Basu, Herbert Y. Meltzer, and Vanja Dukic. Estimating transitions between symptom severity states over time in schizophrenia: a Bayesian meta-analytic approach. Statistics in Medicine, 25(17):2886– 2910, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2006:IPB

[2112] Dan Jackson. The implications of publication bias for meta-analysis’ other parameter. Statistics in Medicine, 25(17):2911–2921, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:2006:CPC

[2113] Edward L. Korn and Boris Freidlin. Conditional power calculations for clinical trials with historical controls. Statistics in Medicine, 25(17): 2922–2931, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frigyesi:2006:EPO

[2114] Attila Frigyesi and Ola Hössjer. Estimating the parameters of the operational model of pharmacological agonism. Statistics in Medicine, 25(17): 2932–2945, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2006:MIM

[2115] Binbing Yu and Ram C. Tiwari. Multiple imputation methods for modelling relative survival data. Statistics in Medicine, 25(17):2946–2955, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2006:BSR

[2116] C. Chen and K. Chaloner. A Bayesian stopping rule for a single arm study: With a case study of stem cell transplantation. Statistics in Medicine, 25(17):2956–2966, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinavaara:2006:PLC

[2117] Sirpa Heinävaara and Timo Hakulinen. Predicting the lung cancer burden: accounting for selection of the patients with respect to general population mortality. Statistics in Medicine, 25(17):2967–2980, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paliwal:2006:EMD

[2118] Prashni Paliwal and Alan E. Gelfand. Estimating measures of diagnostic accuracy when some covariate information is missing. Statistics in Medicine, 25(17):2981–2993, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baughman:2006:MMA

[2119] Andrew L. Baughman, Kristine M. Bisgard, Freyja Lynn, and Bruce D. Meade. Mixture model analysis for establishing a diagnostic cut-off point for pertussis antibody levels. Statistics in Medicine, 25(17):2994–3010, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2006:GCF

[2120] Yin Bun Cheung. Growth and cognitive function of Indonesian children: zero-inflated proportion models. Statistics in Medicine, 25(17): 3011–3022, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2006:EEL

[2121] Haipeng Shen, Lawrence D. Brown, and Hui Zhi. Efficient estimation of log-normal means with application to pharmacokinetic data. Statistics in Medicine, 25(17):3023–3038, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bonett:2006:CIR

[2122] Douglas G. Bonett and Robert M. Price. Confidence intervals for a ratio of binomial proportions based on paired data. Statistics in Medicine, 25 (17):3039–3047, September 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2559, 2855].

Elston:2006:ASH

[2123] Robert C. Elston and M. Anne Spence. Advances in statistical human genetics over the last 25 years. Statistics in Medicine, 25(18):3049–3080, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kelly:2006:SGW

[2124] Patrick Kelly, Nigel Stallard, Yinghui Zhou, John Whitehead, and Clive Bowman. Sequential genome-wide association studies for monitoring adverse events in the clinical evaluation of new drugs. Statistics in Medicine, 25(18):3081–3092, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2006:MLM

[2125] K. F. Cheng. A maximum likelihood method for studying geneenvironment interactions under conditional independence of genotype and exposure. Statistics in Medicine, 25(18):3093–3109, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindstrom:2006:EGE

[2126] L. Lindström, Y. Pawitan, M. Reilly, K. Hemminki, P. Lichtenstein, and K. Czene. Estimation of genetic and environmental factors for melanoma onset using population-based family data. Statistics in Medicine, 25(18): 3110–3123, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karoui:2006:GMD

[2127] Noureddine El Karoui, Wei Zhou, and Alice S. Whittemore. Getting more from digital SNP data. Statistics in Medicine, 25(18):3124–3133, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberg:2006:MHT

[2128] Philip S. Rosenberg, Anney Che, and Bingshu E. Chen. Multiple hypothesis testing strategies for genetic case-control association studies. Statistics in Medicine, 25(18):3134–3149, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2006:EVC

[2129] Gang Zheng and Joseph L. Gastwirth. On estimation of the variance in Cochran–Armitage trend tests for genetic association using case-control studies. Statistics in Medicine, 25(18):3150–3159, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2006:RTT

[2130] Gang Zheng and Xin Tian. Robust trend tests for genetic association using matched case-control design. Statistics in Medicine, 25(18):3160– 3173, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2006:BDT

[2131] Jie Chen and Sanat K. Sarkar. A Bayesian determination of threshold for identifying differentially expressed genes in microarray experiments. Statistics in Medicine, 25(18):3174–3189, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wintrebert:2006:AGE

[2132] Claire M. A. Wintrebert, Aeilko H. Zwinderman, Anneke Maat-Kievit, Raymund A. Roos, and Hans C. van Houwelingen. Assessing genetic effects in survival data by correlating martingale residuals with an application to age at onset of Huntington disease. Statistics in Medicine, 25(18):3190–3200, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanHouwelingen:2006:CVC

[2133] Hans C. van Houwelingen, Tako Bruinsma, Augustinus A. M. Hart, Laura J. van’t Veer, and Lodewyk F. A. Wessels. Cross-validated Cox regression on microarray gene expression data. Statistics in Medicine, 25(18):3201–3216, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hines:2006:LEI

[2134] R. J. O’Hara Hines and W. G. S. Hines. Letter to the Editor: An index of local sensitivity to nonignorable drop-out in longitudinal modelling by G. Ma, A. B. Troxel and D. F. Heitjan, Statistics in Medicine 2005; 24:2129–2150. Statistics in Medicine, 25(18):3217–3218, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1755].

Guoguang:2006:AR

[2135] Ma Guoguang, Andrea B. Troxel, and Daniel F. Heitjan. Authors’ reply. Statistics in Medicine, 25(18):3218–3220, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2006:PIT

[2136] Say-Beng Tan and David Machin. Phase II trial designs in the presence of stratification. Statistics in Medicine, 25(18):3220–3222, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

London:2006:AR

[2137] Wendy B. London and Myron Chang. Authors’ reply. Statistics in Medicine, 25(18):3222–3223, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berger:2006:BRP

[2138] Vance W. Berger. Book review: 1. Permutation, parametric, and bootstrap tests of hypotheses (3rd edn). Phillip Good, Springer, New York, 2005. No. of pages: xix + 315. Price: $84.95. ISBN 0-387-20279-X. Statistics in Medicine, 25(18):3224–3225, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robert:2006:BRG

[2139] Christian P. Robert. Book review: 2. Gaussian Markov Random Fields: Theory and Applications. Håvard Rue and Leonhard Held, Chapman and Hall/CRC, Boca Raton, 2005. No. of pages: 280 (hardcover). Price: $79.95. ISBN: 1-58488-432-0. Statistics in Medicine, 25(18):3225–3226, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karlsson:2006:BRR

[2140] Andreas Karlsson. Book review: 3. Regression methods in biostatistics: linear, logistic, survival, and repeated measures models. Eric Vittinghoff, David V. Glidden, Stephen C. Shiboski and Charles E. McCulloch, Springer, Heidelberg, 2005. No. of pages: xv + 340. Price: $79.95 (hardcover). ISBN: 0-387-20275-7. Statistics in Medicine, 25(18):3226– 3227, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:2006:C

[2141] Bernard Rosner and Robert Glynn. Correction. Statistics in Medicine, 25(18):3228, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaufman:2006:C

[2142] Jay S. Kaufman. Correction. Statistics in Medicine, 25(18):3228, September 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2006:P

[2143] Susan S. Ellenberg, Howard Golub, Cyrus Metha, and Ralph D’Agostino. Preface. Statistics in Medicine, 25(19):3229–3230, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Golub:2006:NME

[2144] Howard L. Golub. The need for more efficient trial designs. Statistics in Medicine, 25(19):3231–3235, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsiatis:2006:IBM

[2145] Anastasios A. Tsiatis. Information-based monitoring of clinical trials. Statistics in Medicine, 25(19):3236–3244, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Temple:2006:FPT

[2146] Robert Temple. FDA perspective on trials with interim efficacy evaluations. Statistics in Medicine, 25(19):3245–3249, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2006:AGS

[2147] Cyrus R. Mehta and Nitin R. Patel. Adaptive, group sequential and decision theoretic approaches to sample size determination. Statistics in Medicine, 25(19):3250–3269, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emerson:2006:IUA

[2148] Scott S. Emerson. Issues in the use of adaptive clinical trial designs. Statistics in Medicine, 25(19):3270–3296, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang-Stein:2006:D

[2149] Christy Chuang-Stein. Discussion. Statistics in Medicine, 25(19):3297– 3301, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2006:D

[2150] Ralph B. D’Agostino, Sr. Discussion. Statistics in Medicine, 25(19):3302– 3304, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleming:2006:SVA

[2151] Thomas R. Fleming. Standard versus adaptive monitoring procedures: a commentary. Statistics in Medicine, 25(19):3305–3312, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hung:2006:D

[2152] Hsien-Ming James Hung. Discussion. Statistics in Medicine, 25(19): 3313–3314, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siegel:2006:D

[2153] Jay P. Siegel. Discussion. Statistics in Medicine, 25(19):3315–3319, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turnbull:2006:D

[2154] Bruce Turnbull. Discussion. Statistics in Medicine, 25(19):3320–3325, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:2006:GD

[2155] Susan S. Ellenberg. Group discussion. Statistics in Medicine, 25(19): 3326–3347, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2006:OCS

[2156] John M. Lachin. Operating characteristics of sample size re-estimation with futility stopping based on conditional power. Statistics in Medicine, 25(19):3348–3365, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brannath:2006:EFT

[2157] Werner Brannath, Franz König, and Peter Bauer. Estimation in flexible two stage designs. Statistics in Medicine, 25(19):3366–3381, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Banerjee:2006:ATS

[2158] Anindita Banerjee and Anastasios A. Tsiatis. Adaptive two-stage designs in phase II clinical trials. Statistics in Medicine, 25(19):3382–3395, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwartz:2006:TSS

[2159] Todd A. Schwartz and Jonathan S. Denne. A two-stage sample size recalculation procedure for placebo- and active-controlled non-inferiority trials. Statistics in Medicine, 25(19):3396–3406, October 15, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:2006:LEB

[2160] Say-Beng Tan and David Machin. Letter to the Editor: Bayesian twostage designs for phase II clinical trials by S. B. Tan and D. Machin, Statistics in Medicine 2002; 21:1991–2012. Statistics in Medicine, 25 (19):3407–3408, October 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [823].

Zelen:2006:BBS

[2161] Marvin Zelen. Biostatisticians, biostatistical science and the future. Statistics in Medicine, 25(20):3409–3414, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeMets:2006:TNG

[2162] David L. DeMets, Gary Stormo, Michael Boehnke, Thomas A. Louis, Jeremy Taylor, and Dennis Dixon. Training of the next generation of biostatisticians: a call to action in the U.S. Statistics in Medicine, 25 (20):3415–3429, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2006:CTS

[2163] Stephen Senn. Cross-over trials in Statistics in Medicine: the first ‘25’ years. Statistics in Medicine, 25(20):3430–3442, October 30, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harbord:2006:MTS

[2164] Roger M. Harbord, Matthias Egger, and Jonathan A. C. Sterne. A modified test for small-study effects in meta-analyses of controlled trials with binary endpoints. Statistics in Medicine, 25(20):3443–3457, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wruck:2006:SDE

[2165] Lisa M. Wruck, Constantin T. Yiannoutsos, and Michael D. Hughes. A sequential design to estimate sensitivity and specificity of a diagnostic or screening test. Statistics in Medicine, 25(20):3458–3473, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chambless:2006:ETD

[2166] Lloyd E. Chambless and Guoqing Diao. Estimation of time-dependent area under the ROC curve for long-term risk prediction. Statistics in Medicine, 25(20):3474–3486, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qin:2006:NCI

[2167] Gengsheng Qin, Yu-Sheng Hsu, and Xiao-Hua Zhou. New confidence intervals for the difference between two sensitivities at a fixed level of specificity. Statistics in Medicine, 25(20):3487–3502, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2006:SAU

[2168] Chiu-Hsieh Hsu, Jeremy M. G. Taylor, Susan Murray, and Daniel Commenges. Survival analysis using auxiliary variables via non-parametric multiple imputation. Statistics in Medicine, 25(20):3503–3517, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Box-Steffensmeier:2006:RES

[2169] Janet M. Box-Steffensmeier and Suzanna De Boef. Repeated events survival models: the conditional frailty model. Statistics in Medicine, 25 (20):3518–3533, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Szwarc:2006:MMS

[2170] Suzanne E. Szwarc and Marco Bonetti. Modelling menstrual status during and after adjuvant treatment for breast cancer. Statistics in Medicine, 25(20):3534–3547, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xia:2006:CEA

[2171] Yingcun Xia and Howell Tong. Cumulative effects of air pollution on public health. Statistics in Medicine, 25(20):3548–3559, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bonetto:2006:ACT

[2172] C. Bonetto, S. Giannerini, and A. Giovagnoli. The analysis of contingency tables with ordinal data: an application to monitoring antibiotic resistance. Statistics in Medicine, 25(20):3560–3575, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ren:2006:ICC

[2173] Shiquan Ren, Shuqin Yang, and Shenghan Lai. Intraclass correlation coefficients and bootstrap methods of hierarchical binary outcomes. Statistics in Medicine, 25(20):3576–3588, October 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:2006:BSM

[2174] Deborah Ashby. Bayesian statistics in medicine: a 25 year review. Statistics in Medicine, 25(21):3589–3631, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2006:NNP

[2175] Byron J. Gajewski, Robert Lee, Sarah Thompson, Nancy Dunton, Annette Becker, and Valorie Wells. Non-normal path analysis in the presence of measurement error and missing data: a Bayesian analysis of nursing homes’ structure and outcomes. Statistics in Medicine, 25(21): 3632–3647, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gardiner:2006:DME

[2176] Joseph C. Gardiner, Zhehui Luo, Cathy J. Bradley, Corina M. Sirbu, and Charles W. Given. A dynamic model for estimating changes in health status and costs. Statistics in Medicine, 25(21):3648–3667, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2006:EGD

[2177] Anastasia Ivanova. Escalation, group and A + B designs for dose-finding trials. Statistics in Medicine, 25(21):3668–3678, November 15, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borkowf:2006:CBC

[2178] Craig B. Borkowf. Constructing binomial confidence intervals with near nominal coverage by adding a single imaginary failure or success. Statistics in Medicine, 25(21):3679–3695, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Felder:2006:PBM

[2179] Stefan Felder and Bernt-Peter Robra. A preference-based measure for test performance with an application to prenatal diagnostics. Statistics in Medicine, 25(21):3696–3706, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:2006:BPP

[2180] Yin Bun Cheung and K. F. Lam. Bivariate Poisson-Poisson model of zero-inflated absenteeism data. Statistics in Medicine, 25(21):3707–3717, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bugli:2006:FAR

[2181] Céline Bugli and Philippe Lambert. Functional ANOVA with random functional effects: an application to event-related potentials modelling for electroencephalograms analysis. Statistics in Medicine, 25(21):3718– 3739, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noren:2006:EMU

[2182] G. Niklas Norén, Andrew Bate, Roland Orre, and I. Ralph Edwards. Extending the methods used to screen the WHO drug safety database towards analysis of complex associations and improved accuracy for rare events. Statistics in Medicine, 25(21):3740–3757, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nishikawa:2006:LET

[2183] Masako Nishikawa and Tango Toshiro. Letter to the Editor: Testing treatment effects in the presence of competing risks by Boris Freidlin and Edward L. Korn, Statistics in Medicine 2005; 24:1703–1712. Statistics in Medicine, 25(21):3758–3760, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1719].

Freidlin:2006:AR

[2184] Boris Freidlin and Edward L. Korn. Authors’ reply. Statistics in Medicine, 25(21):3761–3763, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2006:AR

[2185] Toshiro Tango. Author’s reply. Statistics in Medicine, 25(21):3763–3764, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Okamoto:2006:LEL

[2186] Etsuji Okamoto. Letter to the Editor: Linear equations with random variables by Toshiro Tango, Statistics in Medicine 2005; 24:3213–3222.

Statistics in Medicine, 25(21):3764–3767, November 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1836].

Harel:2006:MIC

[2187] Ofer Harel and Xiao-Hua Zhou. Multiple imputation for correcting verification bias. Statistics in Medicine, 25(22):3769–3786, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2437].

Marshall:2006:CMR

[2188] Roger J. Marshall. Comparison of misclassification rates of search partition analysis and other classification methods. Statistics in Medicine, 25(22):3787–3797, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowden:2006:UPD

[2189] Jack Bowden, John R. Thompson, and Paul Burton. Using pseudo-data to correct for publication bias in meta-analysis. Statistics in Medicine, 25(22):3798–3813, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2006:CCT

[2190] Li Yin, Rolf Sundberg, Xiaoqin Wang, and Donald B. Rubin. Control of confounding through secondary samples. Statistics in Medicine, 25(22): 3814–3825, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2006:MLD

[2191] Song Wu, Jie Yang, and Rongling Wu. Multilocus linkage disequilibrium mapping of epistatic quantitative trait loci that regulate HIV dynamics: a simulation approach. Statistics in Medicine, 25(22):3826–3849, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xue:2006:SPA

[2192] Hongqi Xue, K. F. Lam, Benjamin J. Cowling, and Frank de Wolf. Semiparametric accelerated failure time regression analysis with application to interval-censored HIV/AIDS data. Statistics in Medicine, 25(22):3850– 3863, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2006:MLR

[2193] Kui Wang, Andy H. Lee, Greg Hamilton, and Kelvin K. W. Yau. Multilevel logistic regression modelling with correlated random effects: application to the smoking cessation for youth study. Statistics in Medicine, 25(22):3864–3876, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yun:2006:REM

[2194] Sungcheol Yun and Youngjo Lee. Robust estimation in mixed linear models with non-monotone missingness. Statistics in Medicine, 25(22): 3877–3892, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marler:2006:STC

[2195] Matthew R. Marler, Philip Gehrman, Jennifer L. Martin, and Sonia Ancoli-Israel. The sigmoidally transformed cosine curve: a mathematical model for circadian rhythms with symmetric non-sinusoidal shapes. Statistics in Medicine, 25(22):3893–3904, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Erkanli:2006:BSP

[2196] Alaattin Erkanli, Minje Sung, E. Jane Costello, and Adrian Angold. Bayesian semi-parametric ROC analysis. Statistics in Medicine, 25(22): 3905–3928, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulldorff:2006:ESS

[2197] Martin Kulldorff, Lan Huang, Linda Pickle, and Luiz Duczmal. An elliptic spatial scan statistic. Statistics in Medicine, 25(22):3929–3943, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2006:LEP

[2198] Stephen Senn. Letter to the Editor: Probabilistic index: an intuitive nonparametric approach to measuring the size of the treatment effects by L. Acion, J. J. Peterson, S. Temple and S. Arndt, Statistics in Medicine 2006; 25 (4):591–602. Statistics in Medicine, 25(22):3944–3946, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Acion:2006:AR

[2199] Laura Acion, John J. Peterson, Scott Temple, and Stephan Arndt. Authors’ reply. Statistics in Medicine, 25(22):3946–3948, November 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mazumdar:2006:SDI

[2200] Madhu Mazumdar, Donsheng Tu, and Xi Kathy Zhou. Some design issues of strata-matched non-randomized studies with survival outcomes. Statistics in Medicine, 25(23):3949–3959, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2006:IAB

[2201] Ying C. MacNab, Andrew Kmetic, Paul Gustafson, and Sam Sheps. An innovative application of Bayesian disease mapping methods to patient safety research: a Canadian adverse medical event study. Statistics in Medicine, 25(23):3960–3980, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nishikawa:2006:NPI

[2202] Masako Nishikawa, Toshiro Tango, and Makiko Ogawa. Non-parametric inference of adverse events under informative censoring. Statistics in Medicine, 25(23):3981–4003, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graham:2006:OBS

[2203] Gordon Graham and Leon Aarons. Optimum blood sampling time windows for parameter estimation in population pharmacokinetic experiments. Statistics in Medicine, 25(23):4004–4019, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hou:2006:FME

[2204] Wei Hou, Cynthia W. Garvan, Ramon C. Littell, Marylou Behnke, Fonda Davis Eyler, and Rongling Wu. A framework to monitor environment-induced major genes for developmental trajectories: implication for a prenatal cocaine exposure study. Statistics in Medicine, 25 (23):4020–4035, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rondeau:2006:NFM

[2205] V. Rondeau, L. Filleul, and P. Joly. Nested frailty models using maximum penalized likelihood estimation. Statistics in Medicine, 25(23):

4036–4052, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harezlak:2006:ESF

[2206] Jaroslaw Harezlak and Wanzhu Tu. Estimation of survival functions in interval and right censored data using STD behavioural diaries. Statistics in Medicine, 25(23):4053–4064, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:2006:ICC

[2207] Robert H. Lyles, Andrew S. Allen, W. Dana Flanders, Lawrence L. Kupper, and Deborah L. Christensen. Inference for case-control studies when exposure status is both informatively missing and misclassified. Statistics in Medicine, 25(23):4065–4080, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCaffrey:2006:IHT

[2208] Daniel F. McCaffrey and Robert M. Bell. Improved hypothesis testing for coefficients in generalized estimating equations with small samples of clusters. Statistics in Medicine, 25(23):4081–4098, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2006:IER

[2209] Yea-Hung Chen and Xiao-Hua Zhou. Interval estimates for the ratio and difference of two lognormal means. Statistics in Medicine, 25(23):4099– 4113, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shults:2006:ARB

[2210] Justine Shults, Carissa A. Mazurick, and J. Richard Landis. Analysis of repeated bouts of measurements in the framework of generalized estimating equations. Statistics in Medicine, 25(23):4114–4128, December 15, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:2006:P

[2211] Michael Schemper. Preface. Statistics in Medicine, 25(24):4129, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2006:ISC

[2212] Anonymous. International Society for Clinical Biostatistics. Statistics in Medicine, 25(24):4130–4131, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leffondre:2006:MSH

[2213] Karen Leffondré, Michal Abrahamowicz, Yongling Xiao, and Jack Siemiatycki. Modelling smoking history using a comprehensive smoking index: application to lung cancer. Statistics in Medicine, 25(24):4132–4146, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salanti:2006:BSE

[2214] Georgia Salanti, Julian P. T. Higgins, and Ian R. White. Bayesian synthesis of epidemiological evidence with different combinations of exposure groups: application to a gene-gene-environment interaction. Statistics in Medicine, 25(24):4147–4163, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Touloumi:2006:SCA

[2215] G. Touloumi, E. Samoli, M. Pipikou, A. Le Tertre, R. Atkinson, and K. Katsouyanni. Seasonal confounding in air pollution and health timeseries studies: effect on air pollution effect estimates. Statistics in Medicine, 25(24):4164–4178, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heisterkamp:2006:ADI

[2216] Simon H. Heisterkamp, Arnold L. M. Dekkers, and Janneke C. M. Heijne. Automated detection of infectious disease outbreaks: hierarchical time series models. Statistics in Medicine, 25(24):4179–4196, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosthoj:2006:EOD

[2217] Susanne Rosthøj, Catherine Fullwood, Robin Henderson, and Syd Stewart. Estimation of optimal dynamic anticoagulation regimes from observational data: a regret-based approach. Statistics in Medicine, 25(24): 4197–4215, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinze:2006:CIM

[2218] Georg Heinze. A comparative investigation of methods for logistic regression with separated or nearly separated data. Statistics in Medicine, 25 (24):4216–4226, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valenta:2006:POL

[2219] Zdenek Valenta, Jan Pitha, and Rudolf Poledne. Proportional odds logistic regression-effective means of dealing with limited uncertainty in dichotomizing clinical outcomes. Statistics in Medicine, 25(24):4227– 4234, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:2006:DOR

[2220] Robert G. Newcombe. A deficiency of the odds ratio as a measure of effect size. Statistics in Medicine, 25(24):4235–4240, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsonaka:2006:PSS

[2221] Roula Tsonaka, Dimitris Rizopoulos, and Emmanuel Lesaffre. Power and sample size calculations for discrete bounded outcome scores. Statistics in Medicine, 25(24):4241–4252, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perperoglou:2006:RGF

[2222] Aris Perperoglou, Hans C. van Houwelingen, and Robin Henderson. A relaxation of the gamma frailty (Burr) model. Statistics in Medicine, 25 (24):4253–4266, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katsahian:2006:AMC

[2223] Sandrine Katsahian, Matthieu Resche-Rigon, Sylvie Chevret, and Raphaël Porcher. Analysing multicentre competing risks data with a mixed proportional hazards model for the subdistribution. Statistics in Medicine, 25(24):4267–4278, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burton:2006:DSS

[2224] Andrea Burton, Douglas G. Altman, Patrick Royston, and Roger L. Holder. The design of simulation studies in medical statistics. Statistics in Medicine, 25(24):4279–4292, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2499].

Baksh:2006:CCC

[2225] M. Fazil Baksh, Gerco Haars, Susan Todd, Paulus A. H. Van Noord, and John Whitehead. Comparing correlations of continuous observations from two independent populations using a sequential approach. Statistics in Medicine, 25(24):4293–4310, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zohar:2006:ODE

[2226] Sarah Zohar and John O’Quigley. Optimal designs for estimating the most successful dose. Statistics in Medicine, 25(24):4311–4320, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mittlbock:2006:SSC

[2227] M. Mittlböck and H. Heinzl. A simulation study comparing properties of heterogeneity measures in meta-analyses. Statistics in Medicine, 25(24): 4321–4333, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2006:CBA

[2228] Stephen Senn. Change from baseline and analysis of covariance revisited. Statistics in Medicine, 25(24):4334–4344, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2006:A

[2229] Anonymous. Acknowledgement. Statistics in Medicine, 25(24):4345– 4350, December 30, 2006. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2007:BNY

[2230] Ralph B. D’Agostino, Sr., Michael J. Campbell, and Joel B. Greenhouse. Beginning the next 25 years. Statistics in Medicine, 26(1):1, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:2007:DCR

[2231] M. J. Campbell, A. Donner, and N. Klar. Developments in cluster randomized trials and Statistics in Medicine. Statistics in Medicine, 26(1): 2–19, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:2007:DVA

[2232] Donald B. Rubin. The design versus the analysis of observational studies for causal effects: parallels with the design of randomized trials. Statistics in Medicine, 26(1):20–36, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Viechtbauer:2007:CIA

[2233] Wolfgang Viechtbauer. Confidence intervals for the amount of heterogeneity in meta-analysis. Statistics in Medicine, 26(1):37–52, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bradburn:2007:MAA

[2234] Michael J. Bradburn, Jonathan J. Deeks, Jesse A. Berlin, and A. Russell Localio. Much ado about nothing: a comparison of the performance of meta-analytical methods with rare events. Statistics in Medicine, 26(1): 53–77, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2007:EBR

[2235] R. D. Riley, K. R. Abrams, P. C. Lambert, A. J. Sutton, and J. R. Thompson. An evaluation of bivariate random-effects meta-analysis for the joint synthesis of two correlated outcomes. Statistics in Medicine, 26(1):78–97, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rutter:2007:HNH

[2236] Carolyn M. Rutter, Onchee Yu, and Diana L. Miglioretti. A hierarchical non-homogeneous Poisson model for meta-analysis of adenoma counts. Statistics in Medicine, 30(4):98–109, January 15, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:2007:BRE

[2237] Andrew C. Leon, Donald Hedeker, and Jedediah J. Teres. Bias reduction in effectiveness analyses of longitudinal ordinal doses with a mixed-effects propensity adjustment. Statistics in Medicine, 30(4):110–123, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spliid:2007:MMPa

[2238] Henrik Spliid. Monitoring medical procedures by exponential smoothing. Statistics in Medicine, 30(4):124–138, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2007:SRE

[2239] Lei Liu, Robert A. Wolfe, and John D. Kalbfleisch. A shared random effects model for censored medical costs and mortality. Statistics in Medicine, 30(4):139–155, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2007:MAC

[2240] Stephen A. Roberts. Models for assisted conception data with embryospecific covariates. Statistics in Medicine, 30(4):156–170, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glass:2007:PCN

[2241] K. Glass, N. Becker, and M. Clements. Predicting case numbers during infectious disease outbreaks when some cases are undiagnosed. Statistics in Medicine, 30(4):171–183, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graham:2007:IQP

[2242] P. L. Graham, P. M. Kuhnert, D. A. Cook, and K. Mengersen. Improving the quality of patient care using reliability measures: a classification tree approach. Statistics in Medicine, 30(4):184–196, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2007:ERR

[2243] Yongming Qu, Pandurang M. Kulkarni, and Todd M. Sanger. Estimating the response rate in the presence of measurement error. Statistics in Medicine, 30(4):197–211, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauleau:2007:APC

[2244] Erik A. Sauleau, Andrea Hennerfeind, Antoine Buemi, and Leonhard Held. Age, period and cohort effects in Bayesian smoothing of spatial cancer survival with geoadditive models. Statistics in Medicine, 30(4): 212–229, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2007:LEI

[2245] Anonymous. Letter to the Editor: An introductory note to the CHMP guidelines: choice of the non-inferiority margin and data monitoring committees by David Brown, Peter Volkers and Simon Day, Statistics in Medicine 2006; 25:1623–1627. Statistics in Medicine, 30(4):230–233, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2019].

Brown:2007:AR

[2246] David Brown and Simon Day. Authors’ reply. Statistics in Medicine, 30(4):234–236, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spliid:2007:MMPb

[2247] Henrik Spliid. Monitoring medical procedures by exponential smoothing. Statistics in Medicine, 30(4):236, January 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Todd:2007:BRY

[2248] Susan Todd. Book review: A 25-year review of sequential methodology in clinical studies. Statistics in Medicine, 26(2):237–252, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kieser:2007:PAT

[2249] Meinhard Kieser and Tim Friede. Planning and analysis of three-arm non-inferiority trials with binary endpoints. Statistics in Medicine, 26 (2):253–273, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [3385].

Turner:2007:ACR

[2250] Rebecca M. Turner, Ian R. White, and Tim Croudace. Analysis of cluster randomized cross-over trial data: a comparison of methods. Statistics in Medicine, 26(2):274–289, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nedelman:2007:DCP

[2251] Jerry R. Nedelman, Donald B. Rubin, and Lewis B. Sheiner. Diagnostics for confounding in PK/PD models for oxcarbazepine. Statistics in Medicine, 26(2):290–308, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zuma:2007:AIC

[2252] K. Zuma, M. Lurie, and M. Jorgensen. Analysis of interval-censored data from circular migrant and non-migrant sexual partnerships using the EM algorithm. Statistics in Medicine, 26(2):309–319, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sakarovitch:2007:EIH

[2253] Charlotte Sakarovitch, Ahmadou Alioum, Didier K. Ekouevi, Philippe Msellati, Valériane Leroy, and François Dabis. Estimating incidence of HIV infection in childbearing age African women using serial prevalence data from antenatal clinics. Statistics in Medicine, 26(2):320–335, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shoukri:2007:BMC

[2254] M. M. Shoukri and A. Donner. Bivariate models for co-aggregation of dichotomous traits in twins. Statistics in Medicine, 26(2):336–351, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bulaevskaya:2007:PLA

[2255] Vera L. Bulaevskaya and Gary W. Oehlert. A penalized likelihood approach to magnetic resonance image reconstruction. Statistics in Medicine, 26(2):352–374, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2007:CTC

[2256] Kejian Liu, Peihua Qiu, and Jun Sheng. Comparing two crossing hazard rates by Cox proportional hazards modelling. Statistics in Medicine, 26 (2):375–391, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abrahamowicz:2007:JET

[2257] Michal Abrahamowicz and Todd A. MacKenzie. Joint estimation of time-dependent and non-linear effects of continuous covariates on survival. Statistics in Medicine, 26(2):392–408, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2007:MEM

[2258] Changyu Shen and Sujuan Gao. A mixed-effects model for cognitive decline with non-monotone non-response from a two-phase longitudinal study of dementia. Statistics in Medicine, 26(2):409–425, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siannis:2007:MSM

[2259] F. Siannis, V. T. Farewell, and J. Head. A multi-state model for joint modelling of terminal and non-terminal events with application to Whitehall II. Statistics in Medicine, 26(2):426–442, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:2007:RRB

[2260] Yu-Kang Tu and Mark S. Gilthorpe. Revisiting the relation between change and initial value: a review and evaluation. Statistics in Medicine, 26(2):443–457, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2450].

Erbas:2007:FAS

[2261] Bircan Erbas, Rob J. Hyndman, and Dorota M. Gertig. Forecasting agespecific breast cancer mortality using functional data models. Statistics in Medicine, 26(2):458–470, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2007:LAM

[2262] Anonymous. Life as an academic medical statistician and how to survive it — a south African perspective. Statistics in Medicine, 26(2):471–472, January 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:2007:CY

[2263] T. Colton, L. S. Freedman, and A. L. Johnson. A celebration of 25 years. Statistics in Medicine, 26(3):473–483, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mogg:2007:AAI

[2264] Robin Mogg and Devan V. Mehrotra. Analysis of antiretroviral immunotherapy trials with potentially non-normal and incomplete longitudinal data. Statistics in Medicine, 26(3):484–497, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wakana:2007:MTD

[2265] Akira Wakana, Isao Yoshimura, and Chikuma Hamada. A method for therapeutic dose selection in a phase II clinical trial using contrast statistics. Statistics in Medicine, 26(3):498–511, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gorst-Rasmussen:2007:MIS

[2266] A. Gorst-Rasmussen, D. J. Spiegelhalter, and C. Bull. Monitoring the introduction of a surgical intervention with long-term consequences. Statistics in Medicine, 26(3):512–531, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fung:2007:CMA

[2267] Karen Y. Fung, Shahedul Khan, Daniel Krewski, and Tim Ramsay. A comparison of methods for the analysis of recurrent health outcome data with environmental covariates. Statistics in Medicine, 26(3):532–545, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clarke:2007:NPE

[2268] Paul S. Clarke and Azra C. Ghani. A note on parameter estimation for variant Creutzfeldt–Jakob disease epidemic models. Statistics in Medicine, 26(3):546–552, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salanti:2007:BMA

[2269] Georgia Salanti, Julian P. T. Higgins, Thomas A. Trikalinos, and John P. A. Ioannidis. Bayesian meta-analysis and meta-regression for genedisease associations and deviations from Hardy-Weinberg equilibrium. Statistics in Medicine, 26(3):553–567, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salazar:2007:SRE

[2270] Juan C. Salazar, Frederick A. Schmitt, Lei Yu, Marta M. Mendiondo, and Richard J. Kryscio. Shared random effects analysis of multi-state Markov models: application to a longitudinal study of transitions to dementia. Statistics in Medicine, 26(3):568–580, February 10, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Plevritis:2007:NHM

[2271] Sylvia K. Plevritis, Peter Salzman, Bronislava M. Sigal, and Peter W. Glynn. A natural history model of stage progression applied to breast cancer. Statistics in Medicine, 26(3):581–595, February 10, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hillis:2007:CDD

[2272] Stephen L. Hillis. A comparison of denominator degrees of freedom methods for multiple observer ROC analysis. Statistics in Medicine, 26(3):

596–619, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:2007:CVA

[2273] Jun mo Nam. Comparison of validity of assessment methods using indices of adjusted agreement. Statistics in Medicine, 26(3):620–632, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:2007:IER

[2274] Bernard Rosner and Robert J. Glynn. Interval estimation for rank correlation coefficients based on the probit transformation with extension to measurement error correction of correlated ranked data. Statistics in Medicine, 26(3):633–646, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valet:2007:LLN

[2275] Fabien Valet, Christiane Guinot, and Jean Yves Mary. Log-linear nonuniform association models for agreement between two ratings on an ordinal scale. Statistics in Medicine, 26(3):647–662, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Falcaro:2007:FMM

[2276] Milena Falcaro and Andrew Pickles. A flexible model for multivariate interval-censored survival times with complex correlation structure. Statistics in Medicine, 26(3):663–680, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2007:BAL

[2277] Xin-Yuan Song and Sik-Yum Lee. Bayesian analysis of latent variable models with non-ignorable missing outcomes from exponential family. Statistics in Medicine, 26(3):681–693, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2007:DRT

[2278] Wonkuk Kim, Hongshik Ahn, and Hojin Moon. A dose-response test via closed-form solutions for constrained MLEs in survival/sacrifice experiments. Statistics in Medicine, 26(3):694–708, February 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deutsch:2007:REB

[2279] Reena Deutsch, Shelley Hurwitz, Janine Janosky, and Robert Oster. The role of education in biostatistical consulting. Statistics in Medicine, 26 (4):709–720, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwarzer:2007:TPB

[2280] Guido Schwarzer, Gerd Antes, and Martin Schumacher. A test for publication bias in meta-analysis with sparse binary data. Statistics in Medicine, 26(4):721–733, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2007:CAD

[2281] Peter C. Austin, Paul Grootendorst, and Geoffrey M. Anderson. A comparison of the ability of different propensity score models to balance measured variables between treated and untreated subjects: a Monte Carlo study. Statistics in Medicine, 26(4):734–753, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2007:CPS

[2282] Peter C. Austin, Paul Grootendorst, Sharon-Lise T. Normand, and Geoffrey M. Anderson. Conditioning on the propensity score can result in biased estimation of common measures of treatment effect: a Monte Carlo study. Statistics in Medicine, 26(4):754–768, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2452].

Hsu:2007:MII

[2283] Chiu-Hsieh Hsu, Jeremy M. G. Taylor, Susan Murray, and Daniel Commenges. Multiple imputation for interval censored data with auxiliary variables. Statistics in Medicine, 26(4):769–781, February 20, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demirtas:2007:GBQ

[2284] Hakan Demirtas and Donald Hedeker. Gaussianization-based quasiimputation and expansion strategies for incomplete correlated binary responses. Statistics in Medicine, 26(4):782–799, February 20, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dang:2007:UTB

[2285] Qianyu Dang, Sati Mazumdar, Stewart J. Anderson, Patricia R. Houck, and Charles F. Reynolds. Using trajectories from a bivariate growth curve as predictors in a Cox regression model. Statistics in Medicine, 26(4):800–811, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2007:TET

[2286] Yunchan Chi, Wei-Yann Tsai, and Chia-Ling Chiang. Testing the equality of two survival functions with right truncated data. Statistics in Medicine, 26(4):812–827, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2007:ISR

[2287] B. M. Brown and You-Gan Wang. Induced smoothing for rank regression with censored survival times. Statistics in Medicine, 26(4):828–836, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dignam:2007:MDA

[2288] James J. Dignam, Kelly Wieand, and Paul J. Rathouz. A missing data approach to semi-competing risks problems. Statistics in Medicine, 26 (4):837–856, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2007:NPE

[2289] Feng Gao, Amita K. Manatunga, and Shande Chen. Non-parametric estimation for baseline hazards function and covariate effects with timedependent covariates. Statistics in Medicine, 26(4):857–868, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2007:TCTa

[2290] Jason Kunming Huang and John Jen Tai. Twin concordances test for ascertained trichotomous traits data. Statistics in Medicine, 26(4):869– 894, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:2007:SMD

[2291] Barbra A. Richardson, James P. Hughes, and Sarah Benki. Statistical methods for determining the accuracy of quantitative polymerase chain reaction-based tests. Statistics in Medicine, 26(4):895–902, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bull:2007:CIM

[2292] Shelley B. Bull, Juan Pablo Lewinger, and Sophia S. F. Lee. Confidence intervals for multinomial logistic regression in sparse data. Statistics in Medicine, 26(4):903–918, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cantoni:2007:LVS

[2293] E. Cantoni, C. Field, J. Mills Flemming, and E. Ronchetti. Longitudinal variable selection by cross-validation in the case of many covariates. Statistics in Medicine, 26(4):919–930, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Norman:2007:MIR

[2294] Robert G. Norman and Marc A. Scott. Measurement of inter-rater agreement for transient events using Monte Carlo sampled permutations. Statistics in Medicine, 26(4):931–942, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rogatko:2007:BRS

[2295] André Rogatko. Book review: Statistical Concepts and Applications in Clinical Medicine. John Aitchison, Jim W. Kay and Ian J. Lauder, Chapman Hall/CRC, Boca Raton, 2005. No. of pages: 339. Price: $79.95. ISBN 1-58488-208-5. Statistics in Medicine, 26(4):943–944, February 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2007:ISM

[2296] Lili Tian. Inferences on standardized mean difference: the generalized variable approach. Statistics in Medicine, 26(5):945–953, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bang:2007:ETE

[2297] Heejung Bang and Clarence E. Davis. On estimating treatment effects under non-compliance in randomized clinical trials: are intent-to-treat or instrumental variables analyses perfect solutions? Statistics in Medicine, 26(5):954–964, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3136].

Latouche:2007:MRM

[2298] A. Latouche, V. Boisson, S. Chevret, and R. Porcher. Misspecified regression model for the subdistribution hazard of a competing risk. Statistics in Medicine, 26(5):965–974, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2344].

Kitchen:2007:HVF

[2299] Christina M. R. Kitchen, Robert E. Weiss, Gang Liu, and Terri Wrin. HIV-1 viral fitness estimation using exchangeable on subsets priors and prior model selection. Statistics in Medicine, 26(5):975–990, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2007:RMA

[2300] Honghu Liu, Loren G. Miller, Carol E. Golin, Ron D. Hays, Tongtong Wu, Neil S. Wenger, and Andrew H. Kaplan. Repeated measures analyses of dose timing of antiretroviral medication and its relationship to HIV virologic outcomes. Statistics in Medicine, 26(5):991–1007, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tournoud:2007:APT

[2301] M. Tournoud and R. Ecochard. Application of the promotion time cure model with time-changing exposure to the study of HIV/AIDS and other infectious diseases. Statistics in Medicine, 26(5):1008–1021, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Judkins:2007:VSR

[2302] David R. Judkins, David Morganstein, Paul Zador, Andrea Piesse, Brandon Barrett, and Pushpal Mukhopadhyay. Variable selection and raking in propensity scoring. Statistics in Medicine, 26(5):1022–1033, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Todem:2007:LVM

[2303] David Todem, KyungMann Kim, and Emmanuel Lesaffre. Latentvariable models for longitudinal data with bivariate ordinal outcomes. Statistics in Medicine, 26(5):1034–1054, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Litman:2007:IME

[2304] Heather J. Litman, Nicholas J. Horton, Bernardo Hernández, and Nan M. Laird. Incorporating missingness for estimation of marginal regression models with multiple source predictors. Statistics in Medicine, 26(5): 1055–1068, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zwane:2007:ACR

[2305] Eugene N. Zwane and Peter G. M. van der Heijden. Analysing capturerecapture data when some variables of heterogeneous catchability are not collected or asked in all registrations. Statistics in Medicine, 26(5):1069– 1089, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kowalski:2007:NNP

[2306] Jeanne Kowalski, Amanda Blackford, Changyong Feng, Adam J. Mamelak, and Daniel N. Sauder. Nested, non-parametric, correlative analysis of microarrays for heterogeneous phenotype characterization. Statistics in Medicine, 26(5):1090–1101, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lusa:2007:ASR

[2307] Lara Lusa, Lisa M. McShane, Michael D. Radmacher, Joanna H. Shih, George W. Wright, and Richard Simon. Appropriateness of some resampling-based inference procedures for assessing performance of prognostic classifiers derived from microarray data. Statistics in Medicine, 26 (5):1102–1113, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2007:ACD

[2308] Hao Liu, Laurel A. Beckett, and Gerald L. DeNardo. On the analysis of count data of birth-and-death process type: with application to molecularly targeted cancer therapy. Statistics in Medicine, 26(5):1114– 1135, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salim:2007:ACC

[2309] Agus Salim, Niels G. Becker, Jette Möller, and Anne-Charlotte Hessén-Söderman. An analysis for a cross-over cohort study with an application to the study of triggers of Menière’s disease. Statistics in Medicine, 26 (5):1136–1149, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wraith:2007:ACE

[2310] D. Wraith and K. Mengersen. Assessing the combined effect of asbestos exposure and smoking on lung cancer: a Bayesian approach. Statistics in Medicine, 26(5):1150–1169, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Braun:2007:CSS

[2311] Thomas M. Braun and Zheng Yuan. Comparing the small sample performance of several variance estimators under competing risks. Statistics in Medicine, 26(5):1170–1180, February 28, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang-Stein:2007:CMC

[2312] Christy Chuang-Stein, Paul Stryszak, Alex Dmitrienko, and Walter Offen. Challenge of multiple co-primary endpoints: a new approach. Statistics in Medicine, 26(6):1181–1192, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bloch:2007:CSN

[2313] Daniel A. Bloch, Tze Leung Lai, Zheng Su, and Pascale Tubert-Bitter. A combined superiority and non-inferiority approach to multiple endpoints in clinical trials. Statistics in Medicine, 26(6):1193–1207, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chervoneva:2007:MTP

[2314] Inna Chervoneva, Terry Hyslop, and Walter W. Hauck. A multivariate test for population bioequivalence. Statistics in Medicine, 26(6):1208– 1223, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2007:SPB

[2315] Pulak Ghosh and Gary L. Rosner. A semi-parametric Bayesian approach to average bioequivalence. Statistics in Medicine, 26(6):1224– 1236, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nixon:2007:UMT

[2316] R. M. Nixon, N. Bansback, and A. Brennan. Using mixed treatment comparisons and meta-regression to perform indirect comparisons to estimate the efficacy of biologic treatments in rheumatoid arthritis. Statistics in Medicine, 26(6):1237–1254, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2007:BRE

[2317] Pulak Ghosh, Marcia D. Branco, and Hrishikesh Chakraborty. Bivariate random effect model using skew-normal distribution with application to HIV-RNA. Statistics in Medicine, 26(6):1255–1267, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Panhard:2007:IMI

[2318] Xavière Panhard, Anne-Marie Taburet, Christophe Piketti, and France Mentré. Impact of modelling intra-subject variability on tests based on non-linear mixed-effects models in cross-over pharmacokinetic trials with application to the interaction of tenofovir on atazanavir in HIV patients. Statistics in Medicine, 26(6):1268–1284, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:2007:IPP

[2319] Robert H. Lyles, Amita K. Manatunga, Reneé H. Moore, F. DuBois Bowman, and Curtiss B. Cook. Improving point predictions of random effects for subjects at high risk. Statistics in Medicine, 26(6):1285–1300, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Das:2007:ETR

[2320] Kalyan Das and Malay Naskar. Effectiveness of tibolone on the reduction of menopausal problems — a Bayesian semi-parametric interpretation. Statistics in Medicine, 26(6):1301–1317, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arostegui:2007:ASF

[2321] Inmaculada Arostegui, Vicente Núñez-Antón, and José M. Quintana. Analysis of the short form-36 (SF-36): the beta-binomial distribution approach. Statistics in Medicine, 26(6):1318–1342, March 15, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2007:CTS

[2322] Michael J. Pencina, Martin G. Larson, and Ralph B. D’Agostino. Choice of time scale and its effect on significance of predictors in longitudinal studies. Statistics in Medicine, 26(6):1343–1359, March 15, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3190].

Pintilie:2007:AIC

[2323] Melania Pintilie. Analysing and interpreting competing risk data. Statistics in Medicine, 26(6):1360–1367, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [2489, 2323, 2476].

Bernaards:2007:RMN

[2324] Coen A. Bernaards, Thomas R. Belin, and Joseph L. Schafer. Robustness of a multivariate normal approximation for imputation of incomplete binary data. Statistics in Medicine, 26(6):1368–1382, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harezlak:2007:LTB

[2325] Jaroslaw Harezlak, Elena Naumova, and Nan M. Laird. LongCriSP: a test for bump hunting in longitudinal data. Statistics in Medicine, 26 (6):1383–1397, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mitra:2007:SHR

[2326] Nandita Mitra and Daniel F. Heitjan. Sensitivity of the hazard ratio to nonignorable treatment assignment in an observational study. Statistics in Medicine, 26(6):1398–1414, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2007:CSE

[2327] Anonymous. A comment on: Statistical evaluation of biomarkers as surrogate endpoints: a literature review by C. J. Weir and R. J. Walley, Statistics in Medicine 2006; 25:183–203. Statistics in Medicine, 26(6): 1415–1416, March 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1915].

Senn:2007:TPA

[2328] Stephen Senn. Trying to be precise about vagueness. Statistics in Medicine, 26(7):1417–1430, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2676].

Emerson:2007:BEG

[2329] Scott S. Emerson, John M. Kittelson, and Daniel L. Gillen. Bayesian evaluation of group sequential clinical trial designs. Statistics in Medicine, 26(7):1431–1449, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jahn-Eimermacher:2007:PAS

[2330] Antje Jahn-Eimermacher and Gerhard Hommel. Performance of adaptive sample size adjustment with respect to stopping criteria and time of interim analysis. Statistics in Medicine, 26(7):1450–1461, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ayanlowo:2007:SCP

[2331] A. O. Ayanlowo and D. T. Redden. Stochastically curtailed phase II clinical trials. Statistics in Medicine, 26(7):1462–1472, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bond:2007:IVI

[2332] Simon J. Bond, Ian R. White, and A. Sarah Walker. Instrumental variables and interactions in the causal analysis of a complex clinical trial. Statistics in Medicine, 26(7):1473–1496, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiong:2007:SPM

[2333] Xiaoping Xiong, Ming Tan, and James Boyett. A sequential procedure for monitoring clinical trials against historical controls. Statistics in Medicine, 26(7):1497–1511, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:2007:SAD

[2334] Shein-Chung Chow and Jun Shao. Stability analysis for drugs with multiple active ingredients. Statistics in Medicine, 26(7):1512–1517, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guttner:2007:MTE

[2335] Achim Güttner, Jürgen Kübler, and Iris Pigeot. Multivariate time-to-event analysis of multiple adverse events of drugs in integrated analyses. Statistics in Medicine, 26(7):1518–1531, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Woodard:2007:PAR

[2336] D. B. Woodard, A. E. Gelfand, W. E. Barlow, and J. G. Elmore. Performance assessment for radiologists interpreting screening mammography. Statistics in Medicine, 26(7):1532–1551, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2007:FRA

[2337] Xiaowei Yang, Qing Shen, Hongquan Xu, and Steven Shoptaw. Functional regression analysis using an F test for longitudinal data with large numbers of repeated measures. Statistics in Medicine, 26(7):1552–1566, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2007:WLR

[2338] Chiu-Hsieh Hsu, Sylvan B. Green, and Yulei He. A weighted logistic regression model for estimation of recurrence of adenomas. Statistics in Medicine, 26(7):1567–1578, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2007:ULK

[2339] J. Brooke Marshall, Dan J. Spitzner, and William H. Woodall. Use of the local Knox statistic for the prospective monitoring of disease occurrences in space and time. Statistics in Medicine, 26(7):1579–1593, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2007:SSS

[2340] Inkyung Jung, Martin Kulldorff, and Ann C. Klassen. A spatial scan statistic for ordinal data. Statistics in Medicine, 26(7):1594–1607, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xiang:2007:STO

[2341] Liming Xiang, Andy H. Lee, Kelvin K. W. Yau, and Geoffrey J. McLachlan. A score test for overdispersion in zero-inflated Poisson mixed regression model. Statistics in Medicine, 26(7):1608–1622, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsou:2007:SEW

[2342] Tsung-Shan Tsou. A simple and exploratory way to determine the meanvariance relationship in generalized linear models. Statistics in Medicine, 26(7):1623–1631, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:2007:PAC

[2343] Robert H. Lyles, Hung-Mo Lin, and John M. Williamson. A practical approach to computing power for generalized linear models with nominal, count, or ordinal responses. Statistics in Medicine, 26(7):1632–1648, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2007:LEM

[2344] Anonymous. Letter to the Editor: Misspecified regression model for the subdistribution hazard of a competing risk by Latouche A, Boisson V, Chevret S and Porcher R. Statistics in Medicine 2006; DOI:

10.2002/sim.2600. Statistics in Medicine, 26(7):1649–1651, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2298].

Senn:2007:BRS

[2345] Stephen Senn. Book review: Statistics in the pharmaceutical industry. (3rd edn). C. Ralph Buncher and Jia-Yeong Tsay (eds), Chapman and Hall/CRC, New York, 2006. No. of pages: xxiii + 473. Price: $119.95. ISBN: 0-8247-5469-7. Statistics in Medicine, 26(7):1651–1652, March 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2007:TBC

[2346] Anonymous. Tenth Biennial CDC and ATSDR Symposium on Statistical Methods: Symposium Committees. Statistics in Medicine, 26(8):1653– 1654, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Madans:2007:TBC

[2347] Jennifer Madans and L. Pallos. Tenth Biennial CDC and ATSDR Symposium on Statistical Methods: opening remarks. Statistics in Medicine, 26(8):1655–1656, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kalsbeek:2007:DSP

[2348] William D. Kalsbeek, Walter R. Boyle, Robert P. Agans, and John E. White. Disproportionate sampling for population subgroups in telephone surveys. Statistics in Medicine, 26(8):1657–1674, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kalton:2007:SRM

[2349] Graham Kalton and Andrea Piesse. Survey research methods in evaluation and case-control studies. Statistics in Medicine, 26(8):1675–1687, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murphy:2007:MMC

[2350] Joe Murphy, Robert M. Brackbill, Lisa Thalji, Melissa Dolan, Paul Pulliam, and Deborah J. Walker. Measuring and maximizing coverage in the world trade center health registry. Statistics in Medicine, 26(8): 1688–1701, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strickland:2007:MEI

[2351] Pamela A. Ohman Strickland and Benjamin F. Crabtree. Modelling effectiveness of internally heterogeneous organizations in the presence of survey non-response: an application to the ULTRA study. Statistics in Medicine, 26(8):1702–1711, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:2007:SPM

[2352] Andrew C. Leon, Moonseong Heo, Jedediah J. Teres, and Toshihiko Morikawa. Statistical power of multiplicity adjustment strategies for correlated binary endpoints. Statistics in Medicine, 26(8):1712–1723, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2007:ESC

[2353] Zhen Zhao. Early stopping clinical trials of binomial response with an exact group sequential method. Statistics in Medicine, 26(8):1724–1729, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Purkayastha:2007:SMI

[2354] D. Das Purkayastha and Jie Yin. A statistical model for inter-temporal causal dynamics of glucose production and disposal with insulin level in diabetic patients. Statistics in Medicine, 26(8):1730–1744, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tarima:2007:PEW

[2355] Sergey S. Tarima, Svetla S. Slavova, Travis A. Fritsch, and Laura M. Hall. Probability estimation when some observations are grouped. Statistics in Medicine, 26(8):1745–1761, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burt:2007:MPL

[2356] Catharine W. Burt and Esther Hing. Making patient-level estimates from medical encounter records using a multiplicity estimator. Statistics in Medicine, 26(8):1762–1774, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2007:ERB

[2357] Rongwei Fu, Emily L. Harris, Mark Helfand, and Heidi D. Nelson. Estimating risk of breast cancer in carriers of BRCA 1 and BRCA 2 mutations: a meta-analytic approach. Statistics in Medicine, 26(8):1775–1787, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sirken:2007:LSH

[2358] Monroe G. Sirken and Iris M. Shimizu. Linked surveys of health services utilization. Statistics in Medicine, 26(8):1788–1801, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schenker:2007:CIM

[2359] Nathaniel Schenker and Trivellore E. Raghunathan. Combining information from multiple surveys to enhance estimation of measures of health. Statistics in Medicine, 26(8):1802–1811, April 15, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scanlon:2007:SUU

[2360] Joseph Scanlon. Sampling an unknown universe: problems of researching mass casualty incidents (a history of ECRU’s field research). Statistics in Medicine, 26(8):1812–1823, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulldorff:2007:MSS

[2361] Martin Kulldorff, Farzad Mostashari, Luiz Duczmal, W. Katherine Yih, Ken Kleinman, and Richard Platt. Multivariate scan statistics for disease surveillance. Statistics in Medicine, 26(8):1824–1833, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rolka:2007:IAS

[2362] Henry Rolka, Howard Burkom, Gregory F. Cooper, Martin Kulldorff, David Madigan, and Weng-Keen Wong. Issues in applied statistics for public health bioterrorism surveillance using multiple data streams: research needs. Statistics in Medicine, 26(8):1834–1856, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burr:2007:CAD

[2363] Tom Burr, Frederick Koster, Rick Picard, Dave Forslund, Doug Wokoun, Ed Joyce, Judith Brillman, Phil Froman, and Jack Lee. Computer-aided diagnosis with potential application to rapid detection of disease outbreaks. Statistics in Medicine, 26(8):1857–1874, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wun:2007:MRP

[2364] Lap-Ming Wun, Trena M. Ezzati-Rice, Nuria Diaz-Tena, and Janet Greenblatt. On modelling response propensity for dwelling unit (DU) level non-response adjustment in the medical expenditure panel survey (MEPS). Statistics in Medicine, 26(8):1875–1884, April 15, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zaslavsky:2007:UHM

[2365] Alan M. Zaslavsky. Using hierarchical models to attribute sources of variation in consumer assessments of health care. Statistics in Medicine, 26(8):1885–1900, April 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:2007:MMB

[2366] Brent A. Johnson and Stephen M. Rappaport. On modelling metabolismbased biomarkers of exposure: a comparative analysis of nonlinear models with few repeated measurements. Statistics in Medicine, 26(9):1901– 1919, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scarpa:2007:BMS

[2367] Bruno Scarpa and David B. Dunson. Bayesian methods for searching for optimal rules for timing intercourse to achieve pregnancy. Statistics in Medicine, 26(9):1920–1936, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morgan:2007:CAA

[2368] Caroline C. Morgan and D. Stephen Coad. A comparison of adaptive allocation rules for group-sequential binary response clinical trials. Statistics in Medicine, 26(9):1937–1954, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sarkar:2007:QTE

[2369] Somnath Sarkar and Yongming Qu. Quantifying the treatment effect explained by markers in the presence of measurement error. Statistics in Medicine, 26(9):1955–1963, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sidik:2007:CHV

[2370] Kurex Sidik and Jeffrey N. Jonkman. A comparison of heterogeneity variance estimators in combining results of studies. Statistics in Medicine, 26(9):1964–1981, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jewell:2007:NPE

[2371] Nicholas P. Jewell, Xiudong Lei, Azra C. Ghani, Christl A. Donnelly, Gabriel M. Leung, Lai-Ming Ho, Benjamin J. Cowling, and Anthony J. Hedley. Non-parametric estimation of the case fatality ratio with competing risks data: an application to severe acute respiratory syndrome (SARS). Statistics in Medicine, 26(9):1982–1998, April 30, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lopez-Fidalgo:2007:ODR

[2372] J. López-Fidalgo and J. Villarroel. Optimal designs for radiation retention with Poisson correlated response. Statistics in Medicine, 26(9): 1999–2016, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Panageas:2007:PAM

[2373] Katherine S. Panageas, Deborah Schrag, A. Russell Localio, E. S. Venkatraman, and Colin B. Begg. Properties of analysis methods that account for clustering in volume-outcome studies when the primary predictor is cluster size. Statistics in Medicine, 26(9):2017–2035, April 30, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:2007:MBM

[2374] Allan Donner, Monica Taljaard, and Neil Klar. The merits of breaking the matches: a cautionary tale. Statistics in Medicine, 26(9):2036–2051, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diaz:2007:IDD

[2375] Francisco J. Diaz, Tulia E. Rivera, Richard C. Josiassen, and Jose de Leon. Individualizing drug dosage by using a random intercept linear model. Statistics in Medicine, 26(9):2052–2073, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2007:RCM

[2376] Pulak Ghosh and Florin Vaida. Random changepoint modelling of HIV immunologic responses. Statistics in Medicine, 26(9):2074–2087, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ohlssen:2007:FRE

[2377] D. I. Ohlssen, L. D. Sharples, and D. J. Spiegelhalter. Flexible randomeffects models using Bayesian semi-parametric models: applications to institutional comparisons. Statistics in Medicine, 26(9):2088–2112, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Funatogawa:2007:ALM

[2378] Ikuko Funatogawa, Takashi Funatogawa, and Yasuo Ohashi. An autoregressive linear mixed effects model for the analysis of longitudinal data which show profiles approaching asymptotes. Statistics in Medicine, 26 (9):2113–2130, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2007:LES

[2379] Anonymous. Letter to the Editor: A semi-parametric estimate of extra-Poisson variation for vital rates by M. Fay and E. Feuer, Statistics in Medicine 1997; 16:2389–2401. Statistics in Medicine, 26(9):2131–2133, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [8].

Fay:2007:AR

[2380] Michael P. Fay and Eric J. Feuer. Authors’ reply. Statistics in Medicine, 26(9):2133–2135, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2007:TCTb

[2381] Jason Kunming Huang and John Jen Tai. Twin concordances test for ascertained trichotomous traits data. Statistics in Medicine, 26(9):2136, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAngelo:2007:CMD

[2382] Gina D’Angelo and Lisa Weissfeld. Covariates missing by design: comparison of the efficient score to other weighted methods. Statistics in Medicine, 26(10):2137–2153, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2007:TBM

[2383] Xiaogang Su. Tree-based model checking for logistic regression. Statistics in Medicine, 26(10):2154–2169, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mercaldo:2007:CIP

[2384] Nathaniel D. Mercaldo, Kit F. Lau, and Xiao H. Zhou. Confidence intervals for predictive values with an emphasis to case-control studies. Statistics in Medicine, 26(10):2170–2183, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shardell:2007:SCE

[2385] Michelle Shardell, Daniel O. Scharfstein, and Samuel A. Bozzette. Survival curve estimation for informatively coarsened discrete event-time data. Statistics in Medicine, 26(10):2184–2202, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2007:CDC

[2386] Yuh-Ing Chen and Yu-Mei Chang. Covariates-dependent confidence intervals for the difference or ratio of two median survival times. Statistics in Medicine, 26(10):2203–2213, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Remontet:2007:OSB

[2387] L. Remontet, N. Bossard, A. Belot, J. Estève, and the French network of cancer registries Francim. An overall strategy based on regression models to estimate relative survival and model the effects of prognostic factors in cancer survival studies. Statistics in Medicine, 26(10):2214– 2228, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proust-Lima:2007:NLC

[2388] Cécile Proust-Lima, Luc Letenneur, and Hélène Jacqmin-Gadda. A nonlinear latent class model for joint analysis of multivariate longitudinal data and a binary outcome. Statistics in Medicine, 26(10):2229–2245, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berglund:2007:CRD

[2389] Lars Berglund, Hans Garmo, Johan Lindbäck, and Björn Zethelius. Correction for regression dilution bias using replicates from subjects with extreme first measurements. Statistics in Medicine, 26(10):2246–2257, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2007:IPI

[2390] Yang Xie, Wei Pan, Kyeong S. Jeong, and Arkady Khodursky. Incorporating prior information via shrinkage: a combined analysis of genomewide location data and gene expression data. Statistics in Medicine, 26 (10):2258–2275, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2007:C

[2391] Jun Xie and Chaofeng Liu. Correction. Statistics in Medicine, 26(10): 2276, May 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuhnert:2007:CTD

[2392] Ronny Kuhnert and Dankmar Böhning. A comparison of three different models for estimating relative risk in meta-analysis of clinical trials under unobserved heterogeneity. Statistics in Medicine, 26(11):2277–2296, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:2007:STA

[2393] David M. Murray, Jonathan L. Blitstein, Peter J. Hannan, William L. Baker, and Leslie A. Lytle. Sizing a trial to alter the trajectory of health behaviours: methods, parameter estimates, and their application. Statistics in Medicine, 26(11):2297–2316, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mandrekar:2007:APD

[2394] Sumithra J. Mandrekar, Yue Cui, and Daniel J. Sargent. An adaptive phase I design for identifying a biologically optimal dose for dual agent drug combinations. Statistics in Medicine, 26(11):2317–2330, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCandless:2007:BSA

[2395] Lawrence C. McCandless, Paul Gustafson, and Adrian Levy. Bayesian sensitivity analysis for unmeasured confounding in observational studies. Statistics in Medicine, 26(11):2331–2347, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2007:BAS

[2396] Xin-Yuan Song, Sik-Yum Lee, Maggie C. Y. Ng, Wing-Yee So, and Juliana C. N. Chan. Bayesian analysis of structural equation models with multinomial variables and an application to type 2 diabetic nephropathy. Statistics in Medicine, 26(11):2348–2369, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harel:2007:MIC

[2397] Ofer Harel and Xiao-Hua Zhou. Multiple imputation for the comparison of two screening tests in two-phase Alzheimer studies. Statistics in Medicine, 26(11):2370–2388, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Putter:2007:TBC

[2398] H. Putter, M. Fiocco, and R. B. Geskus. Tutorial in biostatistics: competing risks and multi-state models. Statistics in Medicine, 26(11):2389– 2430, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Houseman:2007:BRU

[2399] E. Andres Houseman. Book review: Using R for introductory statistics. John Verzani, Chapman and Hall/CRC, Boca Raton, FL, 2005. No. of pages: xvi + 414. Price: $44.95. ISBN 1-58488-4509. Statistics in Medicine, 26(11):2431–2432, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2007:SMA

[2400] Zonghui Hu and Dean Follmann. Statistical methods for active extension trials. Statistics in Medicine, 26(12):2433–2448, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Timmesfeld:2007:ISS

[2401] Nina Timmesfeld, Helmut Schäfer, and Hans-Helge Müller. Increasing the sample size during clinical trials with t-distributed test statistics without inflating the type I error rate. Statistics in Medicine, 26(12): 2449–2464, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2007:TSG

[2402] Alex Dmitrienko, Brian L. Wiens, Ajit C. Tamhane, and Xin Wang. Tree-structured gatekeeping tests in clinical trials with hierarchically ordered multiple objectives. Statistics in Medicine, 26(12):2465–2478, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sutton:2007:EBS

[2403] Alexander J. Sutton, Nicola J. Cooper, David R. Jones, Paul C. Lambert, John R. Thompson, and Keith R. Abrams. Evidence-based sample size calculations based upon updated meta-analysis. Statistics in Medicine, 26(12):2479–2500, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deCastro:2007:SPM

[2404] B. Rey deCastro and Donna Neuberg. The statistical performance of an MCF-7 cell culture assay evaluated using generalized linear mixed models and a score test. Statistics in Medicine, 26(12):2501–2518, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2007:REM

[2405] Jinhui Li, Xiaowei Yang, Yingnian Wu, and Steven Shoptaw. A randomeffects Markov transition model for Poisson-distributed repeated measures with non-ignorable missing values. Statistics in Medicine, 26(12): 2519–2532, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2007:RAF

[2406] Zhigang Zhang, Liuquan Sun, Jianguo Sun, and Dianne M. Finkelstein. Regression analysis of failure time data with informative interval censoring. Statistics in Medicine, 26(12):2533–2546, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beath:2007:IGM

[2407] Ken J. Beath. Infant growth modelling using a shape invariant model with random effects. Statistics in Medicine, 26(12):2547–2564, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2007:SSR

[2408] Shuwen Wan and Biao Zhang. Smooth semiparametric receiver operating characteristic curves for continuous diagnostic tests. Statistics in Medicine, 26(12):2565–2586, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:2007:BRA

[2409] Michael Jones. Book review: 1. Analysis of clinical trials using SAS(R): a practical guide. Alex Dmitrienko, Geert Molenberghs, Christy Chuang-Stein and Walter Offen, SAS Press, Cary, 2005. No. of pages: vii + 419. Price: $69.95. ISBN 1-59047-504-6. Statistics in Medicine, 26(12): 2587–2588, May 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBreukelen:2007:REU

[2410] Gerard J. P. van Breukelen, Math J. J. M. Candel, and Martijn P. F. Berger. Relative efficiency of unequal versus equal cluster sizes in cluster randomized and multicentre trials. Statistics in Medicine, 26(13):2589– 2603, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2007:OTS

[2411] Myron N. Chang, Meenakshi Devidas, and James Anderson. One- and two-stage designs for phase II window studies. Statistics in Medicine, 26(13):2604–2614, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taljaard:2007:AEA

[2412] Monica Taljaard, Allan Donner, and Neil Klar. Accounting for expected attrition in the planning of community intervention trials. Statistics in Medicine, 26(13):2615–2628, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacobusse:2007:CAT

[2413] Gert Jacobusse and Stef van Buuren. Computerized adaptive testing for measuring development of young children. Statistics in Medicine, 26(13): 2629–2638, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graubard:2007:EAN

[2414] Barry I. Graubard, Katherine M. Flegal, David F. Williamson, and Mitchell H. Gail. Estimation of attributable number of deaths and standard errors from simple and complex sampled cohorts. Statistics in Medicine, 26(13):2639–2649, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:2007:CME

[2415] Andrew C. Leon and Donald Hedeker. A comparison of mixed-effects quantile stratification propensity adjustment strategies for longitudinal treatment effectiveness analyses of continuous outcomes. Statistics in Medicine, 26(13):2650–2665, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perperoglou:2007:AML

[2416] Aris Perperoglou, Antonis Keramopoullos, and Hans C. van Houwelingen. Approaches in modelling long-term survival: an application to breast cancer. Statistics in Medicine, 26(13):2666–2685, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lehr:2007:PAT

[2417] Stephan Lehr and Michael Schemper. Parsimonious analysis of timedependent effects in the Cox model. Statistics in Medicine, 26(13):2686– 2698, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2007:EPO

[2418] Dawei Xie, Trivellore E. Raghunathan, and James M. Lepkowski. Estimation of the proportion of overweight individuals in small areas — a robust extension of the Fay–Herriot model. Statistics in Medicine, 26 (13):2699–2715, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:2007:ECG

[2419] Adele H. Marshall, Barry Shaw, and Sally I. McClean. Estimating the costs for a group of geriatric patients using the Coxian phase-type distribution. Statistics in Medicine, 26(13):2716–2729, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fung:2007:ALI

[2420] Wing K. Fung, Hong Gu, Liming Xiang, and Kelvin K. W. Yau. Assessing local influence in principal component analysis with application to haematology study data. Statistics in Medicine, 26(13):2730–2744, June 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siersma:2007:MMM

[2421] V. Siersma, B. Als-Nielsen, W. Chen, J. Hilden, L. L. Gluud, and C. Gluud. Multivariable modelling for meta-epidemiological assessment of the association between trial quality and treatment effects estimated in randomized clinical trials. Statistics in Medicine, 26(14):2745–2758, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2007:PBD

[2422] W. Liu, M. Jamshidian, Y. Zhang, F. Bretz, and X. L. Han. Pooling batches in drug stability study by using constant-width simultaneous confidence bands. Statistics in Medicine, 26(14):2759–2771, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2007:ADM

[2423] Mark Chang. Adaptive design method based on sum of p-values. Statistics in Medicine, 26(14):2772–2784, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murad:2007:SCD

[2424] H. Murad, O. Kalter-Leibovici, A. Chetrit, and L. S. Freedman. A statistical comparison of different family history scores. Statistics in Medicine, 26(14):2785–2798, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daimon:2007:CAO

[2425] Takashi Daimon and Masashi Goto. Curvature-adjusted optimal design of sampling times for the inference of pharmacokinetic compartment models. Statistics in Medicine, 26(14):2799–2812, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elashoff:2007:AJA

[2426] Robert M. Elashoff, Gang Li, and Ning Li. An approach to joint analysis of longitudinal measurements and competing risks failure time data. Statistics in Medicine, 26(14):2813–2835, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2007:PBE

[2427] X. Joan Hu, R. Jason Schroeder, Winfred C. Wang, and James M. Boyett. Pseudoscore-based estimation from biased observations. Statistics in Medicine, 26(14):2836–2852, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulinskaya:2007:CIS

[2428] Elena Kulinskaya and Robert G. Staudte. Confidence intervals for the standardized effect arising in the comparison of two normal populations. Statistics in Medicine, 26(14):2853–2871, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tutz:2007:BAF

[2429] G. Tutz and F. Reithinger. A boosting approach to flexible semiparametric mixed models. Statistics in Medicine, 26(14):2872–2900, June 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pires:2007:MLR

[2430] Ana M. Pires and Isabel M. Rodrigues. Multiple linear regression with some correlated errors: Classical and robust methods. Statistics in Medicine, 26(15):2901–2918, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blume:2007:SEG

[2431] Jeffrey D. Blume, Li Su, Remigio M. Olveda, and Stephen T. McGarvey. Statistical evidence for GLM regression parameters: a robust likelihood approach. Statistics in Medicine, 26(15):2919–2936, July 10, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2007:CRT

[2432] Peter C. Austin. A comparison of regression trees, logistic regression, generalized additive models, and multivariate adaptive regression splines for predicting AMI mortality. Statistics in Medicine, 26(15):2937–2957, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:2007:PAL

[2433] X. M. Tu, J. Zhang, J. Kowalski, J. Shults, C. Feng, W. Sun, and W. Tang. Power analyses for longitudinal study designs with missing data. Statistics in Medicine, 26(15):2958–2981, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simmonds:2007:CHM

[2434] M. C. Simmonds and J. P. T. Higgins. Covariate heterogeneity in metaanalysis: Criteria for deciding between meta-regression and individual patient data. Statistics in Medicine, 26(15):2982–2999, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sung:2007:BAN

[2435] Minje Sung, Refik Soyer, and Nguyen Nhan. Bayesian analysis of nonhomogeneous Markov chains: Application to mental health data. Statistics in Medicine, 26(15):3000–3017, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carstensen:2007:APC

[2436] B. Carstensen. Age–period–cohort models for the Lexis diagram. Statistics in Medicine, 26(15):3018–3045, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [2734,

Hanley:2007:LEM

[2437] James A. Hanley, Nandini Dendukuri, and Colin B. Begg. Letter to the Editor: Multiple imputation for correcting verification bias by Ofer Harel and Xiao-Hua Zhou, Statistics in Medicine 2006; 25:3769–3786. Statistics in Medicine, 26(15):3046–3047, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2187].

Harel:2007:RMI

[2438] Ofer Harel and Xiao-Hua Zhou. Rejoinder to multiple imputation for correcting verification bias. Statistics in Medicine, 26(15):3047–3050, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chappell:2007:LEC

[2439] Rick Chappell and Theodore Karrison. Letter to the Editor: Continuous Bayesian adaptive randomization based on event times with covariates by Cheung et al., Statistics in Medicine 2006; 25:55–70. Statistics in Medicine, 26(15):3050–3052, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1907].

Anonymous:2007:CCB

[2440] Anonymous. Comments on ‘Continuous Bayesian adaptive randomization based on event times with covariates’ by Y. K. Cheung et al., Statistics in Medicine 2006; 25:55—70. Statistics in Medicine, 26(15):3052– 3054, July 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2007:ECI

[2441] G. Y. Zou. Exact confidence interval for Cohen’s effect size is readily available. Statistics in Medicine, 26(15):3054–3056, July 10, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harel:2007:MIR

[2442] Ofer Harel and Xiao-Hua Zhou. Multiple imputation: review of theory, implementation and software. Statistics in Medicine, 26(16):3057–3077, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2007:PDP

[2443] Peter C. Austin. The performance of different propensity score methods for estimating marginal odds ratios. Statistics in Medicine, 26(16):3078– 3094, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [3024, 2910].

King:2007:RMC

[2444] Tonya S. King, Vernon M. Chinchilli, and Josep L. Carrasco. A repeated measures concordance correlation coefficient. Statistics in Medicine, 26 (16):3095–3113, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2007:EPQ

[2445] Chuanpu Hu and Yingwen Dong. Estimating the predictive quality of dose-response after model selection. Statistics in Medicine, 26(16):3114– 3139, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2007:IER

[2446] Kung-Jong Lui. Interval estimation of the risk difference in noncompliance randomized trials with repeated binary measurements. Statistics in Medicine, 26(16):3140–3156, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2007:NEM

[2447] Jiajia Zhang and Yingwei Peng. A new estimation method for the semiparametric accelerated failure time mixture cure model. Statistics in Medicine, 26(16):3157–3171, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haneuse:2007:IEE

[2448] S. Haneuse, J. Wakefield, and L. Sheppard. The interpretation of exposure effect estimates in chronic air pollution studies. Statistics in Medicine, 26(16):3172–3187, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2007:NPB

[2449] Zhihong Cai, Manabu Kuroki, and Tosiya Sato. Non-parametric bounds on treatment effects with non-compliance by covariate adjustment. Statistics in Medicine, 26(16):3188–3204, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Funatogawa:2007:LER

[2450] Ikuko Funatogawa and Takashi Funatogawa. Letter to the Editor: Revisiting the relation between change and initial value: a review and evaluation, by Yu-Kang Tu and Mark S. Gilthorpe, Statistics in Medicine 2006; DOI: 10.1002/sim.2538. Statistics in Medicine, 26(16):3205–3206, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2260].

Tu:2007:AR

[2451] Yu-Kang Tu and Mark S. Gilthorpe. Authors’ reply. Statistics in Medicine, 26(16):3206–3208, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martens:2007:LEC

[2452] Edwin P. Martens, Wiebe R. Pestman, and Olaf H. Klungel. Letter to the Editor: Conditioning on the propensity score can result in biased estimation of common measures of treatment effect: a Monte Carlo study (p n/a) by Peter C. Austin, Paul Grootendorst, Sharon-Lise T. Normand, Geoffrey M. Anderson, Statistics in Medicine, Published Online: 16 June 2006. DOI: 10.1002/sim.2618. Statistics in Medicine, 26(16):3208–3210, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2282].

Austin:2007:AR

[2453] P. C. Austin, P. Grootendorst, S. L. T. Normand, and G. M. Anderson. Authors’ reply. Statistics in Medicine, 26(16):3210–3212, July 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chetty:2007:USC

[2454] V. K. Chetty and B. Bruce Zellner. Use of survey and clinical data for screening and diagnosis. Statistics in Medicine, 26(17):3213–3228, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Natarajan:2007:SMD

[2455] Sundar Natarajan, Stuart R. Lipsitz, and Eric Rimm. A simple method of determining confidence intervals for population attributable risk from complex surveys. Statistics in Medicine, 26(17):3229–3239, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mukherjee:2007:AMC

[2456] Bhramar Mukherjee, Ivy Liu, and Samiran Sinha. Analysis of matched case-control data with multiple ordered disease states: possible choices and comparisons. Statistics in Medicine, 26(17):3240–3257, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2007:PDD

[2457] Jialiang Li, Jason P. Fine, and Nasia Safdar. Prevalence-dependent diagnostic accuracy measures. Statistics in Medicine, 26(17):3258–3273, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullenayegum:2007:SPR

[2458] Eleanor M. Pullenayegum and Andrew R. Willan. Semi-parametric regression models for cost-effectiveness analysis: improving the efficiency of estimation from censored data. Statistics in Medicine, 26(17):3274–3299, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deLeon:2007:LAE

[2459] A. R. de Leon, M. Guo, C. J. Rudnisky, and G. Singh. A likelihood approach to estimating sensitivity and specificity for binocular data: application in ophthalmology. Statistics in Medicine, 26(17):3300–3314, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kvaloy:2007:MSW

[2460] Jan Terje Kvaløy and Eirik Skogvoll. Modelling seasonal and weather dependency of cardiac arrests using the covariate order method. Statistics in Medicine, 26(17):3315–3329, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koru-Sengul:2007:RBT

[2461] Tulay Koru-Sengul, David S. Stoffer, and Nancy L. Day. A residualsbased transition model for longitudinal analysis with estimation in the presence of missing data. Statistics in Medicine, 26(17):3330–3341, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2007:HVD

[2462] Lang Wu. HIV viral dynamic models with dropouts and missing covariates. Statistics in Medicine, 26(17):3342–3357, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eilers:2007:DEM

[2463] Paul H. C. Eilers. Data exploration in meta-analysis with smooth latent distributions. Statistics in Medicine, 26(17):3358–3368, July 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lloyd:2007:EOS

[2464] Chris J. Lloyd and Max V. Moldovan. Exact one-sided confidence limits for the difference between two correlated proportions. Statistics in Medicine, 26(18):3369–3384, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demidenko:2007:SSD

[2465] Eugene Demidenko. Sample size determination for logistic regression revisited. Statistics in Medicine, 26(18):3385–3397, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2007:EME

[2466] Yan Li, Julie A. Schneider, and David A. Bennett. Estimation of the mediation effect with a binary mediator. Statistics in Medicine, 26 (18):3398–3414, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ukoumunne:2007:SSO

[2467] Obioha C. Ukoumunne, John B. Carlin, and Martin C. Gulliford. A simulation study of odds ratio estimation for binary outcomes from cluster randomized trials. Statistics in Medicine, 26(18):3415–3428, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:2007:BEF

[2468] Robert G. Newcombe. Bayesian estimation of false-negative rate in a clinical trial of sentinel node biopsy. Statistics in Medicine, 26(18):3429– 3442, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:2007:ABR

[2469] Jeremy M. G. Taylor, Lu Wang, and Zhiguo Li. Analysis on binary responses with ordered covariates and missing data. Statistics in Medicine, 26(18):3443–3458, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanayama:2007:EAP

[2470] Nobutane Hanayama. An extended age period cohort model for analysing (age, period)-tabulated data. Statistics in Medicine, 26(18):3459–3475, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:2007:MSA

[2471] A. Donner and G. Y. Zou. Methods for the statistical analysis of binary data in split-mouth designs with baseline measurements. Statistics in Medicine, 26(18):3476–3486, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2007:BOR

[2472] Henian Chen, Patricia Cohen, and Sophie Chen. Biased odds ratios from dichotomization of age. Statistics in Medicine, 26(18):3487–3497, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borm:2007:ETA

[2473] George F. Borm, Marten Munneke, Oscar Lemmers, and Martien van Zuijlen. An efficient test for the analysis of dichotomized variables when the reliability is known. Statistics in Medicine, 26(18):3498–3510, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:2007:RRA

[2474] Geoffrey Berry. Relative risk and acceleration in lung cancer. Statistics in Medicine, 26(18):3511–3517, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nadarajah:2007:SDD

[2475] Saralees Nadarajah and Samuel Kotz. Statistical distribution of the difference of two proportions. Statistics in Medicine, 26(18):3518–3521, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wolbers:2007:CAI

[2476] Marcel Wolbers and Michael Koller. Comments on ‘Analysing and interpreting competing risk data’ (original article and author’s reply) by Melania Pintilie, Statistics in Medicine 2007; 26:1360–1367; DOI: 10.1002/sim.2655 (original article) and Statistics in Medicine 2007; DOI: 10.1002/sim.2822 (author’s reply). Statistics in Medicine, 26(18):3521– 3523, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2323, 2489].

Pintilie:2007:ARa

[2477] Melania Pintilie. Author’s reply. Statistics in Medicine, 26(18):3523, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Acion:2007:C

[2478] Laura Acion, John J. Peterson, Scott Temple, and Stephan Arndt. Correction. Statistics in Medicine, 26(18):3524, August 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hong:2007:TOD

[2479] Shengyan Hong and Yanping Wang. A three-outcome design for randomized comparative phase II clinical trials. Statistics in Medicine, 26 (19):3525–3534, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2007:MTP

[2480] Yang Song and George Y. H. Chi. A method for testing a prespecified subgroup in clinical trials. Statistics in Medicine, 26(19):3535–3549, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2007:CSP

[2481] Peter C. Austin. A comparison of the statistical power of different methods for the analysis of cluster randomization trials with binary outcomes. Statistics in Medicine, 26(19):3550–3565, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alexander:2007:RPC

[2482] Neal Alexander, Jeff Bethony, Rodrigo Corrêa-Oliveira, Laura C. Rodrigues, Peter Hotez, and Simon Brooker. Repeatability of paired counts. Statistics in Medicine, 26(19):3566–3577, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2007:PDN

[2483] Sander Greenland. Prior data for non-normal priors. Statistics in Medicine, 26(19):3578–3590, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2007:BMD

[2484] Q. Li, X. Shen, and D. K. Pearl. Bayesian modelling of the dynamics of hepatotoxicity. Statistics in Medicine, 26(19):3591–3611, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2007:MSS

[2485] Jichuan Wang, Robert G. Carlson, Russel S. Falck, Carl Leukefeld, and Brenda M. Booth. Multi-sample standardization and decomposition analysis: an application to comparisons of methamphetamine use among rural drug users in three American states. Statistics in Medicine, 26 (19):3612–3623, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2007:ESC

[2486] Yujun Wu, Weichung J. Shih, and Dirk F. Moore. Estimating a survival curve with unlinked entry and failure times. Statistics in Medicine, 26 (19):3624–3638, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muller:2007:STA

[2487] Keith E. Muller, Lloyd J. Edwards, Sean L. Simpson, and Douglas J. Taylor. Statistical tests with accurate size and power for balanced linear mixed models. Statistics in Medicine, 26(19):3639–3660, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:2007:CSF

[2488] Ian Campbell. Chi-squared and Fisher–Irwin tests of two-by-two tables with small sample recommendations. Statistics in Medicine, 26(19):3661– 3675, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2751].

Latouche:2007:CAI

[2489] A. Latouche, J. Beyersmann, and J. P. Fine. Comments on ‘Analysing and interpreting competing risk data’ by M. Pintilie, Statistics in Medicine 2006. DOI: 10.1002/sim.2655. Statistics in Medicine, 26(19): 3676–3679, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2323, 2476].

Pintilie:2007:ARb

[2490] Melania Pintilie. Author’s reply. Statistics in Medicine, 26(19):3679– 3680, August 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2007:MAM

[2491] G. Lu, A. E. Ades, A. J. Sutton, N. J. Cooper, A. H. Briggs, and D. M. Caldwell. Meta-analysis of mixed treatment comparisons at multiple follow-up times. Statistics in Medicine, 26(20):3681–3699, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2007:DDI

[2492] Lang Li, Menggang Yu, Raymond Chin, Aroonrut Lucksiri, David A. Flockhart, and Stephen D. Hall. Drug-drug interaction prediction: a Bayesian meta-analysis approach. Statistics in Medicine, 26(20):3700– 3721, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Locatelli:2007:CFM

[2493] Isabella Locatelli, Alessandro Rosina, Paul Lichtenstein, and Anatoli I. Yashin. A correlated frailty model with long-term survivors for estimating the heritability of breast cancer. Statistics in Medicine, 26(20): 3722–3734, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Govindarajulu:2007:CST

[2494] Usha S. Govindarajulu, Donna Spiegelman, Sally W. Thurston, Bhaswati Ganguli, and Ellen A. Eisen. Comparing smoothing techniques in Cox models for exposure-response relationships. Statistics in Medicine, 26 (20):3735–3752, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leng:2007:PCM

[2495] Chenlei Leng and Shuangge Ma. Path consistent model selection in additive risk model via lasso. Statistics in Medicine, 26(20):3753–3770, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2007:VSP

[2496] Wenbin Lu and Hao H. Zhang. Variable selection for proportional odds model. Statistics in Medicine, 26(20):3771–3781, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2007:RMM

[2497] Ying Yang, Jian Kang, Kai Mao, and Jie Zhang. Regression models for mixed Poisson and continuous longitudinal data. Statistics in Medicine, 26(20):3782–3800, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boos:2007:NCB

[2498] Dennis D. Boos, David Hoffman, Robert Kringle, and Ji Zhang. New confidence bounds for QT studies. Statistics in Medicine, 26(20):3801– 3817, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3302].

Demirtas:2007:LED

[2499] Hakan Demirtas. Letter to the Editor: The design of simulation studies in medical statistics by Andrea Burton, Douglas G. Altman, Patrick Royston and Roger L. Holder, Statistics in Medicine 2006; 25:4279–4292. Statistics in Medicine, 26(20):3818–3821, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2224].

Cuzick:2007:C

[2500] Jack Cuzick. Correction. Statistics in Medicine, 26(20):3821, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2007:BRS

[2501] Alex Dmitrienko. Book review: Statistical monitoring of clinical trials: a unified approach. Michael A. Proschan, K. K. Gordon Lan and Janet Turk Wittes, Springer, New York, 2006. No. of pages: xiv + 258. Price: $74.95, £50.00, EUR69.50. ISBN: 978-0-387-30059-7. Statistics in Medicine, 26 (20):3822–3824, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robinson:2007:BRA

[2502] Andrew Robinson. Book review: Applied longitudinal analysis. Garrett M. Fitzmaurice, Nan M. Laird, James H. Ware, Wiley-Interscience, Hoboken, NJ, 2004. No. of pages: xix + 506. Price: $105.00. ISBN 0-471-21487-6. Statistics in Medicine, 26(20):3824–3825, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julian:2007:BRS

[2503] Jim A. Julian. Book reviews: Statistical evidence in medical trials: What do the data really tell us? Stephen D. Simon, Oxford University Press, Oxford, 2006. No. of pages: xvii + 197. Price: £65.00, $114.50 hardcover; £25.00, $44.50 paperback. ISBN 0-19-856760-X (hardcover), ISBN 0-19-856761-8 (paperback). Statistics in Medicine, 26(20):3825–3826, September 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:2007:SIS

[2504] Emery N. Brown and Robert E. Kass. Special issue on ‘statistical analysis of neuronal data’. Statistics in Medicine, 26(21):3827–3829, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vakorin:2007:CFM

[2505] Vasily A. Vakorin, Ron Borowsky, and Gordon E. Sarty. Characterizing the functional MRI response using Tikhonov regularization. Statistics in Medicine, 26(21):3830–3844, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2007:CSO

[2506] C. M. Zhang, Y. Jiang, and T. Yu. A comparative study of one-level and two-level semiparametric estimation of hemodynamic response function for fMRI data. Statistics in Medicine, 26(21):3845–3861, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Myers:2007:SSM

[2507] Kary L. Myers, Anthony E. Brockwell, and William F. Eddy. Statespace models for optical imaging. Statistics in Medicine, 26(21):3862– 3874, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bressler:2007:CFN

[2508] Steven L. Bressler, Craig G. Richter, Yonghong Chen, and Mingzhou Ding. Cortical functional network organization from autoregressive modeling of local field potential oscillations. Statistics in Medicine, 26(21): 3875–3885, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagarajan:2007:PAR

[2509] Srikantan S. Nagarajan, Hagai T. Attias, Kenneth E. Hild II, and Kensuke Sekihara. A probabilistic algorithm for robust interference suppression in bioelectromagnetic sensor data. Statistics in Medicine, 26(21): 3886–3910, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thorpe:2007:IWL

[2510] Samuel Garrett Thorpe, Paul L. Nunez, and Ramesh Srinivasan. Identification of wave-like spatial structure in the SSVEP: Comparison of simultaneous EEG and MEG. Statistics in Medicine, 26(21):3911–3926, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2007:MHD

[2511] Xiaofeng Wang, Jiayang Sun, Kenneth J. Gustafson, and Guang H. Yue. Modeling heterogeneity and dependence for analysis of neuronal data. Statistics in Medicine, 26(21):3927–2945, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Winter:2007:CEV

[2512] William R. Winter, Paul L. Nunez, Jian Ding, and Ramesh Srinivasan. Comparison of the effect of volume conduction on EEG coherence with the effect of field spread on MEG coherence. Statistics in Medicine, 26 (21):3946–3957, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Behseta:2007:TES

[2513] Sam Behseta, Robert E. Kass, David E. Moorman, and Carl R. Olson. Testing equality of several functions: Analysis of single-unit firing-rate curves across multiple experimental conditions. Statistics in Medicine, 26(21):3958–3975, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ito:2007:BST

[2514] Hiroyuki Ito. Bootstrap significance test of synchronous spike events — a case study of oscillatory spike trains. Statistics in Medicine, 26(21): 3976–3996, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kruskal:2007:BCM

[2515] Peter B. Kruskal, Jessica J. Stanis, Bruce L. McNaughton, and Peter J. Thomas. A binless correlation measure reduces the variability of memory reactivation estimates. Statistics in Medicine, 26(21):3997–4008, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sharpee:2007:CIV

[2516] Tatyana O. Sharpee. Comparison of information and variance maximization strategies for characterizing neural feature selectivity. Statistics in Medicine, 26(21):4009–4031, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shinomoto:2007:SCB

[2517] Shigeru Shinomoto and Shinsuke Koyama. A solution to the controversy between rate and temporal coding. Statistics in Medicine, 26(21):4032– 4038, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vu:2007:CAE

[2518] Vincent Q. Vu, Bin Yu, and Robert E. Kass. Coverage-adjusted entropy estimation. Statistics in Medicine, 26(21):4039–4060, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petkova:2007:BRM

[2519] Eva Petkova. Book review: Models for discrete data, revised edition. Daniel Zelterman, Oxford University Press Inc., New York, 2006. No. of pages: x + 285. Price: $50.00. ISBN: 0-19-856701-4. Statistics in Medicine, 26(21):4061–4062, September 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hommel:2007:PSC

[2520] Gerhard Hommel, Frank Bretz, and Willi Maurer. Powerful short-cuts for multiple testing procedures with special reference to gatekeeping strategies. Statistics in Medicine, 26(22):4063–4073, September 30, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freidlin:2007:RDO

[2521] Boris Freidlin and Edward L. Korn. Release of data from an ongoing randomized clinical trial for sample size adjustment or planning. Statistics in Medicine, 26(22):4074–4082, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gurka:2007:IPC

[2522] Matthew J. Gurka, Christopher S. Coffey, and Keith E. Muller. Internal pilots for a class of linear mixed models with Gaussian and compound symmetric data. Statistics in Medicine, 26(22):4083–4099, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teerenstra:2007:CMA

[2523] S. Teerenstra, M. Moerbeek, R. J. F. Melis, and G. F. Borm. A comparison of methods to analyse continuous data from pseudo cluster randomized trials. Statistics in Medicine, 26(22):4100–4115, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:2007:CAL

[2524] X. M. Tu, C. Feng, J. Kowalski, W. Tang, H. Wang, C. Wan, and Y. Ma. Correlation analysis for longitudinal data: applications to HIV and psychosocial research. Statistics in Medicine, 26(22):4116–4138, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burnside:2007:AMM

[2525] G. Burnside, C. M. Pine, and P. R. Williamson. The application of multilevel modelling to dental caries data. Statistics in Medicine, 26 (22):4139–4149, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rajeswaran:2007:IEI

[2526] Jeevanantham Rajeswaran and Eugene H. Blackstone. Interval estimation for individual categories in cumulative logit models. Statistics in Medicine, 26(22):4150–4162, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anisimov:2007:ARM

[2527] Vladimir V. Anisimov, Hugo J. Maas, Meindert Danhof, and Oscar Della Pasqua. Analysis of responses in migraine modelling using hidden Markov models. Statistics in Medicine, 26(22):4163–4178, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nofuentes:2007:CLR

[2528] José Antonio Roldán Nofuentes and Juan de Dios Luna del Castillo. Comparison of the likelihood ratios of two binary diagnostic tests in paired designs. Statistics in Medicine, 26(22):4179–4201, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burkom:2007:ATS

[2529] Howard S. Burkom, Sean Patrick Murphy, and Galit Shmueli. Automated time series forecasting for biosurveillance. Statistics in Medicine, 26(22): 4202–4218, September 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2007:SSC

[2530] Yongzhao Shao and Chi-Hong Tseng. Sample size calculation with dependence adjustment for FDR-control in microarray studies. Statistics in Medicine, 26(23):4219–4237, October 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2007:NNR

[2531] Guo-Liang Tian, Jun-Wu Yu, Man-Lai Tang, and Zhi Geng. A new non-randomized model for analysing sensitive questions with binary outcomes. Statistics in Medicine, 26(23):4238–4252, October 15, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kwong:2007:TVC

[2532] Koon Shing Kwong, Siu Hung Cheung, Burt Holland, and Yanhui Wang. Three p-value consistent procedures for multiple comparisons with a control in direction-mixed families. Statistics in Medicine, 26(23):4253–4272, October 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2007:TPM

[2533] Lu Tian and Jie Huang. A two-part model for censored medical cost data. Statistics in Medicine, 26(23):4273–4292, October 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murad:2007:ETI

[2534] Havi Murad and Laurence S. Freedman. Estimating and testing interactions in linear regression models when explanatory variables are subject to classical measurement error. Statistics in Medicine, 26(23):4293–4310, October 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2007:IRS

[2535] Xinhua Liu and Zhezhen Jin. Item reduction in a scale for screening. Statistics in Medicine, 26(23):4311–4327, October 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lydersen:2007:CTA

[2536] S. Lydersen, V. Pradhan, P. Senchaudhuri, and P. Laake. Choice of test for association in small sample unordered r × c tables. Statistics in Medicine, 26(23):4328–4343, October 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Herman:2007:SEM

[2537] Amir Herman, Netta Notzer, Zipi Libman, Rony Braunstein, and David M. Steinberg. Statistical education for medical students — concepts are what remain when the details are forgotten. Statistics in Medicine, 26(23):4344–4351, October 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2787].

Cox:2007:PSA

[2538] Christopher Cox, Haitao Chu, Michael F. Schneider, and Alvaro Muñoz. Parametric survival analysis and taxonomy of hazard functions for the generalized gamma distribution. Statistics in Medicine, 26(23):4352– 4374, October 15, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2007:FVR

[2539] Jonathan J. Shuster, Lynn S. Jones, and Daniel A. Salmon. Fixed vs random effects meta-analysis in rare event studies: The Rosiglitazone link with myocardial infarction and cardiac death. Statistics in Medicine, 26(24):4375–4385, October 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2907].

Bandyopadhyay:2007:CAT

[2540] Uttam Bandyopadhyay, Atanu Biswas, and Rahul Bhattacharya. A covariate adjusted two-stage allocation design for binary responses in randomized clinical trials. Statistics in Medicine, 26(24):4386–4399, October 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Litwin:2007:ESD

[2541] Samuel Litwin, Yu-Ning Wong, and Gary Hudes. Early stopping designs based on progression-free survival at an early time point in the initial cohort. Statistics in Medicine, 26(24):4400–4415, October 30, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shu:2007:SEM

[2542] Y. Shu, A. Liu, and Z. Li. Sequential evaluation of a medical diagnostic test with binary outcomes. Statistics in Medicine, 26(24):4416–4427, October 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:2007:ITM

[2543] Edward L. Korn, Ming-Chung Li, Lisa M. McShane, and Richard Simon. An investigation of two multivariate permutation methods for controlling the false discovery proportion. Statistics in Medicine, 26(24):4428–4440, October 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flandre:2007:PSJ

[2544] Philippe Flandre and John O’Quigley. Predictive strength of Jonckheere’s test for trend: an application to genotypic scores in HIV infection. Statistics in Medicine, 26(24):4441–4454, October 30, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2007:RBS

[2545] Ying C. MacNab and Paul Gustafson. Regression B-spline smoothing in Bayesian disease mapping: with an application to patient safety surveillance. Statistics in Medicine, 26(24):4455–4474, October 30, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yip:2007:EVE

[2546] Paul S. F. Yip, Ray Watson, and Qizhi Chen. Estimation of vaccine efficacy and the vaccination threshold. Statistics in Medicine, 33(27):4475– 4488, October 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ocana-Riola:2007:MCD

[2547] Ricardo Ocaña-Riola. The misuse of count data aggregated over time for disease mapping. Statistics in Medicine, 33(27):4489–4504, October 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klein:2007:ASC

[2548] John P. Klein, Brent Logan, Mette Harhoff, and Per Kragh Andersen. Analyzing survival curves at a fixed point in time. Statistics in Medicine, 33(27):4505–4519, October 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2007:ESM

[2549] Hongwei Zhao, Heejung Bang, Hongkun Wang, and Phillip E. Pfeifer. On the equivalence of some medical cost estimators with censored data. Statistics in Medicine, 33(27):4520–4530, October 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brockwell:2007:SMI

[2550] Sarah E. Brockwell and Ian R. Gordon. A simple method for inference on an overall effect in meta-analysis. Statistics in Medicine, 26(25):4531– 4543, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peters:2007:PTF

[2551] Jaime L. Peters, Alex J. Sutton, David R. Jones, Keith R. Abrams, and Lesley Rushton. Performance of the trim and fill method in the presence of publication bias and between-study heterogeneity. Statistics in Medicine, 26(25):4544–4562, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2007:NIT

[2552] Richard J. Cook, Ker-Ai Lee, and Hongdan Li. Non-inferiority trial design for recurrent events. Statistics in Medicine, 26(25):4563–4577, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:2007:ITR

[2553] Maya L. Petersen, Steven G. Deeks, and Mark J. van der Laan. Individualized treatment rules: Generating candidate clinical trials. Statistics in Medicine, 26(25):4578–4601, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2007:IRR

[2554] Byron J. Gajewski, Sara Hart, Sandra Bergquist-Beringer, and Nancy Dunton. Inter-rater reliability of pressure ulcer staging: ordinal probit Bayesian hierarchical model that allows for uncertain rater response. Statistics in Medicine, 26(25):4602–4618, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2007:IRL

[2555] Brian J. Smith, Lixun Zhang, and R. William Field. Iowa radon leukaemia study: a hierarchical population risk model for spatially correlated exposure measured with error. Statistics in Medicine, 26(25):

4619–4642, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schipper:2007:BGM

[2556] Matthew Schipper, Jeremy M. G. Taylor, and Xihong Lin. Bayesian generalized monotonic functional mixed models for the effects of radiation dose histograms on normal tissue complications. Statistics in Medicine, 26(25):4643–4656, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2007:RMR

[2557] Li Hsu, Malka Gorfine, and Kathleen Malone. On robustness of marginal regression coefficient estimates and hazard functions in multivariate survival analysis of family data when the frailty distribution is mis-specified. Statistics in Medicine, 26(25):4657–4678, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2007:LEC

[2558] Kung-Jong Lui. Letter to the Editor: Correcting for non-compliance of repeated binary outcomes in randomized clinical trials: randomized analysis approach by Yutaka Matsuyama, Statistics in Medicine 2002; 21:675–687. Statistics in Medicine, 26(25):4679–4683, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [719].

Newcombe:2007:CCI

[2559] Robert G. Newcombe. Comments on ‘Confidence intervals for a ratio of binomial proportions based on paired data’ by D. G. Bonett and R. M. Price, Statistics in Medicine 2006; 25:3039–3047. Statistics in Medicine, 26(25):4684–4685, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2122].

Shinomoto:2007:C

[2560] Shigeru Shinomoto and Shinkuke Koyama. Correction. Statistics in Medicine, 26(25):4686, November 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:2007:BFT

[2561] Peter F. Thall, Leiko H. Wooten, Christopher J. Logothetis, Randall E. Millikan, and Nizar M. Tannir. Bayesian and frequentist two-stage treatment strategies based on sequential failure times subject to interval censoring. Statistics in Medicine, 26(26):4687–4702, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quan:2007:ABC

[2562] Hui Quan, Daowen Zhang, Ji Zhang, and Laure Devlamynck. Analysis of a binary composite endpoint with missing data in components. Statistics in Medicine, 26(26):4703–4718, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunn:2007:MTE

[2563] Graham Dunn and Richard Bentall. Modelling treatment-effect heterogeneity in randomized controlled trials of complex interventions (psychological treatments). Statistics in Medicine, 26(26):4719–4745, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2007:MDA

[2564] Ying C. MacNab. Mapping disability-adjusted life years: a Bayesian hierarchical model framework for burden of disease and injury assessment. Statistics in Medicine, 26(26):4746–4769, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sonesson:2007:CFD

[2565] Christian Sonesson. A CUSUM framework for detection of space-time disease clusters using scan statistics. Statistics in Medicine, 26(26):4770– 4789, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ha:2007:MSM

[2566] Il Do Ha, Youngjo Lee, and Gilbert MacKenzie. Model selection for multi-component frailty models. Statistics in Medicine, 26(26):4790– 4807, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Betensky:2007:TAU

[2567] Rebecca A. Betensky, David N. Louis, and J. Gregory Cairncross. Tests of association under misclassification: Application to histological sampling in oncology. Statistics in Medicine, 26(26):4808–4816, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2007:AFT

[2568] Wenqing He, Grace Y. Yi, and Juan Xiong. Accelerated failure time models with covariates subject to measurement error. Statistics in Medicine, 26(26):4817–4832, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2007:EEM

[2569] S. D. Walter, C. C. Hsieh, and Qin Liu. Effect of exposure misclassification on the mean squared error of population attributable risk and prevented fraction estimates. Statistics in Medicine, 26(26):4833–4842, November 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:2007:FSS

[2570] Marc Buyse and Sylvain Nicolas. French Society of Statistics, Biopharmacy and Health Group Fifth International Meeting on Statistical Methods in Biopharmacy ‘Statistical Innovations in Clinical Trials’ Opening remarks. Statistics in Medicine, 26(27):4843–4844, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Travison:2007:GEE

[2571] T. G. Travison and R. Brookmeyer. Global effects estimation for multidimensional outcomes. Statistics in Medicine, 26(27):4845–4859, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samson:2007:SAG

[2572] Adeline Samson, Marc Lavielle, and France Mentré. The SAEM algorithm for group comparison tests in longitudinal data analysis based on non-linear mixed-effects model. Statistics in Medicine, 26(27):4860–4875, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vandenhende:2007:JDI

[2573] F. Vandenhende, P. Eilers, E. Ledent, D. Renard, and F. Tibaldi. Joint detection of important biomarkers and optimal dose-response model using penalties. Statistics in Medicine, 26(27):4876–4888, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sebille:2007:SAL

[2574] Véronique Sébille, Jean-Benoit Hardouin, and Mounir Mesbah. Sequential analysis of latent variables using mixed-effect latent variable models: Impact of non-informative and informative missing data. Statistics in Medicine, 26(27):4889–4904, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benda:2007:SCS

[2575] Norbert Benda and Christoph Gerlinger. Sperm count as a surrogate endpoint for male fertility control. Statistics in Medicine, 26(27):4905– 4913, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grouin:2007:BSS

[2576] Jean-Marie Grouin, Maylis Coste, Pierre Bunouf, and Bruno Lecoutre. Bayesian sample size determination in non-sequential clinical trials: Statistical aspects and some regulatory considerations. Statistics in Medicine, 26(27):4914–4924, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmidli:2007:BPP

[2577] Heinz Schmidli, Frank Bretz, and Amy Racine-Poon. Bayesian predictive power for interim adaptation in seamless phase II/III trials where the endpoint is survival up to some specified timepoint. Statistics in Medicine, 26(27):4925–4938, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:2007:IDT

[2578] Christopher R. Palmer and Harutyun Shahumyan. Implementing a decision-theoretic design in clinical trials: Why and how? Statistics in Medicine, 26(27):4939–4957, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anisimov:2007:MPA

[2579] Vladimir V. Anisimov and Valerii V. Fedorov. Modelling, prediction and adaptive adjustment of recruitment in multicentre trials. Statistics in Medicine, 26(27):4958–4975, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Patel:2007:BAI

[2580] Nitin R. Patel and Suresh Ankolekar. A Bayesian approach for incorporating economic factors in sample size design for clinical trials of individual drugs and portfolios of drugs. Statistics in Medicine, 26(27): 4976–4988, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauerbrei:2007:MEM

[2581] Willi Sauerbrei and Patrick Royston. Modelling to extract more information from clinical trials data: On some roles for the bootstrap. Statistics in Medicine, 26(27):4989–5001, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Desseaux:2007:FTS

[2582] K. Desseaux and R. Porcher. Flexible two-stage design with sample size reassessment for survival trials. Statistics in Medicine, 26(27):5002–5013, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2007:CDT

[2583] John M. Lachin and Naji Younes. A composite design for transition from a preliminary to a full-scale study. Statistics in Medicine, 26(27):5014– 5032, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gicquel:2007:RPC

[2584] Stéphanie Gicquel and Roland Marion-Gallois. Randomization with a posteriori constraints: Description and properties. Statistics in Medicine, 26(27):5033–5045, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2819].

Wu:2007:C

[2585] Yujun Wu, Weichung J. Shih, and Dirk F. Moore. Correction. Statistics in Medicine, 26(27):5046, November 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emerson:2007:FEG

[2586] Scott S. Emerson, John M. Kittelson, and Daniel L. Gillen. Frequentist evaluation of group sequential clinical trial designs. Statistics in Medicine, 26(28):5047–5080, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hardoon:2007:CML

[2587] Sarah L. Hardoon, James D. Lewsey, and Jan H. P. van der Meulen. Continuous monitoring of long-term outcomes with application to hip prostheses. Statistics in Medicine, 26(28):5081–5099, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2007:BPC

[2588] Julia Y. Lin, Thomas R. Ten Have, Hillary R. Bogner, and Michael R. Elliott. Baseline patient characteristics and mortality associated with longitudinal intervention compliance. Statistics in Medicine, 26(28):5100– 5115, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:2007:CMT

[2589] David Edwards and Jesper Madsen. Constructing multiple test procedures for partially ordered hypothesis sets. Statistics in Medicine, 26 (28):5116–5124, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiba:2007:BPR

[2590] Yasutaka Chiba, Tosiya Sato, and Sander Greenland. Bounds on potential risks and causal risk differences under assumptions about confounding parameters. Statistics in Medicine, 26(28):5125–5135, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lloyd:2007:UEO

[2591] Chris J. Lloyd and Max V. Moldovan. Unconditional efficient onesided confidence limits for the odds ratio based on conditional likelihood. Statistics in Medicine, 26(28):5136–5146, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2007:POM

[2592] Man-Hua Chen, Xingwei Tong, and Jianguo Sun. The proportional odds model for multivariate interval-censored failure time data. Statistics in Medicine, 26(28):5147–5161, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Retout:2007:DNM

[2593] Sylvie Retout, Emmanuelle Comets, Adeline Samson, and France Mentré. Design in nonlinear mixed effects models: Optimization using the Fedorov–Wynn algorithm and power of the Wald test for binary covariates. Statistics in Medicine, 26(28):5162–5179, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2007:IMR

[2594] Lili Tian and Jianrong Wu. Inferences on the mean response in a log-regression model: The generalized variable approach. Statistics in Medicine, 26(28):5180–5188, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Solis-Trapala:2007:DMS

[2595] I. L. Solis-Trapala, J. Carthey, V. T. Farewell, and M. R. de Leval. Dynamic modelling in a study of surgical error management. Statistics in Medicine, 26(28):5189–5202, December 10, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mohtashemi:2007:MSA

[2596] Mojdeh Mohtashemi, Ken Kleinman, and W. Katherine Yih. Multisyndrome analysis of time series using PCA: a new concept for outbreak investigation. Statistics in Medicine, 26(29):5203–5224, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hogan:2007:BAR

[2597] William R. Hogan, Gregory F. Cooper, Garrick L. Wallstrom, Michael M. Wagner, and Jean-Marc Depinay. The Bayesian aerosol release detector: an algorithm for detecting and characterizing outbreaks caused by an atmospheric release of Bacillus anthracis. Statistics in Medicine, 26(29): 5225–5252, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crotty:2007:MAE

[2598] Suzanne Crotty and Kingshuk Roy Choudhury. Morphometric analysis for early detection of changes in cellular structure in a toxicological experiment. Statistics in Medicine, 26(29):5253–5266, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Congdon:2007:SSE

[2599] Peter Congdon, Michael Almog, Sarah Curtis, and Raymond Ellerman. A spatial structural equation modelling framework for health count responses. Statistics in Medicine, 26(29):5267–5284, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2007:PLC

[2600] Jun Han, Elizabeth H. Slate, and Edsel A. Peña. Parametric latent class joint model for a longitudinal biomarker and recurrent events. Statistics in Medicine, 26(29):5285–5302, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2007:WNM

[2601] Chaofeng Liu, Jun Xie, and Ying Zhang. Weighted nonparametric maximum likelihood estimate of a mixing distribution in nonrandomized clinical trials. Statistics in Medicine, 26(29):5303–5319, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2007:CBM

[2602] Wenyu Jiang and Richard Simon. A comparison of bootstrap methods and an adjusted bootstrap approach for estimating the prediction error in microarray classification. Statistics in Medicine, 26(29):5320–5334, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:2007:MRR

[2603] Shin-Liang Pan, Hui-Min Wu, Amy Ming-Fang Yen, and Tony Hsiu-Hsi Chen. A Markov regression random-effects model for remission of functional disability in patients following a first stroke: a Bayesian approach. Statistics in Medicine, 26(29):5335–5353, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boscardin:2007:BRS

[2604] W. John Boscardin. Book reviews: SAS(R) for mixed models (2nd edn). Ramon C. Littell, George A. Milliken, Walter W. Stroup, Russell D. Wolfinger and Oliver Schabenberger, SAS Institute, Cary, NC, 2006. No. of pages: xii + 814. Price: $89.95. ISBN 10: 1-59047-500-3; ISBN 13: 978-1-59047-500-3. Statistics in Medicine, 26(29):5354–5355, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edler:2007:P

[2605] Lutz Edler. Preface. Statistics in Medicine, 26(30):5357–5359, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beyersmann:2007:CRA

[2606] Jan Beyersmann, Markus Dettenkofer, Hartmut Bertz, and Martin Schumacher. A competing risks analysis of bloodstream infection after stemcell transplantation using subdistribution hazards and cause-specific hazards. Statistics in Medicine, 26(30):5360–5369, December 30, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Latouche:2007:SSC

[2607] A. Latouche and R. Porcher. Sample size calculations in the presence of competing risks. Statistics in Medicine, 26(30):5370–5380, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foucher:2007:SMM

[2608] Yohann Foucher, Magali Giral, Jean-Paul Soulillou, and Jean-Pierre Daures. A semi-Markov model for multistate and interval-censored data with multiple terminal events. application in renal transplantation. Statistics in Medicine, 26(30):5381–5393, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kvist:2007:NPE

[2609] Kajsa Kvist, Mette Gerster, Per Kragh Andersen, and Lars Vedel Kessing. Non-parametric estimation and model checking procedures for marginal gap time distributions for recurrent events. Statistics in Medicine, 26(30):5394–5410, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deslandes:2007:ASJ

[2610] Emmanuelle Deslandes and Sylvie Chevret. Assessing surrogacy from the joint modelling of multivariate longitudinal data and survival: Application to clinical trial data on chronic lymphocytic leukaemia. Statistics in Medicine, 26(30):5411–5421, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:2007:RCI

[2611] Cyrus R. Mehta, Peter Bauer, Martin Posch, and Werner Brannath. Repeated confidence intervals for adaptive group sequential trials. Statistics in Medicine, 26(30):5422–5433, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghidey:2007:SPM

[2612] W. Ghidey, E. Lesaffre, and T. Stijnen. Semi-parametric modelling of the distribution of the baseline risk in meta-analysis. Statistics in Medicine, 26(30):5434–5444, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickering:2007:ACO

[2613] Ruth M. Pickering and Mark Weatherall. The analysis of continuous outcomes in multi-centre trials with small centre sizes. Statistics in Medicine, 26(30):5445–5456, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Komarek:2007:BTE

[2614] Arnost Komárek, Emmanuel Lesaffre, and Catherine Legrand. Baseline and treatment effect heterogeneity for survival times between centers using a random effects accelerated failure time model with flexible error distribution. Statistics in Medicine, 26(30):5457–5472, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pantazis:2007:RPM

[2615] Nikos Pantazis and Giota Touloumi. Robustness of a parametric model for informatively censored bivariate longitudinal data under misspecification of its distributional assumptions: a simulation study. Statistics in Medicine, 26(30):5473–5485, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:2007:FPM

[2616] Christopher P. Nelson, Paul C. Lambert, Iain B. Squire, and David R. Jones. Flexible parametric models for relative survival, with application in coronary heart disease. Statistics in Medicine, 26(30):5486–5498, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Konig:2007:PEN

[2617] I. R. König, J. D. Malley, C. Weimar, H.-C. Diener, and A. Ziegler. Practical experiences on the necessity of external validation. Statistics in Medicine, 26(30):5499–5511, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauerbrei:2007:SIV

[2618] Willi Sauerbrei, Patrick Royston, and Harald Binder. Selection of important variables and determination of functional form for continuous predictors in multivariable model building. Statistics in Medicine, 26(30): 5512–5528, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2007:SPS

[2619] Stephen Senn, Erika Graf, and Angelika Caputo. Stratification for the propensity score compared with linear regression techniques to assess the effect of treatment or exposure. Statistics in Medicine, 26(30):5529– 5544, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steyerberg:2007:DRP

[2620] Ewout W. Steyerberg, Judith Balmaña, David H. Stockwell, and Sapna Syngal. Data reduction for prediction: a case study on robust coding of age and family history for the risk of having a genetic mutation. Statistics in Medicine, 38(22):5545–5556, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Castelli:2007:CEA

[2621] Christel Castelli, Christophe Combescure, Yohann Foucher, and Jean-Pierre Daures. Cost-effectiveness analysis in colorectal cancer using a semi-Markov model. Statistics in Medicine, 38(22):5557–5571, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reilly:2007:SAD

[2622] Marie Reilly and Robert Szulkin. Statistical analysis of donationtransfusion data with complex correlation. Statistics in Medicine, 38 (22):5572–5585, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bender:2007:EAN

[2623] Ralf Bender, Oliver Kuss, Mandy Hildebrandt, and Ulrich Gehrmann. Estimating adjusted NNT measures in logistic regression analysis. Statistics in Medicine, 38(22):5586–5595, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chervoneva:2007:RQB

[2624] Inna Chervoneva, Yanyan Li, Boris Iglewicz, Scott Waldman, and Terry Hyslop. Relative quantification based on logistic models for individual polymerase chain reactions. Statistics in Medicine, 38(22):5596–5611, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2007:A

[2625] Anonymous. Acknowledgement. Statistics in Medicine, 38(22):5612, December 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2008:CE

[2626] Ralph B. D’Agostino and Joel Greenhouse. Comments from the Editors. Statistics in Medicine, 27(1):1–2, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2008:CRM

[2627] Yang-Jin Kim and Myoungshic Jhun. Cure rate model with interval censored data. Statistics in Medicine, 27(1):3–14, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsai:2008:IEB

[2628] Wei-Yann Tsai, Yunchan Chi, and Chia-Min Chen. Interval estimation of binomial proportion in clinical trials with a two-stage design. Statistics in Medicine, 27(1):15–35, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demidenko:2008:SSO

[2629] Eugene Demidenko. Sample size and optimal design for logistic regression with binary interaction. Statistics in Medicine, 27(1):36–46, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:2008:SSD

[2630] Kung-Jong Lui and Kuang-Chao Chang. Sample size determination for assessing equivalence based on proportion ratio under a randomized trial with non-compliance and missing outcomes. Statistics in Medicine, 27(1): 47–67, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kikuchi:2008:BCB

[2631] Takashi Kikuchi, Hamid Pezeshk, and John Gittins. A Bayesian costbenefit approach to the determination of sample size in clinical trials. Statistics in Medicine, 27(1):68–82, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siddique:2008:MIU

[2632] Juned Siddique and Thomas R. Belin. Multiple imputation using an iterative hot-deck with distance-based donor selection. Statistics in Medicine, 27(1):83–102, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:2008:HDM

[2633] Roderick J. Little, Matheos Yosef, Kevin C. Cain, Bin Nan, and Siobán D. Harlow. A hot-deck multiple imputation procedure for gaps in longitudinal data on recurrent events. Statistics in Medicine, 27(1): 103–120, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mei:2008:CMD

[2634] Y. Mei, L. Wang, and S. E. Holte. A comparison of methods for determining HIV viral set point. Statistics in Medicine, 27(1):121–139, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scott:2008:BMT

[2635] Allison N. Scott, Lawrence Joseph, Patrick Bélisle, Marcel A. Behr, and Kevin Schwartzman. Bayesian modelling of tuberculosis clustering from DNA fingerprint data. Statistics in Medicine, 27(1):140–156, January 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2008:EAP

[2636] Michael J. Pencina, Ralph B. D’Agostino, Sr., Ralph B. D’Agostino, Jr., and Ramachandran S. Vasan. Evaluating the added predictive ability of a new marker: From area under the ROC curve to reclassification and beyond. Statistics in Medicine, 27(2):157–172, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:2008:CEA

[2637] M. S. Pepe, Z. Feng, and J. W. Gu. Comments on ‘Evaluating the added predictive ability of a new marker: From area under the ROC curve to reclassification and beyond’ by M. J. Pencina et al., Statistics in Medicine (DOI: 10.1002/sim.2929). Statistics in Medicine, 27(2):173– 181, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2008:NRT

[2638] Yueh-Yun Chi and Xiao-Hua Zhou. The need for reorientation toward cost-effective prediction: Comments on ‘Evaluating the added predictive ability of a new marker: From area under the ROC curve to reclassification and beyond’ by Pencina et al., Statistics in Medicine (DOI: 10.1002/sim.2929). Statistics in Medicine, 27(2):182–184, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ware:2008:CEA

[2639] J. H. Ware and T. Cai. Comments on ‘Evaluating the added predictive ability of a new marker: From area under the ROC curve to reclassification and beyond’ by M. J. Pencina et al., Statistics in Medicine (DOI: 10.1002/sim.2929). Statistics in Medicine, 27(2):185–187, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2008:CEA

[2640] Philip Greenland. Comments on ‘Evaluating the added predictive ability of a new marker: From area under the ROC curve to reclassification and beyond’ by M. J. Pencina, R. B. D’Agostino, Sr., R. B. D’Agostino, Jr., R. S. Vasan, Statistics in Medicine (DOI: 10.1002/sim.2929). Statistics in Medicine, 27(2):188–190, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2008:CEA

[2641] Nancy R. Cook. Comments on ‘Evaluating the added predictive ability of a new marker: From area under the ROC curve to reclassification and beyond’ by M. J. Pencina et al., Statistics in Medicine (DOI: 10.1002/sim.2929). Statistics in Medicine, 27(2):191–195, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:2008:CEA

[2642] Helena Chmura Kraemer. Comments on ‘Evaluating the added predictive ability of a new marker’ by M. Pencina, R. D’Agostino, R. D’Agostino, Jr., R. Vasan, Statistics in Medicine (DOI: 10.1002/sim.2929). Statistics in Medicine, 27(2):196–198, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2008:NRT

[2643] Sander Greenland. The need for reorientation toward cost-effective prediction: Comments on ‘Evaluating the added predictive ability of a new marker: From area under the ROC curve to reclassification and beyond’ by M. J. Pencina et al., Statistics in Medicine (DOI: 10.1002/sim.2929). Statistics in Medicine, 27(2):199–206, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2008:CID

[2644] Michael J. Pencina, Ralph B. D’Agostino, Sr., Ralph B. D’Agostino, Jr., and Ramachandra S. Vasan. Comments on ‘Integrated discrimination and net reclassification improvements — Practical advice’. Statistics in Medicine, 27(2):207–212, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2008:WBG

[2645] Xiao-Hua Zhou, Sierra M. Li, and Constantine A. Gatsonis. Wilcoxonbased group sequential designs for comparison of areas under two correlated ROC curves. Statistics in Medicine, 27(2):213–223, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2008:EIE

[2646] Chi-Rong Li, Chen-Tuo Liao, and Jen-Pei Liu. On the exact interval estimation for the difference in paired areas under the ROC curves. Statistics in Medicine, 27(2):224–242, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2008:HMR

[2647] Fei Wang and Constantine A. Gatsonis. Hierarchical models for ROC curve summary measures: Design and analysis of multi-reader, multimodality studies of medical tests. Statistics in Medicine, 27(2):243–256, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2008:SI

[2648] Wang Zhou. Statistical inference for P (X < Y). Statistics in Medicine, 27(2):257–279, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vexler:2008:ERC

[2649] Albert Vexler, Enrique F. Schisterman, and Aiyi Liu. Estimation of ROC curves based on stably distributed biomarkers subject to measurement error and pooling mixtures. Statistics in Medicine, 27(2):280–296, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schisterman:2008:YIO

[2650] Enrique F. Schisterman, David Faraggi, Benjamin Reiser, and Jessica Hu. Youden index and the optimal threshold for markers with mass at zero. Statistics in Medicine, 27(2):297–315, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:2008:C

[2651] Sander Greenland. Correction. Statistics in Medicine, 27(2):316, January 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleming:2008:CIN

[2652] Thomas R. Fleming. Current issues in non-inferiority trials. Statistics in Medicine, 27(3):317–332, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:2008:CCI

[2653] Gary G. Koch. Comments on ‘Current issues in non-inferiority trials’ by Thomas R. Fleming, Statistics in Medicine, DOI: 10.1002/sim.2855. Statistics in Medicine, 27(3):333–342, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Powers:2008:NET

[2654] John H. Powers. Noninferiority and equivalence trials: deciphering ‘similarity’ of medical interventions. Statistics in Medicine, 27(3):343–352, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siqueira:2008:ACT

[2655] Arminda Lucia Siqueira, Anne Whitehead, and Susan Todd. Activecontrol trials with binary data: a comparison of methods for testing superiority or non-inferiority using the odds ratio. Statistics in Medicine, 27(3):353–370, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3041].

Snapinn:2008:CTE

[2656] Steven Snapinn and Qi Jiang. Controlling the type 1 error rate in noninferiority trials. Statistics in Medicine, 27(3):371–381, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snapinn:2008:PER

[2657] Steven Snapinn and Qi Jiang. Preservation of effect and the regulatory approval of new treatments on the basis of non-inferiority trials. Statistics in Medicine, 27(3):382–391, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:2008:ANI

[2658] Ping Gao and James H. Ware. Assessing non-inferiority: a combination approach. Statistics in Medicine, 27(3):392–406, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Georgopoulos:2008:NAC

[2659] Apostolos P. Georgopoulos and Elissaios Karageorgiou. Neurostatistics: Applications, challenges and expectations. Statistics in Medicine, 27(3): 407–417, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2008:FPM

[2660] Katherine J. Lee and Simon G. Thompson. Flexible parametric models for random-effects distributions. Statistics in Medicine, 27(3):418–434, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2008:SVC

[2661] Andrew D. A. C. Smith and M. P. Wand. Streamlined variance calculations for semiparametric mixed models. Statistics in Medicine, 27 (3):435–448, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strickland:2008:COR

[2662] Pamela A. Ohman Strickland and Shou-En Lu. Comparing odds ratios measuring improvement or deterioration across repeated exposure or treatment sessions. Statistics in Medicine, 27(3):449–461, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2008:SSC

[2663] Yan D. Zhao, Dewi Rahardja, and Yongming Qu. Sample size calculation for the Wilcoxon–Mann–Whitney test adjusting for ties. Statistics in Medicine, 27(3):462–468, February 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moye:2008:BCT

[2664] Lemuel A. Moyé. Bayesians in clinical trials: Asleep at the switch. Statistics in Medicine, 27(4):469–482, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chaloner:2008:CBC

[2665] Kathryn Chaloner. Comments on ‘Bayesians in clinical trials: Asleep at the switch’ by Lemuel A. Moyé, Statistics in Medicine. Statistics in Medicine, 27(4):483–487, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moye:2008:RNA

[2666] Lemuel A. Moyé. Rejoinder: The new adversary. Statistics in Medicine, 27(4):488–489, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hasler:2008:ANI

[2667] Mario Hasler, Richardus Vonk, and Ludwig A. Hothorn. Assessing noninferiority of a new treatment in a three-arm trial in the presence of heteroscedasticity. Statistics in Medicine, 27(4):490–503, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:2008:HBF

[2668] Yongzhao Shao, Vandana Mukhi, and Judith D. Goldberg. A hybrid Bayesian-frequentist approach to evaluate clinical trial designs for tests of superiority and non-inferiority. Statistics in Medicine, 27(4):504–519, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsong:2008:CMD

[2669] Yi Tsong, Mark Levenson, Joanne Zhang, and Jinglin Zhong. Choice of λ-margin and dependency of non-inferiority trials. Statistics in Medicine, 27(4):520–528, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2008:BAD

[2670] Wei Zhang, Kathryn Chaloner, Mary Kathryn Cowles, Ying Zhang, and Jack T. Stapleton. A Bayesian analysis of doubly censored data using a hierarchical Cox model. Statistics in Medicine, 27(4):529–542, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2008:MSD

[2671] John M. Williamson, Hae-Young Kim, Amita Manatunga, and David G. Addiss. Modeling survival data with informative cluster size. Statistics in Medicine, 27(4):543–555, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holmgren:2008:PST

[2672] Eric Holmgren. Are phase 2 screening trials in oncology obsolete? Statistics in Medicine, 27(4):556–567, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2008:RPI

[2673] Sin-Ho Jung. Randomized phase II trials with a prospective control. Statistics in Medicine, 27(4):568–583, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:EAP

[2674] G. Frank Liu, G. Ray Zhu, and Lu Cui. Evaluating the adaptive performance of flexible sample size designs with treatment difference in an interval. Statistics in Medicine, 27(4):584–596, February 20, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shun:2008:ITS

[2675] Zhenming Shun, K. K. Gordon Lan, and Yuhwen Soo. Interim treatment selection using the normal approximation approach in clinical trials. Statistics in Medicine, 27(4):597–618, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:2008:CTP

[2676] Paul C. Lambert, Alex J. Sutton, Paul R. Burton, Keith R. Abrams, and David R. Jones. Comments on ‘Trying to be precise about vagueness’ by Stephen Senn, Statistics in Medicine 2007; 26:1417–1430. Statistics in Medicine, 27(4):619–622, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2328].

Senn:2008:AR

[2677] Stephen Senn. Authors’ reply. Statistics in Medicine, 27(4):622–624, February 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sutton:2008:RDM

[2678] Alexander J. Sutton and Julian P. T. Higgins. Recent developments in meta-analysis. Statistics in Medicine, 27(5):625–650, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sutton:2008:MAI

[2679] A. J. Sutton, D. Kendrick, and C. A. C. Coupland. Meta-analysis of individual- and aggregate-level data. Statistics in Medicine, 27(5):651– 669, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ishak:2008:IAI

[2680] K. Jack Ishak, Robert W. Platt, Lawrence Joseph, and James A. Hanley. Impact of approximating or ignoring within-study covariances in multivariate meta-analyses. Statistics in Medicine, 27(5):670–686, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zwinderman:2008:WSP

[2681] Aeilko H. Zwinderman and Patrick M. Bossuyt. We should not pool diagnostic likelihood ratios in systematic reviews. Statistics in Medicine, 27(5):687–697, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryan:2008:CDM

[2682] Louise Ryan. Combining data from multiple sources, with applications to environmental risk assessment. Statistics in Medicine, 27(5):698–710, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2008:AUDa

[2683] Ian R. White, Julian P. T. Higgins, and Angela M. Wood. Allowing for uncertainty due to missing data in meta-analysis — Part 1: Twostage methods. Statistics in Medicine, 27(5):711–727, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2008:AUDb

[2684] Ian R. White, Nicky J. Welton, Angela M. Wood, A. E. Ades, and Julian P. T. Higgins. Allowing for uncertainty due to missing data in metaanalysis — Part 2: Hierarchical models. Statistics in Medicine, 27(5): 728–745, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rucker:2008:ATP

[2685] Gerta Rücker, Guido Schwarzer, and James Carpenter. Arcsine test for publication bias in meta-analyses with binary outcomes. Statistics in Medicine, 27(5):746–763, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salanti:2008:MAG

[2686] Georgia Salanti and Julian P. T. Higgins. Meta-analysis of genetic association studies under different inheritance models using data reported as merged genotypes. Statistics in Medicine, 27(5):764–777, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krouwer:2008:WBA

[2687] Jan S. Krouwer. Why Bland–Altman plots should use X, not (Y + X)/2 when X is a reference method. Statistics in Medicine, 27(5):778–780, February 28, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heisterkamp:2008:EPH

[2688] S. H. Heisterkamp, R. de Vries, H. G. Sprenger, G. A. A. Hubben, and M. J. Postma. Estimation and prediction of the HIV–AIDS-epidemic under conditions of HAART using mixtures of incubation time distributions. Statistics in Medicine, 27(6):781–794, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:2008:CFE

[2689] Michael Proschan and Dean Follmann. Cluster without fluster: The effect of correlated outcomes on inference in randomized clinical trials. Statistics in Medicine, 27(6):795–809, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:2008:MPA

[2690] Chris Roberts. Modelling patterns of agreement for nominal scales. Statistics in Medicine, 27(6):810–830, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gandy:2008:GTA

[2691] Axel Gandy, Terry M. Therneau, and Odd O. Aalen. Global tests in the additive hazards regression model. Statistics in Medicine, 27(6):831–844, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2008:HTF

[2692] Xiaowei Yang and Kun Nie. Hypothesis testing in functional linear regression models with Neyman’s truncation and wavelet thresholding for longitudinal data. Statistics in Medicine, 27(6):845–863, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haneuse:2008:GBE

[2693] S. Haneuse and J. Wakefield. Geographic-based ecological correlation studies using supplemental case-control data. Statistics in Medicine, 27(6):864–887, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Argyle:2008:CMM

[2694] Jenny Argyle, Allan H. Seheult, and David A. Wooff. Correlation models for monitoring child growth. Statistics in Medicine, 27(6):888–904, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Charitos:2008:CSI

[2695] Theodore Charitos, Peter R. de Waal, and Linda C. van der Gaag. Computing short-interval transition matrices of a discrete-time Markov chain from partially observed data. Statistics in Medicine, 27(6):905–921, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tai:2008:FEM

[2696] Bee-Choo Tai, Bianca L. De Stavola, Victor de Gruttola, Val Gebski, and David Machin. First-event or marginal estimation of cause-specific hazards for analysing correlated multivariate failure-time data? Statistics in Medicine, 27(6):922–936, March 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2008:SBM

[2697] Yue Li and You-Gan Wang. Smooth bootstrap methods for analysis of longitudinal data. Statistics in Medicine, 27(7):937–953, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hamling:2008:FMA

[2698] Jan Hamling, Peter Lee, Rolf Weitkunat, and Mathias Ambühl. Facilitating meta-analyses by deriving relative effect and precision estimates for alternative comparisons from a set of estimates presented by exposure level or disease category. Statistics in Medicine, 27(7):954–970, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:DVI

[2699] Xuefeng Liu and Jeffrey Roth. Development and validation of an infant morbidity index using latent variable models. Statistics in Medicine, 27(7):971–989, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simonetti:2008:ECP

[2700] Arianna Simonetti, Anna Gigli, Riccardo Capocaccia, and Angela Mariotto. Estimating complete prevalence of cancers diagnosed in childhood. Statistics in Medicine, 27(7):990–1007, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shardell:2008:WEE

[2701] Michelle Shardell and Ram R. Miller. Weighted estimating equations for longitudinal studies with death and non-monotone missing timedependent covariates and outcomes. Statistics in Medicine, 27(7):1008– 1025, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Archer:2008:DCE

[2702] K. J. Archer, C. I. Dumur, G. S. Taylor, M. D. Chaplin, A. Guiseppi-Elie, G. A. Buck, G. Grant, A. Ferreira-Gonzalez, and C. T. Garrett. A disattenuated correlation estimate when variables are measured with error: Illustration estimating cross-platform correlations. Statistics in Medicine, 27(7):1026–1039, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boik:2008:QSA

[2703] John C. Boik, Robert A. Newman, and Robert J. Boik. Quantifying synergism/antagonism using nonlinear mixed-effects modeling: a simulation study. Statistics in Medicine, 27(7):1040–1061, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moger:2008:CCM

[2704] Tron Anders Moger, Yudi Pawitan, and Ørnulf Borgan. Case-cohort methods for survival data on families from routine registers. Statistics in Medicine, 27(7):1062–1074, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:2008:ARE

[2705] Yang-Jin Kim and Myoungshic Jhun. Analysis of recurrent event data with incomplete observation gaps. Statistics in Medicine, 27(7):1075– 1085, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yip:2008:CCM

[2706] Benjamin H. Yip, Camilla Björk, Paul Lichtenstein, Christina M. Hultman, and Yudi Pawitan. Covariance component models for multivariate binary traits in family data analysis. Statistics in Medicine, 27(7):1086– 1105, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsui:2008:CSG

[2707] Shigeyuki Matsui, Takeharu Yamanaka, Bart Barlogie, John D. Shaughnessy, Jr., and John Crowley. Clustering of significant genes in prognostic studies with microarrays: Application to a clinical study for multiple myeloma. Statistics in Medicine, 27(7):1106–1120, March 30, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Virtanen:2008:CPL

[2708] A. Virtanen and E. Uusipaikka. Computation of profile likelihood-based confidence intervals for reference limits with covariates. Statistics in Medicine, 27(7):1121–1132, March 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawlor:2008:MRU

[2709] Debbie A. Lawlor, Roger M. Harbord, Jonathan A. C. Sterne, Nic Timpson, and George Davey Smith. Mendelian randomization: Using genes as instruments for making causal inferences in epidemiology. Statistics in Medicine, 27(8):1133–1163, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salim:2008:SAI

[2710] Agus Salim, Andrew Mackinnon, and Kathleen Griffiths. Sensitivity analysis of intention-to-treat estimates when withdrawals are related to unobserved compliance status. Statistics in Medicine, 27(8):1164–1179, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ruan:2008:SAP

[2711] Ping K. Ruan and Robert J. Gray. Sensitivity analysis of progressionfree survival with dependent withdrawal. Statistics in Medicine, 27(8): 1180–1198, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sambucini:2008:BPT

[2712] Valeria Sambucini. A Bayesian predictive two-stage design for phase II clinical trials. Statistics in Medicine, 27(8):1199–1224, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sego:2008:CSM

[2713] Landon H. Sego, William H. Woodall, and Marion R. Reynolds, Jr. A comparison of surveillance methods for small incidence rates. Statistics in Medicine, 27(8):1225–1247, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shepherd:2008:CCP

[2714] Bryan E. Shepherd. The cost of checking proportional hazards. Statistics in Medicine, 27(8):1248–1260, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2008:DTS

[2715] Xiaobing Zhao and Xian Zhou. Discrete-time survival models with longterm survivors. Statistics in Medicine, 27(8):1261–1281, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2008:MAP

[2716] Jeffrey M. Albert. Mediation analysis via potential outcomes models. Statistics in Medicine, 27(8):1282–1304, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hunsberger:2008:TLR

[2717] Sally Hunsberger, Barry I. Graubard, and Edward L. Korn. Testing logistic regression coefficients with clustered data and few positive outcomes. Statistics in Medicine, 27(8):1305–1324, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Suchindran:2008:BRS

[2718] C. M. Suchindran. Book review: Statistical demography and forecasting. Juha M. Alho, Bruce D. Spencer, Springer Science-Business Media. Inc, 2005. No. of pages: xxvii + 410. Price: $89.85 hardcover, $54.95 softcover. ISBN 978-0-387-23530-1 hardcover, 978-0-387-22538-8 softcover. Statistics in Medicine, 27(8):1325–1326, April 15, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robinson:2008:BRS

[2719] Derek Robinson. Book review: Statistical analysis of medical data using SAS. Geoff Der and Brian S. Everitt, Chapman and Hall/CRC, Boca Raton, FL, 2006. No. of pages: 428. Price: $79.95. ISBN 978-1-58488-469-X. Statistics in Medicine, 27(8):1326–1327, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Todd:2008:BRI

[2720] Susan Todd. Book review: Introduction to randomized controlled clinical trials. (2nd edn). John N. S. Matthews, Chapman and Hall/CRC, Boca Raton, FL, 2006. No. of pages: 283. Price: $59.95. ISBN 978-1-58488-624-2. Statistics in Medicine, 27(8):1328, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teixeira-Pinto:2008:SMC

[2721] Armando Teixeira-Pinto and Sharon-Lise T. Normand. Statistical methodology for classifying units on the basis of multiple-related measures. Statistics in Medicine, 27(9):1329–1350, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Timbie:2008:CMC

[2722] Justin W. Timbie and Sharon-Lise T. Normand. A comparison of methods for combining quality and efficiency performance measures: Profiling the value of hospital care following acute myocardial infarction. Statistics in Medicine, 27(9):1351–1370, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diehr:2008:ASP

[2723] Paula Diehr, David Yanez, Ann Derleth, and Anne B. Newman. Agespecific prevalence and years of healthy life in a system with three health states. Statistics in Medicine, 27(9):1371–1386, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eickhoff:2008:PEA

[2724] Jens C. Eickhoff. Placebo effect-adjusted assessment of quality of life in placebo-controlled clinical trials. Statistics in Medicine, 27(9):1387–1402, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Atienza:2008:AMD

[2725] N. Atienza, J. García-Heras, J. M. Muñoz-Pichardo, and R. Villa. An application of mixture distributions in modelization of length of hospital stay. Statistics in Medicine, 27(9):1403–1420, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salama:2008:EGC

[2726] Ibrahim Salama, Anastasia Ivanova, and Bahjat Qaqish. Efficient generation of constrained block allocation sequences. Statistics in Medicine, 27(9):1421–1428, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2008:TNN

[2727] G. Zhao. Tests of non-null hypothesis on proportions for stratified data. Statistics in Medicine, 27(9):1429–1446, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samuelsen:2008:AFS

[2728] Sven Ove Samuelsen and Geir Egil Eide. Attributable fractions with survival data. Statistics in Medicine, 27(9):1447–1467, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2008:BHM

[2729] Huafeng Zhou, Andrew B. Lawson, James R. Hebert, Elizabeth H. Slate, and Elizabeth G. Hill. A Bayesian hierarchical modeling approach for studying the factors affecting the stage at diagnosis of prostate cancer. Statistics in Medicine, 27(9):1468–1489, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2008:EPL

[2730] Tsung I. Lin and Jack C. Lee. Estimation and prediction in linear mixed models with skew-normal random effects for longitudinal data. Statistics in Medicine, 27(9):1490–1507, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Czajkowski:2008:MSL

[2731] Michal Czajkowski, Ryan Gill, and Grzegorz Rempala. Model selection in logistic joinpoint regression with applications to analyzing cohort mortality patterns. Statistics in Medicine, 27(9):1508–1526, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBuuren:2008:IAW

[2732] Stef van Buuren. Improved accuracy when screening for human growth disorders by likelihood ratios. Statistics in Medicine, 27(9):1527–1538, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnston:2008:UIV

[2733] K. M. Johnston, P. Gustafson, A. R. Levy, and P. Grootendorst. Use of instrumental variables in the analysis of generalized linear models in the presence of unmeasured confounding with applications to epidemiological research. Statistics in Medicine, 27(9):1539–1556, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenbauer:2008:CAP

[2734] Joachim Rosenbauer and Klaus Strassburger. Comments on ‘Age-period-cohort models for the Lexis diagram’ by Carstensen B. Statistics in Medicine 2007; 26:3018–3045. Statistics in Medicine, 27(9):1557–1561, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2436, 2735].

Carstensen:2008:LEA

[2735] B. Carstensen. Letter to the Editor: Age–period–cohort models for the Lexis diagram, Statistics in Medicine 2007; 26:3018–3045. Statistics in Medicine, 27(9):1561–1564, April 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2436, 2734].

Bauer:2008:ADL

[2736] Peter Bauer. Adaptive designs: Looking for a needle in the haystack — a new challenge in medical research. Statistics in Medicine, 27(10): 1565–1580, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2008:ADF

[2737] Anastasia Ivanova, James A. Bolognese, and Inna Perevozskaya. Adaptive dose finding based on t-statistic for dose-response trials. Statistics in Medicine, 27(10):1581–1592, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartroff:2008:EAD

[2738] Jay Bartroff and Tze Leung Lai. Efficient adaptive designs with midcourse sample size adjustment in clinical trials. Statistics in Medicine, 27(10):1593–1611, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koenig:2008:ADT

[2739] Franz Koenig, Werner Brannath, Frank Bretz, and Martin Posch. Adaptive Dunnett tests for treatment selection. Statistics in Medicine, 27(10): 1612–1625, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dawson:2008:SCI

[2740] Ree Dawson and Philip W. Lavori. Sequential causal inference: Application to randomized trials of adaptive treatment strategies. Statistics in Medicine, 27(10):1626–1645, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2008:IIB

[2741] Anne Whitehead, Marina Roshini Sooriyarachchi, John Whitehead, and Kim Bolland. Incorporating intermediate binary responses into interim analyses of clinical trials: a comparison of four methods. Statistics in Medicine, 27(10):1646–1666, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:NSE

[2742] Aiyi Liu, Chengqing Wu, and Enrique F. Schisterman. Nonparametric sequential evaluation of diagnostic biomarkers. Statistics in Medicine, 27(10):1667–1678, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:2008:TNP

[2743] Helena Chmura Kraemer. Toward non-parametric and clinically meaningful moderators and mediators. Statistics in Medicine, 27(10):1679– 1692, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2008:CCL

[2744] G. Y. Zou and A. Donner. Construction of confidence limits about effect measures: a general approach. Statistics in Medicine, 27(10):1693–1702, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3381].

Ryu:2008:MIO

[2745] Euijung Ryu and Alan Agresti. Modeling and inference for an ordinal effect size measure. Statistics in Medicine, 27(10):1703–1717, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tanaka:2008:ETE

[2746] Yukari Tanaka, Yutaka Matsuyama, and Yasuo Ohashi. Estimation of treatment effect adjusting for treatment changes using the intensity score method: Application to a large primary prevention study for coronary events (MEGA study). Statistics in Medicine, 27(10):1718–1733, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dodd:2008:LGS

[2747] Lori E. Dodd and Edward L. Korn. Lack of generalizability of sensitivity and specificity with treatment effects. Statistics in Medicine, 27(10): 1734–1744, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Puggioni:2008:JMS

[2748] Gavino Puggioni, Alan E. Gelfand, and Joann G. Elmore. Joint modeling of sensitivity and specificity. Statistics in Medicine, 27(10):1745–1761, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2008:NIT

[2749] Chi-Rong Li, Chen-Tuo Liao, and Jen-Pei Liu. A non-inferiority test for diagnostic accuracy based on the paired partial areas under ROC curves. Statistics in Medicine, 27(10):1762–1776, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berres:2008:EDS

[2750] M. Berres, A. Zehnder, S. Bläsi, and A. U. Monsch. Evaluation of diagnostic scores with adjustment for covariates. Statistics in Medicine, 27(10):1777–1790, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andres:2008:CCS

[2751] Antonio Martín Andrés. Comments on ‘Chi-squared and Fisher– Irwin tests of two-by-two tables with small sample recommendations’ by I. Campbell, Statistics in Medicine 2007; 26:3661–3675. DOI: 10.1002/sim.2832. Statistics in Medicine, 27(10):1791–1795, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2488].

Campbell:2008:AR

[2752] Ian R. Campbell. Author’s reply. Statistics in Medicine, 27(10):1795– 1796, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:2008:C

[2753] Edward L. Korn, Paul S. Albert, and Lisa M. McShane. Correction. Statistics in Medicine, 27(10):1797, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biesheuvel:2008:BRP

[2754] Egbert Biesheuvel. Book review: Pharmaceutical statistics using SAS(R): a practical guide. Alex Dmitrienko, Christy Chuang-Stein and Ralph D’Agostino (eds), SAS(R) Press, Cary, NC, 2007. No. of pages: ix + 443. Price: $69.95. ISBN: 978-1-59047-886-8. Statistics in Medicine, 27(10):1798–1799, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yasui:2008:BRH

[2755] Yutaka Yasui. Book review: A handbook of statistical analyses using Stata. (2nd edn). Sophia Rabe-Hesketh and Brian S. Everitt, Chapman and Hall/CRC, Boca Raton, FL, 2007. No. of pages: ix + 342. Price: $49.46. ISBN 1-58488-756-7. Statistics in Medicine, 27(10):1799–1800, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:2008:GCT

[2756] Joel B. Greenhouse, Eloise E. Kaizar, Kelly Kelleher, Howard Seltman, and William Gardner. Generalizing from clinical trial data: a case study. the risk of suicidality among pediatric antidepressant users. Statistics in Medicine, 27(11):1801–1813, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gibbons:2008:MEP

[2757] Robert D. Gibbons, Eisuke Segawa, George Karabatsos, Anup K. Amatya, Dulal K. Bhaumik, C. Hendricks Brown, Kush Kapur, Sue M. Marcus, Kwan Hur, and J. John Mann. Mixed-effects Poisson regression analysis of adverse event reports: The relationship between antidepressants and suicide. Statistics in Medicine, 27(11):1814–1833, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chung:2008:LTA

[2758] Hwan Chung, Stephanie T. Lanza, and Eric Loken. Latent transition analysis: Inference and estimation. Statistics in Medicine, 27(11):1834– 1854, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sepulveda:2008:BDT

[2759] R. Sepúlveda, J. L. Vicente-Villardón, and M. P. Galindo. The biplot as a diagnostic tool of local dependence in latent class models. a medical application. Statistics in Medicine, 27(11):1855–1869, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2008:MAC

[2760] Richard D. Riley, Paul C. Lambert, Jan A. Staessen, Jiguang Wang, Francois Gueyffier, Lutgarde Thijs, and Florent Boutitie. Meta-analysis of continuous outcomes combining individual patient data and aggregate data. Statistics in Medicine, 27(11):1870–1893, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rondeau:2008:ITT

[2761] V. Rondeau, S. Michiels, B. Liquet, and J. P. Pignon. Investigating trial and treatment heterogeneity in an individual patient data metaanalysis of survival data by means of the penalized maximum likelihood approach. Statistics in Medicine, 27(11):1894–1910, May 20, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zucker:2008:CSE

[2762] David M. Zucker and Donna Spiegelman. Corrected score estimation in the proportional hazards model with misclassified discrete covariates. Statistics in Medicine, 27(11):1911–1933, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanderWeele:2008:ISA

[2763] Tyler J. VanderWeele. Ignorability and stability assumptions in neighborhood effects research. Statistics in Medicine, 27(11):1934–1943, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:2008:MIT

[2764] Ronglai Shen, Debashis Ghosh, and Jeremy M. G. Taylor. Modeling intra-tumor protein expression heterogeneity in tissue microarray experiments. Statistics in Medicine, 27(11):1944–1959, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tong:2008:PGA

[2765] Tiejun Tong and Hongyu Zhao. Practical guidelines for assessing power and false discovery rate for a fixed sample size in microarray experiments. Statistics in Medicine, 27(11):1960–1972, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Capanu:2008:UHM

[2766] Marinela Capanu, Irene Orlow, Marianne Berwick, Amanda J. Hummer, Duncan C. Thomas, and Colin B. Begg. The use of hierarchical models for estimating relative risks of individual genetic variants: an application to a study of melanoma. Statistics in Medicine, 27(11):1973–1992, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dryden:2008:FRN

[2767] Ian L. Dryden, Neil Oxborrow, and Robert Dickson. Familial relationships of normal spine shape. Statistics in Medicine, 27(11):1993–2003, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2008:FDR

[2768] Wei Chen, Debashis Ghosh, Trivellore E. Raghunathan, and Daniel J. Sargent. A false-discovery-rate-based loss framework for selection of interactions. Statistics in Medicine, 27(11):2004–2021, May 20, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kelly:2008:STG

[2769] Patrick Kelly, Yinghui Zhou, John Whitehead, Nigel Stallard, and Clive Bowman. Sequentially testing for a gene-drug interaction in a genomewide analysis. Statistics in Medicine, 27(11):2022–2034, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nadarajah:2008:CEP

[2770] Saralees Nadarajah. Comment on ‘Estimation and prediction of the HIV–AIDS-epidemic under conditions of HAART using mixtures of incubation time distributions’ by S. H. Heisterkamp, R. de Vries, H. G. Sprenger, G. A. A. Hubben, M. J. Postma, Statistics in Medicine, DOI: 10.1002/sim.2974. Statistics in Medicine, 27(11):2035–2036, May 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2008:CAP

[2771] Peter C. Austin. A critical appraisal of propensity-score matching in the medical literature between 1996 and 2003. Statistics in Medicine, 27(12):

2037–2049, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hansen:2008:CER

[2772] B. B. Hansen. Commentary: The essential role of balance tests in propensity-matched observational studies: Comments on ‘A critical appraisal of propensity-score matching in the medical literature between 1996 and 2003’ by Peter Austin, Statistics in Medicine. Statistics in Medicine, 27(12):2050–2054, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hill:2008:DRU

[2773] Jennifer Hill. Discussion of research using propensity-score matching: Comments on ‘A critical appraisal of propensity-score matching in the medical literature between 1996 and 2003’ by Peter Austin, Statistics in Medicine. Statistics in Medicine, 27(12):2055–2061, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stuart:2008:DPR

[2774] Elizabeth A. Stuart. Developing practical recommendations for the use of propensity scores: Discussion of ‘A critical appraisal of propensity score matching in the medical literature between 1996 and 2003’ by Peter Austin, Statistics in Medicine. Statistics in Medicine, 27(12):2062–2065, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2008:DCA

[2775] Peter C. Austin. Discussion of ‘a critical appraisal of propensity-score matching in the medical literature between 1996 and 2003’. Statistics in Medicine, 27(12):2066–2069, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2008:TDC

[2776] Bo Huang and Rick Chappell. Three-dose-cohort designs in cancer phase I trials. Statistics in Medicine, 27(12):2070–2093, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutson:2008:DCM

[2777] Alan D. Hutson. A discrete-continuous mixture quantile function estimator with a practical application to phase II cancer clinical trials. Statistics in Medicine, 27(12):2094–2109, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scharfstein:2008:AWP

[2778] Daniel O. Scharfstein, Jennifer L. Ryea, and Brian Caffo. Accounting for within-patient correlation in assessing relative sensitivity of an adjunctive diagnostic test: Application to lung cancer. Statistics in Medicine, 27 (12):2110–2126, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diva:2008:PMS

[2779] Ulysses Diva, Dipak K. Dey, and Sudipto Banerjee. Parametric models for spatially correlated survival data for individuals with multiple cancers. Statistics in Medicine, 27(12):2127–2144, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rashid:2008:ESN

[2780] Ivan Rashid, Luigi Marcheselli, and Massimo Federico. Estimating survival in newly diagnosed cancer patients: Use of computer simulations to evaluate performances of different approaches in a wide range of scenarios. Statistics in Medicine, 27(12):2145–2158, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fearn:2008:MEE

[2781] T. Fearn, D. C. Hill, and S. C. Darby. Measurement error in the explanatory variable of a binary regression: Regression calibration and integrated conditional likelihood in studies of residential radon and lung cancer. Statistics in Medicine, 27(12):2159–2176, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Titman:2008:GGF

[2782] Andrew C. Titman and Linda D. Sharples. A general goodness-of-fit test for Markov and hidden Markov models. Statistics in Medicine, 27(12): 2177–2195, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rao:2008:SRS

[2783] R. Sowmya Rao, Mark E. Glickman, and Robert J. Glynn. Stopping rules for surveys with multiple waves of nonrespondent follow-up. Statistics in Medicine, 27(12):2196–2213, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinzl:2008:TUB

[2784] Harald Heinzl, Martina Mittlböck, and Lutz Edler. Technical uncertainty in the back-calculation of occupational exposure to dioxins. Statistics in Medicine, 27(12):2214–2233, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bravata:2008:CTV

[2785] Dena M. Bravata, Kaveh G. Shojania, Ingram Olkin, and Adi Raveh. CoPlot: a tool for visualizing multivariate data in medicine. Statistics in Medicine, 27(12):2234–2247, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schildcrout:2008:ALL

[2786] Jonathan S. Schildcrout, Cathy A. Jenkins, Jack H. Ostroff, Daniel L. Gillen, Frank E. Harrell, and Donald C. Trost. Analysis of longitudinal laboratory data in the presence of common selection mechanisms: a view toward greater emphasis on pre-marketing pharmaceutical safety. Statistics in Medicine, 27(12):2248–2266, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smeeton:2008:CSE

[2787] Nigel C. Smeeton. Comments on ‘Statistical education for medical students — Concepts are what remain when the details are forgotten’ by Amir Herman, Netta Notzer, Zipi Libman, Rony Braunstein and David Steinberg. Statistics in Medicine 2007; 26:4344–4351. Statistics in Medicine, 27(12):2267–2268, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2537].

Herman:2008:SEM

[2788] Amir Herman, Netta Notzer, Zipi Libman, Rony Braunstein, and David M. Steinberg. Statistical education for medical students — concepts are what remain when the details are forgotten. Statistics in Medicine, 27(12):2268–2269, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hazelton:2008:KER

[2789] Martin L. Hazelton. Kernel estimation of risk surfaces without the need for edge correction. Statistics in Medicine, 27(12):2269–2272, May 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2008:BAS

[2790] Kunshan Yin, Pankaj K. Choudhary, Diana Varghese, and Steven R. Goodman. A Bayesian approach for sample size determination in method comparison studies. Statistics in Medicine, 27(13):2273–2289, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brutti:2008:RBS

[2791] Pierpaolo Brutti, Fulvio De Santis, and Stefania Gubbiotti. Robust Bayesian sample size determination in clinical trials. Statistics in Medicine, 27(13):2290–2306, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2008:BSS

[2792] John Whitehead, Elsa Valdés-Márquez, Patrick Johnson, and Gordon Graham. Bayesian sample size for exploratory clinical trials incorporating historical data. Statistics in Medicine, 27(13):2307–2327, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2008:PAC

[2793] Byron J. Gajewski, Stephen D. Simon, and Susan E. Carlson. Predicting accrual in clinical trials with Bayesian posterior predictive distributions. Statistics in Medicine, 27(13):2328–2340, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2008:SBA

[2794] Sik-Yum Lee, Bin Lu, and Xin-Yuan Song. Semiparametric Bayesian analysis of structural equation models with fixed covariates. Statistics in Medicine, 27(13):2341–2360, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2008:BHM

[2795] Jie Chen, Jinglin Zhong, and Lei Nie. Bayesian hierarchical modeling of drug stability data. Statistics in Medicine, 27(13):2361–2380, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silva:2008:HBS

[2796] Giovani L. Silva, C. B. Dean, Théophile Niyonsenga, and Alain Vanasse. Hierarchical Bayesian spatiotemporal analysis of revascularization odds using smoothing splines. Statistics in Medicine, 27(13):2381–2401, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2008:BMM

[2797] Pulak Ghosh and Mithat Gönen. Bayesian modeling of multivariate average bioequivalence. Statistics in Medicine, 27(13):2402–2419, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neuenschwander:2008:CAB

[2798] Beat Neuenschwander, Michael Branson, and Thomas Gsponer. Critical aspects of the Bayesian approach to phase I cancer trials. Statistics in Medicine, 27(13):2420–2439, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [3388].

Stamey:2008:BAA

[2799] James D. Stamey, Dean M. Young, and John W. Seaman, Jr. A Bayesian approach to adjust for diagnostic misclassification between two mortality causes in Poisson regression. Statistics in Medicine, 27(13):2440–2452, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buzoianu:2008:AVB

[2800] Manuela Buzoianu and Joseph B. Kadane. Adjusting for verification bias in diagnostic test evaluation: a Bayesian approach. Statistics in Medicine, 27(13):2453–2473, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Branscum:2008:BSR

[2801] Adam J. Branscum, Wesley O. Johnson, Timothy E. Hanson, and Ian A. Gardner. Bayesian semiparametric ROC curve estimation and disease diagnosis. Statistics in Medicine, 27(13):2474–2496, June 15, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2008:BAE

[2802] Haitao Chu, Lei Nie, and Thomas W. Kensler. A Bayesian approach estimating treatment effects on biomarkers containing zeros with detection limits. Statistics in Medicine, 27(13):2497–2508, June 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:2008:SSE

[2803] John M. Lachin. Sample size evaluation for a multiply matched casecontrol study using the score test from a conditional logistic (discrete Cox PH) regression model. Statistics in Medicine, 27(14):2509–2523, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kang:2008:THT

[2804] Seung-Ho Kang and Chul W. Ahn. Tests for the homogeneity of two binomial proportions in extremely unbalanced 2×2 contingency tables. Statistics in Medicine, 37(10):2524–2535, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3135].

Capanu:2008:MTB

[2805] Marinela Capanu and Brett Presnell. Misspecification tests for binomial and beta-binomial models. Statistics in Medicine, 37(10):2536–2554, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joner:2008:DRI

[2806] Michael D. Joner, Jr., William H. Woodall, and Marion R. Reynolds, Jr. Detecting a rate increase using a Bernoulli scan statistic. Statistics in Medicine, 37(10):2555–2575, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsou:2008:CSM

[2807] Tsung-Shan Tsou and Chien-Hung Chen. Comparing several means of dependent populations of count — a parametric robust approach. Statistics in Medicine, 37(10):2576–2585, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tekle:2008:OCD

[2808] Fetene B. Tekle, Frans E. S. Tan, and Martijn P. F. Berger. D-optimal cohort designs for linear mixed-effects models. Statistics in Medicine, 37(10):2586–2600, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ortega-Azurduy:2008:EDE

[2809] S. A. Ortega-Azurduy, F. E. S. Tan, and M. P. F. Berger. The effect of dropout on the efficiency of D-optimal designs of linear mixed models. Statistics in Medicine, 37(10):2601–2617, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sylvestre:2008:CAG

[2810] Marie-Pierre Sylvestre and Michal Abrahamowicz. Comparison of algorithms to generate event times conditional on time-dependent covariates. Statistics in Medicine, 37(10):2618–2634, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sjolander:2008:IDB

[2811] Arvid Sjölander, Keith Humphreys, and Juni Palmgren. On informative detection bias in screening studies. Statistics in Medicine, 37(10):2635– 2650, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hens:2008:MMD

[2812] N. Hens, M. Aerts, Z. Shkedy, H. Theeten, P. Van Damme, and Ph. Beutels. Modelling multisera data: The estimation of new joint and conditional epidemiological parameters. Statistics in Medicine, 37(10): 2651–2664, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:UGQ

[2813] Lei Liu and Xuelin Huang. The use of Gaussian quadrature for estimation in frailty proportional hazards models. Statistics in Medicine, 37(10): 2665–2683, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bebu:2008:CMV

[2814] Ionut Bebu and Thomas Mathew. Comparing the means and variances of a bivariate log-normal distribution. Statistics in Medicine, 37(10): 2684–2696, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2008:MLV

[2815] Leena Choi, Brian Caffo, Charles Rohde, Themba T. Ndovi, and Craig W. Hendrix. A mechanistic latent variable model for estimating drug concentrations in the male genital tract: a case study in drug kinetics. Statistics in Medicine, 37(10):2697–2714, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2008:ANF

[2816] Richard J. Cook, Bethany J. G. White, Grace Y. Yi, Ker-Ai Lee, and Theodore E. Warkentin. Analysis of a nonsusceptible fraction with current status data. Statistics in Medicine, 37(10):2715–2730, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nadarajah:2008:CPA

[2817] Saralees Nadarajah. Comment on ‘The partial area under the summary ROC curve’ by S. D. Walter, Statistics in Medicine 2005; 24:2025– 2040. Statistics in Medicine, 37(10):2731–2734, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1745].

Walter:2008:AR

[2818] S. D. Walter. Author’s reply. Statistics in Medicine, 37(10):2734–2735, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McEntegart:2008:CRP

[2819] Damian J. McEntegart. Comments on ‘Randomization with a posteriori constraints: description and properties’ by Stephanie Gicquel and Roland Marion-Gallois, Statistics in Medicine 2007; 26:5033–5045. Statistics in Medicine, 37(10):2735–2736, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2584].

Bjork:2008:CPE

[2820] Jonas Björk, Michael Green, and Ulf Ekelund. Comments on ‘Practical experiences on the necessity of external validation’ by I. R. König, J. D. Malley, C. Weimar, H.-C. Diener and A. Ziegler, Statistics in Medicine, DOI: 10.1002/sim.3069. Statistics in Medicine, 37(10):2737–2738, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Konig:2008:AR

[2821] Inke R. König, Andreas Ziegler, and James D. Malley. Authors’ reply. Statistics in Medicine, 37(10):2738–2740, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shrier:2008:LE

[2822] Ian Shrier. Letter to the Editor. Statistics in Medicine, 37(10):2740–2741, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:2008:AR

[2823] Donald B. Rubin. Author’s reply. Statistics in Medicine, 37(10):2741– 2742, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boni:2008:BRB

[2824] Riccardo L. Boni. Book review: Basic statistics and pharmaceutical statistical applications (2nd Edition). James E. De Muth, Chapman and Hall/CRC, Boca Raton, FL, 2006. No. of pages: xix + 714. Price: $89.95.

ISBN: 0-8493-3799-2. Statistics in Medicine, 37(10):2743–2744, June 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wehby:2008:MRE

[2825] George L. Wehby, Robert L. Ohsfeldt, and Jeffrey C. Murray. ‘Mendelian randomization’ equals instrumental variable analysis with genetic instruments. Statistics in Medicine, 27(15):2745–2749, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawlor:2008:MRL

[2826] Debbie A. Lawlor, Frank Windmeijer, and George Davey Smith. Is Mendelian randomization ‘lost in translation?’: Comments on ‘Mendelian randomization equals instrumental variable analysis with genetic instruments’ by Wehby et al. Statistics in Medicine, 27(15):2750– 2755, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2008:AED

[2827] Li Zhang, Bhramar Mukherjee, Malay Ghosh, Stephen Gruber, and Victor Moreno. Accounting for error due to misclassification of exposures in case-control studies of gene-environment interaction. Statistics in Medicine, 27(15):2756–2783, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foulkes:2008:MMM

[2828] A. S. Foulkes, R. Yucel, and M. P. Reilly. Mixed modeling and multiple imputation for unobservable genotype clusters. Statistics in Medicine, 27(15):2784–2801, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wathen:2008:APH

[2829] J. Kyle Wathen, Peter F. Thall, John D. Cook, and Elihu H. Estey. Accounting for patient heterogeneity in phase II clinical trials. Statistics in Medicine, 27(15):2802–2815, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yi:2008:RAC

[2830] Qilong Yi, P. Peter Wang, and Yaohua He. Reliability analysis for continuous measurements: Equivalence test for agreement. Statistics in Medicine, 27(15):2816–2825, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:2008:IBS

[2831] Xiaowei Yang, Jinhui Li, and Steven Shoptaw. Imputation-based strategies for clinical trial longitudinal data with nonignorable missing values. Statistics in Medicine, 27(15):2826–2849, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moerbeek:2008:SSF

[2832] Mirjam Moerbeek and Weng Kee Wong. Sample size formulae for trials comparing group and individual treatments in a multilevel model. Statistics in Medicine, 27(15):2850–2864, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelman:2008:SRI

[2833] Andrew Gelman. Scaling regression inputs by dividing by two standard deviations. Statistics in Medicine, 27(15):2865–2873, July 10, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martinez-Beneito:2008:AAS

[2834] M. A. Martínez-Beneito, A. López-Quilez, and P. Botella-Rocamora. An autoregressive approach to spatio-temporal disease mapping. Statistics in Medicine, 27(15):2874–2889, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Luo:2008:ARE

[2835] Xianghua Luo and Gary S. Sorock. Analysis of recurrent event data under the case-crossover design with applications to elderly falls. Statistics in Medicine, 27(15):2890–2901, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bottai:2008:TUA

[2836] Matteo Bottai, Marco Geraci, and Andrew Lawson. Testing for unusual aggregation of health risk in semiparametric models. Statistics in Medicine, 27(15):2902–2921, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martinez-Montes:2008:EER

[2837] Eduardo Martínez-Montes, Elena R. Cuspineda-Bravo, Wael El-Deredy, José M. Sánchez-Bornot, Agustín Lage-Castellanos, and Pedro A. Valdés-Sosa. Exploring event-related brain dynamics with tests on complex valued time-frequency representations. Statistics in Medicine, 27(15): 2922–2947, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Irigoien:2008:INS

[2838] I. Irigoien and C. Arenas. INCA: New statistic for estimating the number of clusters and identifying atypical units. Statistics in Medicine, 27(15): 2948–2973, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ziegler:2008:CMR

[2839] Andreas Ziegler, Friedrich Pahlke, and Inke R. König. Comments on ‘Mendelian randomization: using genes as instruments for making causal inferences in epidemiology’ by Debbie A. Lawlor, R. M. Harbord, J. A. Sterne, N. Timpson and G. Davey Smith, Statistics in Medicine, DOI: 10.1002/sim.3034. Statistics in Medicine, 27(15):2974–2976, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawlor:2008:CMR

[2840] Debbie A. Lawlor, Nic Timpson, George Davey Smith, Roger M. Harbord, and Jonathan A. C. Sterne. Comments on ‘Mendelian randomization: Using genes as instruments for making causal inference in epidemiology’: Authors’ response. Statistics in Medicine, 27(15):2976–2978, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schaalje:2008:BRL

[2841] G. Bruce Schaalje. Book review: Linear mixed models, a practical guide using statistical software. Brady T. West, Kathleen B. Welch, Andrzej T. Galecki, Chapman and Hall/CRC, Boca Raton, FL, 2007. No. of pages: xix + 353. Price: $71.96. ISBN: 1-58488-480-0. Statistics in Medicine, 27(15):2979–2980, July 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diggle:2008:JMR

[2842] Peter J. Diggle, Inês Sousa, and Amanda G. Chetwynd. Joint modelling of repeated measurements and time-to-event outcomes: The Fourth Armitage Lecture. Statistics in Medicine, 27(16):2981–2998, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2008:LBM

[2843] L. Forsberg White and M. Pagano. A likelihood-based method for realtime estimation of the serial interval and reproductive number of an epidemic. Statistics in Medicine, 27(16):2999–3016, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2008:TLS

[2844] Xin-Yuan Song, Sik-Yum Lee, and Yih-Ing Hser. A two-level structural equation model approach for analyzing multivariate longitudinal responses. Statistics in Medicine, 27(16):3017–3041, July 20, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2008:VCM

[2845] Colin O. Wu, Xin Tian, and Heejung Bang. A varying-coefficient model for the evaluation of time-varying concomitant intervention effects in longitudinal studies. Statistics in Medicine, 27(16):3042–3056, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noren:2008:SMD

[2846] G. Niklas Norén, Rolf Sundberg, Andrew Bate, and I. Ralph Edwards. A statistical methodology for drug-drug interaction surveillance. Statistics in Medicine, 27(16):3057–3070, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2008:EDI

[2847] Hong-Bin Fang, Douglas D. Ross, Edward Sausville, and Ming Tan. Experimental design and interaction analysis of combination studies of drugs with log-linear dose responses. Statistics in Medicine, 27(16):3071– 3083, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2008:RBS

[2848] Peng Huang, Robert F. Woolson, and Peter C. O’Brien. A rank-based sample size method for multiple outcomes in clinical trials. Statistics in Medicine, 27(16):3084–3104, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:LRM

[2849] Lei Liu and Zhangsheng Yu. A likelihood reformulation method in nonnormal random effects models. Statistics in Medicine, 30(5):3105–3124, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Litiere:2008:IMR

[2850] S. Litière, A. Alonso, and G. Molenberghs. The impact of a misspecified random-effects distribution on the estimation and the performance of inferential procedures in generalized linear mixed models. Statistics in Medicine, 30(5):3125–3144, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koyama:2008:PIS

[2851] Tatsuki Koyama and Heidi Chen. Proper inference from Simon’s twostage designs. Statistics in Medicine, 29(20):3145–3154, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2008:LSA

[2852] Hui Xie. A local sensitivity analysis approach to longitudinal non-Gaussian data with non-ignorable dropout. Statistics in Medicine, 29 (20):3155–3177, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2008:MAI

[2853] Xiao Song and Shuangge Ma. Multiple augmentation for intervalcensored data with measurement error. Statistics in Medicine, 29(20): 3178–3190, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moskowitz:2008:EEL

[2854] Chaya S. Moskowitz, Venkatraman E. Seshan, Elyn R. Riedel, and Colin B. Begg. Estimating the empirical Lorenz curve and Gini coefficient in the presence of error with nested data. Statistics in Medicine, 29(20):3191–3208, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:2008:CCI

[2855] Jun mo Nam. Comments on ‘Confidence intervals for a ratio of binomial proportions based on paired data’ by D. G. Bonett and R. M. Price, Statistics in Medicine 2006; 25:3039–3047. Statistics in Medicine, 29 (20):3209–3210, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2122].

Friede:2008:BRT

[2856] Tim Friede. Book review: The theory of response-adaptive randomization in clinical trials. Feifang Hu and William F. Rosenberger, Wiley, Hoboken, NJ, 2006. No. of pages: xiii + 232. Price: $96.00. ISBN 13: 978-0-471-65396-7. Statistics in Medicine, 29(20):3211–3212, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leroux:2008:BRM

[2857] Brian Leroux. Book reviews: Multilevel and longitudinal modeling using Stata. Sophia Rabe-Hesketh and Anders Skrondal, Stata Press, College Station, 2005. No. of pages: xxiii + 317. Price: $54.00. ISBN 10: 1-59718-008-4, ISBN 13: 978-1-59718-008-5. Statistics in Medicine, 29(20): 3212–3213, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peters:2008:BRM

[2858] Gareth Peters. Book reviews: Markov Chain Monte Carlo: stochastic simulation for Bayesian inference. (2nd edn). Dani Gamerman and Hedibert F. Lopes, Chapman and Hall/CRC, Boca Raton, FL, 2006. No. of pages: xvii + 323. Price: $69.95. ISBN 10: 1-58488-587-4, ISBN 13: 978-1-58488-587-0. Statistics in Medicine, 29(20):3213–3214, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robinson:2008:BRI

[2859] Andrew Robinson. Book review: Introduction to nonparametric regression. Kunio Takezawa, John Wiley and Sons Inc., Hoboken, New Jersey, 2006. No. of pages: xviii + 538. Price: $123.00. ISBN: 978-0-471-74583-9. Statistics in Medicine, 29(20):3214–3215, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borghi:2008:C

[2860] Elaine Borghi. Correction. Statistics in Medicine, 29(20):3216, July 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2008:GLR

[2861] Jinlong Huang, Chinsan Lee, and Qiqing Yu. A generalized log-rank test for interval-censored failure time data via multiple imputation. Statistics in Medicine, 27(17):3217–3226, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wood:2008:HSV

[2862] Angela M. Wood, Ian R. White, and Patrick Royston. How should variable selection be performed with multiply imputed data? Statistics in Medicine, 27(17):3227–3246, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Su:2008:BSR

[2863] Li Su and Joseph W. Hogan. Bayesian semiparametric regression for longitudinal binary processes with missing data. Statistics in Medicine, 27(17):3247–3268, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buenconsejo:2008:BHM

[2864] Joan Buenconsejo, Durland Fish, James E. Childs, and Theodore R. Holford. A Bayesian hierarchical model for the estimation of two incomplete surveillance data sets. Statistics in Medicine, 27(17):3269–3285, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2008:UBI

[2865] Peter C. Austin. Using the bootstrap to improve estimation and confidence intervals for regression coefficients selected using backwards variable elimination. Statistics in Medicine, 27(17):3286–3300, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2008:CIR

[2866] M. Tian, M. L. Tang, H. K. T. Ng, and P. S. Chan. Confidence intervals for the risk ratio under inverse sampling. Statistics in Medicine, 27(17): 3301–3324, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2008:DNP

[2867] Huazhen Lin, Paul S. F. Yip, and Richard M. Huggins. A doublenonparametric procedure for estimating the number of delay-reported cases. Statistics in Medicine, 27(17):3325–3339, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2008:NIR

[2868] Yongming Qu and Yanping Wang. Nonparametric inference on the E/O ratio in model validation. Statistics in Medicine, 27(17):3340–3349, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2008:SSC

[2869] Sin-Ho Jung. Sample size calculation for the weighted rank statistics with paired survival data. Statistics in Medicine, 27(17):3350–3365, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lora:2008:BBP

[2870] Mayra Ivanoff Lora and Julio M. Singer. Beta-binomial/Poisson regression models for repeated bivariate counts. Statistics in Medicine, 27(17): 3366–3381, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biswas:2008:RACa

[2871] Pinaki Biswas and John D. Kalbfleisch. A risk-adjusted CUSUM in continuous time based on the Cox model. Statistics in Medicine, 27(17): 3382–3406, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fricker:2008:CSS

[2872] Ronald D. Fricker, Jr., Benjamin L. Hegler, and David A. Dunfee. Comparing syndromic surveillance detection methods: EARS’ versus a CUSUM-based methodology. Statistics in Medicine, 27(17):3407–3429, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eilers:2008:MMS

[2873] Paul H. C. Eilers, Jutta Gampe, Brian D. Marx, and Roland Rau. Modulation models for seasonal time series and incidence tables. Statistics in Medicine, 27(17):3430–3441, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alexander:2008:PRE

[2874] Neal Alexander. Precision of rate estimation under uniform interval censoring. Statistics in Medicine, 27(17):3442–3445, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dmitrienko:2008:NTG

[2875] Alex Dmitrienko, Ajit C. Tamhane, Lingyun Liu, and Brian L. Wiens. A note on tree gatekeeping procedures in clinical trials. Statistics in Medicine, 27(17):3446–3451, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biswas:2008:RACb

[2876] Pinaki Biswas and John D. Kalbfleisch. A risk-adjusted CUSUM in continuous time based on the Cox model. Statistics in Medicine, 27 (17):3452, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2008:C

[2877] Anonymous. Correction. Statistics in Medicine, 27(17):3452, July 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2008:EMM

[2878] Babette Brumback and Arthur Berg. On effect-measure modification: Relationships among changes in the relative risk, odds ratio, and risk difference. Statistics in Medicine, 27(18):3453–3465, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:2008:MEC

[2879] Bernard Rosner, Karin B. Michels, Ya-Hua Chen, and Nicholas E. Day. Measurement error correction for nutritional exposures with correlated measurement error: Use of the method of triads in a longitudinal setting. Statistics in Medicine, 27(18):3466–3489, August 15, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3202].

Chakraborty:2008:MMO

[2880] Arindom Chakraborty and Kalyan Das. Mixed models for ordinal data: a pharmacokinetic study on the effectiveness of drug for the reduction of epileptic seizures. Statistics in Medicine, 27(18):3490–3502, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ramamoorthi:2008:AMR

[2881] R. V. Ramamoorthi, M. G. Rossano, N. Paneth, J. C. Gardiner, M. P. Diamond, E. Puscheck, D. C. Daly, R. C. Potter, and J. J. Wirth. An application of multivariate ranks to assess effects from combining factors: Metal exposures and semen analysis outcomes. Statistics in Medicine, 27 (18):3503–3514, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kawaguchi:2008:AFA

[2882] Atsushi Kawaguchi, Koji Yonemoto, Yumihiro Tanizaki, Yutaka Kiyohara, Takashi Yanagawa, and Young K. Truong. Application of functional ANOVA models for hazard regression to the Hisayama data. Statistics in Medicine, 27(18):3515–3527, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:MLT

[2883] Lei Liu, Jennie Z. Ma, and Bankole A. Johnson. A multi-level twopart random effects model, with application to an alcohol-dependence study. Statistics in Medicine, 27(18):3528–3539, August 15, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lloyd:2008:MPE

[2884] Chris J. Lloyd and Max V. Moldovan. A more powerful exact test of noninferiority from binary matched-pairs data. Statistics in Medicine, 27(18):3540–3549, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsou:2008:PRI

[2885] Tsung-Shan Tsou and Chung-Wei Shen. Parametric robust inferences for correlated ordinal data. Statistics in Medicine, 27(18):3550–3562, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cortese:2008:DRH

[2886] Giuliana Cortese and Thomas H. Scheike. Dynamic regression hazards models for relative survival. Statistics in Medicine, 27(18):3563–3584, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pagano:2008:SMC

[2887] Eva Pagano, Michele Petrinco, Alessandro Desideri, Riccardo Bigi, Franco Merletti, and Dario Gregori. Survival models for cost data: The forgotten additive approach. Statistics in Medicine, 27(18):3585–3597, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crans:2008:HCF

[2888] Gerald G. Crans and Jonathan J. Shuster. How conservative is Fisher’s exact test? A quantitative evaluation of the two-sample comparative binomial trial. Statistics in Medicine, 27(18):3598–3611, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3105].

Zhou:2008:JSS

[2889] Huafeng Zhou, Andrew B. Lawson, James R. Hebert, Elizabeth H. Slate, and Elizabeth G. Hill. Joint spatial survival modeling for the age at diagnosis and the vital outcome of prostate cancer. Statistics in Medicine, 27(18):3612–3628, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lefebvre:2008:IMS

[2890] Geneviève Lefebvre, Joseph A. C. Delaney, and Robert W. Platt. Impact of mis-specification of the treatment model on estimates from a marginal structural model. Statistics in Medicine, 27(18):3629–3642, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chai:2008:ULS

[2891] High Seng Chai and Kent R. Bailey. Use of log-skew-normal distribution in analysis of continuous data with a discrete component at zero. Statistics in Medicine, 27(18):3643–3655, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2008:AIP

[2892] Feng-Chang Xie, Bo-Cheng Wei, and Jin-Guan Lin. Assessing influence for pharmaceutical data in zero-inflated generalized Poisson mixed models. Statistics in Medicine, 27(18):3656–3673, August 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kelley:2008:ZIO

[2893] Mary E. Kelley and Stewart J. Anderson. Zero inflation in ordinal data: Incorporating susceptibility to response through the use of a mixture model. Statistics in Medicine, 27(18):3674–3688, August 15, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bembom:2008:ASR

[2894] Oliver Bembom and Mark J. van der Laan. Analyzing sequentially randomized trials based on causal effect models for realistic individualized treatment rules. Statistics in Medicine, 27(19):3689–3716, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frost:2008:ODC

[2895] Chris Frost, Michael G. Kenward, and Nick C. Fox. Optimizing the design of clinical trials where the outcome is a rate. can estimating a baseline rate in a run-in period increase efficiency? Statistics in Medicine, 27(19):3717–3731, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sankoh:2008:NCC

[2896] Abdul J. Sankoh. A note on the conservativeness of the confidence interval approach for the selection of non-inferiority margin in the two-arm active-control trial. Statistics in Medicine, 27(19):3732–3742, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganju:2008:SSE

[2897] Jitendra Ganju, Allen Izu, and Alessandra Anemona. Sample size for equivalence trials: a case study from a vaccine lot consistency trial. Statistics in Medicine, 27(19):3743–3754, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3108].

Guolo:2008:SBC

[2898] A. Guolo and A. R. Brazzale. A simulation-based comparison of techniques to correct for measurement error in matched case-control studies. Statistics in Medicine, 27(19):3755–3775, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghosh:2008:UMF

[2899] S. Ghosh and D. K. Dey. A unified modeling framework for metabonomic profile development and covariate selection for acute trauma subjects. Statistics in Medicine, 27(19):3776–3788, August 30, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2008:MHS

[2900] Hao Wang and Daniel F. Heitjan. Modeling heaping in self-reported cigarette counts. Statistics in Medicine, 27(19):3789–3804, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Woo:2008:EPS

[2901] Mi-Ja Woo, Jerome P. Reiter, and Alan F. Karr. Estimation of propensity scores using generalized additive models. Statistics in Medicine, 27 (19):3805–3816, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kvist:2008:RET

[2902] Kajsa Kvist, Per Kragh Andersen, Jules Angst, and Lars Vedel Kessing. Repeated events and total time on test. Statistics in Medicine, 27 (19):3817–3832, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lim:2008:DFT

[2903] Johan Lim, Xinlei Wang, Seokho Lee, and Sin-Ho Jung. A distributionfree test of constant mean in linear mixed effects models. Statistics in Medicine, 27(19):3833–3846, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2008:TWG

[2904] Zhaoxia Yu, Liewei Wang, Michelle A. T. Hildebrandt, and Daniel J. Schaid. Testing whether genetic variation explains correlation of quantitative measures of gene expression, and application to genetic network analysis. Statistics in Medicine, 27(19):3847–3867, August 30, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hegarty:2008:BDM

[2905] Avril Hegarty and Daniel Barry. Bayesian disease mapping using product partition models. Statistics in Medicine, 27(19):3868–3893, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:MCA

[2906] Mengling Liu, Wenbin Lu, and Yongzhao Shao. A Monte Carlo approach for change-point detection in the Cox proportional hazards model. Statistics in Medicine, 27(19):3894–3909, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carpenter:2008:CFV

[2907] James Carpenter, Gerta Rücker, and Guido Schwarzer. Comments on ‘Fixed vs random effects meta-analysis in rare event studies: the rosiglitazone link with myocardial infarction and cardiac death’ by J. J. Shuster, L. S. Jones and D. A. Salmon, Statistics in Medicine 2007; 26:4375– 4385. Statistics in Medicine, 27(19):3910–3912, August 30, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2539].

Shuster:2008:RCA

[2908] Jonathan J. Shuster, Lynn S. Jones, and Daniel A. Salmon. Rebuttal to Carpenter et al. comments on ‘fixed vs. random effects meta-analysis in rare event studies: The rosiglitazone link with myocardial infarction and cardiac death’. Statistics in Medicine, 27(19):3912–3914, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3137].

Graf:2008:CPD

[2909] Erika Graf and Martin Schumacher. Comments on ‘The performance of different propensity score methods for estimating marginal odds ratios’ by Peter C. Austin, Statistics in Medicine 2007; 26 (16):3078–3094. Statistics in Medicine, 27(19):3915–3917, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2008:LEP

[2910] Peter C. Austin. Letter to the Editor: The performance of different propensity score methods for estimating marginal odds ratios, Statistics in Medicine 2007; 26:3078–3094. Statistics in Medicine, 27(19):3918– 3920, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2443, 3024].

Sjolander:2008:NMH

[2911] Arvid Sjölander. A note on ‘The method of handling non-compliance in clinical trials suggested by Sato/Matsuyama’ by T. Sato, Statistics in Medicine 2001; 20:2761–2774. Y. Matsuyama, Statistics in Medicine 2002; 21:675–687. Statistics in Medicine, 27(19):3920–3924, August 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [582].

Stephens:2008:WOR

[2912] James W. Stephens. Welcome and opening remarks 2007 CDC–ATSDR Symposium on Statistical Methods. Statistics in Medicine, 27(20):3925– 3926, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kennickell:2008:WDE

[2913] Arthur B. Kennickell. What is the difference? Evidence on the distribution of wealth, health, life expectancy, and health insurance coverage. Statistics in Medicine, 27(20):3927–3940, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duan:2008:DDD

[2914] Naihua Duan, Xiao-Li Meng, Julia Y. Lin, Chih nan Chen, and Margarita Alegria. Disparities in defining disparities: Statistical conceptual frameworks. Statistics in Medicine, 27(20):3941–3956, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2008:MIS

[2915] Jarvis T. Chen, Brent A. Coull, Pamela D. Waterman, Joel Schwartz, and Nancy Krieger. Methodologic implications of social inequalities for analyzing health disparities in large spatiotemporal data sets: an example using breast cancer incidence data (northern and southern California, 1988–2002). Statistics in Medicine, 27(20):3957–3983, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curtis:2008:HSH

[2916] Andrew Curtis. From Healthy Start to Hurricane Katrina: Using GIS to eliminate disparities in perinatal health. Statistics in Medicine, 27(20): 3984–3997, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Young:2008:AAB

[2917] Linda J. Young, Carol A. Gotway, Jie Yang, Greg Kearney, and Chris DuClos. Assessing the association between environmental impacts and health outcomes: a case study from Florida. Statistics in Medicine, 27 (20):3998–4015, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elliott:2008:SDM

[2918] Marc N. Elliott, Brian K. Finch, David Klein, Sai Ma, D. Phuong Do, Megan K. Beckett, Nathan Orr, and Nicole Lurie. Sample designs for measuring the health of small racial/ethnic subgroups. Statistics in Medicine, 27(20):4016–4029, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barker:2008:AEP

[2919] Lawrence Barker and Betsy L. Cadwell. An analysis of eight 95 per cent confidence intervals for a ratio of Poisson parameters when events are rare. Statistics in Medicine, 27(20):4030–4037, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malec:2008:SAR

[2920] Donald Malec and Jerry Maples. Small area random effects models for capture/recapture methods with applications to estimating coverage error in the U.S. Decennial Census. Statistics in Medicine, 27(20):4038– 4056, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kleinman:2008:AUP

[2921] Ken P. Kleinman and Allyson M. Abrams. Assessing the utility of public health surveillance using specificity, sensitivity, and lives saved. Statistics in Medicine, 27(20):4057–4068, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wheeler:2008:MED

[2922] David C. Wheeler, Lance A. Waller, and John O. Elliott. Modeling epilepsy disparities among ethnic groups in Philadelphia, PA. Statistics in Medicine, 27(20):4069–4085, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demirtas:2008:ISI

[2923] Hakan Demirtas and Donald Hedeker. An imputation strategy for incomplete longitudinal ordinal data. Statistics in Medicine, 27(20):4086–4093, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caudill:2008:MRQ

[2924] Samuel P. Caudill, Rosemary L. Schleicher, and James L. Pirkle. Multirule quality control for the age-related eye disease study. Statistics in Medicine, 27(20):4094–4106, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2008:RED

[2925] Philip J. Smith and John Stevenson. Racial/ethnic disparities in vaccination coverage by 19 months of age: an evaluation of the impact of missing data resulting from record scattering. Statistics in Medicine, 27 (20):4107–4118, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Du:2008:FAE

[2926] Hongyan Du, Veronica Valenzuela, Patricia Diaz, David Cella, and Elizabeth A. Hahn. Factors affecting enrollment in literacy studies for Englishand Spanish-speaking cancer patients. Statistics in Medicine, 27(20): 4119–4131, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tillerson:2008:ERD

[2927] Kristen Tillerson. Explaining racial disparities in HIV/AIDS incidence among women in the U.S.: a systematic review. Statistics in Medicine, 27(20):4132–4143, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malec:2008:C

[2928] Donald J. Malec. Correction. Statistics in Medicine, 27(20):4144, September 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zehetmayer:2008:OMS

[2929] Sonja Zehetmayer, Peter Bauer, and Martin Posch. Optimized multistage designs controlling the false discovery or the family-wise error rate. Statistics in Medicine, 27(21):4145–4160, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wolbers:2008:TSR

[2930] Marcel Wolbers and Jeffrey D. Helterbrand. Two-stage randomization designs in drug development. Statistics in Medicine, 27(21):4161–4174, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crawford:2008:TPM

[2931] S. B. Crawford and J. J. Hanfelt. Testing for the presence of multiple sources of informative dropout in longitudinal data. Statistics in Medicine, 27(21):4175–4189, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yan:2008:IET

[2932] Guofen Yan and Tom Greene. Investigating the effects of ties on measures of concordance. Statistics in Medicine, 27(21):4190–4206, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2008:IER

[2933] Sudhir R. Paul and Tasneem Zaihra. Interval estimation of risk difference for data sampled from clusters. Statistics in Medicine, 27(21):4207–4220, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:2008:CIN

[2934] Lili Tian. Confidence intervals for P (Y_{1} > Y_{2}) with normal outcomes in linear models. Statistics in Medicine, 27(21):4221–4237, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fagerland:2008:MGF

[2935] Morten W. Fagerland, David W. Hosmer, and Anna M. Bofin. Multinomial goodness-of-fit tests for logistic regression models. Statistics in Medicine, 27(21):4238–4253, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zimmerman:2008:SCF

[2936] Dale L. Zimmerman, Xiangming Fang, and Soumya Mazumdar. Spatial clustering of the failure to geocode and its implications for the detection of disease clustering. Statistics in Medicine, 27(21):4254–4266, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Copas:2008:RVT

[2937] John B. Copas and Paul F. Malley. A robust P -value for treatment effect in meta-analysis with publication bias. Statistics in Medicine, 27(21): 4267–4278, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Trikalinos:2008:MMA

[2938] Thomas A. Trikalinos and Ingram Olkin. A method for the meta-analysis of mutually exclusive binary outcomes. Statistics in Medicine, 27(21): 4279–4300, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2008:GDU

[2939] Christopher Cox. The generalized F distribution: an umbrella for parametric survival analysis. Statistics in Medicine, 27(21):4301–4312, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lau:2008:ECA

[2940] Bryan Lau, Stephen R. Cole, Richard D. Moore, and Stephen J. Gange. Evaluating competing adverse and beneficial outcomes using a mixture model. Statistics in Medicine, 27(21):4313–4327, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rahgozar:2008:PTS

[2941] Mehdi Rahgozar, Soghrat Faghihzadeh, Gholamreza Babaee Rouchi, and Yingwei Peng. The power of testing a semi-parametric shared gamma frailty parameter in failure time data. Statistics in Medicine, 27(21):

4328–4339, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fiocco:2008:RRP

[2942] Marta Fiocco, Hein Putter, and Hans C. van Houwelingen. Reducedrank proportional hazards regression and simulation-based prediction for multi-state models. Statistics in Medicine, 27(21):4340–4358, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2008:MML

[2943] Keunbaik Lee and Michael J. Daniels. Marginalized models for longitudinal ordinal data with application to quality of life studies. Statistics in Medicine, 27(21):4359–4380, September 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arends:2008:MAS

[2944] Lidia R. Arends, M. G. Myriam Hunink, and Theo Stijnen. Meta-analysis of summary survival curve data. Statistics in Medicine, 27(22):4381– 4396, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berglund:2008:MLE

[2945] Lars Berglund, Hans Garmo, Johan Lindbäck, Kurt Svärdsudd, and Björn Zethelius. Maximum likelihood estimation of correction for dilution bias in simple linear regression using replicates from subjects with extreme first measurements. Statistics in Medicine, 27(22):4397–4407, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faes:2008:HDJ

[2946] Christel Faes, Marc Aerts, Geert Molenberghs, Helena Geys, Greet Teuns, and Luc Bijnens. A high-dimensional joint model for longitudinal outcomes of different nature. Statistics in Medicine, 27(22):4408–4427, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2008:NSM

[2947] Kuo-Chin Lin, Yi-Ju Chen, and Yu Shyr. A nonparametric smoothing method for assessing GEE models with longitudinal binary data. Statistics in Medicine, 27(22):4428–4439, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beaudoin:2008:ACM

[2948] D. Beaudoin and L. Lakhal-Chaieb. Archimedean copula model selection under dependent truncation. Statistics in Medicine, 27(22):4440–4454, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martinez-Beneito:2008:BMS

[2949] Miguel A. Martínez-Beneito, David Conesa, Antonio López-Quílez, and Aurora López-Maside. Bayesian Markov switching models for the early detection of influenza epidemics. Statistics in Medicine, 27(22):4455– 4468, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Menten:2008:BLC

[2950] Joris Menten, Marleen Boelaert, and Emmanuel Lesaffre. Bayesian latent class models with conditionally dependent diagnostic tests: a case study. Statistics in Medicine, 27(22):4469–4488, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2008:NLT

[2951] Zonghui Hu, Dean Follmann, Jing Qin, Robin L. Dewar, and Phumelele Sangweni. A nonparametric likelihood test for detecting discordance between two measurements with application to censored viral load determinations. Statistics in Medicine, 27(22):4489–4501, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAngelo:2008:IAC

[2952] Gina D’Angelo, Lisa Weissfeld, and on behalf of the Genims Investigators. An index approach for the Cox model with left censored covariates. Statistics in Medicine, 27(22):4502–4514, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brenning:2008:EER

[2953] Alexander Brenning and Berthold Lausen. Estimating error rates in the classification of paired organs. Statistics in Medicine, 27(22):4515–4531, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Katki:2008:MDC

[2954] Hormuzd A. Katki, Amanda Blackford, Sining Chen, and Giovanni Parmigiani. Multiple diseases in carrier probability estimation: Accounting for surviving all cancers other than breast and ovary in BRCAPRO.

Statistics in Medicine, 27(22):4532–4548, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2008:NMA

[2955] Xiaoshu Lu and Esa-Pekka Takala. A novel mathematical approach for modelling time-dependent musculoskeletal outcomes for office staff. Statistics in Medicine, 27(22):4549–4568, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:2008:EAW

[2956] Philip J. Smith and Lawrence C. Marsh. Evaluating assumptions of weighting class methods for partial response using a selection model. Statistics in Medicine, 27(22):4569–4580, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:2008:OTA

[2957] Weng Kee Wong and Wei Zhu. Optimum treatment allocation rules under a variance heterogeneity model. Statistics in Medicine, 27(22): 4581–4595, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galta:2008:TGA

[2958] R. El Galta, T. Stijnen, and J. J. Houwing-Duistermaat. Testing for genetic association: a powerful score test. Statistics in Medicine, 27(22): 4596–4609, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lumley:2008:BRA

[2959] Thomas Lumley. Book review: Analysis of correlated data with SAS and R (3rd edn). Mohamed M. Shoukri and Mohammad A. Chaudhary, Chapman and Hall/CRC, Boca Raton, 2007. No. of pages: 295. Price: $89.95. ISBN: 1-58488-619-6. Statistics in Medicine, 27(22):4610–4611, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:2008:BRA

[2960] Frank Miller and Stig Johan Wiklund. Book reviews: Adaptive design methods in clinical trials. Shein-Chung Chow and Mark Chang, Chapman and Hall/CRC, Boca Raton, FL, 2007. No. of pages: 277. Price: $89.95. ISBN 10: 1-58488-776-1, ISBN 13: 978-1-58488-776-8. Statistics in Medicine, 27(22):4611–4612, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Post:2008:BRU

[2961] Wendy Post. Book review: Using and understanding medical statistics (4th edn). David E. Matthews and Vernon T. Farewell, Karger, Basel, 2007. No. of pages: xx + 322. Price: CHF 49.00, EUR 35.00, $44.00. ISBN: 978-3-8055-8189-9. Statistics in Medicine, 27(22):4612– 4613, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harel:2008:C

[2962] Ofer Harel and X. H. (Andrew) Zhou. Correction. Statistics in Medicine, 27(22):4614–4615, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2008:C

[2963] Stephen Senn, Erika Graf, and Angelika Caputo. Correction. Statistics in Medicine, 27(22):4615, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Young:2008:C

[2964] Linda J. Young. Correction. Statistics in Medicine, 27(22):4616, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karon:2008:EHI

[2965] John M. Karon, Ruiguang Song, Ron Brookmeyer, Edward H. Kaplan, and H. Irene Hall. Estimating HIV incidence in the United States from HIV/AIDS surveillance data and biomarker HIV test results. Statistics in Medicine, 27(23):4617–4633, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gezmu:2008:P

[2966] Misrak Gezmu. Preface. Statistics in Medicine, 27(23):4634, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dixon:2008:SMH

[2967] Dennis Dixon. Statistical methods in HIV/AIDS. Statistics in Medicine, 27(23):4635, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:2008:A

[2968] Anonymous. Acknowledgement. Statistics in Medicine, 27(23):4636, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2008:ATS

[2969] Peng Zhang and Stephen W. Lagakos. Analysis of time to a silent event whose occurrence is monitored with error, with application to motherto-child HIV transmission. Statistics in Medicine, 27(23):4637–4646, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perelson:2008:EDE

[2970] Alan S. Perelson and Ruy M. Ribeiro. Estimating drug efficacy and viral dynamic parameters: HIV and HCV. Statistics in Medicine, 27 (23):4647–4657, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsiatis:2008:CAT

[2971] Anastasios A. Tsiatis, Marie Davidian, Min Zhang, and Xiaomin Lu. Covariate adjustment for two-sample treatment comparisons in randomized clinical trials: a principled yet flexible approach. Statistics in Medicine, 27(23):4658–4677, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:2008:EEO

[2972] James Robins, Liliana Orellana, and Andrea Rotnitzky. Estimation and extrapolation of optimal treatment and testing strategies. Statistics in Medicine, 27(23):4678–4721, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2008:ISC

[2973] Yangxin Huang, Hua Liang, and Hulin Wu. Identifying significant covariates for anti-HIV treatment response: Mechanism-based differential equation models and empirical semiparametric regression models. Statistics in Medicine, 27(23):4722–4739, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schumi:2008:RBA

[2974] Jennifer Schumi and Victor DeGruttola. Resampling-based analyses of the effects of combinations of HIV genetic mutations on drug susceptibility. Statistics in Medicine, 27(23):4740–4757, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2008:ESE

[2975] Peter B. Gilbert, Li Qin, and Steven G. Self. Evaluating a surrogate endpoint at three levels, with application to vaccine development. Statistics in Medicine, 27(23):4758–4778, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pond:2008:ESP

[2976] S. L. Kosakovsky Pond, A. F. Y. Poon, S. Zárate, D. M. Smith, S. J. Little, S. K. Pillai, R. J. Ellis, J. K. Wong, A. J. Leigh Brown, D. D. Richman, and S. D. W. Frost. Estimating selection pressures on HIV-1 using phylogenetic likelihood models. Statistics in Medicine, 27(23):4779–4789, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahn:2008:PTL

[2977] C. Ahn, F. Seillier-Moiseiwitsch, and G. G. Koch. Predictive tests for linked changes. Statistics in Medicine, 27(23):4790–4804, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wick:2008:MEC

[2978] W. David Wick. On modeling the effects of T-cell vaccines on HIV acquisition and disease. Statistics in Medicine, 27(23):4805–4816, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gran:2008:GRE

[2979] Jon Michael Gran, Linn Wasmuth, Ellen J. Amundsen, Bo H. Lindqvist, and Odd O. Aalen. Growth rates in epidemic models: Application to a model for HIV/AIDS progression. Statistics in Medicine, 27(23):4817– 4834, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taffe:2008:JBC

[2980] Patrick Taffé, Margaret May, and the Swiss Hiv Cohort Study. A joint back calculation model for the imputation of the date of HIV infection in a prevalent cohort. Statistics in Medicine, 27(23):4835–4853, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:2008:C

[2981] C. P. Farrington. Correction. Statistics in Medicine, 27(23):4854–4855, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snapinn:2008:C

[2982] Steven Snapinn and Qi Jiang. Correction. Statistics in Medicine, 27 (23):4855–4856, October 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Imai:2008:VIE

[2983] Kosuke Imai. Variance identification and efficiency analysis in randomized experiments under the matched-pair design. Statistics in Medicine, 27(24):4857–4873, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:2008:SLA

[2984] Leena Choi, Brian Caffo, and Charles Rohde. A survey of the likelihood approach to bioequivalence trials. Statistics in Medicine, 27(24):4874– 4894, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2008:DSF

[2985] Yisheng Li, B. Nebiyou Bekele, Yuan Ji, and John D. Cook. Doseschedule finding in phase I/II clinical trials using a Bayesian isotonic transformation. Statistics in Medicine, 27(24):4895–4913, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strassburger:2008:CSL

[2986] Klaus Strassburger and Frank Bretz. Compatible simultaneous lower confidence bounds for the Holm procedure and other Bonferroni-based closed tests. Statistics in Medicine, 27(24):4914–4927, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holmgren:2008:QUP

[2987] Eric Holmgren. Quantifying the usefulness of PD biomarkers in phase 2 screening trials of oncology drugs. Statistics in Medicine, 27(24):4928– 4938, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2008:SDC

[2988] Mei-Jie Zhang and Jason Fine. Summarizing differences in cumulative incidence functions. Statistics in Medicine, 27(24):4939–4949, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mukherjee:2008:FSP

[2989] Bhramar Mukherjee, Jaeil Ahn, Ivy Liu, Paul J. Rathouz, and Brisa N. Sánchez. Fitting stratified proportional odds models by amalgamating conditional likelihoods. Statistics in Medicine, 27(24):4950–4971, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:2008:EWA

[2990] Babette A. Brumback, Larry H. Winner, George Casella, Malay Ghosh, Allyson Hall, Jianyi Zhang, Lorna Chorba, and Paul Duncan. Estimating a weighted average of stratum-specific parameters. Statistics in Medicine, 27(24):4972–4991, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horton:2008:MLL

[2991] Nicholas J. Horton, Kevin Roberts, Louise Ryan, Shakira Franco Suglia, and Rosalind J. Wright. A maximum likelihood latent variable regression model for multiple informants. Statistics in Medicine, 27(24):4992–5004, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

French:2008:ALD

[2992] Benjamin French and Patrick J. Heagerty. Analysis of longitudinal data to evaluate a policy change. Statistics in Medicine, 27(24):5005–5025, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barus:2008:FRM

[2993] José Ignacio Muñoz Barús, Manuel Febrero-Bande, and Carmen Cadarso-Suárez. Flexible regression models for estimating postmortem interval (PMI) in forensic medicine. Statistics in Medicine, 27(24):5026– 5038, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McClelland:2008:ERF

[2994] Robyn L. McClelland, Richard A. Kronmal, Jeffrey Haessler, Roger S. Blumenthal, and David C. Goff, Jr. Estimation of risk factor associations when the response is influenced by medication use: an imputation approach. Statistics in Medicine, 27(24):5039–5053, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zheng:2008:CTP

[2995] Gang Zheng, Mark Meyer, Wentian Li, and Yaning Yang. Comparison of two-phase analyses for case-control genetic association studies. Statistics in Medicine, 27(24):5054–5075, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:2008:AFA

[2996] Abigail G. Matthews, Dianne M. Finkelstein, and Rebecca A. Betensky. Analysis of familial aggregation studies with complex ascertainment schemes. Statistics in Medicine, 27(24):5076–5092, October 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mielke:2008:ANI

[2997] Matthias Mielke, A. Munk, and A. Schacht. The assessment of noninferiority in a gold standard design with censored, exponentially distributed endpoints. Statistics in Medicine, 27(25):5093–5110, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2008:ESM

[2998] Lan Huang, Linda W. Pickle, and Barnali Das. Evaluating spatial methods for investigating global clustering and cluster detection of cancer cases. Statistics in Medicine, 27(25):5111–5142, November 10, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ukoumunne:2008:CRD

[2999] Obioha C. Ukoumunne, Andrew B. Forbes, John B. Carlin, and Martin C. Gulliford. Comparison of the risk difference, risk ratio and odds ratio scales for quantifying the unadjusted intervention effect in cluster randomized trials. Statistics in Medicine, 27(25):5143–5155, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dragalin:2008:TSD

[3000] Vladimir Dragalin, Valerii V. Fedorov, and Yuehui Wu. Two-stage design for dose-finding that accounts for both efficacy and safety. Statistics in Medicine, 27(25):5156–5176, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peng:2008:EMS

[3001] Yingwei Peng and Jiajia Zhang. Estimation method of the semiparametric mixture cure gamma frailty model. Statistics in Medicine, 27(25): 5177–5194, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:2008:CRC

[3002] Laurence S. Freedman, Douglas Midthune, Raymond J. Carroll, and Victor Kipnis. A comparison of regression calibration, moment reconstruction and imputation for adjusting for covariate measurement error in regression. Statistics in Medicine, 27(25):5195–5216, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deCastro:2008:HTE

[3003] Mário de Castro, Manuel Galea, and Heleno Bolfarine. Hypothesis testing in an errors-in-variables model with heteroscedastic measurement errors. Statistics in Medicine, 27(25):5217–5234, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seppanen:2008:PIM

[3004] Johanna Seppänen, Sirpa Heinävaara, and Timo Hakulinen. Predicting impacts of mass-screening policy changes on breast cancer mortality. Statistics in Medicine, 27(25):5235–5251, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthaus:2008:BMT

[3005] L. Matthäus, P. Trillenberg, T. Fadini, M. Finke, and A. Schweikard. Brain mapping with transcranial magnetic stimulation using a refined correlation ratio and Kendall’s τ. Statistics in Medicine, 27(25):5252– 5270, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2008:SPM

[3006] Xiao-Hua Zhou and Huazhen Lin. Semi-parametric maximum likelihood estimates for ROC curves of continuous-scale tests. Statistics in Medicine, 27(25):5271–5290, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2008:NSI

[3007] Yaohua He and Michael Escobar. Nonparametric statistical inference method for partial areas under receiver operating characteristic curves, with application to genomic studies. Statistics in Medicine, 27(25):5291– 5308, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perme:2008:CHR

[3008] Maja Pohar Perme and Per Kragh Andersen. Checking hazard regression models using pseudo-observations. Statistics in Medicine, 27(25):5309– 5328, November 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerke:2008:OPD

[3009] Oke Gerke and Harald Siedentop. Optimal phase I dose-escalation trial designs in oncology — a simulation study. Statistics in Medicine, 27(26): 5329–5344, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shu:2008:DED

[3010] Jianfen Shu and John O’Quigley. Dose-escalation designs in oncology: ADEPT and the CRM. Statistics in Medicine, 27(26):5345–5353, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerke:2008:ARD

[3011] Oke Gerke and Harald Siedentop. Authors’ rejoinder to ‘dose-escalation designs in oncology: ADEPT and the CRM’. Statistics in Medicine, 27 (26):5354–5355, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quan:2008:CBI

[3012] Hui Quan, Qiankun Sun, Ji Zhang, and Weichung J. Shih. Comparisons between ITT and treatment emergent adverse event analyses. Statistics in Medicine, 27(26):5356–5376, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2008:EMA

[3013] Jianjun (David) Li and Devan V. Mehrotra. An efficient method for accommodating potentially underpowered primary endpoints. Statistics in Medicine, 27(26):5377–5391, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ng:2008:NHC

[3014] Tie-Hua Ng. Noninferiority hypotheses and choice of noninferiority margin. Statistics in Medicine, 27(26):5392–5406, November 20, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gu:2008:BBE

[3015] Jiezhun Gu, Subhashis Ghosal, and Anindya Roy. Bayesian bootstrap estimation of ROC curve. Statistics in Medicine, 27(26):5407–5420, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doehler:2008:SIS

[3016] Kirsten Doehler and Marie Davidian. ‘Smooth’ inference for survival functions with arbitrarily censored data. Statistics in Medicine, 27(26): 5421–5439, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vandenHout:2008:MSA

[3017] Ardo van den Hout and Fiona E. Matthews. Multi-state analysis of cognitive ability data: a piecewise-constant model and a Weibull model. Statistics in Medicine, 27(26):5440–5455, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Magaret:2008:IVS

[3018] Amalia S. Magaret. Incorporating validation subsets into discrete proportional hazards models for mismeasured outcomes. Statistics in Medicine, 27(26):5456–5470, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2008:GAF

[3019] Hye-Seung Lee, Myunghee Cho Paik, and Joseph H. Lee. Genotypeadjusted familial correlation analysis using three generalized estimating equations. Statistics in Medicine, 27(26):5471–5483, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berhane:2008:UTP

[3020] Kiros Berhane, Michael Hauptmann, and Bryan Langholz. Using tensor product splines in modeling exposure-time-response relationships: Application to the Colorado plateau uranium miners cohort. Statistics in Medicine, 27(26):5484–5496, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Price:2008:CIR

[3021] Robert M. Price and Douglas G. Bonett. Confidence intervals for a ratio of two independent binomial proportions. Statistics in Medicine, 27(26): 5497–5508, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:2008:EMF

[3022] Liming Cai, Nathaniel Schenker, James Lubitz, Paula Diehr, Alice Arnold, and Linda P. Fried. Evaluation of a method for fitting a semi-Markov process model in the presence of left-censored spells using the Cardiovascular Health Study. Statistics in Medicine, 27(26):5509–5524, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:2008:EEL

[3023] Charlotte L. Nelson, Jie L. Sun, Anastasios A. Tsiatis, and Daniel B. Mark. Empirical estimation of life expectancy from large clinical trials: Use of left-truncated, right-censored survival analysis methodology. Statistics in Medicine, 27(26):5525–5555, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Forbes:2008:IPW

[3024] Andrew Forbes and Susan Shortreed. Inverse probability weighted estimation of the marginal odds ratio: Correspondence regarding ‘The performance of different propensity score methods for estimating marginal odds ratios’ by P. Austin, Statictics in Medicine, 2007; 26:3078–3094. Statistics in Medicine, 27(26):5556–5559, November 20, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2443, 2910].

Austin:2008:IPW

[3025] Peter C. Austin. Inverse probability weighted estimation of the marginal odds ratio. Statistics in Medicine, 27(26):5560–5563, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gerke:2008:C

[3026] Oke Gerke and Harald Siedentop. Correction. Statistics in Medicine, 27(26):5564, November 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jo:2008:ITA

[3027] Booil Jo, Tihomir Asparouhov, and Bengt O. Muthén. Intention-to-treat analysis in cluster randomized trials with noncompliance. Statistics in Medicine, 27(27):5565–5577, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giraudeau:2008:SSC

[3028] B. Giraudeau, P. Ravaud, and A. Donner. Sample size calculation for cluster randomized cross-over trials. Statistics in Medicine, 27(27):5578– 5585, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wathen:2008:BAM

[3029] J. Kyle Wathen and Peter F. Thall. Bayesian adaptive model selection for optimizing group sequential clinical trials. Statistics in Medicine, 27 (27):5586–5604, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dias:2008:BFR

[3030] Sofia Dias, Roseanne McNamee, and Andy Vail. Bias in frequently reported analyses of subfertility trials. Statistics in Medicine, 27(27):5605– 5619, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Welton:2008:MTC

[3031] N. J. Welton, N. J. Cooper, A. E. Ades, G. Lu, and A. J. Sutton. Mixed treatment comparison with multiple outcomes reported inconsistently across trials: Evaluation of antivirals for treatment of influenza A and B. Statistics in Medicine, 27(27):5620–5639, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becque:2008:RPL

[3032] T. Becque and I. R. White. Regaining power lost by non-compliance via full probability modelling. Statistics in Medicine, 27(27):5640–5663, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yuan:2008:SCR

[3033] Ying Yuan and Guosheng Yin. Sequential continual reassessment method for two-dimensional dose finding. Statistics in Medicine, 27(27):5664– 5678, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:JAM

[3034] Lei Liu, Jennie Z. Ma, and John O’Quigley. Joint analysis of multi-level repeated measures data and survival: an application to the end stage renal disease (ESRD) data. Statistics in Medicine, 27(27):5679–5691, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lai:2008:LTS

[3035] Xin Lai and Kelvin K. W. Yau. Long-term survivor model with bivariate random effects: Applications to bone marrow transplant and carcinoma study data. Statistics in Medicine, 27(27):5692–5708, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ruan:2008:ACI

[3036] Ping K. Ruan and Robert J. Gray. Analyses of cumulative incidence functions via non-parametric multiple imputation. Statistics in Medicine, 27(27):5709–5724, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kirkham:2008:CHP

[3037] Jamie J. Kirkham. A comparison of hospital performance with nonignorable missing covariates: an application to trauma care data. Statistics in Medicine, 27(27):5725–5744, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sanchez:2008:CDP

[3038] B. N. Sánchez, T. E. Raghunathan, A. V. Diez Roux, Y. Zhu, and O. Lee. Combining data from primary and ancillary surveys to assess the association between neighborhood-level characteristics and health outcomes: the multi-ethnic study of artherosclerosis. Statistics in Medicine, 27(27): 5745–5763, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Preisser:2008:FSA

[3039] John S. Preisser, Bing Lu, and Bahjat F. Qaqish. Finite sample adjustments in estimating equations and covariance estimators for intracluster correlations. Statistics in Medicine, 27(27):5764–5785, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dominicus:2008:RCP

[3040] Annica Dominicus, Samuli Ripatti, Nancy L. Pedersen, and Juni Palmgren. A random change point model for assessing variability in repeated measures of cognitive function. Statistics in Medicine, 27(27):5786–5798, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andres:2008:CAC

[3041] A. Martín Andrés and M. Álvarez Hernández. Comments on ‘Activecontrol trials with binary data: a comparison of methods for testing superiority or non-inferiority using the odds ratio’ by A. L. Siqueira, A. Whitehead and S. Todd, Statistics in Medicine 2008; 27:353–370. Statistics in Medicine, 27(27):5799–5800, November 29, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2655].

Xu:2008:TLD

[3042] Jinfeng Xu, Yaning Yang, Zhiliang Ying, and Jurg Ott. Testing linkage disequilibrium from pooled DNA: a contingency table perspective. Statistics in Medicine, 27(28):5801–5815, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lunceford:2008:SPC

[3043] Jared K. Lunceford and Nancy Liu. Statistical performance of cladistic strategies for haplotype grouping in pharmacogenetics. Statistics in Medicine, 27(28):5816–5833, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lusa:2008:CCM

[3044] Lara Lusa, Edward L. Korn, and Lisa M. McShane. A class comparison method with filtering-enhanced variable selection for high-dimensional data sets. Statistics in Medicine, 27(28):5834–5849, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2008:TEP

[3045] Honghua Jiang, James Symanowski, Sofia Paul, Yongming Qu, Anthony Zagar, and Shengyan Hong. The type I error and power of non-parametric logrank and Wilcoxon tests with adjustment for covariates — a simulation study. Statistics in Medicine, 27(28):5850–5860, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shardell:2008:ICI

[3046] Michelle Shardell, Daniel O. Scharfstein, David Vlahov, and Noya Galai. Inference for cumulative incidence functions with informatively coarsened discrete event-time data. Statistics in Medicine, 27(28):5861–5879, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deGroot:2008:MIC

[3047] J. A. H. de Groot, K. J. M. Janssen, A. H. Zwinderman, K. G. M. Moons, and J. B. Reitsma. Multiple imputation to correct for partial verification bias revisited. Statistics in Medicine, 27(28):5880–5889, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Amorim:2008:RST

[3048] Leila D. Amorim, Jianwen Cai, Donglin Zeng, and Maurício L. Barreto. Regression splines in the time-dependent coefficient rates model for recurrent event data. Statistics in Medicine, 27(28):5890–5906, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2008:ESB

[3049] Huafeng Zhou and Andrew B. Lawson. EWMA smoothing and Bayesian spatial modeling for health surveillance. Statistics in Medicine, 27(28): 5907–5928, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yin:2008:BTC

[3050] Guosheng Yin. Bayesian transformation cure frailty models with multivariate failure time data. Statistics in Medicine, 27(28):5929–5940, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bogaerts:2008:ELG

[3051] Kris Bogaerts and Emmanuel Lesaffre. Estimating local and global measures of association for bivariate interval censored data with a smooth estimate of the density. Statistics in Medicine, 27(28):5941–5955, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2008:GCE

[3052] S. D. Walter, A. Gafni, and S. Birch. A geometric confidence ellipse approach to the estimation of the ratio of two variables. Statistics in Medicine, 27(28):5956–5974, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3288].

Lang:2008:SPL

[3053] Joseph B. Lang. Score and profile likelihood confidence intervals for contingency table parameters. Statistics in Medicine, 27(28):5975–5990, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saarela:2008:NCC

[3054] Olli Saarela, Sangita Kulathinal, Elja Arjas, and Esa Läärä. Nested case-control data utilized for multiple outcomes: a likelihood approach and alternatives. Statistics in Medicine, 27(28):5991–6008, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maringwa:2008:ACD

[3055] John T. Maringwa, Helena Geys, Ziv Shkedy, Christel Faes, Geert Molenberghs, Marc Aerts, Karel Van Ammel, Ard Teisman, and Luc Bijnens. Analysis of cross-over designs with serial correlation within periods using semi-parametric mixed models. Statistics in Medicine, 27(28):6009–6033, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nie:2008:BRO

[3056] Lei Nie. Book review: Optimum experimental designs, with SAS. A. C. Atkinson, A. N. Donev and R. D. Tobias Oxford University Press, Oxford, 2007. No. of pages: xvi + 511. Price: £35.00 (paperback), £75.00 (hardback). ISBN 13: 978-0-19-929660-6, 978-0-19-929659-0. Statistics in Medicine, 27(28):6034–6035, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hayen:2008:BRB

[3057] Andrew Hayen. Book review: Bayesian biostatistics and diagnostic medicine. Lyle D. Broemeling, Chapman and Hall/CRC, Boca Raton, FL, 2007. No. of pages: xiii + 198. Price: $79.95. ISBN 13: 97-1-584-88767-6. Statistics in Medicine, 27(28):6035–6036, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ferrari:2008:BMM

[3058] Pietro Ferrari, Raymond J. Carroll, Paul Gustafson, and Elio Riboli. A Bayesian multilevel model for estimating the diet/disease relationship in a multicenter study with exposures measured with error: The EPIC study. Statistics in Medicine, 27(29):6037–6054, December 20, 2008. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olsen:2008:SAM

[3059] Maren K. Olsen, Elizabeth R. DeLong, Eugene Z. Oddone, and Hayden B. Bosworth. Strategies for analyzing multilevel cluster-randomized studies with binary outcomes collected at varying intervals of time. Statistics in Medicine, 27(29):6055–6071, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Higgins:2008:MAS

[3060] Julian P. T. Higgins, Ian R. White, and Judith Anzures-Cabrera. Metaanalysis of skewed data: Combining results reported on log-transformed or raw scales. Statistics in Medicine, 27(29):6072–6092, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biggerstaff:2008:EDC

[3061] Brad J. Biggerstaff and Dan Jackson. The exact distribution of Cochran’s heterogeneity statistic in one-way random effects meta-analysis. Statistics in Medicine, 27(29):6093–6110, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Riley:2008:MAD

[3062] Richard D. Riley, Susanna R. Dodd, Jean V. Craig, John R. Thompson, and Paula R. Williamson. Meta-analysis of diagnostic test studies using individual patient data and aggregate data. Statistics in Medicine, 27 (29):6111–6136, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:2008:SFE

[3063] Lloyd J. Edwards, Keith E. Muller, Russell D. Wolfinger, Bahjat F. Qaqish, and Oliver Schabenberger. An R^{2} statistic for fixed effects in the linear mixed model. Statistics in Medicine, 27(29):6137–6157, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Swartz:2008:FFI

[3064] Michael D. Swartz, Robert K. Yu, and Sanjay Shete. Finding factors influencing risk: Comparing Bayesian stochastic search and standard variable selection methods applied to logistic regression models of cases and controls. Statistics in Medicine, 27(29):6158–6174, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:2008:CTS

[3065] Yunchan Chi and Chia-Min Chen. Curtailed two-stage designs in phase II clinical trials. Statistics in Medicine, 27(29):6175–6189, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:2008:AHD

[3066] Yujun Wu and Weichung J. Shih. Approaches to handling data when a phase II trial deviates from the pre-specified Simon’s two-stage design. Statistics in Medicine, 27(29):6190–6208, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2008:GSD

[3067] Nigel Stallard and Tim Friede. A group-sequential design for clinical trials with treatment selection. Statistics in Medicine, 27(29):6209–6227, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Putter:2008:JAM

[3068] Hein Putter, Tineke Vos, Hanneke de Haes, and Hans van Houwelingen. Joint analysis of multiple longitudinal outcomes: Application of a latent class model. Statistics in Medicine, 27(29):6228–6249, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:2008:MMI

[3069] M. Paul, L. Held, and A. M. Toschke. Multivariate modelling of infectious disease surveillance data. Statistics in Medicine, 27(29):6250–6267, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunning:2008:CES

[3070] Andrew J. Dunning. Comment on ‘Evaluating a surrogate endpoint at three levels, with application to vaccine development’ by Peter B. Gilbert, Li Qin and Steven G. Self, Statistics in Medicine, DOI: 10.1002/sim.3122.

Statistics in Medicine, 27(29):6268–6270, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Noren:2008:C

[3071] G. Niklas Norén, Rolf Sundberg, Andrew Bate, and I. Ralph Edwards. Correction. Statistics in Medicine, 27(29):6271–6272, December 20, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kenward:2008:P

[3072] Mike Kenward and Giota Touloumi. Preface. Statistics in Medicine, 27 (30):6273–6275, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Philipson:2008:CRM

[3073] Peter M. Philipson, Weang Kee Ho, and Robin Henderson. Comparative review of methods for handling drop-out in longitudinal studies. Statistics in Medicine, 27(30):6276–6298, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2008:MIL

[3074] Li C. Liu. A model for incomplete longitudinal multivariate ordinal data. Statistics in Medicine, 27(30):6299–6309, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giorgi:2008:PMI

[3075] Roch Giorgi, Aurélien Belot, Jean Gaudart, and Guy Launoy. The performance of multiple imputation for missing covariate data within the context of regression relative survival analysis. Statistics in Medicine, 27 (30):6310–6331, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Messer:2008:MLM

[3076] Karen Messer and Loki Natarajan. Maximum likelihood, multiple imputation and regression calibration for measurement error adjustment. Statistics in Medicine, 27(30):6332–6350, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Funatogawa:2008:ALM

[3077] Takashi Funatogawa, Ikuko Funatogawa, and Masahiro Takeuchi. An autoregressive linear mixed effects model for the analysis of longitudinal data which include dropouts and show profiles approaching asymptotes. Statistics in Medicine, 27(30):6351–6366, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Funatogawa:2008:BAL

[3078] Ikuko Funatogawa, Takashi Funatogawa, and Yasuo Ohashi. A bivariate autoregressive linear mixed effects model for the analysis of longitudinal data. Statistics in Medicine, 27(30):6367–6378, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bogaerts:2008:MAB

[3079] Kris Bogaerts and Emmanuel Lesaffre. Modeling the association of bivariate interval-censored data using the copula approach. Statistics in Medicine, 27(30):6379–6392, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nikoloulopoulos:2008:MLC

[3080] Aristidis K. Nikoloulopoulos and Dimitris Karlis. Multivariate logit copula model with an application to dental data. Statistics in Medicine, 27 (30):6393–6406, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ambrogi:2008:ECU

[3081] Federico Ambrogi, Elia Biganzoli, and Patrizia Boracchi. Estimates of clinically useful measures in competing risks survival analysis. Statistics in Medicine, 27(30):6407–6425, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2008:JML

[3082] P. R. Williamson, R. Kolamunnage-Dona, P. Philipson, and A. G. Marson. Joint modelling of longitudinal and competing risks data. Statistics in Medicine, 27(30):6426–6438, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beyersmann:2008:ITD

[3083] Jan Beyersmann, Martin Wolkewitz, and Martin Schumacher. The impact of time-dependent bias in proportional hazards modelling. Statistics in Medicine, 27(30):6439–6454, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinze:2008:AIE

[3084] Georg Heinze and Daniela Dunkler. Avoiding infinite estimates of timedependent effects in small-sample survival studies. Statistics in Medicine, 27(30):6455–6469, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Binquet:2008:ESD

[3085] C. Binquet, M. Abrahamowicz, A. Mahboubi, V. Jooste, J. Faivre, C. Bonithon-Kopp, and C. Quantin. Empirical study of the dependence of the results of multivariable flexible survival analyses on model selection strategy. Statistics in Medicine, 27(30):6470–6488, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elgmati:2008:FMC

[3086] E. Elgmati, R. Fiaccone, R. Henderson, and M. Mohammadi. Frailty modelling for clustered recurrent incidence of diarrhoea. Statistics in Medicine, 27(30):6489–6504, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sypsa:2008:MVD

[3087] Vana Sypsa and Angelos Hatzakis. Modelling of viral dynamics in hepatitis B and hepatitis C clinical trials. Statistics in Medicine, 27(30): 6505–6521, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wolkewitz:2008:SEM

[3088] M. Wolkewitz, M. Dettenkofer, H. Bertz, M. Schumacher, and J. Huebner. Statistical epidemic modeling with hospital outbreak data. Statistics in Medicine, 27(30):6522–6531, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Calle:2008:ISD

[3089] M. L. Calle, V. Urrea, G. Vellalta, N. Malats, and K. V. Steen. Improving strategies for detecting genetic patterns of disease susceptibility in association studies. Statistics in Medicine, 27(30):6532–6546, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jorgensen:2008:MQP

[3090] Andrea L. Jorgensen and Paula R. Williamson. Methodological quality of pharmacogenetic studies: Issues of concern. Statistics in Medicine, 27 (30):6547–6569, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:2008:MAM

[3091] Tom M. Palmer, John R. Thompson, and Martin D. Tobin. Meta-analysis of Mendelian randomization studies incorporating all three genotypes.

Statistics in Medicine, 27(30):6570–6582, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2008:SID

[3092] S. D. Walter, A. S. Ismaila, and P. J. Devereaux. Statistical issues in the design and analysis of expertise-based randomized clinical trials. Statistics in Medicine, 27(30):6583–6596, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuroki:2008:FTB

[3093] Manabu Kuroki and Zhihong Cai. Formulating tightest bounds on causal effects in studies with unmeasured confounders. Statistics in Medicine, 27(30):6597–6611, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3219].

Molas:2008:CTR

[3094] Marek Molas and Emmanuel Lesaffre. A comparison of three random effects approaches to analyze repeated bounded outcome scores with an application in a stroke revalidation study. Statistics in Medicine, 27(30): 6612–6633, December 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:2009:UOA

[3095] Ralph B. D’Agostino, Vern Farewell, Joel B. Greenhouse, and John Whitehead. Updating our aims and scope for the future. Statistics in Medicine, 28(1):1–2, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alosh:2009:FST

[3096] Mohamed Alosh and Mohammad F. Huque. A flexible strategy for testing subgroups and overall population. Statistics in Medicine, 28(1):3–23, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganju:2009:RES

[3097] Jitendra Ganju and Biao Xing. Re-estimating the sample size of an on-going blinded trial based on the method of randomization block sums. Statistics in Medicine, 28(1):24–38, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moore:2009:CAR

[3098] K. L. Moore and M. J. van der Laan. Covariate adjustment in randomized trials with binary outcomes: Targeted maximum likelihood estimation. Statistics in Medicine, 28(1):39–64, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2009:APT

[3099] Lingqi Tang, Naihua Duan, Ruth Klap, Joan Rosenbaum Asarnow, and Thomas R. Belin. Applying permutation tests with adjustment for covariates and attrition weights to randomized trials of health-services interventions. Statistics in Medicine, 28(1):65–74, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sottas:2009:EHB

[3100] Pierre-Edouard Sottas, Jérome Lavoué, Raffaella Bruzzi, David Vernez, Nicole Charrière, and Pierre-Olivier Droz. An empirical hierarchical Bayesian unification of occupational exposure assessment methods. Statistics in Medicine, 28(1):75–93, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCandless:2009:BPS

[3101] Lawrence C. McCandless, Paul Gustafson, and Peter C. Austin. Bayesian propensity score analysis for observational data. Statistics in Medicine, 28(1):94–112, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2009:SBM

[3102] Li Zhang, Bhramar Mukherjee, Bo Hu, Victor Moreno, and Kathleen A. Cooney. Semiparametric Bayesian modeling of random genetic effects in family-based association studies. Statistics in Medicine, 28(1):113– 139, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Houwing-Duistermaat:2009:WSA

[3103] J. J. Houwing-Duistermaat, A. Callegaro, M. Beekman, R. G. Westendorp, P. E. Slagboom, and J. C. van Houwelingen. Weighted statistics for aggregation and linkage analysis of human longevity in selected families: The Leiden longevity study. Statistics in Medicine, 28(1):140–151, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bembom:2009:BDU

[3104] Oliver Bembom, Maya L. Petersen, Soo-Yon Rhee, W. Jeffrey Fessel, Sandra E. Sinisi, Robert W. Shafer, and Mark J. van der Laan. Biomarker discovery using targeted maximum-likelihood estimation: Application to the treatment of antiretroviral-resistant HIV infection. Statistics in Medicine, 28(1):152–172, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andres:2009:CHC

[3105] A. Martín Andrés and I. Herranz Tejedor. Comments on ‘How conservative is Fisher’s exact test? A quantitative evaluation of the two-sample comparative binomial trial’ by G. G. Crans and J. J. Shuster, Statistics in Medicine 2008; 27:3598–3611. Statistics in Medicine, 28(1):173–174, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2888].

Tang:2009:CSS

[3106] Yongqiang Tang. Comments on ‘Sample size evaluation for a multiply matched case-control study using the score test from a conditional logistic (discrete Cox PH) regression model’ by J. M. Lachin, Statistics in Medicine 2008; 27 (14):2509–2523. Statistics in Medicine, 28(1):175– 177, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2803, 1780, 12].

Lachin:2009:AR

[3107] John M. Lachin. Author’s reply. Statistics in Medicine, 28(1):177, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kohberger:2009:CSS

[3108] Robert C. Kohberger. Comments on ‘Sample size for equivalence trials: a case study from a vaccine lot consistency trial’ by J. Ganju, A. Izu and A. Anemona, Statistics in Medicine 2008; 27:3743–3754. Statistics in Medicine, 28(1):177–178, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2897].

Ganju:2009:AR

[3109] Jitendra Ganju, Allen Izu, and Alessandra Anemona. Authors’ reply. Statistics in Medicine, 28(1):178–179, January 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2009:AAR

[3110] Yichuan Zhao, Yue Zhou, and Meng Zhao. Analysis of additive risk model with high-dimensional covariates using partial least squares. Statistics in Medicine, 28(2):181–193, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:2009:CCA

[3111] Yannan Jiang, Alastair Scott, and Chris J. Wild. Case-control analysis with a continuous outcome variable. Statistics in Medicine, 28(2):194– 204, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cumming:2009:IER

[3112] Geoff Cumming. Inference by eye: Reading the overlap of independent confidence intervals. Statistics in Medicine, 28(2):205–220, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gardiner:2009:FER

[3113] Joseph C. Gardiner, Zhehui Luo, and Lee Anne Roman. Fixed effects, random effects and GEE: What are the differences? Statistics in Medicine, 28(2):221–239, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roca-Pardinas:2009:SVN

[3114] Javier Roca-Pardiñas, Carmen Cadarso-Suárez, Pablo G. Tahoces, and María J. Lado. Selecting variables in non-parametric regression models for binary response. an application to the computerized detection of breast cancer. Statistics in Medicine, 28(2):240–259, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Soulakova:2009:CST

[3115] Julia N. Soulakova. Comparison of several testing strategies for combination drug efficacy trials based on the closure principle. Statistics in Medicine, 28(2):260–273, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klingenberg:2009:PCD

[3116] B. Klingenberg. Proof of concept and dose estimation with binary responses under model uncertainty. Statistics in Medicine, 28(2):274–292, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wall:2009:MIH

[3117] Melanie M. Wall and Ran Li. Multiple indicator hidden Markov model with an application to medical utilization data. Statistics in Medicine, 28(2):293–310, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2009:CUF

[3118] K. F. Cheng. Combining unrelated family studies to improve the power of genetic association test. Statistics in Medicine, 28(2):311–325, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mathur:2009:RBP

[3119] Sunil K. Mathur. A run-based procedure to identify time-lagged gene clusters in microarray experiments. Statistics in Medicine, 28(2):326– 337, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jackson:2009:REQ

[3120] Dan Jackson and Jack Bowden. A re-evaluation of the ‘quantile approximation method’ for random effects meta-analysis. Statistics in Medicine, 28(2):338–348, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Du:2009:TIN

[3121] Pang Du and Liansheng Tang. Transformation-invariant and nonparametric monotone smooth estimation of ROC curves. Statistics in Medicine, 28(2):349–359, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenblad:2009:BRS

[3122] Andreas Rosenblad. Book review: Sample size calculations in clinical research. (2nd edn). Shein-Chung Chow, Jun Shao and Hansheng Wang, Chapman and Hall/CRC, Boca Raton, FL, 2008. No. of pages: xiv + 465 (hardcover). Price: $89.95. ISBN 1-58488-982-9. Statistics in Medicine, 28(2):360, January 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:2009:DEA

[3123] Hua He, Jeffrey M. Lyness, and Michael P. McDermott. Direct estimation of the area under the receiver operating characteristic curve in the presence of verification bias. Statistics in Medicine, 28(3):361–376, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2009:STE

[3124] Nigel Stallard. Simple tests for the external validation of mortality prediction scores. Statistics in Medicine, 28(3):377–388, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collaboration:2009:MAP

[3125] The Fibrinogen Studies Collaboration. Measures to assess the prognostic ability of the stratified Cox proportional hazards model. Statistics in Medicine, 28(3):389–411, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2009:GDC

[3126] Ivy Liu, Bhramar Mukherjee, Thomas Suesse, David Sparrow, and Sung Kyun Park. Graphical diagnostics to check model misspecification for the proportional odds regression model. Statistics in Medicine, 28(3):412–429, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shoukri:2009:BMI

[3127] Mohamed M. Shoukri and Allan Donner. Bivariate modeling of interobserver agreement coefficients. Statistics in Medicine, 28(3):430–440, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dendukuri:2009:MCD

[3128] Nandini Dendukuri, Alula Hadgu, and Liangliang Wang. Modeling conditional dependence between diagnostic tests: a multiple latent variable model. Statistics in Medicine, 28(3):441–461, February 10, 2009. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsu:2009:NCT

[3129] Chiu-Hsieh Hsu and Jeremy M. G. Taylor. Nonparametric comparison of two survival functions with dependent censoring via nonparametric multiple imputation. Statistics in Medicine, 28(3):462–475, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frisen:2009:ROS

[3130] M. Frisén, E. Andersson, and L. Schiöler. Robust outbreak surveillance of epidemics in Sweden. Statistics in Medicine, 28(3):476–493, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chikobvu:2009:BBR

[3131] Perpetual Chikobvu, Carl J. Lombard, Alan J. Flisher, Gary King, Loraine Townsend, and Martie Muller. Bias in a binary risk behaviour model subject to inconsistent reports and dropout in a south African high school cohort study. Statistics in Medicine, 28(3):494–509, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Corbiere:2009:PLA

[3132] Fabien Corbière, Daniel Commenges, Jeremy M. G. Taylor, and Pierre Joly. A penalized likelihood approach for mixture cure models. Statistics in Medicine, 28(3):510–524, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vickers:2009:NRI

[3133] Andrew J. Vickers, Elena B. Elkin, and Ewout Steyerberg. Net reclassification improvement and decision theory. Statistics in Medicine, 28 (3):525–526, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pencina:2009:RNR

[3134] Michael J. Pencina, Ralph B. D’Agostino, Sr., Ralph B. D’Agostino, Jr., and Ramachandran S. Vasan. Response to ‘net reclassification improvement and decision theory’ by Vickers et al. Statistics in Medicine, 28 (3):526–528, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andres:2009:CTH

[3135] A. Martín Andrés and I. Herranz Tejedor. Comments on ‘Tests for the homogeneity of two binomial proportions in extremely unbalanced 2 × 2 contingency tables’, by S.-H. Kang and C. W. Ahn, Statistics in Medicine 2008; 27:2524–2535. Statistics in Medicine, 28(3):528–531, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2804].

Lui:2009:CET

[3136] Kung-Jong Lui. Comments on ‘On estimating treatment effects under non-compliance in randomized clinical trials: Are intent-to-treat or instrumental variables analyses perfect solutions?’ by Heejung Bang and Clarence E. Davis, Statistics in Medicine 2007; 26:954–964. Statistics in Medicine, 28(3):531–534, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2297].

Waksman:2009:CRC

[3137] Joel Waksman and Christine Kollar. Comments on ‘Rebuttal to Carpenter et al. Comments on ‘Fixed vs random effects meta-analysis in rare event studies: The rosiglitazone link with myocardial infarction and cardiac death’ by J. J. Shuster, L. S. Jones and D. A. Salmon, Statistics in Medicine 2008; 27:3912–3914’. Statistics in Medicine, 28(3):534–536, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2908].

Schemper:2009:BRM

[3138] Michael Schemper. Book review: Multivariable model-building: a pragmatic approach to regression analysis based on fractional polynomials for modelling continuous variables. Patrick Royston and Willi Sauerbrei, Wiley, Chichester, 2008. No. of pages: 322. Price: $130.00. ISBN: 978-0-470-02842-1. Statistics in Medicine, 28(3):537–538, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bandos:2009:BRA

[3139] Andriy I. Bandos and Howard E. Rockette. Book review: Analyzing receiver operating characteristic curves with SAS(R). Mithat Gönen, SAS Institute Inc., Cary, NC, 2007. No. of pages: x + 134. Price: $31.95. ISBN 978-1-59994-298-8. Statistics in Medicine, 28(3):538–539, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vittinghoff:2009:SSC

[3140] E. Vittinghoff, Ś. Sen, and C. E. McCulloch. Sample size calculations for evaluating mediation. Statistics in Medicine, 28(4):541–557, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sjolander:2009:BND

[3141] Arvid Sjölander. Bounds on natural direct effects in the presence of confounded intermediate variables. Statistics in Medicine, 28(4):558– 571, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaciroti:2009:BSM

[3142] Niko A. Kaciroti, M. Anthony Schork, Trivellore Raghunathan, and Stevo Julius. A Bayesian sensitivity model for intention-to-treat analysis on binary outcomes with dropouts. Statistics in Medicine, 28(4):572– 585, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2009:GAS

[3143] Frank Bretz, Willi Maurer, Werner Brannath, and Martin Posch. A graphical approach to sequentially rejective multiple test procedures. Statistics in Medicine, 28(4):586–604, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gunnes:2009:ACD

[3144] Nina Gunnes, Daniel M. Farewell, Taral G. Seierstad, and Odd O. Aalen. Analysis of censored discrete longitudinal data: Estimation of mean response. Statistics in Medicine, 28(4):605–624, February 20, 2009. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:2009:EAU

[3145] Man-Lai Tang, Man-Ho Ling, and Guo-Liang Tian. Exact and approximate unconditional confidence intervals for proportion difference in the presence of incomplete data. Statistics in Medicine, 28(4):625– 641, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hin:2009:WCS

[3146] Lin-Yee Hin and You-Gan Wang. Working-correlation-structure identification in generalized estimating equations. Statistics in Medicine, 28 (4):642–658, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paddock:2009:SPD

[3147] Susan M. Paddock and Patricia Ebener. Subjective prior distributions for modeling longitudinal continuous outcomes with non-ignorable dropout. Statistics in Medicine, 28(4):659–678, February 20, 2009. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:2009:SSD

[3148] Kaifeng Lu, Devan V. Mehrotra, and Guanghan Liu. Sample size determination for constrained longitudinal data analysis. Statistics in Medicine, 28(4):679–699, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2009:NMM

[3149] Donghui Zhang, Chunpeng Fan, Juan Zhang, and Cun-Hui Zhang. Nonparametric methods for measurements below detection limit. Statistics in Medicine, 28(4):700–715, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilbert:2009:RAD

[3150] Peter B. Gilbert, Li Qin, and Steven G. Self. Response to Andrew Dunning’s comment on ‘Evaluating a surrogate endpoint at three levels, with application to vaccine development’. Statistics in Medicine, 28(4):716– 719, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giraudeau:2009:C

[3151] B. Giraudeau, P. Ravaud, and A. Donner. Correction. Statistics in Medicine, 28(4):720, February 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rucker:2009:WAA

[3152] Gerta Rücker, Guido Schwarzer, James Carpenter, and Ingram Olkin. Why add anything to nothing? The arcsine difference as a measure of treatment effect in meta-analysis with zero cells. Statistics in Medicine, 28(5):721–738, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burman:2009:RFC

[3153] C.-F. Burman, C. Sonesson, and O. Guilbaud. A recycling framework for the construction of Bonferroni-based multiple tests. Statistics in Medicine, 28(5):739–761, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:2009:CES

[3154] Margaret Sullivan Pepe, Ziding Feng, Gary Longton, and Joseph Koopmeiners. Conditional estimation of sensitivity and specificity from a phase 2 biomarker study allowing early termination for futility. Statistics in Medicine, 28(5):762–779, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:2009:EDA

[3155] Paul S. Albert. Estimating diagnostic accuracy of multiple binary tests with an imperfect reference standard. Statistics in Medicine, 28(5):780– 797, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2009:EST

[3156] Shu-Chuan Lin, Paul H. Kvam, and Jye-Chyi Lu. Extending the skill test for disease diagnosis. Statistics in Medicine, 28(5):798–813, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crespi:2009:USO

[3157] Catherine M. Crespi, Weng Kee Wong, and Shiraz I. Mishra. Using second-order generalized estimating equations to model heterogeneous intraclass correlation in cluster-randomized trials. Statistics in Medicine, 28(5):814–827, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:2009:OTS

[3158] John Whitehead and Thomas Jaki. One- and two-stage design proposals for a phase II trial comparing three active treatments with control using an ordered categorical endpoint. Statistics in Medicine, 28(5):828–847, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheng:2009:BAA

[3159] Dunlei Cheng, James D. Stamey, and Adam J. Branscum. Bayesian approach to average power calculations for binary regression models with misclassified outcomes. Statistics in Medicine, 28(5):848–863, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:2009:PSS

[3160] Daowen Zhang and Hui Quan. Power and sample size calculation for log-rank test with a time lag in treatment effect. Statistics in Medicine, 28(5):864–879, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:2009:ASR

[3161] T. J. Cole, S. Stanojevic, J. Stocks, A. L. Coates, J. L. Hankinson, and A. M. Wade. Age- and size-related reference ranges: a case study of spirometry through childhood and adulthood. Statistics in Medicine, 28 (5):880–898, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deGroot:2009:MIC

[3162] J. A. H. de Groot, K. J. M. Janssen, A. H. Zwinderman, K. G. M. Moons, and J. B. Reitsma. Multiple imputation to correct for partial verification bias revisited (5880-5889). Statistics in Medicine, 28(5):899, February 28, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Viallon:2009:HEC

[3163] Vivian Viallon, Stéphane Ragusa, Françoise Clavel-Chapelon, and Jacques Bénichou. How to evaluate the calibration of a disease risk prediction tool. Statistics in Medicine, 28(6):901–916, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kimani:2009:DSS

[3164] Peter K. Kimani, Nigel Stallard, and Jane L. Hutton. Dose selection in seamless phase II/III clinical trials based on efficacy and safety. Statistics in Medicine, 28(6):917–936, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seaman:2009:DRG

[3165] Shaun Seaman and Andrew Copas. Doubly robust generalized estimating equations for longitudinal data. Statistics in Medicine, 28(6):937–955, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beyersmann:2009:SCR

[3166] Jan Beyersmann, Aurélien Latouche, Anika Buchholz, and Martin Schumacher. Simulating competing risks data in survival analysis. Statistics in Medicine, 28(6):956–971, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2009:JML

[3167] Lei Liu. Joint modeling longitudinal semi-continuous data and survival, with application to longitudinal medical cost data. Statistics in Medicine, 28(6):972–986, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buzkova:2009:SMR

[3168] Petra Buzková and Thomas Lumley. Semiparametric modeling of repeated measurements under outcome-dependent follow-up. Statistics in Medicine, 28(6):987–1003, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:2009:RFN

[3169] Chung Chang and R. Todd Ogden. Robust fitting for neuroreceptor mapping. Statistics in Medicine, 28(6):1004–1016, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heo:2009:SSR

[3170] Moonseong Heo and Andrew C. Leon. Sample size requirements to detect an intervention by time interaction in longitudinal cluster randomized clinical trials. Statistics in Medicine, 28(6):1017–1027, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:2009:EPP

[3171] Helena Chmura Kraemer. Events per person-time (incidence rate): a misleading statistic? Statistics in Medicine, 28(6):1028–1039, March 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ramasundarahettige:2009:CIC

[3172] Chinthanie F. Ramasundarahettige, Allan Donner, and G. Y. Zou. Confidence interval construction for a difference between two dependent intraclass correlation coefficients. Statistics in Medicine, 28(7):1041–1053, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tunes-da-Silva:2009:RAM

[3173] G. Tunes da Silva and J. P. Klein. Regression analysis of mean quality-adjusted survival time based on pseudo-observations. Statistics in Medicine, 28(7):1054–1066, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collaboration:2009:CMM

[3174] Fibrinogen Studies Collaboration. Correcting for multivariate measurement error by regression calibration in meta-analyses of epidemiological studies. Statistics in Medicine, 28(7):1067–1092, March 30, 2009. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:2009:MAV

[3175] Haonan Wang. Measures of agreement for vectorcardiography data. Statistics in Medicine, 28(7):1093–1107, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gallop:2009:MAP

[3176] Robert Gallop, Dylan S. Small, Julia Y. Lin, Michael R. Elliott, Marshall Joffe, and Thomas R. Ten Have. Mediation analysis with principal stratification. Statistics in Medicine, 28(7):1108–1130, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:2009:GLM

[3177] Inkyung Jung. A generalized linear models approach to spatial scan statistics for covariate adjustment. Statistics in Medicine, 28(7):1131– 1143, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonzo:2009:CTR

[3178] Todd A. Alonzo, Christos T. Nakas, Constantin T. Yiannoutsos, and Sherri Bucher. A comparison of tests for restricted orderings in the three-class case. Statistics in Medicine, 28(7):1144–1158, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lydersen:2009:RTA

[3179] Stian Lydersen, Morten W. Fagerland, and Petter Laake. Recommended tests for association in 2 × 2 tables. Statistics in Medicine, 28(7):1159– 1175, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:2009:BRE

[3180] Thomas A. Louis. Book review: Elementary Bayesian Biostatistics. Lemuel A. Moyé, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: xxi + 377, plus CD. Price: $79.95. ISBN-13: 978-1-58488-724-9 (hardcover). Statistics in Medicine, 28(7):1176–1177, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hothorn:2009:BRS

[3181] Ludwig A. Hothorn. Book review: Statistical issues in drug development. (2nd edn). Stephen Senn, John Wiley and Sons Ltd, Chichester, 2007. No. of pages: xix + 498. Price: $130.00. ISBN 978-0-470-01877-4. Statistics in Medicine, 28(7):1177–1178, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holford:2009:BRS

[3182] Nicholas Hg Holford. Book review: Statistical thinking for nonstatisticians in drug regulation. Richard Kay, Wiley, Chichester, 2007. No. of pages: xviii + 276. Price: $80.00, £39.95, EUR54.00 (hardcover).

ISBN 978-0-470-31971-0. Statistics in Medicine, 28(7):1179, March 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bretz:2009:ADC

[3183] Frank Bretz, Franz Koenig, Werner Brannath, Ekkehard Glimm, and Martin Posch. Adaptive designs for confirmatory clinical trials. Statistics in Medicine, 28(8):1181–1217, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collaboration:2009:SMC

[3184] The Fibrinogen Studies Collaboration. Systematically missing confounders in individual participant data meta-analysis of observational cohort studies. Statistics in Medicine, 28(8):1218–1237, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Timbie:2009:CMH

[3185] Justin W. Timbie, David M. Shahian, Joseph P. Newhouse, Meredith B. Rosenthal, and Sharon-Lise T. Normand. Composite measures for hospital quality using quality-adjusted life years. Statistics in Medicine, 28(8): 1238–1254, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bonetti:2009:SSS

[3186] Marco Bonetti, David Zahrieh, Bernard F. Cole, and Richard D. Gelber. A small sample study of the STEPP approach to assessing treatmentcovariate interactions in survival data. Statistics in Medicine, 28(8): 1255–1268, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2009:NEF

[3187] Rui Feng, Leslie A. McClure, Hemant K. Tiwari, and George Howard. A new estimate of family disease history providing improved prediction of disease risks. Statistics in Medicine, 28(8):1269–1283, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:2009:MRE

[3188] Keunbaik Lee, Yongsung Joo, Jae Keun Yoo, and JungBok Lee. Marginalized random effects models for multivariate longitudinal binary data. Statistics in Medicine, 28(8):1284–1300, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hunsberger:2009:FMS

[3189] Sally Hunsberger, Paul S. Albert, and Wendy B. London. A finite mixture survival model to characterize risk groups of neuroblastoma. Statistics in Medicine, 28(8):1301–1314, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:2009:CCT

[3190] Mitchell H. Gail, Barry Graubard, David F. Williamson, and Katherine M. Flegal. Comments on ‘Choice of time scale and its effect on significance of predictors in longitudinal studies’ by Michael J. Pencina, Martin G. Larson and Ralph B. D’Agostino, Statistics in Medicine 2007; 26:1343–1359. Statistics in Medicine, 28(8):1315–1317, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2322].

Shrier:2009:PS

[3191] Ian Shrier. Propensity scores. Statistics in Medicine, 28(8):1317–1318, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Metcalf:2009:BRP

[3192] Patricia Metcalf. Book review: Presenting medical statistics from proposal to publication. Janet Peacock and Sally Kerry, Oxford University Press, Oxford, 2007. No. of pages: xii + 174. Price: £24.95 (paperback). ISBN: 978-0-19-859966-1. Statistics in Medicine, 28(8):1319–1320, April 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stallard:2009:OCN

[3193] Nigel Stallard, Martin Posch, Tim Friede, Franz Koenig, and Werner Brannath. Optimal choice of the number of treatments to be included in a clinical trial. Statistics in Medicine, 28(9):1321–1338, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2009:BDS

[3194] Lili Zhao and George Woodworth. Bayesian decision sequential analysis with survival endpoint in phase II clinical trials. Statistics in Medicine, 28(9):1339–1352, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Healy:2009:MDS

[3195] Brian C. Healy and David Engler. Modeling disease-state transition heterogeneity through Bayesian variable selection. Statistics in Medicine, 28(9):1353–1368, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MacNab:2009:BMD

[3196] Ying C. MacNab. Bayesian multivariate disease mapping and ecological regression with errors in covariates: Bayesian estimation of DALYs and ‘preventable’ DALYs. Statistics in Medicine, 28(9):1369–1385, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sego:2009:RAM

[3197] Landon H. Sego, Marion R. Reynolds, Jr., and William H. Woodall. Risk-adjusted monitoring of survival times. Statistics in Medicine, 28 (9):1386–1401, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Qu:2009:PSE

[3198] Yongming Qu and Ilya Lipkovich. Propensity score estimation with missing values using a multiple imputation missingness pattern (MIMP) approach. Statistics in Medicine, 28(9):1402–1414, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pearl:2009:RMP

[3199] Judea Pearl. Remarks on the method of propensity score. Statistics in Medicine, 28(9):1415–1416, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sjolander:2009:PSS

[3200] Arvid Sjölander. Propensity scores and M-structures. Statistics in Medicine, 28(9):1416–1420, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:2009:SOS

[3201] Donald B. Rubin. Should observational studies be designed to allow lack of balance in covariate distributions across treatment groups? Statistics in Medicine, 28(9):1420–1423, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bingham:2009:RME

[3202] Sheila Bingham. Response to ‘Measurement error correction for nutritional exposures with correlated measurement error: Use of the method of triads in a longitudinal setting’ by Rosner et al. Statistics in Medicine 2008; 27:3466–3489. Statistics in Medicine, 28(9):1423–1424, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2879].

Chernick:2009:BRD

[3203] Michael R. Chernick. Book review: Design and analysis of bioavailability and bioequivalence studies. (3rd edn). Shein-Chung Chow and Jen-pei Liu, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: 760. Price: $99.95. ISBN 978-1-58488-668-6. Statistics in Medicine, 28(9): 1425–1426, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobbins:2009:BRL

[3204] Timothy Dobbins. Book review: Learning SAS by example: a programmer’s guide. Ron Cody, SAS Institute Inc: Cary, NC, 2007. No. of pages: xxxii + 626 (CD included). Price: $69.95. ISBN: 978-1-59994-165-3. Statistics in Medicine, 28(9):1427–1428, April 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gotte:2009:ADC

[3205] Heiko Götte, Gerhard Hommel, and Andreas Faldum. Adaptive designs with correlated test statistics. Statistics in Medicine, 28(10):1429–1444, May 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brannath:2009:CAD

[3206] Werner Brannath, Emmanuel Zuber, Michael Branson, Frank Bretz, Paul Gallo, Martin Posch, and Amy Racine-Poon. Confirmatory adaptive designs with Bayesian decision tools for a targeted therapy in oncology. Statistics in Medicine, 28(10):1445–1463, May 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jewell:2009:QSI

[3207] Nicholas P. Jewell. Quantifying the source of infection for HIV-infected hemophiliacs in the U.K. from 1979 to 1984. Statistics in Medicine, 28 (10):1464–1472, May 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fu:2009:CBM

[3208] Chong Yau Fu and Ya-Wen Yang. A comparison of binary models dealing with an extraneous effect relating to the main risk factor, but not relating to the outcome variable. Statistics in Medicine, 28(10):1473–1486, May 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fagerland:2009:WMW

[3209] Morten W. Fagerland and Leiv Sandvik. The Wilcoxon–Mann–Whitney test under scrutiny. Statistics in Medicine, 28(10):1487–1497, May 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Egleston:2009:SAI

[3210] Brian L. Egleston and Yu-Ning Wong. Sensitivity analysis to investigate the impact of a missing covariate on survival analyses using cancer registry data. Statistics in Medicine, 28(10):1498–1511, May 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Metcalf:2009:UMF

[3211] Patricia Metcalf and Alastair Scott. Using multiple frames in health surveys. Statistics in Medicine, 28(10):1512–1523, May 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bollen:2009:PAVa

[3212] Kenneth A. Bollen, Richard D. Lennox, and Darren L. Dahly. Practical application of the vanishing tetrad test for causal indicator measurement models: an example from health-related quality of life. Statistics in Medicine, 28(10):1524–1536, May 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Won:2009:COM

[3213] Sungho Won, Nathan Morris, Qing Lu, and Robert C. Elston. Choosing an optimal method to combine P -values. Statistics in Medicine, 28(11): 1537–1553, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [3310, 3386].

Sommen:2009:MAE

[3214] Cécile Sommen, Ahmadou Alioum, and Daniel Commenges. A multistate approach for estimating the incidence of human immunodeficiency virus by using HIV and AIDS French surveillance data. Statistics in Medicine, 28(11):1554–1568, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBuuren:2009:SLD

[3215] Stef van Buuren and Jeroen C. L. Ooms. Stage line diagram: an ageconditional reference diagram for tracking development. Statistics in Medicine, 28(11):1569–1579, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hossain:2009:BAC

[3216] Shahadut Hossain and Paul Gustafson. Bayesian adjustment for covariate measurement errors: a flexible parametric approach. Statistics in Medicine, 28(11):1580–1600, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:2009:BAJ

[3217] Wenhua Hu, Gang Li, and Ning Li. A Bayesian approach to joint analysis of longitudinal measurements and competing risks failure time data. Statistics in Medicine, 28(11):1601–1619, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:2009:SAR

[3218] Xiao-Hua Zhou, Nan Hu, Guizhou Hu, and Martin Root. Synthesis analysis of regression models with a continuous outcome. Statistics in Medicine, 28(11):1620–1635, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sjolander:2009:CFT

[3219] Arvid Sjölander. Comments on ‘Formulating the tightest bounds on causal effects in studies with unmeasured confounders’ by M. Kuroki and Z. Cai, Statistics in Medicine 2008; 27:6597–6611. Statistics in Medicine, 28(11):1636–1638, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3093].

Kuroki:2009:AR

[3220] Manabu Kuroki. Author’s reply. Statistics in Medicine, 28(11):1638– 1641, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2009:LEC

[3221] G. Y. Zou. Letter to the Editor: Confidence intervals for the 50 per cent response dose by D. Faraggi, P. Izikson and B. Reiser, Statistics in Medicine 2003; 22 (12):1977–1988. Statistics in Medicine, 28(11): 1641–1642, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:2009:BRI

[3222] Melissa D. Begg. Book review: An introduction to categorical data analysis. (2nd edn). Alan Agresti, John Wiley and Sons, Inc., Hoboken, New Jersey, 2007. No. of Pages: 400. Price: $100.95. ISBN: 978-0-471-22618-5. Statistics in Medicine, 28(11):1643, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:2009:C

[3223] T. J. Cole, S. Stanojevic, J. Stocks, A. L. Coates, J. L. Hankinson, and A. M. Wade. Correction. Statistics in Medicine, 28(11):1644, May 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:2009:ARM

[3224] Hayley E. Jones and David J. Spiegelhalter. Accounting for regression-to-the-mean in tests for recent changes in institutional performance: Analysis and power. Statistics in Medicine, 28(12):1645–1667, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:2009:NIT

[3225] Jun mo Nam and Deukwoo Kwon. Non-inferiority tests for clustered matched-pair data. Statistics in Medicine, 28(12):1668–1679, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [3382].

Huang:2009:UST

[3226] Xuelin Huang, Jing Ning, Yisheng Li, Elihu Estey, Jean-Pierre Issa, and Donald A. Berry. Using short-term response information to facilitate adaptive randomization for survival clinical trials. Statistics in Medicine, 28(12):1680–1689, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:2009:BAT

[3227] Frank Miller, Tim Friede, and Meinhard Kieser. Blinded assessment of treatment effects utilizing information about the randomization block length. Statistics in Medicine, 28(12):1690–1706, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greco:2009:MCM

[3228] Fedele P. Greco and Carlo Trivisano. A multivariate CAR model for improving the estimation of relative risks. Statistics in Medicine, 28(12): 1707–1724, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cain:2009:IPC

[3229] Lauren E. Cain and Stephen R. Cole. Inverse probability-of-censoring weights for the correction of time-varying noncompliance in the effect of randomized highly active antiretroviral therapy on incident AIDS or death. Statistics in Medicine, 28(12):1725–1738, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomson:2009:MBC

[3230] Andrew Thomson, Richard Hayes, and Simon Cousens. Measures of between-cluster variability in cluster randomized trials with binary outcomes. Statistics in Medicine, 28(12):1739–1751, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schlattmann:2009:BRM

[3231] Peter Schlattmann. Book review: Meta-analysis — a guide to calibrating and combining statistical evidence. Elena Kulinskaya, Stephan Morgenthaler and Robert G. Staudte, John Wiley and Sons Ltd, Chichester, 2008. No. of pages: xic + 260. Price: £37.50; EUR 46.90. ISBN 978-0-470-02864-3. Statistics in Medicine, 28(12):1752, May 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teixeira-Pinto:2009:CBC

[3232] Armando Teixeira-Pinto and Sharon-Lise T. Normand. Correlated bivariate continuous and binary outcomes: Issues and applications. Statistics in Medicine, 28(13):1753–1773, June 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahmed:2009:BPS

[3233] Ismaïl Ahmed, Françoise Haramburu, Annie Fourrier-Réglat, Frantz Thiessard, Carmen Kreft-Jais, Ghada Miremont-Salamé, Bernard Bégaud, and Pascale Tubert-Bitter. Bayesian pharmacovigilance signal detection methods revisited in a multiple comparison setting. Statistics in Medicine, 28(13):1774–1792, June 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jia:2009:CPE

[3234] Juan Jia and Robert E. Weiss. Common predictor effects for multivariate longitudinal data. Statistics in Medicine, 28(13):1793–1804, June 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oron:2009:RDR

[3235] Assaf P. Oron and Peter D. Hoff. The k-in-a-row up-and-down design, revisited. Statistics in Medicine, 28(13):1805–1820, June 15, 2009. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alber:2009:MSA

[3236] Susan A. Alber and Robert E. Weiss. A model selection approach to analysis of variance and covariance. Statistics in Medicine, 28(13):1821– 1840, June 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perkins:2009:GRC

[3237] Neil J. Perkins, Enrique F. Schisterman, and Albert Vexler. Generalized ROC curve inference for a biomarker subject to a limit of detection and measurement error. Statistics in Medicine, 28(13):1841–1860, June 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cooper:2009:ABS

[3238] Nicola J. Cooper, Alex J. Sutton, Danielle Morris, A. E. Ades, and Nicky J. Welton. Addressing between-study heterogeneity and inconsistency in mixed treatment comparisons: Application to stroke prevention treatments in individuals with non-rheumatic atrial fibrillation. Statistics in Medicine, 28(14):1861–1881, June 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simonsen:2009:EII

[3239] J. Simonsen, K. Mølbak, G. Falkenhorst, K. A. Krogfelt, A. Linneberg, and P. F. M. Teunis. Estimation of incidences of infectious diseases based on antibody measurements. Statistics in Medicine, 28(14):1882–1895, June 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2009:HAM

[3240] Qingzhao Yu, Bin Li, and Richard Allen Scribner. Hierarchical additive modeling of nonlinear association with spatial correlations — an application to relate alcohol outlet density and neighborhood assault rates. Statistics in Medicine, 28(14):1896–1912, June 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Callegaro:2009:STA

[3241] A. Callegaro, H. C. van Houwelingen, and J. J. Houwing-Duistermaat. Score test for age at onset genetic linkage analysis in selected sibling pairs. Statistics in Medicine, 28(14):1913–1926, June 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spycher:2009:MMR

[3242] B. D. Spycher, C. E. Minder, and C. E. Kuehni. Multivariate modelling of responses to conditional items: New possibilities for latent class analysis. Statistics in Medicine, 28(14):1927–1939, June 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bazzoli:2009:FIM

[3243] Caroline Bazzoli, Sylvie Retout, and France Mentré. Fisher information matrix for nonlinear mixed effects multiple response models: Evaluation of the appropriateness of the first order linearization using a pharmacokinetic/pharmacodynamic model. Statistics in Medicine, 28(14): 1940–1956, June 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karvanen:2009:VCP

[3244] Juha Karvanen and Frank E. Harrell, Jr. Visualizing covariates in proportional hazards model. Statistics in Medicine, 28(14):1957–1966, June 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bang:2009:AR

[3245] Heejung Bang and Clarence E. Davis. Author’s reply. Statistics in Medicine, 28(14):1967–1968, June 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Swift:2009:DDN

[3246] Louise Swift, Susan Miles, Gill M. Price, Lee Shepstone, and Sam J. Leinster. Do doctors need statistics? Doctors’ use of and attitudes to probability and statistics. Statistics in Medicine, 28(15):1969–1981, July 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:2009:IMC

[3247] Ian R. White and Patrick Royston. Imputing missing covariate values for the Cox model. Statistics in Medicine, 28(15):1982–1998, July 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:2009:PSS

[3248] John M. Williamson, Hung-Mo Lin, and Hae-Young Kim. Power and sample size calculations for current status survival analysis. Statistics in Medicine, 28(15):1999–2011, July 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pradhan:2009:PEQ

[3249] Biswabrata Pradhan and Anup Dewanji. Parametric estimation of quality adjusted lifetime (QAL) distribution in progressive illness-death model. Statistics in Medicine, 28(15):2012–2027, July 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:2009:SST

[3250] Wentao Feng and Abdus S. Wahed. Sample size for two-stage studies with maintenance therapy. Statistics in Medicine, 28(15):2028–2041, July 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meilijson:2009:CPH

[3251] Isaac Meilijson and Sara Meilijson. Coping with participant heterogeneity in dichotomous responses. Statistics in Medicine, 28(15):2042–2053, July 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chefo:2009:SSC

[3252] Solomon Chefo and Alex Tsodikov. Stage-specific cancer incidence: an artificially mixed multinomial logit model. Statistics in Medicine, 28(15): 2054–2076, July 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wahed:2009:NGW

[3253] Abdus S. Wahed, The Minh Luong, and Jong-Hyeon Jeong. A new generalization of Weibull distribution with application to a breast cancer data set. Statistics in Medicine, 28(16):2077–2094, July 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Isaman:2009:IED

[3254] Deanna J. M. Isaman, Jacob Barhak, and Wen Ye. Indirect estimation of a discrete-state discrete-time model using secondary data analysis of regression data. Statistics in Medicine, 28(16):2095–2115, July 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kafadar:2009:ELB

[3255] Karen Kafadar and Philip C. Prorok. Effect of length biased sampling of unobserved sojourn times on the survival distribution when disease is screen detected. Statistics in Medicine, 29(10):2116–2146, July 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kedem:2009:TDS

[3256] Benjamin Kedem, Eun young Kim, Anastasia Voulgaraki, and Barry I. Graubard. Two-dimensional semiparametric density ratio modeling of testicular germ cell data. Statistics in Medicine, 29(10):2147–2159, July 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2009:RMC

[3257] Lei Li and Paul S. Levy. Regression modeling of combined data from multiple sample surveys. Statistics in Medicine, 29(10):2160–2169, July 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:2009:STP

[3258] Jonathan J. Shuster. Student t-tests for potentially abnormal data. Statistics in Medicine, 29(10):2170–2184, July 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brutti:2009:MPD

[3259] Pierpaolo Brutti, Fulvio De Santis, and Stefania Gubbiotti. Mixtures of prior distributions for predictive Bayesian sample size calculations in clinical trials. Statistics in Medicine, 28(17):2185–2201, July 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goel:2009:RDS

[3260] Sharad Goel and Matthew J. Salganik. Respondent-driven sampling as Markov chain Monte Carlo. Statistics in Medicine, 28(17):2202–2229, July 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:2009:MBM

[3261] Eugene M. Laska, Morris Meisner, and Joseph Wanderling. Model-based multiplicity estimation of population size. Statistics in Medicine, 28(17): 2230–2252, July 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Song:2009:BSA

[3262] Xin-Yuan Song, Ye-Mao Xia, and Sik-Yum Lee. Bayesian semiparametric analysis of structural equation models with mixed continuous and unordered categorical variables. Statistics in Medicine, 28(17):2253–2276, July 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baldwin:2009:HP

[3263] Peter Baldwin, Joseph Bernstein, and Howard Wainer. Hip psychometrics. Statistics in Medicine, 28(17):2277–2292, July 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kikuchi:2009:BBM

[3264] Takashi Kikuchi and John Gittins. A behavioral Bayes method to determine the sample size of a clinical trial considering efficacy and safety. Statistics in Medicine, 28(18):2293–2306, August 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Candel:2009:VCS

[3265] Math J. J. M. Candel and Gerard J. P. Van Breukelen. Varying cluster sizes in trials with clusters in one treatment arm: Sample size adjustments when testing treatment effects with linear mixed models. Statistics in Medicine, 28(18):2307–2324, August 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanderWal:2009:SCA

[3266] W. M. van der Wal, M. Prins, B. Lumbreras, and R. B. Geskus. A simple G-computation algorithm to quantify the causal effect of a secondary illness on the progression of a chronic disease. Statistics in Medicine, 28 (18):2325–2337, August 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shults:2009:CSA

[3267] Justine Shults, Wenguang Sun, Xin Tu, Hanjoo Kim, Jay Amsterdam, Joseph M. Hilbe, and Thomas Ten-Have. A comparison of several approaches for choosing between working correlation structures in generalized estimating equation analysis of longitudinal binary data. Statistics in Medicine, 28(18):2338–2355, August 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ma:2009:MHC

[3268] Renjun Ma, M. Tariqul Hasan, and Gary Sneddon. Modelling heterogeneity in clustered count data with extra zeros using compound Poisson random effect. Statistics in Medicine, 28(18):2356–2369, August 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wan:2009:SRS

[3269] Shuwen Wan and Biao Zhang. Semiparametric ROC surfaces for continuous diagnostic tests based on two test measurements. Statistics in Medicine, 28(18):2370–2383, August 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chu:2009:MAD

[3270] Haitao Chu, Lei Nie, Stephen R. Cole, and Charles Poole. Meta-analysis of diagnostic accuracy studies accounting for disease prevalence: Alternative parameterizations and model selection. Statistics in Medicine, 28 (18):2384–2399, August 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McArdle:2009:MSA

[3271] Alison McArdle, Uwe Kruger, and Juergen Hahn. Multivariate statistical analysis applied to an IL6 signal transduction model in hepatocytes. Statistics in Medicine, 28(19):2401–2434, August 30, 2009. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alfo:2009:SML

[3272] Marco Alfò and Antonello Maruotti. A selection model for longitudinal binary responses subject to non-ignorable attrition. Statistics in Medicine, 28(19):2435–2450, August 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2009:EAT

[3273] Hua Yun Chen and Shasha Gao. Estimation of average treatment effect with incompletely observed longitudinal data: Application to a smoking cessation study. Statistics in Medicine, 28(19):2451–2472, August 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:2009:EAH

[3274] Michael Schemper, Samo Wakounig, and Georg Heinze. The estimation of average hazard ratios by weighted Cox regression. Statistics in Medicine, 28(19):2473–2489, August 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ye:2009:GAC

[3275] Jun Ye, Nicole A. Lazar, and Yehua Li. Geostatistical analysis in clustering fMRI time series. Statistics in Medicine, 28(19):2490–2508, August 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2009:SBC

[3276] Guanghan F. Liu, Kaifeng Lu, Robin Mogg, Madhuja Mallick, and Devan V. Mehrotra. Should baseline be a covariate or dependent variable in analyses of change from baseline in clinical trials? Statistics in Medicine, 28(20):2509–2530, September 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [3380].

Cangul:2009:TTE

[3277] M. Z. Cangul, Y. R. Chretien, R. Gutman, and D. B. Rubin. Testing treatment effects in unconfounded studies under model misspecification: Logistic regression, discretization, and their combination. Statistics in Medicine, 28(20):2531–2551, September 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Villarroel:2009:CAU

[3278] Luis Villarroel, Guillermo Marshall, and Anna E. Barón. Cluster analysis using multivariate mixed effects models. Statistics in Medicine, 28(20): 2552–2565, September 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sposto:2009:APH

[3279] Richard Sposto and Paul S. Gaynon. An adjustment for patient heterogeneity in the design of two-stage phase II trials. Statistics in Medicine, 28(20):2566–2579, September 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dang:2009:UAA

[3280] Xin Dang, Stephine Lena Keeton, and Hanxiang Peng. A unified approach for analyzing exchangeable binary data with applications to developmental toxicity studies. Statistics in Medicine, 28(20):2580–2604, September 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nie:2009:EVP

[3281] Lei Nie, Haitao Chu, and Shibao Feng. Estimating variance parameters from multivariate normal variables subject to limit of detection: MLE, REML, or Bayesian approaches? Statistics in Medicine, 28(20):2605– 2616, September 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:2009:TSD

[3282] Richard J. Cook, Pierre-Jerome Bergeron, Jean-Marie Boher, and Yang Liu. Two-stage design of clinical trials involving recurrent events. Statistics in Medicine, 28(21):2617–2638, September 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McNamee:2009:ITP

[3283] Roseanne McNamee. Intention to treat, per protocol, as treated and instrumental variable estimators given non-compliance and effect heterogeneity. Statistics in Medicine, 28(21):2639–2652, September 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chappell:2009:WSR

[3284] F. M. Chappell, G. M. Raab, and J. M. Wardlaw. When are summary ROC curves appropriate for diagnostic meta-analyses? Statistics in Medicine, 28(21):2653–2668, September 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleischer:2009:SMD

[3285] Frank Fleischer, Birgit Gaschler-Markefski, and Erich Bluhmki. A statistical model for the dependence between progression-free survival and overall survival. Statistics in Medicine, 28(21):2669–2686, September 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chakraborty:2009:DMI

[3286] Bibhas Chakraborty, Linda M. Collins, Victor J. Strecher, and Susan A. Murphy. Developing multicomponent interventions using fractional factorial designs. Statistics in Medicine, 28(21):2687–2708, September 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fridley:2009:BHN

[3287] Brooke L. Fridley, Gregory Jenkins, Daniel J. Schaid, and Liewei Wang. A Bayesian hierarchical nonlinear model for assessing the association between genetic variation and drug cytotoxicity. Statistics in Medicine, 28(21):2709–2722, September 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:2009:LEG

[3288] S. D. Walter, A. Gafni, and S. Birch. Letter to the Editor: A geometric confidence ellipse approach to the estimation of the ratio of two variables. Statistics in Medicine 2008; 27:5956–5974. Statistics in Medicine, 28(21): 2723, September 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3052].

Bollen:2009:PAVb

[3289] Kenneth A. Bollen, Richard D. Lennox, and Darren L. Dahly. Practical application of the vanishing tetrad test for causal indicator measurement models: an example from health related quality of life. Statistics in Medicine, 28(21):2724, September 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2009:BII

[3290] Hui Xie. Bayesian inference from incomplete longitudinal data: a simple method to quantify sensitivity to nonignorable dropout. Statistics in Medicine, 28(22):2725–2747, September 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hyun:2009:ACI

[3291] Seunggeun Hyun, Yanqing Sun, and Rajeshwari Sundaram. Assessing cumulative incidence functions under the semiparametric additive risk model. Statistics in Medicine, 28(22):2748–2768, September 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reich:2009:EIP

[3292] Nicholas G. Reich, Justin Lessler, Derek A. T. Cummings, and Ron Brookmeyer. Estimating incubation period distributions with coarse data. Statistics in Medicine, 28(22):2769–2784, September 30, 2009. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hens:2009:CSG

[3293] N. Hens, A. Wienke, M. Aerts, and G. Molenberghs. The correlated and shared gamma frailty model for bivariate current status data: an illustration for cross-sectional serological data. Statistics in Medicine, 28 (22):2785–2800, September 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsui:2009:SSR

[3294] Shigeyuki Matsui and Tomonori Oura. Sample sizes for a robust ranking and selection of genes in microarray experiments. Statistics in Medicine, 28(22):2801–2816, September 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rajicic:2009:ARB

[3295] Natasa Rajicic, Dianne M. Finkelstein, and David A. Schoenfeld. Analysis of the relationship between longitudinal gene expressions and ordered categorical event data. Statistics in Medicine, 28(22):2817–2832, September 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Solomon:2009:SSB

[3296] Patricia J. Solomon. Some statistics in bioinformatics: The Fifth Armitage Lecture. Statistics in Medicine, 28(23):2833–2856, October 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jo:2009:UPS

[3297] Booil Jo and Elizabeth A. Stuart. On the use of propensity scores in principal causal effect estimation. Statistics in Medicine, 28(23):2857– 2875, October 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:2009:FSB

[3298] Yan Lin, Robert G. Newcombe, Stuart Lipsitz, and Rickey E. Carter. Fully specified bootstrap confidence intervals for the difference of two independent binomial proportions based on the median unbiased estimator. Statistics in Medicine, 28(23):2876–2890, October 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBuuren:2009:ERC

[3299] Stef van Buuren, Daniel J. Hayes, D. Mikis Stasinopoulos, Robert A. Rigby, Feiko O. ter Kuile, and Dianne J. Terlouw. Estimating regional centile curves from mixed data sources and countries. Statistics in Medicine, 28(23):2891–2911, October 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2009:GGF

[3300] Peng Chen, Joshua M. Tebbs, and Christopher R. Bilder. Global goodness-of-fit tests for group testing regression models. Statistics in Medicine, 28(23):2912–2928, October 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:2009:EIE

[3301] Jun-Mo Nam. Efficient interval estimation of a ratio of marginal probabilities in matched-pair data: Non-iterative method. Statistics in Medicine, 28(23):2929–2935, October 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsong:2009:CNC

[3302] Yi Tsong, Joanne Zhang, and Jinglin Zhong. Comment on ‘New confidence bounds for QT studies’ by Dennis D. Boos, David Hoffman, Robert Kringle and Ji Zhang, Statistics in Medicine 2007; 26:3801–3817. Statistics in Medicine, 28(23):2936–2938, October 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2498].

Boos:2009:RLT

[3303] Dennis D. Boos, David Hoffman, Robert Kringle, and Ji Zhang. Response to letter by tsong, Zhang, and zhong. Statistics in Medicine, 28(23):2938– 2940, October 15, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ivanova:2009:ADI

[3304] Anastasia Ivanova, Ken Liu, Ellen Snyder, and Duane Snavely. An adaptive design for identifying the dose with the best efficacy/tolerability profile with application to a crossover dose-finding study. Statistics in Medicine, 28(24):2941–2951, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lennox:2009:EED

[3305] Kristin P. Lennox and Michael Sherman. Efficient experimental design for binary matched pairs data. Statistics in Medicine, 28(24):2952–2966, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Todem:2009:HMB

[3306] David Todem and Karen Patricia Williams. A hierarchical model for binary data with dependence between the design and outcome success probabilities. Statistics in Medicine, 28(24):2967–2988, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pacheco:2009:PAM

[3307] Gonzalo Durán Pacheco, Jan Hattendorf, John M. Colford, Jr., Daniel Mäusezahl, and Thomas Smith. Performance of analytical methods for overdispersed counts in cluster randomized trials: Sample size, degree of clustering and imbalance. Statistics in Medicine, 28(24):2989–3011, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paoletti:2009:CMC

[3308] X. Paoletti and A. Kramar. A comparison of model choices for the continual reassessment method in phase I cancer trials. Statistics in Medicine, 28(24):3012–3028, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:2009:SPT

[3309] Zhangsheng Yu, Wanzhu Tu, and Mei-Ling Ting Lee. A semi-parametric threshold regression analysis of sexually transmitted infections in adolescent women. Statistics in Medicine, 28(24):3029–3042, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koziol:2009:CCO

[3310] J. A. Koziol. Comments on ‘Choosing an optimal method to combine P - values’ by S. Won, N. Morris, Q. Liu, R. C. Elston, Statistics in Medicine 2009; 28:1537–1553. Statistics in Medicine, 28(24):3043–3045, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3213].

Graham:2009:BRA

[3311] Petra L. Graham. Book review: Applied nonparametric statistical methods (4th edn). Peter Sprent and Nigel C. Smeeton, Chapman and Hall/ CRC, Boca Raton, 2007. No. of pages: 544. Price: $83.95. ISBN 13: 978-1-58488-701-0. Statistics in Medicine, 28(24):3046–3047, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hamza:2009:BRM

[3312] Taye H. Hamza. Book review: Meta-analysis of binary data using profile likelihood. Dankmar Böhning, Ronny Kuhnert and Sasivimol Rattanasiri, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: 208. Price:

$79.95. ISBN: 978-1-58488-630-3. Statistics in Medicine, 28(24):3047– 3048, October 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lunn:2009:BPE

[3313] David Lunn, David Spiegelhalter, Andrew Thomas, and Nicky Best. The BUGS project: Evolution, critique and future directions. Statistics in Medicine, 28(25):3049–3067, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cowles:2009:CBP

[3314] Mary Kathryn Cowles. Comments on ‘the BUGS project: Evolution, critique and future directions’. Statistics in Medicine, 28(25):3068–3069, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelman:2009:CBP

[3315] Andrew Gelman. Comments on ‘the BUGS project: Evolution, critique and future directions’. Statistics in Medicine, 28(25):3070–3072, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Plummer:2009:CBP

[3316] Martyn Plummer. Comments on ‘the BUGS project: Evolution, critique and future directions’. Statistics in Medicine, 28(25):3073–3074, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Teixeira-Pinto:2009:CBP

[3317] Armando Teixeira-Pinto and Sharon-Lise T. Normand. Comments on ‘the BUGS project: Evolution, critique, and future directions’. Statistics in Medicine, 28(25):3075–3078, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wakefield:2009:CBP

[3318] Jon Wakefield. Comments on ‘the BUGS project: Evolution, critique and future directions’. Statistics in Medicine, 28(25):3079–3080, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lunn:2009:RCB

[3319] David Lunn, David Spiegelhalter, Andrew Thomas, and Nicky Best. Rejoinder to commentaries on ‘the BUGS project: Evolution, critique and future directions’. Statistics in Medicine, 28(25):3081–3082, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:2009:BDC

[3320] Peter C. Austin. Balance diagnostics for comparing the distribution of baseline covariates between treatment groups in propensity-score matched samples. Statistics in Medicine, 28(25):3083–3107, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:2009:IED

[3321] Hsin-Neng Hsieh, Hsiu-Yuan Su, and Xiao-Hua Zhou. Interval estimation for the difference in paired areas under the ROC curves in the absence of a gold standard test. Statistics in Medicine, 28(25):3108–3123, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2009:CSS

[3322] Lingling Li. A conditional sequential sampling procedure for drug safety surveillance. Statistics in Medicine, 28(25):3124–3138, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahn:2009:BIS

[3323] Jaeil Ahn, Bhramar Mukherjee, Mousumi Banerjee, and Kathleen A. Cooney. Bayesian inference for the stereotype regression model: Application to a case-control study of prostate cancer. Statistics in Medicine, 28(25):3139–3157, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartlett:2009:LMM

[3324] Jonathan W. Bartlett, Bianca L. De Stavola, and Chris Frost. Linear mixed models for replication data to efficiently allow for covariate measurement error. Statistics in Medicine, 28(25):3158–3178, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ryu:2009:SCI

[3325] Euijung Ryu. Simultaneous confidence intervals using ordinal effect measures for ordered categorical outcomes. Statistics in Medicine, 28(25): 3179–3188, November 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2009:MCT

[3326] Stephen Senn and Steven Julious. Measurement in clinical trials: a neglected issue for statisticians? Statistics in Medicine, 28(26):3189–3209, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:2009:CMC

[3327] V. T. Farewell and B. D. M. Tom. Commentary on ‘measurement in clinical trials: a neglected issue for statisticians?’. Statistics in Medicine, 28(26):3210–3212, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:2009:CMC

[3328] John A. Lewis. Commentary on ‘measurement in clinical trials: a neglected issue for statisticians?’. Statistics in Medicine, 28(26):3213–3214, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lubsen:2009:CMC

[3329] Jacobus Lubsen. Commentary on ‘measurement in clinical trials: a neglected issue for statisticians?’. Statistics in Medicine, 28(26):3215–3217, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanHouwelingen:2009:CMC

[3330] Hans C. van Houwelingen. Commentary on ‘measurement in clinical trials: a neglected issue for statisticians?’. Statistics in Medicine, 28(26): 3218–3219, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:2009:CMC

[3331] Janet Wittes. Commentary on ‘measurement in clinical trials: a neglected issue for statisticians?’. Statistics in Medicine, 28(26):3220–3222, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2009:ARC

[3332] Stephen Senn and Steven Julious. Authors’ rejoinder to commentaries on ‘measurement in clinical trials: a neglected issue for statisticians?’. Statistics in Medicine, 28(26):3223–3225, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elbert:2009:DED

[3333] Yevgeniy Elbert and Howard S. Burkom. Development and evaluation of a data-adaptive alerting algorithm for univariate temporal biosurveillance data. Statistics in Medicine, 28(26):3226–3248, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiba:2009:BCE

[3334] Yasutaka Chiba. Bounds on causal effects in randomized trials with noncompliance under monotonicity assumptions about covariates. Statistics in Medicine, 28(26):3249–3259, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrison:2009:MAB

[3335] L. Harrison, D. T. Dunn, H. Green, and A. J. Copas. Modelling the association between patient characteristics and the change over time in a disease measure using observational cohort data. Statistics in Medicine, 28(26):3260–3275, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:2009:SAE

[3336] Christopher Cox, Haitao Chu, and Alvaro Muñoz. Survival attributable to an exposure. Statistics in Medicine, 28(26):3276–3293, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2009:RLD

[3337] Yufan Zhao, Michael R. Kosorok, and Donglin Zeng. Reinforcement learning design for cancer clinical trials. Statistics in Medicine, 28(26): 3294–3315, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2009:ITL

[3338] Xianzheng Huang. An improved test of latent-variable model misspecification in structural measurement error models for group testing data. Statistics in Medicine, 28(26):3316–3327, November 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Han:2009:RMU

[3339] Baoguang Han, Nathan H. Enas, and Damian McEntegart. Randomization by minimization for unbalanced treatment allocation. Statistics in Medicine, 28(27):3329–3346, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2009:AGS

[3340] Yiyi Chen and Brian J. Smith. Adaptive group sequential design for phase II clinical trials: a Bayesian decision theoretic approach. Statistics in Medicine, 28(27):3347–3362, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muthen:2009:EDE

[3341] Bengt Muthén and Hendricks C. Brown. Estimating drug effects in the presence of placebo response: Causal inference using growth mixture modeling. Statistics in Medicine, 28(27):3363–3385, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dalen:2009:CMC

[3342] Ingvild Dalen, John P. Buonaccorsi, Joseph A. Sexton, Petter Laake, and Magne Thoresen. Correction for misclassification of a categorized exposure in binary regression using replication data. Statistics in Medicine, 28(27):3386–3410, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:2009:BAM

[3343] Juxin Liu, Paul Gustafson, Nicola Cherry, and Igor Burstyn. Bayesian analysis of a matched case-control study with expert prior information on both the misclassification of exposure and the exposure-disease association. Statistics in Medicine, 28(27):3411–3423, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2009:FMA

[3344] Man-Hua Chen, Xingwei Tong, and Jianguo Sun. A frailty model approach for regression analysis of multivariate current status data. Statistics in Medicine, 28(27):3424–3436, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sylvestre:2009:FMC

[3345] Marie-Pierre Sylvestre and Michal Abrahamowicz. Flexible modeling of the cumulative effects of time-dependent exposures on the hazard. Statistics in Medicine, 28(27):3437–3453, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:2009:STA

[3346] Yun Zhao, Andy H. Lee, Kelvin K. W. Yau, Valerie Burke, and Geoffrey J. McLachlan. A score test for assessing the cured proportion in the long-term survivor mixture model. Statistics in Medicine, 28(27):3454– 3466, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCandless:2009:BRB

[3347] Lawrence McCandless. Book review: Bayesian methods for data analysis. (3rd edn). Bradley P. Carlin and Thomas A. Louis, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: 552. Price: $69.95. ISBN 978-1-58488-697-6. Statistics in Medicine, 28(27):3467–3468, November 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:2009:E

[3348] Emmanuel Lesaffre, Dominique Declerck, and Adriano Decarli. Editorial. Statistics in Medicine, 28(28):3469, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:2009:DAS

[3349] Emmanuel Lesaffre, Bruce Philstrom, Ian Needleman, and Helen Worthington. The design and analysis of split-mouth studies: What statisticians and clinicians should know. Statistics in Medicine, 28(28):3470– 3482, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petracci:2009:SAC

[3350] Elisabetta Petracci, Mauro Farella, Carlotta Galeone, Alessandra Albano, Monica Ferraroni, and Adriano Decarli. Survival analysis with clustered observations of orthodontic brackets. Statistics in Medicine, 28 (28):3483–3491, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bandyopadhyay:2009:BMM

[3351] Dipankar Bandyopadhyay, Brian J. Reich, and Elizabeth H. Slate. Bayesian modeling of multivariate spatial binary data with applications to dental caries. Statistics in Medicine, 28(28):3492–3508, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berchialla:2009:FBG

[3352] Paola Berchialla, Ileana Baldi, Vincenzo Notaro, Sandro Barone-Monfrin, Francesco Bassi, and Dario Gregori. Flexibility of Bayesian generalized linear mixed models for oral health research. Statistics in Medicine, 28(28):3509–3522, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Slate:2009:IMS

[3353] Elizabeth H. Slate and Dipankar Bandyopadhyay. An investigation of the MC-SIMEX method with application to measurement error in periodontal outcomes. Statistics in Medicine, 28(28):3523–3538, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilthorpe:2009:MCD

[3354] Mark S. Gilthorpe, Morten Frydenberg, Yaping Cheng, and Vibeke Baelum. Modelling count data with excessive zeros: The need for class prediction in zero-inflated models and the issue of data generation in choosing between zero-inflated and generic mixture models for dental caries data. Statistics in Medicine, 28(28):3539–3553, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spada:2009:AHD

[3355] Elena Spada, Andrea Carlo Butti, Monica Ferraroni, and Silvano Milani. Adapting Häävikko’s dental age for the assessment of Italian children: use of LMS and other models based on smoothing splines. Statistics in Medicine, 28(28):3554–3561, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neuenschwander:2009:NPP

[3356] Beat Neuenschwander, Michael Branson, and David J. Spiegelhalter. A note on the power prior. Statistics in Medicine, 28(28):3562–3566, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [3379].

Chen:2009:MTB

[3357] Yuh-Ing Chen and Chi-Shen Huang. Multiple testing for bioequivalence with pharmacokinetic data in 2 × 2 crossover designs. Statistics in Medicine, 28(28):3567–3579, December 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Szymczak:2009:DSE

[3358] S. Szymczak, B.-W. Igl, and A. Ziegler. Detecting SNP-expression associations: a comparison of mutual information and median test with standard statistical approaches. Statistics in Medicine, 28(29):3581–3596, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Archer:2009:ORP

[3359] K. J. Archer and V. R. Mas. Ordinal response prediction using bootstrap aggregation, with application to a high-throughput methylation data set. Statistics in Medicine, 28(29):3597–3610, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fang:2009:TFA

[3360] Yixin Fang and Yuanjia Wang. Testing for familial aggregation of functional traits. Statistics in Medicine, 28(29):3611–3625, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leman:2009:MSI

[3361] Scotland C. Leman, Foster Levy, and Elaine S. Walker. Modeling the spread of infectious disease using genetic information within a marked branching process. Statistics in Medicine, 28(29):3626–3642, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munzel:2009:NNI

[3362] Ullrich Munzel. Nonparametric non-inferiority analyses in the three-arm design with active control and placebo. Statistics in Medicine, 28(29): 3643–3656, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nevalainen:2009:MVL

[3363] Jaakko Nevalainen, Michael G. Kenward, and Suvi M. Virtanen. Missing values in longitudinal dietary data: a multiple imputation approach based on a fully conditional specification. Statistics in Medicine, 28(29): 3657–3669, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clegg:2009:EAA

[3364] Limin X. Clegg, Benjamin F. Hankey, Ram Tiwari, Eric J. Feuer, and Brenda K. Edwards. Estimating average annual per cent change in trend analysis. Statistics in Medicine, 28(29):3670–3682, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comment [3383].

Song:2009:QIF

[3365] Peter X.-K. Song, Zhichang Jiang, Eunjoo Park, and Annie Qu. Quadratic inference functions in marginal models for longitudinal data.

Statistics in Medicine, 28(29):3683–3696, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clark:2009:SST

[3366] Robert Graham Clark. Sampling of subpopulations in two-stage surveys. Statistics in Medicine, 28(29):3697–3717, December 20, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailey:2009:DDE

[3367] R. A. Bailey. Designs for dose-escalation trials with quantitative responses. Statistics in Medicine, 28(30):3721–3738, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Atkinson:2009:CDD

[3368] A. C. Atkinson. Commentary on ‘designs for dose-escalation trials with quantitative responses’. Statistics in Medicine, 28(30):3739–3741, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haines:2009:CDD

[3369] Linda M. Haines. Commentary on ‘designs of dose-escalation trials with quantitative responses’. Statistics in Medicine, 28(30):3742–3744, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OQuigley:2009:CDD

[3370] John O’Quigley. Commentary on ‘designs for dose-escalation trials with quantitative responses’. Statistics in Medicine, 28(30):3745–3750, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberger:2009:CDD

[3371] William F. Rosenberger and Zorayr Manukyan. Commentary on ‘designs for dose-escalation trials with quantitative responses’. Statistics in Medicine, 28(30):3751–3753, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:2009:CDD

[3372] Stephen Senn. Commentary on ‘designs for dose-escalation trials with quantitative responses’. Statistics in Medicine, 28(30):3754–3758, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailey:2009:ARC

[3373] R. A. Bailey. Author’s rejoinder to commentaries on ‘designs for doseescalation trials with quantitative responses’. Statistics in Medicine, 28 (30):3759–3760, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:2009:ARP

[3374] G. Y. Zou. Assessment of risks by predicting counterfactuals. Statistics in Medicine, 28(30):3761–3781, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fiocco:2009:MAP

[3375] M. Fiocco, H. Putter, and J. C. van Houwelingen. Meta-analysis of pairs of survival curves under heterogeneity: a Poisson correlated gammafrailty approach. Statistics in Medicine, 28(30):3782–3797, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huang:2009:CAC

[3376] Jian Huang, Agus Salim, Kaibin Lei, Kathleen O’Sullivan, and Yudi Pawitan. Classification of array CGH data using smoothed logistic regression model. Statistics in Medicine, 28(30):3798–3810, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:2009:FML

[3377] Huichao Chen, Amita K. Manatunga, Robert H. Lyles, Limin Peng, and Michele Marcus. Flexible modeling of longitudinal highly skewed outcomes. Statistics in Medicine, 28(30):3811–3828, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

leCessie:2009:ALD

[3378] Saskia le Cessie, Elisabeth G. E. de Vries, Ciska Buijs, and Wendy J. Post. Analyzing longitudinal data with patients in different disease states during follow-up and death as final state. Statistics in Medicine, 28(30): 3829–3843, December 30, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gajewski:2010:CNP

[3379] Byron J. Gajewski. Comments on ‘A note on the power prior’ by Neuenschwander B., Branson M. and Spiegelhalter D. J. Statistics in Medicine; DOI: 10.1002/sim.3722. Statistics in Medicine, 29(6):708–709, March 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3356].

Kenward:2010:LES

[3380] Michael G. Kenward, Ian R. White, and James R. Carpenter. Letter to the Editor: Should baseline be a covariate or dependent variable in analyses of change from baseline in clinical trials? by G. F. Liu, K. Lu, R. Mogg, M. Mallick and D. V. Mehrotra, Statistics in Medicine 2009; 28:2509–2530. Statistics in Medicine, 29(13):1455–1456, June 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3276].

Singer:2010:LEC

[3381] Julia Singer. Letter to the Editor: Construction of confidence limits about effect measures: a general approach, by G. Y. Zou and A. Donner, Statistics in Medicine 2008; 27:1693–1702. Statistics in Medicine, 29(16): 1757–1759, July 20, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2744].

Yang:2010:CNI

[3382] Zhao Yang and Xuezheng Sun. Comments on ‘Non-inferiority tests for clustered matched-pair data’ by J. Nam and D. Kwon, Statistics in Medicine 2009; 28:1668–1679. Statistics in Medicine, 29(17):1857–1858, July 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3225].

Muggeo:2010:CEA

[3383] Vito M. R. Muggeo. Comment on ‘Estimating average annual per cent change in trend analysis’ by Clegg LX, Hankey BF, Tiwari R, Feuer EJ, Edwards BK, Statistics in Medicine 2009; 28:3670–3682. Statistics in Medicine, 29(18):1958–1960, August 15, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3364].

Dibaj:2010:LEM

[3384] Shiva Dibaj, Soghrat Faghihzadeh, and Shohreh Jalaie. Letter to the Editor: Mistake on Quantifying the indirect treatment effect via surrogate markers by Y. Qu and M. Case Statistics in Medicine 2006; 25:223–231. Statistics in Medicine, 29(19):2067, August 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1917].

Li:2011:CPA

[3385] Gang Li. Comments on ‘Planning and analysis of three-arm noninferiority trials with binary endpoints’ by M. Kieser and T. Friede, Statistics in Medicine 2007; 26:253–273. Statistics in Medicine, 30(3): 298–299, February 10, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2249].

Chen:2011:CCO

[3386] Zhongxue Chen and Saralees Nadarajah. Comments on ‘Choosing an optimal method to combine p-values’ by Sungho Won, Nathan Morris, Qing Lu and Robert C. Elston, Statistics in Medicine 2009; 28:1537– 1553. Statistics in Medicine, 30(24):2959–2961, October 30, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [3213].

Noma:2012:ODP

[3387] Hisashi Noma and Shigeyuki Matsui. The optimal discovery procedure in multiple significance testing: an empirical Bayes approach. Statistics in Medicine, 31(2):165–176, January 30, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2020:CCA

[3388] Xiang Li, Kevin Liu, and Hong Tian. Comments on “Critical aspects of the Bayesian approach to phase I cancer trials”. Statistics in Medicine, 39(27):4100, November 30, 2020. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).